A Complete Bibliography of Publications in IEEE Spectrum (1970–1979)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

02 May 2020 Version 1.01

Title word cross-reference

$-efficient [Ele75b]. $100 [Ano74-232, Gin75]. $24 [Her79]. $25 [Fal75f]. $29.95 [Ano75-232]. 3 [Tor76d]. $323 [Her79]. $350 [Sug78c]. $500 [Mau75]. $750 [Fal75f]. $88 [Bea74b]. $9.95 [Ano76-206]. _{2} [All79d, Ano70-296, Wat71]. ± [Ano76-206]. x [Ano77-178]. y [Ano77-178].

-D [Tor76d].

1 [Ano79-188, Ano79-189]. 1-GHz [Ano79-188]. 1-MBd [Ano79-189]. 1.2 [Ano70-296]. 1.2-GHz [Ano70-296]. 1/2 [Ano79-191]. 1/2-digit [Ano79-191].

10.6-nm [Ano70-296]. 10.7-megabyte [Ano79-195]. 1000 [Fri77c]. 1000-kV [Fri77c]. 1100-kg [HL79]. 12 [Ano79-190]. 12-bit [Ano75-235, Ano78-191]. 12-MHz [Ano79-190]. 13 [All74d]. 14 [Dou76]. 16-bit [Ano75-231, Ano79-193]. 16-kb [Ano75-241]. 16-pin [PM77]. 1700-km-long [Eng75]. 1970 [Ano70a, Gra70c]. 1971 [Ano71b]. 1972 [Ano72a]. 1977 [Sug78c, WZ78]. 1980s [Ano79-160, Cla77, VHM77].

2-digit [Ano79-191]. 200th [Fri75d]. 24 [Ano79-159]. 256-kbyte [Ano78-193]. 25th [Tor76b].

3 [Ano79-232]. 30-cm [KLM^{+}79]. 32-pin [Ano79-193].

4-bit [Ano75-238]. 4-watt [Lap74c]. 4.8 [BFM78]. 4.8-MW [BFM78].

5 [Qua79]. 500-kb [Lap74c].

64-kbit [Gos79a, Too78a]. ’65 [Fri76c]. 65-kb [Tor76c].

7 [Hor78]. ’70 [Ano70-297, Ano70-455, Lap73b]. ’70s [Jur71]. ’72 [Bea72a, Ano72-519, Ell72, Ano72-253, Ano72-521, Ano72-520, Ano72-522]. ’73 [Che73, Ano73-303, Ano73-306, Lap73h]. ’74 [All74b, Ano75-153, Chr74n, All74a, All74d, Lap74e]. ’75 [Chr75m, All75g, Ano75-317]. 75th [Mas79d]. ’76 [All76d, Chr76a, And72f]. ’77 [All77a, Chr77l, All77g, All77h]. ’78 [Chr78m, Ano78-165, Ano78-258, Ele78]. ’79 [Chr79m, Ano79-159, Ano79-160].

8-bit [Ano79-194]. ’80s [Lin77, Ano79-159].

92-kbit [Too78a].

A/D [Ano75-235, Ano78-191, Ano79-193, Ele72a, HGB79, SF72a]. abandonment [Wil76b]. ability [Ang74]. aboard [Bla77, Hor78, MF79]. abound [All76e]. abounds [All79e]. above [Fri77c]. abreast [Jen78]. absorption [MVE75]. ac [EH76, Fri76g, Hua70, Kau72]. academia [Rub74f]. Academy [Gre79a]. Accelerating [Woo79]. accepted [Ano79-235]. access [Men75d, VCG71]. accessed [Fet76, Reu78]. accident [Mas79e, Rub79a]. accidents [Bra77]. accommodate [She79]. account [Rub79a]. Accreditation [Rub77a]. accrediting [Rub77a]. accuracy [Bor73, HL79, von79]. accurately [McW77]. achieve [McK74]. achieved [Ano79-188]. achievement [Ang74, Jen78]. Acoustic [And70, KM71, MVE75, Wal78]. acquisition [Ano74-42, Ano74-224, Ano78-201, Bre75]. act [Jur73c]. Action [Rub79c, Jed73]. actions [Wei78b]. activate [Jur75a]. active [Gef70, Mos70b, Mos70c, PM77]. Activities [Wei78b, Ano74-208, Ste75a]. actually [Der79, Wol76a]. actuated [LC75]. actuator [Jur75c]. added [RS77]. additions [Ano77-29]. address [Ano77-179, Gra70c]. addressed [Ano75-304]. addresses [Fri76i, Rub76c]. addressing [TG78]. adequate [FR75, Gol79]. adherence [Kop77]. adjustment [Ano75-42]. Administration [Fri76h]. adopting [Bra77]. Advance [Ano70b, Ano70c, Ano70d, Ano70e, Ano70f, Ano70g, Ano71c, Ano71d, Ano71e, Ano71f, Ano71g, Ano71h, Ano71i, Ano71j, Ano71k, Ano71l, Ano71m, Ano71n, Ano72b, Ano72c, Ano72d, Ano72e, Ano72f, Ano72g, Ano72h, Ano72i, Ano72j, Ano72k, Ano72l, Ano72m, Ano72n, Ano73b, Ano73c, Ano73d, Rub77b]. Advanced [And79, Ano75-232, Fre74, Kap79d, All77e, Jur76a, Kal79, Kap79a, Kop77, Qua79]. Advances [All76c, Ano71o, Ish77, Kap79c, All76a, All76e, Fri76g, HH70, Jur78b, SA78, Tor76c, Tor78b]. advancing [Tor76d]. advantages [Ano75-242, CSS76, Kap77a, Men75d, Mid74]. adversity [IFSK77]. Advertisement [Ano70h, Ano70i, Ano70-34, Ano70-35, Ano70-36, Ano70-37, Ano70-38, Ano70-39, Ano70-40, Ano70-41, Ano70-42, Ano70-43, Ano70-44, Ano70-45, Ano70-46, Ano70-47, Ano70-48, Ano70-49, Ano70-50, Ano70-51, Ano70-52, Ano70-53, Ano70-54, Ano70-55, Ano70-56, Ano70-57, Ano70-58, Ano70-59, Ano70j, Ano70-60, Ano70-61, Ano70-62, Ano70-63, Ano70-64, Ano70-65, Ano70-66, Ano70-67, Ano70-68, Ano70-69, Ano70-70, Ano70-71, Ano70-72, Ano70-73, Ano70-74, Ano70-75, Ano70-76, Ano70-77, Ano70-78, Ano70-79, Ano70-80, Ano70-81, Ano70-82, Ano70-83, Ano70-84, Ano70-85, Ano70k, Ano70-86, Ano70-87, Ano70-88, Ano70-89, Ano70-90, Ano70-91, Ano70-92, Ano70-93, Ano70-94, Ano70-95, Ano70-96, Ano70-97, Ano70-98, Ano70-99, Ano70-100, Ano70-101, Ano70-102, Ano70-103, Ano70-104, Ano70-105, Ano70-106, Ano70-107, Ano70-108, Ano70-109, Ano70-110, Ano70-111, Ano70l, Ano70-112, Ano70-113, Ano70-114, Ano70-115, Ano70-116]. Advertisement [Ano70-117, Ano70-118, Ano70-119, Ano70-120, Ano70-121, Ano70-122, Ano70-123, Ano70-124, Ano70-125, Ano70-126, Ano70-127, Ano70-128, Ano70-129, Ano70-130, Ano70-131, Ano70-132, Ano70-133, Ano70-134, Ano70-135, Ano70-136, Ano70-137, Ano70m, Ano70-138, Ano70-139, Ano70-140, Ano70-141, Ano70-142, Ano70-143, Ano70-144, Ano70-145, Ano70-146, Ano70-147, Ano70-148, Ano70-149, Ano70-150, Ano70-151, Ano70-152, Ano70-153, Ano70-154, Ano70-155, Ano70-156, Ano70-157, Ano70-158, Ano70-159, Ano70-160, Ano70-161, Ano70-162, Ano70-163, Ano70n, Ano70-164, Ano70-165, Ano70-166, Ano70-167, Ano70o, Ano70p, Ano70q, Ano70r, Ano70s, Ano70t, Ano70u, Ano70v, Ano70w, Ano70x, Ano70y, Ano70z, Ano70-27, Ano70-28, Ano70-29, Ano70-30, Ano70-31, Ano70-32, Ano70-33, Ano71p, Ano71-41, Ano71-42, Ano71-43, Ano71-44, Ano71-45, Ano71-46, Ano71-47, Ano71-48, Ano71-49, Ano71-50, Ano71-51, Ano71-52, Ano71-53, Ano71-54, Ano71-55, Ano71-56]. Advertisement [Ano71-57, Ano71-58, Ano71-59, Ano71-60, Ano71-61, Ano71-62, Ano71-63, Ano71-64, Ano71-65, Ano71-66, Ano71q, Ano71-67, Ano71-68, Ano71-69, Ano71-70, Ano71-71, Ano71-72, Ano71-73, Ano71-74, Ano71-75, Ano71-76, Ano71-77, Ano71-78, Ano71-79, Ano71-80, Ano71-81, Ano71-82, Ano71-83, Ano71-84, Ano71-85, Ano71-86, Ano71-87, Ano71-88, Ano71-89, Ano71-90, Ano71-91, Ano71-92, Ano71r, Ano71-93, Ano71-94, Ano71-95, Ano71-96, Ano71-97, Ano71-98, Ano71-99, Ano71-100, Ano71-101, Ano71-102, Ano71-103, Ano71-104, Ano71-105, Ano71-106, Ano71-107, Ano71-108, Ano71-109, Ano71-110, Ano71-111, Ano71s, Ano71t, Ano71u, Ano71v, Ano71w, Ano71x, Ano71y, Ano71z, Ano71-27, Ano71-28, Ano71-29, Ano71-30, Ano71-31, Ano71-32, Ano71-33, Ano71-34, Ano71-35, Ano71-36, Ano71-37, Ano71-38, Ano71-39, Ano71-40, Ano72o, Ano72p, Ano72-40, Ano72-41, Ano72-42, Ano72-43, Ano72-44, Ano72-45, Ano72-46, Ano72-47]. Advertisement [Ano72-48, Ano72-49, Ano72-50, Ano72-51, Ano72-52, Ano72-53, Ano72-54, Ano72-55, Ano72-56, Ano72-57, Ano72-58, Ano72-59, Ano72-60, Ano72-61, Ano72-62, Ano72-63, Ano72-64, Ano72-65, Ano72q, Ano72-66, Ano72-67, Ano72-68, Ano72-69, Ano72-70, Ano72-71, Ano72-72, Ano72-73, Ano72-74, Ano72-75, Ano72-76, Ano72-77, Ano72-78, Ano72-79, Ano72-80, Ano72-81, Ano72-82, Ano72-83, Ano72-84, Ano72-85, Ano72-86, Ano72-87, Ano72-88, Ano72-89, Ano72-90, Ano72-91, Ano72r, Ano72-92, Ano72-93, Ano72-94, Ano72-95, Ano72-96, Ano72-97, Ano72-98, Ano72-99, Ano72-100, Ano72-101, Ano72-102, Ano72-103, Ano72-104, Ano72-105, Ano72-106, Ano72-107, Ano72-108, Ano72-109, Ano72-110, Ano72-111, Ano72-112, Ano72-113, Ano72-114, Ano72-115, Ano72-116, Ano72-117, Ano72s, Ano72-118, Ano72-119, Ano72-120, Ano72-121, Ano72-122, Ano72-123, Ano72-124, Ano72-125, Ano72-126, Ano72-127, Ano72-128, Ano72-129, Ano72-130, Ano72-131, Ano72-132]. Advertisement [Ano72-133, Ano72-134, Ano72-135, Ano72-136, Ano72-137, Ano72-138, Ano72-139, Ano72-140, Ano72-141, Ano72-142, Ano72-143, Ano72t, Ano72-144, Ano72-145, Ano72-146, Ano72-147, Ano72-148, Ano72-149, Ano72-150, Ano72-151, Ano72-152, Ano72-153, Ano72-154, Ano72-155, Ano72-156, Ano72-157, Ano72-158, Ano72-159, Ano72-160, Ano72-161, Ano72-162, Ano72-163, Ano72u, Ano72v, Ano72w, Ano72x, Ano72y, Ano72z, Ano72-27, Ano72-28, Ano72-29, Ano72-30, Ano72-31, Ano72-32, Ano72-33, Ano72-34, Ano72-35, Ano72-36, Ano72-37, Ano72-38, Ano72-39, Ano73e, Ano73f, Ano73-30, Ano73-31, Ano73-32, Ano73-33, Ano73-34, Ano73-35, Ano73-36, Ano73-37, Ano73-38, Ano73-39, Ano73-40, Ano73-41, Ano73-42, Ano73-43, Ano73-44, Ano73-45, Ano73-46, Ano73-47, Ano73-48, Ano73-49, Ano73-50, Ano73-51, Ano73-52, Ano73-53, Ano73-54, Ano73-55, Ano73g, Ano73-56, Ano73-57, Ano73-58, Ano73-59, Ano73-60, Ano73-61, Ano73-62, Ano73-63, Ano73-64]. Advertisement [Ano73-65, Ano73-66, Ano73-67, Ano73-68, Ano73-69, Ano73-70, Ano73-71, Ano73-72, Ano73-73, Ano73-74, Ano73-75, Ano73-76, Ano73-77, Ano73-78, Ano73-79, Ano73-80, Ano73-81, Ano73h, Ano73-82, Ano73-83, Ano73-84, Ano73-85, Ano73-86, Ano73-87, Ano73-88, Ano73-89, Ano73-90, Ano73-91, Ano73-92, Ano73-93, Ano73-94, Ano73-95, Ano73-96, Ano73-97, Ano73-98, Ano73-99, Ano73-100, Ano73-101, Ano73-102, Ano73-103, Ano73-104, Ano73-105, Ano73-106, Ano73-107, Ano73i, Ano73-108, Ano73-109, Ano73-110, Ano73-111, Ano73-112, Ano73-113, Ano73-114, Ano73-115, Ano73-116, Ano73-117, Ano73-118, Ano73-119, Ano73-120, Ano73-121, Ano73-122, Ano73-123, Ano73-124, Ano73-125, Ano73-126, Ano73-127, Ano73-128, Ano73-129, Ano73-130, Ano73-131, Ano73-132, Ano73-133, Ano73j, Ano73-134, Ano73-135, Ano73-136, Ano73-137, Ano73-138, Ano73-139, Ano73-140, Ano73-141, Ano73-142, Ano73-143, Ano73-144, Ano73-145, Ano73-146, Ano73-147, Ano73-148, Ano73-149]. Advertisement [Ano73-150, Ano73-151, Ano73-152, Ano73-153, Ano73-154, Ano73-155, Ano73-156, Ano73k, Ano73l, Ano73m, Ano73n, Ano73o, Ano73p, Ano73q, Ano73r, Ano73s, Ano73t, Ano73u, Ano73v, Ano73w, Ano73x, Ano73y, Ano73z, Ano73-27, Ano73-28, Ano73-29, Ano74w, Ano74x, Ano74y, Ano74z, Ano74-27, Ano74-28, Ano74-29, Ano74-30, Ano74-31, Ano74-32, Ano74a, Ano74b, Ano74c, Ano74d, Ano74e, Ano74f, Ano74g, Ano74h, Ano74i, Ano74j, Ano74k, Ano74l, Ano74m, Ano74n, Ano74o, Ano74p, Ano74q, Ano74r, Ano74s, Ano74t, Ano74u, Ano74v, Ano75w, Ano75x, Ano75y, Ano75z, Ano75-27, Ano75-28, Ano75-29, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano75f, Ano75g, Ano75h, Ano75i, Ano75j, Ano75k, Ano75l, Ano75m, Ano75n, Ano75o, Ano75p, Ano75q, Ano75r, Ano75s, Ano75t, Ano75u, Ano75v, Ano76b, Ano76c, Ano76d]. Advertisement [Ano76e, Ano76f, Ano76g, Ano76h, Ano76i, Ano76j, Ano76k, Ano76l, Ano76m, Ano76n, Ano76o, Ano76p, Ano76q, Ano76r, Ano76s, Ano76t, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano77g, Ano77h, Ano78a, Ano78b, Ano78-27, Ano78-28, Ano78-29, Ano78-30, Ano78-31, Ano78c, Ano78d, Ano78e, Ano78f, Ano78g, Ano78h, Ano78i, Ano78j, Ano78k, Ano78l, Ano78m, Ano78n, Ano78o, Ano78p, Ano78q, Ano78r, Ano78s, Ano78t, Ano78u, Ano78v, Ano78w, Ano78x, Ano78y, Ano78z, Ano79a, Ano79b, Ano79-27, Ano79-28, Ano79-29, Ano79-30, Ano79-31, Ano79-32, Ano79-33, Ano79-34, Ano79-35, Ano79-36, Ano79-37, Ano79-38, Ano79-39, Ano79-40, Ano79-41, Ano79-42, Ano79-43, Ano79c, Ano79d, Ano79e, Ano79f, Ano79g, Ano79h, Ano79i, Ano79j, Ano79k, Ano79l, Ano79m, Ano79n, Ano79o, Ano79p, Ano79q, Ano79r, Ano79s, Ano79t]. Advertisement [Ano79u, Ano79v, Ano79w, Ano79x, Ano79y, Ano79z, Ano70-327, Ano70-455, Ano71-181, Ano71-301, Ano71-380, Ano71-381, Ano71-382, Ano71-420, Ano72-254, Ano72-390, Ano72-477, Ano72-478, Ano72-479, Ano72-480, Ano73-468, Ano73-469, Ano73-470, Ano73-471, Ano73-472, Ano74-108, Ano74-168, Ano74-304, Ano75-105]. Advertisements [Ano71-112, Ano74-33, Ano74-34, Ano74-35, Ano75-30, Ano75-31, Ano75-32, Ano76u, Ano76v, Ano76w, Ano76x, Ano77i, Ano77j, Ano77k, Ano77l, Ano77m, Ano77n, Ano77o, Ano70-388, Ano70-389, Ano70-390, Ano70-391, Ano70-392, Ano70-393, Ano70-394, Ano70-395, Ano70-396, Ano70-397, Ano70-398, Ano71-357, Ano71-358, Ano71-359, Ano71-360, Ano71-361, Ano71-362, Ano71-363, Ano71-364, Ano71-365, Ano71-366, Ano71-367, Ano72-455, Ano72-456, Ano72-457, Ano72-458, Ano72-459, Ano72-460, Ano72-461, Ano72-462, Ano72-463, Ano72-464, Ano73-391, Ano73-443, Ano73-444, Ano73-445, Ano73-446, Ano73-447, Ano73-448, Ano73-449, Ano73-450, Ano73-451, Ano73-452, Ano73-453, Ano77-216, Ano77-217, Ano77-218]. advertisers [Ano70-298, Ano70-299, Ano70-300, Ano70-301, Ano70-302, Ano70-303, Ano70-304, Ano70-305, Ano70-306, Ano70-307, Ano70-308, Ano70-309, Ano71-254, Ano71-255, Ano71-256, Ano71-257, Ano71-258, Ano71-259, Ano71-260, Ano71-261, Ano71-262, Ano71-263, Ano71-264, Ano71-265, Ano72-347, Ano72-348, Ano72-349, Ano72-350, Ano72-351, Ano72-352, Ano72-353, Ano72-354, Ano72-355, Ano72-356, Ano72-357, Ano72-358, Ano73-338, Ano73-339, Ano73-340, Ano73-341, Ano73-342, Ano73-343, Ano73-344, Ano73-345, Ano73-346, Ano73-347, Ano73-348, Ano73-349, Ano74-193, Ano75-199, Ano75-200, Ano76-169, Ano77-124, Ano77-125, Ano77-126, Ano77-127, Ano77-128, Ano77-129, Ano78-158, Ano78-159, Ano78-160, Ano78-161, Ano78-162, Ano78-163, Ano79-158]. advertising [Ano70-232, Ano70-233, Ano70-234, Ano70-235, Ano70-236, Ano70-237, Ano70-238, Ano70-239, Ano70-240, Ano70-241, Ano71-167, Ano71-168, Ano71-169, Ano71-170, Ano71-171, Ano71-172, Ano71-173, Ano71-174, Ano72-228, Ano72-229, Ano72-230, Ano72-231, Ano72-232, Ano72-233, Ano72-234, Ano72-235, Ano72-236, Ano72-237, Ano72-238, Ano72-239, Ano73-219, Ano73-220, Ano73-221, Ano73-222, Ano73-223, Ano73-224, Ano73-225, Ano73-226, Ano73-227, Ano73-228, Ano75-79, Ano75-80, Ano75-81, Ano78-83, Ano79-80, Ano79-81, Ano79-82]. advises [Tuc74]. Advocacy [Ano76-271]. advocate [Chr77i]. Aerospace [Hor78, How70, KRS78, KS79, MF79, Mas79b, Tor78a, Wei77, AB71, Ele75b, Ele76b, Men74e, Tor77d]. affairs [Pon70]. affect [Kap78b]. affecting [Fri76i, Mac78]. affects [Her79, Wei78a]. afford [Pet77]. afraid [Raf78]. Africa [KLH74]. after [Wol76a]. again [Fri76b, Jay76]. against [Fri75e, HB77, Nil76, Rub75a, Wol75b]. Age [Rub75a, Chr77g, Mos75, Rub74f, Yen78]. age-old [Yen78]. agencies [PHC76]. Agency [Ano79-236]. aggravations [Mas79f]. aging [Mad74]. ago [Jur76f]. agree [Ele74f]. agriculture [Rub79c]. ahead [Ano73-304, BDL^{+}78, Bus72, Chr75n, Chr79b, Cos74, FB72, Har77, Kap75b, Kop77, Lin75b, Mar72, Oli74, Oll78, Pat79b, Pun70, Que79, Wil76a]. AIChE [Rub76g]. aid [Jur78b, NS77, Rai76, Ung79]. aided [Ano74-96, Kap75e, Lia78, Sug78a]. aiding [Kap74b]. aids [All76a, Ano71-176, Ano71-177, Ano71-178, Ano71-179, Ano71-180, Ano72-240, Ano72-241, Ano72-242, Ano72-243, Ano72-244, Ano72-245, Ano72-246, Ano72-247, Ano72-248, Ano72-249, Ano72-250, Ano72-251, Ano73-230, Ano73-231, Ano73-232, Ano73-233, Ano73-234, Ano73-235, Ano73-236, Ano73-237, Ano73-238, Ano73-239, Ano73-240, Ano73-241, Ano74-108, Ano74-97, Ano74-98, Ano74-99, Ano74-100, Ano74-101, Ano74-102, Ano74-103, Ano74-104, Ano74-105, Ano74-106, Ano74-107, Ano75-105, Ano75-94, Ano75-95, Ano75-96, Ano75-97, Ano75-98, Ano75-99, Ano75-100, Ano75-101, Ano75-102, Ano75-103, Ano75-104, Ano76-84, Ano76-85, Ano76-86, Ano76-87, Ano76-88, Ano76-89, Ano76-90, Ano76-91, Bea71, Fri70b, Jur76b, Sur72]. aimed [Gan77]. ain’t [Pie79]. Air [Cot75, Kin75, Bib71a, Bib71b, Bib71c, Ele75a, Ele76b, Fri70a, Fri70b, Fri70c, Fri76i, HB77, Jur75a, Kap76e, Kap78c, Men76f, Okr78]. air-pollution [Bib71a, Bib71b, Bib71c]. air-traffic [Fri70a, Fri70b, Fri70c]. airborne [Fri70c]. aircraft [KRS78, LJ70, Mas79b, SA78]. airline [Cot75]. airports [Fri70a]. airspace [Fri70b]. Alfred [Dre74]. alike [Ano79-235]. all-electric [Fri74i]. all-new [All76d]. all-weather [KRS78]. allows [Ano74-231]. almost [Wol76a]. aloft [MF79]. alone [Fal77c]. along [Ano76a, Ano76-273, DTW76, Gal77, Gre79d, Jur78a, Pol70]. alongside [Mun78]. already [KI77, Sch77b, Tor76d]. also [Ano75-234]. alternate [RO77]. alternative [EH76, Fal74b]. alternatives [All79h, Chr74c, Chr74d, Chr74e, Chr74f, Jay76, Rub74f]. Although [Ele75a, Fri76j, Rub77d]. altogether [Bra77]. Aluminum [Fal74f]. always [McW77, Wol76a]. AM [Men78a, Men79b, PM77]. AM/FM [PM77]. AMA [Rub76b]. Amateur [Dun78, Gor75]. Amateurs [Cam77]. Amazing [Chr77b, All78d, All79g, All78d]. ambition [Tor77g]. ambitious [GHNK79]. amendments [Ano76a]. American [Fri74e]. Amid [Jur77a, Jay76]. among [All79h, Fri74j, Men75b]. Amorphous [All77f, Lap73a]. amp [Ano74-223, Ano75-236, Ano78-197, MM74]. Ampère [Fri75d]. amplifier [Bla77, CT70, Lap74c, Mos70c]. amplifiers [Ano75-40, HGB79, LW70, Nay71a, Nay71b, Sch71]. amplitude [Ano75-42]. amps [MM74]. amputees [All76a]. AMTRAK [Fri75c]. anachronistic [Fri74e]. Analog [Hal77, Mor78, NS77, Ano78-34, Ano78-37, Ano79-145, Gre76a, HGB79, Men74d, Nay71a, Nay71b, Pit79]. analyser [Gre76a]. analyses [Gan77]. Analysis [Chr79o, Gre74, All78c, Ano74-293, Ano78-36, Bor73, BHLK71, For71, Hor78, KS79, Rot71, Tri74, Yu75]. Analysis/perspective [Gre74]. Analyst [Gre76a]. analysts [All76i]. analyzer [And78, Ano78-43, Ano78-192, Ano79-54]. analyzers [All74b, All77c, Hil74]. analyzing [Dou76]. Anatomy [WZ78, Joh70b]. and/or [All77c]. Angeles [All74d]. anniversary [Fri75d, Mas79d]. Announcement [Ano70-327, Ano71-253]. Announcing [Chr77b]. annoying [Tor77e]. Annual [Ano70-168, Ano71-113, Ano72-164, Ano73-181, Ano77a, Ano74-208, Ano74-60, Ano75-153, Ano75-57, Ano76-135, Ano77-43, Ano78-62, Dil75, Her79]. answer [BCY75, Kun75, Men74c, Tuc75a]. answers [Lom79, Rub74d]. Antenna [Ano70-375, AFCK79, Reu78]. antennas [Bro79, Fri71c, Mac76]. anticipation [Lin75b]. antique [Tuc75b]. any [Ano75-238, Ano78-199, Chr75i, SR78, TT74]. anyone [Fag78]. applicable [Rub74e]. application [Ano74-47, Ano74-46, Ano76-204, Bia70, Bre71, For70, Jur77a, MM74]. Applications [Ano71-114, Ano71-115, Ano71-116, Ano71-117, Ano71-118, Ano72-165, Ano72-166, Ano72-167, Ano72-168, Ano72-169, Ano72-170, Ano72-171, Ano72-172, Ano72-173, Ano72-174, Ano72-175, Ano72-176, Ano73-157, Ano73-158, Ano73-159, Ano73-160, Ano73-161, Ano73-162, Ano73-163, Ano73-164, Ano73-165, Ano73-166, Ano73-167, Ano73-168, Ano74-40, Ano74-39, Ano74-41, Ano74-42, Ano74-44, Ano74-43, Ano74-45, Ano74-47, Ano74-46, Ano74-36, Ano74-37, Ano74-38, Ano75-38, Ano75-39, Ano75-40, Ano75-41, Ano75-42, Ano75-43, Ano75-33, Ano75-34, Ano75-35, Ano75-36, Ano75-37, Ano76-36, Ano76-35, Ano76y, Ano76z, Ano76-27, Ano76-28, Ano76-29, Ano76-30, Ano76-31, Ano76-32, Ano76-33, Ano76-34, Ano77p, Ano77q, Ano77r, Ano77s, Ano77t, Ano77u, Ano77v, Ano77w, Ano77x, Ano77y, Ano77z, Ano77-27, Ano78-34, Ano78-36, Ano78-35, Ano78-37, Ano78-38, Ano78-39, Ano78-40, Ano78-41, Ano78-42, Ano78-43, Ano78-32]. Applications [Ano78-33, Ano78-44, Ano78-45, Ano78-46, Ano78-47, Ano78-48, Ano79-45, Ano79-44, Ano79-53, Ano79-54, Ano79-51, Ano79-49, Ano79-50, Ano79-52, Chr76a, Fal74a, Fal77b, Jur75a, Kra75, Wal70, All77f, Ano70-328, Ano70-329, Ano70-330, Ano70-331, Ano70-332, Ano70-333, Ano70-334, Ano70-335, Ano70-336, Ano70-337, Ano70-338, Ano70-339, Ano71-311, Ano71-312, Ano71-313, Ano71-314, Ano71-315, Ano71-316, Ano71-317, Ano71-318, Ano71-319, Ano71-320, Ano71-321, Ano71-322, Ano72-406, Ano72-407, Ano72-408, Ano72-409, Ano72-410, Ano72-411, Ano72-412, Ano72-413, Ano72-414, Ano72-415, Ano72-416, Ano72-417, Ano73-392, Ano73-393, Ano73-394, Ano73-395, Ano73-396, Ano73-397, Ano73-398, Ano73-399, Ano73-400, Ano73-401, Ano73-402, Ano73-403, Ano74-230, Ano74-227, Ano74-233, Ano74-235, Ano74-234, Ano74-224, Ano74-226, Ano74-222, Ano74-232, Ano74-229, Ano74-231, Ano74-228, Ano74-225, Ano74-223, Ano75-239, Ano75-232]. applications [Ano75-233, Ano75-237, Ano75-234, Ano75-240, Ano75-231, Ano75-241, Ano75-243, Ano75-238, Ano75-242, Ano75-236, Ano75-235, Ano75-230, Ano76-204, Ano76-194, Ano76-195, Ano76-205, Ano76-196, Ano76-197, Ano76-198, Ano76-199, Ano76-200, Ano76-206, Ano76-201, Ano76-202, Ano76-203, Ano77-178, Ano77-168, Ano77-169, Ano77-170, Ano77-171, Ano77-172, Ano77-173, Ano77-179, Ano77-174, Ano77-175, Ano77-176, Ano77-177, Ano78-49, Ano78-194, Ano78-198, Ano78-201, Ano78-199, Ano78-192, Ano78-200, Ano78-191, Ano78-197, Ano78-195, Ano78-190, Ano78-196, Ano78-193, Ano79-145, Ano79-192, Ano79-193, Ano79-188, Ano79-190, Ano79-191, Ano79-194, Ano79-185, Ano79-186, Ano79-189, Ano79-195, Ano79-187, Dun78, EBC^{+}79, Ele76a, Fag79, Fal74c, Fal76b, Har77, Jur70, Jur71, Kap78a, Men78c, Mor78, Mos75, Nay71a, Nay71b, Pet77, Rei74, Sch71, Shi79b, Sne72a, Sne72b, VHM77, Pol71b]. applied [Gre79b, Mas79c]. Applying [Ano78-36]. appraises [Cot75]. approach [Bib71a, Bib71b, Bib71c, BS71, CC78, Pit79]. approaches [All79c, Pog78]. appropriate [Zuc77]. appropriation [Gin75]. April [Ano79-159]. Arc [Rie70c]. Arc-medium [Rie70c]. archaeologist [Rai76]. Archaeology [Rai76]. architect [Lin76b]. architecture [Fag78]. arcologies [Lin76b].

Area [Fri74d, Ano75-319, Fri70c]. areas [Fal75c, Rub74e]. arena [Tri78]. Arizona [Lin76b]. arm [Tri78]. arms [All76a, Mas79a]. Array [Ano70-375, Ano70-399]. arrestment [Chr76e]. arsenal [All74b]. arsenide [VL77]. art [Cam77, Chr75b, Fin70, Fri73b, Fri75h, Lin75b, Mas79a, Rai76, Sug79c]. artful [Shi79a]. Articles [Lin75b]. arts [All77f, Lom79]. artwork [Ano78-42, Bea71]. ask [Sug78e]. asks [Gre79b]. aspects [GHM71, Kap77f, Mac78, Rub74e]. assembled [Ano74-228]. assembly [Ano78-193, Mit77, Tor77b]. assess [All75f]. Assessing [Chr76d, Man79]. Assessment [Chr76e, Koh71, Fle71, Sug79c]. assessments [Gan77]. assessors [Chr76d]. Assets [Chr77d]. assigned [PW71]. assignment [Kap75e]. assisted [TH77]. associated [Ber79]. Associates [Ano79-232]. assumed [Tor77e]. assumptions [Col78c]. assurance [Spi76b]. assure [Ano74-235]. astronaut [Lin74a]. AT&T [Men77a]. ATC [Fal73d]. ATE [Tor77a]. atmospheric [Wat71]. ATS [TT74]. attack [Ano71-421, Ele74b]. attempt [Ste77]. attendees [All77a]. attention [Rac78]. attitudes [Ano77-36]. audio [All76h, Ber79, Lap74c, Men79b]. audio/video [All76h]. audiotape [Cam77]. Auditor [Ano75-44, Ano76-37, Ano77-28]. aural [Uts77]. Aurelio [Fri76e]. Author [Ano76-38, Ano79-46, Ano70-169, Ano71-120, Bla77, Pet77, Rec75, Zuc77]. authorities [Men78d]. authors [Ano72-319, Ano73-305, Ano74-157, Ano75-152, Ano76-134, Ano77-117, Ano78-147, Ano78-240, Ano79-144]. Auto [McW77, Jur78a, Jur78c, Mit77]. automate [Fal75e]. automated [Fri70a, Fri70b, Men76g]. automates [Fri74h]. Automatic [Cla71, Ele73a, Ele74a, Ele76a, NR74, RRS^{+}74, TT74, Wad76, FK76, Fri70c, Fri74h, Fri75g, McA71, Oll78, Tor77a]. Automating [Tor77b, Sug78b]. Automation [EK71, All75i, All78a, Fri70b, Fri74g, Mok79c, Ura77]. automobile [Jon77a, Jur77a, McW77]. Automotive [Jon77b, Jur75b, Jur75c, Jur78a, KI77, PMG77]. Autos [Kap77a, Jon77b, Jur77a, SK73]. availability [Kap77d, MNS77b, Men78c]. available [Ano77-36, Bea74b, Hil74, Lus78, Zuc77]. availed [Ano75-153]. aviation [KH79]. avidly [Fri75f]. Avionics [Cot75, Ele75a, Pog78, Dou76, Kin75]. Avocation [Gor75, RM75, Tuc75b]. avoid [Jur76a, SK73]. avoidance [Chr79h, Ele75a, Fri70b]. avoided [Bra77]. Avoiding [Lit75, PG76]. await [Har77]. Award [Ano79-47]. Awards [Ano77-29, Ano79-47, Ano79-48]. aware [Ele74b]. away [Nag74, Rub75f].

B.E.T. [Men74c]. bachelor [Men74c]. Back [Ano71-121, Ano74-48, Ano75-45, Ano75-46, Ano77-30, Ano77-31, Ano77-32, Ano77-33, Ano77-34, Ano77-35, Ano71-272, Ano74-194, Ano75-201, Ano75-202, Ano77-130, Ano77-131, Ano77-132, Ano77-133, Ano77-134, Ano77-135, Bea72e, Bus72, Chr79b, Cos74, Fri76b, RF75]. back-to-schoolers [Bea72e]. backing [Gre74]. backup [Chr75k, Fri75g]. bad [And72c]. bag [Ano72-252, Fri74g]. balance [Rin70]. balanced [Kap77a]. Balancing [HB77, Jur73c]. Ballistic [Joh70a]. Ballistic-missile [Joh70a]. balloon [Okr78]. Band [Ano70-375, Ano78-197, Men76c, Men78d]. bandits [Ele72b]. bands [Lus78]. Bandwidth [Ano70-296, JM77, Ano79-188, Cam79, Men76c]. bang [All77b]. banker [SZ74]. bankers [Chr77h]. Banking [Ano77-36, BKS79]. Banking/commerce [Ano77-36]. banner [Fri75h]. Banquet [Gra70c]. barrels [BC77]. barriers [Ano77-36]. BART [Ano72-177, Fri72a, Fri72b, Fri72h, Fri73a, Fri73g, Fri74d, Fri75g, Fri75h]. based [All78b, Bro79, Doi77, Kor77, Lap73h, Rub79a, Tor77h, Wei77]. Basic [Ano74-40, Ano74-39, Ano75-38, Ano76-36, Ano76-35, Gre79b, Jur74g, Dav79, TT74]. basics [Ano79-160]. Batteries [Men76d, PG76, BKS79]. battery [Ano75-234, von79]. battle [Tor77f]. Bay [Ano79-160, Fri74d]. Be [Men74d, All77c, Ano77-36, Ano78-165, Bea75, Bra77, Bro79, Chr79o, EH76, Fal74g, Fal75c, Fey74, Fra78, Fri74b, Fri75c, Gra74, Gre79a, Hea74, Hel79, HGB79, Jon77b, Jur76d, Kap76a, KLH74, Lia78, Lus78, Mas79a, MRR^{+}77, Men74e, Men77e, Men78g, NS77, Nil76, PHC76, Rub75c, Spe78, Sug78a, Til70, Too78b, Tuc74, Ung79, VHM77]. beach [Chr75g]. beam [Bre71, Fet76, Kra75, Mas79c, TC73a]. beaming [AFCK79]. beams [KRS78, Reu78]. Beans [Ele78]. bearings [HL79]. beat [Men78d]. Beating [Ele72b]. Because [Rob74a, Fal74f, Fri74j]. beckons [All77h]. become [Mau75, Men74f, Tor77a]. becomes [HH74]. been [All77f, Ano79-235, Dun78, Ele75a, Pla79a, Rec75, Rub75d, Wol76a]. before [Str71]. began [SOW77, Tor77b]. beginning [Jur76f]. beginnings [Sus70a]. begins [Tor78a]. behalf [Wei78b]. Behind [And73b, Lap73b, Tro74, UG73, Eve77, MT73, Tri78]. being [Ano74-96, Ele75a, Fri77c, Mac76, Ter77]. believed [Rub79a]. believes [Hin79, Rub77c]. Bell [All75f, Ano79-95, Hin79, Ker76, Pie79]. below [Spe78]. beneficiaries [Too78b]. benefit [Bor73, Tor78b]. benefit-cost [Bor73]. benefits [Ano72-252, Mid79]. beset [Tor76d]. Best [Ano78-44, Ano78-45, Ano78-46, Ano78-47, Ano78-48, Ano78-49, Ano79-53, Ano79-54, Ano79-51, Ano79-49, Ano79-50, Ano79-52, Fri75c, Jur74f, Kor77, McW77, Pat79b, Tri74]. bets [SA78]. Better [Mok79c, And70, Cam77, CC78, Fal73d, Hin79, Hor72, Jon77b, Jur76d, Jur77a, Par78b, Rub75c, Tor77d]. Betting [KI77, Mas79c]. between [Fal75b, Men74c, Men78d, von79]. bewilder [Fal74g]. beyond [Chi73]. bias [Ano74-223]. bicentennial [Fri76a]. Big [Shi79a, SZ74, AC75, BC73a, Ele74b, MB79, Men74b, Okr78, Tor76b, Wol76c]. Bigger [Fri73a, All75h, Fet73]. bilked [SR78]. billion [AFCK79, BCH^{+}79]. binary [PP70]. bioelectricity [GH71]. bioelectrode [Ged72]. biological [Jur74a]. Biomechanical [AB71]. biomedic [Jur76b]. Biomedical [All76a, Jur76a, Spe78]. bionic [Spe78]. bipolar [Ano74-228, Ano75-39, Ano75-236, Ano78-199, PM77, Tor77f]. Birth [Ano71-406, Fri75d]. bit [AC75, All75b, Ano75-231, Ano75-243, Ano75-238, Ano75-235, Ano76-206, Ano78-191, Ano79-193, Ano79-194, BCG73, Bre75, Men74a, Sal77]. bit-by-bit [BCG73]. bite [Men74f, RO77]. bits [Ano78-44, Ano78-45, Ano78-46, Ano78-47, Ano78-48, Ano78-49, Ano78-198, Ano79-53, Ano79-54, Ano79-51, Ano79-49, Ano79-50, Ano79-52, BCH^{+}79, Que79]. black [Gra74, MRR^{+}77]. Blackout [Chr77c, Fri76c, Sug78c, WZ78]. blackouts [Ewa78, Par78a]. Blacks [Rub75b]. blast [Lin74a]. blessing [Jur77b]. block [Gre71, Mok79c]. blocks [All75a, Ano74-228, Lap74a]. blooming [Jur72b]. blossomed [Men77b]. blow [Mas79e]. blow-by-blow [Mas79e]. blown [Oll78]. blue [Ele72b, Fri75g]. blue-box [Ele72b]. Blueprint [Fin72a, Rub77b]. blurring [Tor77f]. Board [Wei78b, Ano75-238, Ano78-201, Ano78-199]. Boards [Ano74-208]. bolts [Fri72b]. bonding [Mit77]. Book [Ano70-170, Ano70-171, Ano70-172, Ano70-173, Ano70-174, Ano70-175, Ano70-176, Ano70-177, Ano70-178, Ano70-179, Ano70-180, Ano70-181, Ano70-327, Ano71-122, Ano71-123, Ano71-124, Ano71-125, Ano71-126, Ano71-127, Ano71-128, Ano71-129, Ano71-130, Ano71-131, Ano71-132, Ano71-133, Ano72-178, Ano72-179, Ano72-180, Ano72-181, Ano72-182, Ano72-183, Ano72-184, Ano72-185, Ano72-186, Ano72-187, Ano72-188, Ano72-189, Ano73-169, Ano73-170, Ano73-171, Ano73-172, Ano73-173, Ano73-174, Ano73-175, Ano73-176, Ano73-177, Ano73-178, Ano73-179, Ano73-180, Ano74-49, Ano74-50, Ano74-51, Ano74-52, Ano74-53, Ano74-54, Ano74-55, Ano74-56, Ano74-57, Ano74-58, Ano74-59, Ano75-47, Ano75-48, Ano75-49, Ano75-50, Ano75-51, Ano75-52, Ano75-53, Ano75-54, Ano75-55, Ano75-56, Ano76-39, Ano76-40, Ano76-41, Ano76-42, Ano76-43, Ano76-44, Ano76-45, Ano76-46, Ano76-47, Ano76-48, Ano76-49, Ano76-50, Ano77-37, Ano77-38, Ano77-39, Ano77-40, Ano77-41, Ano77-42]. Book [Ano78-50, Ano78-51, Ano78-52, Ano78-53, Ano78-54, Ano78-55, Ano78-56, Ano78-57, Ano78-58, Ano78-59, Ano78-60, Ano78-61, Ano79-55, Ano79-56, Ano79-57, Ano79-58, Ano79-59, Ano79-60, Ano79-61, Ano79-62, Ano79-63, Ano79-64, Ano79-65, Ano79-66, FS77, GJH77, JE77, KH77, OGG77, SC77, SB75, TSR74, WB75, Bea74b]. books [Ano75-154, Ano75-155, Ano75-156, Ano75-157, Ano75-158, Ano75-159, Ano75-160, Ano75-161, Ano75-162, Ano76-140, Ano76-141, Ano76-142, Ano76-143, Ano76-144, Ano76-145, Ano76-146, Ano76-147, Ano76-148, Lin75b, RM75]. bookshelf [All75i, RM75]. Boom [Ano79-95]. boomingest [Lin77]. boost [All78a, Jur75a, Kap76d, Men75a, Sim72a].

Booths [Ano72-390]. bootleggers [Wol75b]. bootstrap [Tib78]. Born [Jay76, Rub76g]. Boss [Dib72, TDK74]. Boston [Ele78]. Both [Ker76, Mid74, All76h, And79, Fal76c, Fri75b, Jon77a, JF74, Kap77a, MRR^{+}77, Par78a, Rub74f, Wul78]. bottleneck [Ano78-48]. bound [Tri78]. boundaries [Gor75]. box [Ele72b, MRR^{+}77]. bracket [Ano74-71]. braking [Kal75]. brand [Men77b]. brand-new [Men77b]. bread [Ele78]. break [Ano78-48]. Breaker [Ano70-199, Ano70-200, Ano70-201, Ano70-202, Ano70-203, Ano70-204, Ano70-205, Ano70-206, Ano70-207, Ano70-208, Ano70-209, Ano70-210, Ano70-211, Ano70-212, Ano70-213, Ano70-214, Ano70-215, Ano70-216, Ano70-217, Ano71-140, Ano71-141, Ano71-142, Ano71-143, Ano71-144, Ano71-145, Ano71-146, Ano71-147, Ano71-148, Ano71-149, Ano71-150, Ano71-151, Ano71-152, Ano71-153, Ano71-154, Ano72-202, Ano72-203, Ano72-204, Ano72-205, Ano72-206, Ano72-207, Ano72-208, Ano72-209, Ano72-210, Ano72-211, Ano72-212, Ano72-213, Ano72-214, Ano73-193, Ano73-194, Ano73-195, Ano73-196, Ano73-197, Ano73-198, Ano73-199, Ano73-200, Ano73-201, Ano73-202, Ano73-203, Ano73-204, Ano73-205, Mid74]. breakers [Ano74-45, Mid74, Rie70a, Rie70b, Rie70c]. breakfast [Wol75a]. Breaking [Hus74]. breaks [Ano79-53, Dav79]. breakthrough [Ber79, Reu78]. Breakthroughs [Gre79d, Jur74h]. breakup [Ano79-95]. bred [Lin76a]. breed [Ano74-96]. breeder [Fri74b, Kap76b]. Brian [Ele78]. brief [All75g, Ano74-209, Ano75-217, Ano75-317]. briefly [Ano77-29]. brigade [Lap74c]. bright [Ele74e, KI77]. brighter [Fri77a]. brilliance [Fal76c]. bring [Ano73-377, Fri76b]. brings [Ano74-227, SZ74]. bristles [All74b]. Britannia [Fri75a]. Britannica [Ano78-164]. British [Ano70-218, Pla79b, Pla79a]. broad [Lap73h, SA78, Tor76c, Zuc77]. broad-based [Lap73h]. broadband [Man79]. Broadcast [Tel72]. broadcasting [Hav70, Men76b, Rub76i]. broadcasts [CF72]. broaden [Kap78c]. broadly [Bar78]. broke [Pie79]. brought [Fri74d]. brown [Ele78]. Browns [Kap76e]. Brushless [DKR^{+}72]. brushoff [IKT73]. bubble [All75a, Fet76, Too78b, Too78a, Tor76c]. Bubbles [Tor76c, Lap72a]. buck [All77b]. bucket [Lap74c]. budget [Ano74-227, Gre79b, Men74a, Rub76c, Rub76j, Rub76k]. budgetary [Kap75b]. Bugs [Kap75a, Fri73a]. Building [Ger74, All75a, Col79, Gre71, Ker76, Lap74a, Mau75, Yen78]. Built [All74c, All78b, Fal74e, Fri74b]. Built-in [All74c, All78b, Fal74e]. bulk [Ano75-241]. bullets [Ish77]. Bundespost [Kun75]. bureau [Hin79]. bureaucracy [Lom79]. burgeoning [Lin76a, Tor77e]. burn [Fal74f]. Burndy [Fri70d]. Business [Gre76b, Mid79, All74d, All79h, Fal75d, Fri75e, Lap74e, Lin77, Mac78, RO76, Rub76k, SZ74, Tib78]. businessman [Wei78b]. bussed [Ter77]. bust [Ano79-95]. Busy [All79a]. buttons [Ele74c]. buy [RF75]. buyer [Men78e]. buyers [McW77]. buying [Men78e]. bypass [Tor76d]. bytes [Ano78-198]. byword [Kap76b]. bywords [All75d].

Cable [Ano70-326, AC75, Ano79-189, Fri72d, GPPM71, Jur71]. cacophony [Lom79]. CAD [All75i, Kap75e]. cadmium [PG76]. cake [Ano71-301, Lin76c]. calculator [Ano78-196]. calculators [Ano74-233, Ano75-240, Jur74b, Men74b, Men76f]. Calendar [Ano70-219, Ano70-220, Ano70-221, Ano70-222, Ano70-223, Ano70-224, Ano70-225, Ano70-226, Ano70-227, Ano70-228, Ano70-229, Ano70-230, Ano71-155, Ano71-156, Ano71-157, Ano71-158, Ano71-159, Ano71-160, Ano71-161, Ano71-162, Ano71-163, Ano71-164, Ano71-165, Ano71-166, Ano72-215, Ano72-216, Ano72-217, Ano72-218, Ano72-219, Ano72-220, Ano72-221, Ano72-222, Ano72-223, Ano72-224, Ano72-225, Ano72-226, Ano73-206, Ano73-207, Ano73-208, Ano73-209, Ano73-210, Ano73-211, Ano73-212, Ano73-213, Ano73-214, Ano73-215, Ano73-216, Ano73-217, Ano74-72, Ano74-73, Ano74-74, Ano74-75, Ano74-76, Ano74-77, Ano74-78, Ano74-79, Ano74-80, Ano74-81, Ano74-82, Ano74-83, Ano75-67, Ano75-68, Ano75-69, Ano75-70, Ano75-71, Ano75-72, Ano75-73, Ano75-74, Ano75-75, Ano75-76, Ano75-77, Ano75-78, Ano76-60, Ano76-61, Ano76-62, Ano76-63, Ano76-64, Ano76-65, Ano76-66, Ano76-67, Ano76-68, Ano76-69, Ano76-70, Ano76-71, Ano77-55, Ano77-56, Ano77-57, Ano77-58, Ano78-74, Ano78-75, Ano78-76, Ano78-77]. Calendar [Ano78-78, Ano78-79, Ano78-80, Ano78-81, Ano78-82, Ano79-68, Ano79-69, Ano79-70, Ano79-71, Ano79-72, Ano79-73, Ano79-74, Ano79-75, Ano79-76, Ano79-77, Ano79-78, Ano79-79, Ano77-60, Ano77-61, Ano77-62, Ano77-63, Ano77-64, Ano77-65, Ano77-66, Ano77-67, Ano78-95, Ano78-96, Ano78-97]. calibrating [All74c]. calibration [Fal74i, FK76]. California [Ano79-160, Col79, Fri75h, Rub76h]. Call [Ano72-227, CSS76, Joh79, NR74, Wei78a]. calls [Man78]. cam [Ano78-200]. cam-operated [Ano78-200]. CAMAC [HL76]. came [Dil75, GR76]. camera [EW78, Uts77]. cameras [Jur74b, Lap73d, Men78e]. Can [Gra70a, RO76, Sug79a, All77c, Ano74-307, Ano77-178, Ano78-200, Ano78-244, Bra77, Fag78, Fal74g, Fal74e, GH72, Gra79, HGB79, Lia78, Men78c, Nas78, Reu78, RRS^{+}74, Rub76e, Str71, Tor77h, Ung79]. Canada [Gal79]. candid [BDL^{+}78]. Candidate [Ano76a]. candidates [Ano71-407, Ano72-504, Ano73-498, Ano74-209, Ano75-217, Ano76-270]. candle [Kap76e]. canned [Que79]. cannot [Pet77]. can’t [Jur74g]. capabilities [All74b, All76c]. capacitance [Ano75-233]. capacities [Kap79d]. capacity [Ano79-195]. Capital [Ano70-231, Ano73-377, CSS76, Gre74]. Capitol [Gre79a]. Captain [Jur75d]. capture [Men77c]. car [Max77, Sal72a, Tor77b, Wit77]. Card [Ano70-198, Ano70-182, Ano70-183, Ano70-184, Ano70-185, Ano70-186, Ano70-187, Ano70-188, Ano70-189, Ano70-190, Ano70-191, Ano70-192, Ano70-193, Ano70-194, Ano70-195, Ano70-196, Ano70-197, Ano71-134, Ano71-135, Ano71-136, Ano71-137, Ano71-138, Ano71-139, Ano72-190, Ano72-191, Ano72-192, Ano72-193, Ano72-194, Ano72-195, Ano72-196, Ano72-197, Ano72-198, Ano72-199, Ano72-200, Ano72-201, Ano73-182, Ano73-183, Ano73-184, Ano73-185, Ano73-186, Ano73-187, Ano73-188, Ano73-189, Ano73-190, Ano73-191, Ano73-192, Ano74-71, Ano76-59, IFSK77, SZ74]. cardiac [Jur76f, Uts77]. cardiac-output [Uts77]. cards [SZ74]. care [Jur73b]. Career [DTW76, Kop77, RO77, Fal74g, NC71]. Careers [BCY75, Bea74a, Rec75, Rub74a, Rub74b, Rub74c, Rub74d, Rub75a, Rub75b, Rub75c, Rub75d, Rub76a, TDK74]. Careful [Rac78]. carefully [Spi76a]. Carlo [Kap75e]. carpet [Ano72-522]. carrier [Hin79]. Cars [Sch77b, Cla77, Hea74, Jur75b, Tuc75b]. Carter [Ano76-270, Gre79b]. CAS [Fri70c]. Case [Ger74, RL74, Ano74-45, Ano75-303, Dou76, Fin73, Fri72g, Fri74d, Lap73c, Shi79a, Sim72b]. cases [Fra78]. Cash [Rub75e, FR75, RF75, SZ74]. cash-dispensing [SZ74]. cassette [Men78e]. CAT [Jur79]. catch [Tor76a]. catch-up [Tor76a]. categories [All75i, Gre79b]. CATV [Jur72a]. cause [All76i]. Cautious [Pay75, Fri75f]. Caveat [Rub77d]. CB [Men77c]. CCD [Ano75-241, Ano78-192, Jur75e, Too78b, Too78a]. CCDs [All75a, Fet76]. celebrated [Fri75d]. celebrates [Mas79d]. celebration [Ano73-218, Ano79-159]. cell [BFM78]. cells [PG76]. centennial [Ano79-159]. center [Ano78-165, Ano78-164]. centers [Gib76]. cents [Fri74f]. ceramics [Hol70, LH70]. certain [Fra78]. certification [Chr75e]. chaff [Chr79a]. Chairman [Ele78, Gle77]. challenge [Fri73d, Gre79d, Jon77b, Mad74, Mok79c, Pat79b]. challenged [Sug79c]. challenges [All79h, Jur74c, Moo79]. challenging [Men76e]. chance [Raf78, Sug78a]. change [Fin72a, Hin79, Rub77b]. Changes [CF72, Col78b]. Changing [Ter72, Bea73a]. channel [Kap74b]. channels [Lus78]. chapters [RS77]. character [Fre74]. characteristics [Ano74-39, Wol77]. characterize [KRS78]. characters [Ano74-231]. Charge [BS71, Jur74a, Kay78, Too78a]. Charge-coupled [BS71, Jur74a]. charge-coupled-device [Too78a]. charged [Ano75-234]. charlatans [SR78]. cheap [Bea75]. cheaper [Cla77, Fet73, Fet76, Hea74, Lap74a, Nil76]. checking [Jur74g]. checkout [NR74, SZ74]. checkups [Fal74c]. cheek [Rec75]. chess [Men77d]. chief [Hin79]. China [She78]. chip [All75b, And78, Ano75-243, Ano75-236, Ano78-198, Ano78-192, Fag78, Gos79b, Hei79, Lap74d, PM77, Que79]. Chipping [Fal77a]. chips [Ano75-39, Jur78c, MB79, Shi79a, Sug79a]. choices [All75i, Pat79b]. choose [Kap77e, McW77]. chores [Men76a, RRS^{+}74]. chronicle [Fri72a]. Chrysler [Jur78a]. Churchill [Fri71b]. Circa [Fri76a, Lin76b, Lin76a]. circle [Ano74-71]. circles [Chr74h]. Circuit [All75a, HGB79, Lap74a, Rie70a, Rie70b, Rie70c, All77c, All78c, Ano74-44, Ano74-45, Ano74-96, Ano74-222, Ano74-228, Ano75-43, BHLK71, DeF70, Gar70a, Gar70b, Gar70c, Joh70b, Mid74, Sch71, Tor76a, Wol76b]. Circuit/system [All75a, Lap74a]. circuitry [SF72b]. Circuits [Tor76a, All74c, All76f, All77b, Boa73, Fal74e, HGB79, Kle77, Lap74c, Lap74d, Men73b, NS77]. circumstances [Lin76a]. circumvent [Kap77d]. circus [All77a]. cited [Ano75-304, Dre74]. citie [Chr76j]. Cities [All76b, Fri76b, Gib76, Kap76a, Lin76c, Men76a, Nil76, Wil76b, Fri76a]. citizen [All76b]. citizen-city [All76b]. Citizens [Men76c, Han79, Men76a]. city [All76b, Fal73b, Lin76c, Lin76a, Wil76b, Sug78c]. claim [Men75d]. claims [Rac78]. Classic [Kra75, Nas78, Ele74b]. Classified [Ano70-232, Ano70-233, Ano70-234, Ano70-235, Ano70-236, Ano70-237, Ano70-238, Ano70-239, Ano70-240, Ano70-241, Ano71-167, Ano71-168, Ano71-169, Ano71-170, Ano71-171, Ano71-172, Ano71-173, Ano71-174, Ano72-228, Ano72-229, Ano72-230, Ano72-231, Ano72-232, Ano72-233, Ano72-234, Ano72-235, Ano72-236, Ano72-237, Ano72-238, Ano72-239, Ano73-219, Ano73-220, Ano73-221, Ano73-222, Ano73-223, Ano73-224, Ano73-225, Ano73-226, Ano73-227, Ano73-228, Ano75-79, Ano75-80, Ano75-81, Ano78-83, Ano79-80, Ano79-81, Ano79-82]. classroom [Eve77]. clean [HB77]. clear [Kop77]. clearly [Ele74c]. clever [All76e]. climates [Mac78]. clinical [Fri75b]. clock [LC75]. clock-actuated [LC75]. close [Ano74-222, Rub76d]. close-up [Ano74-222]. Closed [Ano74-222, NS77, Chr76k]. Closed-circuit [Ano74-222]. Closed-loop [NS77, Chr76k]. closer [All78d]. closes [Eve77]. cloud [KS79]. clouded [Kap75b]. clout [Gre79a]. Club [Fri76e]. cm [KLM^{+}79]. CMOS [Ano78-195]. co [LPJ76, All79d, Ano70-296, Wat71]. co-op [LPJ76]. Coast [All74d, Fri70f]. coax [Men78b]. coaxial [Ano79-189]. cockpit [Cot75, Fri70c]. Code [Ste75b]. coder [HGB79]. coder-decoder [HGB79]. Coding [For70]. Coexistence [Ver77]. cofounder [Dre74]. Coherent [Yu75, Oli74]. cold [Bea72f, Jur75b]. Coliseum [Ano72-255]. colleagues [Dre74]. collected [Hor76]. collective [Tri74]. collector [Tuc75b]. collision [Ele75a, Fri70b]. collision-avoidance [Fri70b]. collisions [Lit75, SK73]. Color [Gol70, Jur74a, Men75e, TG78]. Columbia [MF79]. combination [Gra74, Mid74]. combinatorial [Hil74]. combined [Ura77]. combined-cycle [Ura77]. combines [Ano75-236]. combining [Tri74]. combustion [Wit77]. come [Chr76b, Men76b]. comeback [Fri72c, Jay76, Rub77g]. comes [Jur74d]. Coming [All76h, Ano78-84, Ano78-85, Ano78-86, Ano78-87, Ano78-88, Ano78-89, Ano78-90, Ano78-91, Ano78-92, Ano78-93, Ano79-83, Ano79-84, Ano79-85, Ano79-86, Ano79-87, Ano79-88, Ano79-89, Ano79-90, Ano79-91, Ano79-92, Ano79-93, Ano79-94, Spi76a]. commands [Mar77]. commerce [Ano77-36]. Commercial [Bar79, Too78a, Mid74, Ste77, Tib78, Sus70a]. Commission [Men74f, Gle77, Mas79e]. committed [Tro74]. committee [Ele78, Rub77c]. Committees [Ano74-208]. common [Ano78-194, Hin79, Fri70a]. communicate [Bel74]. communication [And70, Ano79-53, Bar79, HS70, Kap77f, Ohi70, Reu78]. Communications [AC75, Ano77-59, Ano79-95, BCH^{+}79, Fal74b, Fal75a, Fal75c, Gal77, JM77, LPJ76, Lus78, Mac76, Mar77, Men75a, Men76c, Men76b, Men77a, Men78b, Men79a, Mok79a, Pay75, Reu78, Spi76a, Sug79a, VHM77, Ano75-237, Dun78, Fag79, Fal74a, For70, Fri70b, Her71, Hin79, Ish77, Kap79a, Mas79b, NR74, Pun70, Rob74a, RRS^{+}74, Rub76k, Rub78b, RL74, Wei77]. community [RK79]. commutation [Bat71]. Commutators [IKT73]. commuters [Hor72, Mas79f]. Compact [Ano75-233, All76g, Lin76c]. company [Buc78b, Fri74e, Raf78]. comparators [Nay71b]. compared [Fre74]. comparison [Chr79i, Jur74f, Rub76f]. compatibility [Lou74]. compatible [Ano74-233]. Compensating [Ber71]. compete [Men78b, Rea79]. competence [Ung79]. Competing [Jur74f, Wit77, Men78e, Men78a, Pog78, Roi78]. competition [Ano75-240]. competitive [BKS79, Mac78, Wil70]. complacency [Gol79]. complement [Kap76c]. Complementing [Ray77]. complements [All74a]. Complete [Ura77]. completion [Jur79]. complex [All77b, JF74, Jur78c, Lin75b, NR74, Sug79c, Ter77, Tor77a, Tuc75a]. complexities [All76b]. complexity [Tor79]. Compliance [Mas79a]. Component [Lap74b, Gre79d, Lap74d, Zuc77]. Components [All75b, All75h, All76c, Ano70-242, Men76d, PM77, Tor76a, Tor77a]. Components-Batteries [Men76d]. Comprehensive [Ano78-37]. compression [Cam79]. compromise [HB77]. compromises [Shi79a]. computation [Fal74i]. computations [Kap75e]. Computer [All78a, Bea74b, Bea71, Bel74, Fai74, Fal74c, Fal75b, Fra78, Jur73a, Kap75e, Kap77c, Kor77, Men74a, Rei74, Rob74a, Tor77h, Tro74, Ura77, ATK72, Ano74-96, Ano74-61, Ano74-227, Bea72c, Chr78e, Doi77, Fag78, Fal75e, Gre79d, Har72, Kay78, Lap74d, Lav78, Lia78, Men75d, Oll78, RS77, She79, Sug78a, Sug78e, Sug79a, TE74, TH77, War72, All78d]. Computer-aided [Kap75e, Ano74-96, Sug78a]. computer-assisted [TH77]. Computer-based [Kor77, Tor77h, Doi77]. Computer-implemented [Fra78]. computer-integrated [Oll78]. Computer-watching [Jur73a]. computerised [Obe77]. Computerization [Bec71]. Computers [Ana79, Ano78-94, Ano79-51, Bre75, Ele74d, Fal74d, Fal74e, Fal75d, Fal76a, Fal77b, Fet76, IY77, Kor77, Men78c, Mor78, Que79, SL78, She79, Sug78b, Sug78a, Sug79b, SZ74, TE74, Ano78-37, Ano78-193, Bel74, Fal75e, Fal76b, FCR^{+}70, Fri72b, Kap74b, Kay78, Mac76, Mar77, Men78f, Shi79a, Too78b]. Computing [Ana79, Men77b, Rob74a, Str72]. conceive [Bla77]. concept [Ano71-406, Kra75, Lin76b, Man72, Rob74a]. concepts [Ano76-36, Ano76-35, Kap74b, KLV75]. concern [Chr74h, Mid71]. concerned [Sim71]. concerning [BCY75, Jur77a]. Concerns [Ano74-95, Ano79-235]. conclusion [Bra79]. condition [Han79]. conditioners [Men76f]. conditioning [Hus74, Men74a]. conditions [FC78, Fle71, KRS78]. Conference [Ano77-60, Ano77-61, Ano77-62, Ano77-63, Ano77-64, Ano77-65, Ano77-66, Ano77-67, Ano78-95, Ano78-96, Ano78-97, Lap74e]. Conferences [Ano70-247]. conferencing [TH77]. confidence [Ung79]. confirmed [Rub77b]. Conflict [Chr75c, Ung79]. confluence [Ull70]. confuse [New76]. confusion [Jur72d, Rub74b]. congested [Fri70a]. congestion [Tor77h]. Congress [And72a, Gle77, Jur74a, Rub76d, Tuc74, Wei78b]. Congressman [Tuc74].

Congressman-cum-engineer [Tuc74]. conjecture [Mac76]. connection [Fri75c]. consequences [Gan77]. Conservation [Cle73, Fri73c, Jur74d, Jur74e, JF74, SS71]. conserving [CSS76, Hea74, Ker76]. considerations [Rie70b, Rie70c]. considered [Wit77]. consistent [Nag74]. Consolidated [BFM78]. Consolidating [Ano78-94]. conspiracy [Fri75e]. conspiring [Rub75a]. constant [Wal70]. constant-current [Wal70]. constellations [Fal77c]. constitute [Men78d]. constitutional [Ano76a]. Constraining [Kap77b]. constraint [Ele74c]. Constraints [Kap77d, Kap75b, Kap76a, Rub76k]. constructing [Fri76b]. construction [KLH74]. consulting [Ano70-388, Ano70-389, Ano70-390, Ano70-391, Ano70-392, Ano70-393, Ano70-394, Ano70-395, Ano70-396, Ano70-397, Ano70-398, Ano71-357, Ano71-358, Ano71-359, Ano71-360, Ano71-361, Ano71-362, Ano71-363, Ano71-364, Ano71-365, Ano71-366, Ano71-367, Ano72-455, Ano72-456, Ano72-457, Ano72-458, Ano72-459, Ano72-460, Ano72-461, Ano72-462, Ano72-463, Ano72-464, Ano73-443, Ano73-444, Ano73-445, Ano73-446, Ano73-447, Ano73-448, Ano73-449, Ano73-450, Ano73-451, Ano73-452, Ano73-453, Ano77-216, Ano77-217, Ano77-218]. Consumer [Ber79, Fal74f, Har77, Jur74b, Men74f, Men75b, Men76e, Men76f, Men77b, Men78d, Men78e, Men79b, PM77, Ano77-36, IFSK77, Jur76d, Kap77b, McW77, Men73a, Men73b, Men77c, PMG77]. Consumerism [Ano73-229]. Consumers [Men78f, McW77]. consumption [Jur77a]. contact [Yu70]. contained [Ano78-198, Fal74e, Men77d]. containing [Ano75-238]. contains [Ano78-198]. contend [Pog78]. contenders [Wit77]. Contents [Ano71-175, Ano74-84, Ano74-85, Ano74-86, Ano74-87, Ano74-88, Ano74-89, Ano74-90, Ano74-91, Ano74-92, Ano74-93, Ano74-94, Ano75-82, Ano75-83, Ano75-84, Ano75-85, Ano75-86, Ano75-87, Ano75-88, Ano75-89, Ano75-90, Ano75-91, Ano75-92, Ano75-93, Ano76-72, Ano76-73, Ano76-74, Ano76-75, Ano76-76, Ano76-77, Ano76-78, Ano76-79, Ano76-80, Ano76-81, Ano76-82, Ano76-83, Ano77-68, Ano77-69, Ano77-70, Ano77-71, Ano77-72, Ano77-73, Ano77-74, Ano77-75, Ano77-76, Ano77-77, Ano77-78, Ano77-79, Ano78-98, Ano78-99, Ano78-100, Ano78-101, Ano78-102, Ano78-103, Ano78-104, Ano78-105, Ano78-106, Ano78-107, Ano78-108, Ano79-96, Ano79-97, Ano79-98, Ano79-99, Ano79-100, Ano79-101, Ano79-102, Ano79-103, Ano79-104, Ano79-105, Ano79-106, Ano79-107, Ano70b, Ano70c, Ano70d, Ano70e, Ano70f, Ano70g, Ano70-436, Ano70-437, Ano70-438, Ano70-439, Ano70-440, Ano70-441, Ano70-442, Ano70-443, Ano70-444, Ano70-445, Ano70-446, Ano70-447, Ano71d]. contents [Ano71e, Ano71f, Ano71o, Ano71g, Ano71h, Ano71i, Ano71j, Ano71k, Ano71l, Ano71m, Ano71n, Ano71-408, Ano71-409, Ano71-410, Ano71-411, Ano71-412, Ano71-413, Ano71-414, Ano71-415, Ano71-416, Ano71-417, Ano71-418, Ano72c, Ano72d, Ano72e, Ano72f, Ano72g, Ano72h, Ano72i, Ano72j, Ano72k, Ano72l, Ano72m, Ano72n, Ano72-505, Ano72-506, Ano72-507, Ano72-508, Ano72-509, Ano72-510, Ano72-511, Ano72-512, Ano72-513, Ano72-514, Ano72-515, Ano72-516, Ano73b, Ano73c, Ano73d, Ano73-331, Ano73-332, Ano73-333, Ano73-334, Ano73-335, Ano73-336, Ano73-337, Ano73-499, Ano73-500, Ano73-501, Ano73-502, Ano73-511, Ano73-503, Ano73-512, Ano73-504, Ano73-505, Ano73-506, Ano73-507, Ano73-508, Ano73-509, Ano73-510, Ano74-180, Ano74-169, Ano74-170, Ano74-171, Ano74-172, Ano74-173, Ano74-174, Ano74-175, Ano74-176, Ano74-177, Ano74-178, Ano74-179, Ano74-306, Ano75-182, Ano75-183, Ano75-184]. contents [Ano75-185, Ano75-175, Ano75-176, Ano75-177, Ano75-178, Ano75-179, Ano75-180, Ano75-181, Ano75-186, Ano76-161, Ano76-162, Ano76-163, Ano76-158, Ano76-164, Ano76-159, Ano76-165, Ano76-166, Ano76-160, Ano78-154, Ano78-155, Ano78-148, Ano78-149, Ano78-150, Ano78-151, Ano78-152, Ano78-153, Ano79-146, Ano79-147, Ano79-148, Ano79-149, Ano79-150, Ano79-151, Ano79-156, Ano79-157, Ano79-152, Ano79-153, Ano79-154, Ano79-155]. Contest [All79g, Chr77b, Men75b, All78d]. continental [DBB71]. continue [Fal77b]. Continuing [Fal74g, Rub76j]. Continuous [Rei74, Fri74h]. contract [Rub76a]. contractor [Fal74f]. Contradictory [HB77]. contrast [MVE75]. contrasts [Fri74e]. Control [HL79, Ano79-192, Bib71a, Bib71b, Bib71c, Car79, Chi73, Cot75, Fal74i, FC78, Fri70a, Fri70b, Fri70c, Fri70f, Fri74h, Fri76f, Fri77a, Kap74b, Kap77c, Kay78, Kin75, Kor77, Kun75, LC75, Mas79a, Moo73, NS77, Obe77, Par78a, PMG77, SL78, Sch78, Ura77]. controlled [HL79, Mil71]. controller [Ano75-41, Ano76-204, Kor77]. Controllers [Kap75c, Fri70a, Jur74f]. Controlling [TE74, TC73a]. controls [Ano76-204, Ele76b, Jur75a, Kal79, Kap74b, Kus72, Pla79b, PHC76, SS73]. Controversial [McK72a]. controversy [GH71, Jur77a, Mas79c, Rub77d]. Convention [All74a, All77g, Gra70c, Tor76b, Ano70-297]. conventional [Mar77]. conversational [Ohi70]. Conversations [Wei77]. Conversion [Ull70, Ano74-42, Fal74i, Hus74, RGCS79]. convert [RGCS79]. converter [Ano76-206, Ano79-193, SF72b, Zuc77]. converters [Ano74-235, Ano75-235, Ano78-191, Ele72a, HGB79, Nay71a, SF72a]. cooking [Men75c]. cool [Ano78-38]. Cooling [And73c]. cooperation [BB71]. coordination [Gib70, IY77, Ura77]. cop [Fal74a]. cope [Lap74b]. copes [All76b]. copies [Man78]. copper [Eng75]. copy [Fal76b, Men74d]. corporate [Bea75]. Corporation [Mad74, Wes73]. corporations [Rub74a]. Correction [Ano70-243, Ano70-244, Jur72b]. Corrections [Ano70-245, Ano70-246]. correlates [Jen78]. corresponding [Ano74-71]. Cost [Lap73c, Wel79, All76g, Ano74-95, Ano75-238, Ano75-242, BJ78, Bor73, BCH^{+}79, Jon77b, Jur75e, Kap77f, Lap72b, Lou74, Men74a, Men78b, Pit79, Sal77, Sug78c, TG78, Wei78a]. cost-effective [Jon77b]. cost-efficient [All76g]. cost-per-bit [Men74a]. Cost-performance [Lap73c]. costing [Mau75]. costly [Tor77a]. costs [AC75, Alb78, All76f, All78c, Ano74-95, Ano75-232, Ano76-205, Ano77-178, Ano79-159, Dor72, JM77, Jur75b, Jur79, Lev78, MNS77b, Mit77, PHC76, Rob74a, Sug78a, VHM77]. Cotellessa [Rub76h]. could [BC77, CC78, Fal74f, Gre79a, Hea74, Lin77, Lus78, Mau75, RGCS79, She79, Wei78b]. coulometer [Ano74-229]. counter [Ano78-195]. counterpart [Ano78-200]. counterviews [Ano76a]. countries [Par78b, Rea79]. country [RG79]. country-by-country [RG79]. counts [All77e]. coupled [BS71, Jur74a, Pla79b, Too78a]. course [Kap77b]. courses [Ano74-96, Fal75f, Mil72]. court [Ano75-303, Chr79l]. courts [Fra78]. cover [Ano70-272, Ano70-273, Ano70-274, Ano70-275, Ano70-276, Ano70-277, Ano70-278, Ano70-279, Ano70-280, Ano70-281, Ano70-282, Ano70-283, Ano70-284, Ano70-285, Ano70-286, Ano70-287, Ano70-288, Ano70-289, Ano70-290, Ano70-291, Ano70-292, Ano70-293, Ano70-294, Ano70-295, Ano71-121, Ano71-206, Ano71-207, Ano71-208, Ano71-209, Ano71-210, Ano71-211, Ano71-212, Ano71-213, Ano71-214, Ano71-215, Ano71-216, Ano71-217, Ano71-218, Ano71-219, Ano71-220, Ano71-221, Ano71-222, Ano71-223, Ano71-224, Ano71-225, Ano71-226, Ano71-227, Ano71-228, Ano71-272, Ano71-273, Ano72-280, Ano72-281, Ano72-282, Ano72-283, Ano72-284, Ano72-285, Ano72-286, Ano72-287, Ano72-288, Ano72-289, Ano72-290, Ano72-291, Ano72-292, Ano72-293, Ano72-294, Ano72-295, Ano72-296, Ano72-297, Ano72-298, Ano72-299, Ano72-300, Ano72-301, Ano72-302, Ano72-303, Ano73-267, Ano73-268, Ano73-269, Ano73-270, Ano73-271, Ano73-272, Ano73-273, Ano73-274, Ano73-275, Ano73-276, Ano73-277, Ano73-278, Ano73-279, Ano73-280]. cover [Ano73-281, Ano73-282, Ano73-283, Ano73-284, Ano73-285, Ano73-286, Ano73-287, Ano73-288, Ano73-289, Ano73-290, Ano74-48, Ano74-145, Ano74-194, Ano74-195, Ano75-45, Ano75-46, Ano75-138, Ano75-139, Ano75-201, Ano75-202, Ano75-203, Ano75-204, Ano77-30, Ano77-31, Ano77-32, Ano77-33, Ano77-34, Ano77-35, Ano77-110, Ano77-111, Ano77-112, Ano77-113, Ano77-114, Ano77-115, Ano77-116, Ano77-130, Ano77-131, Ano77-132, Ano77-133, Ano77-134, Ano77-135, Ano77-136, Ano77-137, Ano77-138, Ano77-139, Ano77-140, Ano77-141, Ano77-142, Hus74, KS79]. covers [GB73]. CPU [Lap74d]. crack [Sug79a]. create [Par78b]. creating [Ano78-42, Men77b, Rub76g]. creative [LC75, Mas79d, Wol77]. credit [SZ74]. cries [Men74a]. crime [Sug79a]. crippled [Ele75b]. crisi [Fri73e]. Crisis [Fri73d, And73a, Chr74c, Chr74d, Chr74e, Chr74f, Edd70, Hus74, Kap77e, Lap74b, Nag74, Rub74e]. critical [Chr74a, Jur73c]. critics [Ano79-95]. critique [TG78]. crossroads [Fri70a, Fri70b, Fri70c]. crosstalk [DeF70]. crown [Ang74]. CRT [Ano79-192, Men74d, TG78]. CRTs [Jur74f]. crucial [Sug78b]. crunch [FR75, Rub75e, RF75]. crystal [All75c, GM72]. CTD [Ano78-192]. cubbyhole [Men77e]. cultural [Chr77j]. cum [Tuc74]. cure [SR78]. current [Ano74-122, Ano74-123, Ano74-124, Ano74-125, Ano74-126, Ano74-127, Ano74-128, Ano74-129, Ano74-130, Ano74-131, Ano74-132, Ano74-180, Ano75-118, Ano75-119, Ano75-120, Ano75-121, Ano75-122, Ano75-123, Ano75-124, Ano75-125, Ano75-126, Ano75-127, Ano75-182, Ano75-183, Ano75-184, Ano75-185, Ano75-186, Ano76-104, Ano76-105, Ano76-106, Ano76-107, Ano76-108, Ano76-109, Ano76-110, Ano76-161, Ano76-162, Ano76-163, Ano76-164, Ano76-165, Ano76-166, Ano78-154, Ano78-155, Ano79-156, Ano79-157, Fal76a, GH71, Har77, Mac78, Pat79b, Pol70, Wal70]. currents [Ano74-223]. curricula [Col78b, Ele74b, Fal74j]. curriculum [Ele74b, Flo71]. curve [Gos79b, MM74]. curves [Con70]. cushion [Ele76b]. customer [Fri76f]. customers [IFSK77]. cut [Ano77-178]. cuts [All78a, Ano76-205, Dor72, Mit77]. Cutting [Ano74-95, All79d, Jur74d]. cybernetics [Men74g]. cycle [Kap75a, TG78, Ura77].

D [Ano75-235, Ano76-206, Ano78-191,

Ano79-193, Ele72a, HGB79, SF72a, Tor76d]. D/A [Ano75-235, Ano76-206, Ano78-191, Ele72a, HGB79, SF72a]. daily [BC77, Fal75d, Men76a]. Dale [Ele78]. Dallas [Ano78-165]. damage [All75f]. damping [Mos75]. dangerous [Chr75b]. dark [Chr76j]. darkest [KLH74]. Dartmouth [Man70, Man72]. Dartmouth/Tuskegee [Man70]. Data [Ano70-399, Chr79a, Her71, Hor76, Rob74a, RRS^{+}74, Rud70, Ano74-42, Ano74-224, Ano74-231, Ano78-201, Ano79-145, Bra79, Bre75, BCH^{+}79, CC78, Fag79, Fal77c, JM77, KLM^{+}79, Kra75, KS79, Men75d, Ohi70, Sal77, Sug79a, TC73b, Zuc77]. data-acquisition [Ano78-201]. data-entry [Ano74-231]. dates [All74d]. day [Fal75f, Kap79c, Oll78]. days [Kap75b]. dc [BFM78, Hua70, Kau72, Tuc75b, Wal70]. deal [Cou79]. Dear [Dib72]. dearly [Sug78c]. debatable [Pat79b, Rub77a]. debate [Col78b, Sug79a]. debug [Ano78-43]. debut [Fal74d, Men77d, Men79b]. decade [Ano73-304, Pog78, VHM77]. decades [Joh71, Kap76a]. deception [Wal78]. Decision [Tri74, Fal74i, Lom79, Sug78b, Sur72, TH77]. decision-makers [Lom79]. decision-making [Fal74i, TH77]. decisions [CC78, Dou76, Rei74]. declining [Ano75-304]. decoder [HGB79]. decreasing [MNS77b, Men74a]. decreed [Lin76b]. dedicated [Gre74, Men77d]. dedication [Tor77g]. deduct [RO76]. defense [Ano74-95, Ele75b, Joh70a, Mas79c]. defer [Fal74g]. defined [Kap79c, Kap79d]. Defining [Ano78-242, Moo79]. degree [Kop77]. degrees [Ana79]. delay [GH70]. delayed [Ano76-273, Fri76j]. delegates [Jur74a]. deliberate [Ele74f]. delight [All76g]. delineation [Gos79b]. deliver [Bro79, EBC^{+}79, Spe78, Sug78a]. delivery [Jur73b]. Delta [Sch70]. Demand [PW71, Fri76f, Men77b, Mor78, Pla79b, Rub74b]. Demand-assigned [PW71]. demands [Kap76a, KLV75, Qua79]. demonstrate [Col79]. demonstrated [Shi79a]. demonstration [BFM78, Ste77]. Demystifying [Shi79a]. denies [Ano79-236]. denigrate [Rea79]. denigrated [Rec75]. densities [All75b, Ano79-159]. density [All77e, Bre75, HGB79, Kap77d, Lap74a, Sal77, Tor76a]. Department [Mas79c, Esp77, GHNK79]. dependence [Nag74]. deployment [RJD73]. depths [EW78, LZ73, RGCS79]. derived [Kap75b]. describe [Rub75b]. described [Ano77-29]. descriptive [KLV75]. Desensitization [Hua70]. Design [All75i, Ano74-41, Dou76, Fal75e, Kap75e, Rei74, Rie70b, Tri74, Ano74-44, Ano74-96, Ano76-205, Ano77-178, Ano78-39, Ano78-94, Bea71, Chr75j, Chr79o, Ele74c, Fin73, Fri76f, Ged72, Gre79d, HGB79, Jur72b, Kap74b, Ker76, KLV75, Lap74d, Lin76c, MNS77a, Shi79a, War76]. designed [Ano79-191, Esp77, Jur75a, Men77d, PM77, Tor77b]. Designer [All75i, Ano78-46, Men74b, All74b, Fri74g, Gol72]. Designers [SL78, Ano74-96, IY77]. Designing [Ano75-39, Boa73, Jur78c, Moo79]. designs [Ano75-241, Bra77, Fre74, Lap74d, New76, Wel79]. Desk [Ano78-193]. Desk-top [Ano78-193]. Despite [All77f, Fri76g, Men78e, Han79]. detailed [Dil75, Kra75]. Details [Mas79e, All78d]. detector [Jur76d]. detectors [Men78d, Nay71b]. détente [Fri72i]. determines [Mas79a]. Detroit [Jur78a, Jur78c, Sch77b]. Deutsche [Kun75]. develop [Ang74, Rub76f]. developed [Sug78a]. Developers [BKS79, Men75b]. developing [RRS^{+}74]. development [All77c, Ano74-226, Ano76-205, Ano78-194, Ano79-194, DTW76, Fri71b, Gre79b, Hus74, Kop77, Mit77, Rud70]. Developmental [Tor76c]. developments [Ano74-122, Ano74-123, Ano74-124, Ano74-125, Ano74-126, Ano74-127, Ano74-128, Ano74-129, Ano74-130, Ano74-131, Ano74-132, Ano75-118, Ano75-119, Ano75-120, Ano75-121, Ano75-122, Ano75-123, Ano75-124, Ano75-125, Ano75-126, Ano75-127, Ano76-104, Ano76-105, Ano76-106, Ano76-107, Ano76-108, Ano76-109, Ano76-110, Ano76-273, Fal76a, Jur74a, Kal79, Kap78a, Lap74c]. device [All76c, All76i, BS71, Jur73c, Spe78, Too78a, War76, Yu70]. Devices [EBC^{+}79, Jur76a, All75a, All75h, All77f, All77e, Ano76-35, BS71, Fri76g, Gar70a, Gar70b, Gar70c, Har77, Hor78, Jur74a, KM71, KS79, Lap73a, Lap74a, Lou74, Mar77, ME77, Shi72, Sim72a, Tor76a, Tor77f, Tor78b, Tor79, Zuc77]. Devoted [Gor75]. devouring [Rub76b]. diabetics [Spe78]. diagnoses [ME77]. diagnosis [All77c]. Diagnostic [Jur78b]. Diagnostics [Sus73, All79e, Uts77]. diagram [Pol70]. diameter [Bro79, Har77]. diaphragm [Spe78]. diary [Ano72-403]. did [Rub76f, Spi76b]. didn’t [Sug78a]. dielectric [Dor72]. different [Ano79-50, RGCS79]. difficult [Hil74, Tor77a]. difficulties [Fri76g, Rub75e]. digit [Ano78-195, Ano79-191]. Digital [All75c, All79b, And79, Ano75-234, Ano77-59, Nay71a, Nay71b, All74b, All74c, All77c, All78c, ATK72, Ano74-234, Ano75-233, Bea74b, BCH^{+}79, Fal75c, Fal77a, Gal77, Gar70a, Gar70b, Gar70c, Hil74, Kap74b, Lia78, Man79, Mas79b, Men74d, Men75a, Men76f, Men76b, Men77a, Men78b, Men79b, Mos70b, NS77, Pit79, Rot71]. Digital-transmission-line [Ano75-234]. Digitally [Ber79]. digitized [Bre75]. digits [Ber79]. dilemma [Chr74j, Ele74f, Her73, HH71, Jur76a]. Dillard [Rub76c]. dimensional [Fri70c, Jur74h]. dimensions [Kin72]. dims [RF75]. diodes [All75c, Bat71, BCH^{+}79, KLE73]. DIP [Ano75-242]. direct [Ano79-145, Kap74b]. direct-writing [Ano79-145]. directing [Fal74a]. direction [Rud70, Tuc74]. directly [Ano74-229, Men75d, Nas78, Wei78a]. Director [Dre74]. directory [Tuc75a]. dirty [Mun78]. disabled [All76a, Jur76a]. disagreement [Mas79d]. disaster [TT74]. disasters [Fal76c]. discharge [All75c]. discipline [Men74g, Whi72]. disciplines [All74a]. Disconnecting [Chr75d]. discovery [GH71, VT72, Sus70b]. discrete [Rei74, Zuc77]. discrete-component [Zuc77]. discrete-event [Rei74]. discretes [Bea73b]. discrimination [Rub75b]. discuss [Ano76-270]. discussed [Gle77]. discussion [MRR^{+}77]. Dish [Wol76c]. disk [Ano79-195, BBF^{+}78, BCH^{+}79, Fet76, KLM^{+}79, RL74]. disk- [Fet76]. disks [Ber79, BCH^{+}79, Men75b, Men79b]. dislike [Jur74g]. dislocations [Rub74d]. dispatch [TE74]. dispel [IFSK77]. Dispelling [Mid74]. dispensing [SZ74]. displace [Wel79]. displacing [Fai74, TG78]. display [Dou76, Jur74f, RL74]. Displays [Kap74b, TG78, All76f, Ano72-524, Ano74-222, Fri70c, GM72, Jur72c, Jur75e, MNS77a, MNS77b]. displeasures [Rub74c]. Dispos [Wil76b]. Dispos-a-city [Wil76b]. disposable [Rub74a]. disposal [Ras75]. dissident [Tri74]. distance [EH76, Qua79]. distinct [DTW76]. Distributed [Ano78-47, Mok79b]. diverging [Rub77a]. diverse [EBC^{+}79, Lin76a, Rub75e, Tri74]. diversity [Rub74f]. Diverted [HH74]. division [Reu78]. DMM [Ano78-196]. DMM/calculator [Ano78-196]. do [Alb78, All77b, Ano72-318, Ano74-307, Ano78-244, Chr76l, Chr78i, Chr79o, Jur74c, McW77, Mun78, Rub76a]. Doctor [Dre74]. DOE [BC77]. does [Bra79, Gre79a, Jur77b, Rub76e]. doesn’t [RF75]. doing [Men77b, RO76, Rub76f]. dollar [Ele74a, Fri75g, Her79]. Dollars [Men74e, Rub76k, Gre79b, Her79, Kap77b, Rub76b]. domain [Ano74-43, Ele74d, Gib76, Lap72a]. dome [Lin76b]. domestic [Bar79]. domination [All75c]. done [All76f, Ano79-95, Jur75d, TDK74]. don’t [Chr74i, Fal74j, McW77, Pie79, Tuc74]. door [Eve77]. double [Roi78]. doughnut [Gib76]. dour [Rub76j]. down [All76e, All76f, Ano78-195, Fal74f, JM77, Lap74b, Lap74e, Max77, Rub75c, Spi76a]. downs [Chr78g]. drain [CT70, Sug78c]. Dramatic [KI77]. drilling [All79d]. drive [Ano79-195, BKS79]. driver [Jon77a, Jur75c]. Driving [Clo73, Fri72d]. drop [Sug78a]. drops [All75b]. Dry [Har72]. DTL [Gar70a]. dual [Ano79-193, Ano79-190, Men78d]. dual-band [Men78d]. dual-in-line [Ano79-193]. dual-trace [Ano79-190]. due [Doi77]. dues [Rub76b]. dumping [Rea79]. duplications [Ano74-294, Ano76-261, Ano76-262, Ano78-257]. duration [Par78a]. durations [EBC^{+}79]. duress [Ung79]. during [Cla77]. duty [TG78]. dwindling [Kap78b]. Dynamic [Kal75, And79].

each [Fal76b, Lin76a]. ear [Mac76]. Early [Shi70, Sus70a, Sus70b]. earpiece [Uts77]. Earth [Ele76b, Fal73c, Hor78, Wei77]. Earth-based [Wei77]. Earthquake [Fri73b]. ease [Fal74e, RO77]. easing [All74c, All76b]. East [Fri70f, Rub76i]. Eastern [Gre79c]. easy [BBG75, Men76a, Tor77a, TH77]. Ecap [BHLK71]. ECC [Ell72]. eccentric [Fri72l]. echo [GH70]. ECL [Gar70c]. ECM [Men78d]. Econometric [Sur72]. Economic [Fle71, TE74, CC78, HH74, Jay76, Kap76c, Mac78, Tor76d]. economics [Cou79, Rub77g, Wad76]. economies [Ano77-36]. economy [Ano79-95, BC77, HB77, Rub75e]. Eddy [Jur75d]. edged [Ano72-519]. Edison [BFM78, Der79]. Edisonian [Chr78b]. edited [Mas79e]. edition [Ano79-235]. Editor [Dre74]. Education [Ano74-96, Chr75e, Col78b, Col78a, Fal75f, Jen78, Jur74c, Men74c, Pat79b, Pat79a, Rub77a, Woo79, Fal74g]. Education/Energy [Jur74c]. Educational [Ano71-176, Ano71-177, Ano71-178, Ano71-179, Ano71-180, Ano72-240, Ano72-241, Ano72-242, Ano72-243, Ano72-244, Ano72-245, Ano72-246, Ano72-247, Ano72-248, Ano72-249, Ano72-250, Ano72-251, Ano73-230, Ano73-231, Ano73-232, Ano73-233, Ano73-234, Ano73-235, Ano73-236, Ano73-237, Ano73-238, Ano73-239, Ano73-240, Ano73-241, Ano74-108, Ano74-97, Ano74-98, Ano74-99, Ano74-100, Ano74-101, Ano74-102, Ano74-103, Ano74-104, Ano74-105, Ano74-106, Ano74-107, Ano75-105, Ano75-94, Ano75-95, Ano75-96, Ano75-97, Ano75-98, Ano75-99, Ano75-100, Ano75-101, Ano75-102, Ano75-103, Ano75-104, Ano76-84, Ano76-85, Ano76-86, Ano76-87, Ano76-88, Ano76-89, Ano76-90, Ano76-91, Ano75-153, Dun78, Kat75]. educator [Men74c]. EE [All74a, Ano72-518, Ano73-229, Ano77-80, DTW76, Fri75e, Gra74, Her72, Lin74a, Men77e, Pat79a, Pet77, Rub74c, Rub75a, Rub75b, Rub75d, Rub75f]. EEs [And73d, Ano72-252, Ano72-523, Ano77-81, Ano77-82, Ano77-83, Ano78-109, Ano78-110, Ano78-111, Ano78-112, Ano78-113, Ano78-114, Ano79-108, Ano79-109, Ano79-110, Ano79-111, Ano79-112, Ano79-113, Cor72, Fal74j, Jur73b, Jur75d, Mil72, RO76, Rub74a, Rub74b, Rub74d, Rub74f, Rub75c, Rub76d, Sal73]. effect [Ano74-40, BCY75, Bar79, Rub75e]. Effective [Rot71, Jon77b, Sal77]. effectiveness [Men78b]. effects [Ano75-43, Hol70, Jur74a, LH70, Sug78c]. effervescent [Tro74]. Efficiencies [Lev78]. efficiency [Fri74b, Fri74h, Fri75b, Fri76h, Kap79c, Men75a, TG78, War76]. efficient [All76g, CSS76, Yen78]. effort [Fri77c]. efforts [Men74e, Rub74b]. EHV [Jan70]. EIA [And72c]. eight [Ano75-243]. eight-bit [Ano75-243]. elapsed [Ano74-229]. Election [Ano76a]. Elections [Ano76a]. Electric [Ben70, Dal72, Gle77, Nag74, Der79, Esp77, Fri74e, Fri74i, Hea74, Kap75b, KLH74, LJ70, LC75, ME70, RG79, Rub75e, RGCS79, Sal72a, Sch77b, BDL^{+}78]. Electrical [Ano70-312, Ano70-313, Ano70-314, Ano70-315, Ano70-316, Ano70-317, Ano70-318, Ano70-319, Ano70-320, Ano70-321, Ano70-322, Ano70-323, Ano71-286, Ano71-287, Ano71-288, Ano71-289, Ano71-290, Ano71-291, Ano71-292, Ano71-293, Ano71-294, Ano71-295, Ano71-296, Ano72-375, Ano72-376, Ano72-377, Ano72-378, Ano72-379, Ano72-380, Ano72-381, Ano72-382, Ano72-383, Ano72-384, Ano72-385, Ano72-386, Ano73-365, Ano73-366, Ano73-367, Ano73-368, Ano73-369, Ano73-370, Ano73-371, Ano73-372, Ano73-373, Ano73-374, Ano73-375, Ano73-376, Lee71, Pla79b, Pla79a, BFM78, FC78, Fle71, Flo71, Fri75d, Rub74c, Rub77a, Soa70]. Electricity [Fri71a, GH71, Jay76, Ker76, SR78]. electrics [Wit77]. electrification [Nag74]. electrifying [Fri74i]. Electro [All76d, All77a, Ele78, Ano79-159]. electrodynamics [Fri75d]. electromagnetic [Ber71]. electromagnets [HL79]. electromechanical [Wel79, Mos70a]. electron [Bre71, Fal73b, Fet76, Shi72, Sim72a, Sus70b]. electron-beam-accessed [Fet76]. electron-emission [Sim72a]. electron/ion [Bre71]. Electronic [Ano77-36, Joh71, Jur75a, Man78, Men76e, Ols71, Shi76, Wal78, von79, All76a, All78d, Ano78-38, BHLK71, Ele73b, Fal75e, Fal76b, Fri74h, Jur75b, Kun75, Men77d, Rub79c, Spe78].

Electronica [Ano72-253]. Electronically [Ano70-375, HL79]. electronicker [Ano74-96]. Electronics [All78a, All79c, Ano70-218, Ano70-231, Ano70-312, Ano70-313, Ano70-314, Ano70-315, Ano70-316, Ano70-317, Ano70-318, Ano70-319, Ano70-320, Ano70-321, Ano70-322, Ano70-323, Ano71-286, Ano71-287, Ano71-288, Ano71-289, Ano71-290, Ano71-291, Ano71-292, Ano71-293, Ano71-294, Ano71-295, Ano71-296, Ano71b, Ano72b, Ano72-375, Ano72-376, Ano72-377, Ano72-378, Ano72-379, Ano72-380, Ano72-381, Ano72-382, Ano72-383, Ano72-384, Ano72-385, Ano72-386, Ano73-365, Ano73-366, Ano73-367, Ano73-368, Ano73-369, Ano73-370, Ano73-371, Ano73-372, Ano73-373, Ano73-374, Ano73-375, Ano73-376, Chr72, Chr78c, Ele73c, Fal76b, Fri74g, Jur76b, Jur77b, Jur78b, Jur79, Kap77c, Kap78c, Kap79a, Lap73d, Men76b, ME77, Qua79, SR78, All76a, All76h, Ano72-389, Ano73-377, Ano78-36, Ano79-160, Ber79, Cam77, Chr75n, Clo73, Fin72b, Fle71, IFSK77, Jon77b, Jur74b, Jur74h, Jur75b, Jur75c]. electronics [Jur75d, Jur77a, Jur78a, Kap75c, Kap76d, KI77, Men73b, Men75b, Men76e, Men76f, Men77b, Men78d, Men79b, NR74, PM77, PMG77, Rea79, Rub77g, SOW77, Sur70, Sus70a, Sus70b]. Electrooptic [LH70]. Electrooptics [Wal78]. electroquackery [SR78]. Electrotechnical [Gle77]. Electrotechnology [She78, Sug78d, Tor77c]. Elementary [Ano74-42]. elements [AFCK79, TT74]. Eleven [Bea74b]. elimination [Fri71a]. elite [Ano77-29]. else [Ano79-232]. Embedded [Rub74e]. emerge [BBF^{+}78]. emerged [Wei73]. emergency [FC78, RJD73]. Emeritus [Dre74, Pie79]. eminent [Pie79]. emission [Sim72a]. emissions [Fri70f, HB77, NS77]. emitters [Jur74f]. emitting [All75c]. emotion [Chr76f]. Emphasis [All78b]. emphasizes [Ano74-95]. employed [Par78b]. employee [Rub74a]. employees [Ano79-235, Fri74d]. employers [Ano79-235]. Employment [Ano71-241, Ano73-242, Ano79-235, Gra70a, Kin71]. emptor [Rub77d]. emulates [Ano78-199]. emulation [All78c]. enable [BCH^{+}79]. encoded [Ber79]. encore [Ano78-165]. encourage [Bar78, Esp77, Rab78, Tor77e]. Encyclopaedia [Ano78-164]. end [Ano75-231, Bea75, Eve77, Mas79f, Sug78c]. end-users [Ano75-231]. endemic [Ano77-59]. Energy [Ano74-109, Ano74-110, Ano74-111, Ano74-112, Ano74-113, Ano74-114, Ano74-115, Ano74-116, Ano74-117, Ano74-118, Ano74-119, Ano74-120, Ano75-106, Ano75-107, Ano75-108, Ano75-109, Ano75-110, Ano75-111, Ano75-112, Ano75-113, Ano75-114, Ano75-115, Ano75-116, Ano75-117, Ano76-92, Ano76-93, Ano76-94, Ano76-95, Ano76-96, Ano76-97, Ano76-98, Ano76-99, Ano76-100, Ano76-101, Ano76-102, Ano76-103, Ano77-84, Ano77-85, Ano77-86, Ano77-87, Ano77-88, Ano77-89, Ano77-90, Ano77-91, Ano77-92, Ano77-93, Ano77-94, Ano77-95, Ano78-115, Ano78-116, Ano78-117, Ano78-118, Ano78-119, Ano78-120, Ano78-121, Ano78-122, Ano78-123, Ano78-124, Ano78-125, Ano78-126, Ano79-114, Ano79-115, Ano79-116, Ano79-117, Ano79-118, Ano79-119, Ano79-120, Ano79-121, Ano79-122, Ano79-123, Ano79-124, Ano79-125, Bro79, CSS76, Chr74c, Chr74d, Chr74e, Chr74f, Fel73, Fri73d, Fri73c, Fri74b, FR75, Fri75f, Fri76h, HH74, Jur74d]. Energy [Jur74e, Jur74c, Kap76a, Kap76b, Kap76c, Kay78, Lin75a, Nag74, Rub75e, All78a, AFCK79, BDL^{+}78, BFM78, Bea75, BJ78, Bro73, Col79, Cou79, CC78, Dar70, Ele74e, Fri73f, Fri73h, Fri74c, Fri75b, Fri76f, Fri76g, Fri77b, Fri77c, GHNK79, Gle77, Hea74, Hus74, Joh79, JF74, Kap77b, Kap77c, Kap77e, Kap78a, Kap78b, Kap79c, KLH74, LC75, Lap74b, Lia78, Mic78, Nil76, Okr78, Pla79b, PMG77, RK79, RG79, Rub74e, Rub76i, Rub78b, Rub79c, RGCS79, Sim72b, Sug78c, Esp77, GHNK79]. energy-storage [Fri76g]. energy-supply [Fri76f]. Energy/power [Fri74b]. enforcement [Ele73c, Men78d]. enforcement-minded [Men78d]. Engine [Wit77, Jur75a, NS77]. Engineer [McK71, Men74c, Tuc75a, All76g, Ano76-270, Ano77-96, Ano77-97, Ano77-98, Ano77-99, Ano77-100, Ano77-101, Ano78-127, Ano78-128, Ano78-129, Ano78-130, Ano78-131, Ano78-132, Ano79-126, Ano79-127, Ano79-128, Ano79-129, Ano79-130, Ano79-131, Bea74b, Chr78i, Eve77, Gra74, Her73, Jor70, Pon70, RO77, Rub78b, Sim71, Tri78, Tuc74, Wol77]. Engineering [All76a, Ano72b, Ano72-454, Bar73, Chr76f, Col78b, Fal73a, Fin77, Men75d, Ung79, Wil70, BB71, Der79, Ele74b, Flo71, Han79, Jur74g, Kat75, Kin71, Man70, NC71, Rie70a, Rie70b, Rie70c, Rub74f, Rub76f, Rub77a, TH77, Whi72, Woo79, Gre79a]. Engineers [Ano70-312, Ano70-313, Ano70-314, Ano70-315, Ano70-316, Ano70-317, Ano70-318, Ano70-319, Ano70-320, Ano70-321, Ano70-322, Ano70-323, Ano71-286, Ano71-287, Ano71-288, Ano71-289, Ano71-290, Ano71-291, Ano71-292, Ano71-293, Ano71-294, Ano71-295, Ano71-296, Ano72-375, Ano72-376, Ano72-377, Ano72-378, Ano72-379, Ano72-380, Ano72-381, Ano72-382, Ano72-383, Ano72-384, Ano72-385, Ano72-386, Ano73-365, Ano73-366, Ano73-367, Ano73-368, Ano73-369, Ano73-370, Ano73-371, Ano73-372, Ano73-373, Ano73-374, Ano73-375, Ano73-376, Eve77, Far73, Kap77e, Sco73, Wal74, BCY75, Chr79c, Cou79, Fri74d, GH72, Jen78, Kap77d, Rub74c, Rub74d, Ste77]. engines [Jur75b, Wit77]. English [Bec71, Que79]. English-language [Que79]. enhance [Kau72, SL78]. Enhanced [KS79, Cam77]. Enhancing [HGB79]. ENIAC [Mau75]. Enormous [Joh79]. enough [Bea75, McK73]. enrollments [Men74c]. enter [BJ78]. enters [Jur72a, Pog78, Tri78]. entertain [Men76e]. enticed [MNS77b]. entrepreneur [Jor70, Men78c]. entries [All79g]. entry [Ano74-231]. environment [And73d, Flo71, Fri70e, Fri70f, Gal77, Spo71, Wol77]. Environmental [And72b, CC78, Fri70f]. environments [Wil70]. EO [Wal78]. EPA [Sto77]. epidemic [Men76c]. epitome [McK72b]. EPRI [BDL^{+}78, Lin75a]. equals [Bel74]. equation [BJ78]. equations [Rec75]. Equipment [Pat79a, All75d, All77d, All78c, Ano78-38, Ele74d, Ele76a, Fal75e, Fal76b, GHM71, IY77, Jur74g, Lap74b, Men74a, Mor78, Oli74, Wel79]. era [All76h, All79c, Ana79, Jur72a]. erode [Rub75e]. erratic [Cou79]. error [Ano76-206, Chi73]. errors [Ber71]. Esclangon [Pol70]. especially [Rub76h]. ESS [Joe76, Kun75]. Essay [Mad74, Wal74]. establishing [Rac78]. estimation [Sor70]. Eta [Mas79d]. ethic [Kop77]. ethical [Ung79]. ethics [Ano76-271, Fri74d, Ung79, Ste75b]. Europe [Fri74i, Fri74j, Fri75c, GK76, Gre79c, Kap76b, Mid79, Rip79, Sch77b]. European [Fal74h, Mac78, ME77]. evaluate [New76]. Evaluating [Alb78, Rub76k]. even [Jur74e, Tor76c]. event [Rei74]. events [Tuc74, WZ78]. ever [Ano76-273, Mas79f, Mau75]. every [All74a]. everybody [Men77b]. everywhere [Lap73f]. evidence [Mid71]. evolved [Sal77]. EVR [Gol70, GR76]. EWS [Kun75]. ex [Fri74d]. ex-employees [Fri74d]. examination [Ano74-222, Ewa78, Kra75]. example [Dou76, War76]. examples [Nas78]. excel [Fet76]. excels [Kap77c]. except [Fra78]. excerpted [Mas79e]. excessive [Nag74]. exchange [Tro74, TH77]. excitation [DKR^{+}72]. exciting [All74a, Jur74c, VHM77]. exclusive [Fri76e, Rei74]. executive [Pie79]. exercise [Fri75f]. Exhibition [Ano70-297, All74a, Rub77g]. exhibitor [Ano70a]. Exhibitors [Ano71-181, Ano72-254, Ano72-390, All77a, Ano70-297]. Exhibits [Ano70-455]. existing [Lus78]. exists [Kap77b]. expand [Fal77b]. Expanding [Jur74c, Ele76a]. expands [Sug79b]. expansion [RF75]. expectations [Ern77b, Fal74c]. expected [All77a, EBC^{+}79]. expects [Jur74e]. expenditures [Jur74d]. expense [Her79, Nil76]. expenses [RO76]. Experience [Con70, Chr79e, MRR^{+}77, Pla79b, Pla79a]. experienced [Cot75]. experiment [Col78a]. experimental [RK79]. experimenters [Wei77]. experiments [All76h, Esp77, KI77, MF79]. expert [All75i, Fri75f, Pet77]. Experts [All79c, Col78c, Oll78, RO77, All76i, All79h, Ano71-297, Ano71-298, BCY75, Chr79o, Ele74f, Rub79a, Spi76b, Sug79a]. explained [Ano78-39, Ano79-45, RO76]. explodes [Men77c]. Exploding [Buc78a, Jur74d]. exploit [CSS76, Col79, Kap77a]. exploited [All78c, Hor76, Mor78, SR78]. Exploiting [Cou79, Tor77e]. exploits [All77b, Fal76b, MVE75]. explosion [Men76c, Tea74]. exponential [Edd70]. exported [Fri75c]. exposure [Gol79]. express [Rub77a]. expression [Ele74f]. extending [Rac78]. extent [PHC76]. external [Ano75-243, Wit77]. extortion [Rub76b]. Extra [Ano72-491]. extraterrestrial [NR74]. eye [Hor76].

F [Dou76]. F-14 [Dou76]. FAA [Fri70a, Kin75]. fabricated [HGB79]. fabrication [Hei79, Joh79, Tor79]. face [Lin77, MRR^{+}77, TG78]. faces [Ano75-240]. facilitate [Kap79d]. facilities [Fri76f]. Fact [Chr76g, EW78, Sal72a, RF75]. factor [All77d]. Factors [MNS77a, Mac78, Ele74c, Ewa78, Fri74b, Fri77a, Ger74, Jur74g, Max77, Sug79c]. factory [All79d, Oll78]. facts [Jur74e]. fail [Alb78]. failed [Rub77b]. failure [All76i, Ewa78, Jur76f]. failures [Gan77]. Fall [Ano70-247, Ano74-209, Ano75-217]. fallacious [Jur74d]. Falls [Fri71b]. false [Pay75]. famed [Fal76c]. families [EBC^{+}79, Gar70a, Gar70b, Gar70c]. family [Ano74-226, Ano74-223, Ano76-204, Ano78-195, Ele72c, Gar70a, Lap73g]. famous [Ano75-240]. FAMs [Lap73e]. fancy [Men77c, Sal72a]. fantasy [EW78]. far [Ago76, Ano79-188, Rac78]. far-out [Ago76]. fare [Ele78]. Fast [Mit77, Fal75c, Fri74b, Fri74f, Opp70]. fast-breeder [Fri74b]. fast-moving [Fal75c]. faster [All77b, Fet73, Fet76, Gar77, Hor72, Lap74a, Yu75]. fastest [All79g]. fatal [Bra77]. fate [Fal76c, Tor76d]. Father [Fri75d, McK74]. fault [Dal70, Fal76a]. fault-tolerant [Fal76a]. faults [All76i]. favor [Pit79]. favorable [Wol77]. favorite [Cam77]. Fax [Kap74a, Cos74]. FCC [Hin79, Men77a]. FEA [Fri76h]. feasible [Rob74a]. feature [Ano78-191, Ano78-193, Fal77c, Gos79b]. features [AC75, All74c, Ano74-225, Ano75-242, Ano78-195, Ano79-192, Ano79-195, Car79, Gar70a, KLM^{+}79]. Federal [Fri76h, Jan70]. Feedback [Rub77b, Bla77]. Feedforward [Jur72b]. Feeding [SESJ76]. feeling [Chr76g]. Fees [Ano72-316]. Fellows [Ano77-29]. Ferrites [Sne72a, Sne72b]. ferroelectric [Hol70, LH70]. ferrofluids [Mos75]. Ferry [Bla77, Kap76e]. FET [Ano74-223, Ano78-197]. FET-IC [Ano74-223]. FET-input [Ano78-197]. fever [Men76c]. few [Buc78a, Pay75, Spi76a]. fewer [Fri75f]. fi [Ber79]. Fiber [Fag79, Fal75c, Men75a, Ano77-59, Men77a, Men79a]. Fiber-optics [Fag79]. fibers [Men78b, Oli74]. fidelity [Ber79]. Field [Ano79-47, All78b, All79e, Ano74-40, Bea72d]. fifty [Ele74e]. fifty-fold [Ele74e]. fight [Ano79-51]. fighter [Dou76]. fights [Sto77]. file [Fet76]. files [Men75d]. filling [Fal75b]. fills [HH74]. film [KLM^{+}79]. filters [Ano75-38, Ano78-192, Gef70, HGB79, Mos70a, Mos70b, Oli74]. finance [Lom79, RF75]. Finances [Her79, Rub76g]. Financial [Rub76b, Rub76c, Rub75e]. Financing [Tan72, Jor70]. find [All75c, Cam77, Fal74b, Gup74, Lin74b, MM74, Tuc75b]. finds [Mos75]. fine [IFSK77, War76]. fine-tuned [IFSK77]. finely [Gre79d]. finest [Fri74i, Fri74j]. fingerprint [Ele73a]. Fire [All75f]. fires [Chr78l, Mas79d]. firms [Mac78]. First [GHNK79, Ano78-165, Joh71, Men73a, Moo79, PG76, Rub74c, Spi76a, Spi76b, Tor77b, War76]. first-hand [Rub74c]. firsts [RS77]. fits [Ano78-47]. Five [MRR^{+}77, All75i, Ano78-194, Ano78-196, Max77, Men78a, Rub76d, Rub76f, Yu75]. fix [KH79, Pie79]. fixed [Ano79-195]. Fixing [Fri75g]. flash [Bla77]. Flat [TG78, Men79b]. Flat-panel [TG78]. Flexibility [Lou74, New76, Pit79]. flight [Fri70c]. flight-control [Fri70c]. Flights [Ano71a]. Floating [Fri73e, RGCS79]. Floor [Ano72-255]. floppy [RL74]. Florida [Col79]. flow [Fal74a]. flows [Dar70]. fluctuations [Hua70]. fly [And73f, Fal74h]. Flying [Tho73b]. FM [PM77]. Focal [Ano70-248, Ano70-249, Ano70-250, Ano70-251, Ano70-252, Ano70-253, Ano70-254, Ano70-255, Ano70-256, Ano70-257, Ano70-258, Ano70-259, Ano71-182, Ano71-183, Ano71-184, Ano71-185, Ano71-186, Ano71-187, Ano71-188, Ano71-189, Ano71-190, Ano71-191, Ano71-192, Ano71-193, Ano72-256, Ano72-257, Ano72-258, Ano72-259, Ano72-260, Ano72-261, Ano72-262, Ano72-263, Ano72-264, Ano72-265, Ano72-266, Ano72-267, Ano73-243, Ano73-244, Ano73-245, Ano73-246, Ano73-247, Ano73-248, Ano73-249, Ano73-250, Ano73-251, Ano73-252, Ano73-253, Ano73-254, Ano74-121, Ano74-122, Ano74-123, Ano74-124, Ano74-125, Ano74-126, Ano74-127, Ano74-128, Ano74-129, Ano74-130, Ano74-131, Ano74-132, Ano75-118, Ano75-119, Ano75-120, Ano75-121, Ano75-122, Ano75-123, Ano75-124, Ano75-125, Ano75-126, Ano75-127, Ano76-104, Ano76-105, Ano76-106, Ano76-107, Ano76-108, Ano76-109, Ano76-110]. Focus [Ste77, Gol73]. Focused [Reu78]. focusing [Wei78a]. foes [Kap77a]. foiling [Sug79a]. fold [Ele74e]. following [Ano74-71]. fools [LC75]. foot [Tor78a]. football [Men77d]. force [Gre74]. forced [Pla79b]. Ford [Fri72e, Ano75-304, Ano76-270, Jur78a, Tor77b]. Forecast [Bea72a]. forecasting [All75h, Ano75-319, Fal75c, Fal75d, Fri75f, Fri77a, Jur75e, Koh71, Lin74b, Lin75b, Wil76a]. forecasting-tv [Wil76a]. forecasts [Ano75-319, Lin74b, Rub74b]. forerunner [NS77]. form [Bea74b, Dav79]. formal [Lin74b]. formats [Cam77, Kle77, Men78e, Roi78]. former [Hin79]. forms [Nil76]. Formula [Tor77g, Buc78a]. formulate [Ang74]. fortune [Ker76]. Forum [Ano70-260, Ano70-261, Ano70-262, Ano70-263, Ano70-264, Ano70-265, Ano70-266, Ano70-267, Ano70-268, Ano70-269, Ano70-270, Ano70-271, Ano71-194, Ano71-195, Ano71-196, Ano71-197, Ano71-198, Ano71-199, Ano71-200, Ano71-201, Ano71-202, Ano71-203, Ano71-204, Ano71-205, Ano72-268, Ano72-269, Ano72-270, Ano72-271, Ano72-272, Ano72-273, Ano72-274, Ano72-275, Ano72-276, Ano72-277, Ano72-278, Ano72-279, Ano73-255, Ano73-256, Ano73-257, Ano73-258, Ano73-259, Ano73-260, Ano73-261, Ano73-262, Ano73-263, Ano73-264, Ano73-265, Ano73-266, Ano74-133, Ano74-134, Ano74-135, Ano74-136, Ano74-137, Ano74-138, Ano74-139, Ano74-140, Ano74-141, Ano74-142, Ano74-143, Ano74-144, Ano75-128, Ano75-129, Ano75-130, Ano75-131, Ano75-132, Ano75-133, Ano75-134, Ano75-135, Ano75-136, Ano75-137, Ano76-111, Ano76-112, Ano76-113, Ano76-114, Ano76-115, Ano76-116, Ano76-117, Ano76-118, Ano76-119, Ano76-120, Ano76-121, Ano76-122, Ano77-102, Ano77-103, Ano77-104, Ano77-105, Ano77-106, Ano77-107, Ano77-108, Ano77-109, Ano78-133, Ano78-134]. Forum [Ano78-135, Ano78-136, Ano78-137, Ano78-138, Ano78-139, Ano78-140, Ano78-141, Ano78-142, Ano78-143, Ano78-144, Ano79-132, Ano79-133, Ano79-134, Ano79-135, Ano79-136, Ano79-137, Ano79-138, Ano79-139, Ano79-140, Ano79-141, Ano79-142, Ano79-143, BBW^{+}77, BLD^{+}77, Fan75, HS77, Lap73e, RPR^{+}77, Wil75]. forward [RL74]. fossil [Jay76]. fossil-fuel [Jay76]. found [Kap76e, Mas79d, Rub76d]. Founded [Mas79d]. Founder [Rub76f]. Founders [Rub76g]. Four [Ano78-195, Col79, DTW76, Ele74e, Mas79c]. Four-digit [Ano78-195]. Fourier [Opp70]. France [Fri75c, Kap77e]. Francisco [Fri74f]. Franklin [Okr78]. fraught [Rub77d]. free [Ele74f]. Freedom [Gol74, Rob74b, Mos74]. Freeman [Mic78]. French [Fri75c, Kap76b]. frequencies [Fal74b, Oli74, Reu78].

Frequency [Ano70-326, Ano74-39, Ano74-43, CF72, Ewa78, Fri77a, Tor76a]. frequently [Ano77-179, Lin74b]. freshman [BB71]. Friends [Kap77a, Dun78]. frills [Ano74-95]. fringe [Ano72-252]. Front [Ano70-272, Ano70-273, Ano70-274, Ano70-275, Ano70-276, Ano70-277, Ano70-278, Ano70-279, Ano70-280, Ano70-281, Ano70-282, Ano70-283, Ano70-284, Ano70-285, Ano70-286, Ano70-287, Ano70-288, Ano70-289, Ano70-290, Ano70-291, Ano70-292, Ano70-293, Ano70-294, Ano70-295, Ano71-206, Ano71-207, Ano71-208, Ano71-209, Ano71-210, Ano71-211, Ano71-212, Ano71-213, Ano71-214, Ano71-215, Ano71-216, Ano71-217, Ano71-218, Ano71-219, Ano71-220, Ano71-221, Ano71-222, Ano71-223, Ano71-224, Ano71-225, Ano71-226, Ano71-227, Ano71-228, Ano72-280, Ano72-281, Ano72-282, Ano72-283, Ano72-284, Ano72-285, Ano72-286, Ano72-287, Ano72-288, Ano72-289, Ano72-290, Ano72-291, Ano72-292, Ano72-293, Ano72-294, Ano72-295, Ano72-296, Ano72-297, Ano72-298, Ano72-299, Ano72-300, Ano72-301, Ano72-302, Ano72-303, Ano73-267, Ano73-268, Ano73-269, Ano73-270, Ano73-271, Ano73-272, Ano73-273, Ano73-274, Ano73-275, Ano73-276, Ano73-277, Ano73-278, Ano73-279, Ano73-280, Ano73-281, Ano73-282, Ano73-283]. Front [Ano73-284, Ano73-285, Ano73-286, Ano73-287, Ano73-288, Ano73-289, Ano73-290, Ano74-145, Ano75-138, Ano75-139, Ano76-123, Ano76-124, Ano77-110, Ano77-111, Ano77-112, Ano77-113, Ano77-114, Ano77-115, Ano77-116, Ano71-273, Ano74-195, Ano75-203, Ano75-204, Ano77-136, Ano77-137, Ano77-138, Ano77-139, Ano77-140, Ano77-141, Ano77-142, Rub79c, SA78, Tor76c, UG73]. frugal [Dun78]. fuel [BFM78, BC77, HB77, Hus74, Jay76, Jur75a, Jur77a, Kap75a, Kap77b]. fuel-cell [BFM78]. fueled [Fri74c, Rub76k]. fuels [Fri72e, Ker76]. Fujitsu [Ano74-307]. full [KH79, Oll78, Ura77]. full-blown [Oll78]. fully [Men74f]. Fun [Men76f, And73e, Jur72e, Men77c]. function [All75b, Lap74a, Men76f]. Functional [All77e]. functions [Lia78]. fundamental [Moo79, Ano74-96]. Fundamentals [Ano75-40, Rie70a, Ano74-41]. funded [All76h]. funding [Ano75-304, Ele74e]. funds [Ano77-36, Dil75]. Further [Rub75a, Ano78-241]. fuse [Ano74-229, Mid74]. fuse-breaker [Mid74]. Fuses [Mid74]. fusion [Kap76b, Mil71, Ste77]. Future [Ano71-229, Ano71-230, Ano71-231, Ano71-232, Ano71-233, Ano71-234, Ano71-235, Ano71-236, Ano71-237, Ano71-238, Ano71-239, Ano71-240, Ano72-304, Ano72-305, Ano72-306, Ano72-307, Ano72-308, Ano72-309, Ano72-310, Ano72-311, Ano72-312, Ano72-313, Ano72-314, Ano72-315, Ano73-291, Ano73-292, Ano73-293, Ano73-294, Ano73-295, Ano73-296, Ano73-297, Ano73-298, Ano73-299, Ano73-300, Ano73-301, Ano73-302, Ano74-146, Ano74-147, Ano74-148, Ano74-149, Ano74-150, Ano74-151, Ano74-152, Ano74-153, Ano74-154, Ano74-155, Ano74-156, Ano75-140, Ano75-141, Ano75-142, Ano75-143, Ano75-144, Ano75-145, Ano75-146, Ano75-147, Ano75-148, Ano75-149, Ano75-150, Ano75-151, Ano76-125, Ano76-126, Ano76-127, Ano76-128, Ano76-129, Ano76-130, Ano76-131, Ano76-132, Ano76-133, Ano78-145, Ano78-146, All76b, All79h, Ano73-473, Ano74-181, Ano74-182, Ano74-183, Ano74-184, Ano74-185, Ano74-186, Ano74-187, Ano74-188, Ano74-189, Ano74-190, Ano74-191, Ano74-192]. future [Ano74-180, Ano75-187, Ano75-188, Ano75-189, Ano75-190, Ano75-191, Ano75-192, Ano75-193, Ano75-194, Ano75-195, Ano75-196, Ano75-197, Ano75-198, Ano75-182, Ano75-183, Ano75-184, Ano75-185, Ano75-186, Ano76-167, Ano76-168, Ano76-161, Ano76-162, Ano76-163, Ano76-164, Ano76-165, Ano76-166, Ano77-119, Ano77-120, Ano77-121, Ano77-122, Ano77-123, Ano78-156, Ano78-157, Ano78-154, Ano78-155, Ano79-156, Ano79-157, Bar79, Bea73b, Buc78a, Chr77d, Cot75, Ele74e, Fal75c, Fal75d, Fin77, Fri70a, Fri72e, Fri75f, Fri76a, Fri76e, Fri77a, Gra74, Kap75b, Kap76a, Kap76c, KI77, ME70, Mok79c, NKL72, Oli74, Rub79b, Rud70, Spo71, Tan72, Tel73, Wil76b, Yu70]. futures [Rub76g]. Futurism [And73e].

gages [Ano78-40]. gain [War76]. gained [Doi77]. gains [Ano78-94, Fri75b, Wal78]. Gallium [VL77]. galore [All76c, Ano72-520, Jur74b, Kap78c]. Galvani [GH71]. game [Ano78-242, Fal75a, Fri75f, Rub77g]. games [Men76e, Men77c, Men77d, Men78f]. gamesmanship [Men76e]. gap [Chr75k]. gaps [Fal75b]. Garriott [Lin74a]. gas [All75c, BKS79, EBC^{+}79, Nag74]. gas-discharge [All75c]. Gasdynamic [Ger70]. gasoline [BKS79, Ker76, Tor77h]. Gate [All77h, Ano79-160]. gates [Ano79-145, Fai74, Gos79b, Hei79]. Gauss [Sor70]. gave [Bea75]. GE [Rub78b]. gear [All76g, All77b, All78b, Ber79]. General [Ano72-316, All75i, All77c]. generated [Jay76]. generating [Ben70, Lev78, Yu75]. generation [All75h, Ano75-303, Bea71, Gre72, Jur74h, Kal79, Kap77d, Kap79c, Okr78, Qua79]. generators [Jur75e, Tuc75b]. generic [Sug79c]. genesis [Wol76b]. genius [Fri72l, Jur75d]. geographical [Gor75]. Geographically [Bea74a]. geometries [Gos79b]. Geoscience [Ano71b]. German [Fri74h, Fri74j]. Germany [Fri74h, Fri74j, Jan70, Kun75]. Get [MM74, All75a, Doi77, IKT73, Jur74g, Kap75c]. gets [Chr74k, Chr74l, Chr74m, Fal74e, GK76, Tho73a]. Getting [Lap74e, Man70, SS73, Wad76, Gar77, Jur74h, Jur75d, Men74b]. GHz [Ano70-296, Ano79-188]. giant [BFM78, Fri74e, Sug78b]. giants [Men79a]. gigabit [Kap79d]. gigaflop [Fal76c]. Gigahertz [Ano70-296]. give [All76c, Ano74-209, Ano75-217, Ano75-237]. given [Par78b]. gives [Fri76e]. gladiator [Tri78]. glamor [SR78]. glance [Ano72-387]. glass [Tuc75b]. Global [Fri74c, KH79, Gle77, PW71]. glutton [SESJ76]. gluttons [Kap76a]. GM [Jur78a]. go [Chr76l, Chr79o, Fri75h, Gre74, Hea74, JF74, Jur75b, Kap77e, Lom79, Men79a, Pit79, Rac78, SL78, Spi76b]. goals [Gol73]. gobbler [Kap77b]. goes [Chr73, Gal77, Kap75c, Men76f, Rub79c]. going [Fri72f, Rub77c, SOW77]. Golden [All77h, Ano79-160]. Goldsmith [Dre74]. Good [Hor72, And72c, Fal74b, Fri77d, HB77, Ker76, Lou74, Ste75a]. goods [SZ74]. got [Ano76-273]. Gov. [Fri75e]. Government [Ano78-244, Ele74f, Nas78, PHC76, Rub76d, All76h, Fri70f, IY77, PMG77, Tri74, Rub76a]. government- [All76h]. grade [Ano77-29, Ano79-190]. grand [Ano72-317, Ber73]. granfalloon [Wes73]. graphic [All77f]. graphic-arts [All77f]. Graphics [MNS77a, MNS77b]. gray [TG78]. Great [Fal74c, Fri76c, Han79, Rub74f, All75f]. greater [MNS77b, Men78g, Par78b, Pit79]. greatness [McK74]. Green [Ish77, Kal79, Obe77]. greet [Kat75]. ground [AFCK79, Bro79, Dal70, Fri70a, Fri70b, Gib70, Gib76, GK76]. ground-based [Bro79]. ground-fault [Dal70]. ground-instrument-landing [Fri70b]. grounded [Men74e]. group [Wad76, Ano71a]. groups [Rub77f, Ano71-119]. grow [Har77]. growing [All79d, Rub77d, Ste75a]. grows [All75d, Jur74b, Tor77a]. growth [AB70, Buc72]. guidance [NR74]. Guide [Ano74-157, Ano75-152, Ano76-134, Ano77-117, Ano78-147, Ano79-144, Ano70-169, Ano71-120, Ano72-319, Ano73-305, Kap77f, Rub76e, Tuc75b, Tuc75a]. guided [Men78g]. guideline [Ano79-235]. Guidelines [GHNK79, Ano73-242, Ano79-235, Lin74b, Ste75b]. gymnasium [Col79].

habits [Jen78]. hails [Rub74f]. halted [Fri76j]. ham [Gor75]. Hamburg [ME70]. hammer [HH74]. Hampshire [Col79]. Hand [Ano78-196, Men74b, Men75d, Rub74c]. Hand-held [Ano78-196, Men74b]. hand-made [Men75d]. handful [Tro74]. handicapped [Tri78]. hands [Kal79]. hands-off [Kal79]. handwritten [Ano74-231]. Hansen [Gin75]. happen [Fal74j, Str71]. happened [Ano76-273, Rub79a, Wol76a, Wol79]. happening [All74d]. happy [War76]. Hard [Fai74, Fal74b, Men74d]. Hard-soft [Fai74]. hardened [Men78g]. Hardware [All76b, Ano77-118, Ano77-45, Chr75f, Fal75b, Fal77b, Par78a, Ano74-230, Ano74-62, Ano74-63, Ano74-64, Ano74-65, Ano74-66, Ano74-67, Ano74-68, Ano74-69, Ano74-70, Ano74-71, Ano75-58, Ano75-59, Ano75-60, Ano75-61, Ano75-62, Ano75-63, Ano75-64, Ano75-65, Ano75-66, Ano76-51, Ano76-52, Ano76-53, Ano76-54, Ano76-55, Ano76-56, Ano76-57, Ano76-58, Ano76-59, Ano77-44, Ano77-46, Ano77-47, Ano77-48, Ano77-49, Ano77-50, Ano77-51, Ano77-52, Ano77-53, Ano77-54, Ano78-63, Ano78-64, Ano78-65, Ano78-66, Ano78-67, Ano78-68, Ano78-69, Ano78-70, Ano78-71, Ano78-72, Ano78-73, Ano79-45, Ano79-67, Ele76a, Fal74e, Fri72b, Fri72h, Jur76e, Kap75c, MNS77b, Men77a, Mok79b, Sug78a]. Hardware/software [Fal75b, Kap75c]. Hare [Buc78b]. Harmonic [SK73]. Harnessing [Yen78]. haul [Hea74]. having [Doi77, Fag78]. Hawthorne [Ano75-318]. hazard [Dal70, Dal72]. hazards [Ele73d, Fri71a, Gol79, Jur77a, Kin75]. hazy [Fri75f]. head [Dou76, Fri76i]. head-up [Dou76]. headache [Bea75]. Headquarters [Rub76e]. Health [Jur73b, Her72]. heart [All76b, Ano71-421, Jur76f, Kun75]. heat [All79d, Fal74f, RGCS79]. heat-treating [All79d]. heaters [Pla79a]. heating [Ano78-46]. Heavy [Jen78, Joh79]. heavy-lift [Joh79]. held [Ano78-165, Ano78-196, Men74b]. helix [Roi78]. Hellman [Bra79, Dav79, Tuc79]. help [Fal74j, Gin75, Gra70a, KI77, Mas79f, Men78c, Rub74d, Rub76e, Tuc74, Tuc75a, Ung79]. helped [Bla77]. Helping [Men78d]. helps [Kap77d, SK73, Tib78, Zuc77]. Henry [Tor77b]. herald [Ana79]. Here [RL74, Chr76l, Chr79o, All77e, Lin74b, Men77e, Rub76a]. herewith [Pat79b]. Heterodyne [Ano70-296]. Hew [Ano79-145]. hierarchical [SL78, Sch78]. High [AC75, Ano70-296, Ano78-197, Bre75, BCH^{+}79, Chr78j, DTW76, EH76, HGB79, Tor76a, Uts77, All74b, And79, Ano74-95, Ano74-234, Ano75-231, Ano75-236, Ano75-235, Ano79-145, Ano79-160, Ber79, Boa73, Fri77b, Gef70, Gib76, Lin75a, Okr78, Sal77, War76, Wel79]. High-density [Bre75, HGB79, Sal77]. High-frequency [Tor76a]. High-power [BCH^{+}79]. High-resolution [Uts77]. High-Sensitivity [Ano70-296]. High-speed [Ano78-197, All74b, Gib76]. high-technology [Ano79-160].

High-voltage [EH76, Ano75-236]. high-volume [Ano75-231]. higher [All75h, Ano79-159, Lap74a, Men78g, Pit79, Reu78, Ber79]. highest [Ano79-188]. highly [Mas79c, Sal77]. highs [Pla79b]. hijackers [Ano79-51]. Hill [Gre79a]. him [Jur74e, Rub76a, Tri78]. himself [Rub75a]. hinge [Fal74c, Wad76]. historic [Fri74d]. history [Dou76, Fin73, Fri70d, Fri72g, Ger74, Lap73c, Lap73g, RL74, Shi79a, Soa70, Sus70a, Sus70b, Pol71a]. Hitherto [All78d]. hitters [Jen78]. hobbyist [RM75]. hold [Nay71b]. holding [Alb78, Lap74b]. holds [Ele75b, Gre79b]. holiday [Men77d]. home [Car79, Cos74, Men77b, Men77c, Mid74, RO76, Ano78-164]. honed [Gre79d]. honorary [Gre79a]. hopes [Chr78l, Okr78]. horizon [Ano77-59]. horse [Gre74]. hospitals [Fri71a]. host [All79d, Ano78-195, Mos75]. hostile [Gal77]. hot [Okr78]. hot-air [Okr78]. hours [Fri74f, Rub75f]. house [Fal74f]. household [Car79]. hows [MT73]. HTGR [Fri74b]. HTL [Gar70a]. Human [Ger74, Ele74c, Han79, Jur74g, Lin76a, Men76e, RRS^{+}74, Spe78, Sug78b, Sug79c]. humanitarian [Dun78]. humankind [Fri76e]. humans [Mun78]. hunches [Fey74]. hundreds [Bea74b, BC77]. hunting [Bea73a]. hurdles [Kop77, TG78]. HVDC [EH76, Eng75, KG73, KLH74]. Hybrid [Ano75-235, Ano78-37, Mid74, Sch77b, Zuc77]. hydroelectric [KLH74].

I-squared [War76]. I/Design [Kap75e]. I/O [Ano75-243, Ano79-192]. I/Survey [Rub75f]. I2L [Tor77e, All75a, Gar77, Kle77, Ray77, Sac77, Ver77]. IC [Ano74-223, Ano75-236, Ano78-39, Bea71, Gos79b, PMG77, Tor79, Van71, Wel79]. ICE [Tel73]. ICs [All75h, All76c, All76e, All76f]. idea [Bla77, Chr76b, Gin75, McK74, Tib78, Buc78b]. ideal [TT74]. ideas [Kat75]. identification [Ele73a, Hor76]. identifying [Rub76e]. IEEE [Ano71b, Ano70-169, Ano70-311, Ano70-297, Ano70-352, Ano70-353, Ano70-354, Ano70-355, Ano70-356, Ano70-357, Ano70-358, Ano70-359, Ano70-360, Ano70-361, Ano70-362, Ano70-455, Ano71-120, Ano71-253, Ano71-241, Ano71-274, Ano71-275, Ano71-276, Ano71-277, Ano71-278, Ano71-279, Ano71-280, Ano71-281, Ano71-282, Ano71-283, Ano71-284, Ano71-285, Ano71-242, Ano71-243, Ano71-244, Ano71-245, Ano71-246, Ano71-247, Ano71-248, Ano71-249, Ano71-250, Ano71-251, Ano71-252, Ano71-119, Ano71-407, Ano72b, Ano72-319, Ano72-363, Ano72-364, Ano72-365, Ano72-366, Ano72-367, Ano72-368, Ano72-369, Ano72-370, Ano72-371, Ano72-372, Ano72-373, Ano72-374, Ano72-320, Ano72-321, Ano72-322, Ano72-323, Ano72-324, Ano72-325, Ano72-326, Ano72-327, Ano72-328, Ano72-329, Ano72-330, Ano72-331, Ano72-332, Ano72-333, Ano72-334, Ano72-335, Ano72-336, Ano72-337, Ano72-338, Ano72-339, Ano72-340, Ano72-341, Ano72-342, Ano72-343, Ano72-344, Ano72-345, Ano72-346, Ano72-504, Ano73-303]. IEEE [Ano73-304, Ano73-305, Ano73-306, Ano73-354, Ano73-355, Ano73-356, Ano73-357, Ano73-358, Ano73-359, Ano73-360, Ano73-361, Ano73-362, Ano73-363, Ano73-364, Ano73-307, Ano73-308, Ano73-309, Ano73-310, Ano73-311, Ano73-312, Ano73-313, Ano73-314, Ano73-315, Ano73-316, Ano73-317, Ano73-318, Ano73-319, Ano73-320, Ano73-321, Ano73-322, Ano73-323, Ano73-324, Ano73-325, Ano73-326, Ano73-327, Ano73-328, Ano73-329, Ano73-330, Ano73-331, Ano73-332, Ano73-333, Ano73-334, Ano73-335, Ano73-336, Ano73-337, Ano73-442, Ano73-498, Ano74-196, Ano74-197, Ano74-198, Ano74-199, Ano74-200, Ano74-201, Ano74-202, Ano74-203, Ano74-204, Ano74-205, Ano74-206, Ano74-207, Ano74-208, Ano74-209, Ano74-168, Ano74-158, Ano74-159, Ano74-160, Ano74-161, Ano74-162, Ano74-163, Ano74-164, Ano74-165, Ano74-166, Ano74-167, Ano74-180, Ano74-169, Ano74-170, Ano74-171, Ano74-172, Ano74-173, Ano74-174, Ano74-175, Ano74-176, Ano74-177, Ano74-178, Ano74-179, Ano75-153, Ano75-205]. IEEE [Ano75-206, Ano75-207, Ano75-208, Ano75-209, Ano75-210, Ano75-211, Ano75-212, Ano75-213, Ano75-214, Ano75-215, Ano75-216, Ano75-154, Ano75-155, Ano75-156, Ano75-157, Ano75-158, Ano75-159, Ano75-160, Ano75-161, Ano75-162, Ano75-217, Ano75-163, Ano75-164, Ano75-165, Ano75-166, Ano75-167, Ano75-168, Ano75-169, Ano75-170, Ano75-171, Ano75-172, Ano75-173, Ano75-174, Ano75-182, Ano75-183, Ano75-184, Ano75-185, Ano75-175, Ano75-176, Ano75-177, Ano75-178, Ano75-179, Ano75-180, Ano75-181, Ano75-186, Ano75-304, Ano76-135, Ano76-170, Ano76-171, Ano76-172, Ano76-173, Ano76-174, Ano76-175, Ano76-176, Ano76-177, Ano76-178, Ano76-179, Ano76-180, Ano76-181, Ano76-140, Ano76-141, Ano76-142, Ano76-143, Ano76-144, Ano76-145, Ano76-146, Ano76-147, Ano76-148, Ano76-136, Ano76-137, Ano76-138, Ano76-139, Ano76-149, Ano76-150, Ano76-151, Ano76-152, Ano76-153, Ano76-154, Ano76-155, Ano76-156, Ano76-157, Ano76-161, Ano76-162, Ano76-163, Ano76-158, Ano76-164]. IEEE [Ano76-159, Ano76-165, Ano76-166, Ano76-160, Ano77-29, Ano77-143, Ano77-144, Ano77-145, Ano77-146, Ano77-147, Ano77-160, Ano77-161, Ano77-162, Ano77-163, Ano77-164, Ano77-165, Ano77-166, Ano77-167, Ano78-154, Ano78-155, Ano78-148, Ano78-149, Ano78-150, Ano78-151, Ano78-152, Ano78-153, Ano78-178, Ano78-179, Ano78-180, Ano78-181, Ano78-182, Ano78-183, Ano78-184, Ano78-185, Ano78-186, Ano78-187, Ano78-188, Ano78-189, Ano79-47, Ano79-48, Ano79-146, Ano79-147, Ano79-148, Ano79-149, Ano79-150, Ano79-151, Ano79-156, Ano79-157, Ano79-152, Ano79-153, Ano79-154, Ano79-155, Ano79-173, Ano79-174, Ano79-175, Ano79-176, Ano79-177, Ano79-178, Ano79-179, Ano79-180, Ano79-181, Ano79-182, Ano79-183, Ano79-184, Bea74a, Chr77i, Dil75, Dre74, Gol73, Gol79, Gra70a, Gra70b, Kin72, MRR^{+}77, Mul71, Rub74d, Rub75a, Rub75c, Rub76a, Rub76b, Rub76c, Rub76d, Rub76e, Rub76f, Rub76g, Rub76h]. IEEE [Rub76j, Rub77a, Rub77b, Rub77c, Ste75b, Wei78b, Ano70-182, Ano70-183, Ano70-184, Ano70-185, Ano70-186, Ano70-187, Ano70-188, Ano70-189, Ano70-190, Ano70-191, Ano70-192, Ano70-193, Ano70-194, Ano70-195, Ano70-196, Ano70-197, Ano71-113, Ano71-134, Ano71-135, Ano71-136, Ano71-137, Ano71-138, Ano71-139, Ano72-164, Ano72-190, Ano72-191, Ano72-192, Ano72-193, Ano72-194, Ano72-195, Ano72-196, Ano72-197, Ano72-198, Ano72-199, Ano72-200, Ano72-201, Ano73-181, Ano73-182, Ano73-183, Ano73-184, Ano73-185, Ano73-186, Ano73-187, Ano73-188, Ano73-189, Ano73-190, Ano73-191, Ano73-192, Ano74-60, Ano74-61, Ano75-57, Ano75-319, Ano77-43, Ano78-62]. If [Sim71, Kap76c, Rub76a, Tuc74, War76, Ano70-296].

IFR [Fri70a]. ifs [Gan77]. Ignition [Jur75b]. ignitions [Jur75b]. II [Cot75, Ern77a, Fri76b, Sug78d, Tuc74, Bel74, Bib71b, BHLK71, Chr74d, Chr74l, Fal74i, Fal75d, Fal75e, Fri70b, Fri70f, Fri76d, Gar70b, Gor75, Hei79, Jur75d, Mas79a, Men76g, Mos70b, Nay71b, Pay75, Pie79, Pla79b, Pol71b, PMG77, Rie70b, Rub76g, Shi79b, Sne72b]. II/Avocation [Gor75]. II/Design [Fal75e]. III [All75i, Ano70-296, Ano71-297, Ano75-319, Ano78-243, Bib71c, Chr74e, Chr74m, Dav79, Ern77b, FK76, Fri70c, Gar70c, Lin76b, Lou74, Mid79, Rie70c, Rob74a, Tor76d, Tuc75b]. III/Avocation [Tuc75b]. III/Design [All75i]. ill [RK79, SR78]. Illiac [Fal76c]. illuminate [KLH74]. I’m [Dib72]. Image [ATK72, And79, MVE75]. Images [And79, Ano74-222, ME77]. imagined [Jur76a]. Imaging [Hor78, ME77, Jur74h, Jur78b, MVE75]. immersed [Ano78-46]. imminent [And78, Fag78]. immunity [Boa73, JM77, SS73]. impact [Ano74-96, Ano75-303, Ano75-304, Col78c, Han79, Kap78c, Man79, Max77, Que79, Rub74e, Wil76b]. impacts [Fag78]. impasse [Ano79-53]. Impediments [Mos74, Str71]. imperiled [Fri70e, Fri70f]. implantable [Jur76f]. implanted [Spe78]. implementation [Fal74e, Man79]. implemented [Fra78]. implications [Rub76c, Rub76j, Too78b]. import [Ull70]. Important [Ano71-253, Fri77a, Kap76c, Mit77, Tor77g, Wul78]. impressed [Kap76a, KLM^{+}79]. impressive [All75f]. improve [Bat71, Jur75a, Jur75b, KI77, Rab78, Tor77h]. improved [Fri70b, HB77, Hus74, Tor76a, Tor78a]. improves [von79]. Improving [Bor73, Fri76h, Rub74b]. In-field [All79e]. in-plant [All79e]. in-process [All79e]. in-prototype [Ano78-194]. Inc.

[Ano70-312, Ano70-313, Ano70-314, Ano70-315, Ano70-316, Ano70-317, Ano70-318, Ano70-319, Ano70-320, Ano70-321, Ano70-322, Ano70-323, Ano71-286, Ano71-287, Ano71-288, Ano71-289, Ano71-290, Ano71-291, Ano71-292, Ano71-293, Ano71-294, Ano71-295, Ano71-296, Ano72-375, Ano72-376, Ano72-377, Ano72-378, Ano72-379, Ano72-380, Ano72-381, Ano72-382, Ano72-383, Ano72-384, Ano72-385, Ano72-386, Ano73-365, Ano73-366, Ano73-367, Ano73-368, Ano73-369, Ano73-370, Ano73-371, Ano73-372, Ano73-373, Ano73-374, Ano73-375, Ano73-376]. incentives [Kap75b, Par78b]. incident [Kap76e]. include [Ano78-192, Fag79, LC75, Mac78, Pat79b, Too78a]. includes [Ano74-226, Ano75-239, Fri76f]. including [Jur74f]. income [RO77, Rub74a, Sug78c]. increase [All75b, Ele74e, Ewa78, Nag74, Tor77h, von79]. increased [Tor76a]. increasing [Alb78, Ele75a, Pla79a]. Independent [PHC76, Rub77d]. independently [Lou74]. Index [Ano70-298, Ano70-299, Ano70-300, Ano70-301, Ano70-302, Ano70-303, Ano70-304, Ano70-305, Ano70-306, Ano70-307, Ano70-308, Ano70-309, Ano71-113, Ano71-254, Ano71-255, Ano71-256, Ano71-257, Ano71-258, Ano71-259, Ano71-260, Ano71-261, Ano71-262, Ano71-263, Ano71-264, Ano71-265, Ano72-164, Ano72-347, Ano72-348, Ano72-349, Ano72-350, Ano72-351, Ano72-352, Ano72-353, Ano72-354, Ano72-355, Ano72-356, Ano72-357, Ano72-358, Ano73-338, Ano73-339, Ano73-340, Ano73-341, Ano73-342, Ano73-343, Ano73-344, Ano73-345, Ano73-346, Ano73-347, Ano73-348, Ano73-349, Ano73-181, Ano74-193, Ano75-199, Ano75-200, Ano76-169, Ano77-124, Ano77-125, Ano77-126, Ano77-127, Ano77-128, Ano77-129, Ano78-158, Ano78-159, Ano78-160, Ano78-161, Ano78-162, Ano78-163, Ano79-158, Ano70-168, Ano74-60, Ano75-57, Ano76-38, Ano77-43, Ano78-62, Ano79-46]. indicates [Ano74-229]. indications [Ano74-209, Ano75-217]. indirectly [Nas78]. Indonesia [Rub77f]. induction [Pol71a, Pol71b]. inductively [Spe78]. Inductorless [Mos70a, Mos70b]. Industrial [Kap75c, Kap76d, Kay78, Chr73, Fri74c, Lee71, SS73]. industries [Fle71]. Industry [All79d, Ano76-209, Ano78-205, Ano78-211, Ano78-213, Ano79-196, Ano79-197, Ano79-198, Ano79-199, Ano79-200, Ano79-201, Ano79-204, Ano79-205, Ano79-206, BB71, Fal76b, Kap77b, Ano73-404, Ano73-405, Ano73-406, Ano73-407, Ano73-408, Ano73-409, Ano73-410, Ano74-236, Ano74-237, Ano74-238, Ano74-239, Ano74-240, Ano74-241, Ano74-242, Ano74-243, Ano74-244, Ano74-245, Ano74-246, Ano75-244, Ano75-245, Ano75-246, Ano75-247, Ano75-248, Ano75-249, Ano75-250, Ano75-251, Ano75-252, Ano75-253, Ano75-254, Ano75-255, Ano76-207, Ano76-208, Ano76-210, Ano76-211, Ano76-212, Ano76-213, Ano76-214, Ano76-215, Ano76-216, Ano76-217, Ano76-218, Ano77-180, Ano77-181, Ano77-182, Ano77-183, Ano77-184, Ano77-185, Ano77-186, Ano77-187, Ano77-188, Ano77-189, Ano77-190, Ano77-191, Ano78-202, Ano78-203, Ano78-204, Ano78-206, Ano78-207, Ano78-208, Ano78-209, Ano78-210, Ano78-212, Ano79-202, Ano79-203, Ano79-207, Bea75, Fri75b, Hin79, Jur74c, KI77, Lin75a]. industry [Lom79, Rub77d, Sch77b, Tri74]. Industry-university [BB71]. inertia [Kap75b]. inevitably [Kap78b]. inexpensive [Mor78, Jur76d]. infant [Kap75e]. inferior [New76]. inflation [Ano74-232, Ele75b, Her79, Rub75a, Rub76b, Wei78a]. inflation-crippled [Ele75b]. inflationary [Rub74a, RF75]. influence [Clo73, Nas78]. Information [Ano70-310, Ano71-266, Ano71-267, Ano71-268, Ano71-269, Ano71-270, Ano71-271, Ano72-316, Ano72-359, Ano72-360, Ano72-361, Ano72-362, Ano73-350, Ano73-351, Ano73-352, Ano73-353, Chr78j, Gup74, Jur72c, Lav78, Men75d, Tuc75a, TH77, Ano74-71, Jen78, Mos74, Que79, Tea74, Tro74]. information-retrieval [Jen78]. informed [Mac76]. Infrared [Ano70-296, HS70, Oli74]. ingenor [Far73]. Ingenuity [All79e, Gre79d]. Initial [Gal79, Lin75a]. initiating [Spi76a]. Initiative [Rub76h]. inject [Kor77]. injection [Jur75a, KS70, Tor77e, War76]. injectors [Jur75a]. innards [All79g]. Innovation [Der79, Wol75a, Wol75b, Wol76a, Bar78, Gro72, Jur72c, Nas78]. innovations [Har77]. Innovative [Ele74b, Fre74, Wul78]. innovator [Tib78]. innovators [Buc78b, Raf78]. Input [Ano74-293, Ano75-237, Ano74-223, Ano78-197, Fal74e, Mar77, She79]. Input-output [Ano74-293, Ano75-237, Fal74e]. inputs [Tri74]. inroads [Doi77, Kap75c]. insecure [Bea72c, Hel79]. insidious [Chr77j]. Insight [KLV75]. inspiration [JF74]. inspirational [Men74g]. installation [Ele75a]. Institute [Ano74-209, Ste75a, All74d, Ano70-312, Ano70-313, Ano70-314, Ano70-315, Ano70-316, Ano70-317, Ano70-318, Ano70-319, Ano70-320, Ano70-321, Ano70-322, Ano70-323, Ano71-286, Ano71-287, Ano71-288, Ano71-289, Ano71-290, Ano71-291, Ano71-292, Ano71-293, Ano71-294, Ano71-295, Ano71-296, Ano72-375, Ano72-376, Ano72-377, Ano72-378, Ano72-379, Ano72-380, Ano72-381, Ano72-382, Ano72-383, Ano72-384, Ano72-385, Ano72-386, Ano73-365, Ano73-366, Ano73-367, Ano73-368, Ano73-369, Ano73-370, Ano73-371, Ano73-372, Ano73-373, Ano73-374, Ano73-375, Ano73-376, Ano74-208, Ano75-217, Ano75-304, Ano77-219, Ano77-220, Ano77-221, Ano77-222, Ano77-223, Ano77-224, Ano78-245, Ano78-246, Ano78-247, Ano78-248, Ano78-249, Ano78-250, Ano78-251, Ano78-252, Ano78-253, Ano78-254, Ano78-255, Ano78-256, Ano79-237, Ano79-238, Ano79-239, Ano79-240, Ano79-241, Ano79-242, Ano79-243, Ano79-244, Ano79-245, Ano79-246, Ano79-247, Ano79-248, BCY75, BDL^{+}78, Okr78]. Institute [Rub74c, Rub76c, Rub76e, Rub76h, Rub77a, Rub77b]. institutional [Chr79o, Kap75b]. institutions [Han79]. instruction [Ano78-201, Sug78a]. Instructional [PG70, RRS^{+}74]. instructions [Ano75-232]. Instrument [All76f, Ano74-229, Fri70b, Men74d, Rub77e]. Instrumentation [All74b, All74c, All75c, All76e, All76f, All76g, All77b, All78b, All78c, All79e, Ele74c, Ele76a, Fal74i, FK76, Hil74, Jur74f, Jur74g, Jur75e, Kap74b, KLV75, Lou74, Men74d, MM74, New76, Oli74, TT74, Yu75, All75d, Ano76-35]. Instruments [All75d, Ano70-324, Ele74d, All76g, All78c, Ano74-233, Ano79-190, New76, Rai76, Ter77, Buc78b]. insulators [Tuc75b]. insulin [Spe78]. integral [Ano79-192]. Integrated [And78, Ano74-224, Gar70a, Gar70b, Gar70c, War76, All76f, All77b, Ano74-96, Ano74-228, Fai74, Joh70b, Lap74d, Mas79b, Mok79c, Oll78, Sur70, Tor77e, Wol76b]. Integrated-circuit [Gar70a, Gar70b, Gar70c, Joh70b]. integrates [CC78]. Integrating [SA78]. integration [HH70, Mac78, Moo79, Tor78b]. integrity [Men75d]. intellect [Men76e]. intelligence [Fag78, Kay78]. Intelligent [FFCT74, Too78b]. Intelsat [PW71, Qua79]. intensities [EBC^{+}79]. Interactive [All76h]. INTERCOM [Lap74e]. INTERCON [Ano72-387, Ano72-388, Ano73-303, Ano73-306, Ano73-377, Ano73-442, Ano75-317]. interconnect [Men79a]. interconnecting [Lou74]. interconnection [GPPM71, McK72b]. interconnections [DeF70]. Intercontinental [Mac78]. interdisciplinary [Ele74b]. interest [Ano76-270, Chr75c, Fri74d, Jur76d, Rub77f]. interests [Ano71-119, Fri75e]. Interface [Ged72, Rub76i, All76b, Ano75-239, Ano77-178, Bar73, Fal74a, KLV75, Lou74, Rub76d, Wel79]. interfaces [Ele76a, Ern77a, HFL74, RL74]. Interfacing [RRS^{+}74, Ano78-191, KLV75, Mor78]. interior [Eng75]. interlace [New76]. internal [Wit77]. International [Ano70-297, Ano71b, Dar70, Fal74j, Gre79d, Rea79, Pog78, Rub79c, Ano71-247]. internationally [KH79]. interpreter [Ano79-53]. interrupter [Dal70]. Intervals [von79]. interview [Fri76e]. interviews [Ano71-297, Ano71-298, Mas79e, Rub74c]. intriguing [Lin75b]. intro [Rub79b, Tor77c]. Introducing [Ano78-164]. Introduction [Fri71c, OE70, Rub78a, Ano74-40, Ano79-232, Fri70b, Van71]. invade [Kap77c, Mor78]. invades [All79f]. invaluable [Alb78]. invented [Gin75]. Inventing [Bla77, Wol75a, Wol75b]. Invention [All78e]. inventions [Fra78]. inventor [Der79, Par78b, Rab78]. Inventors [Rac78]. invest [Rub75f]. investigators [Mas79e]. investment [Gre74]. invitation [TT74]. invited [Mas79f]. involved [Fri74b, Man70]. involving [Joh79]. inward [BDL^{+}78]. ion [Bre71]. Iowa [Col79]. IR&D [PHC76]. IRAC [Bus72]. IRE [Dre74, Rub76g].

Island [GPPM71, Mas79e, Rub79b]. isolation [Ano75-40]. Israel [Kap77e]. issue [Rub77a, Rub79b, Tor77c]. issued [GHNK79]. Issues [Ano72-454, Ano78-146, Ano70-400, Ano70-401, Ano70-402, Ano70-403, Ano70-404, Ano70-405, Ano70-406, Ano70-407, Ano70-408, Ano70-409, Ano70-410, Ano70-411, Ano71-229, Ano71-230, Ano71-231, Ano71-232, Ano71-233, Ano71-234, Ano71-235, Ano71-236, Ano71-237, Ano71-238, Ano71-239, Ano71-240, Ano71-368, Ano71-369, Ano71-370, Ano71-371, Ano71-372, Ano71-373, Ano71-374, Ano71-375, Ano71-376, Ano71-377, Ano71-378, Ano71-379, Ano72-304, Ano72-305, Ano72-306, Ano72-307, Ano72-308, Ano72-309, Ano72-310, Ano72-311, Ano72-312, Ano72-313, Ano72-314, Ano72-315, Ano72-465, Ano72-466, Ano72-467, Ano72-468, Ano72-469, Ano72-470, Ano72-471, Ano72-472, Ano72-473, Ano72-474, Ano72-475, Ano72-476, Ano73-291, Ano73-292, Ano73-293, Ano73-294, Ano73-295, Ano73-296, Ano73-297, Ano73-298, Ano73-299, Ano73-300, Ano73-301, Ano73-302, Ano73-456, Ano73-457, Ano73-458, Ano73-459, Ano73-460, Ano73-461, Ano73-462, Ano73-463, Ano73-464, Ano73-465, Ano73-466, Ano73-467]. issues [Ano74-181, Ano74-182, Ano74-183, Ano74-184, Ano74-185, Ano74-186, Ano74-187, Ano74-188, Ano74-189, Ano74-190, Ano74-191, Ano74-192, Ano74-146, Ano74-147, Ano74-148, Ano74-149, Ano74-150, Ano74-151, Ano74-152, Ano74-153, Ano74-154, Ano74-155, Ano74-156, Ano74-209, Ano74-293, Ano74-282, Ano74-283, Ano74-284, Ano74-285, Ano74-286, Ano74-287, Ano74-288, Ano74-289, Ano74-290, Ano74-291, Ano74-292, Ano75-187, Ano75-188, Ano75-189, Ano75-190, Ano75-191, Ano75-192, Ano75-193, Ano75-194, Ano75-195, Ano75-196, Ano75-197, Ano75-198, Ano75-140, Ano75-141, Ano75-142, Ano75-143, Ano75-144, Ano75-145, Ano75-146, Ano75-147, Ano75-148, Ano75-149, Ano75-150, Ano75-151, Ano75-217, Ano75-303, Ano75-292, Ano75-293, Ano75-294, Ano75-295, Ano75-296, Ano75-297, Ano75-298, Ano75-299, Ano75-300, Ano75-301, Ano75-302, Ano76-167, Ano76-168, Ano76-125, Ano76-126, Ano76-127, Ano76-128, Ano76-129, Ano76-130, Ano76-131, Ano76-132, Ano76-133, Ano76-253, Ano76-254]. issues [Ano76-255, Ano76-256, Ano76-257, Ano76-258, Ano76-259, Ano76-270, Ano77-119, Ano77-120, Ano77-121, Ano77-122, Ano77-123, Ano78-156, Ano78-157, Ano78-145, Ele74f, Fri76i, Hef70, JF74, Man79]. item [Men78c]. IV [Ano71-298, Ano78-244, Chr74f, Fal74c, Fal75f, Fal76c, Jur75e, Kap76e, Lin76c, Men74d, RM75, Tuc79]. IV/Avocation [RM75]. IV/Education [Fal75f]. IX [All74b, Wol76c].

jack [Fal74i]. jack-of-all-trades [Fal74i]. jamming [Wal78]. Japan [Ano70-247, Ano70-325, Ano71-297, Ano71-298, Ano72-389, Chr73, Chr77d, Gre74, Gre79d, IFSK77, IY77, Kap79b, Rea79, Sch77b, Shi79b, Shi76, SOW77, Tor77c, Tor77g]. Japanese [Gre79d, KI77, Mac78, Tor77g]. jewel [Ang74]. jewels [McK74]. JFETs [Ano75-236]. Job [Gol72, Bea74a, Bea73a, Jur73b, Men78g, Pat79b, Wei78a]. job-wise [Bea74a]. jobs [Fri76b, Ter77]. John [Men76f]. journals [Ano70-449, Ano70-450, Ano70-451, Ano70-452, Ano70-453, Ano70-454]. judges [Chr75h]. Judging [Chr75h]. July [Sug78c, WZ78]. jump [Tor79]. Jupiter [Ano78-49]. just [AC75, Bea75, Fal74j, Fri74f, Spi76a, Wes73].

Kalman [Sor70]. Kappa [Mas79d]. kb [Ano75-241, Lap74c, Tor76c]. kbit [Gos79a, Too78a]. kbyte [Ano78-193]. keep [Ish77, Jen78, Jur75b, Lap74b, Pat79b]. Keeping [Ano74-305, Ano79-52, Bea72d, Her73, Ste73, Fal74g, Pat79b, Wei77]. Kelvin [Ana79]. kept [Fri74k, Qua79]. Key [Rub77a, Ano75-304, Ano75-319, Col78b, Fri74h, JF74, Oli74, PHC76, SA78]. keyboard [Ano77-178, Ano79-192]. kg [HL79]. Kilometer [Bro79, Fri74k]. Kilometer-diameter [Bro79]. Kilowatt [Fri72c]. Kilowatthours [Sal75]. kilowatts [Sch77b]. kind [Spi76b]. kits [Men77b]. klystron [Gin75]. km [Eng75]. knack [Gre74]. knobs [Ele74c]. know [HFL74, Jur74c, MM74]. knowledge [Alb78]. Kongress [Ano72-253]. kV [Fri77c]. kWh [Lev78].

L [War76]. lab [Tor76c]. label [And73b, Mas79c]. labor [Doi77, Gre74]. laboratory [Ano78-194, Ano79-188, Jur75c, Mos75]. Labrador [Fri71b]. labs [Pat79a, Rub77d, Pie79]. Lackawanna [Bla77]. lags [Par78b]. Laid [Pel70, Tho73a]. Land [Kap77d, SS71, HH74, Lus78]. Lander [Spi76a]. Landing [KRS78, Ele76b, Fri70b, Mas79b, Pog78, SA78]. landings [NR74]. Landsat [Hor76]. language [Ano78-193, KLV75, Que79]. languages [Ele76a]. Large [Fai74, Kun75, Mac78, And79, Ano77-96, Ano77-97, Ano77-98, Ano77-99, Ano77-100, Ano77-101, Ano78-127, Ano78-128, Ano78-129, Ano78-130, Ano78-131, Ano78-132, Ano79-126, Ano79-127, Ano79-128, Ano79-129, Ano79-130, Ano79-131, Fal75b, Har77, HH70, Moo79, Tor78b, Lap74d]. large-diameter [Har77]. Large-scale [Fai74, Mac78, HH70, Tor78b, Lap74d]. largely [Rub77b, Ste75a]. larger [Ano74-229, Tor76c]. largest [KLH74]. Laser [Ano70-296, Ele73d, BCH^{+}79, Jur74a, Jur78b, KLM^{+}79, Ken71, KLE73]. Lasers [All79d, EBC^{+}79, Ele72c, Ele72d, Ger70, Kap77c, Men76b]. late [Jur72b]. late-blooming [Jur72b]. lately [TDK74]. latest [Ano77-29, Rub74b, Rub79c, Tor76a]. launched [Jur75d, Wei77]. launches [Lap72b]. launching [Dun78, Wul78]. laundry [Col79]. law [BCY75, Ele73c, Ele73f, Fin77, Par78b, RO76]. laws [Fin77, Jur73c]. lay [Tro74]. Layer [Lin76c, Ano71-301]. layman [Jur74e]. layouts [Ano77-178]. lead [Lin77]. leaders [Lin77]. Leakage [Ano70-326]. leaks [Kap76e]. learn [Fal74j, GH72, Rea79]. Learning [Ano78-164, Chr74g, Gra79, Jur76f, Eve77, Gos79b, Man72]. Least [Sor70, Ano71-119]. Least-squares [Sor70]. led [WZ78]. LEDs [All75c, Bar72, Men77d, NKL72]. Lee [Fri74a]. leeway [Mas79a]. legal [Kap77f, Man79]. legend [Der79]. legislation [Cla77, Rub76i]. legislative [Rub76d]. leisure [RM75, Rub75f]. lens [Lap73b]. less [Cla77, Hea74, Men77e, SL78]. Lessons [RS77, Rub74e]. Let [Bea75]. lethal [Fri71a]. level [Fri77c, Pog78, Rab78, Sug78a]. levels [Gol79, RGCS79]. liabilities [Chr77d]. library [Bea74b, Fri70d]. life [Fal75d, Jur76d, KLM^{+}79, Spi76a, Wei78a, von79]. life-saving [Jur76d]. lifetime [Jur76f]. lift [Joh79]. Light [Ano79-145, Bea75, Fal74b, Ish77, Jur74f, Kal79, KS70, Kra75, Mid74, All75c]. light-beam [Kra75]. Light-emitting [All75c]. Light-polarizing [Ano79-145]. lighting [Bea75, Der79]. lightning [Sug78c]. Like [Tri78, Ano71-421, Ano79-232, Jur74g, Lin74a, Men77e, Tuc74]. liked [Rub76d]. likelihood [Ewa78]. limit [Ano78-200]. Limited [Kay78, Tor76c]. line [Alb78, All77e, All79f, Ano75-234, Ano79-193, Eng75, Fir75, Gre79b, Ish77, Men76b, Men79a, Mun78, Pol70, SZ74, Tor77b, Lin74b]. Linear [Pol71a, Pol71b, Sch71, All76c, Mor78, Mos70c, Sne72a, Sne72b, Mos70b]. linearity [Ano79-145]. lines [Ano70-424, Ano70-425, Ano70-426, Ano70-427, Ano70-428, Ano70-429, Ano70-430, Ano70-431, Ano70-432, Ano70-433, Ano70-434, Ano70-435, Ano71-406, Ano71-395, Ano71-396, Ano71-397, Ano71-398, Ano71-399, Ano71-400, Ano71-401, Ano71-402, Ano71-403, Ano71-404, Ano71-405, Ano72-492, Ano72-493, Ano72-494, Ano72-495, Ano72-496, Ano72-497, Ano72-498, Ano72-499, Ano72-500, Ano72-501, Ano72-502, Ano72-503, Ano73-486, Ano73-487, Ano73-488, Ano73-489, Ano73-490, Ano73-491, Ano73-492, Ano73-493, Ano73-494, Ano73-495, Ano73-496, Ano73-497, Ano75-318, Ano76-271, Ano77-226, Chr74c, Chr74d, Chr74e, Chr74f, Chr74g, Chr74h, Chr74b, Chr74i, Chr74j, Chr74k, Chr74l, Chr74m, Chr75c, Chr75d, Chr75b, Chr75e, Chr75f, Chr75h, Chr75i, Chr75l, Chr75k, Chr75j, Chr75g, Chr76h, Chr76e, Chr76d, Chr76k, Chr76j, Chr76f, Chr76g, Chr76b, Chr76i]. lines [Chr76c, Chr76l, Chr77a, Chr77b, Chr77g, Chr77h, Chr77c, Chr77i, Chr77j, Chr77d, Chr77f, Chr77e, Chr77k, Chr78a, Chr78c, Chr78b, Chr78e, Chr78j, Chr78d, Chr78k, Chr78g, Chr78l, Chr78h, Chr78f, Chr78i, Chr79a, Chr79b, Chr79c, Chr79e, Chr79d, Chr79f, Chr79g, Chr79j, Chr79h, Chr79i, Chr79k, Chr79l, Jan70, Tor77f]. link [Ano79-189, Men74c]. linked [Gib76, Lia78]. Linking [Fal74a]. links [Fag79, Kau72]. Liquid [GM72, Mos75, All75c, PP70]. liquid-crystal [All75c]. liquid-metal [PP70]. LIRR [Mas79f]. List [Ano70a, Ano71-299, Ano71-300, Ano72-390, All76g, Ano74-232, Ano78-37, Jur74b, Moo79, Zuc77]. listing [Ano73-454]. lists [Ano77-179]. literature [Ano71-114, Ano71-115, Ano71-116, Ano71-117, Ano71-118, Ano72-165, Ano72-166, Ano72-167, Ano72-168, Ano72-169, Ano72-170, Ano72-171, Ano72-172, Ano72-173, Ano72-174, Ano72-175, Ano72-176, Ano73-157, Ano73-158, Ano73-159, Ano73-160, Ano73-161, Ano73-162, Ano73-163, Ano73-164, Ano73-165, Ano73-166, Ano73-167, Ano73-168, Ano74-40, Ano74-39, Ano74-41, Ano74-42, Ano74-44, Ano74-43, Ano74-45, Ano74-47, Ano74-46, Ano74-36, Ano74-37, Ano74-38, Ano75-38, Ano75-39, Ano75-40, Ano75-41, Ano75-42, Ano75-43, Ano75-33, Ano75-34, Ano75-35, Ano75-36, Ano75-37, Ano76-36, Ano76-35, Ano76y, Ano76z, Ano76-27, Ano76-28, Ano76-29, Ano76-30, Ano76-31, Ano76-32, Ano76-33, Ano76-34, Ano77p, Ano77q, Ano77r, Ano77s, Ano77t, Ano77u, Ano77v, Ano77w, Ano77x, Ano77y, Ano77z, Ano77-27, Ano78-34, Ano78-36, Ano78-35, Ano78-37, Ano78-38, Ano78-39, Ano78-40, Ano78-41, Ano78-42, Ano78-43, Ano78-32]. literature [Ano78-33, Ano79-45, Ano79-44]. liters [Sal75]. little [Lin74b]. live [Gra79]. living [Her79, Wei78a]. LMFBR [Fri74b]. load [Chr79a, Fri76f, Fri77a, Kap77f, LC75, Pla79b, Pla79a, Ura77]. loading [Ura77]. loads [Pla79a]. Local [Pla79b, Kay78, SL78, Sug78a]. locally [Kor77]. located [RG79]. Loch [EW78]. locked [Gre71, Moo73]. locomotives [Fri74i]. LOCOS [Ano71-301]. Logic [All77c, All76c, Ano78-43, Ano79-54, Boa73, DeF70, Gar70a, Gar70b, Gar70c, HW71, Hil74, Sus73, Tor77e, War76]. logistics [Tea74]. London [Fri74f]. Long [DBB71, Bar79, Buc78b, EH76, Eng75, Hea74, Jur74c, Kap78a, KLM^{+}79, Okr78, Qua79, Rub77b, Rub77c, GPPM71]. long-distance [EH76, Qua79]. long-life [KLM^{+}79]. long-range [Rub77b]. long-shot [Buc78b]. Long-term [DBB71, Bar79, Jur74c, Kap78a]. longer [Jur75b, Kap75e]. look [All78d, Ano72-405, Ano74-230, Ano79-235, Bar79, Chr78h, Chr79b, Col78c, Fin77, Lin75b, McA71, Oll78, Rub76d, Wil76a]. look-ahead [Wil76a]. look-up [Ano74-230]. Looking [Chr75n, Chr79b, Har77, Max77, Oli74, VHM77, Cos74, Mar72, Tuc75a]. looks [BDL^{+}78, Bus72, Jur75c]. loop [Chr76k, Gre71, NS77]. loops [Moo73]. losing [Mac73]. loss [Ano78-41, BB71]. lost [Sug78c]. lots [Jur77b, Sug79a]. lottery [BC73a]. Low [Ano70-326, Ano75-238, Ano77-178, Ano79-191, AC75, Ano74-227, Ano75-38, Ano75-233, Ano76-206, Ber79, BCH^{+}79, CT70, Fri74e, Jur75e, Lap72b, Pay75, Rob74a, Sal77, Tho73b]. low-budget [Ano74-227]. Low-cost [Ano75-238, BCH^{+}79, Jur75e, Lap72b]. low-cost-per-bit [Sal77]. low-pass [Ano75-38]. low-power-drain [CT70]. Low-price [Ano79-191]. Low-profile [Ano77-178]. low-risk [Pay75]. Lower [Pit79, HB77, VHM77]. lowest [Ano75-235]. LSI [Ano79-159, Fal75b, HGB79, Lap72b, Sug79a, Wel79]. lubricating [Mos75]. lumens [Bea75]. luminous [TG78]. lurk [Kin75]. luxury [PMG77].

m [Ano79-189]. M.I.T. [Mas79c]. M.S. [Gra74]. machine [Ano76-204, Chr77j, Kor77, Ohi70, RK79, RRS^{+}74, SZ74]. machine-tool [Kor77]. Machines [Fre74, Fal75b, FFCT74, Men76f, Pol71a, Pol71b, Yen78]. made [BBG75, Lus78, Men75d, Moo79, TH77]. madness [Ano74-293]. mag [KLM^{+}79]. magazine [Fal76b]. magazines [RM75]. Maglev [Tho73b]. Magnetic [Sal77, SZ74, All75a, Ano76-36, Cam77, Fet76, HL79, Lap72a, Too78a]. magnetic-bubble [Fet76, Too78a]. Magnetic-stripe [SZ74]. magnetics [Mos75]. magnetohydrodynamic [Lev78]. magnetometers [Rai76]. maiden [Ano79-249, MF79]. Mailgram [Ano77-179]. Mainframe [Too78b]. mainline [Fri76j]. maintainability [SL78]. maintenance [Fal74i, Jur75b]. major [All75i, Ano78-258, Ano79-48, Fri70f, Fri76i, Fri77c, Gle77, Max77, Rub76f, Wil76b]. majority [Men76a]. make [Ano72-318, Cla77, Gre74, Kap75c, Lev78, Men74b, Men76a, Men77d, TT74]. makers [Gos79b, Jur78c, Lap74b, Lom79, McW77, Sur72]. makes [Fal74d, Jay76, Lou74]. Making [Ano78-40, Tri74, All76a, Fal74i, Gre76b, Rub76h, Sug78b, Tri78, TH77]. maligned [Rub76d]. Man [Ohi70, Ano71-344, Gra79, Jur75d, Kap74b, Men76e, RRS^{+}74]. Man-machine [Ohi70]. Management [Ano78-243, Gra79, Kat75, Rec75, Bea74a, Fri70c, Fri76f, Kap77f, LC75, Pla79b, Pla79a, RS77, RO77, Wil71, Wol79]. manager [Raf78, Wol77]. managers [Lin74b, Rec75]. Managing [Fri72g, Wol77]. mandate [Ste75b]. mandated [Ele75a]. manipulates [Ano78-196]. Manpower [Gra74, Rub74b]. manufactured [Lou74]. Manufacturers [Jur76a, IY77, Jur78c, Lap74b]. Manufacturing [Whi72, Ano78-36, Kap77c]. Many [Bra77, All74a, Fai74, Fri75f, Kor77, Lus78, Men74e, RS77, Rub79a]. map [Fri70c]. mapping [Hor76]. March [Ano70-297]. Marconi [McK74, Ang74]. Marine [KH79]. Marine/aviation [KH79]. Market [Ano70-218, Bea72f, Gre79d, Men77b, Men79b]. Marketing [NC71, Ano71-344, Gre79d, Wul78]. markets [All79h, Lap74e, Rea79]. marks [Fri77b]. marriage [Wol76a]. marries [Lin76c]. Mars [NR74]. marshaled [All75f]. Martian [Spi76a]. mask [Bea71]. mass [Rub74d]. master [Fri75d]. Mastering [Gos79b, Hei79]. masters [Rub76b]. matching [Fri76f]. Materials [Ele72d, BKS79, Joh79]. math [Mil72]. Mathematical [Bib71c]. mathematician [Fri75d]. mathematics [FK76]. Matrix [TG78]. matter [Ano76-123, Ano76-124, Fre74, MF79, Sug79a]. matters [Chr79o, Fri70f]. maturation [Ano78-94]. maturity [Buc72]. Mauchly [Mau75]. maximum [Ano76-206]. may [Ano78-48, Chr76b, Ele74e, EH76, Fal75c, Fra78, Fri74d, Gra74, Hea74, Jon77b, Jur74d, Jur74e, Jur74h, Jur75c, KI77, Lev78, Mar77, Men74e, Men75b, Men78g, NS77, Nil76, Oli74, Rei74, Rub75c, Rub77b, Sug78a, Til70, Tuc75a, Yen78]. Maze [All78d, All79g, Chr77b]. MBd [Ano79-189]. McGovern [Ano72-454]. me [TDK74]. mean [Reu78]. means [And79, Mas79a]. meant [Col78b, Jur75d]. measure [Hor78]. measurement [All74b, Ano79-192, Ter77]. measurements [Ano78-40, Ele74d, Hil74, Rot71]. measures [Ano75-233, Col78c]. Measuring [Ano74-43, Sta72, Wul78, Ano78-41, Ber71]. mechanical [Jur78a]. mechanics [Cam77]. medalists [Ano79-48]. media [Jur74f]. medic [Ano72-404, Jur76e]. Medical [All76h, Jur73c, Jur74h, Uts77, Ger74, ME77]. medicine [Jur74h, Jur77b, Jur78b, Jur79, ME77, SR78]. Medium [Ano74-225, Rie70c, Sal77]. Medium-speed [Ano74-225]. meet [Fri76f, Men78c]. Meeting [Ano78-165, Ano78-258, Ano79-159, Ano79-160, Ele78, TH77, Ell72, Ano72b]. Meetings [Ano71-302, Ano71-303, Ano71-304, Ano71-305, Ano71-306, Ano71-307, Ano71-308, Ano71-309, Ano71-310, Ano72-391, Ano72-392, Ano72-393, Ano72-394, Ano72-395, Ano72-396, Ano72-397, Ano72-398, Ano72-399, Ano72-400, Ano72-401, Ano72-402, Ano73-379, Ano73-380, Ano73-381, Ano73-382, Ano73-383, Ano73-384, Ano73-385, Ano73-386, Ano73-387, Ano73-388, Ano73-389, Ano73-390, Ano74-210, Ano74-211, Ano74-212, Ano74-213, Ano74-214, Ano74-215, Ano74-216, Ano74-217, Ano74-218, Ano74-219, Ano74-220, Ano74-221, Ano75-218, Ano75-219, Ano75-220, Ano75-221, Ano75-222, Ano75-223, Ano75-224, Ano75-225, Ano75-226, Ano75-227, Ano75-228, Ano75-229, Ano76-182, Ano76-183, Ano76-184, Ano76-185, Ano76-186, Ano76-187, Ano76-188, Ano76-189, Ano76-190, Ano76-191, Ano76-192, Ano76-193, Ano77-148, Ano77-149, Ano77-150, Ano77-151, Ano77-152, Ano77-153, Ano77-154, Ano77-155, Ano77-156, Ano77-157, Ano77-158, Ano77-159, Ano78-166, Ano78-167, Ano78-168, Ano78-169, Ano78-170, Ano78-171, Ano78-172, Ano78-173, Ano78-174, Ano78-175, Ano78-176]. Meetings [Ano78-177, Ano79-161, Ano79-162, Ano79-163, Ano79-164, Ano79-165, Ano79-166, Ano79-167, Ano79-168, Ano79-169, Ano79-170, Ano79-171, Ano79-172]. megabyte [Ano79-195]. megastructures [Fri76b]. Megawatts [Mac75, Fri72g]. member [BCY75, Bea74a, Gol79, Rub75c, Rub76c, Ste75b]. Members [Ano71-253, Ano75-153, MRR^{+}77, Rub75b, Rub76f]. Membership [Rub75a, Ste75a]. Memories [Fet73, Fet76, KLM^{+}79, Tor76c, Tor78b, All75b, All76c, Ano74-234, Ano75-241, Fai74, Fal76a, Too78b, Too78a, TC73a, VCG71]. Memory [Chi73, All77f, Ano74-230, Ano75-239, Ano78-193, Fal74e, Hil74, PG76]. men [McK74]. mere [Gor75, Kap75e]. merely [Gre79a, Tor77e]. merger [Rub76g, War76]. messages [Fal74a]. metal [Mit77, PP70, Yu70]. metal-semiconductor [Yu70]. metals [All79d]. meteorology [Hor76]. meters [All74b]. method [Ano74-293]. methodology [Ano78-39]. methods [Ano78-41, Fal74j, Joh79, Yu75]. metric [And73g]. metropolis [Fri76b, Lin76b, Lin76c]. metropolis-under-one-roof [Fri76b]. Mexico [RK79]. MHD [Lev78, PP70]. MHz [Men76c, Ano79-190]. Micro [All78d, All79g, Ano78-48, Ano79-160, Chr77b, Ano78-165]. Micro-Mouse [All78d, All79g, Chr77b]. Micro-pyramid [Ano78-48]. microcircuit [Ano78-43, Men76e]. microcircuits [Men77d]. Microcomputer [Ano74-226, Ano75-239, Ano79-45, Fal74d, All79f, Ano75-237, Ano75-231, Ano78-201, Ano78-199, Jur78c]. Microcomputers [Ano79-232, Jur78c, Men78c, Mor78, SL78, Ano74-228, Ano79-189, Fal74e, Fal75b, Men77b, Que79]. microelectronic [CT70]. microelectronics [Bre71]. microencounter [Ano78-258]. microfiche [All77f]. microfilm [Men75d]. microfine [All77e]. micrographics [Men75d]. micrometer [HL79]. micromice [All78d, All79g]. Microprocessor [All78b, Ano74-227, Ano76-204, Ano76-205, Fal75f, Ano75-39, Ano75-232, Ano75-243, Ano75-238, Ano78-165, Ano78-198, Ano78-191, Ano79-194, Fal74e, FK76, Hal77, Jur75c, Jur79, New76, Pet77, PMG77, Shi79a, Wic79]. Microprocessor-based [All78b]. Microprocessors [All76c, All78d, All79g, Fag78, Fal76a, Fal77c, Kap78c, Ano78-44, Ano78-45, Ano78-46, Ano78-47, Ano78-48, Ano78-49, Ano78-194, Ano79-53, Ano79-54, Ano79-51, Ano79-49, Ano79-50, Ano79-52, Fal74a, Fal74i, Jur78b, Kap79a, Que79, Ter77, Tor78b]. Microprogrammed [Ano74-228]. Micros [Fal77c, Ano78-49]. microscope [Ano74-222]. microscopy [MVE75]. microuniverse [Sug79b]. Microwave [Ago76, Ano74-44, Gol79, KS79, Men75c, Pog78, SS72, All75i, AFCK79, Ano76-35, Fal75c, HS70, Kap79d, KRS78, Men75a, Men77a, Men78b, Men79a]. Microwaves [Bro79, Fal74b, FB72]. mid [Fal74g]. mid-career [Fal74g]. midair [Lit75]. MIDCON [Ano78-165]. middle [Rab78, Rub76i]. Midwest [Fri70f]. Midyear [Ano77-226]. Mikroelektronik [Ano72-253]. Mile [Mas79e, Rub79b, Okr78]. mile-long [Okr78]. Mileage [Jur75a, HB77, Tor77h]. milestones [Ano74-61]. Military [Dou76, Ele75b, Ele76b, Mas79c, Men74e, SA78, Tor77d, Wal78, Gal77, Mas79b, Tor78a]. Military/avionics [Dou76]. million [Fri75g, Que79, Sug78c]. million-dollar [Fri75g]. mills [HH74, Lev78]. mills/kWh [Lev78]. mind [Tri74]. minded [Men78d]. mini [Ano78-199, Rub75f]. mini- [Ano78-199]. mini-study [Rub75f]. Miniature [Ano74-229, Lin76b, Men78d]. Minicomputer [Ano74-230, Fal74c, HFL74, Jur70, Ano74-224, Ano74-225]. minicomputers [Ano79-189]. minimize [Ano75-243, Par78a]. minimonster [Joe76]. minis [Fal75b]. minorities [Rub75b]. minority [Col78c]. minutia [Gos79b, Hei79]. miracle [Hei79]. mirrors [All76f]. MIS [BS71]. Miscellaneous [Ano73-391]. missile [Joh70a]. mission [Spi76b]. missions [NR74]. mitigate [Ung79]. mix [Ano71-301]. mixed [Ano72-252, Jur77b]. MLS [Pog78]. mm [Tor76a]. Mobile [And73f, SS71, All76a, Lus78]. mobilengineer [Bea74a]. mobility [Fri76i]. mode [Bel74]. model [PM77, Sur72, Lin74a]. Modeling [Fri77a]. models [Bib71c]. modem [Ano75-239]. modems [Ano79-189]. moderate [Har77]. moderate-current [Har77]. modernized [ME70]. modes [Cou79]. modular [HL76]. modulation [Sch70]. module [Ano74-227, Ano74-224]. modules [Nay71b]. molded [Ano74-45]. momentary [Hua70]. Money [RF75, Fal74f]. mongrel [Lin76a]. monitor [Ger74, Rub77f]. monitoring [Car79, Fag79]. Monolithic [Gre72, Kes70, Gre71, Zuc77]. monster [EW78]. Monte [Kap75e]. month [Fal76b, Lin74a, Mas79d]. months [MF79, Rub75d]. mood [Ano72-521]. morality [Ano72-405, Oli72]. MOS [Fet76, Gar70c, HGB79, Lap72b, Tor77f]. MOS/LSI [HGB79, Lap72b]. Moscow [Ano72-403]. Most [McK74, Men78g, Pat79a, PM77, Rub77c, Zuc77]. moths [HS70]. Motivating [Chr79c]. motivation [Tor77g]. motor [Bra77, Kap77a, Kus72, Moo73]. motor-speed [Kus72, Moo73]. motoring [Men78d]. motors [Bia70]. mounted [Fri77c]. mounts [Ano74-224]. Mouse [All78d, All79g, Chr77b]. move [How70]. mover [Bea74a]. movers [Men76a]. Moving [Men78d, Ano74-305, Fal75c, Fri70c, Fri75h, Kop77, Nil76]. Mozambique [KLH74]. Ms. [Gra74]. much [Ano72-318, Ano77-36, Jur75e, Kap76e, Rub76d]. much-maligned [Rub76d]. multichannel [Hil74, Kra75]. multifacets [Tor77e]. multimeter [Ano75-233, Ano75-234]. multimeters [Ano79-191]. multiplex [Bar72]. Multiplexers [Nay71a]. multipliers [Ano78-34]. multipurpose [All74b, Fri76b]. Multispectral [Hor76]. municipal [HH74, Mac75]. municipality [RK79]. museum [Men74b]. mushware [Chr75f]. music [ASM71, Ols71]. must [Ano77-36, Cou79, Ele74c, Mas79a, Rea79, Str71]. musts [Lou74]. mutually [Rei74]. MW [BFM78]. mystery [SR78]. myths [Buc78a, Jur74d, Mid74].

N [Dre74]. nagging [Tuc75a]. name [Chr75i]. named [Ano77-29]. narrative [Rub79a]. NASA [Ano72-404, Ele75b, Fro79, GK76, GHNK79, Rub76k, Spi76a]. nation [Fal74j]. nation-by-nation [Fal74j]. national [Fri73h, Gre79a]. nations [All79c, Fri74c]. nationwide [Bib71a, Bib71b, Bib71c]. natural [Wic79]. nature [GH72]. navigating [KH79]. navigation [Fri70c]. navigational [Fri70b]. Navy [Jur75d, SA78, Wol76c]. NBS [Jur76c]. Nd [All79d]. near [Bar79, Kap78a]. nearing [KH79]. nearly [Rub77d]. nears [Jur79]. necessarily [Rei74, Rub76a]. necessary [Nag74]. need [Bar78, Ele74f, Ele76a, Gup74, Jur74g, Rub77b]. needed [Gos79b, Kay78, Men74c, Rub77c, Rub77d, Hei79]. needs [Bac72, Cor72, Lin76a, ME70, RK79, Rub74d, Rub76d, Ter72]. negative [Bla77]. neglect [TT74]. negligible [MVE75]. neighbors [Jur74e]. NEREM [Fal73b]. Ness [EW78]. network [Ano78-47, ME70, PW71, PG70]. networking [Bel74, Sug78a]. networks [Ano77-59, Bar79, DBB71, Men78f, Mos70c, Sus73]. neutralization [Wal78]. neutron [Til70]. never [Chr76b, Mas79d]. News [Ano70-352, Ano70-353, Ano70-354, Ano70-355, Ano70-356, Ano70-357, Ano70-358, Ano70-359, Ano70-360, Ano70-361, Ano70-362, Ano70-340, Ano70-341, Ano70-342, Ano70-343, Ano70-344, Ano70-345, Ano70-346, Ano70-347, Ano70-348, Ano70-349, Ano70-350, Ano70-351, Ano71-323, Ano71-324, Ano71-325, Ano71-326, Ano71-327, Ano71-328, Ano71-329, Ano71-330, Ano71-331, Ano71-332, Ano71-333, Ano71-334, Ano72-419, Ano72-420, Ano72-421, Ano72-422, Ano72-423, Ano72-424, Ano72-425, Ano72-426, Ano72-427, Ano72-428, Ano72-429, Ano72-430, Ano73-404, Ano73-405, Ano73-406, Ano73-407, Ano73-408, Ano73-409, Ano73-410, Ano73-413, Ano73-411, Ano73-414, Ano73-412, Ano73-416, Ano73-415, Ano73-417, Ano73-418, Ano73-419, Ano73-420, Ano73-421, Ano73-422, Ano73-423, Ano73-424, Ano73-425, Ano73-426, Ano73-427, Ano73-428, Ano74-236, Ano74-237, Ano74-238, Ano74-239, Ano74-240, Ano74-241, Ano74-242, Ano74-243, Ano74-244, Ano74-245, Ano74-246, Ano74-256, Ano74-249, Ano74-247, Ano74-250, Ano74-251]. News [Ano74-252, Ano74-254, Ano74-253, Ano74-248, Ano74-255, Ano74-257, Ano74-258, Ano74-259, Ano74-260, Ano74-261, Ano74-262, Ano74-263, Ano74-264, Ano74-265, Ano74-266, Ano74-267, Ano74-268, Ano75-244, Ano75-245, Ano75-246, Ano75-247, Ano75-248, Ano75-249, Ano75-250, Ano75-251, Ano75-252, Ano75-253, Ano75-254, Ano75-255, Ano75-265, Ano75-256, Ano75-257, Ano75-258, Ano75-259, Ano75-260, Ano75-264, Ano75-261, Ano75-262, Ano75-263, Ano75-267, Ano75-266, Ano75-268, Ano75-269, Ano75-270, Ano75-271, Ano75-272, Ano75-273, Ano75-274, Ano75-275, Ano75-276, Ano75-277, Ano75-278, Ano75-279, Ano76-207, Ano76-208, Ano76-209, Ano76-210, Ano76-211, Ano76-212, Ano76-213, Ano76-214, Ano76-215, Ano76-216, Ano76-217, Ano76-218, Ano76-219, Ano76-220, Ano76-225, Ano76-221, Ano76-226, Ano76-227, Ano76-222, Ano76-223, Ano76-228, Ano76-224, Ano76-230, Ano76-229, Ano76-231, Ano76-232, Ano76-233, Ano76-234, Ano76-235, Ano76-236, Ano76-237, Ano76-238, Ano76-239, Ano76-240, Ano76-241]. News [Ano76-242, Ano77-180, Ano77-181, Ano77-182, Ano77-183, Ano77-184, Ano77-185, Ano77-186, Ano77-187, Ano77-188, Ano77-189, Ano77-190, Ano77-191, Ano77-199, Ano77-192, Ano77-193, Ano77-194, Ano77-198, Ano77-195, Ano77-200, Ano77-196, Ano77-201, Ano77-202, Ano77-197, Ano77-203, Ano77-204, Ano77-205, Ano77-206, Ano77-207, Ano77-208, Ano77-209, Ano77-210, Ano77-211, Ano77-212, Ano77-213, Ano77-214, Ano77-215, Ano78-202, Ano78-203, Ano78-204, Ano78-205, Ano78-206, Ano78-207, Ano78-208, Ano78-209, Ano78-210, Ano78-211, Ano78-212, Ano78-213, Ano78-219, Ano78-214, Ano78-220, Ano78-215, Ano78-216, Ano78-217, Ano78-221, Ano78-222, Ano78-223, Ano78-224, Ano78-218, Ano78-225, Ano78-226, Ano78-227, Ano78-228, Ano78-229, Ano78-230, Ano78-231, Ano78-232, Ano78-233, Ano78-234, Ano78-235, Ano78-236, Ano78-237, Ano79-196, Ano79-197, Ano79-198, Ano79-199, Ano79-200, Ano79-201, Ano79-202, Ano79-203, Ano79-204, Ano79-205, Ano79-206, Ano79-207, Ano79-213, Ano79-214, Ano79-208]. News [Ano79-215, Ano79-216, Ano79-217, Ano79-209, Ano79-212, Ano79-218, Ano79-210, Ano79-211, Ano79-219, Ano79-220, Ano79-221, Ano79-222, Ano79-223, Ano79-224, Ano79-225, Ano79-226, Ano79-227, Ano79-228, Ano79-229, Ano79-230, Ano79-231, And72c, Ano74-208, Ano74-209, Ano74-281, Fri77d, Hor72, Ste75a]. Next [Fri74f, All75h, Jur74h]. NHTSA [Cla77]. niche [Mas79d]. niches [All75c]. nickel [PG76]. nickel-cadmium [PG76]. night [Tor78a]. nighttime [Pla79b]. Nimbus [Hor78]. Nimbus-7 [Hor78]. Nixon [Ano72-454]. nm [Ano70-296]. NMOS [Ver77]. No [Chr79d, Kap75e, RK79, Ano74-229, Lom79, Rub76c, Spi76b, Tuc79]. no-go [Lom79]. Noise [Ken71, LW70, Ber79, Boa73, JM77, Joh71, SS73, Sta72]. nomads [How70]. none [And70]. nonmetals [All79d]. nonobvious [Fra78]. Nontraditional [Kap77e]. nor [McW77]. Northeast [Ano72b]. northwest [KLH74]. nostalgia [Tuc75b]. notable [Fri76g]. notebooks [Der79]. Notes [Jur74a]. nothing [Ano79-232, Spi76b]. Novel [Kle77]. NSF [And72e, Ele74e, Tib78]. Nu [Mas79d]. Nuclear [Chr79e, Fri76d, Gre79c, KJ76, Kap79b, Ras75, RG79, Rip79, Rub76h, Ste77, Sug79c, Ben70, Chr75l, Chr79f, Chr79o, Gan77, Kap75a, Kap75d, Kap76e, Lom79, Rub77f, Rub79b, Sim72b, Mas79e]. number [Ano75-237]. numbers [Ano74-71, Fey74, Rec75]. numerical [Kap77c, Kor77]. nurture [Qua79].

O [Ano75-243, Ano79-192]. Oakland [Fri74f]. Obituaries [Ano70-363, Ano70-364, Ano70-365, Ano70-366, Ano70-367, Ano70-368, Ano70-369, Ano70-370, Ano70-371, Ano70-372, Ano70-373, Ano70-374, Ano71-335, Ano71-336, Ano71-337, Ano71-338, Ano71-339, Ano71-340, Ano71-341, Ano71-342, Ano71-343, Ano72-431, Ano72-432, Ano72-433, Ano72-434, Ano72-435, Ano72-436, Ano72-437, Ano72-438, Ano72-439, Ano72-440, Ano72-441, Ano73-429]. object [Gos79b, Men76a]. objective [Ang74, Bib71a]. objectivity [Chr78d]. objects [MVE75]. observe [Ele74c]. obsolete [Pat79a]. obviate [Rub74d]. obvious [Jur74d, JF74]. occurred [Rub79a]. occurrence [Par78a]. ocean [Fal73d, LZ73, RGCS79]. Octave [Ano70-375]. odds [Wol75b]. OEMs [Ano75-231]. off [All76i, All78d, Ano74-45, Ano74-47, Gib76, GK76, Kal79, Lin74a, Lin76a, MRR^{+}77, Mos75, Pel70, Pla79b, Pla79a, Rub75c, Rub76b, Rub76k, SL78, Sug78c, Tho73a, Wil76b]. off-peak [Pla79b, Pla79a]. offer [Ano74-234, Ano75-235, Gal77, Jur74c, Ung79]. offered [Ano75-241, Ano79-193, RM75]. offering [Rub76a]. offers [Ano74-223, Ano75-232, Ano78-195, Kap75e]. Office [Mok79c, Ano74-209, Ano75-217, Col79, Men77e]. officers [Rub77a]. offing [Jur75e, SZ74]. offspring [Lin75a]. often [Alb78, Par78b, Tri74]. Oil [Fri73f, BKS79, BC77, Nag74]. old [Bla77, Men74f, Rai76, Rub77d, Tuc75b, VT72, Yen78, Yu70]. older [Wil76b, Yu75]. Omega [KH79]. ominous [Ano75-304]. omnipotent [Fri73f]. omnipresent [Fal75d]. On-line [Fir75, Men79a]. on-the-scene [Fri74k]. On-the-spot [FK76]. onboard [NS77]. once [Tor77e]. One [Ano71-344, Mar77, McK73, MVE75, Ano71-119, Ano74-46, Ano75-236, Col78b, Fal75f, Fri76b, Gol72, KLH74, Lin76a, McK74, Pat79b, PM77, RK79, Rub75d, Rub77c, RL74, TT74, Tri78, Gol79]. one-shot [Ano74-46]. oneself [Nil76]. onland [Kap78c]. only [Ele75a, Fai74, Jen78, Jon77b, Rec75]. onto [FR75]. Onward [Fri77c]. op [Ano74-223, Ano75-236, Ano78-197, LPJ76, MM74]. open [Tro74]. opening [Ano79-159]. opens [Ano79-160]. operated [Ano74-230, Ano78-200, Men77b]. operates [Ano78-201, Ano78-197, Fin77]. Operating [Ano75-41, FC78, Chr79o, Rub76j]. operation [Ano75-232, Kal79, Pol71a, TE74]. operational [KH79, Mos70c, Nay71a, Nay71b, Sch71]. operations [Ano71-242, Ano71-243, Ano71-244, Ano71-245, Ano71-246, Ano72-320, Ano72-321, Ano72-322, Ano72-323, Ano72-324, Ano72-325, Ano72-326, Ano72-327, Ano72-328, Ano72-329, Ano72-330, Ano72-331, Ano73-307, Ano73-308, Ano73-309, Ano73-310, Ano73-311, Ano73-312, Ano73-313, Ano73-314, Ano73-315, Ano73-316, Ano73-317, Ano73-318]. operator [Ele74c]. operators [Dun78]. Opinions [Chr79f]. opportunities [Jur74c, Ull70]. opposing [Rub76h]. Optical [BBF^{+}78, Fag79, Fre74, JM77, KS70, Men78b, Men79b, Oli74, Str72, TC73a, TC73b, All75h, And78, Ano77-59, KLM^{+}79, Men76b, MVE75]. optically [Yu75]. optics [Ano77-59, Dou76, Fag79, Fal75c, Mas79b, Men75a, Men76b, Men77a, Men79a, Yu75]. optimize [Kap78b]. optimized [Ano76-204]. optimizing [TE74]. opting [All75b]. option [Bea72e]. Options [FR75, Hus74, Men78e, Ano75-237, Dou76, Fal74g, Fel73, Jur74c, LC75, MB79, Men75b, Men78a, Spo71, Wit77]. orbit [KS79]. orbiter [Wei77]. orbiter-launched [Wei77]. Order [Chr78j]. ordinary [MM74]. Organization [Rub76e, Gra79]. organized [Ele78]. organizers [Lap74e]. Oriental [Rub75b]. oriented [Ano73-306, Ano75-243, MF79]. originals [Men75d]. origins [Shi72]. Orphaned [Men74a]. orphans [Men75b]. Osborne [Ano79-232]. OSCAR [Dun78]. oscillators [Ano74-47]. oscillograph [Kra75]. oscillographs [Kra75]. oscilloscope [Ano79-188]. oscilloscopes [Ano79-190]. OTA [And72a]. Other [Fri74e, Ano75-42, Chr75i, Fal75e, Fri70f, Gre79b, Nil76, Par78b, Rai76, Rea79, Rub74e, Tor79]. otherwise [Fal75c]. ought [MRR^{+}77]. Our [Chr74h, Spo71, Ano74-293, Kap76a, Sug79b, Tan72, Wil76b]. outcome [Bel74, Fri74d]. outlaws [Men78d]. outlined [Pat79b]. Outlook [And72d, BC77, PP70, All79c, Ang74, Ele74e, Gre76b, JF74, KI77, Ste75a]. outnumber [Lin75b]. output [Ano74-293, Ano75-237, Fal74e, Fri76f, Uts77]. ovens [Men75c]. over-ocean [Fal73d]. overall [Kay78]. overcome [Ano77-36]. overdue [Fri74c]. overlook [Pet77]. overseas [Fri77c]. overview [Ano74-208].

Owen [Lin74a]. own [Fal74e, Kap77e]. owner [Men77b]. owner-operated [Men77b]. oxides [Cam77].

PABX [McK72b]. pace

[Ano72-177, Fal76a, Qua79]. paced [Ele76a]. pacemaker [Jur76f]. pacing [Men78c]. package [Ano79-193, Men74d]. Packaged [Ano70-296]. packages [Ano75-238, Fal74c, Men78c]. packaging [All76e, Ana79, Har77, Tor76a]. packet [Rob74a]. packing [Ano79-159]. pacts [All78e]. page [Ano70-199, Ano70-200, Ano70-201, Ano70-202, Ano70-203, Ano70-204, Ano70-205, Ano70-206, Ano70-207, Ano70-208, Ano70-209, Ano70-210, Ano70-211, Ano70-212, Ano70-213, Ano70-214, Ano70-215, Ano70-216, Ano70-217, Ano71-140, Ano71-141, Ano71-142, Ano71-143, Ano71-144, Ano71-145, Ano71-146, Ano71-147, Ano71-148, Ano71-149, Ano71-150, Ano71-151, Ano71-152, Ano71-153, Ano71-154, Ano72-202, Ano72-203, Ano72-204, Ano72-205, Ano72-206, Ano72-207, Ano72-208, Ano72-209, Ano72-210, Ano72-211, Ano72-212, Ano72-213, Ano72-214, Ano73-193, Ano73-194, Ano73-195, Ano73-196, Ano73-197, Ano73-198, Ano73-199, Ano73-200, Ano73-201, Ano73-202, Ano73-203, Ano73-204, Ano73-205]. Painful [Kat75]. pair [Rea79]. pancreas [Spe78]. panel [TG78, UG73, Fri75g]. Paper [Ano79-47, Ano75-304]. Papers [Ano72-227]. parasitics [Tor77e]. Paris [Fri74f]. parity [Chi73]. parries [Men77a]. part [Mar77, Pat79a, Pla79b, RG79, Ano70-296, Chr74l, Chr74m]. partially [Rec75]. participants [Fri75f]. participated [All78d]. participative [Wol79]. particle [Mas79c]. particle-beam [Mas79c]. partly [Fri74j]. partner [Lav78]. parts [Ano75-243]. Pascal [Wic79]. pass [Ano75-38]. passing [Ano74-96]. past [Cot75, Fin72b]. pastiche [Kap78a]. patent [Par78b, Rab78, Rac78]. patentability [Fra78]. patentable [Fra78]. Patents [All78e, Fra78, Par78b]. path [FR75]. Pattern [Ano74-231, Mac76]. patterns [Fal75d]. pave [Car79]. payloads [Wei77]. payment [SZ74]. Payoff [Lev78]. payrolls [Men78c]. PC [Ano75-238, Ano78-42]. pea [Rub77g]. peak [Nay71b, Pla79b, Pla79a]. peaking [CSS76]. Peccei [Fri76e]. peculiarities [Fal74j]. penalties [Fal74g]. penetrate [KS79]. Pension [Bac72, Fin70, BC73b, BCY75]. Pensions [BC73a, BCY75]. People [Ano70-376, Ano70-377, Ano70-378, Ano70-379, Ano70-380, Ano70-381, Ano70-382, Ano70-383, Ano70-384, Ano70-385, Ano70-386, Ano70-387, Ano71-345, Ano71-346, Ano71-347, Ano71-348, Ano71-349, Ano71-350, Ano71-351, Ano71-352, Ano71-353, Ano71-354, Ano71-355, Ano71-356, Ano72-442, Ano72-443, Ano72-444, Ano72-445, Ano72-446, Ano72-447, Ano72-448, Ano72-449, Ano72-450, Ano72-451, Ano72-452, Ano72-453, Ano73-430, Ano73-431, Ano73-432, Ano73-433, Ano73-434, Ano73-435, Ano73-436, Ano73-437, Ano73-438, Ano73-439, Ano73-440, Ano73-441, Ano74-269, Ano74-270, Ano74-271, Ano74-272, Ano74-273, Ano74-274, Ano74-275, Ano74-276, Ano74-277, Ano74-278, Ano74-279, Ano74-280, Ano75-280, Ano75-281, Ano75-282, Ano75-283, Ano75-284, Ano75-285, Ano75-286, Ano75-287, Ano75-288, Ano75-289, Ano75-290, Ano75-291, Ano76-243, Ano76-244, Ano76-245, Ano76-246, Ano76-247, Ano76-248, Ano76-249, Ano76-250, Mau75, Men76a, Ano76-136, Ano76-137, Ano76-138, Ano76-139, Fri72f, Fri75h, Fri76b, Gib76, Kat75]. people [Rub74f, Tor77g]. perceive [Rub75b]. percent [Ano76-206, Her79, Lev78, RK79]. perennial [Eve77]. perfect [Gos79b, Rub78b]. Performance [All76e, All75h, All76f, Ano74-234, Ano75-242, Ano75-235, Fre74, Gre79d, Jon77b, Jur75a, Kap77f, Kau72, Lap73c, Men74d, Pit79, Tor76a, Sne72b]. performer [Sal77]. performers [DTW76]. performs [Rub77d]. perhaps [Fri74j]. perils [Chr78k]. peripherals [Ano74-226, Men77b]. peritoneum [Spe78]. permit [Rub75a]. perseverance [Tro74]. persistence [Rub73b]. persistent [Bla77]. Personal [Ano74-232, Kop77, Men78c, Men79b, RO76, Men78f, Wad76]. personality [Tri78]. Perspective [Ele74e, Fri76e, Jur74a, Men74f, PHC76, Rub76f, Rub76g, Gre74, Ras75, Sur70, Wat71]. persuasive [Lom79]. pervasive [Fal77b, Pet77]. PET [Ano79-54]. petroleum [Hea74]. Ph.D. [Ter72]. Phase [Moo73, Gre71, MVE75, Tib78]. Phase-locked [Moo73, Gre71]. Phased [Ano70-375, Ano70-399]. Phenix [Kap76b]. phenomenon [PG76]. philatelic [Soa70]. philosophies [All78c]. philosophy [MRR^{+}77]. phoenix [Fri74j]. photographic [Fal76b]. photography [Rai76]. Physical [Rie70a, Rie70b, Rie70c, Fal75e]. physicists [Ste77]. physiology [AB71]. picks [Hol72]. picture [Dil75]. piece [Men74b]. piecing [RRS^{+}74]. Piezoelectric [Hol70]. piezoreslstive [Ano78-40]. pilot [Cot75]. pilots [Fri70c]. pin [Ano79-193, PM77]. pinches [Ker76]. pioneer [Ano75-240, McK72a]. pioneers [Tro74]. place [All74d, Her72, Kor77, Mun78, Rub75a, Rub76g]. Placing [Wat71]. planets [Tor78a]. planned [BFM78, Spi76a]. planner [Kat75]. planners [Lin74b]. Planning [Fri77a, Gib70, Kap76c, Lin74b, MF79, Con70, For71, Kat75, Lin76c, Rub77b, Rub77c]. plans [Ano72-255, Fin70, Fri70a, Fri70f, Ish77]. plant [All79e, BFM78, Chr79o, CC78, Fri72i, Fri74b, Fri76d, Fri76h, Kap74b, KJ76, Kap76e, Ura77]. plants [Gan77, Kap75d, Lee71, PP70, RG79, RGCS79, Ura77]. plastic [Mit77]. platter [Men75b]. play [All75h, JF74, Ker76, Men78c, Rub77a, Tor76a, Tor77g]. playback [Ber79]. played [Fri75f, Rub77g]. players [Ano78-242, Men78e]. playing [IFSK77]. plea [Rin70]. Pleasing [IFSK77]. pleasure [Lin76b]. pleasure-dome [Lin76b]. pleasures [Men76a, Rub74c]. plentiful [Men79b]. plotters [Ano78-35]. Plumbing [LZ73]. plummeting [Rub74a]. plunge [Men78f, Sch77b]. plus [Fri76f, New76, Par78b]. plutonium [Ano79-52]. Pocket [Ano79-53]. pocketbook [Rub76c]. point [Ano70-325, Ano71-297, Ano71-298, GHNK79]. points [Ano70-248, Ano70-249, Ano70-250, Ano70-251, Ano70-252, Ano70-253, Ano70-254, Ano70-255, Ano70-256, Ano70-257, Ano70-258, Ano70-259, Ano71-182, Ano71-183, Ano71-184, Ano71-185, Ano71-186, Ano71-187, Ano71-188, Ano71-189, Ano71-190, Ano71-191, Ano71-192, Ano71-193, Ano72-256, Ano72-257, Ano72-258, Ano72-259, Ano72-260, Ano72-261, Ano72-262, Ano72-263, Ano72-264, Ano72-265, Ano72-266, Ano72-267, Ano73-243, Ano73-244, Ano73-245, Ano73-246, Ano73-247, Ano73-248, Ano73-249, Ano73-250, Ano73-251, Ano73-252, Ano73-253, Ano73-254, Ano74-121, Ano74-122, Ano74-123, Ano74-124, Ano74-125, Ano74-126, Ano74-127, Ano74-128, Ano74-129, Ano74-130, Ano74-131, Ano74-132, Ano75-118, Ano75-119, Ano75-120, Ano75-121, Ano75-122, Ano75-123, Ano75-124, Ano75-125, Ano75-126, Ano75-127, Ano76-104, Ano76-105, Ano76-106, Ano76-107, Ano76-108, Ano76-109, Ano76-110]. Poised [Fal76a]. polarizing [Ano79-145]. policies [Rub74a]. Policy [Rub76h, Fri73h, Hef70, Hol72, Rub78b, Tri78]. policy-making [Rub76h, Tri78]. Polish [Ano75-240]. Political [Man79, Gor75]. politician [Fri75e]. politicians [Rub78b]. politics [Fri74c, Gre79a, Lom79, Wad76]. polluting [Cla77, Hea74]. pollution [Bib71a, Bib71b, Bib71c, Fri70e, Fri70f, Hor78, Jur75a, PMG77, Sta72, Sto77]. pollution-control [Fri70f]. ponderous [Fal75c]. pool [Ano78-46]. pools [Hap73]. poor [Kap75b, KRS78]. Popov [Jur74a]. popular [Ano79-194]. popularity [All79d]. population [AB70, Kap77d]. portable [All78c, Ano75-234, Ano79-191, FK76, Men74d, PM77]. portables [All75d, All79e]. pose [Chr74j]. position [Ano75-304, BKS79, Cle73, Rub76h]. positions [Ano74-209, Ano75-217, Ano76a, Mac78]. positive [Ano78-197, Bel74, Cle73, Fri70b]. positive-control [Fri70b]. possibilities [Rec75]. possible [All76a, Bre75, Col78b, Fal74f, Mac76, Moo79, Rub77c, Ura77]. Postal [Chr78l]. potential [Col79]. Power [Alb78, All75e, AFCK79, Ano75-303, BDL^{+}78, BFM78, BKS79, Bro79, CSS76, Chr74a, Col79, Cou79, CC78, EH76, Ewa78, Fri70e, Fri70f, Fri71b, Fri72i, Fri74c, Fri75b, Fri76h, Fri76f, Fri76g, Fri77b, Fri77c, Fri77a, GHM71, GHNK79, Hap73, Har77, Jay76, Joh79, JF74, Kap74b, Kap77e, Kap77d, Kap78a, Kap79c, KLH74, LC75, Lev78, Lia78, Men74a, Mic78, Mid74, Par78a, Pla79b, Pla79a, RK79, RF75, RGCS79, Sch78, Sug78c, Yen78, Ago76, All75i, And73a, Ano74-41, Ano74-44, Ano74-43, Ano74-227, Ano75-43, Bel74, Ben70, Bro73, BCH^{+}79, CT70, Chr75l, Chr78c, Chr78e, Chr79e, Chr79f, Ele74e, Eng75, Fag79, Fal74e, FC78, Fri72f, Fri74b, Fri74e, Fri75f, Fri76d, Gre79c, Hua70, Jur74c, Kap75b, Kap75d, Kap76c, KJ76, Kap76e, Kap79b]. power [Ker76, LZ73, Lin75a, Nag74, Okr78, PP70, Ras75, Reu78, RG79, Rip79, Rub79b, SS72, Sug79c, TE74, Ura77, WZ78]. Power-generation [Ano75-303]. power-transfer [Cou79]. Power/Energy [Bro79, AFCK79, BDL^{+}78, BFM78, Col79, Cou79, CC78, Fri74c, Fri75b, Fri77b, Fri77c, GHNK79, Joh79, Kap77e, Kap78a, Kap79c, LC75, Lia78, Mic78, Pla79b, RK79, RGCS79, Sug78c]. Power/finance [RF75]. powerful [All75a, Reu78]. Practical [HL79, Fri77c, Mau75, Men74g]. Practice [Ano71-247, Ang74, Woo79]. practices [Fri74d, Rea79]. practices/ethics [Fri74d]. practitioner [Der79, Fri75d]. praise [Chr78d, Raf78]. precedent [Fri74d]. precise [KLV75]. precision [Ano74-39]. predicted [Too78b]. predicting [Gan77]. prediction [Fri73b, KS79]. Preemphasis [Ano75-42]. preference [PMG77]. pregrooved [BCH^{+}79]. prelude [Gos79a]. premission [NR74]. premium [Men78b]. preparation [Lin74a]. prepared [Men75d]. Pres. [Rub76c]. presage [Ele74e]. prescription [Fri73g]. present [All74c, Cot75, Fri70a, KRS78, Kin71, Sug79a]. presented [Ste75b]. presenting [Buc78a]. presents [Tuc79]. preserves [Kra75]. presettable [Ano78-195]. President [Fri76e, Rub76c, Ste75a, Ano72-454, Ano75-304, Gra70c, Gre79b]. Presidential [Ano76-270]. Press [Ano75-154, Ano75-155, Ano75-156, Ano75-157, Ano75-158, Ano75-159, Ano75-160, Ano75-161, Ano76-140, Ano76-141, Ano76-142, Ano76-143, Ano76-144, Ano76-145, Ano76-146, Ano76-147, Ano76-148, Ano75-162, Chr79n]. pretrigger [Hil74]. prevention [Chr77c]. preview [All74a, Ano78-165, Ano78-258, Ano79-159, Ano79-160, Ele78, Ell72]. prey [Rub77g]. price [All75b, Ano74-225, Ano75-233, Ano78-164, Ano79-191, Jur77b, Men78f]. priced [Ano76-206, Fal75f]. prices [All76e, Ano74-232, Ano75-235]. primary [All76i, Hus74, Too78b]. prime [Dou76]. Primer [Ano75-43]. principal [All74d, Ele74c, Kap77b, TE74]. Principles [Ano74-45, Ano74-42, Ano75-41]. printed [Fre74, She79]. printers [Ano78-35]. prints [Fal76b]. priorities [Ano72-388, Fal74a]. priority [Ele75b, Moo79]. private [Kap77a, Sch77b]. privately [All76h]. privately-funded [All76h]. Prize [Ano79-47]. probe [All76i, All79c, All79h, Ano78-49, Col78c, EW78, Wol77]. probes [BDL^{+}78, Rub76h]. Probing [All79g]. problem [Ano74-232, Bib71a, Chr77k, Fag78, Gra70a, Str71, Tea74, Tuc75a, Ull70, Yen78]. problem-solvers [Ano74-232]. Problems [Ano77a, EBC^{+}79, Fri74c, Fri76g, All74c, AB70, Ele74b, Ern77a, Fri70a, Jur75c, Kap77d, Lap74e, Lin77, Mas79f, Mok79c, Rie70a, Rie70b, Rie70c, Rub76c, Sug79c, Ter77, Tor76d]. procedures [Chr79o, Kin75]. proceeding [Fri76j]. Proceedings [Chr74m]. process [All79e, Ano76-204, Ano77-29, Eve77, Gro72, Rub76d]. processed [And79]. processing [ATK72, And79, Ele72d, Fal76b, KM71]. processor [Ano75-239, Lap72b]. procurement [Ano74-95]. produce [Fal76b]. produced [KLH74, McK74, RS77]. producing [Jur76a]. Product [Ano78-238, Ano78-239, Ano79-233, Ano79-234, Lap74d, Men74f, Ano70-328, Ano70-329, Ano70-330, Ano70-331, Ano70-332, Ano70-333, Ano70-334, Ano70-335, Ano70-336, Ano70-337, Ano70-338, Ano70-339, Ano71-311, Ano71-312, Ano71-313, Ano71-314, Ano71-315, Ano71-316, Ano71-317, Ano71-318, Ano71-319, Ano71-320, Ano71-321, Ano71-322, Ano72-406, Ano72-407, Ano72-408, Ano72-409, Ano72-410, Ano72-411, Ano72-412, Ano72-413, Ano72-414, Ano72-415, Ano72-416, Ano72-417, Ano73-392, Ano73-393, Ano73-394, Ano73-395, Ano73-396, Ano73-397, Ano73-398, Ano73-399, Ano73-400, Ano73-401, Ano73-402, Ano73-403, Ano74-230, Ano74-227, Ano74-233, Ano74-235, Ano74-234, Ano74-224, Ano74-226, Ano74-222, Ano74-232, Ano74-229, Ano74-231, Ano74-228, Ano74-225, Ano74-223, Ano75-239, Ano75-232, Ano75-233, Ano75-237, Ano75-234, Ano75-240, Ano75-231, Ano75-241, Ano75-243, Ano75-238, Ano75-242, Ano75-236, Ano75-235, Ano75-230, Ano76-204, Ano76-194]. product [Ano76-195, Ano76-205, Ano76-196, Ano76-197, Ano76-198, Ano76-199, Ano76-200, Ano76-206, Ano76-201, Ano76-202, Ano76-203, Ano77-178, Ano77-168, Ano77-169, Ano77-170, Ano77-171, Ano77-172, Ano77-173, Ano77-179, Ano77-174, Ano77-175, Ano77-176, Ano77-177, Ano78-194, Ano78-198, Ano78-201, Ano78-199, Ano78-192, Ano78-200, Ano78-191, Ano78-197, Ano78-195, Ano78-190, Ano78-196, Ano78-193, Ano79-145, Ano79-192, Ano79-193, Ano79-188, Ano79-190, Ano79-191, Ano79-194, Ano79-185, Ano79-186, Ano79-189, Ano79-195, Ano79-187]. Production [Ano75-241, All79f, Ano74-95, Fal75e, Jur76a, Lap74b, Men78g, Mun78, RG79, Sug78b, Tor76c, Tor77b]. productive [Hag71]. Productivity [Ano78-240, Ano78-242, Ano78-241, Ano78-243, Ano78-244, Chr78g, Sug78d, All78a, All79a, Bar78, BJ78, Buc78a, Col78c, Fri76h, Kap76d, Kap78b, Men78g, Mun78, Nas78, Wei78a]. Products [Ano73-442, Ano76-251, All77g, Ano72-418, Ano72-520, Ano74-71, Jur78a, Lap74e, Moo79, Wul78]. profession [Mas79d, Rub75b]. Professional [Ano70-388, Ano70-389, Ano70-390, Ano70-391, Ano70-392, Ano70-393, Ano70-394, Ano70-395, Ano70-396, Ano70-397, Ano70-398, Ano71-357, Ano71-358, Ano71-359, Ano71-360, Ano71-361, Ano71-362, Ano71-363, Ano71-364, Ano71-365, Ano71-366, Ano71-367, Ano72-455, Ano72-456, Ano72-457, Ano72-458, Ano72-459, Ano72-460, Ano72-461, Ano72-462, Ano72-463, Ano72-464, Ano72a, Ano73-443, Ano73-444, Ano73-445, Ano73-446, Ano73-447, Ano73-448, Ano73-449, Ano73-450, Ano73-451, Ano73-452, Ano73-453, Ano77-216, Ano77-217, Ano77-218, Ano79-235, Col78b, Fri74d, Ano75-153, Ano79-190, Gol73, Rub76b, Ste75a, Ste75b]. professional-grade [Ano79-190]. professionalism [You73]. professionals [Cam77, How70]. Profile [Fri74e, Ano77-178]. Profiles [Rub73b]. profit [And73e]. prognoses [Fri76e]. prognosticate [Lin77]. Program [Ano70-448, Ano71c, Ano71-419, Ano72b, Ano72-517, Ano73-454, All74a, All74d, Ano72-519, Ano73-306, Ano74-226, BFM78, Bar78, BHLK71, BC77, Ele74e, Esp77, Fal74e, GHNK79, Gle77, Kun75, Lap73h, Lap74e, Man70, Tib78, Ano71a, Ano72a, Ano78-164, Ele78]. Programmable [Ano74-233, Ano74-41, Ano75-41, Ano78-200, Spe78]. programmers [Men78f]. programming [Ano78-193, Que79]. Programs [Ano75-153, BDL^{+}78, Buc78b, Ele74b, Ele76b, Kap75e, Rub74f, RF75, Rub77a, Wic79]. Progress [Fri74f, Kap79c, KLE73, Mas79b, All77e, Ano75-153, Ano77a, Ano78-94, BDL^{+}78, BKS79, Chr79k, Fal75c, Fal76a, Fal77b, Fri74c, Fri75h, Fri76g, IY77, Jur75c, Kap77d, Kap78a, Pay75, SOW77, Ern77b]. progresses [Ele76b]. progressing [Men74g]. progressive [Fri74e, Pet77]. progressively [DTW76]. project [Ano78-46, RS77]. Projecting [Que79]. projection [Max77, RG79]. projections [Kap74a]. projects [Fri76j, Gra79, Men74e]. proliferate [All75c, Doi77]. proliferation [Men75b]. prologue [Fin72b]. promise [Ber79, Cou79, Ele75a, Ele78, Lin75a, Mil71, Reu78, SL78, Uts77]. promised [Sug78a]. promises [Cla77, EBC^{+}79, Kap78a, SR78]. promising [Ano76-273, Jur75c, Lap74c]. promotes [Esp77]. prompted [Ano74-95]. PROMs [All75b]. pronged [BC77]. proof [Chr75j]. propagation [Ber71]. proper [Fal74a, Rin70]. Properties [Sne72a]. property [Sug78c]. proponents [Kap76c]. proportions [Men76c]. proposal [Ano74-95]. propose [SL78]. proposed [Ano76a, Fri70f, Kin75, Okr78]. proprietary [Rin70]. Prospects [Wad76, Ben70, Pet77, Rub74b, RF75]. prostheses [Jur78b]. Protect [Rac78]. protection [Fri75g, Men74f, Par78a, Par78b, Rac78]. protective [Kap79c]. prototype [Ano78-194, BFM78]. prototypes [Tor76c]. prove [Jur74d]. proven [All76h]. provide [And79, Ano79-145, GHNK79, KS79, Lap74a, ME77, Yen78]. provided [Ano74-208, Fal75f]. provides [Ano78-194]. provisions [RO76]. provocative [Rec75]. psychology [Rub77g]. public [All76b, Chr75l, Cor72, Fag79, Fri74d, Fri75e, Gor75, Her72, Kap75b, Kap77a, Kat75, Men78e, Pon70, Rub77f, SR78, Sug78c, Sug79c, Tri78, Men76f]. publications [Ano70-169, Ano70-412, Ano70-413, Ano70-414, Ano70-415, Ano70-416, Ano70-417, Ano70-418, Ano70-419, Ano70-420, Ano70-421, Ano70-422, Ano70-423, Ano71-120, Ano71-242, Ano71-243, Ano71-244, Ano71-245, Ano71-246, Ano71-383, Ano71-384, Ano71-385, Ano71-386, Ano71-387, Ano71-388, Ano71-389, Ano71-390, Ano71-391, Ano71-392, Ano71-393, Ano71-394, Ano72-320, Ano72-321, Ano72-322, Ano72-323, Ano72-324, Ano72-325, Ano72-326, Ano72-327, Ano72-328, Ano72-329, Ano72-330, Ano72-331, Ano72-481, Ano72-482, Ano72-483, Ano72-484, Ano72-485, Ano72-486, Ano72-487, Ano72-488, Ano72-489, Ano72-490, Ano73-307, Ano73-308, Ano73-309, Ano73-310, Ano73-311, Ano73-312, Ano73-313, Ano73-314, Ano73-315, Ano73-316, Ano73-317, Ano73-318, Ano73-474, Ano73-475, Ano73-476, Ano73-477, Ano73-478, Ano73-479, Ano73-480, Ano73-481, Ano73-482, Ano73-483, Ano73-484, Ano73-485, Ano74-180, Ano74-304, Ano74-295, Ano74-296, Ano74-297, Ano74-298, Ano74-299, Ano74-300, Ano74-301]. publications [Ano74-302, Ano74-303, Ano75-182, Ano75-183, Ano75-184, Ano75-185, Ano75-186, Ano75-305, Ano75-306, Ano75-307, Ano75-308, Ano75-309, Ano75-310, Ano75-311, Ano75-312, Ano75-313, Ano75-314, Ano75-315, Ano75-316, Ano76-161, Ano76-162, Ano76-163, Ano76-164, Ano76-165, Ano76-166, Ano76-263, Ano76-264, Ano76-265, Ano76-266, Ano76-267, Ano76-268, Ano76-269, Ano77-160, Ano77-161, Ano77-162, Ano77-163, Ano77-164, Ano77-165, Ano77-166, Ano77-167, Ano78-154, Ano78-155, Ano78-178, Ano78-179, Ano78-180, Ano78-181, Ano78-182, Ano78-183, Ano78-184, Ano78-185, Ano78-186, Ano78-187, Ano78-188, Ano78-189, Ano79-156, Ano79-157, Ano79-173, Ano79-174, Ano79-175, Ano79-176, Ano79-177, Ano79-178, Ano79-179, Ano79-180, Ano79-181, Ano79-182, Ano79-183, Ano79-184]. published [Chr74k, Chr74l, Chr74m]. publishing [Chr78k]. Pulling [Rub76c]. Pulsars [Til70]. pulse [Ano74-235, EBC^{+}79]. pump [Jur75a]. punches [Rub76c]. purpose [All77c]. pursue [Men76a]. push [Ele74c, Fri77c]. pushes [Fri75b]. pushing [Bac72]. put [Gos79b, Ras75, Tor79]. Putting [Fey74, Joh79]. pyramid [Ano78-48].

Q [Men76f]. Q&A [BCY75]. quad [Jur72d]. quality [Ele74a, Wei78a]. quantified [Ras75]. quantum [Tor79]. quarter [Rub75d]. quartz [von79]. question [PHC76, Tuc75a]. questionable [Han79, Mas79c]. Questions [Ano72-454, BCY75, Col78b, Lap74d]. quick [Rub76e]. quietly [Mun78]. quit [Chr74i]. quitting [Dib72].

R&D [And72d, Ano75-304, Gre79b, PHC76, RS77, Rub79c, Wol75b]. race [Men78d]. rack [Rub74a]. Radar [Ano70-296, Men78d, SK73, Tor76d]. Radars [Tor77d, Joh70a]. radiation [Hor78, Jur74a, Tel72]. radical [Col78a]. Radio [McK74, SS71, And73f, Dun78, Fal77c, Fri71c, Gor75, Kap79d, LC75, Lus78, Men78b]. radios [Jur78a, Men77a, PM77]. Rail [Fri74g, Fri74h, Fri74i, Fri74j, Fri75c, Fri74f, Fri74k, Fri75e]. Railroad [Fri72j, Chr79d, Mas79f]. Railroads [Ish77, Fri75c]. rails [Fri74h, Fri74i, Fri75a]. railway [Fri74h]. railways [Fri74g]. raise [Mun78]. raiser [Ele74a]. raises [Lap74d, Rec75]. RAM [Ano78-198, Gos79a]. RAMs [All75b, Fet76]. random [VCG71]. random-access [VCG71]. range [And79, Ano79-191, EBC^{+}79, Jur75e, Rub77b, Rub77c]. ranges [All74a, Fri74g]. ranging [Fal75f, MRR^{+}77]. rap [Men78d]. Rapid [Fri74d, She79, Hor72, Nag74, Wad76]. rapidly [Ele74c, Rub77d]. Rasmussen [Chr76i]. rate [Ano74-235, LC75, Mil72]. rates [AC75, Bre75, Fri74j, Kap79c, Pla79b, Pla79a]. rather [Men78c]. rational [CC78]. RCA [Jen78]. reach [Jur76f]. reaches [Men78b]. Reaching [Fal76c, Fal74b]. reactions [Fri70c]. Reactor [Chr79g, Fri73e, Fri74b, Ras75]. read [Ano78-193, Fai74, Fre74, McW77]. read-only [Fai74]. read/write [Ano78-193]. Reader [Ano70-198, Ano70-182, Ano70-183, Ano70-184, Ano70-185, Ano70-186, Ano70-187, Ano70-188, Ano70-189, Ano70-190, Ano70-191, Ano70-192, Ano70-193, Ano70-194, Ano70-195, Ano70-196, Ano70-197, Ano71-134, Ano71-135, Ano71-136, Ano71-137, Ano71-138, Ano71-139, Ano72-190, Ano72-191, Ano72-192, Ano72-193, Ano72-194, Ano72-195, Ano72-196, Ano72-197, Ano72-198, Ano72-199, Ano72-200, Ano72-201, Ano73-182, Ano73-183, Ano73-184, Ano73-185, Ano73-186, Ano73-187, Ano73-188, Ano73-189, Ano73-190, Ano73-191, Ano73-192, Ano74-71, Ano76-59, Rub75a, RM75]. readers [Jen78, Rub75f]. reading [Ano78-241]. readout [Men74d]. readouts [All75c]. ready [Spi76b]. Real [Ano79-192, Chr74c, Chr74d, Chr74e, Chr74f, Ele74b, Fal74a, Fal76c, Gre79b, Jur74e, Jur75c, Jur76a, Lav78, Man72, Rub74a]. Real-time [Ano79-192, Lav78]. real-world [Fal76c]. realize [Cou79]. really [Hea74, SOW77, Tuc74]. reasonable [Ano74-225, Lou74]. reasons [Jur76a]. Receiver [Ano70-296]. Receivers [Ano70-296, Tuc75b]. recession [Rub76b]. recognition [Ano74-231, Cla71, Fre74]. recognizing [Mac76]. recommend [Ano79-95]. recommendation [Cor72]. Recommended [Ano71-247, Lin75b, Mid74]. recommends [Fri76h, Wei78b]. reconsidered [Chr79g]. record [All75f, Pla79b]. recorded [RRS^{+}74]. recorder [Ano79-145]. recorders [Jur74b]. recording [Bre75, BCH^{+}79, Hil74, Ken71, Kra75, Men78e, Sal77]. recordings [Ols71]. recovery [Hus74]. rectify [Lin74b]. rectifying [AFCK79, Bro79]. red [Ano72-522, Gra79]. Reddy [Fri72c]. redesign [Fri73c]. reduce [All78c, Jur75b]. Reduced [Men75d, Pla79a, Ano78-164, Ano79-159]. reduces [Ano76-205, Dal70]. reducing [Fri70f]. reduction [Men74a]. reductions [Rub76j]. redundancy [Kap74b]. reentry [Rub76k]. reevaluating [Rub74a]. reexamined [Ano75-303]. reference [Bea74b, Fir75, GHNK79, KRS78, Rub76e]. references [Lin75b]. Refining [Eng75]. Reflection [DeF70]. reflects [Ano72-388, Pie79]. reform [Bac72, Par78b]. Refuse [HH74]. Region [Ano73-413, Ano73-411, Ano73-414, Ano73-412, Ano73-416, Ano73-415, Ano74-256, Ano74-249, Ano74-247, Ano74-250, Ano74-251, Ano74-252, Ano74-254, Ano74-253, Ano74-248, Ano74-255, Ano75-265, Ano75-256, Ano75-257, Ano75-258, Ano75-259, Ano75-260, Ano75-264, Ano75-261, Ano75-262, Ano75-263, Ano76-219, Ano76-220, Ano76-221, Ano76-222, Ano76-223, Ano76-224, Ano77-192, Ano77-193, Ano77-194, Ano77-198, Ano77-195, Ano77-196, Ano77-197, Ano78-219, Ano78-214, Ano78-215, Ano78-216, Ano78-217, Ano78-218, Ano79-208, Ano79-209, Ano79-212, Ano79-210, Ano79-211]. Regional [Ano74-281]. regions [Ano75-266, Ano76-225, Ano76-226, Ano76-227, Ano76-228, Ano76-229, Ano77-199, Ano77-200, Ano77-201, Ano77-202, Ano77-203, Ano78-220, Ano78-221, Ano78-222, Ano78-223, Ano78-224, Ano78-225, Ano79-213, Ano79-214, Ano79-215, Ano79-216, Ano79-217, Ano79-218, Ano79-219]. registers [Ano78-196]. Registration [Chr78h, Chr75e, Fin77]. regulations [Ano77-36]. regulators [Kes70, Men76c]. Regulatory [Mas79e]. reimbursed [PHC76]. rejecting [Wol76a]. related [AB70, Ele74f, Fle71, Lia78]. relationships [Ste75b]. relatively [Fri74e]. Relay [Ano70-399, Ano75-242]. reliability [AC75, Alb78, All76i, Fri77a, Gal77, Gan77, Jur74g, Jur78c, Kap79c, New76, SL78, Wel79]. reliable [FR75, NR74, Sac77, TE74, VHM77]. relieve [RRS^{+}74]. rely [Fre74]. remains [Rub77d, Sch77b]. remarriage [Chr78c]. remember [Dre74]. reminded [Eve77]. remote [Hor76, Hor78, Pla79b]. rendezvous [Tor78a]. Rent [Rub77e]. Rent-an-instrument [Rub77e]. repair [All75f, Cla77]. repetitive [RRS^{+}74]. replace [Ano78-200, KRS78]. replacement [von79]. replaces [Ano75-238, Jur78a, KLM^{+}79]. Report [Ano76-252, Chr77f, Dil75, Sch77a, Ste75a, All74d, All75f, All75g, All77g, All77e, Ano74-109, Ano74-110, Ano74-111, Ano74-112, Ano74-113, Ano74-114, Ano74-115, Ano74-116, Ano74-117, Ano74-118, Ano74-119, Ano74-120, Ano74-305, Ano75-44, Ano75-106, Ano75-107, Ano75-108, Ano75-109, Ano75-110, Ano75-111, Ano75-112, Ano75-113, Ano75-114, Ano75-115, Ano75-116, Ano75-117, Ano75-217, Ano75-317, Ano76-37, Ano76-92, Ano76-93, Ano76-94, Ano76-95, Ano76-96, Ano76-97, Ano76-98, Ano76-99, Ano76-100, Ano76-101, Ano76-102, Ano76-103, Ano76-270, Ano76-272, Ano77-28, Ano77a, Ano77-84, Ano77-85, Ano77-86, Ano77-87, Ano77-88, Ano77-89, Ano77-90, Ano77-91, Ano77-92, Ano77-93, Ano77-94, Ano77-95, Ano77-226, Ano78-115, Ano78-116, Ano78-117, Ano78-118, Ano78-119, Ano78-120, Ano78-121, Ano78-122, Ano78-123, Ano78-124, Ano78-125, Ano78-126, Ano79-114, Ano79-115, Ano79-116, Ano79-117, Ano79-118, Ano79-119, Ano79-120, Ano79-121, Ano79-122]. report [Ano79-123, Ano79-124, Ano79-125, Bea74b, Bel74, BKS79, Chr74n, Chr75m, Chr75a, Chr79k, Col78c, Fai74, Fal74c, Fri74k, Fri75e, Fri75g, Fri76a, Gre79a, Jur74a, Kap78a, Lap74e, Lin74a, Men74a, Mok79a, Rob74a, Rub77b, Rub77c, Tro74]. reported [Gal79]. Reports [Ano72-332, Gra70b, Mul71, Ano74-208, Ano75-153, Ano76-135, Dil75, Gle77, Rub79a]. Representatives [All79c]. reproduction [ASM71]. Republic [Jan70]. require [AFCK79, Fin77, HB77, Shi79a]. requirements [Ano75-243, Bib71b, PMG77, Sne72b, Gar70a]. requires [Wei77]. rescue [Ano78-243, Sug78d]. Research [Ano72b, BDL^{+}78, Gre79b, Hus74, Esp77, Fal76c, Jur75c, Kap79c, Lav78, Mas79b, Mos75, Okr78]. Researching [Tuc75a]. resident [EW78]. residential [RO76]. resides [Fri74h]. residual [All77f]. resolution [And79, Uts77]. resource [HH74, Hor76, Hus74]. resources [All75f, All79h, CSS76, Ele76b]. Respond [Ano72-454]. responded [Rub75f]. responds [Gol79]. response [RRS^{+}74, Ste75b]. responsibilities [Rob74b]. restrictions [Mas79a]. result [CC78, Tor76c]. results [Ele74c, Gal79, Kap77f, Rub75a, Rub75b]. retail [SZ74]. retinal [Uts77]. retirement [Her79]. retracking [FR75]. Retrieval [BBG75, BCH^{+}79, Gup74, Jen78, Men75d, Tuc75a]. retrieves [Ano78-196]. retriggerable [Ano74-46]. Retrodirective [Ano70-399]. retrofit [Dou76]. Retrospective [Dre74, Fri75d, Gin75, Mau75, Wol76b, Tro74]. return [Ano78-41, Lin74a]. Returns [Rub75c]. Revamping [And72e]. reveal [Der79, Rub74c]. revealing [ME77]. reveals [Chr79o, Fri74k, Jen78, Jur74f]. review [All75b, All76i, All79c, Ano74-122, Ano74-123, Ano74-124, Ano74-125, Ano74-126, Ano74-127, Ano74-128, Ano74-129, Ano74-130, Ano74-131, Ano74-132, Ano74-62, Ano74-63, Ano74-64, Ano74-65, Ano74-66, Ano74-67, Ano74-68, Ano74-69, Ano74-70, Ano74-71, Ano75-118, Ano75-119, Ano75-120, Ano75-121, Ano75-122, Ano75-123, Ano75-124, Ano75-125, Ano75-126, Ano75-127, Ano75-58, Ano75-59, Ano75-60, Ano75-61, Ano75-62, Ano75-63, Ano75-64, Ano75-65, Ano75-66, Ano76-104, Ano76-105, Ano76-106, Ano76-107, Ano76-108, Ano76-109, Ano76-110, Ano76-51, Ano76-52, Ano76-53, Ano76-54, Ano76-55, Ano76-56, Ano76-57, Ano76-58, Ano76-59, Ano77-118, Ano77-44, Ano77-45, Ano77-46, Ano77-47, Ano77-48, Ano77-49, Ano77-50, Ano77-51, Ano77-52, Ano77-53, Ano77-54, Ano78-63, Ano78-64, Ano78-65, Ano78-66, Ano78-67, Ano78-68, Ano78-69, Ano78-70, Ano78-71, Ano78-72, Ano78-73, Ano79-67, Chr76a, Fag79]. review [Fet76, Fin77, Kap76a, Sel71]. reviewed [Ano77a]. reviews [Ano70-170, Ano70-171, Ano70-172, Ano70-173, Ano70-174, Ano70-175, Ano70-176, Ano70-177, Ano70-178, Ano70-179, Ano70-180, Ano70-181, Ano71-122, Ano71-123, Ano71-124, Ano71-125, Ano71-126, Ano71-127, Ano71-128, Ano71-129, Ano71-130, Ano71-131, Ano71-132, Ano71-133, Ano72-178, Ano72-179, Ano72-180, Ano72-181, Ano72-182, Ano72-183, Ano72-184, Ano72-185, Ano72-186, Ano72-187, Ano72-188, Ano72-189, Ano73-169, Ano73-170, Ano73-171, Ano73-172, Ano73-173, Ano73-174, Ano73-175, Ano73-176, Ano73-177, Ano73-178, Ano73-179, Ano73-180, Ano74-49, Ano74-50, Ano74-51, Ano74-52, Ano74-53, Ano74-54, Ano74-55, Ano74-56, Ano74-57, Ano74-58, Ano74-59, Ano75-47, Ano75-48, Ano75-49, Ano75-50, Ano75-51, Ano75-52, Ano75-53, Ano75-54, Ano75-55, Ano75-56, Ano76-39, Ano76-40, Ano76-41, Ano76-42, Ano76-43, Ano76-44, Ano76-45, Ano76-46, Ano76-47, Ano76-48, Ano76-49, Ano76-50, Ano77-37, Ano77-38, Ano77-39, Ano77-40, Ano77-41, Ano77-42, Ano78-50]. reviews [Ano78-51, Ano78-52, Ano78-53, Ano78-54, Ano78-55, Ano78-56, Ano78-57, Ano78-58, Ano78-59, Ano78-60, Ano78-61, Ano79-55, Ano79-56, Ano79-57, Ano79-58, Ano79-59, Ano79-60, Ano79-61, Ano79-62, Ano79-63, Ano79-64, Ano79-65, Ano79-66, FS77, GJH77, JE77, KH77, OGG77, SC77, SB75, TSR74, WB75]. revisited [All77f, Ano75-318, Ano79-235, Chr76i, EW78, Ele74f]. revival [Fri72j]. revolution [Ano74-96, Tro74]. revolutionary [Ano78-164]. RF [Tor76a, Wal78]. Ribbon [Fri75g]. ride [Fri74f]. rider [Men76a]. rides [Fri75a]. ridiculed [Wol76a]. ridiculous [Wol76a]. Riding [Fri74i, Fri74k, Lin74a, Mas79f]. right [Fri72j, Gre74]. rights [Fri74i, Rin70]. rights-of-way [Fri74i]. ring [All77a, Gib76]. Ripple [Sug78c, LC75, Rub75e]. rise [Tor76c]. rises [Fri74j]. Risk [Chr79j, Chr79h, Chr79i, Gan77, Pay75, Sug79c, Wil71, Wul78]. risks [Ras75, Tri74]. road [Max77, Obe77, Sal72b, You73]. robotic [Ano79-50]. Robots [Doi77, Mun78, All79a]. rocketed [Rub76k]. rocketry [Qua79]. role [All75h, Ele73e, Fri70f, HGB79, Ker76, Lia78, Nag74, Pon70, Rub77a, Tor77g]. roles [Fai74, JF74, VT72]. rolling [Jur78a]. rolls [Obe77]. ROM [Ano78-198]. romance [Tuc75b]. Rome [Fri76e]. ROMs [All75b]. roof [Fri76b]. Room [Fri70a]. roost [Tor77d]. rooting [Lin75a]. roots [Rub76g]. rough [Ano75-240]. roundup [Men74b]. route [Fri70b]. routes [EH76]. RTL [Gar70a]. rubble [Fri74j]. Rugged [Sac77]. rule [Jur72c, Men74b]. Rules [Ano71-247]. rumble [Lin74a]. run [Chr79d, EH76, Har72]. runaway [All78c]. running [Ano74-209, Ano75-217, Fri74k]. Russell [Gin75].

sabotage [Ras75]. safe

[Chr74j, Gol79, Tel72]. safer [Bra77, Cla77]. Safety [Cla77, Men73a, Men74f, All77d, Ano75-303, Bra77, Chr79g, Cot75, Ele75a, Fri74b, Fri76d, Jur77a, Kap75d, KJ76, Kap79a, Kin75, Lee71, Ras75, Sug79c, Tor77h, Wad76]. safety-i [KJ76]. saga [Fri74i]. sagging [Men74c]. Salaries [Ano73-455, McK71, Rub74a, Rub75c, Sal73]. salary [Rub75a]. SALT [Mas79a]. Sam [Rub76a]. same [Men74d]. Sample [Nay71b]. Sample/hold [Nay71b]. sampler [Kap77e]. samples [Yu75]. sampling [Fal75f]. San [Fri74f]. Sarnoff [McK72a]. Satellite [Bro73, GB73, Joh79, Okr78, Rub76i, AC75, Fal77c, Hor76, Kap79d, Rai76, Reu78, Rub76k, AFCK79, GHNK79]. Satellite/energy [Okr78]. Satellites [AC75, Ano70-399, Bro79, GK76, Rub76i, Bar79, Dun78, Fal74h, McK72c, Qua79, Rub76k, VHM77]. satisfaction [Jur74g, Men78g]. satisfy [PMG77]. Saunders [Rub76h]. save [BC77, Hea74, HFL74, Kap77b]. saved [Fal74f, Kap77c]. saver [Ele74a]. savers [Ano71-420]. saving [Jur76d]. saw [Fri75h]. say [Fra78, Rea79]. says [Kin75, Rub76a]. SB [Ano79-236]. scale [Ano77-36, Fai74, HH70, Lap74d, Mac78, Moo79, TG78, Tor78b]. scanned [Rub76g]. scanner [Jur79, Lap74c]. Scanning [Ano70-400, Ano70-401, Ano70-402, Ano70-403, Ano70-404, Ano70-405, Ano70-406, Ano70-407, Ano70-408, Ano70-409, Ano70-410, Ano70-411, Ano71-368, Ano71-369, Ano71-370, Ano71-371, Ano71-372, Ano71-373, Ano71-374, Ano71-375, Ano71-376, Ano71-377, Ano71-378, Ano71-379, Ano72-465, Ano72-466, Ano72-467, Ano72-468, Ano72-469, Ano72-470, Ano72-471, Ano72-472, Ano72-473, Ano72-474, Ano72-475, Ano72-476, Ano73-456, Ano73-457, Ano73-458, Ano73-459, Ano73-460, Ano73-461, Ano73-462, Ano73-463, Ano73-464, Ano73-465, Ano73-466, Ano73-467, Ano74-293, Ano74-282, Ano74-283, Ano74-284, Ano74-285, Ano74-286, Ano74-287, Ano74-288, Ano74-289, Ano74-290, Ano74-291, Ano74-292, Ano75-303, Ano75-292, Ano75-293, Ano75-294, Ano75-295, Ano75-296, Ano75-297, Ano75-298, Ano75-299, Ano75-300, Ano75-301, Ano75-302, Ano76-253, Ano76-254, Ano76-255, Ano76-256, Ano76-257, Ano76-258, Ano76-259, Ano76-260, Ano77-219, Ano77-220, Ano77-221, Ano77-222, Ano77-223, Ano77-224, Ano77-225, Ano78-245]. Scanning [Ano78-246, Ano78-247, Ano78-248, Ano78-249, Ano78-250, Ano78-251, Ano78-252, Ano78-253, Ano78-254, Ano78-255, Ano78-256, Ano79-237, Ano79-238, Ano79-239, Ano79-240, Ano79-241, Ano79-242, Ano79-243, Ano79-244, Ano79-245, Ano79-246, Ano79-247, Ano79-248, Fal73c, KRS78, Uts77]. scenario [Fal75d]. scene [Fri74k]. schedule [Ish77]. scheme [Ano72-317, Dav79, Fri76b, Lia78, Mas79c]. schemes [SL78, Shi70, Sug79a, Yen78]. Schneider [Rub76j]. school [Col78b]. schoolers [Bea72e]. schools [Pat79a]. Schottky [Ray77]. Science [Ano71-380, Ano71-381, Ano71-382, Ano72-477, Ano72-478, Ano72-479, Ano72-480, Ano73-468, Ano73-469, Ano73-470, Ano73-471, Ano73-472, Ele73f, Han79, Hef70, Sal73, Chr79l, Ele74f, Fri70d, Fri75d, Soa70, EW78]. Science/Scope [Ano71-380, Ano71-381, Ano71-382, Ano72-477, Ano72-478, Ano72-479, Ano72-480, Ano73-468, Ano73-469, Ano73-470, Ano73-471, Ano73-472]. scientific [Ano74-232, Ele74f]. scientists [Mas79c]. scoop [Spi76b]. Scope [Ano71-380, Ano71-381, Ano71-382, Ano72-477, Ano72-478, Ano72-479, Ano72-480, Ano73-468, Ano73-469, Ano73-470, Ano73-471, Ano73-472, Rac78]. scopes [All74b, OP76]. Scoping [All79h]. Scoreboard [Che73]. Scores [Rub75f, Hal77, Kap76b]. SCPC [Ano74-307]. scrambling [Sug79a]. screen [Men79b]. scribing [All79d]. scrutiny [Jur74d]. sea [Kap78c, RGCS79]. sealed [Tor76d]. search [Bla77, EW78, Fri74c, Jay76, Spi76a]. searching [Fir75]. seas [Cou79]. season [Men77d]. seasons [DTW76]. seat [Jur75c, Lin74a]. second [Ano72-405, Ano79-235, Chr78h, Kal79, Kap79d, RO77, Sug78a]. second-generation [Kal79]. seconds [Fri75c]. Secrecy [Ele74f, Ele73f, Gol74, Sal73]. secret [Mas79d]. Secretary [Fri76i]. section [Rac78]. security [EK71, Fri77a, Lia78, TE74, Ano79-236]. security-related [Lia78]. See [All79g, Ano73-303, Chr76c, Kap76c, Mau75, Tor77d]. Seeking [Chr74c, Chr74d, Chr74e, Chr74f, Fri75b, Kap75d, Kap76e, Rub74d]. seeks [Ele75b, Tri74]. seem [Tuc74]. seemingly [Jur74d]. seems [Fal74b, Kap76b]. seen [Fri75c, Kin75, Que79]. segments [Rub75e]. select [Zuc77]. selected [Bea74b]. selecting [Dou76]. selection [Ano77-29, Bia70, Pay75, RM75]. selector [Ano78-238, Ano78-239, Ano79-233, Ano79-234]. Self [Ano78-198, Fal74e, Men77d, All74c, Jur74g]. self-checking [Jur74g]. Self-contained [Ano78-198, Fal74e, Men77d]. self-testing [All74c]. Semiconductor [All75b, Jur78c, VCG71, GHM71, SOW77, War76, Yu70]. Semiconductors [All77e, Lin77, All75e, All77f, Buc72, Har77]. Senator [Ano72-454]. senior [BDL^{+}78]. sense [Jay76]. sensing [Hor78, KS79]. Sensitive [Mac76, Chr74b, Wol77]. Sensitivity [Ano70-296]. sensor [Ano79-192, Jur75c]. sensor/actuator [Jur75c]. Sensors [Jur76d, All76f, Hor76, Jur75e, Kap79a, Rub76i]. sentiments [Pie79]. September [All74d]. sequence [Jur75e]. series [Ano79-195, DTW76, WZ78]. serious [Lap74e, Mok79c]. served [Ano71-119]. serves [Fal74a, Rub76e]. Service [Ano70-198, Ano79-47, Chr78l, All78c, Ano70-182, Ano70-183, Ano70-184, Ano70-185, Ano70-186, Ano70-187, Ano70-188, Ano70-189, Ano70-190, Ano70-191, Ano70-192, Ano70-193, Ano70-194, Ano70-195, Ano70-196, Ano70-197, Ano71-134, Ano71-135, Ano71-136, Ano71-137, Ano71-138, Ano71-139, Ano72-190, Ano72-191, Ano72-192, Ano72-193, Ano72-194, Ano72-195, Ano72-196, Ano72-197, Ano72-198, Ano72-199, Ano72-200, Ano72-201, Ano73-182, Ano73-183, Ano73-184, Ano73-185, Ano73-186, Ano73-187, Ano73-188, Ano73-189, Ano73-190, Ano73-191, Ano73-192, Ano74-71, Ano76-59, Gor75, PW71, Reu78, Rub76a, Rub77d, VHM77]. serviceability [All74c]. Services [SS71, Ano70-388, Ano70-389, Ano70-390, Ano70-391, Ano70-392, Ano70-393, Ano70-394, Ano70-395, Ano70-396, Ano70-397, Ano70-398, Ano71-357, Ano71-358, Ano71-359, Ano71-360, Ano71-361, Ano71-362, Ano71-363, Ano71-364, Ano71-365, Ano71-366, Ano71-367, Ano72-455, Ano72-456, Ano72-457, Ano72-458, Ano72-459, Ano72-460, Ano72-461, Ano72-462, Ano72-463, Ano72-464, Ano73-443, Ano73-444, Ano73-445, Ano73-446, Ano73-447, Ano73-448, Ano73-449, Ano73-450, Ano73-451, Ano73-452, Ano73-453, Ano75-153, Ano77-216, Ano77-217, Ano77-218, Car79, Lav78, Man79, Sug78c]. servicing [All74c]. session [BDL^{+}78]. sessions [All77a, Ano76-260, Ano77-225, Ele78]. set [Ano75-232, Ano75-234, Fri74d, RF75, Ste75b]. sets [Ano72-177, Rub76h]. setting [Fal74a]. settles [Mic78]. Seven [Lin77, TT74, Lin74a, Lus78]. seventies [Ano70-325, Hef70, Jur70]. Several [Lia78, Lap74c, Pat79b]. severe [Fal74g]. sewing [Men76f]. shafts [HL79]. shape [Fal74j, Fri76f]. shaped [Ano74-96, Fal76c]. shapes [Ano73-473]. shaping [Fal75d]. Shapp [Fri75e]. share [Ano78-48]. sheets [MM74]. shelf [Bea74b]. shell [Rub77g]. Shield [Ano70-326]. shielding [Ano76-36]. shift [Ano72-388]. Shinkansen [Ish77]. shipping [Ras75]. shock [All76b, Dal70, Dal72]. shopping [All76g, Jur74b, Zuc77]. short [Ano75-43, Buc78b, Fal75f, Pat79a, SA78]. short-circuit [Ano75-43]. short-term [Buc78b]. shortages [Doi77]. shortcomings [Mid74]. shortcut [Tuc79]. shortfall [Pat79a]. shot [Ano74-46, Buc78b]. should [PHC76, Rub74d, Rub77a]. shouldn’t [Eve77, Rub79a]. Show [Ano72-491, All76d, Ano72-389, Ano76-251, Nas78, Rub74b, Uts77]. showing [RG79]. shows [Rub77g]. shrink [All75d]. shrinks [Wel79]. shutdown [Ura77, WZ78]. Shuttle [And72f, Ano79-249, Ele75b, Fal73a, MF79, Wei77]. Shuttles [Ele76b]. side [Chr76j, New76]. sideband [Lus78]. sidetracked [Ano76-273]. Signal [KM71, CF72, Jur75e, Nay71a, Nay71b, Rot71]. signal-sequence [Jur75e]. signaling [Gal77]. signals [ASM71, BCG73, Fri74g, Fri74h, Ish77, Mac76]. Signature [All78c]. Sigurd [Gin75]. Silicon [All75h]. silver [Men75b]. Simplicity [All75d]. simplify [Mok79c]. simply [Men74c]. simulation [Fri77a, Rei74]. Simultaneous [Hil74]. Since [Dun78]. Single [Ano78-199, Lus78, PM77, Que79, Ano75-232, Ano78-201, Ano78-197, Ano79-193, Ano79-190]. single- [Ano79-190]. Single-board [Ano78-199]. Single-chip [Que79]. Single-sideband [Lus78]. Single-slice [PM77]. single-step [Ano75-232]. site [CC78]. site-specific [CC78]. sites [Col79]. siting [CC78, Fri72i]. situation [Lin74b]. Six [Bla77, Rub75d, TG78]. size [Ano75-242, JM77, Jur77a, Men75d, Wel79]. sizes [Men78e]. skepticism [All77f]. skeptics [Men78g]. skill [Men76e]. skilled [Der79]. skills [Gre79d]. sky [AC75, Hor76, LPJ76]. slice [Hei79, PM77]. slick [Fri74k]. slide [Men74b]. slipping [Sco73]. slow [Bea72f]. slowly [All79b]. small [Ano72-524, Kap76c, Men74b, Raf78, Tib78]. small-business [Tib78]. Smaller [All76e, Fet76, Jon77b, Jur78c, All75a, Gar77, Men74b]. smallest [Ano79-190]. smart [All74b, All78d]. smarter [All76e, All79e, Doi77, Jur78c, Kap75c, Men74b]. smartest [All79g]. smoke [Jur76d]. smoothly [Fri74k]. smug [Gol79]. snuggled [Ele74d]. soar [Her79, Jur79, MF79, Tor77d]. soaring [Men74e]. Social [Ern77a, Han79, Rob74b]. socially [Ele74b]. Societal [Rub74e, Str71]. Societie [Ano72-454, Rub76g]. societies [Rub76f, Ung79, Rub76f]. Society [Ano70-310, Ano71-266, Ano71-267, Ano71-268, Ano71-269, Ano71-270, Ano71-271, Ano72-359, Ano72-360, Ano72-361, Ano72-362, Ano73-350, Ano73-351, Ano73-352, Ano73-353, Chr75j, Chr76k, Gre79a, Hag71, Jon77a, Lap74e, Mas79d, Rub76b, Rub76i, Rub77f, Rub78b, Rub79c]. Society-proof [Chr75j]. sociotechnical [Bar73, Rub74e]. Sociotechnology [Ano75-304, Ele74f, Gol79, Han79, Ste75b, Tuc74]. Soft [Chr75k, Fai74]. Software [Jur76e, All76b, Ano79-45, Chr75f, Ele76a, Fai74, Fal74d, Fal74c, Fal74e, Fal75b, Fal77b, Fra78, Kap75c, Lia78, MNS77b, Men77b, Men77a, Men78c, Men78e, Mok79b, Par78a, Que79, Ter77, Wic79]. Solar [Bro79, Col79, Joh79, Ano78-46, Ele74e, Kap75b, Kap76c, Okr78, AFCK79, GHNK79]. solar-heating [Ano78-46]. soldier [Tor78a]. Soleri [Lin76b]. Solid [All77f, All79h, Ano73-473, Ano78-94, Ano78-200, Gos79b, Kal79, Kus72, LJ70, Lap73f, Mac78, Men76b, Moo79, Shi79a, SOW77, Too78b, Too78a, Tor77f, Tor78b, War76, All77e, Ano74-47, Jur75b, Lap74c, Men75a, Sim72a, Tor79]. Solid-state [All77f, Ano78-94, Ano78-200, Kal79, Kus72, LJ70, Men76b, Tor77f, Tor78b, All77e, Ano74-47, Lap74c, Sim72a, Tor79]. solution [Fag78, Fri73e, Jon77b]. solutions [FK76, HB77, JF74, Tuc79, Yen78]. solve [Mas79f]. solved [Ter77]. solvent [Chr77k]. solvers [Ano74-232]. Solving [Ele74b, Lap74e, Str71]. Some [Chr74i, Moo79, Rea79, Ung79, All78d, Ano75-240, Col78c, Ele74f, Fal74f, Fri70c, Fri70f, Fri74i, Fri76j, Fri77d, Jur74d, Men75b, Men77e, Mid74, Mok79c, New76, Rub74a, Rub76j]. Something [All74a, Mau75]. sometimes [Bea74a]. somewhere [Ano76-273]. sonar [EW78]. soon [Ano77-36, Mas79d, Men76e]. Sophisticated [Gan77, Ele76a, Fri74g, Kap75e]. sophistication [All75d]. sort [Rub74b]. sought [Bar78]. sound [KLV75, Pie79, Ste75a, GPPM71]. source [Bro73, Eng75, Fri73f, LZ73, Rub74b, Tuc75a, Wal70]. sources [Ano74-39, BKS79, Dar70, Fri74c, Jur75e, Kap76c, Kap77e, KS70, Oli74, SS72, Tri74]. southeastward [Eng75]. Soviet [Jur74a].

Space [Hor76, NR74, For70, Hav70, Hor78, MF79, Mas79b, Mos75, Okr78, Pol79, And72f, Ano79-249, Ele75b, Ele76b, Fal73a, MF79, Wei77]. space-age [Mos75]. space-oriented [MF79]. Spanish [Rub75b]. Spanish-surnamed [Rub75b]. spare [Rub75f]. spark [Jur74h]. sparks [And73f]. spawn [All78b, Men75b]. spawned [Lap74d]. spawns [VL77]. speak [Ano76-270]. speaks [Ano75-304, Fri74a, Fri76i, Rub74c, She79]. spec [MM74]. Special [All74d, All75f, All75g, All77g, Ano70-412, Ano70-413, Ano70-414, Ano70-415, Ano70-416, Ano70-417, Ano70-418, Ano70-419, Ano70-420, Ano70-421, Ano70-422, Ano70-423, Ano71-383, Ano71-384, Ano71-385, Ano71-386, Ano71-387, Ano71-388, Ano71-389, Ano71-390, Ano71-391, Ano71-392, Ano71-393, Ano71-394, Ano72-481, Ano72-482, Ano72-483, Ano72-484, Ano72-485, Ano72-486, Ano72-487, Ano72-488, Ano72-489, Ano72-490, Ano72-491, Ano73-474, Ano73-475, Ano73-476, Ano73-477, Ano73-478, Ano73-479, Ano73-480, Ano73-481, Ano73-482, Ano73-483, Ano73-484, Ano73-485, Ano74-294, Ano74-304, Ano74-295, Ano74-296, Ano74-297, Ano74-298, Ano74-299, Ano74-300, Ano74-301, Ano74-302, Ano74-303, Ano74-305, Ano75-305, Ano75-306, Ano75-307, Ano75-308, Ano75-309, Ano75-310, Ano75-311, Ano75-312, Ano75-313, Ano75-314, Ano75-315, Ano75-316, Ano75-317, Ano76-261, Ano76-262, Ano76-263, Ano76-264, Ano76-265, Ano76-266, Ano76-267, Ano76-268, Ano76-269, Ano76-270, Ano78-257, Fri75e]. Special [Lap74e, Lin74a, Rub79b, Tor77c, All75c, All77c, Ano71-229, Ano71-230, Ano71-231, Ano71-232, Ano71-233, Ano71-234, Ano71-235, Ano71-236, Ano71-237, Ano71-238, Ano71-239, Ano71-240, Ano72-304, Ano72-305, Ano72-306, Ano72-307, Ano72-308, Ano72-309, Ano72-310, Ano72-311, Ano72-312, Ano72-313, Ano72-314, Ano72-315, Ano73-291, Ano73-292, Ano73-293, Ano73-294, Ano73-295, Ano73-296, Ano73-297, Ano73-298, Ano73-299, Ano73-300, Ano73-301, Ano73-302, Ano74-146, Ano74-147, Ano74-148, Ano74-149, Ano74-150, Ano74-151, Ano74-152, Ano74-153, Ano74-154, Ano74-155, Ano74-156, Ano75-140, Ano75-141, Ano75-142, Ano75-143, Ano75-144, Ano75-145, Ano75-146, Ano75-147, Ano75-148, Ano75-149, Ano75-150, Ano75-151, Ano76-125, Ano76-126, Ano76-127, Ano76-128, Ano76-129, Ano76-130, Ano76-131, Ano76-132, Ano76-133, Ano76-272, Ano78-145, Ano78-146, Chr74a, Chr75a, Mok79a]. specific [CC78]. specifications [HB77]. Specs [Kap74a]. Spectral [Ano70-424, Ano70-425, Ano70-426, Ano70-427, Ano70-428, Ano70-429, Ano70-430, Ano70-431, Ano70-432, Ano70-433, Ano70-434, Ano70-435, Ano71-406, Ano71-395, Ano71-396, Ano71-397, Ano71-398, Ano71-399, Ano71-400, Ano71-401, Ano71-402, Ano71-403, Ano71-404, Ano71-405, Ano72-492, Ano72-493, Ano72-494, Ano72-495, Ano72-496, Ano72-497, Ano72-498, Ano72-499, Ano72-500, Ano72-501, Ano72-502, Ano72-503, Ano73-486, Ano73-487, Ano73-488, Ano73-489, Ano73-490, Ano73-491, Ano73-492, Ano73-493, Ano73-494, Ano73-495, Ano73-496, Ano73-497, Ano75-318, Ano76-271, Ano77-226, Chr74c, Chr74d, Chr74e, Chr74f, Chr74g, Chr74h, Chr74b, Chr74i, Chr74j, Chr74k, Chr74l, Chr74m, Chr74n, Chr75c, Chr75d, Chr75b, Chr75e, Chr75f, Chr75h, Chr75i, Chr75l, Chr75k, Chr75j, Chr75g, Chr75m, Chr76h, Chr76e, Chr76d, Chr76k, Chr76j, Chr76f, Chr76g]. Spectral [Chr76b, Chr76i, Chr76c, Chr76l, Chr77a, Chr77b, Chr77g, Chr77h, Chr77c, Chr77i, Chr77j, Chr77d, Chr77f, Chr77e, Chr77k, Chr78a, Chr78c, Chr78b, Chr78e, Chr78j, Chr78d, Chr78k, Chr78g, Chr78l, Chr78h, Chr78f, Chr78i, Chr79a, Chr79b, Chr79c, Chr79e, Chr79d, Chr79f, Chr79g, Chr79j, Chr79h, Chr79i, Chr79k, Chr79l, Hor78]. spectrograms [Opp70, Yu75]. Spectrum [Chr79k, Fal76b, SS71, And78, Ano78-192, Fal74b, Gre76a, All78d, Ano70-182, Ano70-183, Ano70-184, Ano70-185, Ano70-186, Ano70-187, Ano70-188, Ano70-189, Ano70-190, Ano70-191, Ano70-192, Ano70-193, Ano70-194, Ano70-195, Ano70-196, Ano70-197, Ano70-198, Ano71-113, Ano71-134, Ano71-135, Ano71-136, Ano71-137, Ano71-138, Ano71-139, Ano72-164, Ano72-190, Ano72-191, Ano72-192, Ano72-193, Ano72-194, Ano72-195, Ano72-196, Ano72-197, Ano72-198, Ano72-199, Ano72-200, Ano72-201, Ano73-181, Ano73-182, Ano73-183, Ano73-184, Ano73-185, Ano73-186, Ano73-187, Ano73-188, Ano73-189, Ano73-190, Ano73-191, Ano73-192, Ano74-60, Ano74-61, Ano74-62, Ano74-63, Ano74-64, Ano74-65, Ano74-66, Ano74-67, Ano74-68, Ano74-69, Ano74-70, Ano74-71, Ano75-57, Ano75-58, Ano75-59, Ano75-60, Ano75-61, Ano75-62, Ano75-63, Ano75-64, Ano75-65, Ano75-66, Ano75-319, Ano76-51, Ano76-52, Ano76-53, Ano76-54, Ano76-55]. Spectrum [Ano76-56, Ano76-57, Ano76-58, Ano76-59, Ano77-43, Ano77-44, Ano77-45, Ano77-46, Ano77-47, Ano77-48, Ano77-49, Ano77-50, Ano77-51, Ano77-52, Ano77-53, Ano77-54, Ano78-84, Ano78-85, Ano78-86, Ano78-87, Ano78-88, Ano78-89, Ano78-90, Ano78-91, Ano78-92, Ano78-93, Ano78-62, Ano78-63, Ano78-64, Ano78-65, Ano78-66, Ano78-67, Ano78-68, Ano78-69, Ano78-70, Ano78-71, Ano78-72, Ano78-73, Ano79-83, Ano79-84, Ano79-85, Ano79-86, Ano79-87, Ano79-88, Ano79-89, Ano79-90, Ano79-91, Ano79-92, Ano79-93, Ano79-94, Ano79-67, Chr75a, Fri76a, Rub75f, Rub76h]. Spectrum/Computer [All78d]. Speech [Opp70, She79, Yu75]. speed [All74b, Ano74-225, Ano78-197, Gib76, Kap79a, Kus72, Moo73, Tor77h, VL77]. speeders [Men78d]. Speeding [RJD73]. speeds [Hor78, Men78e, Mit77, SZ74, Yu75]. spend [Men74e]. spin [Mos75]. spin-off [Mos75]. spinning [HL79]. Spirit [And72f]. spite [Fri74e]. sponsored [KH79]. spooks [PG76]. Spot [Reu78, FK76]. sprang [Fri74j]. sprawling [Wil76b]. SPS [AFCK79]. spur [All79a, IY77]. spurs [All78a]. squared [War76]. squares [Sor70]. squelch [Jur75a]. stability [Fri77a, Gef70, TE74, Wei78a]. stack [Fri70f]. staff [Ano71-299, Ano71-300, BDL^{+}78, Rub76e]. staffed [Men74f]. stage [Ano78-165, Jay76, Rub77g]. stages [DTW76]. stagnate [DTW76]. staid [Pay75]. stakes [Chr78j]. stall [Jur75c]. stamp [Chr78l]. stamps [Ano73-218]. Standard [Jur75b, Ano74-233, Ano75-236, CF72, Ele76a, HL76, Kap75c]. standardization [New76, Pog78]. Standards [Mid71, Ano74-168, Ano74-158, Ano74-159, Ano74-160, Ano74-161, Ano74-162, Ano74-163, Ano74-164, Ano74-165, Ano74-166, Ano74-167, Ano75-163, Ano75-164, Ano75-165, Ano75-166, Ano75-167, Ano75-168, Ano75-169, Ano75-170, Ano75-171, Ano75-172, Ano75-173, Ano75-174, Ano76-149, Ano76-150, Ano76-151, Ano76-152, Ano76-153, Ano76-154, Ano76-155, Ano76-156, Ano76-157, Gol79, Ano71-248, Ano71-249, Ano71-250, Ano71-251, Ano71-252, Ano72-333, Ano72-334, Ano72-335, Ano72-336, Ano72-337, Ano72-338, Ano72-339, Ano72-340, Ano72-341, Ano72-342, Ano72-343, Ano72-344, Ano72-345, Ano72-346, Ano73-319, Ano73-320, Ano73-321, Ano73-322, Ano73-323, Ano73-324, Ano73-325, Ano73-326, Ano73-327, Ano73-328, Ano73-329, Ano73-330]. Standing [Ano74-208]. Stanford [PG70]. star [Fal77c, Men78c]. stars [Mac76, Til70]. start [Bea72f]. starts [Jur75b, Pay75]. startup [Ura77]. State [All79h, Cam77, Moo79, All77f, All77e, Ano73-473, Ano74-47, Ano78-94, Ano78-200, Fin77, Fin70, Gos79b, Hil74, Jur75b, Jur77c, Kal79, Kus72, LJ70, Lap73f, Lap74c, Mac78, Mas79a, Men75a, Men76b, Shi79a, Sim72a, SOW77, Too78b, Too78a, Tor77f, Tor78b, Tor79, War76]. state-of-the-art [Mas79a]. state/computers [Shi79a]. stately [Lin76b]. Statements [Ano71-407, Ano72-504, Ano73-498, Ano74-209, Ano75-217, Ano76a]. States [AC70, Fel73, HH71]. statesmanship [Dre74]. stations [All76a, AFCK79, Bro73]. statistics [All77a, Rub74b]. status [KH79, Kin71, Rub75b]. statutory [Par78b]. steady [Fal77b]. Steerable [Ano70-375]. steering [Tuc74]. stem [BB71]. stems [Fri75g]. step [Ano75-232, BFM78, CF72, DTW76, Pat79b, PG76, Tuc74, Wul78]. Stepper [Bia70]. steps [Chr79o, Fal75e, GHNK79]. stereo [Men78a, Men79b]. still [Fra78, Jur79, Sug78b]. stimulate [Buc78b]. stimulating [Bar78]. stock [Fro79]. STOL [SA78]. STOLports [Fri70b]. stop [SL78]. storage [Ano79-195, CSS76, Fri76g, Hus74, Pla79a, TC73b]. Store [RL74]. Stored [Kun75, And79, RGCS79]. Stored-program [Kun75]. stores [Ano77-179, Ano78-196]. story [Rub76b]. strain [Ano78-40, FC78]. straining [Men76c]. Strategic [Mas79a]. strategies [Sch78]. stress [Bra77, FC78]. stressed [Ano79-159]. stretcher [Fal74b]. strides [Sug78b]. stripe [SZ74]. strive [Lap74b]. stroke [Sug78c]. Strong [Jur74h]. stronger [Jur74h]. structure [Hin79, Mas79b, Rub77c]. structures [All79h, BS71, Joh79]. struggle [Fri75b, Jur74f, Rub78b]. struggles [Kap76c]. student [Eve77]. Students [Rub74f, BB71, Jur74c, Man70]. studies [AB71]. study [Bea74b, Cam79, Fri74b, Fri74e, Rub75f]. style [Bea73a]. styling [Bra77]. submarine [GPPM71]. submicron [Gos79b]. subordinate [Bel74]. subpicoampere [Ano74-223]. substandard [Pat79a]. substantially [Ano78-164]. substitute [Mar77]. subsystems [Tor77a]. suburb [Wil76b]. suburb-cities [Wil76b]. suburbs [Gib76]. succeed [Str71]. succeeds [Gre74]. success [Ber73, Jur75d, Tib78, Tor77g]. Successful [Fin73, Fri75e, SA78, Wul78]. successfully [All77f]. Such [Alb78]. suffice [Kay78, Kop77]. suggest [Rub75c]. suggested [Fag78]. suggestions [Men77e]. suggests [Fal75d, Rub77c]. suit [Fri74d]. summary [Gle77]. sun [Col79, Kap75b]. sun-derived [Kap75b]. Sunbelt [Wil76b]. Sunny [Kap75b, Ele74e]. Super [IY77, Bro79, Jon77a, Kap75d, OP76]. super-driver [Jon77a]. Superbatteries [BKS79, Bea72b]. Supercommunications [Fri72k]. Superconducting [Ana79]. Superconductivity [VT72]. supergrowth [Lin77]. superhet [PM77]. superior [Bel74]. superior-subordinate [Bel74]. superiority [Jur74f]. superpools [Hap73]. supersnobs [Wal74]. supertechnician [Men74c]. supplies [Ano74-41, Hua70, Men74a, RK79]. supply [Ano78-197, Fal74e, FR75, Fri76f, GHM71, Kap78b, Pat79a, Rub74b]. supply/demand [Rub74b]. support [Bra79, Esp77, Fri76b, HL79, Wei77]. supported [Bar78]. supports [Ano79-194]. supposed [Jur76f]. suppression [GH70]. Supurbia [Lin76c]. surface [Fri76i, KM71]. surface-wave [KM71]. surnamed [Rub75b]. surprise [Jur74e]. surprises [Kat75]. surprising [MM74]. Survey [Ano71-241, Rub75b, Rub75f, Bar79, Chr77g, Ele72c, Jen78, ME77, Mos70a, Mos70b, Par78a, Rub75c, Rub75d, Rub75a]. survive [Ung79]. survived [Bra77]. suspects [Fri75e]. Sweden [Fri74k]. Swedish [Chr79e]. swim [Chr75g]. Swiss [Fri74g, Fri74i]. Switch [Dor72, Ano78-200, Fal75a, Jed73, Kap74b]. switches [LC75]. Switching [Pol79, Sel71, Ano77-178, Gal77, Kun75, LJ70, Pay75, Pit79, Shi76]. Switzerland [Fri74g, Fri74i]. SX [Lap73b]. SX-70 [Lap73b]. symbol [Ano78-37]. Symposium [Ano71b]. synchronous [Ano74-235]. Synonym [SOW77]. synthesis [Ols71, She79]. Synthetic [FCR^{+}70]. System [Ano71-247, Lia78, Ago76, Alb78, All75a, All77c, Ang74, Ano74-226, Ano74-222, Ano74-231, Ano75-41, Ano75-239, Ano76-205, Ano78-201, Ano79-192, Ano79-194, BFM78, Bib71a, Bro79, CSS76, Chr75d, DBB71, Der79, EK71, Ewa78, Fal74e, Fin77, Fin73, Fri74g, Fri74j, Fri74k, Fri75g, Gre79d, Kap77a, Kap77f, KH79, Kau72, Kay78, KRS78, Kun75, Lap74a, Lou74, Mas79f, Par78a, Pie79, Rab78, RK79, RL74, Ste73, Sug79c, TE74, Too78b, Wei77, All75f, Ano79-95, Ker76, Rub76d]. Systems [Bib71a, Bib71b, Bib71c, For71, Men74g, AC75, Alb78, All75d, All75i, All76i, All78b, AFCK79, Ang74, Ano74-95, Ano74-234, Ano75-39, Ano75-43, Ano78-43, Ano79-194, ASM71, BBF^{+}78, Ber71, Chr74j, Chr78f, Der79, Doi77, Ele75a, Ele76b, Fag79, Fal75b, Fal76a, Fal77c, FC78, Fri70a, Fri70b, Fri70c, Fri70f, Fri76j, Ged72, Jur71, Jur74h, Jur75b, Jur76d, Jur78b, Jur78c, Kal79, Kap77a, Kap79c, Kap79d, KRS78, KLH74, KLV75, Kor77, Kun75, LJ70, LC75, Lev78, Lin76c, MNS77a, Mas79b, NR74, Ohi70, Pog78, SL78, Sch78, Ter77, TE74, TT74, Tor77h, Uts77, Fre74].

Table [Ano70-436, Ano70-437, Ano70-438, Ano70-439, Ano70-440, Ano70-441, Ano70-442, Ano70-443, Ano70-444, Ano70-445, Ano70-446, Ano70-447, Ano71-408, Ano71-409, Ano71-410, Ano71-411, Ano71-412, Ano71-413, Ano71-414, Ano71-415, Ano71-416, Ano71-417, Ano71-418, Ano72-505, Ano72-506, Ano72-507, Ano72-508, Ano72-509, Ano72-510, Ano72-511, Ano72-512, Ano72-513, Ano72-514, Ano72-515, Ano72-516, Ano73-499, Ano73-500, Ano73-501, Ano73-502, Ano73-503, Ano73-504, Ano73-505, Ano73-506, Ano73-507, Ano73-508, Ano73-509, Ano73-510, Ano74-306, Ano74-230, PM77]. Tables [Ano73-511, Ano73-512, Ano70b, Ano70c, Ano70d, Ano70e, Ano70f, Ano70g, Ano71d, Ano71e, Ano71f, Ano71o, Ano71g, Ano71h, Ano71i, Ano71j, Ano71k, Ano71l, Ano71m, Ano71n, Ano72c, Ano72d, Ano72e, Ano72f, Ano72g, Ano72h, Ano72i, Ano72j, Ano72k, Ano72l, Ano72m, Ano72n, Ano73b, Ano73c, Ano73d, Ano73-331, Ano73-332, Ano73-333, Ano73-334, Ano73-335, Ano73-336, Ano73-337, Ano74-180, Ano74-169, Ano74-170, Ano74-171, Ano74-172, Ano74-173, Ano74-174, Ano74-175, Ano74-176, Ano74-177, Ano74-178, Ano74-179, Ano75-182, Ano75-183, Ano75-184, Ano75-185, Ano75-175, Ano75-176, Ano75-177, Ano75-178, Ano75-179, Ano75-180, Ano75-181, Ano75-186, Ano76-161, Ano76-162, Ano76-163, Ano76-158, Ano76-164, Ano76-159, Ano76-165, Ano76-166, Ano76-160, Ano78-154, Ano78-155, Ano78-148, Ano78-149, Ano78-150, Ano78-151, Ano78-152, Ano78-153, Ano79-146]. tables [Ano79-147, Ano79-148, Ano79-149, Ano79-150, Ano79-151, Ano79-156, Ano79-157, Ano79-152, Ano79-153, Ano79-154, Ano79-155]. tabs [Ano79-52]. tackle [Ano74-232, Ele74d]. take [Mun78, Raf78, Wei78b]. taken [Sch77b]. takeoff [SA78]. takeover [FK76]. takes [Ano78-165, Fro79, Jur79, Kap77c, Kat75, Kay78, Rub76g]. taking [All79b, Chr79j, Tuc74]. tale [Rub76j]. talented [RS77]. Talk [Nil76, Fal77c, Mar77]. talks [Ano79-54]. tampering [Ano79-236]. tangle [Lom79]. tape [Cam77, Fet76, Gra79, Men78e, Roi78, Sal77]. tape-file [Fet76]. tapes [KLM^{+}79]. Tapping [RGCS79]. target [Lev78]. tasks [Ele74d, Mok79c, Moo79, TE74]. Tax [RO76, RO77]. Taxes [RO77]. Teacher [Ano79-54, Fri75d, Sug78e]. teaching [Col78a]. team [RS77, Ste77]. Technical [Ano70-448, Ano71-419, Ano72-454, Ano72-517, Ano78-42, All74a, All74d, All79h, Ano75-153, AB70, Bib71b, Ele78, Jay76, Mas79a, Mas79e, Raf78, Ste75a, Tor76d, Tuc75a, Wul78]. technician [Men74c]. technicians [Jur75d]. technique [Bre75, CC78, Gre74, MVE75, RRS^{+}74, Tor77e]. Techniques [Kap78b, All75i, All76e, Ano75-42, Fal75e, Gos79b, HGB79, JF72, Kap75e, Kap79d, Men75a, Men76b, ME77, Mor78, Par78a, Rai76, Rei74, Tor76a, Tri74, TC73b, Uts77]. Technological [HH70, Tor77g, All76b, Ano72-388, Gre74, HH71, SA78]. technologically [Jur76a]. technologies [Jur74f, Lin76a, Pet77, Tor79, Wal78]. technologist [Wei78b]. Technology [All75h, And73h, Ano75-319, Chr74n, Chr75m, Chr76a, Chr77l, Chr78m, Chr79m, Fal75c, Fal75d, Fri75f, Hol72, Jur75e, Jur77b, Kap77d, Lin74b, Lin75b, Mos75, Rub77f, Rub78b, Rub79c, Wil76a, All78a, Ana79, And79, Ano76-270, Ano77-36, Ano79-160, Bre71, Chr75k, Chr75n, Chr77h, Ele73e, Fal75a, Fre74, Fri77b, Gre79a, Gre79d, Han79, Hor76, Joh70b, Jon77a, Jur77c, Kap77b, Koh71, Mas79a, Men74c, Men76c, Men77a, Men78b, Men79a, ME77, Pay75, Pog78, Rub74e, SOW77, Tor76d, Wad76]. teeth [Men74f]. telecommunication [Fri74g]. Telecommunications [TH77, Ano77-59, Rub79c]. telecommuting [Nil76]. telediagnosis [All76h]. telemedicine [All76h]. Telephone [All75f, Fal75a, Fal77c, Kun75, Shi76, Wel79, Wol76a]. telephone-interface [Wel79]. Telephones [Pay75, Pit79, Fal77a]. telephony [All79b]. Teletext [Car79, Men79a]. Television [Men75b, PG70, Ano74-222, DBB71, GB73, Jur71, Shi70]. tell [Jur74e, RO77, Tor77d]. telling [JF72]. tells [Pet77, Tuc74]. tellurium [KLM^{+}79]. temperatures [Lap74b]. tempered [Ano78-94, Jur74g]. tempers [Lap74b]. Ten [Ano76-206, BCH^{+}79, Hel79, Lus78]. Ten-bit [Ano76-206]. term [Bar79, Buc78b, DBB71, Jur74c, Kap78a]. terminal [Ano74-307, Ano77-179, Fri70b]. terminals [Too78b]. terms [Rub75a]. terrain [KS79]. Tesla [Fri72l]. Test [All77b, ASM71, All75d, All76g, All77d, All78b, All78c, Ano75-234, BFM78, Ele74d, Ele76a, Fal74i, Jur74g, Oli74, RK79, Rub77d, TT74, Tor77d]. tested [Chr79l, Men79a]. testing [All74c, All79e, Ano78-194, Ele74a, Ele76a, Fal75e, FK76, McA71, NR74, Tor77a, Van71]. tests [Hal77, Kap77f, Men79b]. Texas [Buc78b]. text [She79]. texts [Ano77-179, RS77]. theft [Ras75]. their [Ano74-209, Ano75-217, Ano76-270, Bar79, Cam77, Fre74, FR75, Fri76b, Gan77, Han79, Jen78, Kap77e, Men76a, Men77d, Mun78, Rub74c, Rub75b, Rub75f, Rub76c, Rub76f]. them [Ano79-232, Eve77, Fal74g, Fal74j, Kop77, McW77]. themselves [Ano75-153]. Theory [Ano74-47, Ano74-46, Ang74, Ano75-38, Pol71a]. There [Spi76b, Bea75, Hin79, Jur76d, Jur77a, Ano79-232, Rub74f]. there’d [Jur76d]. Thermal [RGCS79, Ano78-36]. Thermoelectric [Ano75-242]. Thermoluminescence [Rai76]. thermonuclear [Tel73]. These [Ano76-273, Jur74e]. They’re [Raf78]. they’ve [All77f]. thief [Sug79a]. things [Jur75d]. thinking [Ano75-240, Chr78f]. third [Gre76b, Pog78]. Thomas [Der79]. Thorough [KLV75]. Those [Ano74-209, Ano75-217, Rub76b, Rub75d]. thousand [Jur75d]. thousands [BC77]. threaten [Rub75e]. Three [All78d, Ano74-234, Kap79d, PHC76, All77a, BC77, Fri70c, Fri74d, Gra74, Jur74h, McK74, Tib78, Mas79e, Rub79b]. three-dimensional [Fri70c, Jur74h]. three-phase [Tib78]. three-pronged [BC77]. three-ring [All77a]. Threshold [HW71, Ano73a, Lap73a, Tor79]. thrived [SR78]. thrusts [Eng75]. thyristor [KG73]. thyristors [Bat71, Kap75c]. thyself [Rac78]. tie [Gra79]. Tilting [Kap76b]. Time [Chr75n, KRS78, All75f, Ano71-420, Ano74-229, Ano76-205, Ano78-48, Ano79-192, CF72, Chr76h, Chr76b, Gib76, Kap79c, Lav78, Nil76, Reu78, Rub75d, Hin79, JF72]. time-of-day [Kap79c]. Time-reference [KRS78]. time-share [Ano78-48]. Time-telling [JF72]. timely [Lap73h]. timers [Men78e]. times [Lus78, Yu75, von79]. timing [DBB71]. tinkerer [Der79, Gor75]. tiny [PMG77]. tips [Ano78-42, Jur74e, RO76]. TM990 [Ano78-201]. TMI [Chr79n, Sug79c]. Today [Kap76b, Chr77d, Gor75, Jon77b, Men76d, Rub75c, Rub76e, Rub76g, Tor77c]. Toeing [All77e]. Together [IY77, Fri76b, Joh79, Lia78, Rei74, RRS^{+}74]. Tokamak [Ste77]. tolerance [Kap75e]. tolerant [Fal76a]. Tom [Fri74a]. tomato [Rub78b]. Tomorrow [Gra74, Kap76b, Max77, Men76d, Wit77]. tongue [Rec75]. Too [Ano77-36, All74d, Chr78e, Ele78, Lap74e, Men76c, Men76b, Pat79a]. tool [Con70, For71, Kor77, Lin74b]. Tools [All75i, Ano77-227, Ano77-228, Fal75e, Fal75f, Kap75e, Mar72, Men74g, Men75d, Tuc75a, Ano77-80, Ano77-81, Ano77-82, Ano77-83, Ano78-109, Ano78-110, Ano78-111, Ano78-112, Ano78-113, Ano78-114, Ano79-108, Ano79-109, Ano79-110, Ano79-111, Ano79-112, Ano79-113, Joh79, Men74b, Mok79c, Rai76]. toothsome [Ele78]. top [Ano78-193, Bea72d, Ele75b]. topic [Chr74b]. topical [RM75]. topics [Rub76i]. Total [Ano79-193, RG79]. totally [All76a, Hel79]. touch [Spi76a, Wei77]. Tougher [Ter77]. tower [Kin75]. town [Wad76]. Toy [Lin74b]. Toys [Gor75, RM75, Rub75f, Tuc75b, Ano77-80, Ano77-81, Ano77-82, Ano77-83, Ano77-227, Ano77-228, Ano78-109, Ano78-110, Ano78-111, Ano78-112, Ano78-113, Ano78-114, Ano79-108, Ano79-109, Ano79-110, Ano79-111, Ano79-112, Ano79-113, Men77d]. trace [Ano79-190]. traced [Rub76g]. tracer [MM74]. tracers [All76i]. track [Col78b, Fri72j]. tracks [Mas79f]. trade [And73c, Men77a, Rea79]. tradeoffs [Cam79, Fai74, Lap73c]. traders [Fri72k]. trades [Fal74i]. traditional [Col78c, Fai74, FR75, Kap76c, Rub74a, Tor77f]. traffic [Cot75, Ele75a, Fal74a, Fri70a, Fri70b, Fri70c, Kin75, Obe77, Qua79, SL78, Tor77h]. traffic-control [SL78]. train [Fri74h, Fri75c, Fri75g, Kal79]. training [Fal74j, Jur75d, Tri78]. trains [Ano74-235]. Transactions [Chr74l]. transbay [Fri75h]. transcending [Gor75]. transfer [Ano77-36, Cou79, Koh71]. Transform [Opp70]. Transistor [Lap73g, Ano74-44, Wei73]. Transistorized [Dal70]. transistors [Ano74-40, Ano75-236]. transit [Hor72, Kap77a, Wad76, Fri74d]. transition [Jor70, Nag74]. translate [Ang74]. Translated [Ano70-449, Ano70-450, Ano70-451, Ano70-452, Ano70-453, Ano70-454]. Transmission [GH70, Ano75-234, AC70, EH76, Eng75, Fag79, Fri76g, Fri77c, Gal77, Hus74, JM77, Kap79c, KG73, KS70, Lus78, Man79, Men75a, Pol70, Rud70]. transmit [Bro79]. transmitted [And79]. Transpo [Fri76i]. transport [Gib70, Gib76, Kap78c]. Transportation [Fri74k, Fri75g, Fri75h, Fri76i, Fri76j, Fri77d, Hea74, Kal75, Kal79, Kap78c, Mas79f, Wad76, Fri74f, Fri74g, Fri74h, Fri74i, Fri74j, Fri75c, Fri75e, Kap77a, Fri76i]. trauma [Ung79]. travail [Fri76j]. travel [Men74d]. Treasure [Ano76-252]. treasurer [Sch77a, Dil75, Rub76j]. treating [All79d]. treaty [Mas79a]. Trends [All79e, Pet77, All79c, Cot75, Fal74j, Gle77, Gra74, Men74c]. TRF [Tuc75b]. trials [Mau75]. tricks [Fri74g]. tried [Rub76d]. triggering [Hil74]. trim [Men74a]. trip [Lin74a]. trolley [Tuc75b]. troubled [Mas79f]. Troubleshooting [Kra75, Men75e]. trump [IFSK77]. truth [Wil70]. try [Men76g]. TTL [Ano75-238, Gar70b, Ray77]. tubes [Mau75, Tuc75b]. tunable [Ano78-192, Oli74]. tuned [Ele74b, IFSK77, Jur75b, Jur75e, Mac76]. tuners [Jur78a]. Tuneup [Jur78a]. tuning [NS77]. turf [Men79a]. turnaround [Ano76-205]. turned [Fri75e]. turning [Ano70-325, Ano71-297, Ano71-298]. turnkey [Men78f]. turns [Gre79d, HH74]. Tuskegee [Man70]. tv [Wil76a, Ano77-59, Jur74a, Men75e, Men76b, Men78f, Men79b, RL74]. TVA [Kap76e, Mic78]. twist [Ele74c]. Two [Ano75-243, Ano78-201, Jur71, Jur74h, Kap77f, Sal72b, All74c, Ano72-519, Col78b, Fal75f, Joh71, Lap74c, Lin74a, McK74]. Two- [Jur74h]. Two-board [Ano78-201]. Two-chip [Ano75-243]. two-edged [Ano72-519]. two-month [Lin74a]. two-track [Col78b]. Two-way [Jur71, Kap77f]. type [SZ74]. typewriter [Ano78-47].

U.K. [Chr77f, Kap77e, Pay75]. U.N.

[And73h]. U.S [Rub75a]. U.S.

[And72d, Ano74-95, Ano75-266, Ano76-229, BCY75, Bar79, Bar78, DBB71, Esp77, Fle71, FR75, Fri75e, Fri76h, Fri76i, Fri77c, Gal79, Gol79, Hol72, Jur75b, Jur75d, Lin77, Mac78, Mas79c, Men74f, Nag74, Par78b, Rab78, Rea79, RO77, Rub75c, Rub75e, Rub76d, Rub77f, Rub77g, SA78, Ste77, Wei78b]. U.S.S.R. [Jur73a, Jur74a]. UHF [Lus78]. UHV [Fri76g, Fri77c]. UK [Ano70-455]. UL [And73b]. ultimate [Jur79, War72]. ultrafast [Ana79]. ultrareliable [Ana79]. Ultrasonic [MVE75, Jur74h]. ultrasonically [MVE75]. Ultrasonics [Jur74h]. unambiguous [KLV75]. unambiguously [Ele74c]. unanimity [Jur77a]. Unchecked [Wil76b]. Uncle [Rub76a]. uncommitted [Sch77b]. unconventional [Ana79]. under-30 [Rub74c]. Undergraduate [Pat79a]. Underground [AC70, Fri72d, Fri72f]. underlie [All75h, Gre79d]. underscored [Sug79c]. Understanding [SF72b, SF72a, KLV75, PG76]. undertaken [Chr79o]. underway [BFM78, Kap77f, Rub74b]. undesirable [PG76]. undisclosed [All78d]. undoing [Chr74g]. unduly [Gra79]. unemployed [Ano72-518, Rub75d]. unfavorable [Cou79]. unformatted [Ano79-195]. uniformity [TG78]. uniformity/gray [TG78]. Union [Wol76a]. unique [All74c, Cou79, Gre74, Joh79, Lin76a]. unit [Ano74-95]. United [AC70, Fel73, HH71]. Units [Ano71-247, Ano78-192, Too78a, Ano71-247]. Universal [Ano79-194, Chr77k, Ano78-191]. Universe [Mac76]. university [BB71]. Unjamming [Tor77h]. Unlimbering [Spi76b]. unlimited [Ano75-237, She79]. Unmanned [NR74]. Untangling [Jur72d]. untracked [Ano76-273]. Unusual [Shi79b, Fri74i, Rob74a]. Unwanted [PG76]. up/down [Ano78-195]. update [Too78a]. uphold [Lin75a]. uplink [Reu78]. upon [Kap76a, KLM^{+}79, Lin76c]. ups [Chr78g]. upstaged [Men78f]. uptime [Alb78]. upturn [Ano72-521]. upward [Bea74a, Fri77c]. urban [For71, Fri76j, Gib76, SL78, Wad76, Wil76b]. USAB [Ano75-267, Ano76-230, Rub76a]. Use [Ano71-247, Ano74-231, Fag78, Fri75b, Gra79, Kap78b, Lin74b, MVE75, Mid74, MM74, Zuc77]. used [All77c, All77f, Ano77-179, Bro79, Col78c, Fal74i, Fey74, RF75]. Useful [Dun78, Car79, Fra78, Jur76f, RM75]. usefulness [Fal74c]. user [Ano73-306, Bea75, Fal74e, New76]. user-oriented [Ano73-306]. Users [MNS77b, All75b, All76c, Ano75-231, Jur74g]. uses [Ano75-237, EW78]. Using [Ano78-43, Bat71, Kau72, Bre75, Fal74f, Opp70, PP70, Rot71]. usually [Bel74, Fal74j]. utilitie [RF75]. utilities [Fri70f, FR75, Kap77f, LC75, Rub75e]. utility [EK71, Fri76f, Lin75a, Pla79a, RK79]. utilizing [KS70]. UV [EBC^{+}79].

V [Gib76, Men75d, All74c, All75h, All76f, Bra79, Fal76c, Men74a, SA78, Sus70a]. V/Information [Men75d]. V/STOL [SA78]. vacuum [Sel71, Tuc75b]. validation [Chr75e]. validity [Rac78]. values [Ano78-196]. valve [KG73]. valves [Jur75a]. vanguard [Hor78]. Varian [Gin75]. variety [All77c, Car79]. vary [Rub74b]. VCRs [Men78f]. vehicle [Bra77]. vehicles [Bra77, Esp77, Ish77, Joh79, Kap77a, RJD73, Sch77b]. verifiable [Mas79a]. versatile [All76g, Doi77, Gre71]. versatility [AC75, Ano74-225, von79]. versions [Wel79]. vertical [SA78]. vertical/short [SA78]. Very [Bre75, Chr74j, Moo79]. very-large-scale [Moo79]. veteran [Kin75]. VHF [CT70, Lus78]. VI [All76g, Jur74g, Jur76f, Lin75b, Sus70b, Tuc75a, Wil76b]. VI/Information [Tuc75a]. via [DBB71]. viability [Kap76c]. viable [All76h, EH76, Lev78]. Video [Cam79, Men78e, All76h, Fag79, Jur74b, Men75b, Men76e, Men77c, Men78f, Men79b]. Video-bandwidth [Cam79]. vie [Jur76d, Men79a]. view [Ano70-325, Ano71-344, Cla77, Cot75, Fri70a, Fri70b, Fri70c]. viewpoint [Ano79-50]. viewpoints [PHC76]. views [Ano76a, Ano76-270, Fri76e, Gol72, Rea79, Rub77a]. vignettes [Chr78b]. VII [All76b, Ele74c, Rub76k, Tro74]. VIII [Bea74b, Kap74b, Kap76a]. Viking [Spi76a, Spi76b]. Vikings [Spi76a, Tor77d]. virtual [Ano74-230]. vis [BKS79]. vis-a-vis [BKS79]. Visible [EBC^{+}79]. vision [Tor78a, Tro74]. visionary [Lin76b]. visit [Lin74a]. Visual [Uts77]. Visual/aural [Uts77]. vital [All77a, Ano74-209, Ano75-217, Eng75]. VLF [KH79]. VLSI [All79c, All79h, Fag78, Gos79a, Gos79b, Hei79, Moo79, Que79, Tor79]. vocabulary [She79]. Voice [BCG73, Mar77, JM77, RRS^{+}74]. voices [FCR^{+}70, Lom79]. Volta [GH71]. Voltage [Ano74-235, Ano75-236, EH76, Jur75e, Kes70, Pol70]. voltage-tuned [Jur75e]. Volume [Ano79-232, Ano75-57, Ano75-231, Ano77-43]. volumes [Bea74b]. voter [Ano76-270]. voyage [MF79]. Voyager [Tor78a]. VPs [Rub76h]. vs [Chr76g, Chr77h, Chr79a, Col78b, Fri74d, Men76e, Rub76k, Sal75, Tor77f]. VTRs [Roi78].

wafers [Har77]. wage [Rub74a]. want [Ano72-523, Rub73a, Rub76a]. Wanted [Bea72b]. wanting [MM74]. wants [Rub76a]. war [And73c]. warehouse [Col79]. warfare [Wal78]. warns [Gol79]. wars [Men78c]. was [Ang74, Der79, Fal76c, Jur76f, Lin74a, Mas79d, Spi76b, Tor76d, Wol76a]. Washington [Ano70-340, Ano70-341, Ano70-342, Ano70-343, Ano70-344, Ano70-345, Ano70-346, Ano70-347, Ano70-348, Ano70-349, Ano70-350, Ano70-351, Ano71-323, Ano71-324, Ano71-325, Ano71-326, Ano71-327, Ano71-328, Ano71-329, Ano71-330, Ano71-331, Ano71-332, Ano71-333, Ano71-334, Ano72-419, Ano72-420, Ano72-421, Ano72-422, Ano72-423, Ano72-424, Ano72-425, Ano72-426, Ano72-427, Ano72-428, Ano72-429, Ano72-430, Ano73-417, Ano73-418, Ano73-419, Ano73-420, Ano73-421, Ano73-422, Ano73-423, Ano73-424, Ano73-425, Ano73-426, Ano73-427, Ano73-428, Ano74-257, Ano74-258, Ano74-259, Ano74-260, Ano74-261, Ano74-262, Ano74-263, Ano74-264, Ano74-265, Ano74-266, Ano74-267, Ano74-268, Ano75-268, Ano75-269, Ano75-270, Ano75-271, Ano75-272, Ano75-273, Ano75-274, Ano75-275, Ano75-276, Ano75-277, Ano75-278, Ano75-279, Ano76-231, Ano76-232, Ano76-233, Ano76-234, Ano76-235, Ano76-236, Ano76-237, Ano76-238, Ano76-239, Ano76-240, Ano76-241, Ano76-242, Ano77-204, Ano77-205, Ano77-206, Ano77-207]. Washington [Ano77-208, Ano77-209, Ano77-210, Ano77-211, Ano77-212, Ano77-213, Ano77-214, Ano77-215, Ano78-226, Ano78-227, Ano78-228, Ano78-229, Ano78-230, Ano78-231, Ano78-232, Ano78-233, Ano78-234, Ano78-235, Ano78-236, Ano78-237, Ano79-220, Ano79-221, Ano79-222, Ano79-223, Ano79-224, Ano79-225, Ano79-226, Ano79-227, Ano79-228, Ano79-229, Ano79-230, Ano79-231]. wasn’t [Wol76a]. Waste [Rub73a, HH74, Mac75, Ras75]. watch [Ele73b, von79]. watchdog [And72b, Men74f]. watchdogs [Man78]. watches [Men76f, von79]. watching [Jur73a]. watt [Bea75, Lap74c]. wave [Cou79, KM71, Obe77]. waveform [Gre72]. waveguide [Men76b]. wavelengths [EBC^{+}79]. waves [Kap79a, Tor76a]. way [And73g, Ano76-273, Bac72, BB71, Car79, Chr79d, DTW76, FFCT74, Fri74i, Hin79, Hol72, Jur71, Kap77f, Mar77, MVE75, Mos75, Rub77c, Spo71]. Ways [JF74, All74c, Fal74g, FR75, Kap77e]. wayside [Fri74g]. weapons [Mas79c]. weather [KRS78, KS79]. wedding [Lin76c]. Week [Ano73-377, All75g, Ano75-317]. weeks [Fal75f, Spi76a]. weighed [Nil76]. weight [JM77]. welding [All79d, Mit77]. well [All79h, Gra74, Kap76b, Kap78b, PMG77, Rei74, Wit77]. went [All76i, Ano76-272, Ano76-273, Dil75, Fal76c, GR76, Jur76f, Kap76e, Men76g, Rub76k, Spi76b, Tor76d, Wol76c]. were [Fri76g, Fri76j, Rub75c, Spi76b, SR78]. WESCON [Ano72a, All74a, All75g, All77g, Ano70a, Ano72-521, Ano72-520, Ano72-522, Ano72-519, Ano78-258, Ano79-160, Lap73h, Tor76b, All74d, All77h]. West [All74d, Fri70f, Fri74h, Fri74j]. Western [Rip79, Wol76a]. Whatever [Wol79]. Wheat [Chr79a]. wheel [McK73]. wheels [Kap79a]. Where [Ano71-421, Chr76l, Chr79o, Jed73, Jur76d, Lin75b, Men77e, RG79, Tuc75b, Zuc77, Dil75, Fey74, Fri74b, Gup74, MRR^{+}77, Rub74d, SOW77]. wherefores [Ewa78]. Wherein [RM75]. Whether [Bel74]. which [Fal75a, Jur74f, KLM^{+}79, Zuc77]. While [Ele74f, Rub77b, Rub78b, Sch77b, Alb78, CSS76, DTW76, Har77, IFSK77, Jur78c, Lap74b, Rub75f, Rub77f, Wel79]. Whirlwind [RS77]. Who [Chr74k, Chr74l, Chr74m, Rub74d, Rub76e, Rub76f, Tho73a, Ele74c, Jur75d, Lap74d, Lin74b, Rub74f, Rub76d, SR78]. whose [Chr76b]. Whys [Ewa78, Fal74g, MT73]. wide [Ano78-197, Ano79-191, Jur75e, Mon76]. wide-band [Ano78-197]. wide-range [Jur75e]. Wild [Buc78b]. Will [Kop77, Mas79f, All75h, Fri74b, Har77, Hea74, Hel79, Her79, JF74, Kap76a, Kap77f, Kap79d, Kay78, KRS78, KLH74, MF79, Man79, Max77, Men74d, PMG77, Rac78, Spe78, Spi76a, Ste77, VHM77, Wil76b, von79]. William [Gin75]. win [IFSK77]. Winchester [Ano79-195]. wind [Hor78, Jay76, Yen78]. wind-generated [Jay76]. Windmills [Jay76]. window [Fri70d]. winds [RK79]. wings [Kap79a, Men79b]. winners [Ano79-47]. Winning [Mac73]. winter [Pla79a]. wire [Fal74f]. wired [Car79, Rub77f]. wisdom [Dre74]. wise [Bea74a]. wisely [Gra79, Men74e]. wish [Lin74b]. within [Hel79, Rub74e]. without [Rub77g]. woman [Gra79]. women [Rub75b]. won [All79g]. woo [IFSK77]. word [Cla71]. words [RRS^{+}74]. Work [Ele76b, All76a, Kop77, Mac73, Mun78, Rei74, Rub75d]. workable [Lou74]. workbench [Men77e]. worker [Kap76e]. workers [Men78g]. Working [Jur72e]. works [Men77e]. World [AB70, Chr78j, All79c, Ano73-377, Ele74b, Fal74a, Fal74j, Fal76c, Fri72k, GB73, Gre76a, Gre76b, Gre79d, Jur75c, Kap78b, Man72, Men74a, Mon76, Mor78, RG79, Gle77]. worry [Jur78c]. worse [Jur77a]. worth [Lin74a]. would [AFCK79, Ano71-421, Bro79, CSS76, Kap77a, Spi76b]. write [Ano78-193]. writing [Ano79-145]. wrong [All76i, Ano76-272, Ano76-273, Fal76c, Fri75e, GR76, Jur76f, Kap76e, Men76g, Rub76k, Spi76b, Tor76d, Wol76c]. WW [Jur75d].

X [Men76a, Spi76b]. XI [Nil76, GR76]. XII [Ano76-273].

YAG [All79d]. Year [Fri76j, Ano73a, Fri75b, Fri75h, Hel79, Men74f, Ste77]. years [Bla77, Chr74a, Ele74e, Ele75a, Her79, Jur76f, Kin75, Lin74a, RF75, Rub77d, Tro74]. yield [Kal79]. yielding [FK76]. yields [Lom79, Wel79]. York [All75f, Ano72-255, Sug78c, Ano79-159, BFM78, WZ78]. you’d [Men77e]. young [Rub74c]. you’re [Chr78i]. yours [Ano73-455].

Zaire [Eng75]. zip [Men76g].

References

Austin:1970:WPG

[AB70] A. L. Austin and J. W. Brewer. World population growth and related technical problems. IEEE Spectrum, 7(12):43–54, December 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anliker:1971:BSA

[AB71] M. Anliker and J. Billingham. Biomechanical studies in aerospace physiology. IEEE Spectrum, 8(3):64–72, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Avila:1970:UTU

[AC70] C. F. Avila and A. F. Corry. Underground transmission in the United States. IEEE Spectrum, 7(3):42–48, March 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Abramson:1975:CSJ

[AC75] N. Abramson and E. R. Cacciamani. Communications: Satellites: Not just a big cable in the sky: High bit rates, high reliability, low costs, and versatility are features of satellite communications systems. IEEE Spectrum, 12(9):36–40, September 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Andryczyk:1979:PES

[AFCK79] R. Andryczyk, P. Foldes, J. Chestek, and B. M. Kaupang. Power/ energy: Solar Power Satellite ground stations: The ground systems for microwave beaming from the SPS would require a rectifying antenna with over 13 billion elements. IEEE Spectrum, 16(7):51–55, July 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Agosto:1976:MPF

[Ago76] W. N. Agosto. Microwave power:

a ‘far-out’ system? IEEE Spectrum, 13(5):48–52, May 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Albrecht:1978:PES

[Alb78] P. F. Albrecht. Power: Evaluating system reliability: When, why, and how often do power systems fail? Such knowledge is invaluable in increasing uptime while holding costs in line. IEEE Spectrum, 15 (8):43–47, August 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1974:CPW

[All74a] R. Allan. Convention preview: WESCON ’74: Something for every EE: The technical program ranges over many disciplines and complements an exciting exhibition. IEEE Spectrum, 11(9):91, September 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1974:IIN

[All74b] R. Allan. Instrumentation IX: New measurement capabilities: The ’74 designer’s arsenal bristles with smart’ scopes, multipurpose meters, and high-speed digital analyzers. IEEE Spectrum, 11(11):83–87, November 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1974:IVT

[All74c] R. Allan. Instrumentation V: Toward serviceability: Built-in calibrating and self-testing features are two ways of easing servicing problems. But digital circuits present unique problems. IEEE Spectrum, 11(11):64–68, November 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1974:SRW

[All74d] R. Allan. Special report: Wescon ’74 program: technical and business, too: Los Angeles is the place, September 10 through 13 the dates, for the Institute’s principal West Coast ‘happening’. IEEE Spectrum, 11 (8):60–61, August 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1975:CSB

[All75a] R. Allan. Circuit/system building blocks: CCDs, I2L, and magnetic bubble devices get smaller and more powerful. IEEE Spectrum, 12(1):49–53, January 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1975:CRS

[All75b]  R. Allan.  Components re-

view: Semiconductor memories: More users are opting for RAMs, ROMs, and PROMs as chip function densities increase and price per bit drops. IEEE Spectrum, 12(8):40–45, August 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1975:ILD

[All75c] R. Allan. Instrumentation LEDs:

Digital domination? Lightemitting diodes proliferate as gas-discharge and liquid-crystal digital readouts find special niches. IEEE Spectrum, 12(2): 40–42, February 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1975:ITE

[All75d]  R. Allan.  Instruments and

test equipment: Simplicity and sophistication are bywords as portables shrink and systems instrumentation grows. IEEE Spectrum, 12(1):90–94, January 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1975:PS

[All75e] R. Allan. Power semiconductors. IEEE Spectrum, 12(11): 37–45, November 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1975:SRG

[All75f] R. Allan. Special report: The Great New York Telephone Fire: The Bell System marshaled its impressive resources to assess and repair damage in record time. IEEE Spectrum, 12 (6):34–40, June 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1975:SRW

[All75g]  R. Allan.  Special report:

WESCON ’75: The week in brief. IEEE Spectrum, 12(9): 52–55, September 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1975:TFV

[All75h] R. Allan. Technology forecasting V: Components: Silicon technology will underlie the next generation of higher performance ICs, and optical devices will play a bigger role. IEEE Spectrum, 12(4):56–59, April 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1975:TID

[All75i] R. Allan. Tools III/design techniques: Designer’s bookshelf: The experts’ choices in five major categories: CAD, design automation, and design techniques for microwave, power, and general systems. IEEE Spectrum, 12 (10):54–57, October 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1976:BEE

[All76a] R. Allan. Biomedical electronics: Electronics aids the disabled: Engineering advances are making possible electronic arms for amputees and mobile work stations for the totally disabled. IEEE Spectrum, 13(11):37–44, November 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1976:CVC

[All76b]  R. Allan.  Cities VII: The

citizen-city interface: Hardware and software are at the heart of easing future shock as the public copes with technological complexities. IEEE Spectrum, 13(7):65–72, July 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1976:CMG

[All76c] R. Allan. Components: Microprocessors galore: Advances in logic, memories, microprocessors, and linear ICs give new capabilities to device users. IEEE Spectrum, 13(1):50–56, January 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1976:EAN

[All76d] R. Allan. Electro ’76: an all-new show. IEEE Spectrum, 13(5):94, May 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1976:ISS

[All76e]  R. Allan.  Instrumentation:

Smaller and smarter: Performance is up and prices are down as advances in ICs and clever packaging techniques abound. IEEE Spectrum, 13 (1):68–72, January 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1976:IVI

[All76f] R. Allan. Instrumentation V: ICs: From sensors to displays: Instrument performance is up, costs are down. It’s done, not with mirrors, but with integrated circuits. IEEE Spectrum, 13 (4):58–62, April 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1976:IVW

[All76g] R. Allan. Instrumentation vi: What’s new in test gear: The shopping list of versatile, compact, and cost-efficient instruments is an engineer’s delight. IEEE Spectrum, 13(4):63–66, April 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1976:MEC

[All76h] R. Allan. Medical electronics:

Coming: The era of telemedicine: Interactive audio/video telediagnosis is proven viable in both government- and privatelyfunded experiments. IEEE Spectrum, 13(12):31–36, December 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1976:WWW

[All76i] R. Allan. What went wrong I: The failure tracers: a review off how reliability experts and analysts probe for the primary cause off device and systems faults. IEEE Spectrum, 13 (10):33–40, October 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1977:ETR

[All77a] R. Allan. Electro ’77: a threering circus: The vital statistics: 300 exhibitors, 42 sessions, and an expected 25 000 attendees. IEEE Spectrum, 14 (4):56–66, April 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1977:IMB

[All77b] R. Allan. Instrumentation: More bang for the buck: Test gear exploits complex integrated circuits to do more, faster. IEEE Spectrum, 14(1):54–57, January 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1977:LAV

[All77c] R. Allan. Logic analyzers: a variety of general and special purpose digital circuit analyzers can be used for system development and/or diagnosis. IEEE Spectrum, 14(8):34–39, August 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1977:SFT

[All77d] R. Allan. The ‘safety factor’ in test equipment. IEEE Spectrum, 14(4):43–47, April 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1977:STM

[All77e] R. Allan. Semiconductors: Toeing the (microfine) line: Functional density is what counts in advanced solid-state devices. Here’s a report on recent progress. IEEE Spectrum, 14 (12):34–42, December 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1977:SSD

[All77f] R. Allan. Solid-state devices: Amorphous semiconductors revisited: Despite residual skepticism, they’ve been successfully used in memory, microfiche, and graphic-arts applications. IEEE Spectrum, 14(5):41–45, May 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1977:SRC

[All77g]  R. Allan.  Special report:

Convention new products at WESCON ’77. IEEE Spectrum, 14(9):78–80P, September 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1977:WGG

[All77h] R. Allan. Wescon ’77: Golden Gate beckons. IEEE Spectrum, 14(8):64, August 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1978:EBP

[All78a] R. Allan. Electronics to boost productivity: Computer technology spurs automation, energy cuts. IEEE Spectrum, 15 (1):45–48, January 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1978:IEF

[All78b] R. Allan. Instrumentation: Emphasis on field: Microprocessorbased systems spawn built-in test gear. IEEE Spectrum, 15 (1):41–44, January 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1978:INP

[All78c] R. Allan. Instrumentation: New philosophies for portable digital instruments: Signature analysis and circuit emulation are exploited in new test equipment to reduce runaway service costs. IEEE Spectrum, 15(6):32–37, June 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1978:MTA

[All78d]  R. Allan.  Microprocessors:

Three amazing micromice: Hitherto undisclosed details: a closer look at some off the “smart” electronic micromice that have participated in the Spectrum/ Computer Amazing Micro-Mouse Maze Contest. IEEE Spectrum, 15(11):62–65, November 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allen:1978:PIP

[All78e] K. R. Allen. Patents: Invention pacts. IEEE Spectrum, 15 (3):54–59, March 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1979:BRS

[All79a] R. Allan. Busy robots spur productivity. IEEE Spectrum, 16 (9):31–36, September 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1979:DTS

[All79b] R. Allan. Digital telephony slowly taking over. IEEE Spectrum, 16(10):56, October 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1979:EEP

[All79c] R. Allan. Electronics: Experts probe world outlook: Representatives of the world’s nations review trends as the VLSI era approaches. IEEE Spectrum, 16(7):56–59, July 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1979:ILF

[All79d] R. Allan. Industry: Lasers in the factory: Nd: YAG and CO_{2} lasers are growing in popularity for cutting, drilling, welding, scribing, and heat-treating a host of metals and nonmetals. IEEE Spectrum, 16 (5):42–51, May 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1979:IIA

[All79e] R. Allan. Instrumentation: Ingenuity abounds: Trends: Infield diagnostics, in-plant and in-process testing, and smarter portables. IEEE Spectrum, 16 (1):48–52, January 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1979:MIP

[All79f]  R. Allan.  The microcom-

puter invades the production line. IEEE Spectrum, 16(1): 53–57, January 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1979:MAM

[All79g] R. Allan. Microprocessors: The amazing micromice: See how they won: Probing the innards of the smartest and fastest entries in the amazing micro-mouse maze contest. IEEE Spectrum, 16 (9):62–65, September 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Allan:1979:SSV

[All79h] R. Allan. Solid state: VLSI: Scoping its future: The experts probe the alternatives among markets, business structures, and resources, as well as technical challenges. IEEE Spectrum, 16 (4):30–37, April 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anacker:1979:CCD

[Ana79] W. Anacker. Computers: Computing at 4 degrees Kelvin: Superconducting technology and unconventional packaging herald an era of ultrafast, ultrareliable computers. IEEE Spectrum, 16(5):26–37, May 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:1970:ACB

[And70]  V. C. Anderson.  Acoustic

communication is better than none. IEEE Spectrum, 7(10): 63–68, October 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Andover:1972:CO

[And72a] J. J. Andover. Congress and the OTA. IEEE Spectrum, 9 (9):32–33, September 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Andover:1972:EW

[And72b] J. J. Andover. Environmental watchdog. IEEE Spectrum, 9 (12):66–67, December 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Andover:1972:EBG

[And72c] J. J. Andover. For EIA: bad and good news. IEEE Spectrum, 9 (11):58–59, November 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Andover:1972:ORU

[And72d] J. J. Andover. Outlook for R&D in the U.S. IEEE Spectrum, 9(7):40–41, July 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Andover:1972:RN

[And72e] J. J. Andover. Revamping NSF. IEEE Spectrum, 9(10): 78–79, October 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Andover:1972:SSS

[And72f] J. J. Andover. Space Shuttle:

Spirit of ’76. IEEE Spectrum, 9(8):61–62, August 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Andover:1973:APC

[And73a] J. J. Andover. About that power crisis. IEEE Spectrum, 10(3):72–73, March 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Andover:1973:BUL

[And73b] J. J. Andover. Behind the UL label. IEEE Spectrum, 10 (1):65–66, January 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Andover:1973:CTW

[And73c] J. J. Andover. Cooling the trade war. IEEE Spectrum, 10 (2):41–42, February 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Andover:1973:EE

[And73d] J. J. Andover. EEs and the environment. IEEE Spectrum, 10(4):45–46, April 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Andover:1973:FFP

[And73e] J. J. Andover. Futurism: for fun and profit. IEEE Spectrum, 10(7):37–38, July 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Andover:1973:MRS

[And73f] J. J. Andover. Mobile radio: sparks fly. IEEE Spectrum, 10 (9):69, September 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Andover:1973:WM

[And73g] J. J. Andover. On the way to metric. IEEE Spectrum, 10 (8):39–40, August 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Andover:1973:TU

[And73h] J. J. Andover. Technology and the U.N. IEEE Spectrum, 10 (10):54–55, October 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anderson:1978:IOS

[And78] D. B. Anderson. Integrated optical spectrum analyzer: an imminent ‘chip’. IEEE Spectrum, 15 (12):22–29, December 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Andrews:1979:ATD

[And79] H. C. Andrews. Advanced technology: Digital image processing: Images that provide both high resolution and large dynamic range are transmitted, stored, and processed by digital means. IEEE Spectrum, 16(4):38–49, April 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Angelini:1974:MAS

[Ang74]  A. M. Angelini.  Marconi’s

achievement: a systems outlook: Yet another jewel in Marconi’s crown was his ability to formulate an objective, to translate theory to practice, to develop a system. IEEE Spectrum, 11 (12):47–52, December 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:WEL

[Ano70a] Anonymous. 1970 WESCON exhibitor list. IEEE Spectrum, 7 (7):114–119, July 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:ATCa

[Ano70b] Anonymous. Advance tables of contents. IEEE Spectrum, 7 (1):94–97, January 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:ATCb

[Ano70c] Anonymous. Advance tables of contents. IEEE Spectrum, 7(2): 103–108, February 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:ATCc

[Ano70d] Anonymous. Advance tables of contents. IEEE Spectrum, 7 (3):100–108, March 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:ATCd

[Ano70e] Anonymous. Advance tables of contents. IEEE Spectrum, 7(4):93–97, April 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:ATCe

[Ano70f] Anonymous. Advance tables of contents. IEEE Spectrum, 7(5):90–93, May 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:ATCf

[Ano70g] Anonymous. Advance tables of contents. IEEE Spectrum, 7 (6):96–100, June 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aa

[Ano70h] Anonymous. [advertisement].

IEEE Spectrum, 7(1):1, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ab

[Ano70i] Anonymous. [advertisement].

IEEE Spectrum, 7(1):1, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ac

[Ano70j] Anonymous. [advertisement].

IEEE Spectrum, 7(1):1, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ad

[Ano70k] Anonymous. [advertisement].

IEEE Spectrum, 7(1):2, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ae

[Ano70l] Anonymous. [advertisement].

IEEE Spectrum, 7(1):7–14, January 1970. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Af

[Ano70m] Anonymous. [advertisement]. IEEE Spectrum, 7(1):16a–16b, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ag

[Ano70n] Anonymous. [advertisement]. IEEE Spectrum, 7(1):16d–16d, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ah

[Ano70o] Anonymous. [advertisement]. IEEE Spectrum, 7(1):16e–16f, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ai

[Ano70p] Anonymous. [advertisement]. IEEE Spectrum, 7(1):16g–16h, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aj

[Ano70q] Anonymous. [advertisement]. IEEE Spectrum, 7(1):16k–16k, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ak

[Ano70r] Anonymous. [advertisement]. IEEE Spectrum, 7(1):16m– 16p, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Al

[Ano70s] Anonymous. [advertisement].

IEEE Spectrum, 7(1):19, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Am

[Ano70t] Anonymous. [advertisement].

IEEE Spectrum, 7(1):21, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:An

[Ano70u] Anonymous. [advertisement].

IEEE Spectrum, 7(1):29–30, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ao

[Ano70v] Anonymous. [advertisement].

IEEE Spectrum, 7(1):63–66, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ap

[Ano70w] Anonymous. [advertisement].

IEEE Spectrum, 7(1):102–103, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aq

[Ano70x] Anonymous. [advertisement].

IEEE Spectrum, 7(1):107–111, January 1970. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ar

[Ano70y] Anonymous. [advertisement]. IEEE Spectrum, 7(1):115, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:As

[Ano70z] Anonymous. [advertisement]. IEEE Spectrum, 7(1):123, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:At

[Ano70-27] Anonymous. [advertisement]. IEEE Spectrum, 7(2):nil1, February 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Au

[Ano70-28] Anonymous. [advertisement]. IEEE Spectrum, 7(2):1a–1a, February 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Av

[Ano70-29] Anonymous. [advertisement]. IEEE Spectrum, 7(2):1, February 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aw

[Ano70-30] Anonymous. [advertisement]. IEEE Spectrum, 7(2):8–10, February 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ax

[Ano70-31] Anonymous. [advertisement].

IEEE Spectrum, 7(2):13, February 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ay

[Ano70-32] Anonymous. [advertisement].

IEEE Spectrum, 7(2):27–38, February 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Az

[Ano70-33] Anonymous. [advertisement].

IEEE Spectrum, 7(2):67–70, February 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aba

[Ano70-34] Anonymous. [advertisement].

IEEE Spectrum, 7(2):107–109, February 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Abb

[Ano70-35] Anonymous. [advertisement].

IEEE Spectrum, 7(2):111, February 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Abc

[Ano70-36] Anonymous. [advertisement].

IEEE Spectrum, 7(2):113, February 1970. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Abd

[Ano70-37] Anonymous. [advertisement]. IEEE Spectrum, 7(2):119–123, February 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Abe

[Ano70-38] Anonymous. [advertisement]. IEEE Spectrum, 7(2):128–133, February 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Abf

[Ano70-39] Anonymous. [advertisement]. IEEE Spectrum, 7(2):nil2, February 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Abg

[Ano70-40] Anonymous. [advertisement]. IEEE Spectrum, 7(3):nil1, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Abh

[Ano70-41] Anonymous. [advertisement]. IEEE Spectrum, 7(3):1a–1b, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Abi

[Ano70-42] Anonymous. [advertisement]. IEEE Spectrum, 7(3):1, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Abj

[Ano70-43] Anonymous. [advertisement].

IEEE Spectrum, 7(3):7–14, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Abk

[Ano70-44] Anonymous. [advertisement].

IEEE Spectrum, 7(3):17–19, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Abl

[Ano70-45] Anonymous. [advertisement].

IEEE Spectrum, 7(3):21–25, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Abm

[Ano70-46] Anonymous. [advertisement].

IEEE Spectrum, 7(3):29–30, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Abn

[Ano70-47] Anonymous. [advertisement].

IEEE Spectrum, 7(3):65–68, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Abo

[Ano70-48] Anonymous. [advertisement].

IEEE Spectrum, 7(3):97, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Abp

[Ano70-49] Anonymous. [advertisement]. IEEE Spectrum, 7(3):101–109, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Abq

[Ano70-50] Anonymous. [advertisement]. IEEE Spectrum, 7(3):110–111, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Abr

[Ano70-51] Anonymous. [advertisement]. IEEE Spectrum, 7(3):112–113, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Abs

[Ano70-52] Anonymous. [advertisement]. IEEE Spectrum, 7(3):115–119, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Abt

[Ano70-53] Anonymous. [advertisement]. IEEE Spectrum, 7(3):123–129, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Abu

[Ano70-54] Anonymous. [advertisement]. IEEE Spectrum, 7(3):217–218, March 1970. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Abv

[Ano70-55] Anonymous. [advertisement].

IEEE Spectrum, 7(3):nil2, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Abw

[Ano70-56] Anonymous. [advertisement].

IEEE Spectrum, 7(4):nil1, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Abx

[Ano70-57] Anonymous. [advertisement].

IEEE Spectrum, 7(4):1a–1b, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aby

[Ano70-58] Anonymous. [advertisement].

IEEE Spectrum, 7(4):1, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Abz

[Ano70-59] Anonymous. [advertisement].

IEEE Spectrum, 7(4):9–10, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aca

[Ano70-60] Anonymous. [advertisement].

IEEE Spectrum, 7(4):13–15, April 1970. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Acb

[Ano70-61] Anonymous. [advertisement]. IEEE Spectrum, 7(4):22, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Acc

[Ano70-62] Anonymous. [advertisement]. IEEE Spectrum, 7(4):63–66, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Acd

[Ano70-63] Anonymous. [advertisement]. IEEE Spectrum, 7(4):99, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ace

[Ano70-64] Anonymous. [advertisement]. IEEE Spectrum, 7(4):100–101, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Acf

[Ano70-65] Anonymous. [advertisement]. IEEE Spectrum, 7(4):109–120, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Acg

[Ano70-66] Anonymous. [advertisement]. IEEE Spectrum, 7(4):122, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ach

[Ano70-67] Anonymous. [advertisement].

IEEE Spectrum, 7(4):129–130, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aci

[Ano70-68] Anonymous. [advertisement].

IEEE Spectrum, 7(4):nil2, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Acj

[Ano70-69] Anonymous. [advertisement].

IEEE Spectrum, 7(5):nil1, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ack

[Ano70-70] Anonymous. [advertisement].

IEEE Spectrum, 7(5):1a–1b, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Acl

[Ano70-71] Anonymous. [advertisement].

IEEE Spectrum, 7(5):1, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Acm

[Ano70-72] Anonymous. [advertisement].

IEEE Spectrum, 7(5):5, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Acn

[Ano70-73] Anonymous. [advertisement]. IEEE Spectrum, 7(5):20, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aco

[Ano70-74] Anonymous. [advertisement]. IEEE Spectrum, 7(5):59–62, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Acp

[Ano70-75] Anonymous. [advertisement]. IEEE Spectrum, 7(5):87, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Acq

[Ano70-76] Anonymous. [advertisement]. IEEE Spectrum, 7(5):95, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Acr

[Ano70-77] Anonymous. [advertisement]. IEEE Spectrum, 7(5):96–97, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Acs

[Ano70-78] Anonymous. [advertisement]. IEEE Spectrum, 7(5):101–103, May 1970. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Act

[Ano70-79] Anonymous. [advertisement].

IEEE Spectrum, 7(5):107–112, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Acu

[Ano70-80] Anonymous. [advertisement].

IEEE Spectrum, 7(5):115, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Acv

[Ano70-81] Anonymous. [advertisement].

IEEE Spectrum, 7(5):121–122, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Acw

[Ano70-82] Anonymous. [advertisement].

IEEE Spectrum, 7(5):nil2, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Acx

[Ano70-83] Anonymous. [advertisement].

IEEE Spectrum, 7(6):nil1, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Acy

[Ano70-84] Anonymous. [advertisement].

IEEE Spectrum, 7(6):1, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Acz

[Ano70-85] Anonymous. [advertisement]. IEEE Spectrum, 7(6):5, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ada

[Ano70-86] Anonymous. [advertisement]. IEEE Spectrum, 7(6):9–10, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Adb

[Ano70-87] Anonymous. [advertisement]. IEEE Spectrum, 7(6):13, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Adc

[Ano70-88] Anonymous. [advertisement]. IEEE Spectrum, 7(6):15–19, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Add

[Ano70-89] Anonymous. [advertisement]. IEEE Spectrum, 7(6):24, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ade

[Ano70-90] Anonymous. [advertisement]. IEEE Spectrum, 7(6):59–62, June 1970. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Adf

[Ano70-91] Anonymous. [advertisement].

IEEE Spectrum, 7(6):97–99, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Adg

[Ano70-92] Anonymous. [advertisement].

IEEE Spectrum, 7(6):119–120, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Adh

[Ano70-93] Anonymous. [advertisement].

IEEE Spectrum, 7(6):121–122, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Adi

[Ano70-94] Anonymous. [advertisement].

IEEE Spectrum, 7(6):nil4, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Adj

[Ano70-95] Anonymous. [advertisement].

IEEE Spectrum, 7(7):nil1, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Adk

[Ano70-96] Anonymous. [advertisement].

IEEE Spectrum, 7(7):1, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Adl

[Ano70-97] Anonymous. [advertisement]. IEEE Spectrum, 7(7):5, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Adm

[Ano70-98] Anonymous. [advertisement]. IEEE Spectrum, 7(7):7–10, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Adn

[Ano70-99] Anonymous. [advertisement]. IEEE Spectrum, 7(7):13, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ado

[Ano70-100] Anonymous. [advertisement]. IEEE Spectrum, 7(7):19, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Adp

[Ano70-101] Anonymous. [advertisement]. IEEE Spectrum, 7(7):22, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Adq

[Ano70-102] Anonymous. [advertisement]. IEEE Spectrum, 7(7):59–62, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Adr

[Ano70-103] Anonymous. [advertisement].

IEEE Spectrum, 7(7):93, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ads

[Ano70-104] Anonymous. [advertisement].

IEEE Spectrum, 7(7):94–95, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Adt

[Ano70-105] Anonymous. [advertisement].

IEEE Spectrum, 7(7):97, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Adu

[Ano70-106] Anonymous. [advertisement].

IEEE Spectrum, 7(7):121–122, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Adv

[Ano70-107] Anonymous. [advertisement].

IEEE Spectrum, 7(7):nil4, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Adw

[Ano70-108] Anonymous. [advertisement].

IEEE Spectrum, 7(8):nil1, August 1970. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Adx

[Ano70-109] Anonymous. [advertisement]. IEEE Spectrum, 7(8):1, August 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ady

[Ano70-110] Anonymous. [advertisement]. IEEE Spectrum, 7(8):5–7, August 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Adz

[Ano70-111] Anonymous. [advertisement]. IEEE Spectrum, 7(8):9–14, August 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aea

[Ano70-112] Anonymous. [advertisement]. IEEE Spectrum, 7(8):17, August 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aeb

[Ano70-113] Anonymous. [advertisement]. IEEE Spectrum, 7(8):19–22, August 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aec

[Ano70-114] Anonymous. [advertisement]. IEEE Spectrum, 7(8):25–28, August 1970. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aed

[Ano70-115] Anonymous. [advertisement].

IEEE Spectrum, 7(8):63–66, August 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aee

[Ano70-116] Anonymous. [advertisement].

IEEE Spectrum, 7(8):108–112, August 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aef

[Ano70-117] Anonymous. [advertisement].

IEEE Spectrum, 7(8):116–119, August 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aeg

[Ano70-118] Anonymous. [advertisement].

IEEE Spectrum, 7(8):123, August 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aeh

[Ano70-119] Anonymous. [advertisement].

IEEE Spectrum, 7(8):129–130, August 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aei

[Ano70-120] Anonymous. [advertisement].

IEEE Spectrum, 7(8):nil4, August 1970. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aej

[Ano70-121] Anonymous. [advertisement]. IEEE Spectrum, 7(9):nil1, September 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aek

[Ano70-122] Anonymous. [advertisement]. IEEE Spectrum, 7(9):1, September 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ael

[Ano70-123] Anonymous. [advertisement]. IEEE Spectrum, 7(9):5, September 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aem

[Ano70-124] Anonymous. [advertisement]. IEEE Spectrum, 7(9):7–9, September 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aen

[Ano70-125] Anonymous. [advertisement]. IEEE Spectrum, 7(9):13, September 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aeo

[Ano70-126] Anonymous. [advertisement]. IEEE Spectrum, 7(9):20, September 1970. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aep

[Ano70-127] Anonymous. [advertisement].

IEEE Spectrum, 7(9):59–62, September 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aeq

[Ano70-128] Anonymous. [advertisement].

IEEE Spectrum, 7(9):94–95, September 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aer

[Ano70-129] Anonymous. [advertisement].

IEEE Spectrum, 7(9):101, September 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aes

[Ano70-130] Anonymous. [advertisement].

IEEE Spectrum, 7(9):121–122, September 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aet

[Ano70-131] Anonymous. [advertisement].

IEEE Spectrum, 7(9):nil4, September 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aeu

[Ano70-132] Anonymous. [advertisement].

IEEE Spectrum, 7(10):nil1, October 1970. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aev

[Ano70-133] Anonymous. [advertisement]. IEEE Spectrum, 7(10):1, October 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aew

[Ano70-134] Anonymous. [advertisement]. IEEE Spectrum, 7(10):5, October 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aex

[Ano70-135] Anonymous. [advertisement]. IEEE Spectrum, 7(10):7–10, October 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aey

[Ano70-136] Anonymous. [advertisement]. IEEE Spectrum, 7(10):13, October 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aez

[Ano70-137] Anonymous. [advertisement]. IEEE Spectrum, 7(10):15, October 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Afa

[Ano70-138] Anonymous. [advertisement]. IEEE Spectrum, 7(10):20, October 1970. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Afb

[Ano70-139] Anonymous. [advertisement].

IEEE Spectrum, 7(10):59–62, October 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Afc

[Ano70-140] Anonymous. [advertisement].

IEEE Spectrum, 7(10):96–97, October 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Afd

[Ano70-141] Anonymous. [advertisement].

IEEE Spectrum, 7(10):115, October 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Afe

[Ano70-142] Anonymous. [advertisement].

IEEE Spectrum, 7(10):121– 122, October 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aff

[Ano70-143] Anonymous. [advertisement].

IEEE Spectrum, 7(10):nil4, October 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Afg

[Ano70-144] Anonymous. [advertisement].

IEEE Spectrum, 7(11):nil1, November 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Afh

[Ano70-145] Anonymous. [advertisement]. IEEE Spectrum, 7(11):1, November 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Afi

[Ano70-146] Anonymous. [advertisement]. IEEE Spectrum, 7(11):5–7, November 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Afj

[Ano70-147] Anonymous. [advertisement]. IEEE Spectrum, 7(11):9–14, November 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Afk

[Ano70-148] Anonymous. [advertisement]. IEEE Spectrum, 7(11):17, November 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Afl

[Ano70-149] Anonymous. [advertisement]. IEEE Spectrum, 7(11):19–23, November 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Afm

[Ano70-150] Anonymous. [advertisement]. IEEE Spectrum, 7(11):26, November 1970. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Afn

[Ano70-151] Anonymous. [advertisement].

IEEE Spectrum, 7(11):28–30, November 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Afo

[Ano70-152] Anonymous. [advertisement].

IEEE Spectrum, 7(11):59–62, November 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Afp

[Ano70-153] Anonymous. [advertisement].

IEEE Spectrum, 7(11):110, November 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Afq

[Ano70-154] Anonymous. [advertisement].

IEEE Spectrum, 7(11):112–116, November 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Afr

[Ano70-155] Anonymous. [advertisement].

IEEE Spectrum, 7(11):119, November 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Afs

[Ano70-156] Anonymous. [advertisement].

IEEE Spectrum, 7(11):121–122, November 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aft

[Ano70-157] Anonymous. [advertisement]. IEEE Spectrum, 7(11):nil4, November 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Afu

[Ano70-158] Anonymous. [advertisement]. IEEE Spectrum, 7(12):nil1, December 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Afv

[Ano70-159] Anonymous. [advertisement]. IEEE Spectrum, 7(12):1, December 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Afw

[Ano70-160] Anonymous. [advertisement]. IEEE Spectrum, 7(12):5, December 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Afx

[Ano70-161] Anonymous. [advertisement]. IEEE Spectrum, 7(12):7–12, December 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Afy

[Ano70-162] Anonymous. [advertisement]. IEEE Spectrum, 7(12):15, December 1970. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Afz

[Ano70-163] Anonymous. [advertisement].

IEEE Spectrum, 7(12):17, December 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Aga

[Ano70-164] Anonymous. [advertisement].

IEEE Spectrum, 7(12):23, December 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Agb

[Ano70-165] Anonymous. [advertisement].

IEEE Spectrum, 7(12):55–58, December 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Agc

[Ano70-166] Anonymous. [advertisement].

IEEE Spectrum, 7(12):113–114, December 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Agd

[Ano70-167] Anonymous. [advertisement].

IEEE Spectrum, 7(12):nil4, December 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:AIa

[Ano70-168] Anonymous. Annual index.

IEEE Spectrum, 7(12):96–110, December 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:AGI

[Ano70-169] Anonymous. An author’s guide to IEEE publications. IEEE Spectrum, 7(1):17–18, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BRa

[Ano70-170] Anonymous. Book reviews. IEEE Spectrum, 7(1):104–111, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BRb

[Ano70-171] Anonymous. Book reviews. IEEE Spectrum, 7(2):114–116, February 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BRc

[Ano70-172] Anonymous. Book reviews. IEEE Spectrum, 7(3):114–119, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BRd

[Ano70-173] Anonymous. Book reviews. IEEE Spectrum, 7(4):102–107, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BRe

[Ano70-174] Anonymous. Book reviews. IEEE Spectrum, 7(5):98–102, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BRf

[Ano70-175] Anonymous. Book reviews.

IEEE Spectrum, 7(6):104–109, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BRg

[Ano70-176] Anonymous. Book reviews.

IEEE Spectrum, 7(7):96–101, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BRh

[Ano70-177] Anonymous. Book reviews.

IEEE Spectrum, 7(8):105–107, August 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BRi

[Ano70-178] Anonymous. Book reviews.

IEEE Spectrum, 7(9):96–103, September 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BRj

[Ano70-179] Anonymous. Book reviews.

IEEE Spectrum, 7(10):98–105, October 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BRk

[Ano70-180] Anonymous. Book reviews.

IEEE Spectrum, 7(11):90–93, November 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BRl

[Ano70-181] Anonymous. Book reviews. IEEE Spectrum, 7(12):86–88, December 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:ISRa

[Ano70-182] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 7(1):129–130, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:ISRb

[Ano70-183] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 7(2):1a–1b, February 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:ISRc

[Ano70-184] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 7(2):137–138, February 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:ISRd

[Ano70-185] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 7(3):1a–1b, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:ISRe

[Ano70-186] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 7(3):217–218, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:ISRf

[Ano70-187] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 7(4):1a–1b, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:ISRg

[Ano70-188] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 7(4):129–130, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:ISRh

[Ano70-189] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 7(5):1a–1b, May 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:ISRi

[Ano70-190] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 7(5):121–122, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:ISRj

[Ano70-191] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 7(6):121–122, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:ISRk

[Ano70-192] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 7(7):121–122, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:ISRl

[Ano70-193] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 7(8):129–130, August 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:ISRm

[Ano70-194] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 7(9):121–122, September 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:ISRn

[Ano70-195] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 7(10):121–122, October 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:ISRo

[Ano70-196] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 7(11):121–122, November 1970. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:ISRp

[Ano70-197] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 7(12):113–114, December 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SRS

[Ano70-198] Anonymous. Spectrum reader service card. IEEE Spectrum, 7 (1):1a–1b, January 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BPa

[Ano70-199] Anonymous. [breaker page].

IEEE Spectrum, 7(1):4, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BPb

[Ano70-200] Anonymous. [breaker page].

IEEE Spectrum, 7(2):4, February 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BPc

[Ano70-201] Anonymous. [breaker page].

IEEE Spectrum, 7(3):4, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BPd

[Ano70-202] Anonymous. [breaker page].

IEEE Spectrum, 7(4):4, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BPe

[Ano70-203] Anonymous. [breaker page]. IEEE Spectrum, 7(5):4, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BPf

[Ano70-204] Anonymous. [breaker page]. IEEE Spectrum, 7(6):nil2, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BPg

[Ano70-205] Anonymous. [breaker page]. IEEE Spectrum, 7(6):4, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BPh

[Ano70-206] Anonymous. [breaker page]. IEEE Spectrum, 7(7):nil2, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BPi

[Ano70-207] Anonymous. [breaker page]. IEEE Spectrum, 7(7):4, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BPj

[Ano70-208] Anonymous. [breaker page]. IEEE Spectrum, 7(8):nil2, August 1970. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BPk

[Ano70-209] Anonymous. [breaker page].

IEEE Spectrum, 7(8):4, August 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BPl

[Ano70-210] Anonymous. [breaker page].

IEEE Spectrum, 7(9):nil2, September 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BPm

[Ano70-211] Anonymous. [breaker page].

IEEE Spectrum, 7(9):4, September 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BPn

[Ano70-212] Anonymous. [breaker page].

IEEE Spectrum, 7(10):nil2, October 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BPo

[Ano70-213] Anonymous. [breaker page].

IEEE Spectrum, 7(10):4, October 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BPp

[Ano70-214] Anonymous. [breaker page].

IEEE Spectrum, 7(11):nil2, November 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BPq

[Ano70-215] Anonymous. [breaker page]. IEEE Spectrum, 7(11):4, November 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BPr

[Ano70-216] Anonymous. [breaker page]. IEEE Spectrum, 7(12):nil2, December 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BPs

[Ano70-217] Anonymous. [breaker page]. IEEE Spectrum, 7(12):4, December 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:BEU

[Ano70-218] Anonymous. British electronics in the US market. IEEE Spectrum, 7(1):16e–16f, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Cb

[Ano70-219] Anonymous. Calendar. IEEE Spectrum, 7(1):26–28, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Cd

[Ano70-220] Anonymous. Calendar. IEEE Spectrum, 7(2):18–20, February 1970. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Cf

[Ano70-221] Anonymous. Calendar. IEEE Spectrum, 7(3):26–28, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Cg

[Ano70-222] Anonymous. Calendar. IEEE Spectrum, 7(4):19–21, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ch

[Ano70-223] Anonymous. Calendar. IEEE Spectrum, 7(5):17–19, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Cj

[Ano70-224] Anonymous. Calendar. IEEE Spectrum, 7(6):22–23, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Cl

[Ano70-225] Anonymous. Calendar. IEEE Spectrum, 7(7):20–21, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Cn

[Ano70-226] Anonymous. Calendar. IEEE Spectrum, 7(8):23–24, August 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Co

[Ano70-227] Anonymous. Calendar. IEEE Spectrum, 7(9):18–19, September 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Cp

[Ano70-228] Anonymous. Calendar. IEEE Spectrum, 7(10):18–19, October 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Cq

[Ano70-229] Anonymous. Calendar. IEEE Spectrum, 7(11):24–25, November 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Cr

[Ano70-230] Anonymous. Calendar. IEEE Spectrum, 7(12):21–22, December 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:CE

[Ano70-231] Anonymous. Capital electronics. IEEE Spectrum, 7(1): 16g–16h, January 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:CAa

[Ano70-232] Anonymous. Classified advertising. IEEE Spectrum, 7(1):

98–99, January 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:CAb

[Ano70-233] Anonymous. Classified advertising. IEEE Spectrum, 7(2): 115–117, February 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:CAc

[Ano70-234] Anonymous. Classified advertising. IEEE Spectrum, 7(4): 103–107, April 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:CAd

[Ano70-235] Anonymous. Classified advertising. IEEE Spectrum, 7(6):102, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:CAe

[Ano70-236] Anonymous. Classified advertising. IEEE Spectrum, 7 (7):105–111, July 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:CAf

[Ano70-237] Anonymous. Classified advertising. IEEE Spectrum, 7(8):122, August 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:CAg

[Ano70-238] Anonymous. Classified advertising. IEEE Spectrum, 7(9):

107–113, September 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:CAh

[Ano70-239] Anonymous. Classified advertising. IEEE Spectrum, 7(10): 99–105, October 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:CAi

[Ano70-240] Anonymous. Classified advertising. IEEE Spectrum, 7(11): 95–106, November 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:CAj

[Ano70-241] Anonymous. Classified advertising. IEEE Spectrum, 7 (12):93, December 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ca

[Ano70-242] Anonymous. Components. IEEE Spectrum, 7(1):16j–16l, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Cc

[Ano70-243] Anonymous. Correction. IEEE Spectrum, 7(2):9, February 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Cm

[Ano70-244] Anonymous. Correction. IEEE Spectrum, 7(8):10, August 1970.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ci

[Ano70-245] Anonymous. Corrections. IEEE Spectrum, 7(6):9, June 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ck

[Ano70-246] Anonymous. Corrections. IEEE Spectrum, 7(7):9, July 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FCJa

[Ano70-247] Anonymous. Fall conferences in Japan. IEEE Spectrum, 7(8):108, August 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FPa

[Ano70-248] Anonymous. Focal points. IEEE Spectrum, 7(1):20–25, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FPb

[Ano70-249] Anonymous. Focal points. IEEE Spectrum, 7(2):14–17, February 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FPc

[Ano70-250] Anonymous. Focal points. IEEE Spectrum, 7(3):20–25, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FPd

[Ano70-251] Anonymous. Focal points. IEEE Spectrum, 7(4):16–18, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FPe

[Ano70-252] Anonymous. Focal points. IEEE Spectrum, 7(5):11–16, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FPf

[Ano70-253] Anonymous. Focal points. IEEE Spectrum, 7(6):14–21, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FPg

[Ano70-254] Anonymous. Focal points. IEEE Spectrum, 7(7):14–19, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FPh

[Ano70-255] Anonymous. Focal points. IEEE Spectrum, 7(8):18–21, August 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FPi

[Ano70-256] Anonymous. Focal points. IEEE Spectrum, 7(9):14–17, September 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FPj

[Ano70-257] Anonymous. Focal points. IEEE Spectrum, 7(10):14–17, October 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FPk

[Ano70-258] Anonymous. Focal points. IEEE Spectrum, 7(11):18–23, November 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FPl

[Ano70-259] Anonymous. Focal points. IEEE Spectrum, 7(12):16–20, December 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Fa

[Ano70-260] Anonymous. Forum. IEEE Spectrum, 7(1):6–13, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Fb

[Ano70-261] Anonymous. Forum. IEEE Spectrum, 7(2):6–9, February 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Fd

[Ano70-262] Anonymous. Forum. IEEE Spectrum, 7(3):6–11, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Fe

[Ano70-263] Anonymous. Forum. IEEE Spectrum, 7(4):6–8, April 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ff

[Ano70-264] Anonymous. Forum. IEEE Spectrum, 7(5):6–8, May 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Fg

[Ano70-265] Anonymous. Forum. IEEE Spectrum, 7(6):6–9, June 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Fh

[Ano70-266] Anonymous. Forum. IEEE Spectrum, 7(7):6–9, July 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Fi

[Ano70-267] Anonymous. Forum. IEEE Spectrum, 7(8):8–10, August 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Fj

[Ano70-268] Anonymous. Forum. IEEE Spectrum, 7(9):6–10, September 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Fk

[Ano70-269] Anonymous. Forum. IEEE Spectrum, 7(10):6–8, October 1970.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Fl

[Ano70-270] Anonymous. Forum. IEEE Spectrum, 7(11):8–11, November 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Fm

[Ano70-271] Anonymous. Forum. IEEE Spectrum, 7(12):6–10, December 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FCa

[Ano70-272] Anonymous. [front cover]. IEEE Spectrum, 7(1):c1, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FCb

[Ano70-273] Anonymous. [front cover]. IEEE Spectrum, 7(1):c2, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FCc

[Ano70-274] Anonymous. [front cover].

IEEE Spectrum, 7(2):c1, February 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FCd

[Ano70-275] Anonymous. [front cover].

IEEE Spectrum, 7(2):c2, February 1970. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FCe

[Ano70-276] Anonymous. [front cover]. IEEE Spectrum, 7(3):c1, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FCf

[Ano70-277] Anonymous. [front cover]. IEEE Spectrum, 7(3):c2, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FCg

[Ano70-278] Anonymous. [front cover]. IEEE Spectrum, 7(4):c1, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FCh

[Ano70-279] Anonymous. [front cover]. IEEE Spectrum, 7(4):c2, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FCi

[Ano70-280] Anonymous. [front cover]. IEEE Spectrum, 7(5):c1, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FCj

[Ano70-281] Anonymous. [front cover]. IEEE Spectrum, 7(5):c2, May 1970.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FCk

[Ano70-282] Anonymous. [front cover]. IEEE Spectrum, 7(6):c1, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FCl

[Ano70-283] Anonymous. [front cover]. IEEE Spectrum, 7(6):c2, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FCm

[Ano70-284] Anonymous. [front cover]. IEEE Spectrum, 7(7):c1, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FCn

[Ano70-285] Anonymous. [front cover]. IEEE Spectrum, 7(7):c2, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FCo

[Ano70-286] Anonymous. [front cover]. IEEE Spectrum, 7(8):c1, August 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FCp

[Ano70-287] Anonymous. [front cover]. IEEE Spectrum, 7(8):c2, August 1970.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FCq

[Ano70-288] Anonymous. [front cover]. IEEE Spectrum, 7(9):c1, September 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FCr

[Ano70-289] Anonymous. [front cover]. IEEE Spectrum, 7(9):c2, September 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FCs

[Ano70-290] Anonymous. [front cover]. IEEE Spectrum, 7(10):c1, October 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FCt

[Ano70-291] Anonymous. [front cover]. IEEE Spectrum, 7(10):c2, October 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FCu

[Ano70-292] Anonymous. [front cover]. IEEE Spectrum, 7(11):c1, November 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FCv

[Ano70-293] Anonymous. [front cover]. IEEE Spectrum, 7(11):c2, November 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FCw

[Ano70-294] Anonymous. [front cover]. IEEE Spectrum, 7(12):c1, December 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:FCx

[Ano70-295] Anonymous. [front cover]. IEEE Spectrum, 7(12):c2, December 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:HSI

[Ano70-296] Anonymous. High-sensitivity infrared heterodyne receivers with gigahertz IF bandwidth Part III: Packaged 10.6-nm heterodyne 1.2-GHz bandwidth receiver for CO_{2} laser radar. IEEE Spectrum, 7(1):5, January 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:IIC

[Ano70-297] Anonymous. IEEE International convention and Exhibition, March 13–26, ’70 — Exhibitors. IEEE Spectrum, 7 (3):134–207, March 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:IAa

[Ano70-298] Anonymous. Index to advertisers. IEEE Spectrum, 7(1): 126–128, January 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:IAb

[Ano70-299] Anonymous. Index to advertisers. IEEE Spectrum, 7(2): 134–136, February 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:IAc

[Ano70-300] Anonymous. Index to advertisers. IEEE Spectrum, 7(3): 212–216, March 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:IAd

[Ano70-301] Anonymous. Index to advertisers. IEEE Spectrum, 7(4): 126–128, April 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:IAe

[Ano70-302] Anonymous. Index to advertisers. IEEE Spectrum, 7(5):120, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:IAf

[Ano70-303] Anonymous. Index to advertisers. IEEE Spectrum, 7 (6):119–120, June 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:IAg

[Ano70-304] Anonymous. Index to advertisers. IEEE Spectrum, 7(7):120, July 1970. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:IAh

[Ano70-305] Anonymous. Index to advertisers. IEEE Spectrum, 7(8): 126–128, August 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:IAi

[Ano70-306] Anonymous. Index to advertisers. IEEE Spectrum, 7(9): 119–120, September 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:IAj

[Ano70-307] Anonymous. Index to advertisers. IEEE Spectrum, 7(10): 118–120, October 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:IAk

[Ano70-308] Anonymous. Index to advertisers. IEEE Spectrum, 7(11): 118–120, November 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:IAl

[Ano70-309] Anonymous. Index to advertisers. IEEE Spectrum, 7(12): 110–112, December 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:IAS

[Ano70-310] Anonymous. Information about the society. IEEE Spectrum, 7 (3):208–211, March 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:II

[Ano70-311] Anonymous. Inside IEEE. IEEE Spectrum, 7(12):89–92, December 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:IEEa

[Ano70-312] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 7(1):4, January 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:IEEb

[Ano70-313] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 7(2):4, February 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:IEEc

[Ano70-314] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 7(3):4, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:IEEd

[Ano70-315] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 7(4): 4, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:IEEe

[Ano70-316] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 7(5): 4, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:IEEf

[Ano70-317] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 7(6): 4, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:IEEg

[Ano70-318] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 7(7): 4, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:IEEh

[Ano70-319] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 7(8):4, August 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:IEEi

[Ano70-320] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 7 (9):4, September 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:IEEj

[Ano70-321] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 7(10):4, October 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:IEEk

[Ano70-322] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 7 (11):4, November 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:IEEl

[Ano70-323] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 7 (12):4, December 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:I

[Ano70-324] Anonymous. Instruments. IEEE Spectrum, 7(1):16i–16i, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:JTP

[Ano70-325] Anonymous. Japan at a turning point: a view of the seventies. IEEE Spectrum, 7(11): 27–117, November 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:LFC

[Ano70-326] Anonymous. Low frequency cable shield leakage. IEEE Spectrum, 7(3):5, March 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NBA

[Ano70-327] Anonymous. New book announcement [advertisement]. IEEE Spectrum, 7(6):93, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NPAa

[Ano70-328] Anonymous. New product applications. IEEE Spectrum, 7 (1):86–91, January 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NPAb

[Ano70-329] Anonymous. New product applications. IEEE Spectrum, 7 (2):95–100, February 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NPAc

[Ano70-330] Anonymous. New product applications. IEEE Spectrum, 7 (3):90–96, March 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NPAd

[Ano70-331] Anonymous. New product applications. IEEE Spectrum, 7 (4):84–89, April 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NPAe

[Ano70-332] Anonymous. New product applications. IEEE Spectrum, 7 (5):81–86, May 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NPAf

[Ano70-333] Anonymous. New product applications. IEEE Spectrum, 7 (6):87–92, June 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NPAg

[Ano70-334] Anonymous. New product applications. IEEE Spectrum, 7 (7):84–89, July 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NPAh

[Ano70-335] Anonymous. New product applications. IEEE Spectrum, 7 (8):95–100, August 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NPAi

[Ano70-336] Anonymous. New product applications. IEEE Spectrum, 7 (9):85–90, September 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NPAj

[Ano70-337] Anonymous. New product applications. IEEE Spectrum, 7 (10):87–92, October 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NPAk

[Ano70-338] Anonymous. New product applications. IEEE Spectrum, 7 (11):80–85, November 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NPAl

[Ano70-339] Anonymous. New product applications. IEEE Spectrum, 7 (12):76–81, December 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NWa

[Ano70-340] Anonymous. News from Washington. IEEE Spectrum, 7(1): 15–16, January 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NWb

[Ano70-341] Anonymous. News from Washington. IEEE Spectrum, 7(2): 11–12, February 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NWc

[Ano70-342] Anonymous. News from Washington. IEEE Spectrum, 7 (3):15–16, March 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NWd

[Ano70-343] Anonymous. News from Washington. IEEE Spectrum, 7 (4):11–12, April 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NWe

[Ano70-344] Anonymous. News from Washington. IEEE Spectrum, 7 (5):9–10, May 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NWf

[Ano70-345] Anonymous. News from Washington. IEEE Spectrum, 7 (6):11–12, June 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NWg

[Ano70-346] Anonymous. News from Washington. IEEE Spectrum, 7(7):11– 12, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NWh

[Ano70-347] Anonymous. News from Washington. IEEE Spectrum, 7 (8):15–16, August 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NWi

[Ano70-348] Anonymous. News from Washington. IEEE Spectrum, 7(9): 11–12, September 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NWj

[Ano70-349] Anonymous. News from Washington. IEEE Spectrum, 7(10): 11–12, October 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NWk

[Ano70-350] Anonymous. News from Washington. IEEE Spectrum, 7(11):

15–16, November 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NWl

[Ano70-351] Anonymous. News from Washington. IEEE Spectrum, 7(12): 13–14, December 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NIa

[Ano70-352] Anonymous. News of the IEEE. IEEE Spectrum, 7(1): 112–121, January 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NIb

[Ano70-353] Anonymous. News of the IEEE. IEEE Spectrum, 7(2): 118–124, February 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NIc

[Ano70-354] Anonymous. News of the IEEE.

IEEE Spectrum, 7(3):120–127, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NId

[Ano70-355] Anonymous. News of the IEEE.

IEEE Spectrum, 7(4):108–119, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NIe

[Ano70-356] Anonymous. News of the IEEE.

IEEE Spectrum, 7(5):104–110, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NIf

[Ano70-357] Anonymous. News of the IEEE. IEEE Spectrum, 7(6):110–114, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NIg

[Ano70-358] Anonymous. News of the IEEE. IEEE Spectrum, 7(7):102–111, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NIh

[Ano70-359] Anonymous. News of the IEEE. IEEE Spectrum, 7(8):113–121, August 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NIi

[Ano70-360] Anonymous. News of the IEEE. IEEE Spectrum, 7(9): 104–113, September 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NIj

[Ano70-361] Anonymous. News of the IEEE. IEEE Spectrum, 7(10): 106–113, October 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:NIk

[Ano70-362] Anonymous. News of the IEEE. IEEE Spectrum, 7(11):

94–107, November 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Oa

[Ano70-363] Anonymous. Obituaries. IEEE Spectrum, 7(1):124–125, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ob

[Ano70-364] Anonymous. Obituaries. IEEE Spectrum, 7(2):126–127, February 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Oc

[Ano70-365] Anonymous. Obituaries. IEEE Spectrum, 7(3):132–133, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Od

[Ano70-366] Anonymous. Obituaries. IEEE Spectrum, 7(4):124–125, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Oe

[Ano70-367] Anonymous. Obituaries. IEEE Spectrum, 7(5):117, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Of

[Ano70-368] Anonymous. Obituaries. IEEE Spectrum, 7(6):116–117, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Og

[Ano70-369] Anonymous. Obituaries. IEEE Spectrum, 7(7):113, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Oh

[Ano70-370] Anonymous. Obituaries. IEEE Spectrum, 7(8):125, August 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Oi

[Ano70-371] Anonymous. Obituaries. IEEE Spectrum, 7(9):117, September 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Oj

[Ano70-372] Anonymous. Obituaries. IEEE Spectrum, 7(10):117, October 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ok

[Ano70-373] Anonymous. Obituaries. IEEE Spectrum, 7(11):109, November 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ol

[Ano70-374] Anonymous. Obituaries. IEEE Spectrum, 7(12):95, December 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:OBE

[Ano70-375] Anonymous. An octave band electronically steerable phased array antenna. IEEE Spectrum, 7(4):5, April 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Pa

[Ano70-376] Anonymous. People. IEEE Spectrum, 7(1):122–124, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Pb

[Ano70-377] Anonymous. People. IEEE Spectrum, 7(2):125–126, February 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Pc

[Ano70-378] Anonymous. People. IEEE Spectrum, 7(3):130–132, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Pd

[Ano70-379] Anonymous. People. IEEE Spectrum, 7(4):121–124, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Pe

[Ano70-380] Anonymous. People. IEEE Spectrum, 7(5):113–117, May 1970.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Pf

[Ano70-381] Anonymous. People. IEEE Spectrum, 7(6):115, June 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Pg

[Ano70-382] Anonymous. People. IEEE Spectrum, 7(7):112–113, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Ph

[Ano70-383] Anonymous. People. IEEE Spectrum, 7(8):122–125, August 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Pi

[Ano70-384] Anonymous. People. IEEE Spectrum, 7(9):114–117, September 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Pj

[Ano70-385] Anonymous. People. IEEE Spectrum, 7(10):114–117, October 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Pk

[Ano70-386] Anonymous. People. IEEE Spectrum, 7(11):108–109, November 1970. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:Pl

[Ano70-387] Anonymous. People. IEEE Spectrum, 7(12):94–95, December 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:PCSa

[Ano70-388] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 7(1):127, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:PCSb

[Ano70-389] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 7(2):135, February 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:PCSc

[Ano70-390] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 7(3):215, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:PCSd

[Ano70-391] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 7(4):127–128, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:PCSe

[Ano70-392] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 7(5):118–119, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:PCSf

[Ano70-393] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 7(6):118, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:PCSg

[Ano70-394] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 7(8):127, August 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:PCSh

[Ano70-395] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 7(9):118, September 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:PCSi

[Ano70-396] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 7(10):119, October 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:PCSj

[Ano70-397] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 7(11):117, November 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:PCSk

[Ano70-398] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 7(12):111, December 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:RPA

[Ano70-399] Anonymous. A retrodirective phased array for the data relay satellites. IEEE Spectrum, 7(2):5, February 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SIa

[Ano70-400] Anonymous. Scanning the issues.

IEEE Spectrum, 7(1):92–93, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SIb

[Ano70-401] Anonymous. Scanning the issues. IEEE Spectrum, 7(2): 101–102, February 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SIc

[Ano70-402] Anonymous. Scanning the issues. IEEE Spectrum, 7(3): 98–99, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SId

[Ano70-403] Anonymous. Scanning the issues. IEEE Spectrum, 7(4): 90–92, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SIe

[Ano70-404] Anonymous. Scanning the issues. IEEE Spectrum, 7(5):88–89, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SIf

[Ano70-405] Anonymous. Scanning the issues. IEEE Spectrum, 7(6): 94–95, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SIg

[Ano70-406] Anonymous. Scanning the issues. IEEE Spectrum, 7(7):90– 92, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SIh

[Ano70-407] Anonymous. Scanning the issues. IEEE Spectrum, 7(8): 101–103, August 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SIi

[Ano70-408] Anonymous. Scanning the issues. IEEE Spectrum, 7(9):

91–93, September 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SIj

[Ano70-409] Anonymous. Scanning the issues. IEEE Spectrum, 7(10): 93–95, October 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SIk

[Ano70-410] Anonymous. Scanning the issues. IEEE Spectrum, 7(11): 86–88, November 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SIl

[Ano70-411] Anonymous. Scanning the issues. IEEE Spectrum, 7(12): 82–84, December 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SPa

[Ano70-412] Anonymous. Special publications. IEEE Spectrum, 7(1): 102–103, January 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SPb

[Ano70-413] Anonymous. Special publications. IEEE Spectrum, 7 (2):112, February 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SPc

[Ano70-414] Anonymous. Special publications. IEEE Spectrum, 7(3):

112–113, March 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SPd

[Ano70-415] Anonymous. Special publications. IEEE Spectrum, 7(4): 100–101, April 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SPe

[Ano70-416] Anonymous. Special publications. IEEE Spectrum, 7 (5):96–97, May 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SPf

[Ano70-417] Anonymous. Special publications. IEEE Spectrum, 7(6):103, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SPg

[Ano70-418] Anonymous. Special publications. IEEE Spectrum, 7(7):94– 95, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SPh

[Ano70-419] Anonymous. Special publications. IEEE Spectrum, 7(8):104, August 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SPi

[Ano70-420] Anonymous. Special publications. IEEE Spectrum, 7(9):

94–95, September 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SPj

[Ano70-421] Anonymous. Special publications. IEEE Spectrum, 7(10): 96–97, October 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SPk

[Ano70-422] Anonymous. Special publications. IEEE Spectrum, 7 (11):89, November 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SPl

[Ano70-423] Anonymous. Special publications. IEEE Spectrum, 7 (12):85, December 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SLa

[Ano70-424] Anonymous. Spectral lines.

IEEE Spectrum, 7(1):31, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SLb

[Ano70-425] Anonymous. Spectral lines.

IEEE Spectrum, 7(2):39, February 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SLc

[Ano70-426] Anonymous. Spectral lines.

IEEE Spectrum, 7(3):31, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SLd

[Ano70-427] Anonymous. Spectral lines. IEEE Spectrum, 7(4):23, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SLe

[Ano70-428] Anonymous. Spectral lines. IEEE Spectrum, 7(5):21, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SLf

[Ano70-429] Anonymous. Spectral lines. IEEE Spectrum, 7(6):25, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SLg

[Ano70-430] Anonymous. Spectral lines. IEEE Spectrum, 7(7):23, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SLh

[Ano70-431] Anonymous. Spectral lines. IEEE Spectrum, 7(8):29, August 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SLi

[Ano70-432] Anonymous. Spectral lines. IEEE Spectrum, 7(9):21, September 1970. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SLj

[Ano70-433] Anonymous. Spectral lines.

IEEE Spectrum, 7(10):21, October 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SLk

[Ano70-434] Anonymous. Spectral lines.

IEEE Spectrum, 7(11):31, November 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:SLl

[Ano70-435] Anonymous. Spectral lines.

IEEE Spectrum, 7(12):24, December 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:TCa

[Ano70-436] Anonymous. Table of contents.

IEEE Spectrum, 7(1):2–3, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:TCb

[Ano70-437] Anonymous. Table of contents.

IEEE Spectrum, 7(2):2–3, February 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:TCc

[Ano70-438] Anonymous. Table of contents. IEEE Spectrum, 7(3):2–3, March 1970. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:TCd

[Ano70-439] Anonymous. Table of contents. IEEE Spectrum, 7(4):2–3, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:TCe

[Ano70-440] Anonymous. Table of contents. IEEE Spectrum, 7(5):2– 3, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:TCf

[Ano70-441] Anonymous. Table of contents. IEEE Spectrum, 7(6):2–3, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:TCg

[Ano70-442] Anonymous. Table of contents. IEEE Spectrum, 7(7):2– 3, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:TCh

[Ano70-443] Anonymous. Table of contents. IEEE Spectrum, 7(8):2–3, August 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:TCi

[Ano70-444] Anonymous. Table of contents. IEEE Spectrum, 7(9):

2–3, September 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:TCj

[Ano70-445] Anonymous. Table of contents.

IEEE Spectrum, 7(10):2–3, October 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:TCk

[Ano70-446] Anonymous. Table of contents. IEEE Spectrum, 7(11): 2–3, November 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:TCl

[Ano70-447] Anonymous. Table of contents.

IEEE Spectrum, 7(12):2–3, December 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:TP

[Ano70-448] Anonymous. Technical program. IEEE Spectrum, 7(2): 21–37, February 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:TJa

[Ano70-449] Anonymous. Translated journals. IEEE Spectrum, 7(1): 99–101, January 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:TJb

[Ano70-450] Anonymous. Translated journals. IEEE Spectrum, 7(2):

110–111, February 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:TJc

[Ano70-451] Anonymous. Translated journals. IEEE Spectrum, 7(3):110, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:TJd

[Ano70-452] Anonymous. Translated journals. IEEE Spectrum, 7(4):98–99, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:TJe

[Ano70-453] Anonymous. Translated journals. IEEE Spectrum, 7(5):94, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:TJf

[Ano70-454] Anonymous. Translated journals. IEEE Spectrum, 7(6): 101–102, June 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1970:UEI

[Ano70-455] Anonymous. UK exhibits at IEEE ’70 [advertisement]. IEEE Spectrum, 7(1):16m–16o, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:GFP

[Ano71a] Anonymous. 1971 Group Flights Program. IEEE Spectrum, 8 (3):127, March 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IIG

[Ano71b] Anonymous. 1971 IEEE International Geoscience Electronics Symposium. IEEE Spectrum, 8(7):5, July 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:APa

[Ano71c] Anonymous. Advance program. IEEE Spectrum, 8(9): 21–28, September 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:ATCa

[Ano71d] Anonymous. Advance tables of contents. IEEE Spectrum, 8 (1):84–87, January 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:ATCb

[Ano71e] Anonymous. Advance tables of contents. IEEE Spectrum, 8(2): 104–107, February 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:ATCc

[Ano71f] Anonymous. Advance tables of contents. IEEE Spectrum, 8 (3):98–101, March 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:ATCe

[Ano71g] Anonymous. Advance tables of contents. IEEE Spectrum, 8(5):96–98, May 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:ATCf

[Ano71h] Anonymous. Advance tables of contents. IEEE Spectrum, 8(6):77–79, June 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:ATCg

[Ano71i] Anonymous. Advance tables of contents. IEEE Spectrum, 8(7):76–79, July 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:ATCh

[Ano71j] Anonymous. Advance tables of contents. IEEE Spectrum, 8 (8):84–87, August 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:ATCi

[Ano71k] Anonymous. Advance tables of contents. IEEE Spectrum, 8(9): 92–96, September 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:ATCj

[Ano71l] Anonymous. Advance tables of contents. IEEE Spectrum, 8 (10):81–85, October 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:ATCk

[Ano71m] Anonymous. Advance tables of contents. IEEE Spectrum, 8(11):

91–92, November 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:ATCl

[Ano71n] Anonymous. Advance tables of contents. IEEE Spectrum, 8(12): 81–82, December 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:ATCd

[Ano71o] Anonymous. Advances tables of contents. IEEE Spectrum, 8(4):79–83, April 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Ab

[Ano71p] Anonymous. [advertisement].

IEEE Spectrum, 8(2):nil1–1, February 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Ac

[Ano71q] Anonymous. [advertisement].

IEEE Spectrum, 8(2):5, February 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Ad

[Ano71r] Anonymous. [advertisement].

IEEE Spectrum, 8(2):7–8, February 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Ae

[Ano71s] Anonymous. [advertisement].

IEEE Spectrum, 8(2):13, February 1971. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Af

[Ano71t] Anonymous. [advertisement]. IEEE Spectrum, 8(2):21–40, February 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Ag

[Ano71u] Anonymous. [advertisement]. IEEE Spectrum, 8(2):61–68, February 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Ah

[Ano71v] Anonymous. [advertisement]. IEEE Spectrum, 8(2):108–109, February 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Ai

[Ano71w] Anonymous. [advertisement]. IEEE Spectrum, 8(2):119, February 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Aj

[Ano71x] Anonymous. [advertisement]. IEEE Spectrum, 8(2):129–130, February 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Ak

[Ano71y] Anonymous. [advertisement]. IEEE Spectrum, 8(2):nil4, February 1971. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Al

[Ano71z] Anonymous. [advertisement].

IEEE Spectrum, 8(3):nil1, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Am

[Ano71-27] Anonymous. [advertisement].

IEEE Spectrum, 8(3):1, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:An

[Ano71-28] Anonymous. [advertisement].

IEEE Spectrum, 8(3):5–7, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Ao

[Ano71-29] Anonymous. [advertisement].

IEEE Spectrum, 8(3):9, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Ap

[Ano71-30] Anonymous. [advertisement].

IEEE Spectrum, 8(3):13, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Aq

[Ano71-31] Anonymous. [advertisement].

IEEE Spectrum, 8(3):15–17, March 1971. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Ar

[Ano71-32] Anonymous. [advertisement]. IEEE Spectrum, 8(3):21–36, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:As

[Ano71-33] Anonymous. [advertisement]. IEEE Spectrum, 8(3):73–76, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:At

[Ano71-34] Anonymous. [advertisement]. IEEE Spectrum, 8(3):102–103, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Au

[Ano71-35] Anonymous. [advertisement]. IEEE Spectrum, 8(3):122–126, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Av

[Ano71-36] Anonymous. [advertisement]. IEEE Spectrum, 8(3):nil4, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Aw

[Ano71-37] Anonymous. [advertisement]. IEEE Spectrum, 8(4):nil1–1, April 1971. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Ax

[Ano71-38] Anonymous. [advertisement].

IEEE Spectrum, 8(4):5–6, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Ay

[Ano71-39] Anonymous. [advertisement].

IEEE Spectrum, 8(4):12–14, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Az

[Ano71-40] Anonymous. [advertisement].

IEEE Spectrum, 8(4):51–54, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Aba

[Ano71-41] Anonymous. [advertisement].

IEEE Spectrum, 8(4):83, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Abb

[Ano71-42] Anonymous. [advertisement].

IEEE Spectrum, 8(4):84–85, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Abc

[Ano71-43] Anonymous. [advertisement].

IEEE Spectrum, 8(4):93–97, April 1971. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Abd

[Ano71-44] Anonymous. [advertisement]. IEEE Spectrum, 8(4):105–106, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Abe

[Ano71-45] Anonymous. [advertisement]. IEEE Spectrum, 8(4):nil4, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Abf

[Ano71-46] Anonymous. [advertisement]. IEEE Spectrum, 8(5):nil1–1, May 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Abg

[Ano71-47] Anonymous. [advertisement]. IEEE Spectrum, 8(5):5, May 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Abh

[Ano71-48] Anonymous. [advertisement]. IEEE Spectrum, 8(5):10–11, May 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Abi

[Ano71-49] Anonymous. [advertisement]. IEEE Spectrum, 8(5):13–18, May 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Abj

[Ano71-50] Anonymous. [advertisement].

IEEE Spectrum, 8(5):21–30, May 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Abk

[Ano71-51] Anonymous. [advertisement].

IEEE Spectrum, 8(5):59–62, May 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Abl

[Ano71-52] Anonymous. [advertisement].

IEEE Spectrum, 8(5):99–100, May 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Abm

[Ano71-53] Anonymous. [advertisement].

IEEE Spectrum, 8(5):104, May 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Abn

[Ano71-54] Anonymous. [advertisement].

IEEE Spectrum, 8(5):111–113, May 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Abo

[Ano71-55] Anonymous. [advertisement].

IEEE Spectrum, 8(5):121–122, May 1971. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Abp

[Ano71-56] Anonymous. [advertisement]. IEEE Spectrum, 8(5):nil4, May 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Abq

[Ano71-57] Anonymous. [advertisement]. IEEE Spectrum, 8(6):nil1, June 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Abr

[Ano71-58] Anonymous. [advertisement]. IEEE Spectrum, 8(6):1, June 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Abs

[Ano71-59] Anonymous. [advertisement]. IEEE Spectrum, 8(6):47–50, June 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Abt

[Ano71-60] Anonymous. [advertisement]. IEEE Spectrum, 8(6):83, June 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Abu

[Ano71-61] Anonymous. [advertisement]. IEEE Spectrum, 8(6):85, June 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Abv

[Ano71-62] Anonymous. [advertisement].

IEEE Spectrum, 8(6):97–98, June 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Abw

[Ano71-63] Anonymous. [advertisement].

IEEE Spectrum, 8(6):nil4, June 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Abx

[Ano71-64] Anonymous. [advertisement].

IEEE Spectrum, 8(7):nil1, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Aby

[Ano71-65] Anonymous. [advertisement].

IEEE Spectrum, 8(7):1, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Abz

[Ano71-66] Anonymous. [advertisement].

IEEE Spectrum, 8(7):4, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Aca

[Ano71-67] Anonymous. [advertisement].

IEEE Spectrum, 8(7):6, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Acb

[Ano71-68] Anonymous. [advertisement]. IEEE Spectrum, 8(7):15–16, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Acc

[Ano71-69] Anonymous. [advertisement]. IEEE Spectrum, 8(7):49–52, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Acd

[Ano71-70] Anonymous. [advertisement]. IEEE Spectrum, 8(7):85, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Ace

[Ano71-71] Anonymous. [advertisement]. IEEE Spectrum, 8(7):101–102, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Acf

[Ano71-72] Anonymous. [advertisement]. IEEE Spectrum, 8(8):4–5, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Acg

[Ano71-73] Anonymous. [advertisement]. IEEE Spectrum, 8(8):6, August 1971. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Ach

[Ano71-74] Anonymous. [advertisement].

IEEE Spectrum, 8(8):11–12, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Aci

[Ano71-75] Anonymous. [advertisement].

IEEE Spectrum, 8(8):53–56, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Acj

[Ano71-76] Anonymous. [advertisement].

IEEE Spectrum, 8(8):97–101, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Ack

[Ano71-77] Anonymous. [advertisement].

IEEE Spectrum, 8(8):109–110, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Acl

[Ano71-78] Anonymous. [advertisement].

IEEE Spectrum, 8(8):nil4, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Acm

[Ano71-79] Anonymous. [advertisement].

IEEE Spectrum, 8(9):nil1–1, September 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Acn

[Ano71-80] Anonymous. [advertisement]. IEEE Spectrum, 8(9):4–5, September 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Aco

[Ano71-81] Anonymous. [advertisement]. IEEE Spectrum, 8(9):6, September 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Acp

[Ano71-82] Anonymous. [advertisement]. IEEE Spectrum, 8(9):14–18, September 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Acq

[Ano71-83] Anonymous. [advertisement]. IEEE Spectrum, 8(9):59–62, September 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Acr

[Ano71-84] Anonymous. [advertisement]. IEEE Spectrum, 8(9):103, September 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Acs

[Ano71-85] Anonymous. [advertisement]. IEEE Spectrum, 8(9):106, September 1971. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Act

[Ano71-86] Anonymous. [advertisement].

IEEE Spectrum, 8(9):111–115, September 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Acu

[Ano71-87] Anonymous. [advertisement].

IEEE Spectrum, 8(9):123–124, September 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Acv

[Ano71-88] Anonymous. [advertisement].

IEEE Spectrum, 8(9):nil4, September 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Acw

[Ano71-89] Anonymous. [advertisement].

IEEE Spectrum, 8(10):4, October 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Acx

[Ano71-90] Anonymous. [advertisement].

IEEE Spectrum, 8(10):5–6, October 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Acy

[Ano71-91] Anonymous. [advertisement].

IEEE Spectrum, 8(10):15–16, October 1971. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Acz

[Ano71-92] Anonymous. [advertisement]. IEEE Spectrum, 8(10):51–54, October 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Ada

[Ano71-93] Anonymous. [advertisement]. IEEE Spectrum, 8(10):107– 108, October 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Adb

[Ano71-94] Anonymous. [advertisement]. IEEE Spectrum, 8(10):nil2, October 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Adc

[Ano71-95] Anonymous. [advertisement]. IEEE Spectrum, 8(11):nil1–1, November 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Add

[Ano71-96] Anonymous. [advertisement]. IEEE Spectrum, 8(11):4, November 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Ade

[Ano71-97] Anonymous. [advertisement]. IEEE Spectrum, 8(11):5–6, November 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Adf

[Ano71-98] Anonymous. [advertisement].

IEEE Spectrum, 8(11):15–18, November 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Adg

[Ano71-99] Anonymous. [advertisement].

IEEE Spectrum, 8(11):21–22, November 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Adh

[Ano71-100] Anonymous. [advertisement].

IEEE Spectrum, 8(11):55–58, November 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Adi

[Ano71-101] Anonymous. [advertisement].

IEEE Spectrum, 8(11):nil2, November 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Adj

[Ano71-102] Anonymous. [advertisement].

IEEE Spectrum, 8(12):nil1, December 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Adk

[Ano71-103] Anonymous. [advertisement].

IEEE Spectrum, 8(12):1, December 1971. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Adl

[Ano71-104] Anonymous. [advertisement]. IEEE Spectrum, 8(12):4, December 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Adm

[Ano71-105] Anonymous. [advertisement]. IEEE Spectrum, 8(12):5–6, December 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Adn

[Ano71-106] Anonymous. [advertisement]. IEEE Spectrum, 8(12):11–13, December 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Ado

[Ano71-107] Anonymous. [advertisement]. IEEE Spectrum, 8(12):59–62, December 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Adp

[Ano71-108] Anonymous. [advertisement]. IEEE Spectrum, 8(12):90, December 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Adq

[Ano71-109] Anonymous. [advertisement]. IEEE Spectrum, 8(12):95, December 1971. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Adr

[Ano71-110] Anonymous. [advertisement].

IEEE Spectrum, 8(12):106–108, December 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Ads

[Ano71-111] Anonymous. [advertisement].

IEEE Spectrum, 8(12):121–122, December 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Aa

[Ano71-112] Anonymous. [advertisements].

IEEE Spectrum, 8(1):5, January 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:AII

[Ano71-113] Anonymous. Annual index — IEEE Spectrum. IEEE Spectrum, 8(12):109–119, December 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:ALa

[Ano71-114] Anonymous. Applications literature. IEEE Spectrum, 8(8):80, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:ALb

[Ano71-115] Anonymous. Applications literature. IEEE Spectrum, 8 (9):89, September 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:ALc

[Ano71-116] Anonymous. Applications literature. IEEE Spectrum, 8 (10):78, October 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:ALd

[Ano71-117] Anonymous. Applications literature. IEEE Spectrum, 8 (11):88, November 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:ALe

[Ano71-118] Anonymous. Applications literature. IEEE Spectrum, 8 (12):78, December 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:LOY

[Ano71-119] Anonymous. At least one of your interests is now served by one of IEEE’s Groups. IEEE Spectrum, 8(3):31, March 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:AGI

[Ano71-120] Anonymous. An author’s guide to IEEE publications. IEEE Spectrum, 8(1):93–94, January 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:BC

[Ano71-121] Anonymous. [back cover]. IEEE Spectrum, 8(1):c4, January 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:BRa

[Ano71-122] Anonymous. Book reviews.

IEEE Spectrum, 8(1):90–92, January 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:BRb

[Ano71-123] Anonymous. Book reviews.

IEEE Spectrum, 8(2):111–112, February 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:BRc

[Ano71-124] Anonymous. Book reviews.

IEEE Spectrum, 8(3):105–111, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:BRd

[Ano71-125] Anonymous. Book reviews.

IEEE Spectrum, 8(4):87, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:BRe

[Ano71-126] Anonymous. Book reviews.

IEEE Spectrum, 8(5):102–104, May 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:BRf

[Ano71-127] Anonymous. Book reviews. IEEE Spectrum, 8(6):82–83, June 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:BRg

[Ano71-128] Anonymous. Book reviews. IEEE Spectrum, 8(7):82–85, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:BRh

[Ano71-129] Anonymous. Book reviews. IEEE Spectrum, 8(8):90–92, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:BRi

[Ano71-130] Anonymous. Book reviews. IEEE Spectrum, 8(9):99–103, September 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:BRj

[Ano71-131] Anonymous. Book reviews. IEEE Spectrum, 8(10):88–89, October 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:BRk

[Ano71-132] Anonymous. Book reviews. IEEE Spectrum, 8(11):98–99, November 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:BRl

[Ano71-133] Anonymous. Book reviews.

IEEE Spectrum, 8(12):89–90, December 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:ISRa

[Ano71-134] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 8(3):149–150, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:ISRb

[Ano71-135] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 8(4):105–106, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:ISRc

[Ano71-136] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 8(5):121–122, May 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:ISRd

[Ano71-137] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 8(7):101–102, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:ISRe

[Ano71-138] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 8(8):109–110, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:ISRf

[Ano71-139] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 8(12):121–122, December 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:BPa

[Ano71-140] Anonymous. [breaker page]. IEEE Spectrum, 8(2):4, February 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:BPb

[Ano71-141] Anonymous. [breaker page]. IEEE Spectrum, 8(3):nil2, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:BPc

[Ano71-142] Anonymous. [breaker page]. IEEE Spectrum, 8(3):4, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:BPd

[Ano71-143] Anonymous. [breaker page]. IEEE Spectrum, 8(4):4, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:BPe

[Ano71-144] Anonymous. [breaker page].

IEEE Spectrum, 8(5):4, May 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:BPf

[Ano71-145] Anonymous. [breaker page].

IEEE Spectrum, 8(6):nil2, June 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:BPg

[Ano71-146] Anonymous. [breaker page].

IEEE Spectrum, 8(6):4, June 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:BPh

[Ano71-147] Anonymous. [breaker page].

IEEE Spectrum, 8(7):nil2, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:BPi

[Ano71-148] Anonymous. [breaker page].

IEEE Spectrum, 8(7):4, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:BPj

[Ano71-149] Anonymous. [breaker page].

IEEE Spectrum, 8(8):4, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:BPk

[Ano71-150] Anonymous. [breaker page]. IEEE Spectrum, 8(9):4, September 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:BPl

[Ano71-151] Anonymous. [breaker page]. IEEE Spectrum, 8(10):4, October 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:BPm

[Ano71-152] Anonymous. [breaker page]. IEEE Spectrum, 8(11):4, November 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:BPn

[Ano71-153] Anonymous. [breaker page]. IEEE Spectrum, 8(12):nil2, December 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:BPo

[Ano71-154] Anonymous. [breaker page]. IEEE Spectrum, 8(12):4, December 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Cb

[Ano71-155] Anonymous. Calendar. IEEE Spectrum, 8(1):19–20, January 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Cc

[Ano71-156] Anonymous. Calendar. IEEE Spectrum, 8(2):15–16, February 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Cd

[Ano71-157] Anonymous. Calendar. IEEE Spectrum, 8(3):19–20, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Ce

[Ano71-158] Anonymous. Calendar. IEEE Spectrum, 8(4):99–100, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Cf

[Ano71-159] Anonymous. Calendar. IEEE Spectrum, 8(5):115–116, May 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Cg

[Ano71-160] Anonymous. Calendar. IEEE Spectrum, 8(6):92–93, June 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Ch

[Ano71-161] Anonymous. Calendar. IEEE Spectrum, 8(7):95–96, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Ci

[Ano71-162] Anonymous. Calendar. IEEE Spectrum, 8(8):102–103, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Cj

[Ano71-163] Anonymous. Calendar. IEEE Spectrum, 8(9):116–117, September 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Ck

[Ano71-164] Anonymous. Calendar. IEEE Spectrum, 8(10):99–100, October 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Cl

[Ano71-165] Anonymous. Calendar. IEEE Spectrum, 8(11):107–108, November 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Cm

[Ano71-166] Anonymous. Calendar. IEEE Spectrum, 8(12):101–102, December 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:CAa

[Ano71-167] Anonymous. Classified advertising. IEEE Spectrum, 8(3):101, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:CAb

[Ano71-168] Anonymous. Classified advertising. IEEE Spectrum, 8(4):91, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:CAc

[Ano71-169] Anonymous. Classified advertising. IEEE Spectrum, 8(6):93, June 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:CAd

[Ano71-170] Anonymous. Classified advertising. IEEE Spectrum, 8(7):87, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:CAe

[Ano71-171] Anonymous. Classified advertising. IEEE Spectrum, 8(8):103, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:CAf

[Ano71-172] Anonymous. Classified advertising. IEEE Spectrum, 8 (10):92, October 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:CAg

[Ano71-173] Anonymous. Classified advertising. IEEE Spectrum, 8(11): 106, November 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:CAh

[Ano71-174] Anonymous. Classified advertising. IEEE Spectrum, 8 (12):86, December 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Ca

[Ano71-175] Anonymous. Contents. IEEE Spectrum, 8(1):2–3, January 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:EAb

[Ano71-176] Anonymous. Educational aids. IEEE Spectrum, 8(8):89, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:EAc

[Ano71-177] Anonymous. Educational aids. IEEE Spectrum, 8(9):98, September 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:EAd

[Ano71-178] Anonymous. Educational aids. IEEE Spectrum, 8(10):87, October 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:EAe

[Ano71-179] Anonymous. Educational aids. IEEE Spectrum, 8(11):97, November 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:EAf

[Ano71-180] Anonymous. Educational aids.

IEEE Spectrum, 8(12):88, December 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:EAa

[Ano71-181] Anonymous. Exhibitors [advertisement]. IEEE Spectrum, 8 (3):130–145, March 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FPa

[Ano71-182] Anonymous. Focal points. IEEE Spectrum, 8(1):16–18, January 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FPb

[Ano71-183] Anonymous. Focal points. IEEE Spectrum, 8(2):11–14, February 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FPc

[Ano71-184] Anonymous. Focal points. IEEE Spectrum, 8(3):14–18, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FPd

[Ano71-185] Anonymous. Focal points. IEEE Spectrum, 8(4):88–91, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FPe

[Ano71-186] Anonymous. Focal points. IEEE Spectrum, 8(5):105–108, May 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FPf

[Ano71-187] Anonymous. Focal points. IEEE Spectrum, 8(6):84–86, June 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FPg

[Ano71-188] Anonymous. Focal points. IEEE Spectrum, 8(7):86–87, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FPh

[Ano71-189] Anonymous. Focal points. IEEE Spectrum, 8(8):93–94, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FPi

[Ano71-190] Anonymous. Focal points. IEEE Spectrum, 8(9):104–106, September 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FPj

[Ano71-191] Anonymous. Focal points. IEEE Spectrum, 8(10):90–92, October 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FPk

[Ano71-192] Anonymous. Focal points.

IEEE Spectrum, 8(11):100–101, November 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FPl

[Ano71-193] Anonymous. Focal points. IEEE Spectrum, 8(12):91–92, December 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Fa

[Ano71-194] Anonymous. Forum. IEEE Spectrum, 8(1):6–12, January 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Fb

[Ano71-195] Anonymous. Forum. IEEE Spectrum, 8(2):6–8, February 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Fd

[Ano71-196] Anonymous. Forum. IEEE Spectrum, 8(3):8–10, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Fe

[Ano71-197] Anonymous. Forum. IEEE Spectrum, 8(4):11–13, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Ff

[Ano71-198] Anonymous. Forum. IEEE Spectrum, 8(5):12–16, May 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Fg

[Ano71-199] Anonymous. Forum. IEEE Spectrum, 8(6):10, June 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Fh

[Ano71-200] Anonymous. Forum. IEEE Spectrum, 8(7):10–12, July 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Fi

[Ano71-201] Anonymous. Forum. IEEE Spectrum, 8(8):9–12, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Fj

[Ano71-202] Anonymous. Forum. IEEE Spectrum, 8(9):13–17, September 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Fk

[Ano71-203] Anonymous. Forum. IEEE Spectrum, 8(10):13–16, October 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Fl

[Ano71-204] Anonymous. Forum. IEEE Spectrum, 8(11):14–17, November 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Fm

[Ano71-205] Anonymous. Forum. IEEE Spectrum, 8(12):14–16, December 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FCa

[Ano71-206] Anonymous. [front cover]. IEEE Spectrum, 8(1):c1, January 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FCb

[Ano71-207] Anonymous. [front cover].

IEEE Spectrum, 8(2):c1, February 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FCc

[Ano71-208] Anonymous. [front cover].

IEEE Spectrum, 8(2):c2, February 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FCd

[Ano71-209] Anonymous. [front cover]. IEEE Spectrum, 8(3):c1, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FCe

[Ano71-210] Anonymous. [front cover]. IEEE Spectrum, 8(3):c2, March 1971.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FCf

[Ano71-211] Anonymous. [front cover]. IEEE Spectrum, 8(4):c1, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FCg

[Ano71-212] Anonymous. [front cover]. IEEE Spectrum, 8(4):c2, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FCh

[Ano71-213] Anonymous. [front cover]. IEEE Spectrum, 8(5):c1, May 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FCi

[Ano71-214] Anonymous. [front cover]. IEEE Spectrum, 8(5):c2, May 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FCj

[Ano71-215] Anonymous. [front cover]. IEEE Spectrum, 8(6):c1, June 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FCk

[Ano71-216] Anonymous. [front cover]. IEEE Spectrum, 8(6):c2, June 1971.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FCl

[Ano71-217] Anonymous. [front cover]. IEEE Spectrum, 8(7):c1, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FCm

[Ano71-218] Anonymous. [front cover]. IEEE Spectrum, 8(7):c2, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FCn

[Ano71-219] Anonymous. [front cover]. IEEE Spectrum, 8(8):c1, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FCo

[Ano71-220] Anonymous. [front cover]. IEEE Spectrum, 8(8):c2, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FCp

[Ano71-221] Anonymous. [front cover]. IEEE Spectrum, 8(9):c1, September 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FCq

[Ano71-222] Anonymous. [front cover]. IEEE Spectrum, 8(9):c2, September 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FCr

[Ano71-223] Anonymous. [front cover]. IEEE Spectrum, 8(10):c1, October 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FCs

[Ano71-224] Anonymous. [front cover]. IEEE Spectrum, 8(10):c2, October 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FCt

[Ano71-225] Anonymous. [front cover]. IEEE Spectrum, 8(11):c1, November 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FCu

[Ano71-226] Anonymous. [front cover]. IEEE Spectrum, 8(11):c2, November 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FCv

[Ano71-227] Anonymous. [front cover]. IEEE Spectrum, 8(12):c1, December 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FCw

[Ano71-228] Anonymous. [front cover]. IEEE Spectrum, 8(12):c2, December 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FSIa

[Ano71-229] Anonymous. Future special issues. IEEE Spectrum, 8 (1):89, January 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FSIb

[Ano71-230] Anonymous. Future special issues. IEEE Spectrum, 8(2): 110, February 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FSIc

[Ano71-231] Anonymous. Future special issues. IEEE Spectrum, 8(3):104, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FSId

[Ano71-232] Anonymous. Future special issues. IEEE Spectrum, 8(4):86, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FSIe

[Ano71-233] Anonymous. Future special issues. IEEE Spectrum, 8(5):101, May 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FSIf

[Ano71-234] Anonymous. Future special issues. IEEE Spectrum, 8(6):81, June 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FSIg

[Ano71-235] Anonymous. Future special issues. IEEE Spectrum, 8(7):81, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FSIh

[Ano71-236] Anonymous. Future special issues. IEEE Spectrum, 8(8):87, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FSIi

[Ano71-237] Anonymous. Future special issues. IEEE Spectrum, 8(9): 96, September 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FSIj

[Ano71-238] Anonymous. Future special issues. IEEE Spectrum, 8 (10):85, October 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FSIk

[Ano71-239] Anonymous. Future special issues. IEEE Spectrum, 8(11): 92–95, November 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:FSIl

[Ano71-240] Anonymous. Future special issues. IEEE Spectrum, 8(12):

82–86, December 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IES

[Ano71-241] Anonymous. IEEE employment survey. IEEE Spectrum, 8(3): 149–150, March 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IPOa

[Ano71-242] Anonymous. IEEE publications operations. IEEE Spectrum, 8(8):4, August 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IPOb

[Ano71-243] Anonymous. IEEE publications operations. IEEE Spectrum, 8 (9):4, September 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IPOc

[Ano71-244] Anonymous. IEEE publications operations. IEEE Spectrum, 8 (10):4, October 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IPOd

[Ano71-245] Anonymous. IEEE publications operations. IEEE Spectrum, 8 (11):4, November 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IPOe

[Ano71-246] Anonymous. IEEE publications operations. IEEE Spectrum, 8 (12):4, December 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IRP

[Ano71-247] Anonymous. IEEE recommended practice: Rules for the use of units of the International System of Units. IEEE Spectrum, 8(3):77–78, March 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:ISa

[Ano71-248] Anonymous. IEEE Standards. IEEE Spectrum, 8(8):88, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:ISb

[Ano71-249] Anonymous. IEEE Standards. IEEE Spectrum, 8(9): 97, September 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:ISc

[Ano71-250] Anonymous. IEEE Standards. IEEE Spectrum, 8(10):86, October 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:ISd

[Ano71-251] Anonymous. IEEE Standards. IEEE Spectrum, 8 (11):96, November 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:ISe

[Ano71-252] Anonymous. IEEE Standards.

IEEE Spectrum, 8(12):87, December 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IAI

[Ano71-253] Anonymous. Important announcement to IEEE members. IEEE Spectrum, 8(2):116, February 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IAa

[Ano71-254] Anonymous. Index to advertisers. IEEE Spectrum, 8(1): 112–114, January 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IAb

[Ano71-255] Anonymous. Index to advertisers. IEEE Spectrum, 8 (2):128, February 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IAc

[Ano71-256] Anonymous. Index to advertisers. IEEE Spectrum, 8(3): 146–148, March 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IAd

[Ano71-257] Anonymous. Index to advertisers. IEEE Spectrum, 8(4):104, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IAe

[Ano71-258] Anonymous. Index to advertisers. IEEE Spectrum, 8(5):120, May 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IAf

[Ano71-259] Anonymous. Index to advertisers. IEEE Spectrum, 8(6):96, June 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IAg

[Ano71-260] Anonymous. Index to advertisers. IEEE Spectrum, 8(7):100, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IAh

[Ano71-261] Anonymous. Index to advertisers. IEEE Spectrum, 8(8):108, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IAj

[Ano71-262] Anonymous. Index to advertisers. IEEE Spectrum, 8(9): 120, September 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IAk

[Ano71-263] Anonymous. Index to advertisers. IEEE Spectrum, 8 (10):104, October 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IAl

[Ano71-264] Anonymous. Index to advertisers. IEEE Spectrum, 8(11): 112, November 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IAm

[Ano71-265] Anonymous. Index to advertisers. IEEE Spectrum, 8(12): 120, December 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IASa

[Ano71-266] Anonymous. Information about the society. IEEE Spectrum, 8 (2):129–130, February 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IASb

[Ano71-267] Anonymous. Information about the society. IEEE Spectrum, 8(6):97–98, June 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IASc

[Ano71-268] Anonymous. Information about the society. IEEE Spectrum, 8(9): 121–122, September 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IASd

[Ano71-269] Anonymous. Information about the society. IEEE Spectrum, 8(9): 123–124, September 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IASe

[Ano71-270] Anonymous. Information about the society. IEEE Spectrum, 8 (10):105–106, October 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IASf

[Ano71-271] Anonymous. Information about the society. IEEE Spectrum, 8 (10):107–108, October 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IBC

[Ano71-272] Anonymous. [inside back cover]. IEEE Spectrum, 8(1):c3, January 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IFC

[Ano71-273] Anonymous. [inside front cover]. IEEE Spectrum, 8(1):c2, January 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IIa

[Ano71-274] Anonymous. Inside IEEE. IEEE Spectrum, 8(1):95–104, January 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IIb

[Ano71-275] Anonymous. Inside IEEE. IEEE Spectrum, 8(2):113–122, February 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IIc

[Ano71-276] Anonymous. Inside IEEE. IEEE Spectrum, 8(3):112–126, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IId

[Ano71-277] Anonymous. Inside IEEE. IEEE Spectrum, 8(4):7–10, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IIe

[Ano71-278] Anonymous. Inside IEEE. IEEE Spectrum, 8(5):6–9, May 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IIf

[Ano71-279] Anonymous. Inside IEEE. IEEE Spectrum, 8(6):5–9, June 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IIh

[Ano71-280] Anonymous. Inside IEEE. IEEE Spectrum, 8(7):7–9, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IIi

[Ano71-281] Anonymous. Inside IEEE. IEEE Spectrum, 8(8):7–8, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IIj

[Ano71-282] Anonymous. Inside IEEE. IEEE Spectrum, 8(9):7–12, September 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IIk

[Ano71-283] Anonymous. Inside IEEE. IEEE Spectrum, 8(10):7–12, October 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IIl

[Ano71-284] Anonymous. Inside IEEE. IEEE Spectrum, 8(11):7–13, November 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IIm

[Ano71-285] Anonymous. Inside IEEE. IEEE Spectrum, 8(12):7–10, December 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IEEa

[Ano71-286] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 8(2):4, February 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IEEb

[Ano71-287] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 8(3):4, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IEEc

[Ano71-288] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 8(4): 4, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IEEd

[Ano71-289] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 8(5): 4, May 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IEEe

[Ano71-290] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 8(6): 4, June 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IEEf

[Ano71-291] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 8(7): 6, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IEEg

[Ano71-292] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 8(8):6, August 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IEEh

[Ano71-293] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 8 (9):6, September 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IEEi

[Ano71-294] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 8(10):6, October 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IEEj

[Ano71-295] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 8 (11):6, November 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:IEEk

[Ano71-296] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 8 (12):6, December 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:JTPa

[Ano71-297] Anonymous. Japan at a turning point: interviews with the experts (III). IEEE Spectrum, 8(2):124, February 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:JTPb

[Ano71-298] Anonymous. Japan at a turning point: interviews with the experts (IV). IEEE Spectrum, 8(3):22–32, March 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:LSa

[Ano71-299] Anonymous. List of staff. IEEE Spectrum, 8(1):4, January 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:LSb

[Ano71-300] Anonymous. [list of staff]. IEEE Spectrum, 8(7):4, July 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:LNL

[Ano71-301] Anonymous. Locos — the new layer cake mix [advertisement]. IEEE Spectrum, 8(7):nil4, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Ma

[Ano71-302] Anonymous. Meetings. IEEE Spectrum, 8(4):92–98, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Mb

[Ano71-303] Anonymous. Meetings. IEEE Spectrum, 8(5):109–114, May 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Mc

[Ano71-304] Anonymous. Meetings. IEEE Spectrum, 8(6):87–90, June 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Md

[Ano71-305] Anonymous. Meetings. IEEE Spectrum, 8(7):88–94, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Me

[Ano71-306] Anonymous. Meetings. IEEE Spectrum, 8(8):95–101, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Mf

[Ano71-307] Anonymous. Meetings. IEEE Spectrum, 8(9):107–115, September 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Mg

[Ano71-308] Anonymous. Meetings. IEEE Spectrum, 8(10):93–98, October 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Mh

[Ano71-309] Anonymous. Meetings. IEEE Spectrum, 8(11):102–106, November 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Mi

[Ano71-310] Anonymous. Meetings. IEEE Spectrum, 8(12):93–100, December 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:NPAa

[Ano71-311] Anonymous. New product applications. IEEE Spectrum, 8 (1):76–81, January 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:NPAb

[Ano71-312] Anonymous. New product applications. IEEE Spectrum, 8 (2):96–101, February 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:NPAc

[Ano71-313] Anonymous. New product applications. IEEE Spectrum, 8 (3):87–95, March 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:NPAd

[Ano71-314] Anonymous. New product applications. IEEE Spectrum, 8 (4):71–76, April 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:NPAe

[Ano71-315] Anonymous. New product applications. IEEE Spectrum, 8 (5):88–93, May 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:NPAf

[Ano71-316] Anonymous. New product applications. IEEE Spectrum, 8 (6):69–74, June 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:NPAg

[Ano71-317] Anonymous. New product applications. IEEE Spectrum, 8 (7):68–73, July 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:NPAh

[Ano71-318] Anonymous. New product applications. IEEE Spectrum, 8 (8):74–79, August 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:NPAi

[Ano71-319] Anonymous. New product applications. IEEE Spectrum, 8 (9):83–88, September 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:NPAj

[Ano71-320] Anonymous. New product applications. IEEE Spectrum, 8 (10):72–77, October 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:NPAk

[Ano71-321] Anonymous. New product applications. IEEE Spectrum, 8 (11):82–87, November 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:NPAl

[Ano71-322] Anonymous. New product applications. IEEE Spectrum, 8 (12):72–77, December 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:NWa

[Ano71-323] Anonymous. News from Washington. IEEE Spectrum, 8(1): 13–15, January 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:NWb

[Ano71-324] Anonymous. News from Washington. IEEE Spectrum, 8(2): 9–10, February 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:NWc

[Ano71-325] Anonymous. News from Washington. IEEE Spectrum, 8 (3):11–12, March 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:NWd

[Ano71-326] Anonymous. News from Washington. IEEE Spectrum, 8 (4):15–16, April 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:NWe

[Ano71-327] Anonymous. News from Washington. IEEE Spectrum, 8 (5):19–20, May 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:NWf

[Ano71-328] Anonymous. News from Washington. IEEE Spectrum, 8 (6):11–12, June 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:NWg

[Ano71-329] Anonymous. News from Washington. IEEE Spectrum, 8(7):13– 14, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:NWh

[Ano71-330] Anonymous. News from Washington. IEEE Spectrum, 8 (8):13–14, August 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:NWi

[Ano71-331] Anonymous. News from Washington. IEEE Spectrum, 8(9): 19–20, September 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:NWj

[Ano71-332] Anonymous. News from Washington. IEEE Spectrum, 8(10): 17–18, October 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:NWk

[Ano71-333] Anonymous. News from Washington. IEEE Spectrum, 8(11): 19–20, November 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:NWl

[Ano71-334] Anonymous. News from Washington. IEEE Spectrum, 8(12): 17–18, December 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Oa

[Ano71-335] Anonymous. Obituaries. IEEE Spectrum, 8(3):129, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Ob

[Ano71-336] Anonymous. Obituaries. IEEE Spectrum, 8(4):103, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Oc

[Ano71-337] Anonymous. Obituaries. IEEE Spectrum, 8(5):119, May 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Od

[Ano71-338] Anonymous. Obituaries. IEEE Spectrum, 8(7):99, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Oe

[Ano71-339] Anonymous. Obituaries. IEEE Spectrum, 8(8):107, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Of

[Ano71-340] Anonymous. Obituaries. IEEE Spectrum, 8(9):119, September 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Og

[Ano71-341] Anonymous. Obituaries. IEEE Spectrum, 8(10):103, October 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Oh

[Ano71-342] Anonymous. Obituaries. IEEE Spectrum, 8(11):111, November 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Oi

[Ano71-343] Anonymous. Obituaries. IEEE Spectrum, 8(12):105, December 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:OMV

[Ano71-344] Anonymous. One man’s view of marketing. IEEE Spectrum, 8(9):82, September 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Pa

[Ano71-345] Anonymous. People. IEEE Spectrum, 8(1):105–111, January 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Pb

[Ano71-346] Anonymous. People. IEEE Spectrum, 8(2):123, February 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Pc

[Ano71-347] Anonymous. People. IEEE Spectrum, 8(3):128–129, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Pd

[Ano71-348] Anonymous. People. IEEE Spectrum, 8(4):102–103, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Pe

[Ano71-349] Anonymous. People. IEEE Spectrum, 8(5):118–119, May 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Pf

[Ano71-350] Anonymous. People. IEEE Spectrum, 8(6):94–95, June 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Pg

[Ano71-351] Anonymous. People. IEEE Spectrum, 8(7):97–99, July 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Ph

[Ano71-352] Anonymous. People. IEEE Spectrum, 8(8):104–107, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Pi

[Ano71-353] Anonymous. People. IEEE Spectrum, 8(9):118–119, September 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Pj

[Ano71-354] Anonymous. People. IEEE Spectrum, 8(10):102, October 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Pk

[Ano71-355] Anonymous. People. IEEE Spectrum, 8(11):109–111, November 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:Pl

[Ano71-356] Anonymous. People. IEEE Spectrum, 8(12):104–105, December 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:PCSa

[Ano71-357] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 8(2):123–127, February 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:PCSb

[Ano71-358] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 8(3):147, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:PCSc

[Ano71-359] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 8(4):101, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:PCSd

[Ano71-360] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 8(5):117, May 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:PCSe

[Ano71-361] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 8(6):91, June 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:PCSf

[Ano71-362] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 8(7):96, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:PCSg

[Ano71-363] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 8(8):105, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:PCSh

[Ano71-364] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 8(9):117, September 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:PCSi

[Ano71-365] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 8(10):100– 101, October 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:PCSj

[Ano71-366] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 8(11):108, November 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:PCSk

[Ano71-367] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 8(12):102–103, December 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SIa

[Ano71-368] Anonymous. Scanning the issues.

IEEE Spectrum, 8(1):82–83, January 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SIb

[Ano71-369] Anonymous. Scanning the issues. IEEE Spectrum, 8(2): 102–103, February 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SIc

[Ano71-370] Anonymous. Scanning the issues. IEEE Spectrum, 8(3): 96–97, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SId

[Ano71-371] Anonymous. Scanning the issues. IEEE Spectrum, 8(4): 77–78, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SIe

[Ano71-372] Anonymous. Scanning the issues.

IEEE Spectrum, 8(5):94–95, May 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SIf

[Ano71-373] Anonymous. Scanning the issues. IEEE Spectrum, 8(6): 75–76, June 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SIg

[Ano71-374] Anonymous. Scanning the issues. IEEE Spectrum, 8(7):74– 75, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SIh

[Ano71-375] Anonymous. Scanning the issues. IEEE Spectrum, 8(8):81–83, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SIi

[Ano71-376] Anonymous. Scanning the issues. IEEE Spectrum, 8(9): 90–91, September 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SIj

[Ano71-377] Anonymous. Scanning the issues. IEEE Spectrum, 8(10): 79–80, October 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SIk

[Ano71-378] Anonymous. Scanning the issues. IEEE Spectrum, 8(11): 89–90, November 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SIl

[Ano71-379] Anonymous. Scanning the issues. IEEE Spectrum, 8(12): 79–80, December 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SSAa

[Ano71-380] Anonymous. Science/scope [advertisement]. IEEE Spectrum, 8 (8):nil1–1, August 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SSAb

[Ano71-381] Anonymous. Science/scope [advertisement]. IEEE Spectrum, 8 (10):nil1–1, October 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SSAc

[Ano71-382] Anonymous. Science/scope [advertisement]. IEEE Spectrum, 8 (12):nil4, December 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SPa

[Ano71-383] Anonymous. Special publications. IEEE Spectrum, 8 (1):88, January 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SPb

[Ano71-384] Anonymous. Special publications. IEEE Spectrum, 8(2): 108–109, February 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SPc

[Ano71-385] Anonymous. Special publications. IEEE Spectrum, 8(3): 102–103, March 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SPd

[Ano71-386] Anonymous. Special publications. IEEE Spectrum, 8 (4):84–85, April 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SPe

[Ano71-387] Anonymous. Special publications. IEEE Spectrum, 8 (5):99–100, May 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SPf

[Ano71-388] Anonymous. Special publications. IEEE Spectrum, 8(6):80, June 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SPg

[Ano71-389] Anonymous. Special publications. IEEE Spectrum, 8(7):80, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SPh

[Ano71-390] Anonymous. Special publications. IEEE Spectrum, 8(8): 88–89, August 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SPi

[Ano71-391] Anonymous. Special publications. IEEE Spectrum, 8(9): 97–98, September 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SPj

[Ano71-392] Anonymous. Special publications. IEEE Spectrum, 8(10): 86–87, October 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SPk

[Ano71-393] Anonymous. Special publications. IEEE Spectrum, 8(11): 96–97, November 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SPl

[Ano71-394] Anonymous. Special publications. IEEE Spectrum, 8(12): 87–88, December 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SLa

[Ano71-395] Anonymous. Spectral lines.

IEEE Spectrum, 8(2):41, February 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SLc

[Ano71-396] Anonymous. Spectral lines.

IEEE Spectrum, 8(3):37, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SLd

[Ano71-397] Anonymous. Spectral lines.

IEEE Spectrum, 8(4):17, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SLe

[Ano71-398] Anonymous. Spectral lines. IEEE Spectrum, 8(5):31, May 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SLf

[Ano71-399] Anonymous. Spectral lines. IEEE Spectrum, 8(6):13, June 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SLg

[Ano71-400] Anonymous. Spectral lines. IEEE Spectrum, 8(7):17, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SLh

[Ano71-401] Anonymous. Spectral lines. IEEE Spectrum, 8(8):21, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SLi

[Ano71-402] Anonymous. Spectral lines. IEEE Spectrum, 8(9):29, September 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SLj

[Ano71-403] Anonymous. Spectral lines. IEEE Spectrum, 8(10):19, October 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SLk

[Ano71-404] Anonymous. Spectral lines.

IEEE Spectrum, 8(11):23, November 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SLl

[Ano71-405] Anonymous. Spectral lines.

IEEE Spectrum, 8(12):19, December 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SLB

[Ano71-406] Anonymous. Spectral lines:

Birth of a concept. IEEE Spectrum, 8(1):21–22, January 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:SICa

[Ano71-407] Anonymous. Statements from IEEE candidates. IEEE Spectrum, 8(9):52–58, September 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:TCa

[Ano71-408] Anonymous. Table of contents.

IEEE Spectrum, 8(2):2–3, February 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:TCb

[Ano71-409] Anonymous. Table of contents. IEEE Spectrum, 8(3):2–3, March 1971. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:TCc

[Ano71-410] Anonymous. Table of contents. IEEE Spectrum, 8(4):2–3, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:TCd

[Ano71-411] Anonymous. Table of contents. IEEE Spectrum, 8(5):2– 3, May 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:TCe

[Ano71-412] Anonymous. Table of contents. IEEE Spectrum, 8(6):2–3, June 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:TCf

[Ano71-413] Anonymous. Table of contents. IEEE Spectrum, 8(7):2– 3, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:TCg

[Ano71-414] Anonymous. Table of contents. IEEE Spectrum, 8(8):2–3, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:TCh

[Ano71-415] Anonymous. Table of contents. IEEE Spectrum, 8(9):

2–3, September 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:TCi

[Ano71-416] Anonymous. Table of contents.

IEEE Spectrum, 8(10):2–3, October 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:TCj

[Ano71-417] Anonymous. Table of contents. IEEE Spectrum, 8(11): 2–3, November 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:TCk

[Ano71-418] Anonymous. Table of contents.

IEEE Spectrum, 8(12):2–3, December 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:TP

[Ano71-419] Anonymous. Technical program. IEEE Spectrum, 8(2): 17–39, February 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:TSA

[Ano71-420] Anonymous. The time savers [advertisement]. IEEE Spectrum, 8(1):1, January 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1971:WWY

[Ano71-421] Anonymous. Where would you like to have your heart attack? IEEE Spectrum, 8 (10):44–50, October 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:WPP

[Ano72a] Anonymous. 1972 WESCON Professional Program. IEEE Spectrum, 9(9):w39–w52, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:API

[Ano72b] Anonymous. Advance program for the IEEE Northeast Electronics Research Engineering Meeting. IEEE Spectrum, 9(10): 27–30, October 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ATCa

[Ano72c] Anonymous. Advance tables of contents. IEEE Spectrum, 9 (1):99–100, January 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ATCb

[Ano72d] Anonymous. Advance tables of contents. IEEE Spectrum, 9 (2):81–82, February 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ATCc

[Ano72e] Anonymous. Advance tables of contents. IEEE Spectrum, 9 (3):101–102, March 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ATCd

[Ano72f] Anonymous. Advance tables of contents. IEEE Spectrum, 9(4):82–83, April 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ATCe

[Ano72g] Anonymous. Advance tables of contents. IEEE Spectrum, 9(5):77–78, May 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ATCf

[Ano72h] Anonymous. Advance tables of contents. IEEE Spectrum, 9(6):80–81, June 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ATCg

[Ano72i] Anonymous. Advance tables of contents. IEEE Spectrum, 9(7):76–77, July 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ATCh

[Ano72j] Anonymous. Advance tables of contents. IEEE Spectrum, 9(8):107, August 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ATCi

[Ano72k] Anonymous. Advance tables of contents. IEEE Spectrum, 9(9): 90–91, September 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ATCj

[Ano72l] Anonymous. Advance tables of contents. IEEE Spectrum, 9(10): 104–105, October 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ATCk

[Ano72m] Anonymous. Advance tables of contents. IEEE Spectrum, 9(11): 86–87, November 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ATCl

[Ano72n] Anonymous. Advance tables of contents. IEEE Spectrum, 9(12): 88–89, December 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aa

[Ano72o] Anonymous. [advertisement]. IEEE Spectrum, 9(1):nil1–1, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ab

[Ano72p] Anonymous. [advertisement]. IEEE Spectrum, 9(1):4, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ac

[Ano72q] Anonymous. [advertisement]. IEEE Spectrum, 9(1):5, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ad

[Ano72r] Anonymous. [advertisement].

IEEE Spectrum, 9(1):6, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ae

[Ano72s] Anonymous. [advertisement].

IEEE Spectrum, 9(1):13–15, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Af

[Ano72t] Anonymous. [advertisement].

IEEE Spectrum, 9(1):20, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ag

[Ano72u] Anonymous. [advertisement].

IEEE Spectrum, 9(1):38, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ah

[Ano72v] Anonymous. [advertisement].

IEEE Spectrum, 9(1):59–62, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ai

[Ano72w] Anonymous. [advertisement].

IEEE Spectrum, 9(1):96, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aj

[Ano72x] Anonymous. [advertisement]. IEEE Spectrum, 9(1):107, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ak

[Ano72y] Anonymous. [advertisement]. IEEE Spectrum, 9(1):120, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Am

[Ano72z] Anonymous. [advertisement]. IEEE Spectrum, 9(1):123, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:An

[Ano72-27] Anonymous. [advertisement]. IEEE Spectrum, 9(1):nil2, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ao

[Ano72-28] Anonymous. [advertisement]. IEEE Spectrum, 9(2):nil1–1, February 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ap

[Ano72-29] Anonymous. [advertisement]. IEEE Spectrum, 9(2):4, February 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aq

[Ano72-30] Anonymous. [advertisement].

IEEE Spectrum, 9(2):5, February 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ar

[Ano72-31] Anonymous. [advertisement].

IEEE Spectrum, 9(2):6, February 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:As

[Ano72-32] Anonymous. [advertisement].

IEEE Spectrum, 9(2):16, February 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:At

[Ano72-33] Anonymous. [advertisement].

IEEE Spectrum, 9(2):51–54, February 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Au

[Ano72-34] Anonymous. [advertisement].

IEEE Spectrum, 9(2):86, February 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Av

[Ano72-35] Anonymous. [advertisement].

IEEE Spectrum, 9(2):91–93, February 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aw

[Ano72-36] Anonymous. [advertisement]. IEEE Spectrum, 9(2):107–108, February 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ax

[Ano72-37] Anonymous. [advertisement]. IEEE Spectrum, 9(3):nil1–1, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ay

[Ano72-38] Anonymous. [advertisement]. IEEE Spectrum, 9(3):4, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Az

[Ano72-39] Anonymous. [advertisement]. IEEE Spectrum, 9(3):5, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aba

[Ano72-40] Anonymous. [advertisement]. IEEE Spectrum, 9(3):6, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Abb

[Ano72-41] Anonymous. [advertisement]. IEEE Spectrum, 9(3):11–20, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Abc

[Ano72-42] Anonymous. [advertisement].

IEEE Spectrum, 9(3):23–s2, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Abd

[Ano72-43] Anonymous. [advertisement].

IEEE Spectrum, 9(3):s5–s7, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Abe

[Ano72-44] Anonymous. [advertisement].

IEEE Spectrum, 9(3):s17–s28, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Abf

[Ano72-45] Anonymous. [advertisement].

IEEE Spectrum, 9(3):s31–s32, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Abg

[Ano72-46] Anonymous. [advertisement].

IEEE Spectrum, 9(3):67–70, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Abh

[Ano72-47] Anonymous. [advertisement].

IEEE Spectrum, 9(3):105, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Abi

[Ano72-48] Anonymous. [advertisement]. IEEE Spectrum, 9(3):111–112, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Abj

[Ano72-49] Anonymous. [advertisement]. IEEE Spectrum, 9(3):115, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Abk

[Ano72-50] Anonymous. [advertisement]. IEEE Spectrum, 9(3):117–121, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Abl

[Ano72-51] Anonymous. [advertisement]. IEEE Spectrum, 9(3):126, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Abm

[Ano72-52] Anonymous. [advertisement]. IEEE Spectrum, 9(3):131–132, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Abn

[Ano72-53] Anonymous. [advertisement]. IEEE Spectrum, 9(3):nil2, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Abo

[Ano72-54] Anonymous. [advertisement].

IEEE Spectrum, 9(4):nil1, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Abp

[Ano72-55] Anonymous. [advertisement].

IEEE Spectrum, 9(4):1, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Abq

[Ano72-56] Anonymous. [advertisement].

IEEE Spectrum, 9(4):4, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Abr

[Ano72-57] Anonymous. [advertisement].

IEEE Spectrum, 9(4):5, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Abs

[Ano72-58] Anonymous. [advertisement].

IEEE Spectrum, 9(4):6, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Abt

[Ano72-59] Anonymous. [advertisement].

IEEE Spectrum, 9(4):13–15, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Abu

[Ano72-60] Anonymous. [advertisement]. IEEE Spectrum, 9(4):51–54, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Abv

[Ano72-61] Anonymous. [advertisement]. IEEE Spectrum, 9(4):85, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Abw

[Ano72-62] Anonymous. [advertisement]. IEEE Spectrum, 9(4):89, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Abx

[Ano72-63] Anonymous. [advertisement]. IEEE Spectrum, 9(4):91–94, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aby

[Ano72-64] Anonymous. [advertisement]. IEEE Spectrum, 9(4):105–106, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Abz

[Ano72-65] Anonymous. [advertisement]. IEEE Spectrum, 9(4):nil4, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aca

[Ano72-66] Anonymous. [advertisement].

IEEE Spectrum, 9(5):nil1–1, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Acb

[Ano72-67] Anonymous. [advertisement].

IEEE Spectrum, 9(5):4, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Acc

[Ano72-68] Anonymous. [advertisement].

IEEE Spectrum, 9(5):5, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Acd

[Ano72-69] Anonymous. [advertisement].

IEEE Spectrum, 9(5):6, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ace

[Ano72-70] Anonymous. [advertisement].

IEEE Spectrum, 9(5):10, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Acf

[Ano72-71] Anonymous. [advertisement].

IEEE Spectrum, 9(5):47–50, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Acg

[Ano72-72] Anonymous. [advertisement]. IEEE Spectrum, 9(5):89, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ach

[Ano72-73] Anonymous. [advertisement]. IEEE Spectrum, 9(5):97–98, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aci

[Ano72-74] Anonymous. [advertisement]. IEEE Spectrum, 9(5):99–100, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Acj

[Ano72-75] Anonymous. [advertisement]. IEEE Spectrum, 9(5):nil2, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ack

[Ano72-76] Anonymous. [advertisement]. IEEE Spectrum, 9(6):nil1–1, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Acl

[Ano72-77] Anonymous. [advertisement]. IEEE Spectrum, 9(6):4, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Acm

[Ano72-78] Anonymous. [advertisement].

IEEE Spectrum, 9(6):5, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Acn

[Ano72-79] Anonymous. [advertisement].

IEEE Spectrum, 9(6):6, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aco

[Ano72-80] Anonymous. [advertisement].

IEEE Spectrum, 9(6):11, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Acp

[Ano72-81] Anonymous. [advertisement].

IEEE Spectrum, 9(6):30, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Acq

[Ano72-82] Anonymous. [advertisement].

IEEE Spectrum, 9(6):49–52, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Acr

[Ano72-83] Anonymous. [advertisement].

IEEE Spectrum, 9(6):77, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Acs

[Ano72-84] Anonymous. [advertisement]. IEEE Spectrum, 9(6):87, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Act

[Ano72-85] Anonymous. [advertisement]. IEEE Spectrum, 9(6):101–102, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Acu

[Ano72-86] Anonymous. [advertisement]. IEEE Spectrum, 9(6):nil4, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Acv

[Ano72-87] Anonymous. [advertisement]. IEEE Spectrum, 9(7):nil1–1, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Acw

[Ano72-88] Anonymous. [advertisement]. IEEE Spectrum, 9(7):4, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Acx

[Ano72-89] Anonymous. [advertisement]. IEEE Spectrum, 9(7):5, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Acy

[Ano72-90] Anonymous. [advertisement].

IEEE Spectrum, 9(7):6, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Acz

[Ano72-91] Anonymous. [advertisement].

IEEE Spectrum, 9(7):9–11, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ada

[Ano72-92] Anonymous. [advertisement].

IEEE Spectrum, 9(7):13–nil5, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Adb

[Ano72-93] Anonymous. [advertisement].

IEEE Spectrum, 9(7):48–49, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Adc

[Ano72-94] Anonymous. [advertisement].

IEEE Spectrum, 9(7):79, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Add

[Ano72-95] Anonymous. [advertisement].

IEEE Spectrum, 9(7):89, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ade

[Ano72-96] Anonymous. [advertisement]. IEEE Spectrum, 9(7):97–98, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Adf

[Ano72-97] Anonymous. [advertisement]. IEEE Spectrum, 9(7):nil6, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Adg

[Ano72-98] Anonymous. [advertisement]. IEEE Spectrum, 9(8):nil1–1, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Adh

[Ano72-99] Anonymous. [advertisement]. IEEE Spectrum, 9(8):4, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Adi

[Ano72-100] Anonymous. [advertisement]. IEEE Spectrum, 9(8):6, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Adj

[Ano72-101] Anonymous. [advertisement]. IEEE Spectrum, 9(8):11–13, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Adk

[Ano72-102] Anonymous. [advertisement].

IEEE Spectrum, 9(8):19–22, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Adl

[Ano72-103] Anonymous. [advertisement].

IEEE Spectrum, 9(8):25–32, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Adm

[Ano72-104] Anonymous. [advertisement].

IEEE Spectrum, 9(8):67–68, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Adn

[Ano72-105] Anonymous. [advertisement].

IEEE Spectrum, 9(8):103–104, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ado

[Ano72-106] Anonymous. [advertisement].

IEEE Spectrum, 9(8):123, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Adp

[Ano72-107] Anonymous. [advertisement].

IEEE Spectrum, 9(8):nil2–nil3, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Adq

[Ano72-108] Anonymous. [advertisement]. IEEE Spectrum, 9(8):nil4, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Adr

[Ano72-109] Anonymous. [advertisement]. IEEE Spectrum, 9(9):nil1, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ads

[Ano72-110] Anonymous. [advertisement]. IEEE Spectrum, 9(9):4, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Adt

[Ano72-111] Anonymous. [advertisement]. IEEE Spectrum, 9(9):5, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Adu

[Ano72-112] Anonymous. [advertisement]. IEEE Spectrum, 9(9):6, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Adv

[Ano72-113] Anonymous. [advertisement]. IEEE Spectrum, 9(9):10–13, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Adw

[Ano72-114] Anonymous. [advertisement].

IEEE Spectrum, 9(9):15–16, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Adx

[Ano72-115] Anonymous. [advertisement].

IEEE Spectrum, 9(9):19–20, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ady

[Ano72-116] Anonymous. [advertisement].

IEEE Spectrum, 9(9):57, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Adz

[Ano72-117] Anonymous. [advertisement].

IEEE Spectrum, 9(9):70, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aea

[Ano72-118] Anonymous. [advertisement].

IEEE Spectrum, 9(9):w1–w2, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aeb

[Ano72-119] Anonymous. [advertisement].

IEEE Spectrum, 9(9):w10–w11, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aec

[Ano72-120] Anonymous. [advertisement]. IEEE Spectrum, 9(9):w17–w38, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aed

[Ano72-121] Anonymous. [advertisement]. IEEE Spectrum, 9(9):w42–w53, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aee

[Ano72-122] Anonymous. [advertisement]. IEEE Spectrum, 9(9):w55–w79, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aef

[Ano72-123] Anonymous. [advertisement]. IEEE Spectrum, 9(9):w80–w88, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aeg

[Ano72-124] Anonymous. [advertisement]. IEEE Spectrum, 9(9):93–94, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aeh

[Ano72-125] Anonymous. [advertisement]. IEEE Spectrum, 9(9):97–99, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aei

[Ano72-126] Anonymous. [advertisement].

IEEE Spectrum, 9(9):101, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aej

[Ano72-127] Anonymous. [advertisement].

IEEE Spectrum, 9(9):103, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aek

[Ano72-128] Anonymous. [advertisement].

IEEE Spectrum, 9(9):107, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ael

[Ano72-129] Anonymous. [advertisement].

IEEE Spectrum, 9(9):117–118, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aem

[Ano72-130] Anonymous. [advertisement].

IEEE Spectrum, 9(9):nil2, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aen

[Ano72-131] Anonymous. [advertisement].

IEEE Spectrum, 9(10):nil1– 1, October 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aeo

[Ano72-132] Anonymous. [advertisement]. IEEE Spectrum, 9(10):4, October 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aep

[Ano72-133] Anonymous. [advertisement]. IEEE Spectrum, 9(10):5, October 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aeq

[Ano72-134] Anonymous. [advertisement]. IEEE Spectrum, 9(10):6, October 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aer

[Ano72-135] Anonymous. [advertisement]. IEEE Spectrum, 9(10):9–11, October 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aes

[Ano72-136] Anonymous. [advertisement]. IEEE Spectrum, 9(10):13–24, October 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aet

[Ano72-137] Anonymous. [advertisement]. IEEE Spectrum, 9(10):nil2– nil3, October 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aeu

[Ano72-138] Anonymous. [advertisement].

IEEE Spectrum, 9(10):56, October 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aev

[Ano72-139] Anonymous. [advertisement].

IEEE Spectrum, 9(10):73, October 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aew

[Ano72-140] Anonymous. [advertisement].

IEEE Spectrum, 9(10):116, October 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aex

[Ano72-141] Anonymous. [advertisement].

IEEE Spectrum, 9(10):123– 124, October 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aey

[Ano72-142] Anonymous. [advertisement].

IEEE Spectrum, 9(11):nil1–1, November 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aez

[Ano72-143] Anonymous. [advertisement].

IEEE Spectrum, 9(11):4, November 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Afa

[Ano72-144] Anonymous. [advertisement]. IEEE Spectrum, 9(11):5, November 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Afb

[Ano72-145] Anonymous. [advertisement]. IEEE Spectrum, 9(11):6, November 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Afc

[Ano72-146] Anonymous. [advertisement]. IEEE Spectrum, 9(11):14, November 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Afd

[Ano72-147] Anonymous. [advertisement]. IEEE Spectrum, 9(11):55–57, November 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Afe

[Ano72-148] Anonymous. [advertisement]. IEEE Spectrum, 9(11):69, November 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aff

[Ano72-149] Anonymous. [advertisement]. IEEE Spectrum, 9(11):95–97, November 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Afg

[Ano72-150] Anonymous. [advertisement].

IEEE Spectrum, 9(11):113–114, November 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Afh

[Ano72-151] Anonymous. [advertisement].

IEEE Spectrum, 9(11):nil4, November 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Afi

[Ano72-152] Anonymous. [advertisement].

IEEE Spectrum, 9(12):nil1–1, December 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Afj

[Ano72-153] Anonymous. [advertisement].

IEEE Spectrum, 9(12):4, December 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Afk

[Ano72-154] Anonymous. [advertisement].

IEEE Spectrum, 9(12):6, December 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Afl

[Ano72-155] Anonymous. [advertisement].

IEEE Spectrum, 9(12):13–15, December 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Afm

[Ano72-156] Anonymous. [advertisement]. IEEE Spectrum, 9(12):18, December 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Afn

[Ano72-157] Anonymous. [advertisement]. IEEE Spectrum, 9(12):21–22, December 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Afo

[Ano72-158] Anonymous. [advertisement]. IEEE Spectrum, 9(12):64–65, December 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Afp

[Ano72-159] Anonymous. [advertisement]. IEEE Spectrum, 9(12):97–101, December 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Afq

[Ano72-160] Anonymous. [advertisement]. IEEE Spectrum, 9(12):111, December 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Afr

[Ano72-161] Anonymous. [advertisement]. IEEE Spectrum, 9(12):127, December 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Afs

[Ano72-162] Anonymous. [advertisement].

IEEE Spectrum, 9(12):129–130, December 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Aft

[Ano72-163] Anonymous. [advertisement].

IEEE Spectrum, 9(12):nil4, December 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:AII

[Ano72-164] Anonymous. Annual index — IEEE Spectrum. IEEE Spectrum, 9(12):114–126, December 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ALa

[Ano72-165] Anonymous. Applications literature. IEEE Spectrum, 9 (1):93, January 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ALb

[Ano72-166] Anonymous. Applications literature. IEEE Spectrum, 9 (2):78, February 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ALc

[Ano72-167] Anonymous. Applications literature. IEEE Spectrum, 9(3):98, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ALd

[Ano72-168] Anonymous. Applications literature. IEEE Spectrum, 9(4):79, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ALe

[Ano72-169] Anonymous. Applications literature. IEEE Spectrum, 9(5):74, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ALf

[Ano72-170] Anonymous. Applications literature. IEEE Spectrum, 9(6):76, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ALg

[Ano72-171] Anonymous. Applications literature. IEEE Spectrum, 9(7):73, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ALh

[Ano72-172] Anonymous. Applications literature. IEEE Spectrum, 9(8):102, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ALi

[Ano72-173] Anonymous. Applications literature. IEEE Spectrum, 9 (9):87, September 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ALj

[Ano72-174] Anonymous. Applications literature. IEEE Spectrum, 9 (10):101, October 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ALk

[Ano72-175] Anonymous. Applications literature. IEEE Spectrum, 9 (11):83, November 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ALl

[Ano72-176] Anonymous. Applications literature. IEEE Spectrum, 9 (12):85, December 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:BSP

[Ano72-177] Anonymous. BART sets the pace. IEEE Spectrum, 9(11): 45–54, November 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:BRa

[Ano72-178] Anonymous. Book reviews.

IEEE Spectrum, 9(1):106–107, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:BRb

[Ano72-179] Anonymous. Book reviews.

IEEE Spectrum, 9(2):89, February 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:BRc

[Ano72-180] Anonymous. Book reviews. IEEE Spectrum, 9(3):108–109, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:BRd

[Ano72-181] Anonymous. Book reviews. IEEE Spectrum, 9(4):88, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:BRe

[Ano72-182] Anonymous. Book reviews. IEEE Spectrum, 9(5):85, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:BRf

[Ano72-183] Anonymous. Book reviews. IEEE Spectrum, 9(6):87, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:BRg

[Ano72-184] Anonymous. Book reviews. IEEE Spectrum, 9(7):83–84, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:BRh

[Ano72-185] Anonymous. Book reviews. IEEE Spectrum, 9(8):115, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:BRi

[Ano72-186] Anonymous. Book reviews.

IEEE Spectrum, 9(9):100, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:BRj

[Ano72-187] Anonymous. Book reviews.

IEEE Spectrum, 9(10):110, October 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:BRk

[Ano72-188] Anonymous. Book reviews.

IEEE Spectrum, 9(11):92, November 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:BRl

[Ano72-189] Anonymous. Book reviews.

IEEE Spectrum, 9(12):95, December 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ISRa

[Ano72-190] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 9(1):123–124, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ISRb

[Ano72-191] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 9(2):107–108, February 1972. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ISRc

[Ano72-192] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 9(3):131–132, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ISRd

[Ano72-193] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 9(4):105–106, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ISRe

[Ano72-194] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 9(5):99–100, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ISRf

[Ano72-195] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 9(6):101–102, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ISRg

[Ano72-196] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 9(7):97–98, July 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ISRh

[Ano72-197] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 9(8):nil2–nil3, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ISRi

[Ano72-198] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 9(9):117–118, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ISRj

[Ano72-199] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 9(10):123–124, October 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ISRk

[Ano72-200] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 9(11):113–114, November 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ISRl

[Ano72-201] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 9(12):129–130, December 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:BPa

[Ano72-202] Anonymous. [breaker page]. IEEE Spectrum, 9(1):4, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:BPb

[Ano72-203] Anonymous. [breaker page]. IEEE Spectrum, 9(2):4, February 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:BPc

[Ano72-204] Anonymous. [breaker page]. IEEE Spectrum, 9(3):4, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:BPd

[Ano72-205] Anonymous. [breaker page]. IEEE Spectrum, 9(4):nil2, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:BPe

[Ano72-206] Anonymous. [breaker page]. IEEE Spectrum, 9(4):4, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:BPf

[Ano72-207] Anonymous. [breaker page]. IEEE Spectrum, 9(5):4, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:BPg

[Ano72-208] Anonymous. [breaker page].

IEEE Spectrum, 9(6):4, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:BPh

[Ano72-209] Anonymous. [breaker page].

IEEE Spectrum, 9(7):4, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:BPi

[Ano72-210] Anonymous. [breaker page].

IEEE Spectrum, 9(8):4, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:BPj

[Ano72-211] Anonymous. [breaker page].

IEEE Spectrum, 9(9):4, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:BPk

[Ano72-212] Anonymous. [breaker page].

IEEE Spectrum, 9(10):4, October 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:BPl

[Ano72-213] Anonymous. [breaker page].

IEEE Spectrum, 9(11):4, November 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:BPm

[Ano72-214] Anonymous. [breaker page]. IEEE Spectrum, 9(12):4, December 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Cb

[Ano72-215] Anonymous. Calendar. IEEE Spectrum, 9(1):116–117, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Cc

[Ano72-216] Anonymous. Calendar. IEEE Spectrum, 9(2):98–99, February 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Cd

[Ano72-217] Anonymous. Calendar. IEEE Spectrum, 9(3):122–123, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ce

[Ano72-218] Anonymous. Calendar. IEEE Spectrum, 9(4):100–101, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Cf

[Ano72-219] Anonymous. Calendar. IEEE Spectrum, 9(5):92–93, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Cg

[Ano72-220] Anonymous. Calendar. IEEE Spectrum, 9(6):96–97, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ch

[Ano72-221] Anonymous. Calendar. IEEE Spectrum, 9(7):91–92, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ci

[Ano72-222] Anonymous. Calendar. IEEE Spectrum, 9(8):126–127, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Cj

[Ano72-223] Anonymous. Calendar. IEEE Spectrum, 9(9):112–113, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ck

[Ano72-224] Anonymous. Calendar. IEEE Spectrum, 9(10):117–118, October 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Cl

[Ano72-225] Anonymous. Calendar. IEEE Spectrum, 9(11):108–109, November 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Cm

[Ano72-226] Anonymous. Calendar. IEEE Spectrum, 9(12):108–109, December 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:CP

[Ano72-227] Anonymous. Call for papers. IEEE Spectrum, 9(9):1, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:CAa

[Ano72-228] Anonymous. Classified advertising. IEEE Spectrum, 9(1): 109–111, January 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:CAb

[Ano72-229] Anonymous. Classified advertising. IEEE Spectrum, 9 (2):97, February 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:CAc

[Ano72-230] Anonymous. Classified advertising. IEEE Spectrum, 9(3):123, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:CAd

[Ano72-231] Anonymous. Classified advertising. IEEE Spectrum, 9(4):99, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:CAe

[Ano72-232] Anonymous. Classified advertising. IEEE Spectrum, 9(5):82, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:CAf

[Ano72-233] Anonymous. Classified advertising. IEEE Spectrum, 9(6):84, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:CAg

[Ano72-234] Anonymous. Classified advertising. IEEE Spectrum, 9(7):84, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:CAh

[Ano72-235] Anonymous. Classified advertising. IEEE Spectrum, 9(8):119, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:CAi

[Ano72-236] Anonymous. Classified advertising. IEEE Spectrum, 9(9): 109–111, September 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:CAj

[Ano72-237] Anonymous. Classified advertising. IEEE Spectrum, 9 (10):118, October 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:CAk

[Ano72-238] Anonymous. Classified advertising. IEEE Spectrum, 9(11): 99–107, November 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:CAl

[Ano72-239] Anonymous. Classified advertising. IEEE Spectrum, 9(12): 105–107, December 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:EAa

[Ano72-240] Anonymous. Educational aids. IEEE Spectrum, 9(1):105, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:EAb

[Ano72-241] Anonymous. Educational aids. IEEE Spectrum, 9(2):88, February 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:EAd

[Ano72-242] Anonymous. Educational aids. IEEE Spectrum, 9(3):107, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:EAe

[Ano72-243] Anonymous. Educational aids. IEEE Spectrum, 9(4):87, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:EAf

[Ano72-244] Anonymous. Educational aids.

IEEE Spectrum, 9(5):84, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:EAg

[Ano72-245] Anonymous. Educational aids.

IEEE Spectrum, 9(6):86, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:EAh

[Ano72-246] Anonymous. Educational aids.

IEEE Spectrum, 9(7):82, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:EAi

[Ano72-247] Anonymous. Educational aids.

IEEE Spectrum, 9(8):114, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:EAj

[Ano72-248] Anonymous. Educational aids.

IEEE Spectrum, 9(9):98, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:EAk

[Ano72-249] Anonymous. Educational aids.

IEEE Spectrum, 9(10):109, October 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:EAl

[Ano72-250] Anonymous. Educational aids. IEEE Spectrum, 9(11):91, November 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:EAm

[Ano72-251] Anonymous. Educational aids. IEEE Spectrum, 9(12):94, December 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:EFB

[Ano72-252] Anonymous. EEs’ fringe benefits: a mixed bag. IEEE Spectrum, 9 (9):67–69, September 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:EKM

[Ano72-253] Anonymous. Electronica 72 and Kongress Mikroelektronik. IEEE Spectrum, 9(10):49, October 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:EAc

[Ano72-254] Anonymous. Exhibitors [advertisement]. IEEE Spectrum, 9 (3):s9–s12, March 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FPN

[Ano72-255] Anonymous. Floor plans of the New York Coliseum. IEEE Spectrum, 9(3):s13–s28, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FPa

[Ano72-256] Anonymous. Focal points. IEEE Spectrum, 9(1):108–111, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FPb

[Ano72-257] Anonymous. Focal points. IEEE Spectrum, 9(2):90–93, February 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FPc

[Ano72-258] Anonymous. Focal points. IEEE Spectrum, 9(3):110–112, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FPd

[Ano72-259] Anonymous. Focal points. IEEE Spectrum, 9(4):90–94, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FPe

[Ano72-260] Anonymous. Focal points. IEEE Spectrum, 9(5):86–87, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FPf

[Ano72-261] Anonymous. Focal points. IEEE Spectrum, 9(6):88–90, June 1972.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FPg

[Ano72-262] Anonymous. Focal points. IEEE Spectrum, 9(7):85–86, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FPh

[Ano72-263] Anonymous. Focal points. IEEE Spectrum, 9(8):116–119, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FPi

[Ano72-264] Anonymous. Focal points. IEEE Spectrum, 9(9):102–104, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FPj

[Ano72-265] Anonymous. Focal points. IEEE Spectrum, 9(10):111–112, October 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FPk

[Ano72-266] Anonymous. Focal points. IEEE Spectrum, 9(11):93–96, November 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FPl

[Ano72-267] Anonymous. Focal points. IEEE Spectrum, 9(12):96–101, December 1972. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Fa

[Ano72-268] Anonymous. Forum. IEEE Spectrum, 9(1):16–19, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Fb

[Ano72-269] Anonymous. Forum. IEEE Spectrum, 9(2):13–15, February 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Fd

[Ano72-270] Anonymous. Forum. IEEE Spectrum, 9(3):10–18, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Fe

[Ano72-271] Anonymous. Forum. IEEE Spectrum, 9(4):10–16, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ff

[Ano72-272] Anonymous. Forum. IEEE Spectrum, 9(5):11–12, May 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Fg

[Ano72-273] Anonymous. Forum. IEEE Spectrum, 9(6):10–14, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Fh

[Ano72-274] Anonymous. Forum. IEEE Spectrum, 9(7):12–16, July 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Fi

[Ano72-275] Anonymous. Forum. IEEE Spectrum, 9(8):14–18, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Fj

[Ano72-276] Anonymous. Forum. IEEE Spectrum, 9(9):14–16, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Fk

[Ano72-277] Anonymous. Forum. IEEE Spectrum, 9(10):12–22, October 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Fl

[Ano72-278] Anonymous. Forum. IEEE Spectrum, 9(11):12–13, November 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Fm

[Ano72-279] Anonymous. Forum. IEEE Spectrum, 9(12):16–17, December 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FCa

[Ano72-280] Anonymous. [front cover]. IEEE Spectrum, 9(1):c1, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FCb

[Ano72-281] Anonymous. [front cover]. IEEE Spectrum, 9(1):c2, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FCc

[Ano72-282] Anonymous. [front cover].

IEEE Spectrum, 9(2):c1, February 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FCd

[Ano72-283] Anonymous. [front cover].

IEEE Spectrum, 9(2):c2, February 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FCe

[Ano72-284] Anonymous. [front cover]. IEEE Spectrum, 9(3):c1, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FCf

[Ano72-285] Anonymous. [front cover]. IEEE Spectrum, 9(3):c2, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FCg

[Ano72-286] Anonymous. [front cover]. IEEE Spectrum, 9(4):c1, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FCh

[Ano72-287] Anonymous. [front cover]. IEEE Spectrum, 9(4):c2, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FCi

[Ano72-288] Anonymous. [front cover]. IEEE Spectrum, 9(5):c1, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FCj

[Ano72-289] Anonymous. [front cover]. IEEE Spectrum, 9(5):c2, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FCk

[Ano72-290] Anonymous. [front cover]. IEEE Spectrum, 9(6):c1, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FCl

[Ano72-291] Anonymous. [front cover]. IEEE Spectrum, 9(6):c2, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FCm

[Ano72-292] Anonymous. [front cover]. IEEE Spectrum, 9(7):c1, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FCn

[Ano72-293] Anonymous. [front cover]. IEEE Spectrum, 9(7):c2, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FCo

[Ano72-294] Anonymous. [front cover]. IEEE Spectrum, 9(8):c1, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FCp

[Ano72-295] Anonymous. [front cover]. IEEE Spectrum, 9(8):c2, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FCq

[Ano72-296] Anonymous. [front cover]. IEEE Spectrum, 9(9):c1, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FCr

[Ano72-297] Anonymous. [front cover]. IEEE Spectrum, 9(9):c2, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FCs

[Ano72-298] Anonymous. [front cover]. IEEE Spectrum, 9(10):c1, October 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FCt

[Ano72-299] Anonymous. [front cover]. IEEE Spectrum, 9(10):c2, October 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FCu

[Ano72-300] Anonymous. [front cover]. IEEE Spectrum, 9(11):c1, November 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FCv

[Ano72-301] Anonymous. [front cover]. IEEE Spectrum, 9(11):c2, November 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FCw

[Ano72-302] Anonymous. [front cover]. IEEE Spectrum, 9(12):c1, December 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FCx

[Ano72-303] Anonymous. [front cover]. IEEE Spectrum, 9(12):c2, December 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FSIa

[Ano72-304] Anonymous. Future special issues. IEEE Spectrum, 9(1): 100–103, January 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FSIb

[Ano72-305] Anonymous. Future special issues. IEEE Spectrum, 9(2): 82–86, February 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FSIc

[Ano72-306] Anonymous. Future special issues. IEEE Spectrum, 9(3): 102–105, March 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FSId

[Ano72-307] Anonymous. Future special issues. IEEE Spectrum, 9 (4):83–85, April 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FSIe

[Ano72-308] Anonymous. Future special issues. IEEE Spectrum, 9 (5):78–82, May 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FSIf

[Ano72-309] Anonymous. Future special issues. IEEE Spectrum, 9 (6):81–84, June 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FSIg

[Ano72-310] Anonymous. Future special issues. IEEE Spectrum, 9(7):78– 80, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FSIh

[Ano72-311] Anonymous. Future special issues. IEEE Spectrum, 9(8): 108–112, August 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FSIi

[Ano72-312] Anonymous. Future special issues. IEEE Spectrum, 9(9): 91–95, September 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FSIj

[Ano72-313] Anonymous. Future special issues. IEEE Spectrum, 9(10): 105–107, October 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FSIk

[Ano72-314] Anonymous. Future special issues. IEEE Spectrum, 9(11): 87–89, November 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:FSIl

[Ano72-315] Anonymous. Future special issues. IEEE Spectrum, 9(12): 89–92, December 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:GIF

[Ano72-316] Anonymous. General information and fees. IEEE Spectrum, 9(3):116, March 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:GS

[Ano72-317] Anonymous. The grand scheme.

IEEE Spectrum, 9(9):35–43, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:HMD

[Ano72-318] Anonymous. How much do you make? IEEE Spectrum, 9 (8):82–87, August 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IGA

[Ano72-319] Anonymous. IEEE guide for authors. IEEE Spectrum, 9 (1):97–98, January 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IPOa

[Ano72-320] Anonymous. IEEE publications operations. IEEE Spectrum, 9(1):4, January 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IPOb

[Ano72-321] Anonymous. IEEE publications operations. IEEE Spectrum, 9 (2):4, February 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IPOc

[Ano72-322] Anonymous. IEEE publications operations. IEEE Spectrum, 9(3):4, March 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IPOd

[Ano72-323] Anonymous. IEEE publications operations. IEEE Spectrum, 9(4):4, April 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IPOe

[Ano72-324] Anonymous. IEEE publications operations. IEEE Spectrum, 9(5):4, May 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IPOf

[Ano72-325] Anonymous. IEEE publications operations. IEEE Spectrum, 9(6):4, June 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IPOg

[Ano72-326] Anonymous. IEEE publications operations. IEEE Spectrum, 9(7):4, July 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IPOh

[Ano72-327] Anonymous. IEEE publications operations. IEEE Spectrum, 9(8):4, August 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IPOi

[Ano72-328] Anonymous. IEEE publications operations. IEEE Spectrum, 9 (9):4, September 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IPOj

[Ano72-329] Anonymous. IEEE publications operations. IEEE Spectrum, 9 (10):4, October 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IPOk

[Ano72-330] Anonymous. IEEE publications operations. IEEE Spectrum, 9 (11):4, November 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IPOl

[Ano72-331] Anonymous. IEEE publications operations. IEEE Spectrum, 9 (12):4, December 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IR

[Ano72-332] Anonymous. IEEE reports for 1971. IEEE Spectrum, 9 (6):58–69, June 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ISa

[Ano72-333] Anonymous. IEEE Standards.

IEEE Spectrum, 9(1):104, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ISb

[Ano72-334] Anonymous. IEEE Standards. IEEE Spectrum, 9(2):87, February 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ISc

[Ano72-335] Anonymous. IEEE Standards. IEEE Spectrum, 9(2): 87–88, February 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ISd

[Ano72-336] Anonymous. IEEE Standards. IEEE Spectrum, 9(3):106, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ISe

[Ano72-337] Anonymous. IEEE Standards. IEEE Spectrum, 9(3):106–107, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ISf

[Ano72-338] Anonymous. IEEE Standards. IEEE Spectrum, 9(4):86, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ISg

[Ano72-339] Anonymous. IEEE Standards. IEEE Spectrum, 9(5):83, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ISh

[Ano72-340] Anonymous. IEEE Standards.

IEEE Spectrum, 9(6):85, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ISi

[Ano72-341] Anonymous. IEEE Standards.

IEEE Spectrum, 9(7):81, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ISj

[Ano72-342] Anonymous. IEEE Standards.

IEEE Spectrum, 9(8):112–113, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ISk

[Ano72-343] Anonymous. IEEE Standards. IEEE Spectrum, 9(9): 95–96, September 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ISl

[Ano72-344] Anonymous. IEEE Standards. IEEE Spectrum, 9(10): 107–108, October 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ISm

[Ano72-345] Anonymous. IEEE Standards. IEEE Spectrum, 9(11): 89–90, November 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:ISn

[Ano72-346] Anonymous. IEEE Standards. IEEE Spectrum, 9(12): 92–93, December 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IAa

[Ano72-347] Anonymous. Index to advertisers. IEEE Spectrum, 9 (1):120, January 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IAb

[Ano72-348] Anonymous. Index to advertisers. IEEE Spectrum, 9 (2):104, February 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IAc

[Ano72-349] Anonymous. Index to advertisers. IEEE Spectrum, 9(3): 126–128, March 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IAd

[Ano72-350] Anonymous. Index to advertisers. IEEE Spectrum, 9(4):104, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IAe

[Ano72-351] Anonymous. Index to advertisers. IEEE Spectrum, 9(5):96, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IAf

[Ano72-352] Anonymous. Index to advertisers. IEEE Spectrum, 9(6):100, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IAg

[Ano72-353] Anonymous. Index to advertisers. IEEE Spectrum, 9(7):96, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IAh

[Ano72-354] Anonymous. Index to advertisers. IEEE Spectrum, 9(8):130, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IAi

[Ano72-355] Anonymous. Index to advertisers. IEEE Spectrum, 9(9): 113–116, September 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IAj

[Ano72-356] Anonymous. Index to advertisers. IEEE Spectrum, 9 (10):122, October 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IAk

[Ano72-357] Anonymous. Index to advertisers. IEEE Spectrum, 9(11): 112, November 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IAl

[Ano72-358] Anonymous. Index to advertisers. IEEE Spectrum, 9(12): 128, December 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IASa

[Ano72-359] Anonymous. Information about the society. IEEE Spectrum, 9 (1):121–122, January 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IASb

[Ano72-360] Anonymous. Information about the society. IEEE Spectrum, 9 (2):105–106, February 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IASc

[Ano72-361] Anonymous. Information about the society. IEEE Spectrum, 9 (3):129–130, March 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IASd

[Ano72-362] Anonymous. Information about the society. IEEE Spectrum, 9 (8):33–34, August 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IIa

[Ano72-363] Anonymous. Inside IEEE. IEEE Spectrum, 9(1):7–12, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IIb

[Ano72-364] Anonymous. Inside IEEE. IEEE Spectrum, 9(2):7–12, February 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IIc

[Ano72-365] Anonymous. Inside IEEE. IEEE Spectrum, 9(3):7–9, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IId

[Ano72-366] Anonymous. Inside IEEE. IEEE Spectrum, 9(4):7–9, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IIe

[Ano72-367] Anonymous. Inside IEEE. IEEE Spectrum, 9(5):7–10, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IIf

[Ano72-368] Anonymous. Inside IEEE. IEEE Spectrum, 9(6):7–9, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IIg

[Ano72-369] Anonymous. Inside IEEE. IEEE Spectrum, 9(7):7–9, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IIh

[Ano72-370] Anonymous. Inside IEEE. IEEE Spectrum, 9(8):7–11, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IIi

[Ano72-371] Anonymous. Inside IEEE. IEEE Spectrum, 9(9):7–10, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IIj

[Ano72-372] Anonymous. Inside IEEE. IEEE Spectrum, 9(10):7–10, October 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IIk

[Ano72-373] Anonymous. Inside IEEE. IEEE Spectrum, 9(11):7–11, November 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IIl

[Ano72-374] Anonymous. Inside IEEE. IEEE Spectrum, 9(12):7–12, December 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IEEa

[Ano72-375] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 9(1):6, January 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IEEb

[Ano72-376] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 9(2):6, February 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IEEc

[Ano72-377] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 9(3):6, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IEEd

[Ano72-378] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 9(4): 6, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IEEe

[Ano72-379] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 9(5): 6, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IEEf

[Ano72-380] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 9(6): 6, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IEEg

[Ano72-381] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 9(7): 6, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IEEh

[Ano72-382] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 9(8):6, August 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IEEi

[Ano72-383] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 9 (9):6, September 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IEEj

[Ano72-384] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 9(10):6, October 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IEEk

[Ano72-385] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 9 (11):6, November 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IEEl

[Ano72-386] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 9 (12):6, December 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IG

[Ano72-387] Anonymous. Intercon at a glance. IEEE Spectrum, 9(2): 24–25, February 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:IRS

[Ano72-388] Anonymous. Intercon reflects a shift in technological priorities. IEEE Spectrum, 9(2): 20–23, February 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:JES

[Ano72-389] Anonymous. Japan electronics show. IEEE Spectrum, 9 (8):19–22, August 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:LEB

[Ano72-390] Anonymous. List of exhibitors and booths [advertisement]. IEEE Spectrum, 9(9): w80–w86, September 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ma

[Ano72-391] Anonymous. Meetings. IEEE Spectrum, 9(1):112–115, January 1972. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Mb

[Ano72-392] Anonymous. Meetings. IEEE Spectrum, 9(2):94–96, February 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Mc

[Ano72-393] Anonymous. Meetings. IEEE Spectrum, 9(3):113–114, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Md

[Ano72-394] Anonymous. Meetings. IEEE Spectrum, 9(4):95–99, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Me

[Ano72-395] Anonymous. Meetings. IEEE Spectrum, 9(5):88–91, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Mf

[Ano72-396] Anonymous. Meetings. IEEE Spectrum, 9(6):91–95, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Mg

[Ano72-397] Anonymous. Meetings. IEEE Spectrum, 9(7):87–90, July 1972.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Mh

[Ano72-398] Anonymous. Meetings. IEEE Spectrum, 9(8):120–125, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Mi

[Ano72-399] Anonymous. Meetings. IEEE Spectrum, 9(9):105–110, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Mj

[Ano72-400] Anonymous. Meetings. IEEE Spectrum, 9(10):113–116, October 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Mk

[Ano72-401] Anonymous. Meetings. IEEE Spectrum, 9(11):98–106, November 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ml

[Ano72-402] Anonymous. Meetings. IEEE Spectrum, 9(12):102–106, December 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:MDa

[Ano72-403] Anonymous. Moscow diary. IEEE Spectrum, 9(9):

74–80, September 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:NNM

[Ano72-404] Anonymous. NASA: the new medic. IEEE Spectrum, 9 (3):71–73, March 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:NMS

[Ano72-405] Anonymous. A new morality: a second look. IEEE Spectrum, 9(5):59–61, May 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:NPAa

[Ano72-406] Anonymous. New product applications. IEEE Spectrum, 9 (1):87–92, January 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:NPAb

[Ano72-407] Anonymous. New product applications. IEEE Spectrum, 9 (2):72–77, February 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:NPAc

[Ano72-408] Anonymous. New product applications. IEEE Spectrum, 9 (3):92–97, March 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:NPAd

[Ano72-409] Anonymous. New product applications. IEEE Spectrum, 9 (4):73–78, April 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:NPAe

[Ano72-410] Anonymous. New product applications. IEEE Spectrum, 9 (5):68–73, May 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:NPAf

[Ano72-411] Anonymous. New product applications. IEEE Spectrum, 9 (6):70–75, June 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:NPAg

[Ano72-412] Anonymous. New product applications. IEEE Spectrum, 9 (7):67–72, July 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:NPAh

[Ano72-413] Anonymous. New product applications. IEEE Spectrum, 9 (8):96–101, August 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:NPAi

[Ano72-414] Anonymous. New product applications. IEEE Spectrum, 9 (9):81–86, September 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:NPAj

[Ano72-415] Anonymous. New product applications. IEEE Spectrum, 9 (10):95–100, October 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:NPAk

[Ano72-416] Anonymous. New product applications. IEEE Spectrum, 9 (11):77–82, November 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:NPAl

[Ano72-417] Anonymous. New product applications. IEEE Spectrum, 9 (12):79–84, December 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:NP

[Ano72-418] Anonymous. New products. IEEE Spectrum, 9(3):s29–s30, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:NWa

[Ano72-419] Anonymous. News from Washington. IEEE Spectrum, 9(1): 21–22, January 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:NWb

[Ano72-420] Anonymous. News from Washington. IEEE Spectrum, 9(2): 17–18, February 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:NWc

[Ano72-421] Anonymous. News from Washington. IEEE Spectrum, 9 (3):21–22, March 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:NWd

[Ano72-422] Anonymous. News from Washington. IEEE Spectrum, 9 (4):17–18, April 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:NWe

[Ano72-423] Anonymous. News from Washington. IEEE Spectrum, 9 (5):13–14, May 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:NWf

[Ano72-424] Anonymous. News from Washington. IEEE Spectrum, 9 (6):15–16, June 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:NWg

[Ano72-425] Anonymous. News from Washington. IEEE Spectrum, 9(7):17– 18, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:NWh

[Ano72-426] Anonymous. News from Washington. IEEE Spectrum, 9 (8):23–24, August 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:NWi

[Ano72-427] Anonymous. News from Washington. IEEE Spectrum, 9(9):

17–18, September 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:NWj

[Ano72-428] Anonymous. News from Washington. IEEE Spectrum, 9(10): 25–26, October 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:NWk

[Ano72-429] Anonymous. News from Washington. IEEE Spectrum, 9(11): 15–16, November 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:NWl

[Ano72-430] Anonymous. News from Washington. IEEE Spectrum, 9(12): 19–20, December 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Oa

[Ano72-431] Anonymous. Obituaries. IEEE Spectrum, 9(1):117–119, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ob

[Ano72-432] Anonymous. Obituaries. IEEE Spectrum, 9(2):103, February 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Oc

[Ano72-433] Anonymous. Obituaries. IEEE Spectrum, 9(3):125, March 1972.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Od

[Ano72-434] Anonymous. Obituaries. IEEE Spectrum, 9(4):103, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Oe

[Ano72-435] Anonymous. Obituaries. IEEE Spectrum, 9(5):95, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Of

[Ano72-436] Anonymous. Obituaries. IEEE Spectrum, 9(7):95, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Og

[Ano72-437] Anonymous. Obituaries. IEEE Spectrum, 9(8):129, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Oh

[Ano72-438] Anonymous. Obituaries. IEEE Spectrum, 9(9):115, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Oi

[Ano72-439] Anonymous. Obituaries. IEEE Spectrum, 9(10):121, October 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Oj

[Ano72-440] Anonymous. Obituaries. IEEE Spectrum, 9(11):111, November 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ok

[Ano72-441] Anonymous. Obituaries. IEEE Spectrum, 9(12):113, December 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Pa

[Ano72-442] Anonymous. People. IEEE Spectrum, 9(1):117, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Pb

[Ano72-443] Anonymous. People. IEEE Spectrum, 9(2):100–102, February 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Pc

[Ano72-444] Anonymous. People. IEEE Spectrum, 9(3):124–125, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Pd

[Ano72-445] Anonymous. People. IEEE Spectrum, 9(4):102–103, April 1972.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Pe

[Ano72-446] Anonymous. People. IEEE Spectrum, 9(5):94–95, May 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Pf

[Ano72-447] Anonymous. People. IEEE Spectrum, 9(6):98–99, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Pg

[Ano72-448] Anonymous. People. IEEE Spectrum, 9(7):94–95, July 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Ph

[Ano72-449] Anonymous. People. IEEE Spectrum, 9(8):128–129, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Pi

[Ano72-450] Anonymous. People. IEEE Spectrum, 9(9):114–115, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Pj

[Ano72-451] Anonymous. People. IEEE Spectrum, 9(10):119–121, October 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Pk

[Ano72-452] Anonymous. People. IEEE Spectrum, 9(11):110–111, November 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:Pl

[Ano72-453] Anonymous. People. IEEE Spectrum, 9(12):110–113, December 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:PNS

[Ano72-454] Anonymous. President Nixon and Senator McGovern respond to engineering societies’ questions on technical issues. IEEE Spectrum, 9(10):nil4–72, October 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:PCSa

[Ano72-455] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 9(1):113–115, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:PCSb

[Ano72-456] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 9(2):101, February 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:PCSc

[Ano72-457] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 9(3):127, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:PCSd

[Ano72-458] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 9(4):101, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:PCSe

[Ano72-459] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 9(5):93, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:PCSf

[Ano72-460] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 9(6):97, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:PCSg

[Ano72-461] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 9(7):92–93, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:PCSh

[Ano72-462] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 9(8):127, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:PCSi

[Ano72-463] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 9(11):109, November 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:PCSj

[Ano72-464] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 9(12):109, December 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SIa

[Ano72-465] Anonymous. Scanning the issues. IEEE Spectrum, 9(1):94–95, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SIb

[Ano72-466] Anonymous. Scanning the issues. IEEE Spectrum, 9(2): 79–80, February 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SIc

[Ano72-467] Anonymous. Scanning the issues. IEEE Spectrum, 9(3):

99–100, March 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SId

[Ano72-468] Anonymous. Scanning the issues. IEEE Spectrum, 9(4): 80–81, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SIe

[Ano72-469] Anonymous. Scanning the issues.

IEEE Spectrum, 9(5):75–76, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SIf

[Ano72-470] Anonymous. Scanning the issues. IEEE Spectrum, 9(6): 78–79, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SIg

[Ano72-471] Anonymous. Scanning the issues. IEEE Spectrum, 9(7):74– 75, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SIh

[Ano72-472] Anonymous. Scanning the issues. IEEE Spectrum, 9(8): 105–106, August 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SIi

[Ano72-473] Anonymous. Scanning the issues. IEEE Spectrum, 9(9):

88–89, September 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SIj

[Ano72-474] Anonymous. Scanning the issues. IEEE Spectrum, 9(10): 102–103, October 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SIk

[Ano72-475] Anonymous. Scanning the issues. IEEE Spectrum, 9(11): 84–85, November 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SIl

[Ano72-476] Anonymous. Scanning the issues. IEEE Spectrum, 9(12): 86–87, December 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SSAa

[Ano72-477] Anonymous. Science/scope [advertisement]. IEEE Spectrum, 9(2):nil2, February 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SSAb

[Ano72-478] Anonymous. Science/scope [advertisement]. IEEE Spectrum, 9(8):5, August 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SSAc

[Ano72-479] Anonymous. Science/scope [advertisement]. IEEE Spectrum, 9 (10):nil8, October 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SSAd

[Ano72-480] Anonymous. Science/scope [advertisement]. IEEE Spectrum, 9(12):5, December 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SPa

[Ano72-481] Anonymous. Special publications. IEEE Spectrum, 9(1): 104–105, January 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SPb

[Ano72-482] Anonymous. Special publications. IEEE Spectrum, 9 (4):86–87, April 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SPc

[Ano72-483] Anonymous. Special publications. IEEE Spectrum, 9 (5):83–84, May 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SPd

[Ano72-484] Anonymous. Special publications. IEEE Spectrum, 9 (6):85–86, June 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SPe

[Ano72-485] Anonymous. Special publications. IEEE Spectrum, 9(7):81– 82, July 1972. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SPf

[Ano72-486] Anonymous. Special publications. IEEE Spectrum, 9(8): 113–114, August 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SPg

[Ano72-487] Anonymous. Special publications. IEEE Spectrum, 9(9): 96–98, September 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SPh

[Ano72-488] Anonymous. Special publications. IEEE Spectrum, 9(10): 108–109, October 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SPi

[Ano72-489] Anonymous. Special publications. IEEE Spectrum, 9(11): 90–91, November 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SPj

[Ano72-490] Anonymous. Special publications. IEEE Spectrum, 9(12): 93–94, December 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SSE

[Ano72-491] Anonymous. Special show extra. IEEE Spectrum, 9(3):s3, March 1972. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SLa

[Ano72-492] Anonymous. Spectral lines.

IEEE Spectrum, 9(1):39, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SLb

[Ano72-493] Anonymous. Spectral lines.

IEEE Spectrum, 9(2):19, February 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SLc

[Ano72-494] Anonymous. Spectral lines.

IEEE Spectrum, 9(3):25, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SLd

[Ano72-495] Anonymous. Spectral lines.

IEEE Spectrum, 9(4):19, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SLe

[Ano72-496] Anonymous. Spectral lines.

IEEE Spectrum, 9(5):15, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SLf

[Ano72-497] Anonymous. Spectral lines.

IEEE Spectrum, 9(6):17, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SLg

[Ano72-498] Anonymous. Spectral lines. IEEE Spectrum, 9(7):19, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SLh

[Ano72-499] Anonymous. Spectral lines. IEEE Spectrum, 9(8):35, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SLi

[Ano72-500] Anonymous. Spectral lines. IEEE Spectrum, 9(9):21, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SLj

[Ano72-501] Anonymous. Spectral lines. IEEE Spectrum, 9(10):31, October 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SLk

[Ano72-502] Anonymous. Spectral lines. IEEE Spectrum, 9(11):17, November 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SLl

[Ano72-503] Anonymous. Spectral lines. IEEE Spectrum, 9(12):23, December 1972. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:SICa

[Ano72-504] Anonymous. Statements from IEEE candidates. IEEE Spectrum, 9(8):90–95, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:TCa

[Ano72-505] Anonymous. Table of contents.

IEEE Spectrum, 9(1):2–3, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:TCb

[Ano72-506] Anonymous. Table of contents.

IEEE Spectrum, 9(2):2–3, February 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:TCc

[Ano72-507] Anonymous. Table of contents. IEEE Spectrum, 9(3):2–3, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:TCd

[Ano72-508] Anonymous. Table of contents. IEEE Spectrum, 9(4):2–3, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:TCe

[Ano72-509] Anonymous. Table of contents. IEEE Spectrum, 9(5):2– 3, May 1972. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:TCf

[Ano72-510] Anonymous. Table of contents. IEEE Spectrum, 9(6):2–3, June 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:TCg

[Ano72-511] Anonymous. Table of contents. IEEE Spectrum, 9(7):2– 3, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:TCh

[Ano72-512] Anonymous. Table of contents. IEEE Spectrum, 9(8):2–3, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:TCi

[Ano72-513] Anonymous. Table of contents. IEEE Spectrum, 9(9): 2–3, September 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:TCj

[Ano72-514] Anonymous. Table of contents. IEEE Spectrum, 9(10):2–3, October 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:TCk

[Ano72-515] Anonymous. Table of contents. IEEE Spectrum, 9(11):

2–3, November 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:TCl

[Ano72-516] Anonymous. Table of contents.

IEEE Spectrum, 9(12):2–3, December 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:TP

[Ano72-517] Anonymous. Technical program.

IEEE Spectrum, 9(1):23–37, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:UE

[Ano72-518] Anonymous. The unemployed EE. IEEE Spectrum, 9(10): 74–77, October 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:WTE

[Ano72-519] Anonymous. WESCON ’72: a ‘two-edged’ program. IEEE Spectrum, 9(9):w12–w18, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:WPG

[Ano72-520] Anonymous. WESCON 72:

products galore. IEEE Spectrum, 9(9):w54–w79, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:WMU

[Ano72-521] Anonymous. WESCON 72: the mood is ‘upturn’. IEEE Spectrum, 9(9):w6–w9, September 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:WRC

[Ano72-522] Anonymous. WESCON 72: the red carpet. IEEE Spectrum, 9 (9):w3–w5, September 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:WEW

[Ano72-523] Anonymous. What EEs want. IEEE Spectrum, 9(3):48–52, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1972:WNS

[Ano72-524] Anonymous. What’s new in small displays. IEEE Spectrum, 9 (11):19–24, November 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:TY

[Ano73a] Anonymous. 1972: a ‘threshold’ year. IEEE Spectrum, 10(7):56– 66, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ATCa

[Ano73b] Anonymous. Advance tables of contents. IEEE Spectrum, 10 (1):91–92, January 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ATCb

[Ano73c] Anonymous. Advance tables of contents. IEEE Spectrum, 10 (2):77–78, February 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ATCc

[Ano73d] Anonymous. Advance tables of contents. IEEE Spectrum, 10 (3):89–91, March 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aa

[Ano73e] Anonymous. [advertisement].

IEEE Spectrum, 10(1):nil1– 1, January 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ab

[Ano73f] Anonymous. [advertisement].

IEEE Spectrum, 10(1):4, January 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ac

[Ano73g] Anonymous. [advertisement].

IEEE Spectrum, 10(1):5, January 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ad

[Ano73h] Anonymous. [advertisement].

IEEE Spectrum, 10(1):6, January 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ae

[Ano73i] Anonymous. [advertisement]. IEEE Spectrum, 10(1):10–11, January 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Af

[Ano73j] Anonymous. [advertisement]. IEEE Spectrum, 10(1):13–18, January 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ag

[Ano73k] Anonymous. [advertisement]. IEEE Spectrum, 10(1):21–22, January 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ah

[Ano73l] Anonymous. [advertisement]. IEEE Spectrum, 10(1):43, January 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ai

[Ano73m] Anonymous. [advertisement]. IEEE Spectrum, 10(1):88, January 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aj

[Ano73n] Anonymous. [advertisement]. IEEE Spectrum, 10(1):97, January 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ak

[Ano73o] Anonymous. [advertisement].

IEEE Spectrum, 10(1):111– 115, January 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Am

[Ano73p] Anonymous. [advertisement].

IEEE Spectrum, 10(1):123– 124, January 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:An

[Ano73q] Anonymous. [advertisement].

IEEE Spectrum, 10(1):nil2, January 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ao

[Ano73r] Anonymous. [advertisement].

IEEE Spectrum, 10(2):4, February 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ap

[Ano73s] Anonymous. [advertisement].

IEEE Spectrum, 10(2):5, February 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aq

[Ano73t] Anonymous. [advertisement].

IEEE Spectrum, 10(2):6, February 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ar

[Ano73u] Anonymous. [advertisement]. IEEE Spectrum, 10(2):15, February 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:As

[Ano73v] Anonymous. [advertisement]. IEEE Spectrum, 10(2):17–20, February 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:At

[Ano73w] Anonymous. [advertisement]. IEEE Spectrum, 10(2):23–24, February 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Au

[Ano73x] Anonymous. [advertisement]. IEEE Spectrum, 10(2):43, February 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Av

[Ano73y] Anonymous. [advertisement]. IEEE Spectrum, 10(2):81, February 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aw

[Ano73z] Anonymous. [advertisement]. IEEE Spectrum, 10(2):91–93, February 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ax

[Ano73-27] Anonymous. [advertisement].

IEEE Spectrum, 10(2):95, February 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ay

[Ano73-28] Anonymous. [advertisement].

IEEE Spectrum, 10(2):107– 108, February 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Az

[Ano73-29] Anonymous. [advertisement].

IEEE Spectrum, 10(2):nil2, February 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aba

[Ano73-30] Anonymous. [advertisement].

IEEE Spectrum, 10(3):nil1–1, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Abb

[Ano73-31] Anonymous. [advertisement].

IEEE Spectrum, 10(3):4, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Abc

[Ano73-32] Anonymous. [advertisement].

IEEE Spectrum, 10(3):5, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Abd

[Ano73-33] Anonymous. [advertisement]. IEEE Spectrum, 10(3):6, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Abe

[Ano73-34] Anonymous. [advertisement]. IEEE Spectrum, 10(3):9–13, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Abf

[Ano73-35] Anonymous. [advertisement]. IEEE Spectrum, 10(3):14–20, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Abg

[Ano73-36] Anonymous. [advertisement]. IEEE Spectrum, 10(3):23–a2, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Abh

[Ano73-37] Anonymous. [advertisement]. IEEE Spectrum, 10(3):a5–a6, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Abi

[Ano73-38] Anonymous. [advertisement]. IEEE Spectrum, 10(3):a9–a15, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Abj

[Ano73-39] Anonymous. [advertisement].

IEEE Spectrum, 10(3):a17–a22, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Abk

[Ano73-40] Anonymous. [advertisement].

IEEE Spectrum, 10(3):a25–a33, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Abl

[Ano73-41] Anonymous. [advertisement].

IEEE Spectrum, 10(3):a35–a44, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Abm

[Ano73-42] Anonymous. [advertisement].

IEEE Spectrum, 10(3):98–101, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Abn

[Ano73-43] Anonymous. [advertisement].

IEEE Spectrum, 10(3):113–114, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Abo

[Ano73-44] Anonymous. [advertisement].

IEEE Spectrum, 10(3):nil2, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Abp

[Ano73-45] Anonymous. [advertisement]. IEEE Spectrum, 10(4):nil1–1b, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Abq

[Ano73-46] Anonymous. [advertisement]. IEEE Spectrum, 10(4):1, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Abr

[Ano73-47] Anonymous. [advertisement]. IEEE Spectrum, 10(4):4, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Abs

[Ano73-48] Anonymous. [advertisement]. IEEE Spectrum, 10(4):5, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Abt

[Ano73-49] Anonymous. [advertisement]. IEEE Spectrum, 10(4):6, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Abu

[Ano73-50] Anonymous. [advertisement]. IEEE Spectrum, 10(4):11–13, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Abv

[Ano73-51] Anonymous. [advertisement].

IEEE Spectrum, 10(4):15–18, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Abw

[Ano73-52] Anonymous. [advertisement].

IEEE Spectrum, 10(4):21–22, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Abx

[Ano73-53] Anonymous. [advertisement].

IEEE Spectrum, 10(4):33, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aby

[Ano73-54] Anonymous. [advertisement].

IEEE Spectrum, 10(4):78, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Abz

[Ano73-55] Anonymous. [advertisement].

IEEE Spectrum, 10(4):97–98, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aca

[Ano73-56] Anonymous. [advertisement].

IEEE Spectrum, 10(5):nil1–1, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Acb

[Ano73-57] Anonymous. [advertisement]. IEEE Spectrum, 10(5):4, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Acc

[Ano73-58] Anonymous. [advertisement]. IEEE Spectrum, 10(5):5, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Acd

[Ano73-59] Anonymous. [advertisement]. IEEE Spectrum, 10(5):6, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ace

[Ano73-60] Anonymous. [advertisement]. IEEE Spectrum, 10(5):12–14, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Acf

[Ano73-61] Anonymous. [advertisement]. IEEE Spectrum, 10(5):20, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Acg

[Ano73-62] Anonymous. [advertisement]. IEEE Spectrum, 10(5):48–49, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ach

[Ano73-63] Anonymous. [advertisement].

IEEE Spectrum, 10(5):65, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aci

[Ano73-64] Anonymous. [advertisement].

IEEE Spectrum, 10(5):79, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Acj

[Ano73-65] Anonymous. [advertisement].

IEEE Spectrum, 10(5):97–98, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ack

[Ano73-66] Anonymous. [advertisement].

IEEE Spectrum, 10(5):99–100, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Acl

[Ano73-67] Anonymous. [advertisement].

IEEE Spectrum, 10(5):nil2, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Acm

[Ano73-68] Anonymous. [advertisement].

IEEE Spectrum, 10(6):nil1–1, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Acn

[Ano73-69] Anonymous. [advertisement]. IEEE Spectrum, 10(6):4, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aco

[Ano73-70] Anonymous. [advertisement]. IEEE Spectrum, 10(6):5, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Acp

[Ano73-71] Anonymous. [advertisement]. IEEE Spectrum, 10(6):6, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Acq

[Ano73-72] Anonymous. [advertisement]. IEEE Spectrum, 10(6):10–11, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Acr

[Ano73-73] Anonymous. [advertisement]. IEEE Spectrum, 10(6):15, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Acs

[Ano73-74] Anonymous. [advertisement]. IEEE Spectrum, 10(6):19–20, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Act

[Ano73-75] Anonymous. [advertisement].

IEEE Spectrum, 10(6):29, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Acu

[Ano73-76] Anonymous. [advertisement].

IEEE Spectrum, 10(6):48–49, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Acv

[Ano73-77] Anonymous. [advertisement].

IEEE Spectrum, 10(6):97–98, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Acw

[Ano73-78] Anonymous. [advertisement].

IEEE Spectrum, 10(6):99–100, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Acx

[Ano73-79] Anonymous. [advertisement].

IEEE Spectrum, 10(6):nil2, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Acy

[Ano73-80] Anonymous. [advertisement].

IEEE Spectrum, 10(7):nil1–1, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Acz

[Ano73-81] Anonymous. [advertisement]. IEEE Spectrum, 10(7):4, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ada

[Ano73-82] Anonymous. [advertisement]. IEEE Spectrum, 10(7):5, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Adb

[Ano73-83] Anonymous. [advertisement]. IEEE Spectrum, 10(7):6, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Adc

[Ano73-84] Anonymous. [advertisement]. IEEE Spectrum, 10(7):9, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Add

[Ano73-85] Anonymous. [advertisement]. IEEE Spectrum, 10(7):11–12, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ade

[Ano73-86] Anonymous. [advertisement]. IEEE Spectrum, 10(7):15–16, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Adf

[Ano73-87] Anonymous. [advertisement].

IEEE Spectrum, 10(7):49, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Adg

[Ano73-88] Anonymous. [advertisement].

IEEE Spectrum, 10(7):55, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Adh

[Ano73-89] Anonymous. [advertisement].

IEEE Spectrum, 10(7):86, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Adi

[Ano73-90] Anonymous. [advertisement].

IEEE Spectrum, 10(7):97, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Adj

[Ano73-91] Anonymous. [advertisement].

IEEE Spectrum, 10(7):101–102, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Adk

[Ano73-92] Anonymous. [advertisement].

IEEE Spectrum, 10(7):nil2, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Adl

[Ano73-93] Anonymous. [advertisement]. IEEE Spectrum, 10(8):4, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Adm

[Ano73-94] Anonymous. [advertisement]. IEEE Spectrum, 10(8):5, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Adn

[Ano73-95] Anonymous. [advertisement]. IEEE Spectrum, 10(8):10–12, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ado

[Ano73-96] Anonymous. [advertisement]. IEEE Spectrum, 10(8):17, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Adp

[Ano73-97] Anonymous. [advertisement]. IEEE Spectrum, 10(8):19–22, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Adq

[Ano73-98] Anonymous. [advertisement]. IEEE Spectrum, 10(8):50, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Adr

[Ano73-99] Anonymous. [advertisement].

IEEE Spectrum, 10(8):79–87, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ads

[Ano73-100] Anonymous. [advertisement].

IEEE Spectrum, 10(8):93, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Adt

[Ano73-101] Anonymous. [advertisement].

IEEE Spectrum, 10(8):99–100, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Adu

[Ano73-102] Anonymous. [advertisement].

IEEE Spectrum, 10(8):nil2, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Adv

[Ano73-103] Anonymous. [advertisement].

IEEE Spectrum, 10(9):nil1–1, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Adw

[Ano73-104] Anonymous. [advertisement].

IEEE Spectrum, 10(9):4, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Adx

[Ano73-105] Anonymous. [advertisement]. IEEE Spectrum, 10(9):5–7, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ady

[Ano73-106] Anonymous. [advertisement]. IEEE Spectrum, 10(9):9–15, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Adz

[Ano73-107] Anonymous. [advertisement]. IEEE Spectrum, 10(9):17–19, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aea

[Ano73-108] Anonymous. [advertisement]. IEEE Spectrum, 10(9):21–25, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aeb

[Ano73-109] Anonymous. [advertisement]. IEEE Spectrum, 10(9):27–32, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aec

[Ano73-110] Anonymous. [advertisement]. IEEE Spectrum, 10(9):107–109, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aed

[Ano73-111] Anonymous. [advertisement].

IEEE Spectrum, 10(9):114–115, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aee

[Ano73-112] Anonymous. [advertisement].

IEEE Spectrum, 10(9):117–118, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aef

[Ano73-113] Anonymous. [advertisement].

IEEE Spectrum, 10(9):nil2, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aeg

[Ano73-114] Anonymous. [advertisement].

IEEE Spectrum, 10(10):nil1– 1, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aeh

[Ano73-115] Anonymous. [advertisement].

IEEE Spectrum, 10(10):4, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aei

[Ano73-116] Anonymous. [advertisement].

IEEE Spectrum, 10(10):5, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aej

[Ano73-117] Anonymous. [advertisement]. IEEE Spectrum, 10(10):6–7, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aek

[Ano73-118] Anonymous. [advertisement]. IEEE Spectrum, 10(10):8a– 8b, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ael

[Ano73-119] Anonymous. [advertisement]. IEEE Spectrum, 10(10):10– 11, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aem

[Ano73-120] Anonymous. [advertisement]. IEEE Spectrum, 10(10):13– 15, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aen

[Ano73-121] Anonymous. [advertisement]. IEEE Spectrum, 10(10):68, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aeo

[Ano73-122] Anonymous. [advertisement]. IEEE Spectrum, 10(10):79– 88, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aep

[Ano73-123] Anonymous. [advertisement].

IEEE Spectrum, 10(10):91, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aeq

[Ano73-124] Anonymous. [advertisement].

IEEE Spectrum, 10(10):103– 104, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aer

[Ano73-125] Anonymous. [advertisement].

IEEE Spectrum, 10(10):105– 106, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aes

[Ano73-126] Anonymous. [advertisement].

IEEE Spectrum, 10(10):nil2, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aet

[Ano73-127] Anonymous. [advertisement].

IEEE Spectrum, 10(11):nil1, November 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aeu

[Ano73-128] Anonymous. [advertisement].

IEEE Spectrum, 10(11):1, November 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aev

[Ano73-129] Anonymous. [advertisement]. IEEE Spectrum, 10(11):4, November 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aew

[Ano73-130] Anonymous. [advertisement]. IEEE Spectrum, 10(11):5, November 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aex

[Ano73-131] Anonymous. [advertisement]. IEEE Spectrum, 10(11):6, November 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aey

[Ano73-132] Anonymous. [advertisement]. IEEE Spectrum, 10(11):7, November 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aez

[Ano73-133] Anonymous. [advertisement]. IEEE Spectrum, 10(11):9, November 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Afa

[Ano73-134] Anonymous. [advertisement]. IEEE Spectrum, 10(11):13, November 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Afb

[Ano73-135] Anonymous. [advertisement].

IEEE Spectrum, 10(11):15, November 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Afc

[Ano73-136] Anonymous. [advertisement].

IEEE Spectrum, 10(11):21–26, November 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Afd

[Ano73-137] Anonymous. [advertisement].

IEEE Spectrum, 10(11):44, November 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Afe

[Ano73-138] Anonymous. [advertisement].

IEEE Spectrum, 10(11):107– 109, November 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aff

[Ano73-139] Anonymous. [advertisement].

IEEE Spectrum, 10(11):111, November 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Afg

[Ano73-140] Anonymous. [advertisement].

IEEE Spectrum, 10(11):117– 118, November 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Afh

[Ano73-141] Anonymous. [advertisement]. IEEE Spectrum, 10(11):nil4, November 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Afi

[Ano73-142] Anonymous. [advertisement]. IEEE Spectrum, 10(12):nil1– 1, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Afj

[Ano73-143] Anonymous. [advertisement]. IEEE Spectrum, 10(12):4, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Afk

[Ano73-144] Anonymous. [advertisement]. IEEE Spectrum, 10(12):5–7, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Afl

[Ano73-145] Anonymous. [advertisement]. IEEE Spectrum, 10(12):13, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Afm

[Ano73-146] Anonymous. [advertisement]. IEEE Spectrum, 10(12):15–16, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Afn

[Ano73-147] Anonymous. [advertisement].

IEEE Spectrum, 10(12):19, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Afo

[Ano73-148] Anonymous. [advertisement].

IEEE Spectrum, 10(12):21–24, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Afp

[Ano73-149] Anonymous. [advertisement].

IEEE Spectrum, 10(12):63, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Afq

[Ano73-150] Anonymous. [advertisement].

IEEE Spectrum, 10(12):90, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Afr

[Ano73-151] Anonymous. [advertisement].

IEEE Spectrum, 10(12):110– 111, December 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Afs

[Ano73-152] Anonymous. [advertisement].

IEEE Spectrum, 10(12):115, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Aft

[Ano73-153] Anonymous. [advertisement]. IEEE Spectrum, 10(12):121, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Afu

[Ano73-154] Anonymous. [advertisement]. IEEE Spectrum, 10(12):135, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Afv

[Ano73-155] Anonymous. [advertisement]. IEEE Spectrum, 10(12):139– 140, December 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Afw

[Ano73-156] Anonymous. [advertisement]. IEEE Spectrum, 10(12):nil2, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ALa

[Ano73-157] Anonymous. Applications literature. IEEE Spectrum, 10 (1):85, January 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ALb

[Ano73-158] Anonymous. Applications literature. IEEE Spectrum, 10 (2):74, February 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ALc

[Ano73-159] Anonymous. Applications literature. IEEE Spectrum, 10(3):86, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ALd

[Ano73-160] Anonymous. Applications literature. IEEE Spectrum, 10(4):74, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ALe

[Ano73-161] Anonymous. Applications literature. IEEE Spectrum, 10(5):72, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ALf

[Ano73-162] Anonymous. Applications literature. IEEE Spectrum, 10(6):71, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ALg

[Ano73-163] Anonymous. Applications literature. IEEE Spectrum, 10(7):73, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ALh

[Ano73-164] Anonymous. Applications literature. IEEE Spectrum, 10(8):75, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ALi

[Ano73-165] Anonymous. Applications literature. IEEE Spectrum, 10 (9):89, September 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ALj

[Ano73-166] Anonymous. Applications literature. IEEE Spectrum, 10 (10):75, October 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ALk

[Ano73-167] Anonymous. Applications literature. IEEE Spectrum, 10 (11):88, November 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ALl

[Ano73-168] Anonymous. Applications literature. IEEE Spectrum, 10 (12):97, December 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:BRa

[Ano73-169] Anonymous. Book reviews. IEEE Spectrum, 10(1):102– 105, January 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:BRb

[Ano73-170] Anonymous. Book reviews. IEEE Spectrum, 10(2):88–89, February 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:BRc

[Ano73-171] Anonymous. Book reviews.

IEEE Spectrum, 10(3):94, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:BRd

[Ano73-172] Anonymous. Book reviews.

IEEE Spectrum, 10(4):85, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:BRe

[Ano73-173] Anonymous. Book reviews.

IEEE Spectrum, 10(5):83, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:BRf

[Ano73-174] Anonymous. Book reviews.

IEEE Spectrum, 10(6):82, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:BRg

[Ano73-175] Anonymous. Book reviews.

IEEE Spectrum, 10(7):84, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:BRh

[Ano73-176] Anonymous. Book reviews.

IEEE Spectrum, 10(8):91, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:BRi

[Ano73-177] Anonymous. Book reviews. IEEE Spectrum, 10(9):110, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:BRj

[Ano73-178] Anonymous. Book reviews. IEEE Spectrum, 10(10):95– 96, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:BRk

[Ano73-179] Anonymous. Book reviews. IEEE Spectrum, 10(11):110– 111, November 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:BRl

[Ano73-180] Anonymous. Book reviews. IEEE Spectrum, 10(12):109– 111, December 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ISAa

[Ano73-181] Anonymous. IEEE Spectrum annual index. IEEE Spectrum, 10(12):122–134, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ISRa

[Ano73-182] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 10(1):123–124, January 1973. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ISRb

[Ano73-183] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 10(3):113–114, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ISRc

[Ano73-184] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 10(4):97–98, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ISRd

[Ano73-185] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 10(5):99–100, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ISRe

[Ano73-186] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 10(6):99–100, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ISRf

[Ano73-187] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 10(7):101–102, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ISRg

[Ano73-188] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 10(8):99–100, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ISRh

[Ano73-189] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 10(9):117–118, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ISRi

[Ano73-190] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 10(10):105–106, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ISRj

[Ano73-191] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 10(11):117–118, November 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ISRk

[Ano73-192] Anonymous. IEEE Spectrum reader service card. IEEE Spectrum, 10(12):139–140, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:BPa

[Ano73-193] Anonymous. [breaker page].

IEEE Spectrum, 10(1):4, January 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:BPb

[Ano73-194] Anonymous. [breaker page].

IEEE Spectrum, 10(2):4, February 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:BPc

[Ano73-195] Anonymous. [breaker page].

IEEE Spectrum, 10(3):4, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:BPd

[Ano73-196] Anonymous. [breaker page].

IEEE Spectrum, 10(4):4, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:BPe

[Ano73-197] Anonymous. [breaker page].

IEEE Spectrum, 10(5):4, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:BPf

[Ano73-198] Anonymous. [breaker page].

IEEE Spectrum, 10(6):4, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:BPg

[Ano73-199] Anonymous. [breaker page]. IEEE Spectrum, 10(7):4, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:BPh

[Ano73-200] Anonymous. [breaker page]. IEEE Spectrum, 10(8):4, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:BPi

[Ano73-201] Anonymous. [breaker page]. IEEE Spectrum, 10(9):4, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:BPj

[Ano73-202] Anonymous. [breaker page]. IEEE Spectrum, 10(10):4, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:BPk

[Ano73-203] Anonymous. [breaker page]. IEEE Spectrum, 10(11):nil2, November 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:BPl

[Ano73-204] Anonymous. [breaker page]. IEEE Spectrum, 10(11):4, November 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:BPm

[Ano73-205] Anonymous. [breaker page].

IEEE Spectrum, 10(12):4, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Cb

[Ano73-206] Anonymous. Calendar. IEEE Spectrum, 10(1):116–117, January 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Cc

[Ano73-207] Anonymous. Calendar. IEEE Spectrum, 10(2):98–99, February 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Cd

[Ano73-208] Anonymous. Calendar. IEEE Spectrum, 10(3):108–109, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ce

[Ano73-209] Anonymous. Calendar. IEEE Spectrum, 10(4):92–93, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Cf

[Ano73-210] Anonymous. Calendar. IEEE Spectrum, 10(5):91–92, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Cg

[Ano73-211] Anonymous. Calendar. IEEE Spectrum, 10(6):91–92, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ch

[Ano73-212] Anonymous. Calendar. IEEE Spectrum, 10(7):96–97, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ci

[Ano73-213] Anonymous. Calendar. IEEE Spectrum, 10(8):9–12, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Cj

[Ano73-214] Anonymous. Calendar. IEEE Spectrum, 10(9):16–17, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ck

[Ano73-215] Anonymous. Calendar. IEEE Spectrum, 10(10):12–13, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Cl

[Ano73-216] Anonymous. Calendar. IEEE Spectrum, 10(11):12–13, November 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Cm

[Ano73-217] Anonymous. Calendar. IEEE Spectrum, 10(12):12–13, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:CS

[Ano73-218] Anonymous. A celebration in stamps. IEEE Spectrum, 10 (7):24–25, July 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:CAa

[Ano73-219] Anonymous. Classified advertising. IEEE Spectrum, 10(1): 105–107, January 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:CAb

[Ano73-220] Anonymous. Classified advertising. IEEE Spectrum, 10(2): 100–101, February 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:CAc

[Ano73-221] Anonymous. Classified advertising. IEEE Spectrum, 10 (3):95–97, March 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:CAd

[Ano73-222] Anonymous. Classified advertising. IEEE Spectrum, 10(4):91, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:CAe

[Ano73-223] Anonymous. Classified advertising. IEEE Spectrum, 10 (5):86–87, May 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:CAf

[Ano73-224] Anonymous. Classified advertising. IEEE Spectrum, 10 (6):87–90, June 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:CAg

[Ano73-225] Anonymous. Classified advertising. IEEE Spectrum, 10(8):92, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:CAh

[Ano73-226] Anonymous. Classified advertising. IEEE Spectrum, 10(10): 94–95, October 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:CAi

[Ano73-227] Anonymous. Classified advertising. IEEE Spectrum, 10(11): 92–105, November 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:CAj

[Ano73-228] Anonymous. Classified advertising. IEEE Spectrum, 10(12): 101–105, December 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:CEa

[Ano73-229] Anonymous. Consumerism and the EE. IEEE Spectrum, 10 (5):46–47, May 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:EAa

[Ano73-230] Anonymous. Educational aids.

IEEE Spectrum, 10(1):101, January 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:EAb

[Ano73-231] Anonymous. Educational aids.

IEEE Spectrum, 10(2):87, February 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:EAc

[Ano73-232] Anonymous. Educational aids.

IEEE Spectrum, 10(3):93, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:EAd

[Ano73-233] Anonymous. Educational aids.

IEEE Spectrum, 10(4):84, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:EAe

[Ano73-234] Anonymous. Educational aids.

IEEE Spectrum, 10(5):82, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:EAf

[Ano73-235] Anonymous. Educational aids. IEEE Spectrum, 10(6):81, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:EAg

[Ano73-236] Anonymous. Educational aids. IEEE Spectrum, 10(7):83, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:EAh

[Ano73-237] Anonymous. Educational aids. IEEE Spectrum, 10(8):90, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:EAi

[Ano73-238] Anonymous. Educational aids. IEEE Spectrum, 10(9):108, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:EAj

[Ano73-239] Anonymous. Educational aids. IEEE Spectrum, 10(10):92, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:EAk

[Ano73-240] Anonymous. Educational aids. IEEE Spectrum, 10(11):108, November 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:EAl

[Ano73-241] Anonymous. Educational aids.

IEEE Spectrum, 10(12):108, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:EG

[Ano73-242] Anonymous. Employment guidelines. IEEE Spectrum, 10 (4):57–60, April 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FPa

[Ano73-243] Anonymous. Focal points. IEEE Spectrum, 10(1):107–109, January 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FPb

[Ano73-244] Anonymous. Focal points. IEEE Spectrum, 10(2):90–93, February 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FPc

[Ano73-245] Anonymous. Focal points. IEEE Spectrum, 10(3):97–100, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FPd

[Ano73-246] Anonymous. Focal points. IEEE Spectrum, 10(4):86–87, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FPe

[Ano73-247] Anonymous. Focal points. IEEE Spectrum, 10(5):84–85, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FPf

[Ano73-248] Anonymous. Focal points. IEEE Spectrum, 10(6):83–85, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FPg

[Ano73-249] Anonymous. Focal points. IEEE Spectrum, 10(7):85–86, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FPh

[Ano73-250] Anonymous. Focal points. IEEE Spectrum, 10(8):15–17, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FPi

[Ano73-251] Anonymous. Focal points. IEEE Spectrum, 10(9):20–23, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FPj

[Ano73-252] Anonymous. Focal points. IEEE Spectrum, 10(10):16, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FPk

[Ano73-253] Anonymous. Focal points. IEEE Spectrum, 10(11):14–16, November 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FPl

[Ano73-254] Anonymous. Focal points. IEEE Spectrum, 10(12):14–16, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Fa

[Ano73-255] Anonymous. Forum. IEEE Spectrum, 10(1):12–18, January 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Fb

[Ano73-256] Anonymous. Forum. IEEE Spectrum, 10(2):16–18, February 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Fd

[Ano73-257] Anonymous. Forum. IEEE Spectrum, 10(3):14–18, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Fe

[Ano73-258] Anonymous. Forum. IEEE Spectrum, 10(4):14–16, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ff

[Ano73-259] Anonymous. Forum. IEEE Spectrum, 10(5):11–14, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Fg

[Ano73-260] Anonymous. Forum. IEEE Spectrum, 10(6):12–16, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Fh

[Ano73-261] Anonymous. Forum. IEEE Spectrum, 10(7):10–12, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Fi

[Ano73-262] Anonymous. Forum. IEEE Spectrum, 10(8):18–22, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Fj

[Ano73-263] Anonymous. Forum. IEEE Spectrum, 10(9):26–30, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Fk

[Ano73-264] Anonymous. Forum. IEEE Spectrum, 10(10):19–20, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Fl

[Ano73-265] Anonymous. Forum. IEEE Spectrum, 10(11):19–24, November 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Fm

[Ano73-266] Anonymous. Forum. IEEE Spectrum, 10(12):20–22, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FCa

[Ano73-267] Anonymous. [front cover].

IEEE Spectrum, 10(1):c1, January 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FCb

[Ano73-268] Anonymous. [front cover].

IEEE Spectrum, 10(1):c2, January 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FCc

[Ano73-269] Anonymous. [front cover]. IEEE Spectrum, 10(2):c1, February 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FCd

[Ano73-270] Anonymous. [front cover]. IEEE Spectrum, 10(2):c2, February 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FCe

[Ano73-271] Anonymous. [front cover]. IEEE Spectrum, 10(3):c1, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FCf

[Ano73-272] Anonymous. [front cover]. IEEE Spectrum, 10(3):c2, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FCg

[Ano73-273] Anonymous. [front cover]. IEEE Spectrum, 10(4):c1, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FCh

[Ano73-274] Anonymous. [front cover]. IEEE Spectrum, 10(4):c2, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FCi

[Ano73-275] Anonymous. [front cover]. IEEE Spectrum, 10(5):c1, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FCj

[Ano73-276] Anonymous. [front cover]. IEEE Spectrum, 10(5):c2, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FCk

[Ano73-277] Anonymous. [front cover]. IEEE Spectrum, 10(6):c1, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FCl

[Ano73-278] Anonymous. [front cover]. IEEE Spectrum, 10(6):c2, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FCm

[Ano73-279] Anonymous. [front cover]. IEEE Spectrum, 10(7):c1, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FCn

[Ano73-280] Anonymous. [front cover]. IEEE Spectrum, 10(7):c2, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FCo

[Ano73-281] Anonymous. [front cover]. IEEE Spectrum, 10(8):c1, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FCp

[Ano73-282] Anonymous. [front cover]. IEEE Spectrum, 10(8):c2, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FCq

[Ano73-283] Anonymous. [front cover]. IEEE Spectrum, 10(9):c1, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FCr

[Ano73-284] Anonymous. [front cover]. IEEE Spectrum, 10(9):c2, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FCs

[Ano73-285] Anonymous. [front cover]. IEEE Spectrum, 10(10):c1, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FCt

[Ano73-286] Anonymous. [front cover]. IEEE Spectrum, 10(10):c2, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FCu

[Ano73-287] Anonymous. [front cover]. IEEE Spectrum, 10(11):c1, November 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FCv

[Ano73-288] Anonymous. [front cover]. IEEE Spectrum, 10(11):c2, November 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FCw

[Ano73-289] Anonymous. [front cover]. IEEE Spectrum, 10(12):c1, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FCx

[Ano73-290] Anonymous. [front cover]. IEEE Spectrum, 10(12):c2, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FSIa

[Ano73-291] Anonymous. Future special issues. IEEE Spectrum, 10(1): 92–99, January 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FSIb

[Ano73-292] Anonymous. Future special issues. IEEE Spectrum, 10(2): 78–85, February 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FSIc

[Ano73-293] Anonymous. Future special issues. IEEE Spectrum, 10(3):91, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FSId

[Ano73-294] Anonymous. Future special issues. IEEE Spectrum, 10 (4):79–82, April 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FSIe

[Ano73-295] Anonymous. Future special issues. IEEE Spectrum, 10 (5):76–80, May 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FSIf

[Ano73-296] Anonymous. Future special issues. IEEE Spectrum, 10(6):78, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FSIg

[Ano73-297] Anonymous. Future special issues. IEEE Spectrum, 10(7):81, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FSIh

[Ano73-298] Anonymous. Future special issues. IEEE Spectrum, 10(8):88, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FSIi

[Ano73-299] Anonymous. Future special issues. IEEE Spectrum, 10(9): 105, September 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FSIj

[Ano73-300] Anonymous. Future special issues. IEEE Spectrum, 10 (10):89, October 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FSIk

[Ano73-301] Anonymous. Future special issues. IEEE Spectrum, 10(11): 105, November 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:FSIl

[Ano73-302] Anonymous. Future special issues. IEEE Spectrum, 10(12): 106, December 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:HSI

[Ano73-303] Anonymous. How to see IEEE INTERCON ’73. IEEE Spectrum, 10(3):a3–a4, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IDA

[Ano73-304] Anonymous. IEEE — the decade ahead. IEEE Spectrum, 10(2): 62–67, February 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IGA

[Ano73-305] Anonymous. IEEE guide for authors. IEEE Spectrum, 10 (1):89–90, January 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IIU

[Ano73-306] Anonymous. IEEE INTERCON ’73: a user-oriented program. IEEE Spectrum, 10(3):a7–a14, March 1973. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IPOa

[Ano73-307] Anonymous. IEEE publications operations. IEEE Spectrum, 10(1):4, January 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IPOb

[Ano73-308] Anonymous. IEEE publications operations. IEEE Spectrum, 10(2):4, February 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IPOc

[Ano73-309] Anonymous. IEEE publications operations. IEEE Spectrum, 10(3):4, March 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IPOd

[Ano73-310] Anonymous. IEEE publications operations. IEEE Spectrum, 10(4):4, April 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IPOe

[Ano73-311] Anonymous. IEEE publications operations. IEEE Spectrum, 10(5):4, May 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IPOf

[Ano73-312] Anonymous. IEEE publications operations. IEEE Spectrum, 10(6):4, June 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IPOg

[Ano73-313] Anonymous. IEEE publications operations. IEEE Spectrum, 10(7):4, July 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IPOh

[Ano73-314] Anonymous. IEEE publications operations. IEEE Spectrum, 10(8):4, August 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IPOi

[Ano73-315] Anonymous. IEEE publications operations. IEEE Spectrum, 10 (9):4, September 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IPOj

[Ano73-316] Anonymous. IEEE publications operations. IEEE Spectrum, 10(10):4, October 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IPOk

[Ano73-317] Anonymous. IEEE publications operations. IEEE Spectrum, 10 (11):4, November 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IPOl

[Ano73-318] Anonymous. IEEE publications operations. IEEE Spectrum, 10 (12):4, December 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ISa

[Ano73-319] Anonymous. IEEE Standards. IEEE Spectrum, 10(1): 99–100, January 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ISb

[Ano73-320] Anonymous. IEEE Standards. IEEE Spectrum, 10(2): 85–86, February 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ISc

[Ano73-321] Anonymous. IEEE Standards. IEEE Spectrum, 10(3):92, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ISd

[Ano73-322] Anonymous. IEEE Standards. IEEE Spectrum, 10(4):83, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ISe

[Ano73-323] Anonymous. IEEE Standards. IEEE Spectrum, 10(5):81, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ISf

[Ano73-324] Anonymous. IEEE Standards. IEEE Spectrum, 10(6):80, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ISg

[Ano73-325] Anonymous. IEEE Standards.

IEEE Spectrum, 10(7):82, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ISh

[Ano73-326] Anonymous. IEEE Standards.

IEEE Spectrum, 10(8):89, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ISi

[Ano73-327] Anonymous. IEEE Standards. IEEE Spectrum, 10(9): 106, September 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ISj

[Ano73-328] Anonymous. IEEE Standards.

IEEE Spectrum, 10(10):90, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ISk

[Ano73-329] Anonymous. IEEE Standards. IEEE Spectrum, 10 (11):106, November 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ISl

[Ano73-330] Anonymous. IEEE Standards.

IEEE Spectrum, 10(12):107, December 1973. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ITCa

[Ano73-331] Anonymous. IEEE tables of contents. IEEE Spectrum, 10 (6):74–77, June 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ITCb

[Ano73-332] Anonymous. IEEE tables of contents. IEEE Spectrum, 10(7):76– 80, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ITCc

[Ano73-333] Anonymous. IEEE tables of contents. IEEE Spectrum, 10 (8):78–87, August 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ITCd

[Ano73-334] Anonymous. IEEE tables of contents. IEEE Spectrum, 10(9): 92–105, September 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ITCe

[Ano73-335] Anonymous. IEEE tables of contents. IEEE Spectrum, 10 (10):78–88, October 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ITCf

[Ano73-336] Anonymous. IEEE tables of contents. IEEE Spectrum, 10(11):

91–103, November 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:ITCg

[Ano73-337] Anonymous. IEEE tables of contents. IEEE Spectrum, 10(12): 100–106, December 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IAa

[Ano73-338] Anonymous. Index to advertisers. IEEE Spectrum, 10 (1):120, January 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IAb

[Ano73-339] Anonymous. Index to advertisers. IEEE Spectrum, 10 (2):104, February 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IAc

[Ano73-340] Anonymous. Index to advertisers. IEEE Spectrum, 10(3): 109–112, March 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IAd

[Ano73-341] Anonymous. Index to advertisers. IEEE Spectrum, 10(4):96, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IAe

[Ano73-342] Anonymous. Index to advertisers. IEEE Spectrum, 10(5):96, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IAf

[Ano73-343] Anonymous. Index to advertisers. IEEE Spectrum, 10(6):96, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IAg

[Ano73-344] Anonymous. Index to advertisers. IEEE Spectrum, 10(7):100, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IAh

[Ano73-345] Anonymous. Index to advertisers. IEEE Spectrum, 10(8):96, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IAi

[Ano73-346] Anonymous. Index to advertisers. IEEE Spectrum, 10(9): 111–116, September 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IAj

[Ano73-347] Anonymous. Index to advertisers. IEEE Spectrum, 10 (10):102, October 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IAk

[Ano73-348] Anonymous. Index to advertisers. IEEE Spectrum, 10(11):

116, November 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IAl

[Ano73-349] Anonymous. Index to advertisers. IEEE Spectrum, 10 (12):136, December 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IASa

[Ano73-350] Anonymous. Information about the society. IEEE Spectrum, 10 (1):9–10, January 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IASb

[Ano73-351] Anonymous. Information about the society. IEEE Spectrum, 10 (1):121–122, January 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IASc

[Ano73-352] Anonymous. Information about the society. IEEE Spectrum, 10 (2):105–106, February 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IASd

[Ano73-353] Anonymous. Information about the society. IEEE Spectrum, 10 (2):107–108, February 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IIa

[Ano73-354] Anonymous. Inside IEEE. IEEE Spectrum, 10(2):7–14, February 1973. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IIb

[Ano73-355] Anonymous. Inside IEEE. IEEE Spectrum, 10(3):7–10, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IIc

[Ano73-356] Anonymous. Inside IEEE. IEEE Spectrum, 10(4):7–11, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IId

[Ano73-357] Anonymous. Inside IEEE. IEEE Spectrum, 10(5):7–10, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IIe

[Ano73-358] Anonymous. Inside IEEE. IEEE Spectrum, 10(6):7–9, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IIf

[Ano73-359] Anonymous. Inside IEEE. IEEE Spectrum, 10(7):7–9, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IIg

[Ano73-360] Anonymous. Inside IEEE. IEEE Spectrum, 10(8):48–93, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IIh

[Ano73-361] Anonymous. Inside IEEE. IEEE Spectrum, 10(9):70–71, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IIi

[Ano73-362] Anonymous. Inside IEEE. IEEE Spectrum, 10(10):52–53, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IIj

[Ano73-363] Anonymous. Inside IEEE.

IEEE Spectrum, 10(11):58–114, November 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IIk

[Ano73-364] Anonymous. Inside IEEE. IEEE Spectrum, 10(12):66–114, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IEEa

[Ano73-365] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 10(1):6, January 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IEEb

[Ano73-366] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 10(2):6, February 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IEEc

[Ano73-367] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 10(3):6, March 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IEEd

[Ano73-368] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 10(4):6, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IEEe

[Ano73-369] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 10(5):6, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IEEf

[Ano73-370] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 10(6):6, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IEEg

[Ano73-371] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 10(7):6, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IEEh

[Ano73-372] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 10(8):5, August 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IEEi

[Ano73-373] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 10 (9):5, September 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IEEj

[Ano73-374] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 10(10):5, October 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IEEk

[Ano73-375] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 10 (11):6, November 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IEEl

[Ano73-376] Anonymous. The Institute of Electrical and Electronics Engineers, Inc. IEEE Spectrum, 10 (12):5, December 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:IWB

[Ano73-377] Anonymous. Intercon week to bring 30,000 to ‘world’s capital’ of electronics. IEEE Spectrum, 10(3):a45–a48, March 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:LTA

[Ano73-378] Anonymous. … and that’s not all. IEEE Spectrum, 10(3):31, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ma

[Ano73-379] Anonymous. Meetings. IEEE Spectrum, 10(1):110–115, January 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Mb

[Ano73-380] Anonymous. Meetings. IEEE Spectrum, 10(2):94–97, February 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Mc

[Ano73-381] Anonymous. Meetings. IEEE Spectrum, 10(3):102–107, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Md

[Ano73-382] Anonymous. Meetings. IEEE Spectrum, 10(4):88–90, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Me

[Ano73-383] Anonymous. Meetings. IEEE Spectrum, 10(5):87–90, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Mf

[Ano73-384] Anonymous. Meetings. IEEE Spectrum, 10(6):86–90, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Mg

[Ano73-385] Anonymous. Meetings. IEEE Spectrum, 10(7):89–94, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Mh

[Ano73-386] Anonymous. Meetings. IEEE Spectrum, 10(8):6–8, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Mi

[Ano73-387] Anonymous. Meetings. IEEE Spectrum, 10(9):8–13, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Mj

[Ano73-388] Anonymous. Meetings. IEEE Spectrum, 10(10):8–9, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Mk

[Ano73-389] Anonymous. Meetings. IEEE Spectrum, 10(11):8–11, November 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ml

[Ano73-390] Anonymous. Meetings. IEEE Spectrum, 10(12):8–10, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:MAa

[Ano73-391] Anonymous. [miscellaneous advertisements]. IEEE Spectrum, 10(9):93–104, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NPAa

[Ano73-392] Anonymous. New product applications. IEEE Spectrum, 10 (1):79–84, January 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NPAb

[Ano73-393] Anonymous. New product applications. IEEE Spectrum, 10 (2):68–73, February 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NPAc

[Ano73-394] Anonymous. New product applications. IEEE Spectrum, 10 (3):80–85, March 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NPAd

[Ano73-395] Anonymous. New product applications. IEEE Spectrum, 10 (4):68–73, April 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NPAe

[Ano73-396] Anonymous. New product applications. IEEE Spectrum, 10 (5):66–71, May 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NPAf

[Ano73-397] Anonymous. New product applications. IEEE Spectrum, 10 (6):65–70, June 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NPAg

[Ano73-398] Anonymous. New product applications. IEEE Spectrum, 10 (7):67–72, July 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NPAh

[Ano73-399] Anonymous. New product applications. IEEE Spectrum, 10 (8):69–74, August 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NPAi

[Ano73-400] Anonymous. New product applications. IEEE Spectrum, 10 (9):83–88, September 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NPAj

[Ano73-401] Anonymous. New product applications. IEEE Spectrum, 10 (10):69–74, October 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NPAk

[Ano73-402] Anonymous. New product applications. IEEE Spectrum, 10 (11):82–87, November 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NPAl

[Ano73-403] Anonymous. New product applications. IEEE Spectrum, 10 (12):91–96, December 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NIa

[Ano73-404] Anonymous. News from industry. IEEE Spectrum, 10(6):79, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NIb

[Ano73-405] Anonymous. News from industry. IEEE Spectrum, 10(7):87, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NIc

[Ano73-406] Anonymous. News from industry. IEEE Spectrum, 10(8):83, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NId

[Ano73-407] Anonymous. News from industry. IEEE Spectrum, 10 (9):99, September 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NIe

[Ano73-408] Anonymous. News from industry. IEEE Spectrum, 10 (10):85, October 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NIf

[Ano73-409] Anonymous. News from industry. IEEE Spectrum, 10 (11):99, November 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NIg

[Ano73-410] Anonymous. News from industry. IEEE Spectrum, 10 (12):119, December 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NRb

[Ano73-411] Anonymous. News from region 10. IEEE Spectrum, 10(8):84, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NRd

[Ano73-412] Anonymous. News from region 2.

IEEE Spectrum, 10(10):86, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NRa

[Ano73-413] Anonymous. News from region 3. IEEE Spectrum, 10(7):88, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NRc

[Ano73-414] Anonymous. News from region 4. IEEE Spectrum, 10 (9):100, September 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NRf

[Ano73-415] Anonymous. News from region 5. IEEE Spectrum, 10(12):120, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NRe

[Ano73-416] Anonymous. News from region 6. IEEE Spectrum, 10 (11):100, November 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NWa

[Ano73-417] Anonymous. News from Washington. IEEE Spectrum, 10 (1):19–20, January 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NWb

[Ano73-418] Anonymous. News from Washington. IEEE Spectrum, 10(2): 21–22, February 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NWc

[Ano73-419] Anonymous. News from Washington. IEEE Spectrum, 10 (3):21–22, March 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NWd

[Ano73-420] Anonymous. News from Washington. IEEE Spectrum, 10 (4):19–20, April 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NWe

[Ano73-421] Anonymous. News from Washington. IEEE Spectrum, 10 (5):15–16, May 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NWf

[Ano73-422] Anonymous. News from Washington. IEEE Spectrum, 10 (6):17–18, June 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NWg

[Ano73-423] Anonymous. News from Washington. IEEE Spectrum, 10 (7):13–14, July 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NWh

[Ano73-424] Anonymous. News from Washington. IEEE Spectrum, 10 (8):13–14, August 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NWi

[Ano73-425] Anonymous. News from Washington. IEEE Spectrum, 10(9): 33–34, September 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NWj

[Ano73-426] Anonymous. News from Washington. IEEE Spectrum, 10 (10):17–18, October 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NWk

[Ano73-427] Anonymous. News from Washington. IEEE Spectrum, 10(11): 17–18, November 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:NWl

[Ano73-428] Anonymous. News from Washington. IEEE Spectrum, 10(12): 17–18, December 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:O

[Ano73-429] Anonymous. Obituaries. IEEE Spectrum, 10(1):119, January 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Pa

[Ano73-430] Anonymous. People. IEEE Spectrum, 10(1):118–119, January 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Pb

[Ano73-431] Anonymous. People. IEEE Spectrum, 10(2):101–103, February 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Pc

[Ano73-432] Anonymous. People. IEEE Spectrum, 10(3):110–111, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Pd

[Ano73-433] Anonymous. People. IEEE Spectrum, 10(4):94–95, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Pe

[Ano73-434] Anonymous. People. IEEE Spectrum, 10(5):93–95, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Pf

[Ano73-435] Anonymous. People. IEEE Spectrum, 10(6):93–95, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Pg

[Ano73-436] Anonymous. People. IEEE Spectrum, 10(7):98–99, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Ph

[Ano73-437] Anonymous. People. IEEE Spectrum, 10(8):94–95, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Pi

[Ano73-438] Anonymous. People. IEEE Spectrum, 10(9):112–113, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Pj

[Ano73-439] Anonymous. People. IEEE Spectrum, 10(10):100–101, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Pk

[Ano73-440] Anonymous. People. IEEE Spectrum, 10(11):112–113, November 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:Pm

[Ano73-441] Anonymous. People. IEEE Spectrum, 10(12):116–118, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:PII

[Ano73-442] Anonymous. Products at IEEE INTERCON. IEEE Spectrum, 10 (3):a34–a42, March 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:PCSa

[Ano73-443] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 10(1):117, January 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:PCSb

[Ano73-444] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 10(2):99, February 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:PCSc

[Ano73-445] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 10(3):103–107, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:PCSd

[Ano73-446] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 10(4):93, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:PCSe

[Ano73-447] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 10(5):92, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:PCSf

[Ano73-448] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 10(6):92, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:PCSg

[Ano73-449] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 10(7):91–95, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:PCSh

[Ano73-450] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 10(8):97–98, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:PCSi

[Ano73-451] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 10(10):97– 99, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:PCSj

[Ano73-452] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 10(11):115, November 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:PCSk

[Ano73-453] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 10(12):137– 138, December 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:PL

[Ano73-454] Anonymous. Program listing.

IEEE Spectrum, 10(3):a16–a24, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SY

[Ano73-455] Anonymous. Salaries are up:

is yours? IEEE Spectrum, 10 (10):48–51, October 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SIa

[Ano73-456] Anonymous. Scanning the issues. IEEE Spectrum, 10(1): 86–87, January 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SIb

[Ano73-457] Anonymous. Scanning the issues. IEEE Spectrum, 10(2): 75–76, February 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SIc

[Ano73-458] Anonymous. Scanning the issues.

IEEE Spectrum, 10(3):87–88, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SId

[Ano73-459] Anonymous. Scanning the issues. IEEE Spectrum, 10(4):75–76, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SIe

[Ano73-460] Anonymous. Scanning the issues. IEEE Spectrum, 10(5):73–74, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SIf

[Ano73-461] Anonymous. Scanning the issues. IEEE Spectrum, 10(6):72–73, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SIg

[Ano73-462] Anonymous. Scanning the issues. IEEE Spectrum, 10(7):74– 75, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SIh

[Ano73-463] Anonymous. Scanning the issues. IEEE Spectrum, 10(8): 76–77, August 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SIi

[Ano73-464] Anonymous. Scanning the issues. IEEE Spectrum, 10(9): 90–91, September 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SIj

[Ano73-465] Anonymous. Scanning the issues. IEEE Spectrum, 10(10): 76–77, October 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SIk

[Ano73-466] Anonymous. Scanning the issues. IEEE Spectrum, 10(11): 89–90, November 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SIl

[Ano73-467] Anonymous. Scanning the issues. IEEE Spectrum, 10(12): 98–99, December 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SSAa

[Ano73-468] Anonymous. Science/scope [advertisement]. IEEE Spectrum, 10 (2):nil1–1, February 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SSAb

[Ano73-469] Anonymous. Science/scope [advertisement]. IEEE Spectrum, 10(4):nil4, April 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SSAc

[Ano73-470] Anonymous. Science/scope [advertisement]. IEEE Spectrum, 10 (8):nil1–1, August 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SSAd

[Ano73-471] Anonymous. Science/scope [advertisement]. IEEE Spectrum, 10(10):93, October 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SSAe

[Ano73-472] Anonymous. Science/scope [advertisement]. IEEE Spectrum, 10 (12):9–11, December 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SSS

[Ano73-473] Anonymous. Solid state shapes the future. IEEE Spectrum, 10 (2):52–53, February 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SPa

[Ano73-474] Anonymous. Special publications. IEEE Spectrum, 10(1): 100–101, January 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SPb

[Ano73-475] Anonymous. Special publications. IEEE Spectrum, 10(2): 86–87, February 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SPc

[Ano73-476] Anonymous. Special publications. IEEE Spectrum, 10 (3):92–93, March 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SPd

[Ano73-477] Anonymous. Special publications. IEEE Spectrum, 10 (4):83–84, April 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SPe

[Ano73-478] Anonymous. Special publications. IEEE Spectrum, 10 (5):81–82, May 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SPf

[Ano73-479] Anonymous. Special publications. IEEE Spectrum, 10 (6):80–81, June 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SPg

[Ano73-480] Anonymous. Special publications. IEEE Spectrum, 10(7):82– 83, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SPh

[Ano73-481] Anonymous. Special publications. IEEE Spectrum, 10 (8):89–90, August 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SPi

[Ano73-482] Anonymous. Special publications. IEEE Spectrum, 10(9): 106–108, September 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SPj

[Ano73-483] Anonymous. Special publications. IEEE Spectrum, 10(10): 90–92, October 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SPk

[Ano73-484] Anonymous. Special publications. IEEE Spectrum, 10(11): 106–108, November 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SPl

[Ano73-485] Anonymous. Special publications. IEEE Spectrum, 10(12): 107–108, December 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SLa

[Ano73-486] Anonymous. Spectral lines. IEEE Spectrum, 10(1):23, January 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SLb

[Ano73-487] Anonymous. Spectral lines. IEEE Spectrum, 10(2):25, February 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SLc

[Ano73-488] Anonymous. Spectral lines. IEEE Spectrum, 10(3):25, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SLd

[Ano73-489] Anonymous. Spectral lines.

IEEE Spectrum, 10(4):23, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SLe

[Ano73-490] Anonymous. Spectral lines.

IEEE Spectrum, 10(5):17, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SLf

[Ano73-491] Anonymous. Spectral lines.

IEEE Spectrum, 10(6):21, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SLg

[Ano73-492] Anonymous. Spectral lines.

IEEE Spectrum, 10(7):17, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SLh

[Ano73-493] Anonymous. Spectral lines.

IEEE Spectrum, 10(8):23, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SLi

[Ano73-494] Anonymous. Spectral lines.

IEEE Spectrum, 10(9):35, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SLj

[Ano73-495] Anonymous. Spectral lines. IEEE Spectrum, 10(10):21, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SLk

[Ano73-496] Anonymous. Spectral lines. IEEE Spectrum, 10(11):27, November 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SLl

[Ano73-497] Anonymous. Spectral lines. IEEE Spectrum, 10(12):25, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:SICa

[Ano73-498] Anonymous. Statements from IEEE candidates. IEEE Spectrum, 10(8):60–68, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:TCa

[Ano73-499] Anonymous. Table of contents. IEEE Spectrum, 10(1):2–3, January 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:TCb

[Ano73-500] Anonymous. Table of contents. IEEE Spectrum, 10 (2):2–3, February 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:TCc

[Ano73-501] Anonymous. Table of contents. IEEE Spectrum, 10(3):2–3, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:TCd

[Ano73-502] Anonymous. Table of contents. IEEE Spectrum, 10(4):2–3, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:TCf

[Ano73-503] Anonymous. Table of contents. IEEE Spectrum, 10(5):2– 3, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:TCh

[Ano73-504] Anonymous. Table of contents. IEEE Spectrum, 10(6):2– 3, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:TCi

[Ano73-505] Anonymous. Table of contents. IEEE Spectrum, 10(7):2– 3, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:TCj

[Ano73-506] Anonymous. Table of contents.

IEEE Spectrum, 10(8):2–3, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:TCk

[Ano73-507] Anonymous. Table of contents. IEEE Spectrum, 10(9): 2–3, September 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:TCl

[Ano73-508] Anonymous. Table of contents. IEEE Spectrum, 10(10):2–3, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:TCm

[Ano73-509] Anonymous. Table of contents. IEEE Spectrum, 10(11): 2–3, November 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:TCn

[Ano73-510] Anonymous. Table of contents. IEEE Spectrum, 10(12):2–3, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:TCe

[Ano73-511] Anonymous. Tables of contents. IEEE Spectrum, 10 (4):77–79, April 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1973:TCg

[Ano73-512] Anonymous. Tables of contents. IEEE Spectrum, 10 (5):75–76, May 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Ac

[Ano74a] Anonymous. [advertisement].

IEEE Spectrum, 11(3):1, March 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Ad

[Ano74b] Anonymous. [advertisement].

IEEE Spectrum, 11(3):4–9, March 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Ae

[Ano74c] Anonymous. [advertisement].

IEEE Spectrum, 11(3):1–7, March 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Af

[Ano74d] Anonymous. [advertisement].

IEEE Spectrum, 11(4):1, April 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Ag

[Ano74e] Anonymous. [advertisement].

IEEE Spectrum, 11(4):4–9, April 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Ah

[Ano74f] Anonymous. [advertisement].

IEEE Spectrum, 11(4):1–7, April 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Ai

[Ano74g] Anonymous. [advertisement]. IEEE Spectrum, 11(5):1, May 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Aj

[Ano74h] Anonymous. [advertisement]. IEEE Spectrum, 11(5):4–9, May 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Am

[Ano74i] Anonymous. [advertisement]. IEEE Spectrum, 11(6):1, June 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:An

[Ano74j] Anonymous. [advertisement]. IEEE Spectrum, 11(6):4–9, June 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Ao

[Ano74k] Anonymous. [advertisement]. IEEE Spectrum, 11(6):1–7, June 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Ap

[Ano74l] Anonymous. [advertisement]. IEEE Spectrum, 11(7):1, July 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Aq

[Ano74m] Anonymous. [advertisement].

IEEE Spectrum, 11(7):4–9, July 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Ar

[Ano74n] Anonymous. [advertisement].

IEEE Spectrum, 11(7):1–2, July 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:As

[Ano74o] Anonymous. [advertisement].

IEEE Spectrum, 11(8):1, August 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:At

[Ano74p] Anonymous. [advertisement].

IEEE Spectrum, 11(8):4–7, August 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Au

[Ano74q] Anonymous. [advertisement].

IEEE Spectrum, 11(8):1–7, August 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Av

[Ano74r] Anonymous. [advertisement].

IEEE Spectrum, 11(9):1–2, September 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Aw

[Ano74s] Anonymous. [advertisement]. IEEE Spectrum, 11(9):1, September 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Ax

[Ano74t] Anonymous. [advertisement]. IEEE Spectrum, 11(9):4–9, September 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Ay

[Ano74u] Anonymous. [advertisement]. IEEE Spectrum, 11(9):1–2, September 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Az

[Ano74v] Anonymous. [advertisement]. IEEE Spectrum, 11(10):1–2, October 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Aba

[Ano74w] Anonymous. [advertisement]. IEEE Spectrum, 11(10):1, October 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Abb

[Ano74x] Anonymous. [advertisement]. IEEE Spectrum, 11(10):4–9, October 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Abc

[Ano74y] Anonymous. [advertisement].

IEEE Spectrum, 11(10):1–4, October 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Abd

[Ano74z] Anonymous. [advertisement].

IEEE Spectrum, 11(11):1–2, November 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Abe

[Ano74-27] Anonymous. [advertisement].

IEEE Spectrum, 11(11):1, November 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Abf

[Ano74-28] Anonymous. [advertisement].

IEEE Spectrum, 11(11):4–7, November 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Abg

[Ano74-29] Anonymous. [advertisement].

IEEE Spectrum, 11(11):1–4, November 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Abh

[Ano74-30] Anonymous. [advertisement].

IEEE Spectrum, 11(12):1, December 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Abi

[Ano74-31] Anonymous. [advertisement]. IEEE Spectrum, 11(12):4–7, December 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Abj

[Ano74-32] Anonymous. [advertisement]. IEEE Spectrum, 11(12):1–4, December 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Aa

[Ano74-33] Anonymous. [advertisements]. IEEE Spectrum, 11(1):1, January 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Ab

[Ano74-34] Anonymous. [advertisements]. IEEE Spectrum, 11(1):4–7, January 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Ak

[Ano74-35] Anonymous. [advertisements]. IEEE Spectrum, 11(5):1–7, May 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ALa

[Ano74-36] Anonymous. Applications literature. IEEE Spectrum, 11 (1):108, January 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ALb

[Ano74-37] Anonymous. Applications literature. IEEE Spectrum, 11 (2):102, February 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ALc

[Ano74-38] Anonymous. Applications literature. IEEE Spectrum, 11 (12):77, December 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ALBb

[Ano74-39] Anonymous. Applications literature: Basic characteristics of precision frequency sources. IEEE Spectrum, 11(8):95, August 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ALBa

[Ano74-40] Anonymous. Applications literature: Basic introduction to field effect transistors. IEEE Spectrum, 11(5):90, May 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ALD

[Ano74-41] Anonymous. Applications literature: Design fundamentals of programmable power supplies. IEEE Spectrum, 11(11): 94–95, November 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ALE

[Ano74-42] Anonymous. Applications literature: Elementary principles of data acquisition and conversion. IEEE Spectrum, 11 (9):99, September 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ALMb

[Ano74-43] Anonymous. Applications literature: Measuring power in the frequency domain. IEEE Spectrum, 11(6):98–98D, June 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ALMa

[Ano74-44] Anonymous. Applications literature: Microwave power transistor design and circuit applications. IEEE Spectrum, 11(3):105, March 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ALP

[Ano74-45] Anonymous. Applications literature: Principles and applications off molded case circuit breakers. IEEE Spectrum, 11(7):89– 90, July 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ALTb

[Ano74-46] Anonymous. Applications literature: Theory and application of a retriggerable one-shot. IEEE Spectrum, 11(10):104, October 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ALTa

[Ano74-47] Anonymous. Applications literature: Theory and application off solid-state oscillators. IEEE Spectrum, 11(4):89, April 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:BC

[Ano74-48] Anonymous. [back cover]. IEEE Spectrum, 11(1):c4, January 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:BRa

[Ano74-49] Anonymous. Book reviews.

IEEE Spectrum, 11(2):118– 121, February 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:BRb

[Ano74-50] Anonymous. Book reviews.

IEEE Spectrum, 11(3):124–125, March 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:BRc

[Ano74-51] Anonymous. Book reviews.

IEEE Spectrum, 11(4):104–106, April 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:BRd

[Ano74-52] Anonymous. Book reviews.

IEEE Spectrum, 11(5):112–117, May 1974. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:BRe

[Ano74-53] Anonymous. Book reviews. IEEE Spectrum, 11(6):117–120, June 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:BRf

[Ano74-54] Anonymous. Book reviews. IEEE Spectrum, 11(7):104–105, July 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:BRg

[Ano74-55] Anonymous. Book reviews. IEEE Spectrum, 11(8):111–113, August 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:BRh

[Ano74-56] Anonymous. Book reviews. IEEE Spectrum, 11(9):120–123, September 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:BRi

[Ano74-57] Anonymous. Book reviews. IEEE Spectrum, 11(10):125– 127, October 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:BRj

[Ano74-58] Anonymous. Book reviews. IEEE Spectrum, 11(11):118– 121, November 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:BRk

[Ano74-59] Anonymous. Book reviews.

IEEE Spectrum, 11(12):92–93, December 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ISAb

[Ano74-60] Anonymous. IEEE Spectrum annual index. IEEE Spectrum, 11(12):101–114, December 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ISC

[Ano74-61] Anonymous. IEEE Spectrum’s computer milestones. IEEE Spectrum, 11(2):80–81, February 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SHRa

[Ano74-62] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 11(1):107, January 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SHRb

[Ano74-63] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 11(2):101, February 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SHRc

[Ano74-64] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 11(3):104, March 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SHRd

[Ano74-65] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 11(4):86, April 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SHRe

[Ano74-66] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 11(5):88–89, May 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SHRf

[Ano74-67] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 11(6):97, June 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SHRg

[Ano74-68] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 11(8):94, August 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SHRh

[Ano74-69] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 11(9):98, September 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SHRi

[Ano74-70] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 11(10):103, October 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SHRj

[Ano74-71] Anonymous. Spectrum’s hardware review: For more information on the following new products, circle numbers on the reader service card corresponding to bracket numbers. IEEE Spectrum, 11(11):93, November 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Cb

[Ano74-72] Anonymous. Calendar. IEEE Spectrum, 11(1):14–16, January 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Cc

[Ano74-73] Anonymous. Calendar. IEEE Spectrum, 11(2):16, February 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Ce

[Ano74-74] Anonymous. Calendar. IEEE Spectrum, 11(3):20–25, March 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Cg

[Ano74-75] Anonymous. Calendar. IEEE Spectrum, 11(4):17–20, April 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Ci

[Ano74-76] Anonymous. Calendar. IEEE Spectrum, 11(5):18–22, May 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Ck

[Ano74-77] Anonymous. Calendar. IEEE Spectrum, 11(6):22–26, June 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Cm

[Ano74-78] Anonymous. Calendar. IEEE Spectrum, 11(7):15–16, July 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Co

[Ano74-79] Anonymous. Calendar. IEEE Spectrum, 11(8):20–23, August 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Cq

[Ano74-80] Anonymous. Calendar. IEEE Spectrum, 11(9):18–24, September 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Cs

[Ano74-81] Anonymous. Calendar. IEEE Spectrum, 11(10):20–23, October 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Cu

[Ano74-82] Anonymous. Calendar. IEEE Spectrum, 11(11):18–21, November 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Cw

[Ano74-83] Anonymous. Calendar. IEEE Spectrum, 11(12):14–16, December 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Ca

[Ano74-84] Anonymous. Contents. IEEE Spectrum, 11(1):2–3, January 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Cd

[Ano74-85] Anonymous. Contents. IEEE Spectrum, 11(3):2–3, March 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Cf

[Ano74-86] Anonymous. Contents. IEEE Spectrum, 11(4):2–3, April 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Ch

[Ano74-87] Anonymous. Contents. IEEE Spectrum, 11(5):2–3, May 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Cj

[Ano74-88] Anonymous. Contents. IEEE Spectrum, 11(6):2–3, June 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Cl

[Ano74-89] Anonymous. Contents. IEEE Spectrum, 11(7):2–3, July 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Cn

[Ano74-90] Anonymous. Contents. IEEE Spectrum, 11(8):2–3, August 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Cp

[Ano74-91] Anonymous. Contents. IEEE Spectrum, 11(9):2–3, September 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Cr

[Ano74-92] Anonymous. Contents. IEEE Spectrum, 11(10):2–3, October 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Ct

[Ano74-93] Anonymous. Contents. IEEE Spectrum, 11(11):2–3, November 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Cv

[Ano74-94] Anonymous. Contents. IEEE Spectrum, 11(12):2–3, December 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:CFC

[Ano74-95] Anonymous. Cutting the frills and the costs: Concerns about the high cost of U.S. defense systems procurement has prompted a proposal that emphasizes unit production cost. IEEE Spectrum, 11(1):96–97, January 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:EDE

[Ano74-96] Anonymous. Education for designers: The electronicker: a passing breed? Fundamental courses are being shaped by the impact of computer-aided design and by the integrated circuit revolution. IEEE Spectrum, 11(5):64–67, May 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:EAb

[Ano74-97] Anonymous. Educational aids.

IEEE Spectrum, 11(2):117, February 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:EAc

[Ano74-98] Anonymous. Educational aids.

IEEE Spectrum, 11(3):123, March 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:EAd

[Ano74-99] Anonymous. Educational aids. IEEE Spectrum, 11(4):103, April 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:EAe

[Ano74-100] Anonymous. Educational aids. IEEE Spectrum, 11(5):110–111, May 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:EAf

[Ano74-101] Anonymous. Educational aids. IEEE Spectrum, 11(6):116, June 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:EAg

[Ano74-102] Anonymous. Educational aids. IEEE Spectrum, 11(7):103, July 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:EAh

[Ano74-103] Anonymous. Educational aids. IEEE Spectrum, 11(8):110, August 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:EAi

[Ano74-104] Anonymous. Educational aids. IEEE Spectrum, 11(9):116–119, September 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:EAj

[Ano74-105] Anonymous. Educational aids.

IEEE Spectrum, 11(10):122– 124, October 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:EAk

[Ano74-106] Anonymous. Educational aids.

IEEE Spectrum, 11(11):117, November 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:EAl

[Ano74-107] Anonymous. Educational aids.

IEEE Spectrum, 11(12):91, December 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:EAA

[Ano74-108] Anonymous. Educational aids [advertisement]. IEEE Spectrum, 11(1):123–125, January 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ERa

[Ano74-109] Anonymous. Energy report.

IEEE Spectrum, 11(1):18–19, January 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ERb

[Ano74-110] Anonymous. Energy report.

IEEE Spectrum, 11(2):14, February 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ERc

[Ano74-111] Anonymous. Energy report. IEEE Spectrum, 11(3):18–19, March 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ERd

[Ano74-112] Anonymous. Energy report. IEEE Spectrum, 11(4):22–23, April 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ERe

[Ano74-113] Anonymous. Energy report. IEEE Spectrum, 11(5):24–25, May 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ERf

[Ano74-114] Anonymous. Energy report. IEEE Spectrum, 11(6):18–21, June 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ERg

[Ano74-115] Anonymous. Energy report. IEEE Spectrum, 11(7):18–19, July 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ERh

[Ano74-116] Anonymous. Energy report. IEEE Spectrum, 11(8):18–19, August 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ERi

[Ano74-117] Anonymous. Energy report.

IEEE Spectrum, 11(9):26–27, September 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ERj

[Ano74-118] Anonymous. Energy report.

IEEE Spectrum, 11(10):14– 19, October 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ERk

[Ano74-119] Anonymous. Energy report.

IEEE Spectrum, 11(11):26–27, November 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ERl

[Ano74-120] Anonymous. Energy report.

IEEE Spectrum, 11(12):18–19, December 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FP

[Ano74-121] Anonymous. Focal points. IEEE Spectrum, 11(2):22, February 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FPRa

[Ano74-122] Anonymous. Focal points:

a review of current developments. IEEE Spectrum, 11 (1):20–22, January 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FPRb

[Ano74-123] Anonymous. Focal points: a review of current developments. IEEE Spectrum, 11(3):26–29, March 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FPRc

[Ano74-124] Anonymous. Focal points: a review of current developments. IEEE Spectrum, 11(4):24–29, April 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FPRd

[Ano74-125] Anonymous. Focal points: a review of current developments. IEEE Spectrum, 11(5):26–28, May 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FPRe

[Ano74-126] Anonymous. Focal points: a review of current developments. IEEE Spectrum, 11(6):27–30, June 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FPRf

[Ano74-127] Anonymous. Focal points: a review of current developments. IEEE Spectrum, 11(7):20–23, July 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FPRg

[Ano74-128] Anonymous. Focal points: a review of current developments. IEEE Spectrum, 11(8):24, August 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FPRh

[Ano74-129] Anonymous. Focal points:

a review of current developments. IEEE Spectrum, 11(9): 28–31, September 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FPRi

[Ano74-130] Anonymous. Focal points:

a review of current developments. IEEE Spectrum, 11 (10):24–25, October 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FPRj

[Ano74-131] Anonymous. Focal points:

a review of current developments. IEEE Spectrum, 11(11): 22–24D, November 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FPRk

[Ano74-132] Anonymous. Focal points:

a review of current developments. IEEE Spectrum, 11(12): 20–23, December 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Fa

[Ano74-133] Anonymous. Forum. IEEE Spectrum, 11(1):23–28, January 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Fb

[Ano74-134] Anonymous. Forum. IEEE Spectrum, 11(2):26–31, February 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Fd

[Ano74-135] Anonymous. Forum. IEEE Spectrum, 11(3):30–42, March 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Fe

[Ano74-136] Anonymous. Forum. IEEE Spectrum, 11(4):30–32, April 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Ff

[Ano74-137] Anonymous. Forum. IEEE Spectrum, 11(5):29–32, May 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Fg

[Ano74-138] Anonymous. Forum. IEEE Spectrum, 11(6):31–36, June 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Fh

[Ano74-139] Anonymous. Forum. IEEE Spectrum, 11(7):24–32, July 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Fj

[Ano74-140] Anonymous. Forum. IEEE Spectrum, 11(8):25–30, August 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Fk

[Ano74-141] Anonymous. Forum. IEEE Spectrum, 11(9):32–42, September 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Fl

[Ano74-142] Anonymous. Forum. IEEE Spectrum, 11(10):34–40, October 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Fm

[Ano74-143] Anonymous. Forum. IEEE Spectrum, 11(11):34–42, November 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Fn

[Ano74-144] Anonymous. Forum. IEEE Spectrum, 11(12):28–42, December 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FC

[Ano74-145] Anonymous. [front cover]. IEEE Spectrum, 11(1):c1, January 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FSIa

[Ano74-146] Anonymous. Future special issues. IEEE Spectrum, 11 (1):121, January 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FSIb

[Ano74-147] Anonymous. Future special issues. IEEE Spectrum, 11 (2):110, February 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FSIc

[Ano74-148] Anonymous. Future special issues. IEEE Spectrum, 11(3):121, March 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FSId

[Ano74-149] Anonymous. Future special issues. IEEE Spectrum, 11 (4):99–101, April 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FSIe

[Ano74-150] Anonymous. Future special issues. IEEE Spectrum, 11 (5):106–107, May 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FSIf

[Ano74-151] Anonymous. Future special issues. IEEE Spectrum, 11(6):113, June 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FSIg

[Ano74-152] Anonymous. Future special issues. IEEE Spectrum, 11 (7):100–101, July 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FSIh

[Ano74-153] Anonymous. Future special issues. IEEE Spectrum, 11(9): 115, September 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FSIi

[Ano74-154] Anonymous. Future special issues. IEEE Spectrum, 11 (10):120, October 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FSIj

[Ano74-155] Anonymous. Future special issues. IEEE Spectrum, 11(11): 113–115, November 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FSIk

[Ano74-156] Anonymous. Future special issues. IEEE Spectrum, 11 (12):89, December 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:GA

[Ano74-157] Anonymous. Guide for authors. IEEE Spectrum, 11(1): 101–102, January 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ISb

[Ano74-158] Anonymous. IEEE standards. IEEE Spectrum, 11 (2):116, February 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ISd

[Ano74-159] Anonymous. IEEE standards. IEEE Spectrum, 11 (3):122, March 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ISe

[Ano74-160] Anonymous. IEEE standards. IEEE Spectrum, 11(4):102, April 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ISf

[Ano74-161] Anonymous. IEEE standards. IEEE Spectrum, 11(5):108, May 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ISg

[Ano74-162] Anonymous. IEEE standards. IEEE Spectrum, 11(6):114, June 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ISh

[Ano74-163] Anonymous. IEEE standards.

IEEE Spectrum, 11(7):102, July 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ISi

[Ano74-164] Anonymous. IEEE standards. IEEE Spectrum, 11(9): 116, September 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ISj

[Ano74-165] Anonymous. IEEE standards.

IEEE Spectrum, 11(10):121, October 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ISk

[Ano74-166] Anonymous. IEEE standards. IEEE Spectrum, 11 (11):116, November 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ISl

[Ano74-167] Anonymous. IEEE standards.

IEEE Spectrum, 11(12):90, December 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ISAa

[Ano74-168] Anonymous. IEEE standards [advertisement]. IEEE Spectrum, 11(1):122, January 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ITCb

[Ano74-169] Anonymous. IEEE tables of contents. IEEE Spectrum, 11(2): 106–109, February 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ITCc

[Ano74-170] Anonymous. IEEE tables of contents. IEEE Spectrum, 11 (3):108–114, March 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ITCd

[Ano74-171] Anonymous. IEEE tables of contents. IEEE Spectrum, 11 (4):92–98, April 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ITCe

[Ano74-172] Anonymous. IEEE tables of contents. IEEE Spectrum, 11 (5):93–96, May 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ITCf

[Ano74-173] Anonymous. IEEE tables of contents. IEEE Spectrum, 11 (6):101–106B, June 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ITCg

[Ano74-174] Anonymous. IEEE tables of contents. IEEE Spectrum, 11(7):95– 99, July 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ITCh

[Ano74-175] Anonymous. IEEE tables of contents. IEEE Spectrum, 11 (8):98–110, August 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ITCi

[Ano74-176] Anonymous. IEEE tables of contents. IEEE Spectrum, 11(9): 102–106, September 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ITCj

[Ano74-177] Anonymous. IEEE tables of contents. IEEE Spectrum, 11(10): 107–116, October 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ITCk

[Ano74-178] Anonymous. IEEE tables of contents. IEEE Spectrum, 11(11): 102–112, November 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ITCl

[Ano74-179] Anonymous. IEEE tables of contents. IEEE Spectrum, 11(12): 80–88, December 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:ITCa

[Ano74-180] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 11 (1):111–114, January 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FIa

[Ano74-181] Anonymous. In future issues. IEEE Spectrum, 11(1):132, January 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FIb

[Ano74-182] Anonymous. In future issues. IEEE Spectrum, 11(2): 124, February 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FIc

[Ano74-183] Anonymous. In future issues. IEEE Spectrum, 11(3):128, March 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FId

[Ano74-184] Anonymous. In future issues. IEEE Spectrum, 11(4):108, April 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FIe

[Ano74-185] Anonymous. In future issues. IEEE Spectrum, 11(5):120, May 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FIf

[Ano74-186] Anonymous. In future issues. IEEE Spectrum, 11(6):124–125, June 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FIg

[Ano74-187] Anonymous. In future issues.

IEEE Spectrum, 11(7):108–110, July 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FIh

[Ano74-188] Anonymous. In future issues.

IEEE Spectrum, 11(8):116–118, August 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FIi

[Ano74-189] Anonymous. In future issues. IEEE Spectrum, 11(9): 128–130, September 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FIj

[Ano74-190] Anonymous. In future issues.

IEEE Spectrum, 11(10):130, October 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FIk

[Ano74-191] Anonymous. In future issues. IEEE Spectrum, 11(11): 123–124, November 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:FIl

[Ano74-192] Anonymous. In future issues. IEEE Spectrum, 11(12): 115–116, December 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:IA

[Ano74-193] Anonymous. Index to advertisers. IEEE Spectrum, 11(1): 132–134, January 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:IBC

[Ano74-194] Anonymous. [inside back cover]. IEEE Spectrum, 11(1):c3, January 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:IFC

[Ano74-195] Anonymous. [inside front cover]. IEEE Spectrum, 11(1):c2, January 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:IIa

[Ano74-196] Anonymous. Inside IEEE. IEEE Spectrum, 11(1):92–95, January 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:IIb

[Ano74-197] Anonymous. Inside IEEE. IEEE Spectrum, 11(2):94–95, February 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:IIc

[Ano74-198] Anonymous. Inside IEEE. IEEE Spectrum, 11(3):74–75, March 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:IId

[Ano74-199] Anonymous. Inside IEEE. IEEE Spectrum, 11(4):80–81, April 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:IIe

[Ano74-200] Anonymous. Inside IEEE.

IEEE Spectrum, 11(5):82–83, May 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:IIf

[Ano74-201] Anonymous. Inside IEEE. IEEE Spectrum, 11(6):76–77, June 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:IIg

[Ano74-202] Anonymous. Inside IEEE.

IEEE Spectrum, 11(7):82–83, July 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:IIh

[Ano74-203] Anonymous. Inside IEEE. IEEE Spectrum, 11(8):88, August 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:IIi

[Ano74-204] Anonymous. Inside IEEE. IEEE Spectrum, 11(9):92–93, September 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:IIj

[Ano74-205] Anonymous. Inside IEEE. IEEE Spectrum, 11(10):26–33, October 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:IIk

[Ano74-206] Anonymous. Inside IEEE. IEEE Spectrum, 11(11):28–33, November 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:IIl

[Ano74-207] Anonymous. Inside IEEE. IEEE Spectrum, 11(12):24–27, December 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:INI

[Ano74-208] Anonymous. Institute news: IEEE annual reports: an overview of IEEE activities is provided in the annual reports from Institute Boards and Standing Committees. IEEE Spectrum, 11(6):78–81, June 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:INS

[Ano74-209] Anonymous. Institute news: Statements from IEEE candidates: Those running for IEEE office this fall give brief indications of their positions on vital issues. IEEE Spectrum, 11 (8):81–87, August 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Ma

[Ano74-210] Anonymous. Meetings. IEEE Spectrum, 11(1):8–13, January 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Mb

[Ano74-211] Anonymous. Meetings. IEEE Spectrum, 11(2):10–12, February 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Mc

[Ano74-212] Anonymous. Meetings. IEEE Spectrum, 11(3):10–16, March 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Md

[Ano74-213] Anonymous. Meetings. IEEE Spectrum, 11(4):10–16, April 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Me

[Ano74-214] Anonymous. Meetings. IEEE Spectrum, 11(5):10–17, May 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Mf

[Ano74-215] Anonymous. Meetings. IEEE Spectrum, 11(6):10–16C, June 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Mg

[Ano74-216] Anonymous. Meetings. IEEE Spectrum, 11(7):10–14, July 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Mh

[Ano74-217] Anonymous. Meetings. IEEE Spectrum, 11(8):10–16, August 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Mi

[Ano74-218] Anonymous. Meetings. IEEE Spectrum, 11(9):10–17, September 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Mj

[Ano74-219] Anonymous. Meetings. IEEE Spectrum, 11(10):10–12, October 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Mk

[Ano74-220] Anonymous. Meetings. IEEE Spectrum, 11(11):8–17, November 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Ml

[Ano74-221] Anonymous. Meetings. IEEE Spectrum, 11(12):8–13, December 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NPAh

[Ano74-222] Anonymous. New product applications: Closed-circuit television system displays microscope images for close-up examination. IEEE Spectrum, 11(7): 88, July 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NPAn

[Ano74-223] Anonymous. New product applications: FET-IC op amp family offers subpicoampere input bias currents. IEEE Spectrum, 11 (12):76, December 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NPAf

[Ano74-224] Anonymous. New product applications: Integrated data acquisition module mounts inside minicomputer. IEEE Spectrum, 11(6):92–97, June 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NPAm

[Ano74-225] Anonymous. New product applications: Medium-speed minicomputer features versatility at a reasonable price. IEEE Spectrum, 11(12):72–76, December 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NPAg

[Ano74-226] Anonymous. New product applications: Microcomputer product family now includes peripherals and a program development system. IEEE Spectrum, 11(7):84– 88, July 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NPAb

[Ano74-227] Anonymous. New product applications: Microprocessor module brings computer power to low-budget applications. IEEE Spectrum, 11(2):96–101, February 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NPAl

[Ano74-228] Anonymous. New product applications: Microprogrammed bipolar microcomputers are assembled from integrated circuit blocks. IEEE Spectrum, 11(11): 88–93, November 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NPAj

[Ano74-229] Anonymous. New product applications: Miniature coulometer no larger than an instrument fuse directly indicates elapsed time. IEEE Spectrum, 11(9): 94–98, September 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NPAa

[Ano74-230] Anonymous. New product applications: Minicomputer virtual memory operated by hardware look-up table. IEEE Spectrum, 11(1):103–106, January 1974. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NPAk

[Ano74-231] Anonymous. New product applications: Pattern recognition allows use of handwritten characters in data-entry system. IEEE Spectrum, 11(10):98–102, October 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NPAi

[Ano74-232] Anonymous. New product applications: Personal, scientific problem-solvers tackle inflation — with list prices under $100! IEEE Spectrum, 11(8):89–94, August 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NPAc

[Ano74-233] Anonymous. New product applications: Programmable calculators compatible with standard instruments. IEEE Spectrum, 11 (3):100–104, March 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NPAe

[Ano74-234] Anonymous. New product applications: Three new memories offer high performance for digital systems. IEEE Spectrum, 11(5):84–89, May 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NPAd

[Ano74-235] Anonymous. New product applications: Voltage to pulse rate converters assure synchronous pulse trains. IEEE Spectrum, 11(4):82–86, April 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NIa

[Ano74-236] Anonymous. News from industry. IEEE Spectrum, 11 (1):115, January 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NIb

[Ano74-237] Anonymous. News from industry. IEEE Spectrum, 11 (2):111, February 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NIc

[Ano74-238] Anonymous. News from industry. IEEE Spectrum, 11(3):115, March 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NId

[Ano74-239] Anonymous. News from industry. IEEE Spectrum, 11(4):87, April 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NIe

[Ano74-240] Anonymous. News from industry. IEEE Spectrum, 11(5):97, May 1974. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NIf

[Ano74-241] Anonymous. News from industry. IEEE Spectrum, 11(6):107, June 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NIg

[Ano74-242] Anonymous. News from industry. IEEE Spectrum, 11(7):91, July 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NIh

[Ano74-243] Anonymous. News from industry. IEEE Spectrum, 11 (9):107, September 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NIi

[Ano74-244] Anonymous. News from industry. IEEE Spectrum, 11 (10):117, October 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NIj

[Ano74-245] Anonymous. News from industry. IEEE Spectrum, 11 (11):99, November 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NIk

[Ano74-246] Anonymous. News from industry. IEEE Spectrum, 11(12):

94–100, December 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NRc

[Ano74-247] Anonymous. News from region 1. IEEE Spectrum, 11(3):116–121, March 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NRi

[Ano74-248] Anonymous. News from region 2. IEEE Spectrum, 11(10): 118–119, October 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NRb

[Ano74-249] Anonymous. News from region 3. IEEE Spectrum, 11(2): 112–115, February 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NRd

[Ano74-250] Anonymous. News from region 3. IEEE Spectrum, 11(4):88, April 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NRe

[Ano74-251] Anonymous. News from region 3. IEEE Spectrum, 11(5):98–105, May 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NRf

[Ano74-252] Anonymous. News from region 5. IEEE Spectrum, 11(6):108–112, June 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NRh

[Ano74-253] Anonymous. News from region 5. IEEE Spectrum, 11(9): 108–114, September 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NRg

[Ano74-254] Anonymous. News from region 7. IEEE Spectrum, 11(7):92, July 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NRj

[Ano74-255] Anonymous. News from region 8. IEEE Spectrum, 11(11): 100–101, November 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NRa

[Ano74-256] Anonymous. News from region 9. IEEE Spectrum, 11(1): 116–121, January 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NWa

[Ano74-257] Anonymous. News from Washington. IEEE Spectrum, 11 (1):17, January 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NWb

[Ano74-258] Anonymous. News from Washington. IEEE Spectrum, 11 (2):13, February 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NWc

[Ano74-259] Anonymous. News from Washington. IEEE Spectrum, 11 (3):17, March 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NWd

[Ano74-260] Anonymous. News from Washington. IEEE Spectrum, 11 (4):21, April 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NWe

[Ano74-261] Anonymous. News from Washington. IEEE Spectrum, 11 (5):23, May 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NWf

[Ano74-262] Anonymous. News from Washington. IEEE Spectrum, 11 (6):17, June 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NWg

[Ano74-263] Anonymous. News from Washington. IEEE Spectrum, 11(7): 17, July 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NWh

[Ano74-264] Anonymous. News from Washington. IEEE Spectrum, 11 (8):17, August 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NWi

[Ano74-265] Anonymous. News from Washington. IEEE Spectrum, 11 (9):25, September 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NWj

[Ano74-266] Anonymous. News from Washington. IEEE Spectrum, 11 (10):13, October 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NWk

[Ano74-267] Anonymous. News from Washington. IEEE Spectrum, 11 (11):25, November 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:NWl

[Ano74-268] Anonymous. News from Washington. IEEE Spectrum, 11 (12):17, December 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Pa

[Ano74-269] Anonymous. People. IEEE Spectrum, 11(1):130–131, January 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Pb

[Ano74-270] Anonymous. People. IEEE Spectrum, 11(2):122–123, February 1974. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Pc

[Ano74-271] Anonymous. People. IEEE Spectrum, 11(3):126–127, March 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Pd

[Ano74-272] Anonymous. People. IEEE Spectrum, 11(4):107, April 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Pe

[Ano74-273] Anonymous. People. IEEE Spectrum, 11(5):118–119, May 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Pf

[Ano74-274] Anonymous. People. IEEE Spectrum, 11(6):121–123, June 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Pg

[Ano74-275] Anonymous. People. IEEE Spectrum, 11(7):106–107, July 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Ph

[Ano74-276] Anonymous. People. IEEE Spectrum, 11(8):114–115, August 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Pi

[Ano74-277] Anonymous. People. IEEE Spectrum, 11(9):124–127, September 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Pj

[Ano74-278] Anonymous. People. IEEE Spectrum, 11(10):128–129, October 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Pk

[Ano74-279] Anonymous. People. IEEE Spectrum, 11(11):122, November 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:Pl

[Ano74-280] Anonymous. People. IEEE Spectrum, 11(12):94–98, December 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:RN

[Ano74-281] Anonymous. Regional news.

IEEE Spectrum, 11(12):100, December 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SIa

[Ano74-282] Anonymous. Scanning the issues. IEEE Spectrum, 11(2): 103–105, February 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SIb

[Ano74-283] Anonymous. Scanning the issues. IEEE Spectrum, 11(3):106–107, March 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SIc

[Ano74-284] Anonymous. Scanning the issues. IEEE Spectrum, 11(4):90–91, April 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SId

[Ano74-285] Anonymous. Scanning the issues. IEEE Spectrum, 11(5):91–92, May 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SIe

[Ano74-286] Anonymous. Scanning the issues. IEEE Spectrum, 11(6):99–100, June 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SIf

[Ano74-287] Anonymous. Scanning the issues. IEEE Spectrum, 11(7):93– 94, July 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SIg

[Ano74-288] Anonymous. Scanning the issues. IEEE Spectrum, 11(8): 96–97, August 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SIh

[Ano74-289] Anonymous. Scanning the issues. IEEE Spectrum, 11(9): 100–101, September 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SIj

[Ano74-290] Anonymous. Scanning the issues. IEEE Spectrum, 11(10): 105–106, October 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SIk

[Ano74-291] Anonymous. Scanning the issues. IEEE Spectrum, 11(11): 96–98, November 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SIl

[Ano74-292] Anonymous. Scanning the issues. IEEE Spectrum, 11(12): 78–79, December 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SII

[Ano74-293] Anonymous. Scanning the issues: Input-output analysis: a method for our madness. IEEE Spectrum, 11(1):109–110, January 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SD

[Ano74-294] Anonymous. Special duplications. IEEE Spectrum, 11 (7):102–103, July 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SPb

[Ano74-295] Anonymous. Special publications. IEEE Spectrum, 11(2): 116–117, February 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SPc

[Ano74-296] Anonymous. Special publications. IEEE Spectrum, 11(3): 122–123, March 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SPd

[Ano74-297] Anonymous. Special publications. IEEE Spectrum, 11(4): 102–103, April 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SPe

[Ano74-298] Anonymous. Special publications. IEEE Spectrum, 11 (5):108–110, May 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SPf

[Ano74-299] Anonymous. Special publications. IEEE Spectrum, 11 (6):114–116, June 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SPg

[Ano74-300] Anonymous. Special publications. IEEE Spectrum, 11(9):

116–118, September 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SPh

[Ano74-301] Anonymous. Special publications. IEEE Spectrum, 11(10): 121–122, October 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SPi

[Ano74-302] Anonymous. Special publications. IEEE Spectrum, 11(11): 116–117, November 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SPj

[Ano74-303] Anonymous. Special publications. IEEE Spectrum, 11(12): 90–91, December 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SPA

[Ano74-304] Anonymous. Special publications [advertisement]. IEEE Spectrum, 11(1):122–123, January 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:SRK

[Ano74-305] Anonymous. Special report:

Keeping up, moving up. IEEE Spectrum, 11(5):34–35, May 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:TC

[Ano74-306] Anonymous. Table of contents. IEEE Spectrum, 11 (2):2–3, February 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1974:WCF

[Ano74-307] Anonymous. What can the Fujitsu SCPC terminal do for you? IEEE Spectrum, 11 (3):98–99, March 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Ac

[Ano75a] Anonymous. [advertisement]. IEEE Spectrum, 12(2):1–2, February 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Ad

[Ano75b] Anonymous. [advertisement]. IEEE Spectrum, 12(2):1–5, February 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Ae

[Ano75c] Anonymous. [advertisement]. IEEE Spectrum, 12(3):1–3, March 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Af

[Ano75d] Anonymous. [advertisement]. IEEE Spectrum, 12(3):2–5, March 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Ag

[Ano75e] Anonymous. [advertisement].

IEEE Spectrum, 12(4):1–3, April 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Ah

[Ano75f] Anonymous. [advertisement].

IEEE Spectrum, 12(4):2–5, April 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Ai

[Ano75g] Anonymous. [advertisement].

IEEE Spectrum, 12(5):1–2, May 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Aj

[Ano75h] Anonymous. [advertisement].

IEEE Spectrum, 12(5):1, May 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Ak

[Ano75i] Anonymous. [advertisement].

IEEE Spectrum, 12(5):4–7, May 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Am

[Ano75j] Anonymous. [advertisement].

IEEE Spectrum, 12(5):1–4, May 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:An

[Ano75k] Anonymous. [advertisement]. IEEE Spectrum, 12(6):1–2, June 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Ao

[Ano75l] Anonymous. [advertisement]. IEEE Spectrum, 12(6):1, June 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Ap

[Ano75m] Anonymous. [advertisement]. IEEE Spectrum, 12(6):4–11, June 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Aq

[Ano75n] Anonymous. [advertisement]. IEEE Spectrum, 12(6):1–2, June 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:As

[Ano75o] Anonymous. [advertisement]. IEEE Spectrum, 12(7):1, July 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:At

[Ano75p] Anonymous. [advertisement]. IEEE Spectrum, 12(7):4–5, July 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Au

[Ano75q] Anonymous. [advertisement].

IEEE Spectrum, 12(7):1–3, July 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Av

[Ano75r] Anonymous. [advertisement].

IEEE Spectrum, 12(8):1–2, August 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Aw

[Ano75s] Anonymous. [advertisement].

IEEE Spectrum, 12(8):1, August 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Ax

[Ano75t] Anonymous. [advertisement].

IEEE Spectrum, 12(8):4–7, August 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Ay

[Ano75u] Anonymous. [advertisement].

IEEE Spectrum, 12(8):1–2, August 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Az

[Ano75v] Anonymous. [advertisement].

IEEE Spectrum, 12(9):1–3, September 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Aba

[Ano75w] Anonymous. [advertisement]. IEEE Spectrum, 12(9):4–7, September 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Abb

[Ano75x] Anonymous. [advertisement]. IEEE Spectrum, 12(9):1, September 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Abc

[Ano75y] Anonymous. [advertisement]. IEEE Spectrum, 12(10):1–2, October 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Abd

[Ano75z] Anonymous. [advertisement]. IEEE Spectrum, 12(10):2–7, October 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Abe

[Ano75-27] Anonymous. [advertisement]. IEEE Spectrum, 12(10):1, October 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Abf

[Ano75-28] Anonymous. [advertisement]. IEEE Spectrum, 12(11):2–5, November 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Abg

[Ano75-29] Anonymous. [advertisement].

IEEE Spectrum, 12(11):1–3, November 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Aa

[Ano75-30] Anonymous. [advertisements].

IEEE Spectrum, 12(1):1, January 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Ab

[Ano75-31] Anonymous. [advertisements].

IEEE Spectrum, 12(1):4–9, January 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Abh

[Ano75-32] Anonymous. [advertisements].

IEEE Spectrum, 12(12):2–5, December 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ALa

[Ano75-33] Anonymous. Applications literature. IEEE Spectrum, 12 (1):103, January 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ALb

[Ano75-34] Anonymous. Applications literature. IEEE Spectrum, 12 (2):71, February 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ALc

[Ano75-35] Anonymous. Applications literature. IEEE Spectrum, 12(3):73, March 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ALd

[Ano75-36] Anonymous. Applications literature. IEEE Spectrum, 12(4):82, April 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ALe

[Ano75-37] Anonymous. Applications literature. IEEE Spectrum, 12 (12):61, December 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ALBa

[Ano75-38] Anonymous. Applications literature: Basic theory of low-pass filters. IEEE Spectrum, 12(7):70, July 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ALDa

[Ano75-39] Anonymous. Applications literature: Designing systems with bipolar microprocessor chips. IEEE Spectrum, 12(5):72–73, May 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ALF

[Ano75-40] Anonymous. Applications literature: Fundamentals of isolation amplifiers. IEEE Spectrum, 12 (11):61, November 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ALO

[Ano75-41] Anonymous. Applications literature: Operating principles of a programmable controller system. IEEE Spectrum, 12(6):85, June 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ALPa

[Ano75-42] Anonymous. Applications literature: Preemphasis and other amplitude adjustment techniques. IEEE Spectrum, 12(8):91, August 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ALPb

[Ano75-43] Anonymous. Applications literature: Primer on short-circuit effects in power systems. IEEE Spectrum, 12(9):85, September 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:AR

[Ano75-44] Anonymous. Auditors’ report.

IEEE Spectrum, 12(6):70–72, June 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:BCa

[Ano75-45] Anonymous. [back cover]. IEEE Spectrum, 12(1):c4, January 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:BCb

[Ano75-46] Anonymous. [back cover]. IEEE Spectrum, 12(12):c4, December 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:BRa

[Ano75-47] Anonymous. Book reviews. IEEE Spectrum, 12(2):90–91, February 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:BRb

[Ano75-48] Anonymous. Book reviews. IEEE Spectrum, 12(3):88–89, March 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:BRc

[Ano75-49] Anonymous. Book reviews. IEEE Spectrum, 12(4):106–107, April 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:BRd

[Ano75-50] Anonymous. Book reviews. IEEE Spectrum, 12(5):83–87, May 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:BRe

[Ano75-51] Anonymous. Book reviews. IEEE Spectrum, 12(6):96–97, June 1975. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:BRf

[Ano75-52] Anonymous. Book reviews.

IEEE Spectrum, 12(7):83–85, July 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:BRg

[Ano75-53] Anonymous. Book reviews.

IEEE Spectrum, 12(8):106–107, August 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:BRh

[Ano75-54] Anonymous. Book reviews.

IEEE Spectrum, 12(9):103–104, September 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:BRi

[Ano75-55] Anonymous. Book reviews.

IEEE Spectrum, 12(10):102– 104, October 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:BRj

[Ano75-56] Anonymous. Book reviews.

IEEE Spectrum, 12(11):70–72, November 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ISAa

[Ano75-57] Anonymous. IEEE Spectrum annual index volume 12, 1975. IEEE Spectrum, 12(12):81–91, December 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SHRa

[Ano75-58] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 12(1):102, January 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SHRb

[Ano75-59] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 12(4):81, April 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SHRc

[Ano75-60] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 12(5):71, May 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SHRd

[Ano75-61] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 12(6):82, June 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SHRe

[Ano75-62] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 12(7):69, July 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SHRf

[Ano75-63] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 12(8):90, August 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SHRg

[Ano75-64] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 12(9):84, September 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SHRh

[Ano75-65] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 12 (11):60, November 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SHRi

[Ano75-66] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 12(12):60, December 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Cb

[Ano75-67] Anonymous. Calendar. IEEE Spectrum, 12(1):20–21, January 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Cd

[Ano75-68] Anonymous. Calendar. IEEE Spectrum, 12(2):9–12, February 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Cf

[Ano75-69] Anonymous. Calendar. IEEE Spectrum, 12(3):18–19, March 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Ch

[Ano75-70] Anonymous. Calendar. IEEE Spectrum, 12(4):22–25, April 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Cj

[Ano75-71] Anonymous. Calendar. IEEE Spectrum, 12(5):16–19, May 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Cl

[Ano75-72] Anonymous. Calendar. IEEE Spectrum, 12(6):14–16, June 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Cn

[Ano75-73] Anonymous. Calendar. IEEE Spectrum, 12(7):12–13, July 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Cp

[Ano75-74] Anonymous. Calendar. IEEE Spectrum, 12(8):20–23, August 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Cr

[Ano75-75] Anonymous. Calendar. IEEE Spectrum, 12(9):12–13, September 1975. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Ct

[Ano75-76] Anonymous. Calendar. IEEE Spectrum, 12(10):19–20, October 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Cv

[Ano75-77] Anonymous. Calendar. IEEE Spectrum, 12(11):14–17, November 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Cx

[Ano75-78] Anonymous. Calendar. IEEE Spectrum, 12(12):10–13, December 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:CAa

[Ano75-79] Anonymous. Classified advertising. IEEE Spectrum, 12(2): 87–89, February 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:CAb

[Ano75-80] Anonymous. Classified advertising. IEEE Spectrum, 12 (3):86–87, March 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:CAc

[Ano75-81] Anonymous. Classified advertising. IEEE Spectrum, 12 (6):94–95, June 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Ca

[Ano75-82] Anonymous. Contents. IEEE Spectrum, 12(1):2–3, January 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Cc

[Ano75-83] Anonymous. Contents. IEEE Spectrum, 12(2):2–3, February 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Ce

[Ano75-84] Anonymous. Contents. IEEE Spectrum, 12(3):2–3, March 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Cg

[Ano75-85] Anonymous. Contents. IEEE Spectrum, 12(4):2–3, April 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Ci

[Ano75-86] Anonymous. Contents. IEEE Spectrum, 12(5):2–3, May 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Ck

[Ano75-87] Anonymous. Contents. IEEE Spectrum, 12(6):2–3, June 1975.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Cm

[Ano75-88] Anonymous. Contents. IEEE Spectrum, 12(7):2–3, July 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Co

[Ano75-89] Anonymous. Contents. IEEE Spectrum, 12(8):2–3, August 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Cq

[Ano75-90] Anonymous. Contents. IEEE Spectrum, 12(9):2–3, September 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Cs

[Ano75-91] Anonymous. Contents. IEEE Spectrum, 12(10):1, October 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Cu

[Ano75-92] Anonymous. Contents. IEEE Spectrum, 12(11):1, November 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Cw

[Ano75-93] Anonymous. Contents. IEEE Spectrum, 12(12):1, December 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:EAa

[Ano75-94] Anonymous. Educational aids. IEEE Spectrum, 12(1):118, January 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:EAb

[Ano75-95] Anonymous. Educational aids. IEEE Spectrum, 12(2):86, February 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:EAc

[Ano75-96] Anonymous. Educational aids. IEEE Spectrum, 12(3):85, March 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:EAd

[Ano75-97] Anonymous. Educational aids. IEEE Spectrum, 12(4):100–101, April 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:EAe

[Ano75-98] Anonymous. Educational aids. IEEE Spectrum, 12(5):82, May 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:EAf

[Ano75-99] Anonymous. Educational aids. IEEE Spectrum, 12(6):93, June 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:EAg

[Ano75-100] Anonymous. Educational aids.

IEEE Spectrum, 12(7):82, July 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:EAh

[Ano75-101] Anonymous. Educational aids.

IEEE Spectrum, 12(8):104–105, August 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:EAi

[Ano75-102] Anonymous. Educational aids.

IEEE Spectrum, 12(9):101–102, September 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:EAj

[Ano75-103] Anonymous. Educational aids.

IEEE Spectrum, 12(10):101, October 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:EAk

[Ano75-104] Anonymous. Educational aids.

IEEE Spectrum, 12(11):69, November 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:EAAa

[Ano75-105] Anonymous. Educational aids [advertisement]. IEEE Spectrum, 12(12):70–73, December 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ERa

[Ano75-106] Anonymous. Energy report. IEEE Spectrum, 12(1):18–19, January 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ERb

[Ano75-107] Anonymous. Energy report. IEEE Spectrum, 12(2):14, February 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ERc

[Ano75-108] Anonymous. Energy report. IEEE Spectrum, 12(3):14–17, March 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ERd

[Ano75-109] Anonymous. Energy report. IEEE Spectrum, 12(4):20–21, April 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ERe

[Ano75-110] Anonymous. Energy report. IEEE Spectrum, 12(5):14–15, May 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ERf

[Ano75-111] Anonymous. Energy report. IEEE Spectrum, 12(6):18–21, June 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ERg

[Ano75-112] Anonymous. Energy report.

IEEE Spectrum, 12(7):18, July 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ERh

[Ano75-113] Anonymous. Energy report.

IEEE Spectrum, 12(8):18–19, August 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ERi

[Ano75-114] Anonymous. Energy report.

IEEE Spectrum, 12(9):18–18E, September 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ERj

[Ano75-115] Anonymous. Energy report.

IEEE Spectrum, 12(10):18, October 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ERk

[Ano75-116] Anonymous. Energy report.

IEEE Spectrum, 12(11):22–28, November 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ERl

[Ano75-117] Anonymous. Energy report.

IEEE Spectrum, 12(12):18–19, December 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FPRa

[Ano75-118] Anonymous. Focal points: a review of current developments. IEEE Spectrum, 12 (1):22–24, January 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FPRb

[Ano75-119] Anonymous. Focal points: a review of current developments. IEEE Spectrum, 12(2): 15–17, February 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FPRc

[Ano75-120] Anonymous. Focal points: a review of current developments. IEEE Spectrum, 12(3):20–21, March 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FPRd

[Ano75-121] Anonymous. Focal points: a review of current developments. IEEE Spectrum, 12(4):26–27, April 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FPRe

[Ano75-122] Anonymous. Focal points: a review of current developments. IEEE Spectrum, 12(5):20–21, May 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FPRf

[Ano75-123] Anonymous. Focal points: a review of current developments. IEEE Spectrum, 12(6):22, June 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FPRg

[Ano75-124] Anonymous. Focal points: a review of current developments. IEEE Spectrum, 12(7):14, July 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FPRh

[Ano75-125] Anonymous. Focal points: a review of current developments. IEEE Spectrum, 12(8):24, August 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FPRi

[Ano75-126] Anonymous. Focal points:

a review of current developments. IEEE Spectrum, 12(9): 14–16, September 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FPRj

[Ano75-127] Anonymous. Focal points:

a review of current developments. IEEE Spectrum, 12 (10):21–23, October 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Fa

[Ano75-128] Anonymous. Forum. IEEE Spectrum, 12(2):24–28, February 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Fb

[Ano75-129] Anonymous. Forum. IEEE Spectrum, 12(3):24–32, March 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Fd

[Ano75-130] Anonymous. Forum. IEEE Spectrum, 12(4):32–38, April 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Fe

[Ano75-131] Anonymous. Forum. IEEE Spectrum, 12(5):24–30, May 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Ff

[Ano75-132] Anonymous. Forum. IEEE Spectrum, 12(6):28–32, June 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Fg

[Ano75-133] Anonymous. Forum. IEEE Spectrum, 12(7):19–24, July 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Fh

[Ano75-134] Anonymous. Forum. IEEE Spectrum, 12(8):26–32, August 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Fj

[Ano75-135] Anonymous. Forum. IEEE Spectrum, 12(9):24–34, September 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Fk

[Ano75-136] Anonymous. Forum. IEEE Spectrum, 12(10):26–36, October 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Fl

[Ano75-137] Anonymous. Forum. IEEE Spectrum, 12(11):20, November 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FCa

[Ano75-138] Anonymous. [front cover].

IEEE Spectrum, 12(1):c1, January 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FCb

[Ano75-139] Anonymous. [front cover]. IEEE Spectrum, 12(12):c1, December 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FSIa

[Ano75-140] Anonymous. Future special issues. IEEE Spectrum, 12(1): 113–116, January 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FSIb

[Ano75-141] Anonymous. Future special issues. IEEE Spectrum, 12 (2):84, February 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FSIc

[Ano75-142] Anonymous. Future special issues. IEEE Spectrum, 12(3):83, March 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FSId

[Ano75-143] Anonymous. Future special issues. IEEE Spectrum, 12(4):97, April 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FSIe

[Ano75-144] Anonymous. Future special issues. IEEE Spectrum, 12(5):78, May 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FSIf

[Ano75-145] Anonymous. Future special issues. IEEE Spectrum, 12(6):91, June 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FSIg

[Ano75-146] Anonymous. Future special issues. IEEE Spectrum, 12(7):80, July 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FSIh

[Ano75-147] Anonymous. Future special issues. IEEE Spectrum, 12(8):102, August 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FSIi

[Ano75-148] Anonymous. Future special issues. IEEE Spectrum, 12(9): 98–99, September 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FSIj

[Ano75-149] Anonymous. Future special issues. IEEE Spectrum, 12(10): 98–99, October 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FSIk

[Ano75-150] Anonymous. Future special issues. IEEE Spectrum, 12 (11):67, November 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FSIl

[Ano75-151] Anonymous. Future special issues. IEEE Spectrum, 12(12): 66–67, December 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:GA

[Ano75-152] Anonymous. Guide for authors. IEEE Spectrum, 12(1): 95–96, January 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:IAR

[Ano75-153] Anonymous. IEEE annual reports: ’74: Programs and progress: Members availed themselves of new technical, educational, and professional services. IEEE Spectrum, 12 (6):63–66, June 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:IPBa

[Ano75-154] Anonymous. IEEE press books. IEEE Spectrum, 12(3):85, March 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:IPBb

[Ano75-155] Anonymous. IEEE press books. IEEE Spectrum, 12(4):100, April 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:IPBc

[Ano75-156] Anonymous. IEEE press books. IEEE Spectrum, 12(5):82, May 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:IPBd

[Ano75-157] Anonymous. IEEE press books. IEEE Spectrum, 12(6):93, June 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:IPBe

[Ano75-158] Anonymous. IEEE press books.

IEEE Spectrum, 12(7):82, July 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:IPBf

[Ano75-159] Anonymous. IEEE press books.

IEEE Spectrum, 12(8):104, August 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:IPBg

[Ano75-160] Anonymous. IEEE press books. IEEE Spectrum, 12 (9):101, September 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:IPBh

[Ano75-161] Anonymous. IEEE press books. IEEE Spectrum, 12 (11):69, November 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:IPBi

[Ano75-162] Anonymous. IEEE press books.

IEEE Spectrum, 12(12):70, December 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ISa

[Ano75-163] Anonymous. IEEE standards.

IEEE Spectrum, 12(1):117, January 1975. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ISb

[Ano75-164] Anonymous. IEEE standards. IEEE Spectrum, 12(2):85, February 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ISc

[Ano75-165] Anonymous. IEEE standards. IEEE Spectrum, 12(3):84, March 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ISd

[Ano75-166] Anonymous. IEEE standards. IEEE Spectrum, 12(4):98, April 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ISe

[Ano75-167] Anonymous. IEEE standards. IEEE Spectrum, 12(5):81, May 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ISf

[Ano75-168] Anonymous. IEEE standards. IEEE Spectrum, 12(6):92, June 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ISg

[Ano75-169] Anonymous. IEEE standards. IEEE Spectrum, 12(7):81, July 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ISh

[Ano75-170] Anonymous. IEEE standards.

IEEE Spectrum, 12(8):103, August 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ISi

[Ano75-171] Anonymous. IEEE standards. IEEE Spectrum, 12(9): 100, September 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ISj

[Ano75-172] Anonymous. IEEE standards.

IEEE Spectrum, 12(10):100, October 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ISk

[Ano75-173] Anonymous. IEEE standards. IEEE Spectrum, 12 (11):68, November 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ISl

[Ano75-174] Anonymous. IEEE standards.

IEEE Spectrum, 12(12):68, December 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ITCe

[Ano75-175] Anonymous. IEEE tables of contents. IEEE Spectrum, 12 (5):76–78, May 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ITCf

[Ano75-176] Anonymous. IEEE tables of contents. IEEE Spectrum, 12 (6):88–91, June 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ITCg

[Ano75-177] Anonymous. IEEE tables of contents. IEEE Spectrum, 12(7):75– 80, July 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ITCh

[Ano75-178] Anonymous. IEEE tables of contents. IEEE Spectrum, 12(8):94, August 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ITCi

[Ano75-179] Anonymous. IEEE tables of contents. IEEE Spectrum, 12(9): 90–94, September 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ITCj

[Ano75-180] Anonymous. IEEE tables of contents. IEEE Spectrum, 12 (10):92–94, October 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ITCk

[Ano75-181] Anonymous. IEEE tables of contents. IEEE Spectrum, 12(11):

63–67, November 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ITCa

[Ano75-182] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 12(1):106, January 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ITCb

[Ano75-183] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 12 (2):74–80, February 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ITCc

[Ano75-184] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 12 (3):76–78, March 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ITCd

[Ano75-185] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 12(4):89–94, April 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:ITCl

[Ano75-186] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 12 (12):63–66, December 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FIa

[Ano75-187] Anonymous. In future issues. IEEE Spectrum, 12(1):123, January 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FIb

[Ano75-188] Anonymous. In future issues. IEEE Spectrum, 12(2):93, February 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FIc

[Ano75-189] Anonymous. In future issues. IEEE Spectrum, 12(3):1–8, March 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FId

[Ano75-190] Anonymous. In future issues. IEEE Spectrum, 12(4):110–116, April 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FIe

[Ano75-191] Anonymous. In future issues. IEEE Spectrum, 12(5):91–92, May 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FIf

[Ano75-192] Anonymous. In future issues. IEEE Spectrum, 12(6):99–104, June 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FIg

[Ano75-193] Anonymous. In future issues. IEEE Spectrum, 12(7):88– 90, July 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FIh

[Ano75-194] Anonymous. In future issues.

IEEE Spectrum, 12(8):111–114, August 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FIi

[Ano75-195] Anonymous. In future issues. IEEE Spectrum, 12(9): 110–114, September 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FIj

[Ano75-196] Anonymous. In future issues. IEEE Spectrum, 12(10): 112–115, October 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FIk

[Ano75-197] Anonymous. In future issues. IEEE Spectrum, 12(11): 76–80, November 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:FIl

[Ano75-198] Anonymous. In future issues.

IEEE Spectrum, 12(12):92, December 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:IAa

[Ano75-199] Anonymous. Index to advertisers. IEEE Spectrum, 12(1): 123–126, January 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:IAb

[Ano75-200] Anonymous. Index to advertisers. IEEE Spectrum, 12(12): 92–94, December 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:IBCa

[Ano75-201] Anonymous. [inside back cover]. IEEE Spectrum, 12(1):c3, January 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:IBCb

[Ano75-202] Anonymous. [inside back cover]. IEEE Spectrum, 12(12):c3, December 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:IFCa

[Ano75-203] Anonymous. [inside front cover]. IEEE Spectrum, 12(1):c2, January 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:IFCb

[Ano75-204] Anonymous. [inside front cover]. IEEE Spectrum, 12(12):c2, December 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:IIa

[Ano75-205] Anonymous. Inside IEEE. IEEE Spectrum, 12(1):25–29, January 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:IIb

[Ano75-206] Anonymous. Inside IEEE. IEEE Spectrum, 12(2):18–23, February 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:IIc

[Ano75-207] Anonymous. Inside IEEE. IEEE Spectrum, 12(3):22–23, March 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:IId

[Ano75-208] Anonymous. Inside IEEE. IEEE Spectrum, 12(4):28–31, April 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:IIe

[Ano75-209] Anonymous. Inside IEEE.

IEEE Spectrum, 12(5):22–23, May 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:IIf

[Ano75-210] Anonymous. Inside IEEE. IEEE Spectrum, 12(6):23–27, June 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:IIg

[Ano75-211] Anonymous. Inside IEEE. IEEE Spectrum, 12(7):15–16, July 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:IIh

[Ano75-212] Anonymous. Inside IEEE. IEEE Spectrum, 12(8):25, August 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:IIi

[Ano75-213] Anonymous. Inside IEEE. IEEE Spectrum, 12(9):20–23, September 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:IIj

[Ano75-214] Anonymous. Inside IEEE. IEEE Spectrum, 12(10):24–25, October 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:IIk

[Ano75-215] Anonymous. Inside IEEE. IEEE Spectrum, 12(11):18–19, November 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:IIl

[Ano75-216] Anonymous. Inside IEEE. IEEE Spectrum, 12(12):14–16, December 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:IRS

[Ano75-217] Anonymous. Institute report:

Statements from IEEE candidates: Those running for IEEE office this fall give brief indications of their positions on vital issues. IEEE Spectrum, 12 (8):79–85, August 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Ma

[Ano75-218] Anonymous. Meetings. IEEE Spectrum, 12(1):10–16, January 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Mb

[Ano75-219] Anonymous. Meetings. IEEE Spectrum, 12(2):6–8, February 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Mc

[Ano75-220] Anonymous. Meetings. IEEE Spectrum, 12(3):8–12, March 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Md

[Ano75-221] Anonymous. Meetings. IEEE Spectrum, 12(4):8–18, April 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Me

[Ano75-222] Anonymous. Meetings. IEEE Spectrum, 12(5):8–12, May 1975.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Mf

[Ano75-223] Anonymous. Meetings. IEEE Spectrum, 12(6):10–13, June 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Mg

[Ano75-224] Anonymous. Meetings. IEEE Spectrum, 12(7):6–11, July 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Mh

[Ano75-225] Anonymous. Meetings. IEEE Spectrum, 12(8):8–16, August 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Mi

[Ano75-226] Anonymous. Meetings. IEEE Spectrum, 12(9):8–11, September 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Mj

[Ano75-227] Anonymous. Meetings. IEEE Spectrum, 12(10):8–16, October 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Mk

[Ano75-228] Anonymous. Meetings. IEEE Spectrum, 12(11):006–013, November 1975. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Ml

[Ano75-229] Anonymous. Meetings. IEEE Spectrum, 12(12):6–9, December 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NPAn

[Ano75-230] Anonymous. New product applications. IEEE Spectrum, 12 (12):57–60, December 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NPAg

[Ano75-231] Anonymous. New product applications: a 16-bit microcomputer for OEMs and high-volume end-users. IEEE Spectrum, 12 (5):67–71, May 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NPAb

[Ano75-232] Anonymous. New product applications: Advanced microprocessor costs $29.95; offers set of 60 instructions, single-step operation. IEEE Spectrum, 12 (2):66–69, February 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NPAc

[Ano75-233] Anonymous. New product applications: Compact digital multimeter measures capacitance, is low in price. IEEE Spectrum, 12(2):70, February 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NPAe

[Ano75-234] Anonymous. New product applications: Digital-transmission-line test set is also a multimeter, is battery charged, and is portable. IEEE Spectrum, 12(3):72, March 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NPAm

[Ano75-235] Anonymous. New product applications: Hybrid 12-bit A/D and D/A converters offer high performance at lowest prices. IEEE Spectrum, 12(11):57–60, November 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NPAl

[Ano75-236] Anonymous. New product applications: IC op amp combines high-voltage JFETs with standard bipolar transistors on one chip. IEEE Spectrum, 12 (10):83–89, October 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NPAd

[Ano75-237] Anonymous. New product applications: Input-output options give microcomputer unlimited number of uses in communications. IEEE Spectrum, 12 (3):68–72, March 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NPAj

[Ano75-238] Anonymous. New product applications: Low-cost, 4-bit microprocessor replaces any PC board containing 30 to 90 packages of TTL. IEEE Spectrum, 12 (8):86–90, August 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NPAa

[Ano75-239] Anonymous. New product applications: Microcomputer system includes processor, memory, interface, and modem. IEEE Spectrum, 12(1):97–102, January 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NPAf

[Ano75-240] Anonymous. New product applications: More calculators thinking polish as a famous pioneer faces some rough competition. IEEE Spectrum, 12 (4):77–81, April 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NPAh

[Ano75-241] Anonymous. New product applications: Production 16-kb CCD memories are offered for bulk designs. IEEE Spectrum, 12 (6):78–82, June 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NPAk

[Ano75-242] Anonymous. New product applications: Thermoelectric DIP relay features cost, size, and performance advantages. IEEE Spectrum, 12(9):80–84, September 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NPAi

[Ano75-243] Anonymous. New product applications: Two-chip, eight-bit microprocessor is I/O oriented to minimize external parts requirements. IEEE Spectrum, 12(7):65–69, July 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NIa

[Ano75-244] Anonymous. News from industry. IEEE Spectrum, 12 (1):107, January 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NIb

[Ano75-245] Anonymous. News from industry. IEEE Spectrum, 12 (2):81, February 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NIc

[Ano75-246] Anonymous. News from industry. IEEE Spectrum, 12(3):79, March 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NId

[Ano75-247] Anonymous. News from industry. IEEE Spectrum, 12(4):95, April 1975. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NIe

[Ano75-248] Anonymous. News from industry. IEEE Spectrum, 12(5):79, May 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NIf

[Ano75-249] Anonymous. News from industry. IEEE Spectrum, 12(6):83, June 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NIg

[Ano75-250] Anonymous. News from industry. IEEE Spectrum, 12(7):71, July 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NIh

[Ano75-251] Anonymous. News from industry. IEEE Spectrum, 12(8):95, August 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NIi

[Ano75-252] Anonymous. News from industry. IEEE Spectrum, 12 (9):95, September 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NIj

[Ano75-253] Anonymous. News from industry. IEEE Spectrum, 12 (10):95, October 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NIk

[Ano75-254] Anonymous. News from industry. IEEE Spectrum, 12 (11):55, November 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NIl

[Ano75-255] Anonymous. News from industry. IEEE Spectrum, 12 (12):75, December 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NRb

[Ano75-256] Anonymous. News from region 1. IEEE Spectrum, 12 (2):82–83, February 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NRc

[Ano75-257] Anonymous. News from region 1. IEEE Spectrum, 12(4):96–97, April 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NRd

[Ano75-258] Anonymous. News from region 1. IEEE Spectrum, 12(5):80, May 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NRe

[Ano75-259] Anonymous. News from region 1. IEEE Spectrum, 12(6):84, June 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NRf

[Ano75-260] Anonymous. News from region 1. IEEE Spectrum, 12(7):72, July 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NRh

[Ano75-261] Anonymous. News from region 1. IEEE Spectrum, 12(9): 96–97, September 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NRi

[Ano75-262] Anonymous. News from region 1. IEEE Spectrum, 12 (10):96–98, October 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NRj

[Ano75-263] Anonymous. News from region 2.

IEEE Spectrum, 12(12):76, December 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NRg

[Ano75-264] Anonymous. News from region 7. IEEE Spectrum, 12(8):96–102, August 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NRa

[Ano75-265] Anonymous. News from region 9. IEEE Spectrum, 12(1):

108–113, January 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NUR

[Ano75-266] Anonymous. News from the U.S. regions. IEEE Spectrum, 12(11):56, November 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NU

[Ano75-267] Anonymous. News from USAB. IEEE Spectrum, 12(3):80–83, March 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NWa

[Ano75-268] Anonymous. News from Washington. IEEE Spectrum, 12 (1):17, January 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NWb

[Ano75-269] Anonymous. News from Washington. IEEE Spectrum, 12 (2):13, February 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NWc

[Ano75-270] Anonymous. News from Washington. IEEE Spectrum, 12 (3):13, March 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NWd

[Ano75-271] Anonymous. News from Washington. IEEE Spectrum, 12 (4):19, April 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NWe

[Ano75-272] Anonymous. News from Washington. IEEE Spectrum, 12 (5):13, May 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NWf

[Ano75-273] Anonymous. News from Washington. IEEE Spectrum, 12 (6):17, June 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NWg

[Ano75-274] Anonymous. News from Washington. IEEE Spectrum, 12(7): 17, July 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NWh

[Ano75-275] Anonymous. News from Washington. IEEE Spectrum, 12 (8):17, August 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NWi

[Ano75-276] Anonymous. News from Washington. IEEE Spectrum, 12 (9):17, September 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NWj

[Ano75-277] Anonymous. News from Washington. IEEE Spectrum, 12 (10):17, October 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NWk

[Ano75-278] Anonymous. News from Washington. IEEE Spectrum, 12 (11):21, November 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:NWl

[Ano75-279] Anonymous. News from Washington. IEEE Spectrum, 12 (12):17, December 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Pa

[Ano75-280] Anonymous. People. IEEE Spectrum, 12(1):121–122, January 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Pb

[Ano75-281] Anonymous. People. IEEE Spectrum, 12(2):92, February 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Pc

[Ano75-282] Anonymous. People. IEEE Spectrum, 12(3):90–91, March 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Pd

[Ano75-283] Anonymous. People. IEEE Spectrum, 12(4):108–109, April 1975.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Pe

[Ano75-284] Anonymous. People. IEEE Spectrum, 12(5):88–90, May 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Pf

[Ano75-285] Anonymous. People. IEEE Spectrum, 12(6):98, June 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Pg

[Ano75-286] Anonymous. People. IEEE Spectrum, 12(7):86–87, July 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Ph

[Ano75-287] Anonymous. People. IEEE Spectrum, 12(8):108–111, August 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Pi

[Ano75-288] Anonymous. People. IEEE Spectrum, 12(9):105–109, September 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Pj

[Ano75-289] Anonymous. People. IEEE Spectrum, 12(10):105–111, October 1975. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Pk

[Ano75-290] Anonymous. People. IEEE Spectrum, 12(11):73–75, November 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:Pl

[Ano75-291] Anonymous. People. IEEE Spectrum, 12(12):77–80, December 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SIa

[Ano75-292] Anonymous. Scanning the issues. IEEE Spectrum, 12(2): 72–73, February 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SIb

[Ano75-293] Anonymous. Scanning the issues. IEEE Spectrum, 12(3):74–75, March 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SIc

[Ano75-294] Anonymous. Scanning the issues. IEEE Spectrum, 12(4):87–88, April 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SId

[Ano75-295] Anonymous. Scanning the issues. IEEE Spectrum, 12(5):74–75, May 1975. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SIe

[Ano75-296] Anonymous. Scanning the issues.

IEEE Spectrum, 12(6):86–87, June 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SIf

[Ano75-297] Anonymous. Scanning the issues. IEEE Spectrum, 12(7):73– 74, July 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SIg

[Ano75-298] Anonymous. Scanning the issues. IEEE Spectrum, 12(8): 92–93, August 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SIh

[Ano75-299] Anonymous. Scanning the issues. IEEE Spectrum, 12(9): 86–89, September 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SIi

[Ano75-300] Anonymous. Scanning the issues. IEEE Spectrum, 12(10): 90–91, October 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SIj

[Ano75-301] Anonymous. Scanning the issues. IEEE Spectrum, 12(11): 62, November 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SIk

[Ano75-302] Anonymous. Scanning the issues. IEEE Spectrum, 12(12):62, December 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SIP

[Ano75-303] Anonymous. Scanning the issues: Power-generation safety reexamined: The impact of a court case. IEEE Spectrum, 12(1): 104–105, January 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SIS

[Ano75-304] Anonymous. Sociotechnology: IEEE speaks out on R&D: In a key position paper addressed to President Ford, the Institute cited the ominous impact of declining R&D funding. IEEE Spectrum, 12(9):76–79, September 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SPa

[Ano75-305] Anonymous. Special publications. IEEE Spectrum, 12(1): 117–118, January 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SPb

[Ano75-306] Anonymous. Special publications. IEEE Spectrum, 12(2): 85–86, February 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SPc

[Ano75-307] Anonymous. Special publications. IEEE Spectrum, 12 (3):84–85, March 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SPd

[Ano75-308] Anonymous. Special publications. IEEE Spectrum, 12 (4):98–100, April 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SPe

[Ano75-309] Anonymous. Special publications. IEEE Spectrum, 12 (5):81–82, May 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SPf

[Ano75-310] Anonymous. Special publications. IEEE Spectrum, 12 (6):92–93, June 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SPg

[Ano75-311] Anonymous. Special publications. IEEE Spectrum, 12(7):81, July 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SPh

[Ano75-312] Anonymous. Special publications. IEEE Spectrum, 12(8):

103–104, August 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SPi

[Ano75-313] Anonymous. Special publications. IEEE Spectrum, 12(9): 100–101, September 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SPj

[Ano75-314] Anonymous. Special publications. IEEE Spectrum, 12(10): 100–101, October 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SPk

[Ano75-315] Anonymous. Special publications. IEEE Spectrum, 12 (11):68, November 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SPl

[Ano75-316] Anonymous. Special publications. IEEE Spectrum, 12(12): 68–69, December 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SRI

[Ano75-317] Anonymous. Special report: IN-TERCON ’75: The week in brief. IEEE Spectrum, 12(4):83–A16, April 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:SLH

[Ano75-318] Anonymous. Spectral lines: Hawthorne revisited. IEEE Spectrum, 12(1):35, January 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1975:TFI

[Ano75-319] Anonymous. Technology forecasting III: IEEE Spectrum’s key area forecasts. IEEE Spectrum, 12(4):50–51, April 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:EES

[Ano76a] Anonymous. 1976 elections:

Election statements: Candidates’ positions, along with views and counterviews on proposed constitutional amendments. IEEE Spectrum, 13 (8):54–61, August 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Aa

[Ano76b] Anonymous. [advertisement].

IEEE Spectrum, 13(1):1–2, January 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Ab

[Ano76c] Anonymous. [advertisement].

IEEE Spectrum, 13(1):2–7, January 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Ac

[Ano76d] Anonymous. [advertisement].

IEEE Spectrum, 13(2):1–2, February 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Ad

[Ano76e] Anonymous. [advertisement]. IEEE Spectrum, 13(2):2–5, February 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Ae

[Ano76f] Anonymous. [advertisement]. IEEE Spectrum, 13(3):2–7, March 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Af

[Ano76g] Anonymous. [advertisement]. IEEE Spectrum, 13(4):4–9, April 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Ag

[Ano76h] Anonymous. [advertisement]. IEEE Spectrum, 13(5):1–2, May 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Ah

[Ano76i] Anonymous. [advertisement]. IEEE Spectrum, 13(5):2–7, May 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Ai

[Ano76j] Anonymous. [advertisement]. IEEE Spectrum, 13(6):1, June 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Aj

[Ano76k] Anonymous. [advertisement].

IEEE Spectrum, 13(6):2–7, June 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Ak

[Ano76l] Anonymous. [advertisement].

IEEE Spectrum, 13(6):1–4, June 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Am

[Ano76m] Anonymous. [advertisement].

IEEE Spectrum, 13(7):1–2, July 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:An

[Ano76n] Anonymous. [advertisement].

IEEE Spectrum, 13(7):2–7, July 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Ao

[Ano76o] Anonymous. [advertisement].

IEEE Spectrum, 13(7):1–4, July 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Ap

[Ano76p] Anonymous. [advertisement].

IEEE Spectrum, 13(8):2–7, August 1976. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Aq

[Ano76q] Anonymous. [advertisement]. IEEE Spectrum, 13(9):1–2, September 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:As

[Ano76r] Anonymous. [advertisement]. IEEE Spectrum, 13(9):4–9, September 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:At

[Ano76s] Anonymous. [advertisement]. IEEE Spectrum, 13(10):1–2, October 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Au

[Ano76t] Anonymous. [advertisement]. IEEE Spectrum, 13(10):2–7, October 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Av

[Ano76u] Anonymous. [advertisements]. IEEE Spectrum, 13(11):1–4, November 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Aw

[Ano76v] Anonymous. [advertisements]. IEEE Spectrum, 13(11):2–7, November 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Ax

[Ano76w] Anonymous. [advertisements].

IEEE Spectrum, 13(12):1, December 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Ay

[Ano76x] Anonymous. [advertisements].

IEEE Spectrum, 13(12):1–3, December 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ALa

[Ano76y] Anonymous. Applications literature. IEEE Spectrum, 13 (1):100, January 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ALb

[Ano76z] Anonymous. Applications literature. IEEE Spectrum, 13 (2):87, February 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ALc

[Ano76-27] Anonymous. Applications literature. IEEE Spectrum, 13(3):82, March 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ALd

[Ano76-28] Anonymous. Applications literature. IEEE Spectrum, 13(4):88, April 1976. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ALe

[Ano76-29] Anonymous. Applications literature. IEEE Spectrum, 13(5):93, May 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ALf

[Ano76-30] Anonymous. Applications literature. IEEE Spectrum, 13(6):86, June 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ALg

[Ano76-31] Anonymous. Applications literature. IEEE Spectrum, 13(7):120, July 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ALh

[Ano76-32] Anonymous. Applications literature. IEEE Spectrum, 13(8):68, August 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ALi

[Ano76-33] Anonymous. Applications literature. IEEE Spectrum, 13(10): 111–114, October 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ALj

[Ano76-34] Anonymous. Applications literature. IEEE Spectrum, 13(12):

59–75, December 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ALBb

[Ano76-35] Anonymous. Applications literature: Basic concepts of microwave devices and instrumentation. IEEE Spectrum, 13(11): 72–75, November 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ALBa

[Ano76-36] Anonymous. Applications literature: Basic magnetic shielding concepts and applications. IEEE Spectrum, 13(9):63–64, September 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:AR

[Ano76-37] Anonymous. Auditors’ report.

IEEE Spectrum, 13(7):110–112, July 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:AIa

[Ano76-38] Anonymous. Author index.

IEEE Spectrum, 13(12):81–95, December 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:BRa

[Ano76-39] Anonymous. Book reviews.

IEEE Spectrum, 13(1):115– 119, January 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:BRb

[Ano76-40] Anonymous. Book reviews. IEEE Spectrum, 13(2):101– 102, February 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:BRc

[Ano76-41] Anonymous. Book reviews. IEEE Spectrum, 13(3):94–97, March 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:BRd

[Ano76-42] Anonymous. Book reviews. IEEE Spectrum, 13(4):104–107, April 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:BRe

[Ano76-43] Anonymous. Book reviews. IEEE Spectrum, 13(5):112, May 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:BRf

[Ano76-44] Anonymous. Book reviews. IEEE Spectrum, 13(6):93–95, June 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:BRg

[Ano76-45] Anonymous. Book reviews. IEEE Spectrum, 13(7):127–129, July 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:BRh

[Ano76-46] Anonymous. Book reviews.

IEEE Spectrum, 13(8):75–78, August 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:BRi

[Ano76-47] Anonymous. Book reviews.

IEEE Spectrum, 13(9):74–77, September 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:BRj

[Ano76-48] Anonymous. Book reviews.

IEEE Spectrum, 13(10):115– 118, October 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:BRk

[Ano76-49] Anonymous. Book reviews.

IEEE Spectrum, 13(11):76–79, November 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:BRl

[Ano76-50] Anonymous. Book reviews.

IEEE Spectrum, 13(12):76–78, December 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:SHRa

[Ano76-51] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 13(1):99, January 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:SHRb

[Ano76-52] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 13(2):84, February 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:SHRc

[Ano76-53] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 13(3):81, March 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:SHRd

[Ano76-54] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 13(4):87, April 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:SHRe

[Ano76-55] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 13(5):92, May 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:SHRf

[Ano76-56] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 13(8):67, August 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:SHRg

[Ano76-57] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 13(9):62, September 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:SHRh

[Ano76-58] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 13(10):110, October 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:SHRi

[Ano76-59] Anonymous. Spectrum’s hardware review [reader service card]. IEEE Spectrum, 13 (11):71, November 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Cb

[Ano76-60] Anonymous. Calendar. IEEE Spectrum, 13(1):14–16, January 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Cd

[Ano76-61] Anonymous. Calendar. IEEE Spectrum, 13(2):14–19, February 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Cf

[Ano76-62] Anonymous. Calendar. IEEE Spectrum, 13(3):14–16, March 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Ch

[Ano76-63] Anonymous. Calendar. IEEE Spectrum, 13(4):14–16, April 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Cj

[Ano76-64] Anonymous. Calendar. IEEE Spectrum, 13(5):12–15, May 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Cl

[Ano76-65] Anonymous. Calendar. IEEE Spectrum, 13(6):14–15, June 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Cn

[Ano76-66] Anonymous. Calendar. IEEE Spectrum, 13(7):8–11, July 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Cp

[Ano76-67] Anonymous. Calendar. IEEE Spectrum, 13(8):12–13, August 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Cr

[Ano76-68] Anonymous. Calendar. IEEE Spectrum, 13(9):12–13, September 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Ct

[Ano76-69] Anonymous. Calendar. IEEE Spectrum, 13(10):15–17, October 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Cv

[Ano76-70] Anonymous. Calendar. IEEE Spectrum, 13(11):12–16, November 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Cx

[Ano76-71] Anonymous. Calendar. IEEE Spectrum, 13(12):12–16, December 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Ca

[Ano76-72] Anonymous. Contents. IEEE Spectrum, 13(1):1, January 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Cc

[Ano76-73] Anonymous. Contents. IEEE Spectrum, 13(2):1, February 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Ce

[Ano76-74] Anonymous. Contents. IEEE Spectrum, 13(3):1, March 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Cg

[Ano76-75] Anonymous. Contents. IEEE Spectrum, 13(4):1, April 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Ci

[Ano76-76] Anonymous. Contents. IEEE Spectrum, 13(5):1, May 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Ck

[Ano76-77] Anonymous. Contents. IEEE Spectrum, 13(6):2, June 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Cm

[Ano76-78] Anonymous. Contents. IEEE Spectrum, 13(7):1, July 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Co

[Ano76-79] Anonymous. Contents. IEEE Spectrum, 13(8):1, August 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Cq

[Ano76-80] Anonymous. Contents. IEEE Spectrum, 13(9):2, September 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Cs

[Ano76-81] Anonymous. Contents. IEEE Spectrum, 13(10):1, October 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Cu

[Ano76-82] Anonymous. Contents. IEEE Spectrum, 13(11):1, November 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Cw

[Ano76-83] Anonymous. Contents. IEEE Spectrum, 13(12):1, December 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:EAa

[Ano76-84] Anonymous. Educational aids.

IEEE Spectrum, 13(1):114, January 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:EAb

[Ano76-85] Anonymous. Educational aids.

IEEE Spectrum, 13(2):100, February 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:EAc

[Ano76-86] Anonymous. Educational aids.

IEEE Spectrum, 13(3):93, March 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:EAd

[Ano76-87] Anonymous. Educational aids.

IEEE Spectrum, 13(4):103, April 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:EAe

[Ano76-88] Anonymous. Educational aids. IEEE Spectrum, 13(5):109, May 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:EAf

[Ano76-89] Anonymous. Educational aids. IEEE Spectrum, 13(6):92, June 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:EAg

[Ano76-90] Anonymous. Educational aids. IEEE Spectrum, 13(7):126, July 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:EAh

[Ano76-91] Anonymous. Educational aids. IEEE Spectrum, 13(9):71–73, September 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ERa

[Ano76-92] Anonymous. Energy report. IEEE Spectrum, 13(1):18–19, January 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ERb

[Ano76-93] Anonymous. Energy report. IEEE Spectrum, 13(2):22, February 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ERc

[Ano76-94] Anonymous. Energy report.

IEEE Spectrum, 13(3):18–19, March 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ERd

[Ano76-95] Anonymous. Energy report.

IEEE Spectrum, 13(4):18–19, April 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ERe

[Ano76-96] Anonymous. Energy report.

IEEE Spectrum, 13(5):18–21, May 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ERf

[Ano76-97] Anonymous. Energy report.

IEEE Spectrum, 13(6):18–19, June 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ERg

[Ano76-98] Anonymous. Energy report.

IEEE Spectrum, 13(7):18, July 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ERh

[Ano76-99] Anonymous. Energy report.

IEEE Spectrum, 13(8):18, August 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ERi

[Ano76-100] Anonymous. Energy report. IEEE Spectrum, 13(9):16–22, September 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ERj

[Ano76-101] Anonymous. Energy report. IEEE Spectrum, 13(10):10, October 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ERk

[Ano76-102] Anonymous. Energy report. IEEE Spectrum, 13(11):18–19, November 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ERl

[Ano76-103] Anonymous. Energy report. IEEE Spectrum, 13(12):18–19, December 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:FPRa

[Ano76-104] Anonymous. Focal points: a review of current developments. IEEE Spectrum, 13 (1):20–23, January 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:FPRb

[Ano76-105] Anonymous. Focal points: a review of current developments. IEEE Spectrum, 13(2):20, February 1976. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:FPRc

[Ano76-106] Anonymous. Focal points: a review of current developments. IEEE Spectrum, 13(3):20–23, March 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:FPRd

[Ano76-107] Anonymous. Focal points: a review of current developments. IEEE Spectrum, 13(4):20–21, April 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:FPRe

[Ano76-108] Anonymous. Focal points: a review of current developments. IEEE Spectrum, 13(5):16, May 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:FPRf

[Ano76-109] Anonymous. Focal points: a review of current developments. IEEE Spectrum, 13(6):16, June 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:FPRg

[Ano76-110] Anonymous. Focal points: a review of current developments. IEEE Spectrum, 13(7):12–13, July 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Fa

[Ano76-111] Anonymous. Forum. IEEE Spectrum, 13(1):30–40, January 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Fb

[Ano76-112] Anonymous. Forum. IEEE Spectrum, 13(2):26–32, February 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Fc

[Ano76-113] Anonymous. Forum. IEEE Spectrum, 13(3):28–34, March 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Fd

[Ano76-114] Anonymous. Forum. IEEE Spectrum, 13(4):28–36, April 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Fe

[Ano76-115] Anonymous. Forum. IEEE Spectrum, 13(5):26–34, May 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Ff

[Ano76-116] Anonymous. Forum. IEEE Spectrum, 13(6):24–34, June 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Fg

[Ano76-117] Anonymous. Forum. IEEE Spectrum, 13(7):19–26, July 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Fh

[Ano76-118] Anonymous. Forum. IEEE Spectrum, 13(8):19–22, August 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Fi

[Ano76-119] Anonymous. Forum. IEEE Spectrum, 13(9):23–26, September 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Fj

[Ano76-120] Anonymous. Forum. IEEE Spectrum, 13(10):20–28, October 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Fk

[Ano76-121] Anonymous. Forum. IEEE Spectrum, 13(11):23–28, November 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Fl

[Ano76-122] Anonymous. Forum. IEEE Spectrum, 13(12):20–24, December 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:FMa

[Ano76-123] Anonymous. [front matter]. IEEE Spectrum, 13(11):1–2, November 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:FMb

[Ano76-124] Anonymous. [front matter]. IEEE Spectrum, 13(12):2–5, December 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:FSIa

[Ano76-125] Anonymous. Future special issues. IEEE Spectrum, 13 (1):112, January 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:FSIb

[Ano76-126] Anonymous. Future special issues. IEEE Spectrum, 13(2): 94–98, February 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:FSIc

[Ano76-127] Anonymous. Future special issues. IEEE Spectrum, 13 (3):89–91, March 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:FSId

[Ano76-128] Anonymous. Future special issues. IEEE Spectrum, 13(4):101, April 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:FSIe

[Ano76-129] Anonymous. Future special issues. IEEE Spectrum, 13 (5):105–107, May 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:FSIf

[Ano76-130] Anonymous. Future special issues. IEEE Spectrum, 13(6):90, June 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:FSIg

[Ano76-131] Anonymous. Future special issues. IEEE Spectrum, 13(7):124, July 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:FSIh

[Ano76-132] Anonymous. Future special issues. IEEE Spectrum, 13(8):72, August 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:FSIi

[Ano76-133] Anonymous. Future special issues. IEEE Spectrum, 13(9): 69, September 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:GA

[Ano76-134] Anonymous. Guide for authors. IEEE Spectrum, 13(1): 101–103, January 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:IAR

[Ano76-135] Anonymous. IEEE annual reports. IEEE Spectrum, 13 (7):100–106, July 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:IPa

[Ano76-136] Anonymous. IEEE people. IEEE Spectrum, 13(9):77, September 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:IPc

[Ano76-137] Anonymous. IEEE people. IEEE Spectrum, 13(9):78–86, September 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:IPd

[Ano76-138] Anonymous. IEEE people. IEEE Spectrum, 13(10):121–132, October 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:IPe

[Ano76-139] Anonymous. IEEE people. IEEE Spectrum, 13(11):80–84, November 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:IPBa

[Ano76-140] Anonymous. IEEE press books. IEEE Spectrum, 13(1):114, January 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:IPBb

[Ano76-141] Anonymous. IEEE press books. IEEE Spectrum, 13 (2):100, February 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:IPBc

[Ano76-142] Anonymous. IEEE press books.

IEEE Spectrum, 13(3):93, March 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:IPBd

[Ano76-143] Anonymous. IEEE press books.

IEEE Spectrum, 13(4):103, April 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:IPBe

[Ano76-144] Anonymous. IEEE press books.

IEEE Spectrum, 13(5):109–111, May 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:IPBf

[Ano76-145] Anonymous. IEEE press books.

IEEE Spectrum, 13(6):92, June 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:IPBg

[Ano76-146] Anonymous. IEEE press books.

IEEE Spectrum, 13(7):126, July 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:IPBh

[Ano76-147] Anonymous. IEEE press books. IEEE Spectrum, 13(8):74, August 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:IPBi

[Ano76-148] Anonymous. IEEE press books. IEEE Spectrum, 13(9):71–73, September 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ISa

[Ano76-149] Anonymous. IEEE standards. IEEE Spectrum, 13(1):113, January 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ISb

[Ano76-150] Anonymous. IEEE standards. IEEE Spectrum, 13(2):99, February 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ISc

[Ano76-151] Anonymous. IEEE standards. IEEE Spectrum, 13(3):92, March 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ISd

[Ano76-152] Anonymous. IEEE standards. IEEE Spectrum, 13(4):102, April 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ISe

[Ano76-153] Anonymous. IEEE standards.

IEEE Spectrum, 13(5):108, May 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ISf

[Ano76-154] Anonymous. IEEE standards.

IEEE Spectrum, 13(6):91, June 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ISg

[Ano76-155] Anonymous. IEEE standards.

IEEE Spectrum, 13(7):125, July 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ISh

[Ano76-156] Anonymous. IEEE standards.

IEEE Spectrum, 13(8):73, August 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ISi

[Ano76-157] Anonymous. IEEE standards. IEEE Spectrum, 13 (9):70, September 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ITCd

[Ano76-158] Anonymous. IEEE tables of contents. IEEE Spectrum, 13 (4):95–100, April 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ITCf

[Ano76-159] Anonymous. IEEE tables of contents. IEEE Spectrum, 13 (6):88–89, June 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ITCi

[Ano76-160] Anonymous. IEEE tables of contents. IEEE Spectrum, 13(9): 67–69, September 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ITCa

[Ano76-161] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 13(1):106, January 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ITCb

[Ano76-162] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 13 (2):89–93, February 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ITCc

[Ano76-163] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 13(3):84, March 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ITCe

[Ano76-164] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 13 (5):102–104, May 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ITCg

[Ano76-165] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 13 (7):122–124, July 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:ITCh

[Ano76-166] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 13 (8):68–71, August 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:FIa

[Ano76-167] Anonymous. In future issues. IEEE Spectrum, 13(11): 84, November 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:FIb

[Ano76-168] Anonymous. In future issues.

IEEE Spectrum, 13(12):95, December 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:IA

[Ano76-169] Anonymous. Index to advertisers. IEEE Spectrum, 13(12): 96–97, December 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:IIa

[Ano76-170] Anonymous. Inside IEEE. IEEE Spectrum, 13(1):24–29, January 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:IIb

[Ano76-171] Anonymous. Inside IEEE. IEEE Spectrum, 13(2):23–25, February 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:IIc

[Ano76-172] Anonymous. Inside IEEE. IEEE Spectrum, 13(3):24–27, March 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:IId

[Ano76-173] Anonymous. Inside IEEE. IEEE Spectrum, 13(4):22–27, April 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:IIe

[Ano76-174] Anonymous. Inside IEEE. IEEE Spectrum, 13(5):22–25, May 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:IIf

[Ano76-175] Anonymous. Inside IEEE. IEEE Spectrum, 13(6):20–23, June 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:IIg

[Ano76-176] Anonymous. Inside IEEE.

IEEE Spectrum, 13(7):14–16, July 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:IIh

[Ano76-177] Anonymous. Inside IEEE. IEEE Spectrum, 13(8):14–16, August 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:IIi

[Ano76-178] Anonymous. Inside IEEE. IEEE Spectrum, 13(9):14, September 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:IIj

[Ano76-179] Anonymous. Inside IEEE. IEEE Spectrum, 13(10):18–19, October 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:IIk

[Ano76-180] Anonymous. Inside IEEE. IEEE Spectrum, 13(11):20–22, November 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:IIl

[Ano76-181] Anonymous. Inside IEEE. IEEE Spectrum, 13(12):47–50, December 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Ma

[Ano76-182] Anonymous. Meetings. IEEE Spectrum, 13(1):8–13, January 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Mb

[Ano76-183] Anonymous. Meetings. IEEE Spectrum, 13(2):6–13, February 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Mc

[Ano76-184] Anonymous. Meetings. IEEE Spectrum, 13(3):6–13, March 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Md

[Ano76-185] Anonymous. Meetings. IEEE Spectrum, 13(4):8–13, April 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Me

[Ano76-186] Anonymous. Meetings. IEEE Spectrum, 13(5):6–11, May 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Mf

[Ano76-187] Anonymous. Meetings. IEEE Spectrum, 13(6):8–13, June 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Mg

[Ano76-188] Anonymous. Meetings. IEEE Spectrum, 13(7):5–7, July 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Mh

[Ano76-189] Anonymous. Meetings. IEEE Spectrum, 13(8):7–11, August 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Mi

[Ano76-190] Anonymous. Meetings. IEEE Spectrum, 13(9):8–11, September 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Mj

[Ano76-191] Anonymous. Meetings. IEEE Spectrum, 13(10):11–14, October 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Mk

[Ano76-192] Anonymous. Meetings. IEEE Spectrum, 13(11):8–11, November 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Ml

[Ano76-193] Anonymous. Meetings. IEEE Spectrum, 13(12):6–11, December 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NPAb

[Ano76-194] Anonymous. New product applications. IEEE Spectrum, 13 (2):80–83, February 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NPAc

[Ano76-195] Anonymous. New product applications. IEEE Spectrum, 13 (3):78–80, March 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NPAe

[Ano76-196] Anonymous. New product applications. IEEE Spectrum, 13 (5):89–91, May 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NPAf

[Ano76-197] Anonymous. New product applications. IEEE Spectrum, 13 (6):79–82, June 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NPAg

[Ano76-198] Anonymous. New product applications. IEEE Spectrum, 13(6):85, June 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NPAh

[Ano76-199] Anonymous. New product applications. IEEE Spectrum, 13 (7):113–114, July 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NPAi

[Ano76-200] Anonymous. New product applications. IEEE Spectrum, 13 (8):62–64, August 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NPAk

[Ano76-201] Anonymous. New product applications. IEEE Spectrum, 13 (10):107–109, October 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NPAl

[Ano76-202] Anonymous. New product applications. IEEE Spectrum, 13 (11):65–66, November 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NPAm

[Ano76-203] Anonymous. New product applications. IEEE Spectrum, 13 (12):55–58, December 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NPAa

[Ano76-204] Anonymous. New product applications: Microprocessor controller family optimized for application in machine and process controls. IEEE Spectrum, 13 (1):95–98, January 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NPAd

[Ano76-205] Anonymous. New product applications: Microprocessor development system reduces design turnaround time and cuts development costs. IEEE Spectrum, 13(4):84–86, April 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NPAj

[Ano76-206] Anonymous. New product applications: Ten-bit D/A converter has maximum error of ±10.05 percent; is priced as low as $9.95. IEEE Spectrum, 13(9): 59–61, September 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NIa

[Ano76-207] Anonymous. News from industry. IEEE Spectrum, 13 (1):107, January 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NIb

[Ano76-208] Anonymous. News from industry. IEEE Spectrum, 13 (2):85, February 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NIc

[Ano76-209] Anonymous. News from industry. IEEE Spectrum, 13(3):85, March 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NId

[Ano76-210] Anonymous. News from industry. IEEE Spectrum, 13(4):93, April 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NIe

[Ano76-211] Anonymous. News from industry. IEEE Spectrum, 13(5):99, May 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NIf

[Ano76-212] Anonymous. News from industry. IEEE Spectrum, 13(6):83, June 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NIg

[Ano76-213] Anonymous. News from industry. IEEE Spectrum, 13(7):115, July 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NIh

[Ano76-214] Anonymous. News from industry. IEEE Spectrum, 13(8):65, August 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NIi

[Ano76-215] Anonymous. News from industry. IEEE Spectrum, 13 (9):65, September 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NIj

[Ano76-216] Anonymous. News from industry. IEEE Spectrum, 13 (10):119, October 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NIk

[Ano76-217] Anonymous. News from industry. IEEE Spectrum, 13 (11):67, November 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NIl

[Ano76-218] Anonymous. News from industry. IEEE Spectrum, 13 (12):79, December 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NRa

[Ano76-219] Anonymous. News from region 1. IEEE Spectrum, 13(1): 108–112, January 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NRb

[Ano76-220] Anonymous. News from region 1. IEEE Spectrum, 13(2):86, February 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NRd

[Ano76-221] Anonymous. News from region 1. IEEE Spectrum, 13(5):100, May 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NRg

[Ano76-222] Anonymous. News from region 1. IEEE Spectrum, 13 (9):66, September 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NRh

[Ano76-223] Anonymous. News from region 1.

IEEE Spectrum, 13(10):120, October 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NRj

[Ano76-224] Anonymous. News from region 1.

IEEE Spectrum, 13(12):80, December 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NRc

[Ano76-225] Anonymous. News from the regions. IEEE Spectrum, 13(4):94, April 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NRe

[Ano76-226] Anonymous. News from the regions. IEEE Spectrum, 13(6):84, June 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NRf

[Ano76-227] Anonymous. News from the regions. IEEE Spectrum, 13 (7):116–119, July 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NRi

[Ano76-228] Anonymous. News from the regions. IEEE Spectrum, 13(11): 68–70, November 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NUR

[Ano76-229] Anonymous. News from the U.S. regions. IEEE Spectrum, 13(8):66, August 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NU

[Ano76-230] Anonymous. News from USAB. IEEE Spectrum, 13(3):86–89, March 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NWa

[Ano76-231] Anonymous. News from Washington. IEEE Spectrum, 13 (1):17, January 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NWb

[Ano76-232] Anonymous. News from Washington. IEEE Spectrum, 13 (2):21, February 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NWc

[Ano76-233] Anonymous. News from Washington. IEEE Spectrum, 13 (3):17, March 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NWd

[Ano76-234] Anonymous. News from Washington. IEEE Spectrum, 13 (4):17, April 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NWe

[Ano76-235] Anonymous. News from Washington. IEEE Spectrum, 13 (5):17, May 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NWf

[Ano76-236] Anonymous. News from Washington. IEEE Spectrum, 13 (6):17, June 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NWg

[Ano76-237] Anonymous. News from Washington. IEEE Spectrum, 13(7): 17, July 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NWh

[Ano76-238] Anonymous. News from Washington. IEEE Spectrum, 13 (8):17, August 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NWi

[Ano76-239] Anonymous. News from Washington. IEEE Spectrum, 13 (9):15, September 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NWj

[Ano76-240] Anonymous. News from Washington. IEEE Spectrum, 13 (10):9, October 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NWk

[Ano76-241] Anonymous. News from Washington. IEEE Spectrum, 13 (11):17, November 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:NWl

[Ano76-242] Anonymous. News from Washington. IEEE Spectrum, 13 (12):17, December 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Pa

[Ano76-243] Anonymous. People. IEEE Spectrum, 13(1):120–128, January 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Pb

[Ano76-244] Anonymous. People. IEEE Spectrum, 13(2):103–112, February 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Pc

[Ano76-245] Anonymous. People. IEEE Spectrum, 13(3):98–99, March 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Pd

[Ano76-246] Anonymous. People. IEEE Spectrum, 13(4):108–114, April 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Pe

[Ano76-247] Anonymous. People. IEEE Spectrum, 13(5):113–117, May 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Pf

[Ano76-248] Anonymous. People. IEEE Spectrum, 13(6):96–102, June 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Pg

[Ano76-249] Anonymous. People. IEEE Spectrum, 13(7):130–134, July 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:Ph

[Ano76-250] Anonymous. People. IEEE Spectrum, 13(8):79–82, August 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:PS

[Ano76-251] Anonymous. Products at the show. IEEE Spectrum, 13(5):98, May 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:RT

[Ano76-252] Anonymous. Report of the treasure. IEEE Spectrum, 13 (7):107–109, July 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:SIa

[Ano76-253] Anonymous. Scanning the issues. IEEE Spectrum, 13(1): 104–105, January 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:SIb

[Ano76-254] Anonymous. Scanning the issues. IEEE Spectrum, 13(2):88, February 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:SIc

[Ano76-255] Anonymous. Scanning the issues. IEEE Spectrum, 13(3):83, March 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:SId

[Ano76-256] Anonymous. Scanning the issues. IEEE Spectrum, 13(4):89–92, April 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:SIe

[Ano76-257] Anonymous. Scanning the issues. IEEE Spectrum, 13(5):101, May 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:SIf

[Ano76-258] Anonymous. Scanning the issues. IEEE Spectrum, 13(6):87, June 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:SIg

[Ano76-259] Anonymous. Scanning the issues. IEEE Spectrum, 13(7):121, July 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:SS

[Ano76-260] Anonymous. Scanning the sessions. IEEE Spectrum, 13 (5):96–97, May 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:SDa

[Ano76-261] Anonymous. Special duplications. IEEE Spectrum, 13(3):92, March 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:SDb

[Ano76-262] Anonymous. Special duplications. IEEE Spectrum, 13 (5):108–109, May 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:SPa

[Ano76-263] Anonymous. Special publications. IEEE Spectrum, 13 (1):113, January 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:SPb

[Ano76-264] Anonymous. Special publications. IEEE Spectrum, 13 (2):99, February 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:SPc

[Ano76-265] Anonymous. Special publications. IEEE Spectrum, 13(4):102, April 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:SPd

[Ano76-266] Anonymous. Special publications. IEEE Spectrum, 13 (6):91–92, June 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:SPe

[Ano76-267] Anonymous. Special publications. IEEE Spectrum, 13 (7):125–126, July 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:SPf

[Ano76-268] Anonymous. Special publications. IEEE Spectrum, 13 (8):73–74, August 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:SPg

[Ano76-269] Anonymous. Special publications. IEEE Spectrum, 13(9): 70–71, September 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:SRF

[Ano76-270] Anonymous. Special report: Ford and Carter speak out on technology: The Presidential candidates discuss their views on issues of interest to the engineer– voter. IEEE Spectrum, 13(10):

101–106, October 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:SLA

[Ano76-271] Anonymous. Spectral lines: Advocacy and ethics. IEEE Spectrum, 13(5):35, May 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:WWWa

[Ano76-272] Anonymous. What went wrong?

A special report. IEEE Spectrum, 13(10):30–32, October 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1976:WWWb

[Ano76-273] Anonymous. What went wrong XII: What ever happened to … ? These promising developments got sidetracked, untracked, or delayed somewhere along the way. IEEE Spectrum, 13(10): 97–100, October 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ARP

[Ano77a] Anonymous. 1976 annual report: Problems and progress reviewed. IEEE Spectrum, 14 (5):65–72, May 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Aa

[Ano77b] Anonymous. [advertisement].

IEEE Spectrum, 14(1):2–7, January 1977. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Ab

[Ano77c] Anonymous. [advertisement]. IEEE Spectrum, 14(2):1–2, February 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Ac

[Ano77d] Anonymous. [advertisement]. IEEE Spectrum, 14(2):2–7, February 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Ad

[Ano77e] Anonymous. [advertisement]. IEEE Spectrum, 14(3):2–5, March 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Ae

[Ano77f] Anonymous. [advertisement]. IEEE Spectrum, 14(4):1–2, April 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Af

[Ano77g] Anonymous. [advertisement]. IEEE Spectrum, 14(4):2–5, April 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Ag

[Ano77h] Anonymous. [advertisement]. IEEE Spectrum, 14(5):2–5, May 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Ah

[Ano77i] Anonymous. [advertisements].

IEEE Spectrum, 14(6):2–5, June 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Ai

[Ano77j] Anonymous. [advertisements].

IEEE Spectrum, 14(7):2–5, July 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Aj

[Ano77k] Anonymous. [advertisements].

IEEE Spectrum, 14(8):2–8, August 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Ak

[Ano77l] Anonymous. [advertisements].

IEEE Spectrum, 14(9):2–11, September 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Am

[Ano77m] Anonymous. [advertisements].

IEEE Spectrum, 14(10):2–8, October 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:An

[Ano77n] Anonymous. [advertisements].

IEEE Spectrum, 14(11):2–7, November 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Ao

[Ano77o] Anonymous. [advertisements]. IEEE Spectrum, 14(12):2–7, December 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ALa

[Ano77p] Anonymous. Applications literature. IEEE Spectrum, 14 (1):98–102, January 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ALb

[Ano77q] Anonymous. Applications literature. IEEE Spectrum, 14 (2):84–86, February 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ALc

[Ano77r] Anonymous. Applications literature. IEEE Spectrum, 14 (3):97–103, March 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ALd

[Ano77s] Anonymous. Applications literature. IEEE Spectrum, 14 (4):74–75, April 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ALe

[Ano77t] Anonymous. Applications literature. IEEE Spectrum, 14(5):91, May 1977. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ALf

[Ano77u] Anonymous. Applications literature. IEEE Spectrum, 14(6):79, June 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ALg

[Ano77v] Anonymous. Applications literature. IEEE Spectrum, 14(7):75– 78, July 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ALh

[Ano77w] Anonymous. Applications literature. IEEE Spectrum, 14 (8):86–89, August 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ALi

[Ano77x] Anonymous. Applications literature. IEEE Spectrum, 14(9): 102–108, September 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ALj

[Ano77y] Anonymous. Applications literature. IEEE Spectrum, 14 (10):81–84, October 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ALk

[Ano77z] Anonymous. Applications literature. IEEE Spectrum, 14(11):

123–128, November 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ALl

[Ano77-27] Anonymous. Applications literature. IEEE Spectrum, 14 (12):94, December 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:AR

[Ano77-28] Anonymous. Auditors’ report. IEEE Spectrum, 14(7):52–54, July 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ANF

[Ano77-29] Anonymous. Awards: The new Fellows: IEEE’s latest additions to its elite grade are named and the selection process is briefly described. IEEE Spectrum, 14 (2):74–77, February 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:BCa

[Ano77-30] Anonymous. [back cover]. IEEE Spectrum, 14(6):c4, June 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:BCb

[Ano77-31] Anonymous. [back cover]. IEEE Spectrum, 14(7):c4, July 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:BCc

[Ano77-32] Anonymous. [back cover]. IEEE Spectrum, 14(8):c4, August 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:BCd

[Ano77-33] Anonymous. [back cover]. IEEE Spectrum, 14(9):c4, September 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:BCf

[Ano77-34] Anonymous. [back cover]. IEEE Spectrum, 14(10):c4, October 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:BCg

[Ano77-35] Anonymous. [back cover]. IEEE Spectrum, 14(11):c4, November 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:BCE

[Ano77-36] Anonymous. Banking/commerce:

Electronic funds transfer: Too much, too soon? The technology is available but the barriers of regulations, economies of scale, and consumer attitudes must be overcome. IEEE Spectrum, 14(5):51, May 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:BRa

[Ano77-37] Anonymous. Book reviews. IEEE Spectrum, 14(1):103– 108, January 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:BRb

[Ano77-38] Anonymous. Book reviews. IEEE Spectrum, 14(2):91–96, February 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:BRc

[Ano77-39] Anonymous. Book reviews. IEEE Spectrum, 14(3):104–108, March 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:BRd

[Ano77-40] Anonymous. Book reviews. IEEE Spectrum, 14(4):76–80, April 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:BRe

[Ano77-41] Anonymous. Book reviews. IEEE Spectrum, 14(5):79–82, May 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:BRf

[Ano77-42] Anonymous. Book reviews. IEEE Spectrum, 14(10):68– 71, October 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ISA

[Ano77-43] Anonymous. IEEE Spectrum annual index volume 13, 1977. IEEE Spectrum, 14(12):95–106, December 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:SHRa

[Ano77-44] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 14(2):83, February 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:SHRb

[Ano77-45] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 14(3):96, March 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:SHRc

[Ano77-46] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 14(4):73, April 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:SHRd

[Ano77-47] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 14(5):90, May 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:SHRe

[Ano77-48] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 14(6):78, June 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:SHRf

[Ano77-49] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 14(7):74, July 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:SHRg

[Ano77-50] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 14(8):85, August 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:SHRh

[Ano77-51] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 14 (9):101, September 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:SHRi

[Ano77-52] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 14(10):80, October 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:SHRj

[Ano77-53] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 14 (11):122, November 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:SHRk

[Ano77-54] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 14(12):93, December 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Cb

[Ano77-55] Anonymous. Calendar. IEEE Spectrum, 14(1):11–12, January 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Cd

[Ano77-56] Anonymous. Calendar. IEEE Spectrum, 14(2):14–18, February 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Cf

[Ano77-57] Anonymous. Calendar. IEEE Spectrum, 14(3):10–13, March 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Ch

[Ano77-58] Anonymous. Calendar. IEEE Spectrum, 14(4):6, April 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:CDO

[Ano77-59] Anonymous. Communications:

Digital and optical: Digital networks are endemic and fiber optics are on the horizon for TV and telecommunications. IEEE Spectrum, 14(9):57–60, September 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:CCa

[Ano77-60] Anonymous. Conference calendar. IEEE Spectrum, 14(5):6–7, May 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:CCb

[Ano77-61] Anonymous. Conference calendar. IEEE Spectrum, 14 (6):19–21, June 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:CCc

[Ano77-62] Anonymous. Conference calendar. IEEE Spectrum, 14(7):6–8H, July 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:CCd

[Ano77-63] Anonymous. Conference calendar. IEEE Spectrum, 14 (8):11–14, August 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:CCe

[Ano77-64] Anonymous. Conference calendar. IEEE Spectrum, 14(9): 12–14, September 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:CCf

[Ano77-65] Anonymous. Conference calendar. IEEE Spectrum, 14(10): 14–15, October 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:CCg

[Ano77-66] Anonymous. Conference calendar. IEEE Spectrum, 14(11):

8–11, November 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:CCh

[Ano77-67] Anonymous. Conference calendar. IEEE Spectrum, 14(12): 8–11, December 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Ca

[Ano77-68] Anonymous. Contents. IEEE Spectrum, 14(1):1, January 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Cc

[Ano77-69] Anonymous. Contents. IEEE Spectrum, 14(2):1, February 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Ce

[Ano77-70] Anonymous. Contents. IEEE Spectrum, 14(3):1, March 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Cg

[Ano77-71] Anonymous. Contents. IEEE Spectrum, 14(4):1, April 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Ci

[Ano77-72] Anonymous. Contents. IEEE Spectrum, 14(5):1, May 1977.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Cj

[Ano77-73] Anonymous. Contents. IEEE Spectrum, 14(6):1, June 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Ck

[Ano77-74] Anonymous. Contents. IEEE Spectrum, 14(7):1, July 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Cl

[Ano77-75] Anonymous. Contents. IEEE Spectrum, 14(8):1, August 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Cm

[Ano77-76] Anonymous. Contents. IEEE Spectrum, 14(9):1, September 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Cn

[Ano77-77] Anonymous. Contents. IEEE Spectrum, 14(10):1, October 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Co

[Ano77-78] Anonymous. Contents. IEEE Spectrum, 14(11):1, November 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Cp

[Ano77-79] Anonymous. Contents. IEEE Spectrum, 14(12):2–1, December 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ETTa

[Ano77-80] Anonymous. EE’s tools toys.

IEEE Spectrum, 14(6):22–26, June 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ETTb

[Ano77-81] Anonymous. EEs’ tools toys.

IEEE Spectrum, 14(8):20, August 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ETTc

[Ano77-82] Anonymous. EEs’ tools toys.

IEEE Spectrum, 14(10):18– 19, October 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ETTd

[Ano77-83] Anonymous. EEs’ tools toys.

IEEE Spectrum, 14(12):15, December 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ERa

[Ano77-84] Anonymous. Energy report.

IEEE Spectrum, 14(1):22, January 1977. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ERb

[Ano77-85] Anonymous. Energy report. IEEE Spectrum, 14(2):10–13, February 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ERc

[Ano77-86] Anonymous. Energy report. IEEE Spectrum, 14(3):18–19, March 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ERd

[Ano77-87] Anonymous. Energy report. IEEE Spectrum, 14(4):8–9, April 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ERe

[Ano77-88] Anonymous. Energy report. IEEE Spectrum, 14(5):14–14h, May 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ERf

[Ano77-89] Anonymous. Energy report. IEEE Spectrum, 14(6):10–10H, June 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ERg

[Ano77-90] Anonymous. Energy report. IEEE Spectrum, 14(7):10, July 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ERh

[Ano77-91] Anonymous. Energy report.

IEEE Spectrum, 14(8):10–10H, August 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ERi

[Ano77-92] Anonymous. Energy report.

IEEE Spectrum, 14(9):16, September 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ERj

[Ano77-93] Anonymous. Energy report.

IEEE Spectrum, 14(10):10– 13, October 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ERk

[Ano77-94] Anonymous. Energy report.

IEEE Spectrum, 14(11):18–18H, November 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ERl

[Ano77-95] Anonymous. Energy report.

IEEE Spectrum, 14(12):18–18H, December 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ELa

[Ano77-96] Anonymous. The engineer at large. IEEE Spectrum, 14(1):

33–34, January 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ELb

[Ano77-97] Anonymous. The engineer at large. IEEE Spectrum, 14 (3):20–38, March 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ELc

[Ano77-98] Anonymous. The engineer at large. IEEE Spectrum, 14 (5):24–26, May 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ELd

[Ano77-99] Anonymous. The engineer at large. IEEE Spectrum, 14(7):14, July 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ELe

[Ano77-100] Anonymous. The engineer at large. IEEE Spectrum, 14(9): 30–31, September 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:ELf

[Ano77-101] Anonymous. The engineer at large. IEEE Spectrum, 14(11): 19–21, November 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Fa

[Ano77-102] Anonymous. Forum. IEEE Spectrum, 14(1):13–20, January 1977.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Fb

[Ano77-103] Anonymous. Forum. IEEE Spectrum, 14(2):19–20, February 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Fc

[Ano77-104] Anonymous. Forum. IEEE Spectrum, 14(3):14–16, March 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Fd

[Ano77-105] Anonymous. Forum. IEEE Spectrum, 14(4):10–13, April 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Fe

[Ano77-106] Anonymous. Forum. IEEE Spectrum, 14(5):8–12, May 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Ff

[Ano77-107] Anonymous. Forum. IEEE Spectrum, 14(6):6–8, June 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Fg

[Ano77-108] Anonymous. Forum. IEEE Spectrum, 14(7):11–13, July 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Fh

[Ano77-109] Anonymous. Forum. IEEE Spectrum, 14(9):17–29, September 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:FCa

[Ano77-110] Anonymous. [front cover]. IEEE Spectrum, 14(6):c1, June 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:FCb

[Ano77-111] Anonymous. [front cover]. IEEE Spectrum, 14(7):c1, July 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:FCc

[Ano77-112] Anonymous. [front cover]. IEEE Spectrum, 14(8):c1, August 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:FCd

[Ano77-113] Anonymous. [front cover]. IEEE Spectrum, 14(9):c1, September 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:FCe

[Ano77-114] Anonymous. [front cover]. IEEE Spectrum, 14(10):c1, October 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:FCf

[Ano77-115] Anonymous. [front cover]. IEEE Spectrum, 14(11):c1, November 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:FCg

[Ano77-116] Anonymous. [front cover]. IEEE Spectrum, 14(12):c1, December 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:GA

[Ano77-117] Anonymous. Guide for authors. IEEE Spectrum, 14(1): 23–24, January 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:HR

[Ano77-118] Anonymous. Hardware review.

IEEE Spectrum, 14(1):97, January 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:FIa

[Ano77-119] Anonymous. In future issues.

IEEE Spectrum, 14(1):1–5, January 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:FIb

[Ano77-120] Anonymous. In future issues. IEEE Spectrum, 14(2): 97–100, February 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:FIc

[Ano77-121] Anonymous. In future issues. IEEE Spectrum, 14(3):109–111, March 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:FId

[Ano77-122] Anonymous. In future issues. IEEE Spectrum, 14(4):89–92, April 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:FIe

[Ano77-123] Anonymous. In future issues. IEEE Spectrum, 14(5):101–106, May 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:IAa

[Ano77-124] Anonymous. Index to advertisers. IEEE Spectrum, 14 (5):99–100, May 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:IAb

[Ano77-125] Anonymous. Index to advertisers. IEEE Spectrum, 14(7):79– 80, July 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:IAc

[Ano77-126] Anonymous. Index to advertisers. IEEE Spectrum, 14(8):91–92, August 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:IAd

[Ano77-127] Anonymous. Index to advertisers. IEEE Spectrum, 14(9): 109–110, September 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:IAe

[Ano77-128] Anonymous. Index to advertisers. IEEE Spectrum, 14(10): 93–94, October 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:IAf

[Ano77-129] Anonymous. Index to advertisers. IEEE Spectrum, 14(11): 130–132B, November 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:IBCa

[Ano77-130] Anonymous. [inside back cover].

IEEE Spectrum, 14(6):c3, June 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:IBCb

[Ano77-131] Anonymous. [inside back cover].

IEEE Spectrum, 14(7):c3, July 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:IBCc

[Ano77-132] Anonymous. [inside back cover].

IEEE Spectrum, 14(8):c3, August 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:IBCd

[Ano77-133] Anonymous. [inside back cover]. IEEE Spectrum, 14 (9):c3, September 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:IBCe

[Ano77-134] Anonymous. [inside back cover]. IEEE Spectrum, 14(10):c3, October 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:IBCf

[Ano77-135] Anonymous. [inside back cover]. IEEE Spectrum, 14 (11):c3, November 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:IFCa

[Ano77-136] Anonymous. [inside front cover]. IEEE Spectrum, 14(6):c2, June 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:IFCb

[Ano77-137] Anonymous. [inside front cover]. IEEE Spectrum, 14(7):c2, July 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:IFCc

[Ano77-138] Anonymous. [inside front cover]. IEEE Spectrum, 14(8):c2, August 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:IFCd

[Ano77-139] Anonymous. [inside front cover]. IEEE Spectrum, 14 (9):c2, September 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:IFCe

[Ano77-140] Anonymous. [inside front cover].

IEEE Spectrum, 14(10):c2, October 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:IFCf

[Ano77-141] Anonymous. [inside front cover]. IEEE Spectrum, 14 (11):c2, November 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:IFCg

[Ano77-142] Anonymous. [inside front cover].

IEEE Spectrum, 14(12):c2, December 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:IIa

[Ano77-143] Anonymous. Inside IEEE. IEEE Spectrum, 14(1):25–32, January 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:IIb

[Ano77-144] Anonymous. Inside IEEE. IEEE Spectrum, 14(2):21–28b, February 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:IIc

[Ano77-145] Anonymous. Inside IEEE. IEEE Spectrum, 14(4):17–24b, April 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:IId

[Ano77-146] Anonymous. Inside IEEE. IEEE Spectrum, 14(5):15–23, May 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:IIe

[Ano77-147] Anonymous. Inside IEEE. IEEE Spectrum, 14(6):11–18, June 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Ma

[Ano77-148] Anonymous. Meetings. IEEE Spectrum, 14(1):8–10, January 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Mb

[Ano77-149] Anonymous. Meetings. IEEE Spectrum, 14(2):8, February 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Mc

[Ano77-150] Anonymous. Meetings. IEEE Spectrum, 14(3):6–9, March 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Md

[Ano77-151] Anonymous. Meetings. IEEE Spectrum, 14(4):83–88, April 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Me

[Ano77-152] Anonymous. Meetings. IEEE Spectrum, 14(5):73–78, May 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Mf

[Ano77-153] Anonymous. Meetings. IEEE Spectrum, 14(6):62–65, June 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Mg

[Ano77-154] Anonymous. Meetings. IEEE Spectrum, 14(7):62–64, July 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Mh

[Ano77-155] Anonymous. Meetings. IEEE Spectrum, 14(8):76–77, August 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Mi

[Ano77-156] Anonymous. Meetings. IEEE Spectrum, 14(9):88–93, September 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Mj

[Ano77-157] Anonymous. Meetings. IEEE Spectrum, 14(10):90–92, October 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Mk

[Ano77-158] Anonymous. Meetings. IEEE Spectrum, 14(11):108–111, November 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:Ml

[Ano77-159] Anonymous. Meetings. IEEE Spectrum, 14(12):78–83, December 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NRIa

[Ano77-160] Anonymous. New and recent IEEE publications. IEEE Spectrum, 14(5):92–98, May 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NRIb

[Ano77-161] Anonymous. New and recent IEEE publications. IEEE Spectrum, 14(6):68–132, June 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NRIc

[Ano77-162] Anonymous. New and recent IEEE publications. IEEE Spectrum, 14(7):55–59, July 1977.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NRId

[Ano77-163] Anonymous. New and recent IEEE publications. IEEE Spectrum, 14(8):67–68, August 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NRIe

[Ano77-164] Anonymous. New and recent IEEE publications. IEEE Spectrum, 14(9):81–84, September 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NRIf

[Ano77-165] Anonymous. New and recent IEEE publications. IEEE Spectrum, 14(10):65–67, October 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NRIg

[Ano77-166] Anonymous. New and recent IEEE publications. IEEE Spectrum, 14(11):99–105, November 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NRIh

[Ano77-167] Anonymous. New and recent IEEE publications. IEEE Spectrum, 14(12):64–67, December 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NPAb

[Ano77-168] Anonymous. New product applications. IEEE Spectrum, 14 (2):78–82, February 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NPAc

[Ano77-169] Anonymous. New product applications. IEEE Spectrum, 14 (3):89–90, March 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NPAd

[Ano77-170] Anonymous. New product applications. IEEE Spectrum, 14 (4):67–72, April 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NPAe

[Ano77-171] Anonymous. New product applications. IEEE Spectrum, 14 (5):83–86h, May 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NPAf

[Ano77-172] Anonymous. New product applications. IEEE Spectrum, 14 (6):73–74H, June 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NPAg

[Ano77-173] Anonymous. New product applications. IEEE Spectrum, 14 (7):65–70, July 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NPAi

[Ano77-174] Anonymous. New product applications. IEEE Spectrum, 14 (9):94–96, September 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NPAj

[Ano77-175] Anonymous. New product applications. IEEE Spectrum, 14 (10):72–80, October 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NPAk

[Ano77-176] Anonymous. New product applications. IEEE Spectrum, 14(11): 112–114H, November 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NPAl

[Ano77-177] Anonymous. New product applications. IEEE Spectrum, 14(12): 84–90H, December 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NPAa

[Ano77-178] Anonymous. New product applications: Low-profile keyboard design can cut costs for x– y switching and interface layouts. IEEE Spectrum, 14(1): 89–96, January 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NPAh

[Ano77-179] Anonymous. New product applications: Mailgram terminal stores frequently used texts and address lists. IEEE Spectrum, 14 (8):78–82H, August 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NIa

[Ano77-180] Anonymous. News from industry. IEEE Spectrum, 14 (1):87, January 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NIb

[Ano77-181] Anonymous. News from industry. IEEE Spectrum, 14 (2):87, February 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NIc

[Ano77-182] Anonymous. News from industry. IEEE Spectrum, 14(3):91, March 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NId

[Ano77-183] Anonymous. News from industry. IEEE Spectrum, 14(4):81, April 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NIe

[Ano77-184] Anonymous. News from industry. IEEE Spectrum, 14(5):87, May 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NIf

[Ano77-185] Anonymous. News from industry. IEEE Spectrum, 14(6):75, June 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NIg

[Ano77-186] Anonymous. News from industry. IEEE Spectrum, 14(7):71, July 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NIh

[Ano77-187] Anonymous. News from industry. IEEE Spectrum, 14(8):83, August 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NIi

[Ano77-188] Anonymous. News from industry. IEEE Spectrum, 14 (9):97, September 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NIj

[Ano77-189] Anonymous. News from industry. IEEE Spectrum, 14 (10):85, October 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NIk

[Ano77-190] Anonymous. News from industry. IEEE Spectrum, 14 (11):115, November 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NIl

[Ano77-191] Anonymous. News from industry. IEEE Spectrum, 14 (12):91, December 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NRb

[Ano77-192] Anonymous. News from region 1. IEEE Spectrum, 14 (2):88–90, February 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NRc

[Ano77-193] Anonymous. News from region 1. IEEE Spectrum, 14(3):92–95, March 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NRd

[Ano77-194] Anonymous. News from region 1. IEEE Spectrum, 14(4):82, April 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NRf

[Ano77-195] Anonymous. News from region 1. IEEE Spectrum, 14(6):76–78, June 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NRh

[Ano77-196] Anonymous. News from region 1. IEEE Spectrum, 14(8):84–85, August 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NRl

[Ano77-197] Anonymous. News from region 1. IEEE Spectrum, 14(11):

116–122, November 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NRe

[Ano77-198] Anonymous. News from region 2. IEEE Spectrum, 14(5):88–89, May 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NRa

[Ano77-199] Anonymous. News from the regions. IEEE Spectrum, 14 (1):88, January 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NRg

[Ano77-200] Anonymous. News from the regions. IEEE Spectrum, 14(7):72– 74, July 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NRj

[Ano77-201] Anonymous. News from the regions. IEEE Spectrum, 14(9): 98–101, September 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NRk

[Ano77-202] Anonymous. News from the regions. IEEE Spectrum, 14 (10):86–89, October 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NRm

[Ano77-203] Anonymous. News from the regions. IEEE Spectrum, 14(12):

92–93, December 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NWa

[Ano77-204] Anonymous. News from Washington. IEEE Spectrum, 14 (1):21, January 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NWb

[Ano77-205] Anonymous. News from Washington. IEEE Spectrum, 14 (2):9, February 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NWc

[Ano77-206] Anonymous. News from Washington. IEEE Spectrum, 14 (3):17, March 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NWd

[Ano77-207] Anonymous. News from Washington. IEEE Spectrum, 14(4):7, April 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NWe

[Ano77-208] Anonymous. News from Washington. IEEE Spectrum, 14 (5):13, May 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NWf

[Ano77-209] Anonymous. News from Washington. IEEE Spectrum, 14(6):

9, June 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NWg

[Ano77-210] Anonymous. News from Washington. IEEE Spectrum, 14(7): 9, July 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NWh

[Ano77-211] Anonymous. News from Washington. IEEE Spectrum, 14(8):9, August 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NWi

[Ano77-212] Anonymous. News from Washington. IEEE Spectrum, 14 (9):15, September 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NWj

[Ano77-213] Anonymous. News from Washington. IEEE Spectrum, 14 (10):9, October 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NWk

[Ano77-214] Anonymous. News from Washington. IEEE Spectrum, 14 (11):17, November 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:NWl

[Ano77-215] Anonymous. News from Washington. IEEE Spectrum, 14 (12):17, December 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:PCSa

[Ano77-216] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 14(6):80–84, June 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:PCSb

[Ano77-217] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 14(8):90, August 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:PCSc

[Ano77-218] Anonymous. Professional consulting services [advertisements]. IEEE Spectrum, 14(11):129, November 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:SIa

[Ano77-219] Anonymous. Scanning the Institute. IEEE Spectrum, 14(7):15– 16, July 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:SIb

[Ano77-220] Anonymous. Scanning the Institute. IEEE Spectrum, 14 (8):21–22, August 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:SIc

[Ano77-221] Anonymous. Scanning the Institute. IEEE Spectrum, 14(9): 32–32H, September 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:SId

[Ano77-222] Anonymous. Scanning the Institute. IEEE Spectrum, 14(10): 20–20H, October 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:SIe

[Ano77-223] Anonymous. Scanning the Institute. IEEE Spectrum, 14(11): 22–24, November 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:SIf

[Ano77-224] Anonymous. Scanning the Institute. IEEE Spectrum, 14 (12):16, December 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:SS

[Ano77-225] Anonymous. Scanning the sessions. IEEE Spectrum, 14 (8):64–66, August 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:SLM

[Ano77-226] Anonymous. Spectral lines:

Midyear report. IEEE Spectrum, 14(7):17, July 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:TTa

[Ano77-227] Anonymous. Tools toys. IEEE Spectrum, 14(2):29–30, February 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1977:TTb

[Ano77-228] Anonymous. Tools toys. IEEE Spectrum, 14(4):14–16, April 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Aa

[Ano78a] Anonymous. [advertisement]. IEEE Spectrum, 15(1):1, January 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Ab

[Ano78b] Anonymous. [advertisement]. IEEE Spectrum, 15(1):2–7, January 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Ac

[Ano78c] Anonymous. [advertisement]. IEEE Spectrum, 15(2):1–2, February 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Ad

[Ano78d] Anonymous. [advertisement]. IEEE Spectrum, 15(2):2–5, February 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Ae

[Ano78e] Anonymous. [advertisement].

IEEE Spectrum, 15(3):1–2, March 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Af

[Ano78f] Anonymous. [advertisement].

IEEE Spectrum, 15(3):2–7, March 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Ag

[Ano78g] Anonymous. [advertisement].

IEEE Spectrum, 15(4):1–2, April 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Ah

[Ano78h] Anonymous. [advertisement].

IEEE Spectrum, 15(4):2–7, April 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Ai

[Ano78i] Anonymous. [advertisement].

IEEE Spectrum, 15(4):1–2, April 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Aj

[Ano78j] Anonymous. [advertisement].

IEEE Spectrum, 15(5):1, May 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Ak

[Ano78k] Anonymous. [advertisement]. IEEE Spectrum, 15(5):2–5, May 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Al

[Ano78l] Anonymous. [advertisement]. IEEE Spectrum, 15(5):1–3, May 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Am

[Ano78m] Anonymous. [advertisement]. IEEE Spectrum, 15(6):1–2, June 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:An

[Ano78n] Anonymous. [advertisement]. IEEE Spectrum, 15(6):2–7, June 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Ao

[Ano78o] Anonymous. [advertisement]. IEEE Spectrum, 15(6):1–4, June 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Ap

[Ano78p] Anonymous. [advertisement]. IEEE Spectrum, 15(7):2–7, July 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Aq

[Ano78q] Anonymous. [advertisement].

IEEE Spectrum, 15(7):80, July 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Ar

[Ano78r] Anonymous. [advertisement].

IEEE Spectrum, 15(7):1–4, July 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:As

[Ano78s] Anonymous. [advertisement].

IEEE Spectrum, 15(8):1–2, August 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:At

[Ano78t] Anonymous. [advertisement].

IEEE Spectrum, 15(8):2–7, August 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Au

[Ano78u] Anonymous. [advertisement].

IEEE Spectrum, 15(8):84, August 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Av

[Ano78v] Anonymous. [advertisement].

IEEE Spectrum, 15(8):1–4, August 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Aw

[Ano78w] Anonymous. [advertisement]. IEEE Spectrum, 15(9):1–2, September 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Ax

[Ano78x] Anonymous. [advertisement]. IEEE Spectrum, 15(9):2–5, September 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Ay

[Ano78y] Anonymous. [advertisement]. IEEE Spectrum, 15(9):1–4, September 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Az

[Ano78z] Anonymous. [advertisement]. IEEE Spectrum, 15(10):1–2, October 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Aba

[Ano78-27] Anonymous. [advertisement]. IEEE Spectrum, 15(10):2–8, October 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Abb

[Ano78-28] Anonymous. [advertisement]. IEEE Spectrum, 15(11):2–7, November 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Abc

[Ano78-29] Anonymous. [advertisement].

IEEE Spectrum, 15(11):1–4, November 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Abd

[Ano78-30] Anonymous. [advertisement].

IEEE Spectrum, 15(12):2–7, December 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Abe

[Ano78-31] Anonymous. [advertisement].

IEEE Spectrum, 15(12):1–4, December 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ALb

[Ano78-32] Anonymous. Applications literature. IEEE Spectrum, 15(10): 115–123, October 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ALd

[Ano78-33] Anonymous. Applications literature. IEEE Spectrum, 15(11): 107–109, November 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ALAa

[Ano78-34] Anonymous. Applications literature: All about analog multipliers. IEEE Spectrum, 15(6):81–82, June 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ALAc

[Ano78-35] Anonymous. Applications literature: All about printers and plotters. IEEE Spectrum, 15(9): 96–101, September 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ALAb

[Ano78-36] Anonymous. Applications literature: Applying thermal analysis to electronics manufacturing. IEEE Spectrum, 15(8): 78–83, August 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ALC

[Ano78-37] Anonymous. Applications literature: Comprehensive symbol list for analog and hybrid computers. IEEE Spectrum, 15(3):73–81, March 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ALH

[Ano78-38] Anonymous. Applications literature: How to cool electronic equipment. IEEE Spectrum, 15 (4):77–82, April 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ALI

[Ano78-39] Anonymous. Applications literature: IC design methodology and applications explained. IEEE Spectrum, 15(5):78, May 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ALM

[Ano78-40] Anonymous. Applications literature: Making measurements with piezoreslstive strain gages. IEEE Spectrum, 15(2):83, February 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ALN

[Ano78-41] Anonymous. Applications literature: New methods for measuring return loss. IEEE Spectrum, 15(12):78, December 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ALT

[Ano78-42] Anonymous. Applications literature: Technical tips on creating PC artwork. IEEE Spectrum, 15 (1):104–107, January 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ALU

[Ano78-43] Anonymous. Applications literature: Using a logic analyzer to debug microcircuit systems. IEEE Spectrum, 15(7):67– 70, July 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:BBAa

[Ano78-44] Anonymous. Best bits: Applications of microprocessors. IEEE Spectrum, 15(1):14–15, January 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:BBAb

[Ano78-45] Anonymous. Best bits: Applications of microprocessors. IEEE Spectrum, 15(3):24–28, March 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:BBAc

[Ano78-46] Anonymous. Best bits: Applications of microprocessors: Designer immersed in pool solarheating project. IEEE Spectrum, 15(5):22–26, May 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:BBAd

[Ano78-47] Anonymous. Best bits: Applications of microprocessors: Distributed network fits in typewriter. IEEE Spectrum, 15 (7):23–24, July 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:BBAe

[Ano78-48] Anonymous. Best bits: Applications of microprocessors: Micropyramid may break time-share bottleneck. IEEE Spectrum, 15 (9):22–25, September 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:BBAf

[Ano78-49] Anonymous. Best bits: applications of microprocessors: micros to probe Jupiter …. IEEE Spectrum, 15(11):14–18, November 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:BRa

[Ano78-50] Anonymous. Book reviews.

IEEE Spectrum, 15(1):108– 109, January 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:BRb

[Ano78-51] Anonymous. Book reviews.

IEEE Spectrum, 15(2):84–85, February 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:BRc

[Ano78-52] Anonymous. Book reviews.

IEEE Spectrum, 15(3):82, March 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:BRd

[Ano78-53] Anonymous. Book reviews.

IEEE Spectrum, 15(4):59–63, April 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:BRe

[Ano78-54] Anonymous. Book reviews.

IEEE Spectrum, 15(5):60–66, May 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:BRf

[Ano78-55] Anonymous. Book reviews.

IEEE Spectrum, 15(6):59–61, June 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:BRg

[Ano78-56] Anonymous. Book reviews. IEEE Spectrum, 15(7):74–79, July 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:BRh

[Ano78-57] Anonymous. Book reviews. IEEE Spectrum, 15(8):62–65, August 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:BRi

[Ano78-58] Anonymous. Book reviews. IEEE Spectrum, 15(9):78–79, September 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:BRj

[Ano78-59] Anonymous. Book reviews. IEEE Spectrum, 15(10):100– 101, October 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:BRk

[Ano78-60] Anonymous. Book reviews. IEEE Spectrum, 15(11):76–83, November 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:BRl

[Ano78-61] Anonymous. Book reviews. IEEE Spectrum, 15(12):58, December 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ISA

[Ano78-62] Anonymous. IEEE Spectrum annual index. IEEE Spectrum, 15 (12):79–90, December 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:SHRa

[Ano78-63] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 15(1):103, January 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:SHRb

[Ano78-64] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 15(2):80, February 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:SHRc

[Ano78-65] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 15(3):72, March 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:SHRd

[Ano78-66] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 15(4):74, April 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:SHRe

[Ano78-67] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 15(6):78–81, June 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:SHRf

[Ano78-68] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 15(7):66, July 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:SHRg

[Ano78-69] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 15(8):78, August 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:SHRh

[Ano78-70] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 15(9):95, September 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:SHRi

[Ano78-71] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 15 (10):113–114, October 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:SHRj

[Ano78-72] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 15 (11):106, November 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:SHRk

[Ano78-73] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 15(12):77, December 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Cf

[Ano78-74] Anonymous. Calendar. IEEE Spectrum, 15(4):8, April 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Cg

[Ano78-75] Anonymous. Calendar. IEEE Spectrum, 15(5):6–7, May 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Ci

[Ano78-76] Anonymous. Calendar. IEEE Spectrum, 15(6):8, June 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Ck

[Ano78-77] Anonymous. Calendar. IEEE Spectrum, 15(7):16, July 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Cm

[Ano78-78] Anonymous. Calendar. IEEE Spectrum, 15(8):8, August 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Co

[Ano78-79] Anonymous. Calendar. IEEE Spectrum, 15(9):12–19, September 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Cq

[Ano78-80] Anonymous. Calendar. IEEE Spectrum, 15(10):11–13, October 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Ct

[Ano78-81] Anonymous. Calendar. IEEE Spectrum, 15(11):8, November 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Cv

[Ano78-82] Anonymous. Calendar. IEEE Spectrum, 15(12):8–11, December 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:CAa

[Ano78-83] Anonymous. Classified advertising. IEEE Spectrum, 15 (5):79–82, May 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:CSa

[Ano78-84] Anonymous. Coming in Spectrum. IEEE Spectrum, 15 (3):22–23, March 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:CSb

[Ano78-85] Anonymous. Coming in Spectrum. IEEE Spectrum, 15 (4):18–20, April 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:CSc

[Ano78-86] Anonymous. Coming in Spectrum. IEEE Spectrum, 15 (5):20–21, May 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:CSd

[Ano78-87] Anonymous. Coming in Spectrum. IEEE Spectrum, 15 (6):18–22, June 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:CSe

[Ano78-88] Anonymous. Coming in Spectrum. IEEE Spectrum, 15(7):22, July 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:CSf

[Ano78-89] Anonymous. Coming in Spectrum. IEEE Spectrum, 15 (8):16–18, August 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:CSg

[Ano78-90] Anonymous. Coming in Spectrum. IEEE Spectrum, 15(9): 26–28, September 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:CSh

[Ano78-91] Anonymous. Coming in Spectrum. IEEE Spectrum, 15(10): 26–30, October 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:CSi

[Ano78-92] Anonymous. Coming in Spectrum. IEEE Spectrum, 15(11): 25–28, November 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:CSj

[Ano78-93] Anonymous. Coming in Spectrum. IEEE Spectrum, 15(12): 19–20, December 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:CCG

[Ano78-94] Anonymous. Computers: Consolidating gains: Solid-state progress is tempered by design maturation. IEEE Spectrum, 15 (1):26–30, January 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:CCa

[Ano78-95] Anonymous. Conference calendar. IEEE Spectrum, 15 (1):8–9, January 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:CCb

[Ano78-96] Anonymous. Conference calendar. IEEE Spectrum, 15 (2):6–9, February 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:CCc

[Ano78-97] Anonymous. Conference calendar. IEEE Spectrum, 15(3):8–13, March 1978. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Ca

[Ano78-98] Anonymous. Contents. IEEE Spectrum, 15(1):1, January 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Cb

[Ano78-99] Anonymous. Contents. IEEE Spectrum, 15(2):1, February 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Cd

[Ano78-100] Anonymous. Contents. IEEE Spectrum, 15(3):1, March 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Ce

[Ano78-101] Anonymous. Contents. IEEE Spectrum, 15(4):1, April 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Ch

[Ano78-102] Anonymous. Contents. IEEE Spectrum, 15(6):1, June 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Cj

[Ano78-103] Anonymous. Contents. IEEE Spectrum, 15(7):1, July 1978.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Cl

[Ano78-104] Anonymous. Contents. IEEE Spectrum, 15(8):1, August 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Cn

[Ano78-105] Anonymous. Contents. IEEE Spectrum, 15(9):1, September 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Cp

[Ano78-106] Anonymous. Contents. IEEE Spectrum, 15(10):1, October 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Cr

[Ano78-107] Anonymous. Contents. IEEE Spectrum, 15(11):1, November 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Cu

[Ano78-108] Anonymous. Contents. IEEE Spectrum, 15(12):1, December 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ETTa

[Ano78-109] Anonymous. EEs’ tools toys. IEEE Spectrum, 15(2):19, February 1978. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ETTb

[Ano78-110] Anonymous. EEs’ tools toys.

IEEE Spectrum, 15(4):16–17, April 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ETTc

[Ano78-111] Anonymous. EEs’ tools toys.

IEEE Spectrum, 15(6):16–17, June 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ETTd

[Ano78-112] Anonymous. EEs’ tools toys.

IEEE Spectrum, 15(8):14–15, August 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ETTe

[Ano78-113] Anonymous. EEs’ tools toys.

IEEE Spectrum, 15(10):20– 25, October 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ETTf

[Ano78-114] Anonymous. EEs’ tools toys.

IEEE Spectrum, 15(12):16, December 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ERa

[Ano78-115] Anonymous. Energy report.

IEEE Spectrum, 15(1):18, January 1978. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ERb

[Ano78-116] Anonymous. Energy report. IEEE Spectrum, 15(2):18, February 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ERc

[Ano78-117] Anonymous. Energy report. IEEE Spectrum, 15(3):20, March 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ERd

[Ano78-118] Anonymous. Energy report. IEEE Spectrum, 15(4):10, April 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ERe

[Ano78-119] Anonymous. Energy report. IEEE Spectrum, 15(5):14–17, May 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ERf

[Ano78-120] Anonymous. Energy report. IEEE Spectrum, 15(6):10, June 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ERg

[Ano78-121] Anonymous. Energy report. IEEE Spectrum, 15(7):18, July 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ERh

[Ano78-122] Anonymous. Energy report.

IEEE Spectrum, 15(8):10–11, August 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ERi

[Ano78-123] Anonymous. Energy report.

IEEE Spectrum, 15(9):10–11, September 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ERj

[Ano78-124] Anonymous. Energy report.

IEEE Spectrum, 15(10):10, October 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ERk

[Ano78-125] Anonymous. Energy report.

IEEE Spectrum, 15(11):10–13, November 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ERl

[Ano78-126] Anonymous. Energy report.

IEEE Spectrum, 15(12):18, December 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ELa

[Ano78-127] Anonymous. The engineer at large. IEEE Spectrum, 15(1):

16–16B, January 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ELb

[Ano78-128] Anonymous. The engineer at large. IEEE Spectrum, 15(3):21, March 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ELc

[Ano78-129] Anonymous. The engineer at large. IEEE Spectrum, 15 (5):18–19, May 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ELd

[Ano78-130] Anonymous. The engineer at large. IEEE Spectrum, 15(7):19– 21, July 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ELe

[Ano78-131] Anonymous. The engineer at large. IEEE Spectrum, 15(9): 20–21, September 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ELf

[Ano78-132] Anonymous. The engineer at large. IEEE Spectrum, 15 (11):24, November 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Fa

[Ano78-133] Anonymous. Forum. IEEE Spectrum, 15(1):10–13, January 1978.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Fb

[Ano78-134] Anonymous. Forum. IEEE Spectrum, 15(2):10–16b, February 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Fc

[Ano78-135] Anonymous. Forum. IEEE Spectrum, 15(3):14–18, March 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Fd

[Ano78-136] Anonymous. Forum. IEEE Spectrum, 15(4):11–15, April 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Fe

[Ano78-137] Anonymous. Forum. IEEE Spectrum, 15(5):8–8D, May 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Ff

[Ano78-138] Anonymous. Forum. IEEE Spectrum, 15(6):11–15, June 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Fg

[Ano78-139] Anonymous. Forum. IEEE Spectrum, 15(7):8–15, July 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Fh

[Ano78-140] Anonymous. Forum. IEEE Spectrum, 15(8):12–13, August 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Fj

[Ano78-141] Anonymous. Forum. IEEE Spectrum, 15(9):6–8, September 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Fk

[Ano78-142] Anonymous. Forum. IEEE Spectrum, 15(10):14–19, October 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Fl

[Ano78-143] Anonymous. Forum. IEEE Spectrum, 15(11):19–23, November 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Fm

[Ano78-144] Anonymous. Forum. IEEE Spectrum, 15(12):12–15, December 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:FSIa

[Ano78-145] Anonymous. Future special issues. IEEE Spectrum, 15(1):

89–95, January 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:FSIb

[Ano78-146] Anonymous. Future special issues. IEEE Spectrum, 15 (2):71, February 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:GA

[Ano78-147] Anonymous. Guide to authors. IEEE Spectrum, 15(1): 79–81, January 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ITCc

[Ano78-148] Anonymous. IEEE tables of contents. IEEE Spectrum, 15(7):56– 59, July 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ITCd

[Ano78-149] Anonymous. IEEE tables of contents. IEEE Spectrum, 15 (8):58–61, August 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ITCe

[Ano78-150] Anonymous. IEEE tables of contents. IEEE Spectrum, 15(9): 72–77, September 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ITCf

[Ano78-151] Anonymous. IEEE tables of contents. IEEE Spectrum, 15 (10):94–99, October 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ITCg

[Ano78-152] Anonymous. IEEE tables of contents. IEEE Spectrum, 15(11): 71–75, November 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ITCh

[Ano78-153] Anonymous. IEEE tables of contents. IEEE Spectrum, 15(12): 55–57, December 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ITCa

[Ano78-154] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 15 (1):86–88, January 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:ITCb

[Ano78-155] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 15 (2):67–71, February 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:FIa

[Ano78-156] Anonymous. In future issues. IEEE Spectrum, 15(1): 117–120, January 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:FIb

[Ano78-157] Anonymous. In future issues. IEEE Spectrum, 15(2): 99–102, February 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:IAa

[Ano78-158] Anonymous. Index to advertisers. IEEE Spectrum, 15(3):87, March 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:IAb

[Ano78-159] Anonymous. Index to advertisers. IEEE Spectrum, 15(4):83–84, April 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:IAc

[Ano78-160] Anonymous. Index to advertisers. IEEE Spectrum, 15(5):83–84, May 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:IAd

[Ano78-161] Anonymous. Index to advertisers. IEEE Spectrum, 15(6):83, June 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:IAe

[Ano78-162] Anonymous. Index to advertisers. IEEE Spectrum, 15(9): 102, September 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:IAf

[Ano78-163] Anonymous. Index to advertisers. IEEE Spectrum, 15(12): 91–92, December 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:IRN

[Ano78-164] Anonymous. Introducing a revolutionary new Home Learning Center: The new Encyclopaedia Britannica Program at a substantially reduced price! IEEE Spectrum, 15(5):9–12, May 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:MPM

[Ano78-165] Anonymous. Meeting preview: MIDCON ’78: an encore for the micro: The microprocessor takes center stage at the first MID-CON to be held in Dallas. IEEE Spectrum, 15(11):66–67, November 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Ma

[Ano78-166] Anonymous. Meetings. IEEE Spectrum, 15(1):110–116, January 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Mb

[Ano78-167] Anonymous. Meetings. IEEE Spectrum, 15(2):86–98, February 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Mc

[Ano78-168] Anonymous. Meetings. IEEE Spectrum, 15(3):83–86, March 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Md

[Ano78-169] Anonymous. Meetings. IEEE Spectrum, 15(4):64–67, April 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Me

[Ano78-170] Anonymous. Meetings. IEEE Spectrum, 15(5):67–69, May 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Mf

[Ano78-171] Anonymous. Meetings. IEEE Spectrum, 15(6):62–68, June 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Mg

[Ano78-172] Anonymous. Meetings. IEEE Spectrum, 15(7):71–73, July 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Mh

[Ano78-173] Anonymous. Meetings. IEEE Spectrum, 15(8):66–68, August 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Mi

[Ano78-174] Anonymous. Meetings. IEEE Spectrum, 15(9):80–84, September 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Mj

[Ano78-175] Anonymous. Meetings. IEEE Spectrum, 15(10):102–107, October 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Mk

[Ano78-176] Anonymous. Meetings. IEEE Spectrum, 15(11):84–97, November 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:Ml

[Ano78-177] Anonymous. Meetings. IEEE Spectrum, 15(12):59–69, December 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NRIa

[Ano78-178] Anonymous. New and recent IEEE publications. IEEE Spectrum, 15(1):82–85, January 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NRIb

[Ano78-179] Anonymous. New and recent IEEE publications. IEEE Spectrum, 15(2):72–73, February 1978. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NRIc

[Ano78-180] Anonymous. New and recent IEEE publications. IEEE Spectrum, 15(3):65–66, March 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NRId

[Ano78-181] Anonymous. New and recent IEEE publications. IEEE Spectrum, 15(4):56–58, April 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NRIe

[Ano78-182] Anonymous. New and recent IEEE publications. IEEE Spectrum, 15(5):56–57, May 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NRIf

[Ano78-183] Anonymous. New and recent IEEE publications. IEEE Spectrum, 15(6):57–58B, June 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NRIg

[Ano78-184] Anonymous. New and recent IEEE publications. IEEE Spectrum, 15(7):54–55, July 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NRIh

[Ano78-185] Anonymous. New and recent IEEE publications. IEEE Spectrum, 15(8):56–57, August 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NRIi

[Ano78-186] Anonymous. New and recent IEEE publications. IEEE Spectrum, 15(9):70–71, September 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NRIj

[Ano78-187] Anonymous. New and recent IEEE publications. IEEE Spectrum, 15(10):92–93, October 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NRIk

[Ano78-188] Anonymous. New and recent IEEE publications. IEEE Spectrum, 15(11):68–70, November 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NRIl

[Ano78-189] Anonymous. New and recent IEEE publications. IEEE Spectrum, 15(12):53–54, December 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NPAj

[Ano78-190] Anonymous. New product applications. IEEE Spectrum, 15 (10):108–110, October 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NPAg

[Ano78-191] Anonymous. New product applications: 12-bit D/A and A/D converters feature universal microprocessor interfacing. IEEE Spectrum, 15(7):60–62d, July 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NPAe

[Ano78-192] Anonymous. New product applications: CCD and CTD filters include tunable units and a spectrum analyzer on a chip. IEEE Spectrum, 15(5):70–74, May 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NPAl

[Ano78-193] Anonymous. New product applications: Desk-top computers feature a 256-kbyte read/write memory and assembly language programming. IEEE Spectrum, 15(12):70–74, December 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NPAa

[Ano78-194] Anonymous. New product applications: development ‘laboratory’ provides in-prototype testing for five common microprocessors. IEEE Spectrum, 15 (1):96–98, January 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NPAi

[Ano78-195] Anonymous. New product applications: Four-digit, presettable up/down CMOS counter family offers a host of features. IEEE Spectrum, 15(9):85–92, September 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NPAk

[Ano78-196] Anonymous. New product applications: Hand-held DMM/ calculator manipulates, stores, and retrieves values in five registers. IEEE Spectrum, 15(11): 98–100, November 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NPAh

[Ano78-197] Anonymous. New product applications: High-speed wideband FET-input op amp operates from a single positive supply. IEEE Spectrum, 15(8):69–74h, August 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NPAb

[Ano78-198] Anonymous. New product applications: Self-contained microprocessor chip contains 1024 bytes of ROM and 256 bits of RAM. IEEE Spectrum, 15(2):

74–79, February 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NPAd

[Ano78-199] Anonymous. New product applications: Single-board bipolar microcomputer emulates any minior microcomputer. IEEE Spectrum, 15(4):68–73, April 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NPAf

[Ano78-200] Anonymous. New product applications: Solid-state programmable limit switch can replace cam-operated counterpart. IEEE Spectrum, 15(6):69–74, June 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NPAc

[Ano78-201] Anonymous. New product applications: Two-board dataacquisition system for tm990 microcomputer operates with a single instruction. IEEE Spectrum, 15(3):67–68, March 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NIa

[Ano78-202] Anonymous. News from industry. IEEE Spectrum, 15 (1):99, January 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NIb

[Ano78-203] Anonymous. News from industry. IEEE Spectrum, 15 (2):81, February 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NIc

[Ano78-204] Anonymous. News from industry. IEEE Spectrum, 15(3):69, March 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NId

[Ano78-205] Anonymous. News from industry. IEEE Spectrum, 15(4):75, April 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NIe

[Ano78-206] Anonymous. News from industry. IEEE Spectrum, 15(5):75, May 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NIf

[Ano78-207] Anonymous. News from industry. IEEE Spectrum, 15(6):75, June 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NIg

[Ano78-208] Anonymous. News from industry. IEEE Spectrum, 15(7):63, July 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NIh

[Ano78-209] Anonymous. News from industry. IEEE Spectrum, 15(8):75, August 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NIi

[Ano78-210] Anonymous. News from industry. IEEE Spectrum, 15 (9):93, September 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NIj

[Ano78-211] Anonymous. News from industry. IEEE Spectrum, 15 (10):111, October 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NIk

[Ano78-212] Anonymous. News from industry. IEEE Spectrum, 15 (11):101, November 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NIl

[Ano78-213] Anonymous. News from industry. IEEE Spectrum, 15 (12):75, December 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NRb

[Ano78-214] Anonymous. News from region 1. IEEE Spectrum, 15(2): 82–82B, February 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NRd

[Ano78-215] Anonymous. News from region 1. IEEE Spectrum, 15(4):76, April 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NRe

[Ano78-216] Anonymous. News from region 1. IEEE Spectrum, 15(5):76–77, May 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NRf

[Ano78-217] Anonymous. News from region 1. IEEE Spectrum, 15(6):76–77, June 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NRl

[Ano78-218] Anonymous. News from region 1. IEEE Spectrum, 15(11): 102–105, November 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NRa

[Ano78-219] Anonymous. News from region 2. IEEE Spectrum, 15(1): 100–102, January 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NRc

[Ano78-220] Anonymous. News from the regions. IEEE Spectrum, 15 (3):70–71, March 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NRg

[Ano78-221] Anonymous. News from the regions. IEEE Spectrum, 15(7):64– 65, July 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NRh

[Ano78-222] Anonymous. News from the regions. IEEE Spectrum, 15 (8):76–77, August 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NRj

[Ano78-223] Anonymous. News from the regions. IEEE Spectrum, 15 (9):94, September 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NRk

[Ano78-224] Anonymous. News from the regions. IEEE Spectrum, 15 (10):112, October 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NRm

[Ano78-225] Anonymous. News from the regions. IEEE Spectrum, 15 (12):76, December 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NWa

[Ano78-226] Anonymous. News from Washington. IEEE Spectrum, 15 (1):17, January 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NWb

[Ano78-227] Anonymous. News from Washington. IEEE Spectrum, 15 (2):17, February 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NWc

[Ano78-228] Anonymous. News from Washington. IEEE Spectrum, 15 (3):19, March 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NWd

[Ano78-229] Anonymous. News from Washington. IEEE Spectrum, 15(4):9, April 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NWe

[Ano78-230] Anonymous. News from Washington. IEEE Spectrum, 15 (5):13, May 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NWf

[Ano78-231] Anonymous. News from Washington. IEEE Spectrum, 15(6): 9, June 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NWg

[Ano78-232] Anonymous. News from Washington. IEEE Spectrum, 15(7): 17, July 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NWh

[Ano78-233] Anonymous. News from Washington. IEEE Spectrum, 15(8):9, August 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NWi

[Ano78-234] Anonymous. News from Washington. IEEE Spectrum, 15 (9):9, September 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NWj

[Ano78-235] Anonymous. News from Washington. IEEE Spectrum, 15 (10):9, October 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NWk

[Ano78-236] Anonymous. News from Washington. IEEE Spectrum, 15 (11):9, November 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:NWl

[Ano78-237] Anonymous. News from Washington. IEEE Spectrum, 15 (12):17, December 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:PSa

[Ano78-238] Anonymous. Product selector.

IEEE Spectrum, 15(3):88–92, March 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:PSb

[Ano78-239] Anonymous. Product selector. IEEE Spectrum, 15(11): 110, November 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:PAA

[Ano78-240] Anonymous. Productivity about the authors. IEEE Spectrum, 15 (10):89–91, October 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:PFR

[Ano78-241] Anonymous. Productivity for further reading. IEEE Spectrum, 15(10):88–89, October 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:PDG

[Ano78-242] Anonymous. Productivity i: Defining the game and the players. IEEE Spectrum, 15 (10):34, October 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:PIM

[Ano78-243] Anonymous. Productivity III: Management to the rescue. IEEE Spectrum, 15(10):74, October 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:PIW

[Ano78-244] Anonymous. Productivity IV: What can government do? IEEE Spectrum, 15(10):85, October 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:SIa

[Ano78-245] Anonymous. Scanning the Institute. IEEE Spectrum, 15 (1):19–22, January 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:SIb

[Ano78-246] Anonymous. Scanning the Institute. IEEE Spectrum, 15(2): 20–24, February 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:SIc

[Ano78-247] Anonymous. Scanning the Institute. IEEE Spectrum, 15(3):22, March 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:SId

[Ano78-248] Anonymous. Scanning the Institute. IEEE Spectrum, 15(4):18, April 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:SIe

[Ano78-249] Anonymous. Scanning the Institute. IEEE Spectrum, 15(5):20, May 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:SIf

[Ano78-250] Anonymous. Scanning the Institute. IEEE Spectrum, 15(6):18, June 1978. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:SIg

[Ano78-251] Anonymous. Scanning the Institute. IEEE Spectrum, 15(7):22, July 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:SIh

[Ano78-252] Anonymous. Scanning the Institute. IEEE Spectrum, 15(8):16, August 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:SIi

[Ano78-253] Anonymous. Scanning the Institute. IEEE Spectrum, 15 (9):26, September 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:SIj

[Ano78-254] Anonymous. Scanning the Institute. IEEE Spectrum, 15 (10):26, October 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:SIk

[Ano78-255] Anonymous. Scanning the Institute. IEEE Spectrum, 15 (11):25, November 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:SIl

[Ano78-256] Anonymous. Scanning the Institute. IEEE Spectrum, 15 (12):19, December 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:SD

[Ano78-257] Anonymous. Special duplications. IEEE Spectrum, 15 (5):58–59, May 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1978:WMM

[Ano78-258] Anonymous. WESCON ’78: a major ‘microencounter’ [meeting preview]. IEEE Spectrum, 15 (8):53–55, August 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Aa

[Ano79a] Anonymous. [advertisement].

IEEE Spectrum, 16(1):1–2, January 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Ab

[Ano79b] Anonymous. [advertisement].

IEEE Spectrum, 16(1):2–7, January 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Ac

[Ano79c] Anonymous. [advertisement].

IEEE Spectrum, 16(1):121– 124, January 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Ad

[Ano79d] Anonymous. [advertisement].

IEEE Spectrum, 16(1):1–2, January 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Ae

[Ano79e] Anonymous. [advertisement]. IEEE Spectrum, 16(2):1–2, February 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Af

[Ano79f] Anonymous. [advertisement]. IEEE Spectrum, 16(2):2–7, February 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Ag

[Ano79g] Anonymous. [advertisement]. IEEE Spectrum, 16(2):102, February 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Ah

[Ano79h] Anonymous. [advertisement]. IEEE Spectrum, 16(2):1–4, February 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Ai

[Ano79i] Anonymous. [advertisement]. IEEE Spectrum, 16(3):1–2, March 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Aj

[Ano79j] Anonymous. [advertisement]. IEEE Spectrum, 16(3):2–7, March 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Ak

[Ano79k] Anonymous. [advertisement].

IEEE Spectrum, 16(3):108, March 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Al

[Ano79l] Anonymous. [advertisement].

IEEE Spectrum, 16(3):1–4, March 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Am

[Ano79m] Anonymous. [advertisement].

IEEE Spectrum, 16(4):1–2, April 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:An

[Ano79n] Anonymous. [advertisement].

IEEE Spectrum, 16(4):2–7, April 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Ao

[Ano79o] Anonymous. [advertisement].

IEEE Spectrum, 16(4):104–108, April 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Ap

[Ano79p] Anonymous. [advertisement].

IEEE Spectrum, 16(4):1–2, April 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Aq

[Ano79q] Anonymous. [advertisement]. IEEE Spectrum, 16(5):1–2, May 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Ar

[Ano79r] Anonymous. [advertisement]. IEEE Spectrum, 16(5):2–7, May 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:As

[Ano79s] Anonymous. [advertisement]. IEEE Spectrum, 16(5):100, May 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:At

[Ano79t] Anonymous. [advertisement]. IEEE Spectrum, 16(5):1–4, May 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Au

[Ano79u] Anonymous. [advertisement]. IEEE Spectrum, 16(6):1–2, June 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Av

[Ano79v] Anonymous. [advertisement]. IEEE Spectrum, 16(6):2–5, June 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Aw

[Ano79w] Anonymous. [advertisement].

IEEE Spectrum, 16(6):84B–88, June 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Ax

[Ano79x] Anonymous. [advertisement].

IEEE Spectrum, 16(6):1–2, June 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Ay

[Ano79y] Anonymous. [advertisement].

IEEE Spectrum, 16(7):1–2, July 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Az

[Ano79z] Anonymous. [advertisement].

IEEE Spectrum, 16(7):2–5, July 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Aba

[Ano79-27] Anonymous. [advertisement].

IEEE Spectrum, 16(8):1–2, August 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Abb

[Ano79-28] Anonymous. [advertisement].

IEEE Spectrum, 16(8):2–5, August 1979. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Abc

[Ano79-29] Anonymous. [advertisement]. IEEE Spectrum, 16(8):96–100, August 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Abd

[Ano79-30] Anonymous. [advertisement]. IEEE Spectrum, 16(8):1–2, August 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Abe

[Ano79-31] Anonymous. [advertisement]. IEEE Spectrum, 16(9):1–2, September 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Abf

[Ano79-32] Anonymous. [advertisement]. IEEE Spectrum, 16(9):2–3, September 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Abg

[Ano79-33] Anonymous. [advertisement]. IEEE Spectrum, 16(9):4–7, September 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Abh

[Ano79-34] Anonymous. [advertisement]. IEEE Spectrum, 16(9):1–4, September 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Abi

[Ano79-35] Anonymous. [advertisement].

IEEE Spectrum, 16(10):1–2, October 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Abj

[Ano79-36] Anonymous. [advertisement].

IEEE Spectrum, 16(10):2–6, October 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Abk

[Ano79-37] Anonymous. [advertisement].

IEEE Spectrum, 16(10):1–4, October 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Abl

[Ano79-38] Anonymous. [advertisement].

IEEE Spectrum, 16(11):1–2, November 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Abm

[Ano79-39] Anonymous. [advertisement].

IEEE Spectrum, 16(11):2–3, November 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Abn

[Ano79-40] Anonymous. [advertisement].

IEEE Spectrum, 16(11):1–4, November 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Abo

[Ano79-41] Anonymous. [advertisement]. IEEE Spectrum, 16(12):1–2, December 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Abp

[Ano79-42] Anonymous. [advertisement]. IEEE Spectrum, 16(12):2–3, December 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Abq

[Ano79-43] Anonymous. [advertisement]. IEEE Spectrum, 16(12):1–4, December 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ALa

[Ano79-44] Anonymous. Applications literature. IEEE Spectrum, 16 (8):88–90, August 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ALM

[Ano79-45] Anonymous. Applications literature: Microcomputer hardware and software explained. IEEE Spectrum, 16(2):96–101, February 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:AIa

[Ano79-46] Anonymous. Author index. IEEE Spectrum, 16(12):89–104, December 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:AIF

[Ano79-47] Anonymous. Awards: IEEE Field, Service, and Prize Paper Award winners. IEEE Spectrum, 16(8):60–63, August 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:AIM

[Ano79-48] Anonymous. Awards: IEEE major medalists 1979. IEEE Spectrum, 16(4):72–73, April 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:BBAd

[Ano79-49] Anonymous. Best bits: Applications of microprocessors. IEEE Spectrum, 16(7):16–17, July 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:BBAe

[Ano79-50] Anonymous. Best bits: Applications of microprocessors: a different robotic viewpoint. IEEE Spectrum, 16(9):18–19, September 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:BBAc

[Ano79-51] Anonymous. Best bits: Applications of microprocessors: Computers to fight hijackers. IEEE Spectrum, 16(5):20–24, May 1979. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:BBAf

[Ano79-52] Anonymous. Best bits: Applications of microprocessors: Keeping tabs on plutonium. IEEE Spectrum, 16(11):26–28, November 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:BBAa

[Ano79-53] Anonymous. Best bits: Applications of microprocessors: Pocket interpreter breaks communication impasse. IEEE Spectrum, 16(1):24–25, January 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:BBAb

[Ano79-54] Anonymous. Best bits: Applications of microprocessors: Teacher’s PET talks to logic analyzer. IEEE Spectrum, 16 (3):18–21, March 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:BRa

[Ano79-55] Anonymous. Book reviews. IEEE Spectrum, 16(1):95, January 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:BRb

[Ano79-56] Anonymous. Book reviews. IEEE Spectrum, 16(2):74–80, February 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:BRc

[Ano79-57] Anonymous. Book reviews.

IEEE Spectrum, 16(3):80–83, March 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:BRd

[Ano79-58] Anonymous. Book reviews.

IEEE Spectrum, 16(4):82–83, April 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:BRe

[Ano79-59] Anonymous. Book reviews.

IEEE Spectrum, 16(5):76–79, May 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:BRf

[Ano79-60] Anonymous. Book reviews.

IEEE Spectrum, 16(6):61–65, June 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:BRg

[Ano79-61] Anonymous. Book reviews.

IEEE Spectrum, 16(7):71–72, July 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:BRh

[Ano79-62] Anonymous. Book reviews.

IEEE Spectrum, 16(8):72–73, August 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:BRi

[Ano79-63] Anonymous. Book reviews. IEEE Spectrum, 16(9):78–80, September 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:BRj

[Ano79-64] Anonymous. Book reviews. IEEE Spectrum, 16(10):96– 97, October 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:BRk

[Ano79-65] Anonymous. Book reviews. IEEE Spectrum, 16(11):129– 130, November 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:BRl

[Ano79-66] Anonymous. Book reviews. IEEE Spectrum, 16(12):67–68, December 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:SHR

[Ano79-67] Anonymous. Spectrum’s hardware review. IEEE Spectrum, 16(2):96, February 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Cb

[Ano79-68] Anonymous. Calendar. IEEE Spectrum, 16(1):8, January 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Cd

[Ano79-69] Anonymous. Calendar. IEEE Spectrum, 16(2):8, February 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Cf

[Ano79-70] Anonymous. Calendar. IEEE Spectrum, 16(3):8, March 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Ch

[Ano79-71] Anonymous. Calendar. IEEE Spectrum, 16(4):8, April 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Cj

[Ano79-72] Anonymous. Calendar. IEEE Spectrum, 16(5):8, May 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Cl

[Ano79-73] Anonymous. Calendar. IEEE Spectrum, 16(6):6–7, June 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Cn

[Ano79-74] Anonymous. Calendar. IEEE Spectrum, 16(7):6–7, July 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Cp

[Ano79-75] Anonymous. Calendar. IEEE Spectrum, 16(8):6–7, August 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Cr

[Ano79-76] Anonymous. Calendar. IEEE Spectrum, 16(9):8, September 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Cu

[Ano79-77] Anonymous. Calendar. IEEE Spectrum, 16(10):8, October 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Cw

[Ano79-78] Anonymous. Calendar. IEEE Spectrum, 16(11):4–5, November 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Cy

[Ano79-79] Anonymous. Calendar. IEEE Spectrum, 16(12):4–9, December 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:CAa

[Ano79-80] Anonymous. Classified advertising. IEEE Spectrum, 16(7):87– 91, July 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:CAb

[Ano79-81] Anonymous. Classified advertising. IEEE Spectrum, 16 (8):91–95, August 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:CAc

[Ano79-82] Anonymous. Classified advertising. IEEE Spectrum, 16(11): 137–144, November 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:CSa

[Ano79-83] Anonymous. Coming in Spectrum. IEEE Spectrum, 16(1): 26–28, January 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:CSb

[Ano79-84] Anonymous. Coming in Spectrum. IEEE Spectrum, 16(2): 20–22, February 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:CSc

[Ano79-85] Anonymous. Coming in Spectrum. IEEE Spectrum, 16 (3):24–26, March 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:CSd

[Ano79-86] Anonymous. Coming in Spectrum. IEEE Spectrum, 16 (4):26–28, April 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:CSe

[Ano79-87] Anonymous. Coming in Spectrum. IEEE Spectrum, 16(5):19, May 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:CSf

[Ano79-88] Anonymous. Coming in Spectrum. IEEE Spectrum, 16 (6):14–16, June 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:CSg

[Ano79-89] Anonymous. Coming in Spectrum. IEEE Spectrum, 16(7):18– 20, July 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:CSh

[Ano79-90] Anonymous. Coming in Spectrum. IEEE Spectrum, 16 (8):15–16, August 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:CSi

[Ano79-91] Anonymous. Coming in Spectrum. IEEE Spectrum, 16(9): 22–23, September 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:CSj

[Ano79-92] Anonymous. Coming in Spectrum. IEEE Spectrum, 16(10): 20–26, October 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:CSk

[Ano79-93] Anonymous. Coming in Spectrum. IEEE Spectrum, 16 (11):16, November 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:CSl

[Ano79-94] Anonymous. Coming in Spectrum. IEEE Spectrum, 16(12): 22–23, December 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:CBS

[Ano79-95] Anonymous. Communications.

On the Bell System: Boom or ‘bust? For all it’s done for communications and the economy, its critics recommend its breakup. IEEE Spectrum, 16(12):42, December 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Ca

[Ano79-96] Anonymous. Contents. IEEE Spectrum, 16(1):1, January 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Cc

[Ano79-97] Anonymous. Contents. IEEE Spectrum, 16(2):1, February 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Ce

[Ano79-98] Anonymous. Contents. IEEE Spectrum, 16(3):1, March 1979.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Cg

[Ano79-99] Anonymous. Contents. IEEE Spectrum, 16(4):1, April 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Ci

[Ano79-100] Anonymous. Contents. IEEE Spectrum, 16(5):1, May 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Ck

[Ano79-101] Anonymous. Contents. IEEE Spectrum, 16(6):1, June 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Cm

[Ano79-102] Anonymous. Contents. IEEE Spectrum, 16(7):1, July 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Co

[Ano79-103] Anonymous. Contents. IEEE Spectrum, 16(8):1, August 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Cq

[Ano79-104] Anonymous. Contents. IEEE Spectrum, 16(9):1, September 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Ct

[Ano79-105] Anonymous. Contents. IEEE Spectrum, 16(10):1, October 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Cv

[Ano79-106] Anonymous. Contents. IEEE Spectrum, 16(11):1, November 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Cx

[Ano79-107] Anonymous. Contents. IEEE Spectrum, 16(12):1, December 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ETTa

[Ano79-108] Anonymous. EEs’ tools toys.

IEEE Spectrum, 16(2):18–19, February 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ETTb

[Ano79-109] Anonymous. EEs’ tools toys.

IEEE Spectrum, 16(4):22–25, April 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ETTc

[Ano79-110] Anonymous. EEs’ tools toys.

IEEE Spectrum, 16(6):12–13, June 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ETTd

[Ano79-111] Anonymous. EEs’ tools toys. IEEE Spectrum, 16(8):14, August 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ETTe

[Ano79-112] Anonymous. EEs’ tools toys. IEEE Spectrum, 16(10):16– 19, October 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ETTf

[Ano79-113] Anonymous. EEs’ tools toys. IEEE Spectrum, 16(12):24–26, December 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ERa

[Ano79-114] Anonymous. Energy report. IEEE Spectrum, 16(1):10–11, January 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ERb

[Ano79-115] Anonymous. Energy report. IEEE Spectrum, 16(2):10–11, February 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ERc

[Ano79-116] Anonymous. Energy report. IEEE Spectrum, 16(3):10–11, March 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ERd

[Ano79-117] Anonymous. Energy report.

IEEE Spectrum, 16(4):10–10H, April 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ERe

[Ano79-118] Anonymous. Energy report.

IEEE Spectrum, 16(5):10, May 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ERf

[Ano79-119] Anonymous. Energy report.

IEEE Spectrum, 16(6):10–11, June 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ERg

[Ano79-120] Anonymous. Energy report.

IEEE Spectrum, 16(7):10–13, July 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ERh

[Ano79-121] Anonymous. Energy report.

IEEE Spectrum, 16(8):10–13, August 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ERi

[Ano79-122] Anonymous. Energy report.

IEEE Spectrum, 16(9):10–11, September 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ERj

[Ano79-123] Anonymous. Energy report. IEEE Spectrum, 16(10):10– 11, October 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ERk

[Ano79-124] Anonymous. Energy report. IEEE Spectrum, 16(11):18–19, November 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ERl

[Ano79-125] Anonymous. Energy report. IEEE Spectrum, 16(12):18–21, December 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ELa

[Ano79-126] Anonymous. The engineer at large. IEEE Spectrum, 16(1): 22–23, January 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ELb

[Ano79-127] Anonymous. The engineer at large. IEEE Spectrum, 16 (3):22–23, March 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ELc

[Ano79-128] Anonymous. The engineer at large. IEEE Spectrum, 16 (5):16–18, May 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ELd

[Ano79-129] Anonymous. The engineer at large. IEEE Spectrum, 16(7):14– 15, July 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ELe

[Ano79-130] Anonymous. The engineer at large. IEEE Spectrum, 16(9): 16–17, September 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ELf

[Ano79-131] Anonymous. The engineer at large. IEEE Spectrum, 16(11): 22–25, November 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Fa

[Ano79-132] Anonymous. Forum. IEEE Spectrum, 16(1):12–15, January 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Fb

[Ano79-133] Anonymous. Forum. IEEE Spectrum, 16(2):12–17, February 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Fc

[Ano79-134] Anonymous. Forum. IEEE Spectrum, 16(3):12–17, March 1979.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Fd

[Ano79-135] Anonymous. Forum. IEEE Spectrum, 16(4):11–21, April 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Fe

[Ano79-136] Anonymous. Forum. IEEE Spectrum, 16(5):11–15, May 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Ff

[Ano79-137] Anonymous. Forum. IEEE Spectrum, 16(6):8, June 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Fg

[Ano79-138] Anonymous. Forum. IEEE Spectrum, 16(7):8, July 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Fh

[Ano79-139] Anonymous. Forum. IEEE Spectrum, 16(8):8, August 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Fi

[Ano79-140] Anonymous. Forum. IEEE Spectrum, 16(9):12–15, September 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Fj

[Ano79-141] Anonymous. Forum. IEEE Spectrum, 16(10):12–15, October 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Fk

[Ano79-142] Anonymous. Forum. IEEE Spectrum, 16(11):6–15, November 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Fl

[Ano79-143] Anonymous. Forum. IEEE Spectrum, 16(12):10–16, December 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:GA

[Ano79-144] Anonymous. Guide to authors. IEEE Spectrum, 16(1): 87–89, January 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:HPA

[Ano79-145] Anonymous. Hew product applications: Light-polarizing gates provide high linearity for direct-writing analog data recorder. IEEE Spectrum, 16 (9):81–84, September 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ITCa

[Ano79-146] Anonymous. IEEE tables of contents. IEEE Spectrum, 16 (1):92–94, January 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ITCb

[Ano79-147] Anonymous. IEEE tables of contents. IEEE Spectrum, 16 (2):71–73, February 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ITCc

[Ano79-148] Anonymous. IEEE tables of contents. IEEE Spectrum, 16 (3):76–79, March 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ITCd

[Ano79-149] Anonymous. IEEE tables of contents. IEEE Spectrum, 16 (4):80–81, April 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ITCe

[Ano79-150] Anonymous. IEEE tables of contents. IEEE Spectrum, 16 (5):72–75, May 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ITCf

[Ano79-151] Anonymous. IEEE tables of contents. IEEE Spectrum, 16 (6):56–57, June 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ITCi

[Ano79-152] Anonymous. IEEE tables of contents. IEEE Spectrum, 16(9):

72–77, September 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ITCj

[Ano79-153] Anonymous. IEEE tables of contents. IEEE Spectrum, 16 (10):84–95, October 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ITCk

[Ano79-154] Anonymous. IEEE tables of contents. IEEE Spectrum, 16(11): 114–122, November 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ITCl

[Ano79-155] Anonymous. IEEE tables of contents. IEEE Spectrum, 16(12): 60–66, December 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ITCg

[Ano79-156] Anonymous. IEEE tables of contents for current and future publications. IEEE Spectrum, 16(7):56–70, July 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:ITCh

[Ano79-157] Anonymous. IEEE tables of contents or current and future publications. IEEE Spectrum, 16 (8):66–71, August 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:IA

[Ano79-158] Anonymous. Index to advertisers. IEEE Spectrum, 16(9): 106, September 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:MPE

[Ano79-159] Anonymous. Meeting preview: Electro ’79: LSI, the 80s, and a centennial celebration: higher LSI packing densities at reduced costs are stressed at Electro ’79 — opening April 24 in New York. IEEE Spectrum, 16(4):74–77, April 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:MPW

[Ano79-160] Anonymous. Meeting preview: WESCON ’79: Micro basics on the Bay: California opens its Golden Gate to the hightechnology electronics of the 1980s. IEEE Spectrum, 16(9): 66–69, September 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Ma

[Ano79-161] Anonymous. Meetings. IEEE Spectrum, 16(1):96–107, January 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Mb

[Ano79-162] Anonymous. Meetings. IEEE Spectrum, 16(2):81–86, February 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Mc

[Ano79-163] Anonymous. Meetings. IEEE Spectrum, 16(3):84–93, March 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Md

[Ano79-164] Anonymous. Meetings. IEEE Spectrum, 16(4):84–90, April 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Me

[Ano79-165] Anonymous. Meetings. IEEE Spectrum, 16(5):80–86, May 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Mf

[Ano79-166] Anonymous. Meetings. IEEE Spectrum, 16(6):58–60, June 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Mg

[Ano79-167] Anonymous. Meetings. IEEE Spectrum, 16(7):82–86, July 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Mh

[Ano79-168] Anonymous. Meetings. IEEE Spectrum, 16(8):74–79, August 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Mi

[Ano79-169] Anonymous. Meetings. IEEE Spectrum, 16(9):94–106, September 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Mj

[Ano79-170] Anonymous. Meetings. IEEE Spectrum, 16(10):104–106, October 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Mk

[Ano79-171] Anonymous. Meetings. IEEE Spectrum, 16(11):131–136, November 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:Ml

[Ano79-172] Anonymous. Meetings. IEEE Spectrum, 16(12):69–71, December 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NRIa

[Ano79-173] Anonymous. New and recent IEEE publications. IEEE Spectrum, 16(1):90–91, January 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NRIb

[Ano79-174] Anonymous. New and recent IEEE publications. IEEE Spectrum, 16(2):69–70, February 1979. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NRIc

[Ano79-175] Anonymous. New and recent IEEE publications. IEEE Spectrum, 16(3):74–75, March 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NRId

[Ano79-176] Anonymous. New and recent IEEE publications. IEEE Spectrum, 16(4):78–79, April 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NRIe

[Ano79-177] Anonymous. New and recent IEEE publications. IEEE Spectrum, 16(5):70–71, May 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NRIf

[Ano79-178] Anonymous. New and recent IEEE publications. IEEE Spectrum, 16(6):54–55, June 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NRIg

[Ano79-179] Anonymous. New and recent IEEE publications. IEEE Spectrum, 16(7):64–65, July 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NRIh

[Ano79-180] Anonymous. New and recent IEEE publications. IEEE Spectrum, 16(8):64–65, August 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NRIi

[Ano79-181] Anonymous. New and recent IEEE publications. IEEE Spectrum, 16(9):70–71, September 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NRIj

[Ano79-182] Anonymous. New and recent IEEE publications. IEEE Spectrum, 16(10):82–83, October 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NRIk

[Ano79-183] Anonymous. New and recent IEEE publications. IEEE Spectrum, 16(11):112–113, November 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NRIl

[Ano79-184] Anonymous. New and recent IEEE publications. IEEE Spectrum, 16(12):58–59, December 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NPAg

[Ano79-185] Anonymous. New product applications. IEEE Spectrum, 16 (7):73–74, July 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NPAh

[Ano79-186] Anonymous. New product applications. IEEE Spectrum, 16 (8):80–82P, August 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NPAk

[Ano79-187] Anonymous. New product applications. IEEE Spectrum, 16(12): 72–74FF, December 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NPAc

[Ano79-188] Anonymous. New product applications: 1-GHz bandwidth, highest so far, achieved in a laboratory oscilloscope. IEEE Spectrum, 16(3):94–98d, March 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NPAi

[Ano79-189] Anonymous. New product applications: 1-MBd modems link microcomputers and minicomputers at up to 10,440 m over coaxial cable. IEEE Spectrum, 16 (10):98–103, October 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NPAd

[Ano79-190] Anonymous. New product applications: 12-MHz single- and dual-trace oscilloscopes are the smallest professional-grade instruments. IEEE Spectrum, 16 (4):91–94L, April 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NPAe

[Ano79-191] Anonymous. New product applications: Low-price, portable, 3 1/2-digit multimeters designed for a wide range of applications. IEEE Spectrum, 16 (5):87–90, May 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NPAa

[Ano79-192] Anonymous. New product applications: Real-time measurement and control system with keyboard and CRT features integral sensor I/O. IEEE Spectrum, 16 (1):108–110, January 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NPAb

[Ano79-193] Anonymous. New product applications: Total 16-bit A/D converter is offered in a single 32pin dual-in-line package. IEEE Spectrum, 16(2):87–92L, February 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NPAf

[Ano79-194] Anonymous. New product applications: Universal development system supports all popular 8-bit microprocessor systems. IEEE Spectrum, 16 (6):66–74, June 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NPAj

[Ano79-195] Anonymous. New product applications: Winchester fixed disk drive series features 5.3 — and 10.7-megabyte unformatted storage capacity. IEEE Spectrum, 16(11):123–124, November 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NIa

[Ano79-196] Anonymous. News from industry. IEEE Spectrum, 16 (1):111, January 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NIb

[Ano79-197] Anonymous. News from industry. IEEE Spectrum, 16 (2):93, February 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NIc

[Ano79-198] Anonymous. News from industry. IEEE Spectrum, 16(3):99, March 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NId

[Ano79-199] Anonymous. News from industry. IEEE Spectrum, 16(4):95, April 1979. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NIe

[Ano79-200] Anonymous. News from industry. IEEE Spectrum, 16(5):91, May 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NIf

[Ano79-201] Anonymous. News from industry. IEEE Spectrum, 16(6):75, June 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NIg

[Ano79-202] Anonymous. News from industry. IEEE Spectrum, 16(7):75, July 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NIh

[Ano79-203] Anonymous. News from industry. IEEE Spectrum, 16(8):83, August 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NIi

[Ano79-204] Anonymous. News from industry. IEEE Spectrum, 16 (9):85, September 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NIj

[Ano79-205] Anonymous. News from industry. IEEE Spectrum, 16 (10):107, October 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NIk

[Ano79-206] Anonymous. News from industry. IEEE Spectrum, 16 (11):125, November 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NIl

[Ano79-207] Anonymous. News from industry. IEEE Spectrum, 16 (12):75, December 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NRc

[Ano79-208] Anonymous. News from region 1.

IEEE Spectrum, 16(3):100–107, March 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NRg

[Ano79-209] Anonymous. News from region 1. IEEE Spectrum, 16(7):76–81, July 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NRk

[Ano79-210] Anonymous. News from region 1. IEEE Spectrum, 16(10): 108–121, October 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NRl

[Ano79-211] Anonymous. News from region 1. IEEE Spectrum, 16(11):

126–128, November 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NRh

[Ano79-212] Anonymous. News from region 5. IEEE Spectrum, 16(8):84–87, August 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NRa

[Ano79-213] Anonymous. News from the regions. IEEE Spectrum, 16(1): 112–120, January 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NRb

[Ano79-214] Anonymous. News from the regions. IEEE Spectrum, 16(2): 94–95, February 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NRd

[Ano79-215] Anonymous. News from the regions. IEEE Spectrum, 16 (4):96–103, April 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NRe

[Ano79-216] Anonymous. News from the regions. IEEE Spectrum, 16 (5):92–99, May 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NRf

[Ano79-217] Anonymous. News from the regions. IEEE Spectrum, 16 (6):76–83, June 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NRj

[Ano79-218] Anonymous. News from the regions. IEEE Spectrum, 16(9): 86–93, September 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NRm

[Ano79-219] Anonymous. News from the regions. IEEE Spectrum, 16(12): 76–88, December 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NWa

[Ano79-220] Anonymous. News from Washington. IEEE Spectrum, 16 (1):9, January 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NWb

[Ano79-221] Anonymous. News from Washington. IEEE Spectrum, 16 (2):9, February 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NWc

[Ano79-222] Anonymous. News from Washington. IEEE Spectrum, 16(3):9, March 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NWd

[Ano79-223] Anonymous. News from Washington. IEEE Spectrum, 16(4):9, April 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NWe

[Ano79-224] Anonymous. News from Washington. IEEE Spectrum, 16(5): 9, May 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NWf

[Ano79-225] Anonymous. News from Washington. IEEE Spectrum, 16(6): 9, June 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NWg

[Ano79-226] Anonymous. News from Washington. IEEE Spectrum, 16(7): 9, July 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NWh

[Ano79-227] Anonymous. News from Washington. IEEE Spectrum, 16(8):9, August 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NWi

[Ano79-228] Anonymous. News from Washington. IEEE Spectrum, 16 (9):9, September 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NWj

[Ano79-229] Anonymous. News from Washington. IEEE Spectrum, 16 (10):9, October 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NWk

[Ano79-230] Anonymous. News from Washington. IEEE Spectrum, 16 (11):17, November 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:NWl

[Ano79-231] Anonymous. News from Washington. IEEE Spectrum, 16 (12):17, December 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:OAI

[Ano79-232] Anonymous. Osborne & Associates: An introduction to Microcomputers: Volume 2 and volume 3. There’s nothing else like them. IEEE Spectrum, 16 (1):16–21, January 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:PSa

[Ano79-233] Anonymous. Product selector. IEEE Spectrum, 16(7):92– 96, July 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:PSb

[Ano79-234] Anonymous. Product selector.

IEEE Spectrum, 16(10):1, October 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:PCE

[Ano79-235] Anonymous. Professional concerns: Employment guidelines revisited: a look at the guidelines’ second edition; it has been accepted by employers and employees alike. IEEE Spectrum, 16(3):70–73, March 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:SSA

[Ano79-236] Anonymous. SB. Security Agency denies tampering with DES. IEEE Spectrum, 16(7):39, July 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:SIa

[Ano79-237] Anonymous. Scanning the Institute. IEEE Spectrum, 16 (1):26, January 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:SIb

[Ano79-238] Anonymous. Scanning the Institute. IEEE Spectrum, 16 (2):20, February 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:SIc

[Ano79-239] Anonymous. Scanning the Institute. IEEE Spectrum, 16(3):24, March 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:SId

[Ano79-240] Anonymous. Scanning the Institute. IEEE Spectrum, 16(4):26, April 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:SIe

[Ano79-241] Anonymous. Scanning the Institute. IEEE Spectrum, 16(5):19, May 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:SIf

[Ano79-242] Anonymous. Scanning the Institute. IEEE Spectrum, 16(6):14, June 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:SIg

[Ano79-243] Anonymous. Scanning the Institute. IEEE Spectrum, 16(7):18, July 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:SIh

[Ano79-244] Anonymous. Scanning the Institute. IEEE Spectrum, 16(8):15, August 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:SIi

[Ano79-245] Anonymous. Scanning the Institute. IEEE Spectrum, 16 (9):22, September 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:SIj

[Ano79-246] Anonymous. Scanning the Institute. IEEE Spectrum, 16 (10):20, October 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:SIk

[Ano79-247] Anonymous. Scanning the Institute. IEEE Spectrum, 16 (11):16, November 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:SIl

[Ano79-248] Anonymous. Scanning the Institute. IEEE Spectrum, 16 (12):22, December 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Anonymous:1979:SSM

[Ano79-249] Anonymous. Space Shuttle maiden. IEEE Spectrum, 16 (7):48–50, July 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ashley:1971:TSM

[ASM71] J. R. Ashley, T. A. Saponas, and R. C. Matson. Test signals for music reproduction systems. IEEE Spectrum, 8(7):53– 61, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Andrews:1972:IPD

[ATK72] H. C. Andrews, A. G. Tescher, and R. P. Kruger. Image processing by digital computer. IEEE Spectrum, 9(7):20–32, July 1972.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Backe:1972:PRW

[Bac72] R. J. Backe. Pension reform, on the way, needs pushing. IEEE Spectrum, 9(11):62–68, November 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Barton:1972:WML

[Bar72] D. M. Barton. Why multiplex LEDs? IEEE Spectrum, 9(11): 30–32, November 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Barnes:1973:ESI

[Bar73] F. S. Barnes. Engineering at the sociotechnical interface. IEEE Spectrum, 10(10):63–65, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Baruch:1978:SPU

[Bar78] J. J. Baruch. On stimulating productivity in the U.S.: The need to encourage innovation is broadly supported, and a program is sought. IEEE Spectrum, 15(10):47–48, October 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bargellini:1979:CUS

[Bar79] P. L. Bargellini. Commercial U.S. satellites: a survey of communication satellites, their effect on domestic networks, and a look into the near and longterm future. IEEE Spectrum, 16 (10):30–32, October 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bates:1971:UTD

[Bat71] J. J. Bates. Using thyristors and diodes to improve commutation. IEEE Spectrum, 8(1):38–47, January 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Beakley:1971:IUC

[BB71] G. C. Beakley and J. F. Bregar. Industry-university cooperation: a way to stem the loss of freshman engineering students. IEEE Spectrum, 8(2): 92–95, February 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bartolini:1978:ODS

[BBF^{+}78] R. A. Bartolini, A. E. Bell, R. E. Flory, M. Lurie, and F. W. Spong. Optical disk systems emerge. IEEE Spectrum, 15 (8):20–28, August 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bhagat:1975:RME

[BBG75] N. A. Bhagat, J. J. Baruch, and E. K. Gannett. Retrieval made easy. IEEE Spectrum, 12 (2):61–62, February 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Berger:1977:F

[BBW^{+}77] M. Berger, J. A. Brockwell, H. A.

Wheeler, D. J. Comer, W. Pferd, E. C. Wentz, A. E. Emanuel, F. C. Foss, J. S. Krikorian, and P. C. deLuca. Forum. IEEE Spectrum, 14(11):12–16, November 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Backe:1973:PBL

[BC73a] R. J. Backe and F. Cummings.

Pensions: the big lottery. IEEE Spectrum, 10(6):59–61, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Backe:1973:YYP

[BC73b] R. J. Backe and F. Cummings. You and your pension. IEEE Spectrum, 10(5):55–58, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brogan:1977:DOF

[BC77] J. J. Brogan and C. S. Chen.

DOE: Outlook on fuel economy: a three-pronged program that could save hundreds of thousands of barrels of oil daily. IEEE Spectrum, 14(11):63–64, November 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bayless:1973:VSB

[BCG73] J. W. Bayless, S. J. Campanella, and A. J. Goldberg. Voice signals: bit-by-bit. IEEE Spectrum, 10(10):28–34, October 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bulthuis:1979:CTB

[BCH^{+}79] K. Bulthuis, M. G. Carasso, J. P. J. Heemskerk, P. J. Kivits, W. J. Kleuters, and P. Zalm. Communications: Ten billion bits on a disk: High-power laser diodes and pregrooved disks enable low-cost digital data recording and retrieval. IEEE Spectrum, 16(8):26–33, August 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Backe:1975:CQP

[BCY75] R. J. Backe, F. Cummings, and L. Young. Careers: Q&A on pensions: Institute experts answer members’ questions concerning the new U.S. pension law and its effect on engineers. IEEE Spectrum, 12(8):58–61, August 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Balzhiser:1978:PEE

[BDL^{+}78] R. Balzhiser, J. Dougherty, M. Levenson, R. Males, R. Rudman, C. Starr, D. Christiansen, and L. King. Power/energy: EPRI looks inward, and ahead: The Electric Power Research Institute’s senior staff, in a candid session, probes its programs and progress. IEEE Spectrum, 15(2):48–53, February 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Beardsley:1971:CAI

[Bea71] C. W. Beardsley. Computer aids for IC design, artwork, and mask generation. IEEE Spectrum, 8 (9):63–79, September 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Beardsley:1972:F

[Bea72a] C. Beardsley. Forecast for ’72. IEEE Spectrum, 9(3):s4–s8, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Beardsley:1972:WS

[Bea72b] C. Beardsley. Wanted: superbatteries. IEEE Spectrum, 9(7):42, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Beardsley:1972:YCI

[Bea72c] C. W. Beardsley. Is your computer insecure? IEEE Spectrum, 9(1):67–78, January 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Beardsley:1972:KTY

[Bea72d] C. W. Beardsley. Keeping on top of your field. IEEE Spectrum, 9 (12):68–71, December 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Beardsley:1972:NOB

[Bea72e] C. W. Beardsley. A new option for back-to-schoolers. IEEE Spectrum, 9(10):57–59, October 1972. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Beardsley:1972:SSC

[Bea72f] C. W. Beardsley. A slow start in a cold market. IEEE Spectrum, 9 (9):22–31, September 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Beardsley:1973:CSJ

[Bea73a] C. W. Beardsley. The changing style of job hunting. IEEE Spectrum, 10(2):59–61, February 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Beardsley:1973:FD

[Bea73b] C. W. Beardsley. The future of discretes. IEEE Spectrum, 10 (1):36–42, January 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Beam:1974:CMI

[Bea74a] W. R. Beam. Careers: The mobilengineer: The IEEE member is a mover: Geographically, jobwise, and, sometimes, upward in management. IEEE Spectrum, 11(3):88–91, March 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Beam:1974:CRV

[Bea74b] W. R. Beam. Computer report VIII: an $88 book shelf: Eleven volumes, selected from hundreds available, form this digital engineers’ reference and study library. IEEE Spectrum, 11(2):

82–84, February 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Beardstey:1975:LTL

[Bea75] C. W. Beardstey. Let there be light, but just enough: The end of cheap energy gave the lighting industry a corporate headache and the user more lumens per watt. IEEE Spectrum, 12(12): 28–34, December 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Beckmann:1971:CE

[Bec71] P. Beckmann. Computerization of english. IEEE Spectrum, 8 (12):20–27, December 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bell:1974:CRI

[Bel74] C. G. Bell. Computer report II: More power by networking: Whether computers communicate as equals or in a superiorsubordinate mode, the outcome is usually positive. IEEE Spectrum, 11(2):40–45, February 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Benedict:1970:EGP

[Ben70]  M. Benedict.  Electric gen-

erating prospects for nuclear power. IEEE Spectrum, 7(7):24– 30, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bertram:1971:CPE

[Ber71] S. Bertram. Compensating for propagation errors in electromagnetic measuring systems. IEEE Spectrum, 8(3):58–63, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernstein:1973:GS

[Ber73] T. Bernstein. A grand success. IEEE Spectrum, 10(2): 54–58, February 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bernhard:1979:CEH

[Ber79] R. Bernhard. Consumer electronics. Higher fi by digits: Digitally encoded audio disks and associated playback gear promise breakthrough in high fidelity, low noise. IEEE Spectrum, 16(12): 28–32, December 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Barry:1978:PEG

[BFM78] E. P. Barry, R. L. A. Fernandes, and W. A. Messner. Power/ energy: a giant step planned in fuel-cell plant test: a demonstration program for a prototype 4.8-MW dc fuel cell is underway for New York’s Consolidated Edison electrical system. IEEE Spectrum, 15(11):47–55, November 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Branin:1971:EIE

[BHLK71] F. H. Branin, G. R. Hogsett, R. L. Lunde, and L. E. Kugel. Ecap II: An electronic circuit analysis program. IEEE Spectrum, 8(6):14–25, June 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bianculli:1970:SMA

[Bia70] A. J. Bianculli. Stepper motors:

application and selection. IEEE Spectrum, 7(12):25–29, December 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bibbero:1971:SATa

[Bib71a] R. J. Bibbero. Systems approach toward nationwide air-pollution control I. The problem, the system, the objective. IEEE Spectrum, 8(10):20–31, October 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bibbero:1971:SATb

[Bib71b] R. J. Bibbero. Systems approach toward nationwide air-pollution control II. The technical requirements. IEEE Spectrum, 8(11): 73–81, November 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bibbero:1971:SATc

[Bib71c] R. J. Bibbero. Systems approach toward nationwide airpollution control III. Mathematical models. IEEE Spectrum, 8 (12):47–58, December 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Berndt:1978:HEC

[BJ78] E. R. Berndt and D. W. Jorgenson. How energy, and its cost, enter the ‘productivity equation’. IEEE Spectrum, 15(10): 50–52, October 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Birk:1979:PSS

[BKS79] J. R. Birk, K. Klunder, and J. C. Smith. Power sources: Superbatteries: a progress report: Developers are banking on new materials to drive batteries into a competitive position vis-a-vis oil, gas, and gasoline. IEEE Spectrum, 16(3):49–56, March 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Black:1977:INF

[Bla77] H. S. Black. Inventing the negative feedback amplifier: Six years of persistent search helped the author conceive the idea in a flash aboard the old Lackawanna Ferry. IEEE Spectrum, 14(12): 55–60, December 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Burns:1977:F

[BLD^{+}77] J. J. Burns, B. Levine, D. Dies, L. Dwon, M. K. Van Valkenburg, D. W. Glover, and D. P. Morison. Forum. IEEE Spectrum, 14(8):15–20, August 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Boaen:1973:DLC

[Boa73]  V. Boaen.  Designing logic

circuits for high noise immunity. IEEE Spectrum, 10(1): 53–59, January 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bordman:1973:IAB

[Bor73] S. L. Bordman. Improving the accuracy of benefit-cost analysis. IEEE Spectrum, 10(9): 72–76, September 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Braun:1977:TSM

[Bra77] R. L. Braun. Toward safer motor vehicles: Many fatal accidents can be survived, or avoided altogether, by adopting vehicle designs that stress safety over styling. IEEE Spectrum, 14(11): 81–86, November 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Branstad:1979:VHD

[Bra79] D. Branstad. V. ‘Hellman’s data does not support his conclusion’. IEEE Spectrum, 16(7): 41, July 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brewer:1971:AEI

[Bre71] G. R. Brewer. The application of electron/ion beam technology to microelectronics. IEEE Spectrum, 8(1):23–37, January 1971.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Breikss:1975:CHD

[Bre75] I. P. Breikss. Computers: Highdensity data recording: Very high bit rates in the acquisition of digitized data are possible using this new technique. IEEE Spectrum, 12(5):58–62, May 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brown:1973:SPS

[Bro73] W. C. Brown. Satellite power stations: a new source of energy? IEEE Spectrum, 10 (3):38–47, March 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Brown:1979:PES

[Bro79] W. C. Brown. Power/energy: Solar power satellites: Microwaves deliver the power: Kilometerdiameter antennas would be used in this super system to transmit power to ground-based rectifying antennas. IEEE Spectrum, 16(6):36–43, June 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Boyle:1971:CCD

[BS71] W. S. Boyle and G. E. Smith. Charge-coupled devices — a new approach to MIS device structures. IEEE Spectrum, 8(7):18– 27, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bucy:1972:SGM

[Buc72]  J. F. Bucy.  For semicon-

ductors: growth, not maturity. IEEE Spectrum, 9(4):55–58, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bucy:1978:EFM

[Buc78a] J. F. Bucy. Exploding a few myths about productivity and presenting a formula for the future. IEEE Spectrum, 15(10): 45–47, October 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Bucy:1978:TIS

[Buc78b] J. F. Bucy. Texas Instruments’ short-term ‘Idea’ programs and its long-shot ‘Wild Hare’ programs stimulate the company’s innovators. IEEE Spectrum, 15 (10):78–80, October 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Buss:1972:ILB

[Bus72] L. A. Buss. IRAC looks back:

and ahead. IEEE Spectrum, 9(6):56–57, June 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Camras:1977:SAA

[Cam77] M. Camras. State of the audiotape art: Amateurs and professionals find their favorite tape formats enhanced by better electronics, mechanics, and magnetic oxides. IEEE Spectrum, 14 (10):28–35, October 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Camana:1979:VBC

[Cam79] P. Camana. Video-bandwidth compression: a study in tradeoffs. IEEE Spectrum, 16(6): 24–29, June 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Carne:1979:WHT

[Car79] E. B. Carne. The wired household: Teletext services pave the way for a variety of useful monitoring and control features in the home. IEEE Spectrum, 16 (10):61–66, October 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cummings:1978:PER

[CC78] F. H. Cummings and P. B. Clark. Power/energy: a rational approach to power plant siting: a technique that integrates economic, environmental, and sitespecific data could result in better decisions. IEEE Spectrum, 15 (12):48–52, December 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chi:1972:STC

[CF72] A. R. Chi and H. S. Fosque. A step in time changes in standard frequency and time signal broadcasts. IEEE Spectrum, 9 (1):82–86, January 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chestnut:1973:S

[Che73] H. Chestnut. Scoreboard ’73.

IEEE Spectrum, 10(8):57–59, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Chien:1973:MEC

[Chi73] R. T. Chien. Memory error control: beyond parity. IEEE Spectrum, 10(7):18–23, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1972:E

[Chr72] D. Christiansen. Electronics 1985. IEEE Spectrum, 9(7):50– 54, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1973:JGI

[Chr73] D. Christiansen. Japan goes industrial. IEEE Spectrum, 10 (12):44–55, December 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1974:PSC

[Chr74a] D. Christiansen. Power special:

The critical years. IEEE Spectrum, 11(6):38–39, June 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1974:SLSa

[Chr74b] D. Christiansen. Spectral lines: a sensitive topic. IEEE Spectrum, 11(10):41, October 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1974:SLEa

[Chr74c] D. Christiansen. Spectral lines: Energy crisis: Seeking the real alternatives. IEEE Spectrum, 11(1):29, January 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1974:SLEb

[Chr74d] D. Christiansen. Spectral lines: Energy crisis: Seeking the real alternatives, II. IEEE Spectrum, 11(2):33, February 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1974:SLEc

[Chr74e] D. Christiansen. Spectral lines: Energy crisis: Seeking the real alternatives III. IEEE Spectrum, 11(3):43, March 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1974:SLEd

[Chr74f] D. Christiansen. Spectral lines: Energy crisis: Seeking the real alternatives, IV. IEEE Spectrum, 11(4):33, April 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1974:SLL

[Chr74g] D. Christiansen. Spectral lines: Learning by undoing. IEEE Spectrum, 11(5):33, May 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1974:SLO

[Chr74h] D. Christiansen. Spectral lines:

Our circles of concern. IEEE Spectrum, 11(6):37, June 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1974:SLSb

[Chr74i] D. Christiansen. Spectral lines:

Some of us don’t quit. IEEE Spectrum, 11(12):43, December 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1974:SLV

[Chr74j]  D. Christiansen.  Spectral

lines: Very safe systems pose dilemma. IEEE Spectrum, 11 (11):43, November 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1974:SLWa

[Chr74k] D. Christiansen. Spectral lines:

Who gets published? IEEE Spectrum, 11(7):33, July 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1974:SLWb

[Chr74l] D. Christiansen. Spectral lines:

Who gets published? Part II: Transactions. IEEE Spectrum, 11(8):31, August 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1974:SLWc

[Chr74m] D. Christiansen. Spectral lines:

Who gets published? Part III:

Proceedings. IEEE Spectrum, 11 (9):43, September 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1974:SRT

[Chr74n] D. Christiansen. Spectral report: Technology ’74. IEEE Spectrum, 11(1):30–31, January 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1975:SSR

[Chr75a] D. Christiansen. A Spectrum special report. IEEE Spectrum, 12 (10):38–39, October 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1975:SLDb

[Chr75b] D. Christiansen. Spectral lines: a dangerous art. IEEE Spectrum, 12(4):39–40, April 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1975:SLC

[Chr75c] D. Christiansen. Spectral lines: Conflict of interest. IEEE Spectrum, 12(6):33, June 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1975:SLDa

[Chr75d] D. Christiansen. Spectral lines: Disconnecting from the system. IEEE Spectrum, 12(3):33, March 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1975:SLE

[Chr75e] D. Christiansen. Spectral lines:

Education, registration, certification, validation. IEEE Spectrum, 12(8):33, August 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1975:SLH

[Chr75f] D. Christiansen. Spectral lines:

Hardware, software, and mushware. IEEE Spectrum, 12 (10):37, October 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1975:SLSc

[Chr75g] D. Christiansen. Spectral lines:

In the swim, or on the beach? IEEE Spectrum, 12(12):27, December 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1975:SLJ

[Chr75h] D. Christiansen. Spectral lines:

Judging the judges. IEEE Spectrum, 12(11):29, November 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1975:SLL

[Chr75i] D. Christiansen. Spectral lines:

… by any other name? IEEE Spectrum, 12(7):25, July 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1975:SLSb

[Chr75j] D. Christiansen. Spectral lines: Society-proof design. IEEE Spectrum, 12(5):31, May 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1975:SLSa

[Chr75k] D. Christiansen. Spectral lines: ‘soft’ technology and the ‘backup gap’. IEEE Spectrum, 12(2): 29, February 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1975:SLP

[Chr75l]  D. Christiansen.  Spectral

lines: The public and nuclear power. IEEE Spectrum, 12 (9):35, September 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1975:SRT

[Chr75m] D. Christiansen. Spectral report: Technology ’75. IEEE Spectrum, 12(1):36–37, January 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1975:TTL

[Chr75n] D. Christiansen. Time and technology: Looking ahead in electronics. IEEE Spectrum, 12(4):41, April 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1976:ART

[Chr76a] D. Christiansen. Applications review: Technology ’76. IEEE Spectrum, 13(1):42–43, January 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1976:SLI

[Chr76b] D. Christiansen. Spectral lines:

an idea whose time may never come. IEEE Spectrum, 13(11): 29–30, November 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1976:SLS

[Chr76c] D. Christiansen. Spectral lines:

As others see us. IEEE Spectrum, 13(6):35–36, June 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1976:SLAc

[Chr76d] D. Christiansen. Spectral lines:

Assessing the assessors. IEEE Spectrum, 13(10):29, October 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1976:SLAb

[Chr76e]  D. Christiansen.  Spectral

lines: Assessment or arrestment? IEEE Spectrum, 13 (9):27, September 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1976:SLE

[Chr76f] D. Christiansen. Spectral lines:

Engineering and emotion. IEEE Spectrum, 13(4):37, April 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1976:SLF

[Chr76g] D. Christiansen. Spectral lines: Fact vs. feeling. IEEE Spectrum, 13(3):35, March 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1976:SLAa

[Chr76h] D. Christiansen. Spectral lines: It’s about time. IEEE Spectrum, 13(8):23, August 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1976:SLR

[Chr76i] D. Christiansen. Spectral lines: Rasmussen revisited. IEEE Spectrum, 13(1):41, January 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1976:SLCb

[Chr76j] D. Christiansen. Spectral lines: The cities’ dark side. IEEE Spectrum, 13(7):27, July 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1976:SLCa

[Chr76k] D. Christiansen. Spectral lines: The closed-loop society. IEEE Spectrum, 13(2):33, February 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1976:SLW

[Chr76l]  D. Christiansen.  Spectral

lines: Where do we go from here? IEEE Spectrum, 13(12):

25–26, December 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1977:SL

[Chr77a] D. Christiansen. Spectral lines. IEEE Spectrum, 14(2): 31–32, February 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1977:SLAa

[Chr77b] D. Christiansen. Spectral lines:

Announcing the amazing micromouse maze contest. IEEE Spectrum, 14(5):27, May 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1977:SLBb

[Chr77c] D. Christiansen. Spectral lines:

Blackout prevention. IEEE Spectrum, 14(10):21–22, October 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1977:SLJ

[Chr77d] D. Christiansen. Spectral lines:

Japan today: Assets, liabilities, and the future. IEEE Spectrum, 14(9):33, September 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1977:SLUa

[Chr77e] D. Christiansen. Spectral lines:

Over (or under) 40 and out (or in). IEEE Spectrum, 14(4):25–26, April 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1977:SLR

[Chr77f] D. Christiansen. Spectral lines: Report from the U.K. IEEE Spectrum, 14(3):39–40, March 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1977:SLAb

[Chr77g] D. Christiansen. Spectral lines: The age of the survey. IEEE Spectrum, 14(12):19–20, December 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1977:SLBa

[Chr77h] D. Christiansen. Spectral lines: The bankers vs. technology. IEEE Spectrum, 14(1):35, January 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1977:SLIa

[Chr77i] D. Christiansen. Spectral lines: The IEEE as an advocate. IEEE Spectrum, 14(8):23, August 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1977:SLIb

[Chr77j] D. Christiansen. Spectral lines: The insidious and cultural machine. IEEE Spectrum, 14 (11):25, November 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1977:SLUb

[Chr77k] D. Christiansen. Spectral lines:

Universal solvent or universal problem. IEEE Spectrum, 14 (6):27–28, June 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1977:T

[Chr77l] D. Christiansen. Technology ’77. IEEE Spectrum, 14(1): 36–37, January 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1978:SL

[Chr78a] D. Christiansen. Spectral lines.

IEEE Spectrum, 15(1):23, January 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1978:SLEb

[Chr78b] D. Christiansen. Spectral lines:

Edisonian vignettes. IEEE Spectrum, 15(9):29, September 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1978:SLEa

[Chr78c] D. Christiansen. Spectral lines:

Electronics and power: a remarriage. IEEE Spectrum, 15(7):25, July 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1978:SLPa

[Chr78d] D. Christiansen. Spectral lines:

In praise of objectivity. IEEE Spectrum, 15(4):21, April 1978.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1978:SLM

[Chr78e] D. Christiansen. Spectral lines: More power to you and more computer power, too. IEEE Spectrum, 15(2):25, February 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1978:SLS

[Chr78f] D. Christiansen. Spectral lines: On systems thinking. IEEE Spectrum, 15(8):19, August 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1978:SLPc

[Chr78g] D. Christiansen. Spectral lines: Productivity: Its ups and downs. IEEE Spectrum, 15(10):31, October 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1978:SLR

[Chr78h] D. Christiansen. Spectral lines: Registration: a second look. IEEE Spectrum, 15(3):29, March 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1978:SLY

[Chr78i] D. Christiansen. Spectral lines: So you’re an engineer. But what do you do? IEEE Spectrum, 15(6):23, June 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1978:SLN

[Chr78j] D. Christiansen. Spectral lines:

The new world information order: High stakes. IEEE Spectrum, 15(11):29, November 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1978:SLPb

[Chr78k] D. Christiansen. Spectral lines:

The perils of publishing. IEEE Spectrum, 15(5):27, May 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1978:SLPd

[Chr78l] D. Christiansen. Spectral lines:

The postal service hopes to stamp out its fires. IEEE Spectrum, 15(12):21, December 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1978:T

[Chr78m] D. Christiansen. Technology ’78. IEEE Spectrum, 15(1): 24–25, January 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1979:SLD

[Chr79a] D. Christiansen. Spectral lines:

Data load: Wheat vs. chaff. IEEE Spectrum, 16(2):23, February 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1979:SLL

[Chr79b] D. Christiansen. Spectral lines: Looking back to look ahead. IEEE Spectrum, 16(1):29, January 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1979:SLM

[Chr79c] D. Christiansen. Spectral lines: Motivating engineers. IEEE Spectrum, 16(3):27, March 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1979:SLNb

[Chr79d] D. Christiansen. Spectral lines: No way to run a railroad. IEEE Spectrum, 16(12):27, December 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1979:SLNa

[Chr79e] D. Christiansen. Spectral lines: Nuclear power: The Swedish experience. IEEE Spectrum, 16(8):17, August 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1979:SLO

[Chr79f] D. Christiansen. Spectral lines: Opinions on nuclear power. IEEE Spectrum, 16(11):29, November 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1979:SLRa

[Chr79g] D. Christiansen. Spectral lines: Reactor safety reconsidered. IEEE Spectrum, 16(4):29, April 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1979:SLRc

[Chr79h] D. Christiansen. Spectral lines:

Risk avoidance. IEEE Spectrum, 16(6):17, June 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1979:SLRd

[Chr79i] D. Christiansen. Spectral lines:

Risk comparison. IEEE Spectrum, 16(9):25, September 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1979:SLRb

[Chr79j] D. Christiansen. Spectral lines:

Risk taking. IEEE Spectrum, 16(5):25, May 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1979:SLSa

[Chr79k] D. Christiansen. Spectral lines:

Spectrum: a progress report. IEEE Spectrum, 16(7):21, July 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1979:SLSb

[Chr79l] D. Christiansen. Spectral lines:

The science court tested. IEEE Spectrum, 16(10):27, October 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1979:T

[Chr79m] D. Christiansen. Technology ’79. IEEE Spectrum, 16(1): 30–31, January 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1979:TP

[Chr79n] D. Christiansen. TMI and the press. IEEE Spectrum, 16(11): 92–96, November 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Christiansen:1979:WDW

[Chr79o] D. Christiansen. Where do we go from here?: Analysis by the experts reveals steps to be undertaken in nuclear plant design, operating procedures, and institutional matters. IEEE Spectrum, 16(11):110–111, November 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Clapper:1971:AWR

[Cla71] G. L. Clapper. Automatic word recognition. IEEE Spectrum, 8 (8):57–69, August 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Claybrook:1977:SVN

[Cla77] J. Claybrook. Safety: The view from NHTSA: New legislation promises to make cars safer, cheaper to repair, and less polluting during the 1980s. IEEE Spectrum, 14(11):87–88, November 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Clement:1973:CPP

[Cle73] R. C. Clement. Conservation:

a positive position. IEEE Spectrum, 10(8):44–47, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Clorfeine:1973:DUI

[Clo73] A. S. Clorfeine. Driving under the influence of electronics. IEEE Spectrum, 10(5):32–37, May 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Collins:1978:ERE

[Col78a] T. Collins. Education: a radical experiment in teaching. IEEE Spectrum, 15(8):48–52, August 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Collins:1978:EEC

[Col78b] T. Collins. Education: Engineering curricula: ‘professional’ vs. ‘two-track’: In the debate over possible changes in engineering education, one of the key questions is what is meant by ‘professional school’. IEEE Spectrum, 15(5):52–55, May 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Collins:1978:MRE

[Col78c] T. Collins. A minority report: Experts look at some of the assumptions used in traditional measures of productivity and probe productivity’s impact. IEEE Spectrum, 15(10): 40–45, October 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Collins:1979:PES

[Col79] T. Collins. Power/energy: Solar energy: Four sites demonstrate potential: an office building in Florida, a warehouse in New Hampshire, a gymnasium in Iowa, and a laundry in California exploit sun power. IEEE Spectrum, 16(4):60–65, April 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Conley:1970:ECP

[Con70] P. Conley. Experience curves as a planning tool. IEEE Spectrum, 7(6):63–68, June 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cory:1972:EPN

[Cor72] W. E. Cory. EEs and public needs: A recommendation. IEEE Spectrum, 9(2):33–35, February 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Costigan:1974:FHL

[Cos74]  D. M. Costigan.  ‘Fax’ in

the home: looking back and ahead. IEEE Spectrum, 11(9): 76–82, September 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cotton:1975:AIA

[Cot75] W. B. Cotton. Avionics ii: Air safety: The view from the cockpit: an experienced airline pilot appraises the past, present, and future of air safety and air traffic control trends. IEEE Spectrum, 12(8):71–74, August 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Count:1979:PEE

[Cou79] B. M. Count. Power/energy: Exploiting wave power: To realize the promise, engineers must deal with unique power-transfer modes, erratic seas, and unfavorable economics. IEEE Spectrum, 16(9):42–49, September 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Caeazza:1976:PEC

[CSS76] J. A. Caeazza, T. R. Schneider, and V. T. Sulzberger. Power: Energy on call: a more efficient peaking system would exploit the advantages of energy storage, while conserving capital and resources. IEEE Spectrum, 13(6):45–48, June 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Cho:1970:LPD

[CT70] S. Cho and A. G. Thiele. The low-power-drain microelectronic VHF amplifier. IEEE Spectrum, 7(3):49–53, March 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dalziel:1970:TGF

[Dal70] C. F. Dalziel. Transistorized ground-fault interrupter reduces shock hazard. IEEE Spectrum, 7(1):55–62, January 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dalziel:1972:ESH

[Dal72] C. F. Dalziel. Electric shock hazard. IEEE Spectrum, 9(2): 41–50, February 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Darmstadter:1970:IFE

[Dar70] J. Darmstadter. International flows of energy sources. IEEE Spectrum, 7(5):67–72, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Davida:1979:IHS

[Dav79] G. I. Davida. III. ‘Hellman’s scheme breaks DES in its basic form’. IEEE Spectrum, 16(7):39, July 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Davis:1971:LTC

[DBB71] D. D. Davis, B. E. Blair, and J. F. Barnaba. Long-term continental U.S. timing system via television networks. IEEE Spectrum, 8(8):41–52, August 1971.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

DeFalco:1970:RCL

[DeF70] J. A. DeFalco. Reflection and crosstalk in logic circuit interconnections. IEEE Spectrum, 7(7):44–50, July 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Derganc:1979:ITE

[Der79] C. S. Derganc. Innovation:

Thomas Edison and his electric lighting system: The inventor’s notebooks reveal that he was actually a skilled practitioner of systems engineering, not the tinkerer of legend. IEEE Spectrum, 16(2):50–59, February 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dible:1972:DBW

[Dib72] D. M. Dible. Dear boss: Why I’m quitting …. IEEE Spectrum, 9(5):62–67, May 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dillard:1975:IAR

[Dil75] J. K. Dillard. IEEE annual reports: Report of the Treasurer: a detailed picture of where IEEE funds came from in 1974, and where they went. IEEE Spectrum, 12(6):67–70, June 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dillman:1972:BE

[DKR^{+}72] T. L. Dillman, F. W. Keay, C. Raczkowski, J. W. Skooglund, and W. H. South. Brushless excitation. IEEE Spectrum, 9 (3):58–66, March 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Doi:1977:RGS

[Doi77] Y. Doi. Robots get smarter and more versatile: Robots and computer-based systems proliferate, having gained inroads due to labor shortages. IEEE Spectrum, 14(9):65–68, September 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dorrell:1972:SDC

[Dor72] R. F. Dorrell. Switch in dielectric cuts costs. IEEE Spectrum, 9(6):53–55, June 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Doucette:1976:MAD

[Dou76]  A. R. Doucette.  Military/

avionics: Design decisions for a head-up display: This F-14 fighter optics retrofit case history is a prime example of analyzing and selecting design options. IEEE Spectrum, 13 (8):29–32, August 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dreher:1974:RHC

[Dre74] C. Dreher. Retrospective: His colleagues remember ‘the doctor’: Alfred N. Goldsmith, cofounder of IRE and Director and Editor Emeritus of IEEE, is cited for his wisdom and statesmanship. IEEE Spectrum, 11 (8):32–37, August 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dalton:1976:CDE

[DTW76] G. Dalton, P. Thompson, and I. Wilson. Career development: an EE for all seasons: High performers step progressively through a series of four distinct career stages while others stagnate along the way. IEEE Spectrum, 13(12):42–46, December 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Dunkerley:1978:ARO

[Dun78] W. I. Dunkerley. Amateur radio:

OSCAR (and friends): Useful yet frugal: Since 1961, amateur radio operators have been launching satellites for communications, educational, and humanitarian applications. IEEE Spectrum, 15 (12):30–33, December 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eden:1979:DVU

[EBC^{+}79] G. Eden, R. Burnham, L. F.

Champagne, T. Donohue, and N. Djeu. Devices: Visible and UV lasers: Problems and promises: New families of gas lasers are expected to deliver a range of pulse intensities, durations, and wavelengths for diverse applications. IEEE Spectrum, 16(4):50–59, April 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eddy:1970:EC

[Edd70] N. Eddy. The exponential crisis. IEEE Spectrum, 7(8):53–56, August 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ellert:1976:PHL

[EH76] F. J. Ellert and N. G. Hingorani. Power: HVDC for the long run: High-voltage de transmission may be a viable alternative to ac transmission for longdistance routes. IEEE Spectrum, 13(8):37–44, August 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ewart:1971:AUS

[EK71] D. N. Ewart and L. K. Kirchmayer. Automation and utility system security. IEEE Spectrum, 8(7):37–42, July 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eleccion:1972:DDC

[Ele72a] M. Eleccion. A/D and D/A converters. IEEE Spectrum, 9(7):63– 66, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eleccion:1972:BBB

[Ele72b] M. Eleccion. Beating the bluebox bandits. IEEE Spectrum, 9(8):52–58, August 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eleccion:1972:FLS

[Ele72c] M. Eleccion. The family of lasers:

a survey. IEEE Spectrum, 9 (3):26–40, March 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eleccion:1972:MPL

[Ele72d] M. Eleccion. Materials processing with lasers. IEEE Spectrum, 9(4):62–72, April 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eleccion:1973:AFI

[Ele73a] M. Eleccion. Automatic fingerprint identification. IEEE Spectrum, 10(9):36–45, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eleccion:1973:EW

[Ele73b] M. Eleccion. The electronic watch. IEEE Spectrum, 10 (4):24–32, April 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eleccion:1973:ELE

[Ele73c] M. Eleccion. Electronics in law enforcement. IEEE Spectrum, 10 (2):33–40, February 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eleccion:1973:LH

[Ele73d] M. Eleccion. Laser hazards.

IEEE Spectrum, 10(8):32–38, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eleccion:1973:NRT

[Ele73e] M. Eleccion. A new role for technology. IEEE Spectrum, 10(6):50–51, June 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eleccion:1973:SSL

[Ele73f] M. Eleccion. Science, secrecy, and the law. IEEE Spectrum, 10 (11):65–70, November 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eleccion:1974:ATQ

[Ele74a] M. Eleccion. Automatic testing: quality raiser, dollar saver. IEEE Spectrum, 11(8):38–43, August 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eleccion:1974:IPS

[Ele74b] M. Eleccion. Innovative programs: Solving big, real problems: The classic engineering curriculum is under attack. New curricula are interdisciplinary, socially aware, and tuned to the real world. IEEE Spectrum, 11(5):52–53, May 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eleccion:1974:IVH

[Ele74c] M. Eleccion. Instrumentation VII: The human factors: The operator, who must push buttons, twist knobs, and observe results rapidly, clearly, and unambiguously, is a principal design constraint. IEEE Spectrum, 11 (11):73–74, November 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eleccion:1974:ITE

[Ele74d] M. Eleccion. Instruments and test equipment: Computers have snuggled into the measurements domain to tackle new tasks. IEEE Spectrum, 11(1): 82–84, January 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eleccion:1974:PSO

[Ele74e] M. Eleccion. Perspective: a sunny outlook for solar power: At NSF, a fifty-fold increase in funding over four years may presage a bright future for its solar energy program. IEEE Spectrum, 11(4):76–77, April 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eleccion:1974:SSS

[Ele74f] M. Eleccion. Sociotechnology:

Secrecy in science: a dilemma revisited: While government experts agree on the need for free scientific expression, some deliberate on related issues. IEEE Spectrum, 11(4):56–57, April 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eleccion:1975:APA

[Ele75a] M. Eleccion. Avionics: The promise of air safety: Although air traffic has been increasing for years, installation of collision avoidance systems is only now being mandated. IEEE Spectrum, 12(7):26–37, July 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eleccion:1975:MAS

[Ele75b]  M. Eleccion.  Military and

aerospace: Space Shuttle holds top priority at NASA; inflationcrippled military seeks $-efficient defense. IEEE Spectrum, 12 (1):83–89, January 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eleccion:1976:IAT

[Ele76a] M. Eleccion. Instrumentation: Automatic test equipment; hardware and software: The need for standard languages and sophisticated interfaces is paced by the expanding applications for automatic testing. IEEE Spectrum, 13(6):60–64, June 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eleccion:1976:MAW

[Ele76b]  M. Eleccion.  Military and

aerospace: Work progresses on Space Shuttles, air cushion landing systems, Earth resources programs, and controls. IEEE Spectrum, 13(1):80–83, January 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Eleccion:1978:MPE

[Ele78] M. Eleccion. Meeting preview:

ELECTRO ’78: Beans, brown bread, and Boston: 36 technical sessions organized by Program Chairman Brian Dale’s committee promise toothsome technical fare, too. IEEE Spectrum, 15(4):52–55, April 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Elleccion:1972:EMP

[Ell72] M. Elleccion. ECC ’72: a meeting preview. IEEE Spectrum, 9(5):43–44, May 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Engstrom:1975:RCH

[Eng75] L. Engström. Refining copper with HVDC: a 1700-km-long transmission line thrusts southeastward through Zaire to power a vital copper source in the interior. IEEE Spectrum, 12(12): 40–46, December 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ernst:1977:ISP

[Ern77a] M. L. Ernst. Ii social problems and interfaces. IEEE Spectrum, 14(5):54–55, May 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ernst:1977:IPE

[Ern77b] M. L. Ernst. III. Progress and expectations. IEEE Spectrum, 14(5):56–57, May 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Esposito:1977:UPE

[Esp77] V. J. Esposito. The U.S. promotes electric vehicles: an Energy Department program is designed to encourage research and support experiments. IEEE Spectrum, 14(11):69–70, November 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Everitt:1977:EPS

[Eve77] W. L. Everitt. The engineer: a perennial student: Engineers are reminded that the process of learning shouldn’t end as the classroom door closes behind them. IEEE Spectrum, 14(12): 61–63, December 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Edgerton:1978:LNR

[EW78] H. E. Edgerton and C. W. Wyckoff. Loch Ness revisited: Fact or fantasy? Science uses sonar and camera to probe the depths of Loch Ness in search of its resident monster. IEEE Spectrum, 15(2):26–29, February 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ewart:1978:PWW

[Ewa78] D. N. Ewart. Power: Whys and wherefores of power system blackouts: an examination of the factors that increase the likelihood and the frequency of system failure. IEEE Spectrum, 15(4):36–41, April 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Faggin:1978:MHV

[Fag78]  F. Faggin.  Microprocessors:

How VLSI impacts computer architecture: a solution is suggested to the imminent problem of having more computer intelligence on a chip than anyone can use. IEEE Spectrum, 15(5):28–31, May 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fagenbaum:1979:OSR

[Fag79] J. Fagenbaum. Optical systems: a review: Fiber-optics applications include power monitoring, video transmission, data links, and public communications. IEEE Spectrum, 16 (10):70, October 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Faik:1974:CRH

[Fai74] H. Faik. Computer report i:

Hard-soft tradeoffs: Large-scale integrated gates and read-only memories are displacing software from many of its traditional roles. IEEE Spectrum, 11 (2):34–39, February 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Falk:1973:ESS

[Fal73a] H. Falk. Engineering the Space Shuttle. IEEE Spectrum, 10 (5):50–54, May 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Falk:1973:NEC

[Fal73b] H. Falk. Nerem 73: the electron and the city. IEEE Spectrum, 10 (10):66–67, October 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Falk:1973:SE

[Fal73c] H. Falk. Scanning the Earth. IEEE Spectrum, 10(12):84–89, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Falk:1973:TBO

[Fal73d] H. Falk. Toward better overocean ATC. IEEE Spectrum, 10 (12):64–65, December 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Falk:1974:ALM

[Fal74a] H. Falk. Applications: Linking microprocessors to the real world: a proper interface serves as communications traffic cop, setting priorities and directing the flow of messages. IEEE Spectrum, 11(9):59–67, September 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Falk:1974:CMG

[Fal74b] H. Falk. Communications and microwaves: a good spectrum stretcher is hard to find. Reaching toward light frequencies seems a good alternative. IEEE Spectrum, 11(1):32–36, January 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Falk:1974:CRI

[Fal74c] H. Falk. Computer report IV:

a checkups: Minicomputer software: Great expectations for new applications hinge on the usefulness of software packages. IEEE Spectrum, 11(2):52–56, February 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Falk:1974:CMS

[Fal74d] H. Falk. Computers: Microcomputer software makes its debut. IEEE Spectrum, 11(10): 78–84, October 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Falk:1974:CSC

[Fal74e] H. Falk. Computers: Selfcontained microcomputers ease system implementation: The user gets a microprocessor, memory, built-in software, a power supply, and input-output circuits so he can program his own hardware. IEEE Spectrum, 11(12): 53–55, December 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Falk:1974:CAW

[Fal74f] H. Falk. Consumer: Aluminum wire: The heat is on: Is it possible your house could burn down because the contractor saved some money by using aluminum wire? IEEE Spectrum, 11(12): 44–45, December 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Falk:1974:CEW

[Fal74g] H. Falk. Continuing education: Whys and ways for keeping up: The options in midcareer education can bewilder; defer them and the penalties can be severe. IEEE Spectrum, 11(5):36–45, May 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Falk:1974:ESF

[Fal74h] H. Falk. European satellites to ‘fly’. IEEE Spectrum, 11 (8):55–59, August 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Falk:1974:IIM

[Fal74i] H. Falk. Instrumentation II: The microprocessors jack-of-all-trades: It’s used for calibration, control, computation, conversion, decision-making, test, and maintenance. IEEE Spectrum, 11 (11):46–51, November 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Falk:1974:ITH

[Fal74j] H. Falk. International trends: how EEs learn around the world: curricula and training methods usually don’t just happen — nation-by-nation peculiarities help shape them. IEEE Spectrum, 11(5):78–81, May 1974.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Falk:1975:CTT

[Fal75a] H. Falk. Communications I: Telephone technology: The ‘which switch’ game. IEEE Spectrum, 12(6):41–42, June 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Falk:1975:CSH

[Fal75b] H. Falk. Computer systems:

Hardware/software: LSI is filling the gaps between microcomputers, minis, and large machines. IEEE Spectrum, 12 (1):38–43, January 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Falk:1975:TFC

[Fal75c] H. Falk. Technology forecasting I: Communications: Fiber optics and digital microwave may be fast-moving areas in the otherwise ponderous progress of future communications. IEEE Spectrum, 12(4):42–45, April 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Falk:1975:TFI

[Fal75d] H. Falk. Technology forecasting II: Computers: The scenario suggests omnipresent computers shaping future patterns of business and daily life. IEEE Spectrum, 12(4):46–49, April 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Falk:1975:TID

[Fal75e] H. Falk. Tools II/design techniques: Design for production: The computer steps in to automate physical design and testing of computers and other electronic equipment. IEEE Spectrum, 12 (10):48–53, October 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Falk:1975:TIE

[Fal75f] H. Falk. Tools IV/education: Microprocessor short courses: a sampling of courses, ranging from one day to two weeks, and priced from $25 to $750, is provided. IEEE Spectrum, 12 (10):58–61, October 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Falk:1976:CPP

[Fal76a] H. Falk. Computers: Poised for progress: Microprocessors, memories, and fault-tolerant systems pace current developments. IEEE Spectrum, 13 (1):44–49, January 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Falk:1976:IAE

[Fal76b] H. Falk. Industry applications: Electronics prints this magazine: Spectrum exploits computers, photographic processing, and electronic equipment to produce each month’s copy. IEEE Spectrum, 13(2):71–75, February 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Falk:1976:WWW

[Fal76c] H. Falk. What went wrong V:

Reaching for a gigaflop: The fate of the famed Illiac IV was shaped by both research brilliance and real-world disasters. IEEE Spectrum, 13(10): 65–70, October 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Falk:1977:CDT

[Fal77a] H. Falk. ‘Chipping in’ to digital telephones. IEEE Spectrum, 14(2):42, February 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Falk:1977:CAP

[Fal77b] H. Falk. Computers: All pervasive: Hardware and software progress is steady; applications continue to expand. IEEE Spectrum, 14(1):38–42, January 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Falk:1977:MMT

[Fal77c] H. Falk. Microprocessors: Microprocessors talk it up: Micros, alone and in constellations, are star feature of new radio, telephone, data, and satellite systems. IEEE Spectrum, 14 (3):49–53, March 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fanshel:1975:F

[Fan75] S. Fanshel. Forum. IEEE Spectrum, 12(1):30–34, January 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Farris:1973:EI

[Far73] H. W. Farris. Engineers or ‘ingenors’? IEEE Spectrum, 10 (3):74–79, March 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frey:1972:WAM

[FB72] J. Frey and R. Bowers. What’s ahead for microwaves. IEEE Spectrum, 9(3):41–47, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fink:1978:OUS

[FC78] L. H. Fink and K. Carlsen. Operating under stress and strain [electrical power systems control under emergency conditions]. IEEE Spectrum, 15(3):48–53, March 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Flanagan:1970:SVC

[FCR^{+}70] J. L. Flanagan, C. H. Coker, L. R. Rabiner, R. W. Schafer, and N. Umeda. Synthetic voices for computers. IEEE Spectrum, 7(10):22–45, October 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Felix:1973:EOU

[Fel73] F. Felix. Energy options for the United States. IEEE Spectrum, 10(9):63–68, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Feth:1973:MBF

[Fet73] G. C. Feth. Memories are bigger, faster and cheaper. IEEE Spectrum, 10(11):28–35, November 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Feth:1976:CMS

[Fet76] G. C. Feth. Computers: Memories: Smaller, faster, and cheaper: a review of magneticbubble, electron-beam-accessed, disk- and tape-file memories; MOS RAMs and CCDs excel. IEEE Spectrum, 13(6):37–44, June 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fey:1974:PNW

[Fey74] C. F. Fey. Putting numbers where your hunches used to be. IEEE Spectrum, 11(7):34– 40, July 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Firschein:1974:IMW

[FFCT74] O. Firschein, M. A. Fischler, L. S. Coles, and J. M. Tenenbaum. Intelligent machines are on the way. IEEE Spectrum, 11(7):41–49, July 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fink:1970:PPS

[Fin70] D. G. Fink. Pension plans: The state of the art. IEEE Spectrum, 7(11):34–39, November 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fink:1972:BC

[Fin72a]  D. G. Fink.  Blueprint for

change. IEEE Spectrum, 9 (6):38–43, June 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fink:1972:EPP

[Fin72b] D. G. Fink. For electronics, the past is prologue. IEEE Spectrum, 9(1):63–66, January 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Finney:1973:SSD

[Fin73] R. G. Finney. Successful system design: a case history. IEEE Spectrum, 10(9):58–62, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fine:1977:ERL

[Fin77] M. S. Fine. Engineering registration and the law: a review of engineering registration in 1977 how the system operates, what state laws require, and a look into the future. IEEE Spectrum, 14 (10):60–64D, October 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Firschein:1975:LRS

[Fir75] O. Firschein. On-line reference searching. IEEE Spectrum, 12 (10):68–71, October 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Falk:1976:IIM

[FK76] H. Falk and Gadi Kaplan. Instrumentation III: The microprocessor takeover: On-the-spot mathematics, automatic testing, and portable calibration are yielding to microprocessor solutions. IEEE Spectrum, 13 (4):44–49, April 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fleckenstein:1971:ECU

[Fle71] W. O. Fleckenstein. Economic conditions in the U.S. electrical, electronics, and related industries: an assessment. IEEE Spectrum, 8(12):63–71, December 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Flock:1971:EEE

[Flo71] W. L. Flock. The environment and the electrical engineering curriculum. IEEE Spectrum, 8(7):62–67, July 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Forney:1970:CAS

[For70] G. D. Forney. Coding and its application in space communications. IEEE Spectrum, 7(6):47–58, June 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Forrester:1971:SAT

[For71] J. W. Forrester. Systems analysis as a tool for urban planning. IEEE Spectrum, 8(1):48–58, January 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1975:EWA

[FR75] G. D. Friedlander and E. Rubinstein. Energy ways around the cash crunch: Options for retracking the U.S. utilities onto their traditional path of adequate and reliable supply. IEEE Spectrum, 12(4):62–65, April 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frank:1978:PPS

[Fra78] R. J. Frank. Patents: The patentability of software inventions: Computer-implemented inventions that are useful, new, and nonobvious may still not be patentable — except, say courts, in certain cases. IEEE Spectrum, 15(4):42–47, April 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Freedman:1974:ATO

[Fre74] M. D. Freedman. Advanced technology: Optical character recognition: Machines that read printed matter rely on innovative designs. Systems and their performance are compared. IEEE Spectrum, 11(3):44–52, March 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1970:CATa

[Fri70a] G. D. Friedlander. At the crossroads in air-traffic control I. The view from the ground — the congested airports; the Common IFR Room; present and future automated systems; plans of the FAA; problems of the air-traffic controllers. IEEE Spectrum, 7 (6):26–40, June 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1970:CATb

[Fri70b] G. D. Friedlander. At the crossroads in air-traffic control II. The view from the groundinstrument-landing systems, automated ground control, en route automation, introduction to collision-avoidance systems, positive-control airspace, improved communications, STOLports, and terminal navigational aids. IEEE Spectrum, 7(7):69–83, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1970:CATc

[Fri70c] G. D. Friedlander. At the crossroads in air-traffic control III. View from the cockpit — reactions of some pilots, automatic flight management, flightcontrol systems, airborne CAS, ‘moving map’ displays, ‘threedimensional’ area navigation, and others. IEEE Spectrum, 7 (8):76–89, August 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1970:BLW

[Fri70d] G. D. Friedlander. The Burndy Library: window on the history of science. IEEE Spectrum, 7 (3):54–64, March 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1970:PPIa

[Fri70e] G. D. Friedlander. Power, pollution, and the imperiled environment. IEEE Spectrum, 7(11): 40–50, November 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1970:PPIb

[Fri70f] G. D. Friedlander. Power, pollution, and the imperiled environment II. East, Midwest, and West Coast: pollution-control plans of some major utilities; role of government in environmental matters; other proposed systems for reducing stack emissions. IEEE Spectrum, 7(12): 65–75, December 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1971:EHE

[Fri71a] G. D. Friedlander. Electricity in hospitals: elimination of lethal hazards. IEEE Spectrum, 8(9): 40–51, September 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1971:PLC

[Fri71b] G. D. Friedlander. Power from Labrador: the Churchill Falls development. IEEE Spectrum, 8 (2):81–91, February 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friis:1971:IRR

[Fri71c] H. T. Friis. Introduction to radio and radio antennas. IEEE Spectrum, 8(4):55–61, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1972:BC

[Fri72a] G. D. Friedlander. The BART chronicle. IEEE Spectrum, 9 (9):34, September 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1972:BHB

[Fri72b] G. D. Friedlander. BART’s hardware — from bolts to computers. IEEE Spectrum, 9(10): 60–62, October 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1972:CRK

[Fri72c] G. D. Friedlander. A comeback for reddy kilowatt? IEEE Spectrum, 9(4):44–50, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1972:DCU

[Fri72d] G. D. Friedlander. Driving cable underground. IEEE Spectrum, 9 (11):60–61, November 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1972:FFY

[Fri72e] G. D. Friedlander. Ford, fuels, and your future. IEEE Spectrum, 9(8):59–60, August 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1972:PPG

[Fri72f] G. D. Friedlander. Is ‘power to the people’ going underground? IEEE Spectrum, 9 (2):62–71, February 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1972:MMC

[Fri72g] G. D. Friedlander. Managing megawatts: a case history. IEEE Spectrum, 9(5):39–42, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1972:MBH

[Fri72h] G. D. Friedlander. More BART hardware. IEEE Spectrum, 9 (11):41–45, November 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1972:PPS

[Fri72i] G. D. Friedlander. Power plant siting: a détente? IEEE Spectrum, 9(7):38–39, July 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1972:RRR

[Fri72j] G. D. Friedlander. Railroad revival: On the right track. IEEE Spectrum, 9(8):63–66, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1972:SWT

[Fri72k] G. D. Friedlander. Supercommunications for world traders. IEEE Spectrum, 9(3):74–83, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1972:TEG

[Fri72l] G. D. Friedlander. Tesla: eccentric genius. IEEE Spectrum, 9(6):26–29, June 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1973:BBB

[Fri73a] G. D. Friedlander. Bigger bugs in BART? IEEE Spectrum, 10 (3):32–37, March 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1973:EPN

[Fri73b] G. D. Friedlander. Earthquake prediction: a new art. IEEE Spectrum, 10(9):46–57, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1973:ECR

[Fri73c] G. D. Friedlander. Energy conservation by redesign. IEEE Spectrum, 10(11):36–43, November 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1973:ECC

[Fri73d] G. D. Friedlander. Energy: Crisis and challenge. IEEE Spectrum, 10(5):18–27, May 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1973:FRC

[Fri73e] G. D. Friedlander. Floating reactor: ‘crisis’ solution? IEEE Spectrum, 10(2):44–51, February 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1973:OOE

[Fri73f] G. D. Friedlander. Oil: the omnipotent energy source. IEEE Spectrum, 10(7):26–29, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1973:PB

[Fri73g] G. D. Friedlander. A prescription for BART. IEEE Spectrum, 10(4):40–44, April 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1973:TNE

[Fri73h] G. D. Friedlander. Toward a national energy policy. IEEE Spectrum, 10(6):36–43, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1974:TLS

[Fri74a] G. Friedlander. Tom lee speaks out. IEEE Spectrum, 11(6): 46–47, June 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1974:EPF

[Fri74b] G. D. Friedlander. Energy/ power: The fast-breeder reactor: When, where, why, and how? An LMFBR plant will be built and the HTGR is under study; the factors involved are efficiency and safety. IEEE Spectrum, 11 (2):85–89, February 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1974:PEP

[Fri74c]  G. D. Friedlander.  Power/

energy: Problems and progress: Global politics has fueled an overdue search by the industrial nations for new energy sources. IEEE Spectrum, 11 (1):60–65, January 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1974:PPE

[Fri74d] G. D. Friedlander. Professional practices/ethics: The case of the three engineers vs. BART: The outcome of a suit brought by three Bay Area Rapid Transit ex-employees may set a historic precedent in the public interest. IEEE Spectrum, 11 (10):69–77, October 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1974:POE

[Fri74e]  G. D. Friedlander.  Profile:

the “other” electric company: American electric power is a study in contrasts — anachronistic, yet progressive, and a giant in spite of its relatively low profile. IEEE Spectrum, 11(6):48–55, June 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1974:PRT

[Fri74f] G. D. Friedlander. Progress in rail transportation: Next: London to Paris in just over 3 hours, and a fast ride from Oakland to San Francisco for 60 cents. IEEE Spectrum, 11 (1):66–70, January 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1974:RTE

[Fri74g] G. D. Friedlander. Rail transportation: Electronics and Swiss railways: Switzerland’s rail system designers’ bag of tricks ranges from wayside signals to sophisticated telecommunication and automation. IEEE Spectrum, 11(9):68–75, September 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1974:RTG

[Fri74h] G. D. Friedlander. Rail transportation: Germany automates its rails: The key to West German railway efficiency resides in electronic signals and continuous automatic train control. IEEE Spectrum, 11(7):73– 77, July 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1974:RTR

[Fri74i] G. D. Friedlander. Rail transportation: Riding Swiss rails: an electrifying saga: Switzerland has some of the finest all-electric locomotives and unusual rightsof-way in Europe. IEEE Spectrum, 11(8):44–54, August 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1974:RTW

[Fri74j] G. D. Friedlander. Rail transportation: West Germany: The phoenix rises: The German rail system rates among Europe’s finest, perhaps partly because it sprang from the rubble of 1945. IEEE Spectrum, 11(4):62–70, April 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1974:TRS

[Fri74k] G. D. Friedlander. Transportation: Riding Sweden’s slick rail system: an on-the-scene report reveals how this 11400 kilometer system is kept smoothly running. IEEE Spectrum, 11 (3):53–64, March 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1975:BRR

[Fri75a] G. D. Friedlander. Britannia rides the rails. IEEE Spectrum, 12(7):50–59, July 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1975:PEY

[Fri75b]  G. D. Friedlander.  Power/

energy: a year of struggle: Seeking efficiency gains, the industry pushes both new and clinical use. IEEE Spectrum, 12 (1):70–73, January 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1975:RTF

[Fri75c] G. D. Friedlander. Rail transportation the French (train) connection: The railroads of France are seen to be the best in Europe: Its “seconds” are exported to AMTRAK. IEEE Spectrum, 12(3):45–52, March 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1975:RAF

[Fri75d] G. D. Friedlander. Retrospective: Ampère: Father of electrodynamics: The master mathematician, teacher, and practitioner of electrical science is celebrated on the 200th anniversary of his birth. IEEE Spectrum, 12(8):75–78, August 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1975:SRG

[Fri75e] G. D. Friedlander. Special report: Gov. Shapp on what’s wrong with rail transportation: an EE turned successful politician suspects a conspiracy of rail interests in the U.S. against business and the public. IEEE Spectrum, 12(7):60–64, July 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1975:TFE

[Fri75f] G. D. Friedlander. Technology forecasting: Energy’s hazy future: The power forecasting game, played so avidly by so many in 1970, is now a cautious exercise with fewer, more expert participants. IEEE Spectrum, 12(5):32–40, May 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1975:TFB

[Fri75g] G. D. Friedlander. Transportation: Fixing BART: a “milliondollar” backup for the automatic train protection system stems from BART’s “Blue Ribbon Panel” report. IEEE Spectrum, 12(2):43–45, February 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1975:TPB

[Fri75h] G. D. Friedlander. Transportation progress: a banner year in the art of people moving saw California’s BART go transbay. IEEE Spectrum, 12(1): 80–83, January 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1976:CSB

[Fri76a] G. D. Friedlander. Circa 2000: a Spectrum bicentennial report on cities of the future. IEEE Spectrum, 13(7):28–30, July 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1976:CIS

[Fri76b] G. D. Friedlander. Cities ii: In support of megastructures: a scheme to bring people and their jobs back together again by constructing a multipurpose metropolis-under-one-roof. IEEE Spectrum, 13(7):37–42, July 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1976:GB

[Fri76c] G. D. Friedlander. The great blackout of ’65. IEEE Spectrum, 13(10):83–88, October 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1976:NPP

[Fri76d] G. D. Friedlander. Nuclear power plant safety. II. IEEE Spectrum, 13(5):70–75, May 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1976:PPF

[Fri76e] G. D. Friedlander. Perspective: Peccei on the future: Club of Rome President Aurelio Peccei, in an exclusive interview, gives his views and prognoses for humankind’s future. IEEE Spectrum, 13(11):54–57, November 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1976:PMU

[Fri76f]  G. D. Friedlander.  Power:

matching utility output to customer demand: load management includes the design of energy-supply facilities to meet customer load, plus the control of a utility’s “load shape”. IEEE Spectrum, 13(9):51–53, September 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1976:PPP

[Fri76g]  G. D. Friedlander.  Power:

Problems and progress: Despite difficulties, advances were notable in UHV and ac transmission, and energy-storage devices. IEEE Spectrum, 13(1): 64–67, January 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1976:PFP

[Fri76h] G. D. Friedlander. Power: The FEA on productivity: How the U.S. Federal Energy Administration recommends improving power plant efficiency and productivity. IEEE Spectrum, 13 (3):48–54, March 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1976:TTH

[Fri76i] G. D. Friedlander. Transportation: Transpo head speaks out:

The new Secretary of Transportation addresses major issues affecting U.S. surface and air mobility. IEEE Spectrum, 13 (2):76–79, February 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1976:TYT

[Fri76j] G. D. Friedlander. Transportation: Year of travail: Although some projects were halted or delayed, new urban and mainline systems are proceeding. IEEE Spectrum, 13(1): 73–75, January 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1977:PPB

[Fri77a] G. D. Friedlander. Power: Planning for a brighter future: Modeling and simulation, load forecasting, frequency control, stability, security, and reliability are important factors. IEEE Spectrum, 14(3):61–68, March 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1977:PEH

[Fri77b]  G. D. Friedlander.  Power/

energy: high marks for technology. IEEE Spectrum, 14 (1):62–65, January 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1977:PEU

[Fri77c]  G. D. Friedlander.  Power/

energy: UHV: Onward and upward: a major effort is being mounted in the U.S. and overseas to push practical transmission above the 1000-kV level. IEEE Spectrum, 14(2):57–65, February 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Friedlander:1977:TSG

[Fri77d] G. D. Friedlander. Transportation: some good news. IEEE Spectrum, 14(1):70–73, January 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Frosch:1979:NTS

[Fro79] R. A. Frosch. NASA takes stock. IEEE Spectrum, 16 (6):44–49, June 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Fee:1977:BR

[FS77] W. F. Fee and S. Sherr. Book reviews. IEEE Spectrum, 14(7):60– 61, July 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gallagher:1977:CMG

[Gal77] E. F. Gallagher. Communications: The military goes digital: Digital switching and signaling, along with digital transmission, offer reliability in a hostile environment. IEEE Spectrum, 14 (2):42–48, February 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gallawa:1979:UCI

[Gal79] R. Gallawa. I. U.S. and Canada: Initial results reported. IEEE Spectrum, 16(10):71–74, October 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gangloff:1977:WIN

[Gan77] W. C. Gangloff. What ifs? for nuclear plants: Sophisticated reliability analyses and risk assessments are aimed at predicting failures and their consequences. IEEE Spectrum, 14 (6):53–59, June 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Garrett:1970:ICDa

[Gar70a] L. S. Garrett. Integrated-circuit digital I. Requirements and features of a logic family; RTL, DTL, and HTL devices logic families. IEEE Spectrum, 7 (10):46–58, October 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Garrett:1970:ICDb

[Gar70b] L. S. Garrett. Integrated-circuit digital logic families II. TTL devices. IEEE Spectrum, 7(11): 63–72, November 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Garrett:1970:ICDc

[Gar70c] L. S. Garrett. Integrated-circuit digital logic families III. ECL and MOS devices. IEEE Spectrum, 7 (12):30–42, December 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gary:1977:IGF

[Gar77] P. Gary. I2L: It’s getting faster and smaller. IEEE Spectrum, 14(6):30–33, June 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Golding:1973:STC

[GB73] L. S. Golding and J. E. D.

Ball. Satellite television covers the world. IEEE Spectrum, 10 (8):24–31, August 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geddes:1972:IDB

[Ged72] L. A. Geddes. Interface design for bioelectrode systems. IEEE Spectrum, 9(10):41–48, October 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geffe:1970:THS

[Gef70] P. R. Geffe. Toward high stability in active filters. IEEE Spectrum, 7(5):63–66, May 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gerry:1970:GL

[Ger70]  E. T. Gerry.  Gasdynamic

lasers. IEEE Spectrum, 7(11): 51–58, November 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gerakos:1974:HFC

[Ger74] J. Gerakos. Human factors: Case history: Building a medical monitor. IEEE Spectrum, 11(11): 75–77, November 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gould:1970:TDE

[GH70] R. G. Gould and G. K. Helder. Transmission delay and echo suppression. IEEE Spectrum, 7 (4):47–54, April 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Geddes:1971:DBC

[GH71] L. A. Geddes and H. E. Hoff. The discovery of bioelectricity and current electricity: The Galvani–Volta controversy. IEEE Spectrum, 8(12):38–46, December 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gamow:1972:WEC

[GH72] R. I. Gamow and J. F. Harris. What engineers can learn from nature. IEEE Spectrum, 9(8):36–42, August 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gauper:1971:PSA

[GHM71] H. A. Gauper, J. D. Harnden, and A. M. McQuarrie. Power supply aspects of semiconductor equipment. IEEE Spectrum, 8 (10):32–43, October 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Glaser:1979:PEF

[GHNK79] P. E. Glaser, G. M. Hanley, R. H. Nansen, and R. L. Kline. Power/energy: First steps to the Solar Power Satellite: Guidelines issued by the Department of Energy and NASA provide a reference point for an ambitious program. IEEE Spectrum, 16(5):52–58, May 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gibson:1970:PCG

[Gib70] J. E. Gibson. Planning the coordination of ground transport. IEEE Spectrum, 7(10): 79–86, October 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gibson:1976:CVP

[Gib76] J. E. Gibson. Cities v: The people ‘doughnut’: a ring off urban centers and suburbs is linked in the time domain by high-speed ground transport. IEEE Spectrum, 13(7):51–56, July 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ginzton:1975:RIH

[Gin75] E. L. Ginzton. Retrospective:

The $100 idea: How Russell and Sigurd Varian, with the help of William Hansen and a $100 appropriation, invented the klystron. IEEE Spectrum, 12 (2):30–39, February 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldberg:1977:BR

[GJH77] D. L. Goldberg, W. G. Jung, and F. Hill. Book reviews. IEEE Spectrum, 14(9):85–87, September 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gibson:1976:SEN

[GK76] R. Gibson and W. J. Kleen. Satellites — Europe’s ‘NASA’ gets off the ground. IEEE Spectrum, 13(2):66–70, February 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Glebov:1977:EEG

[Gle77] I. A. Glebov. Electric energy: a global summary: The chairman of its program commission reports on major trends discussed at the recent World Electrotechnical Congress. IEEE Spectrum, 14(12):52–54, December 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gurtler:1972:LCD

[GM72] R. W. Gurtler and C. Maze. Liquid crystal displays. IEEE Spectrum, 9(11):25–29, November 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldmark:1970:CE

[Gol70] P. C. Goldmark. Color EVR. IEEE Spectrum, 7(9):22–33, September 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gold:1972:HOD

[Gol72] M. Gold. How one designer views his job. IEEE Spectrum, 9(7):33– 37, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldberg:1973:NFI

[Gol73] H. S. Goldberg. A new focus for IEEE’s professional goals. IEEE Spectrum, 10(6):62–64, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldwater:1974:FS

[Gol74] S. B. Goldwater. Freedom and secrecy. IEEE Spectrum, 11 (4):57–58, April 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldwag:1979:SMH

[Gol79] H. Goldwag. Sociotechnology:

Microwave hazards: Are U.S. standards for safe exposure levels adequate? One IEEE member warns of smug complacency; another responds. IEEE Spectrum, 16(5):66–69, May 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gorski:1975:TIA

[Gor75] J. C. Gorski. Toys II/avocation:

Inside amateur radio: Devoted to public service transcending geographical and political boundaries, today’s ham is not a mere tinkerer. IEEE Spectrum, 12 (10):76–78, October 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gossen:1979:KRP

[Gos79a] R. N. Gossen. I. The 64-kbit RAM: a prelude to VLSI. IEEE Spectrum, 16(3):42–45, March 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gossen:1979:SSG

[Gos79b] R. N. Gossen. Solid state: 100,000+ gates on a chip: Mastering the minutia: The submicron geometries needed for VLSI have put IC makers on a new learning curve; object, to perfect feature delineation techniques. IEEE Spectrum, 16(3):42, March 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gazzana-Priaroggia:1971:LIS

[GPPM71] P. Gazzana-Priaroggia, J. H. Piscioneri, and S. W. Margolin. The Long Island Sound submarine cable interconnection. IEEE Spectrum, 8(10):63–71, October 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Goldmark:1976:WWW

[GR76] P. C. Goldmark and E. Rubinstein. What went wrong XI: Then came EVR. IEEE Spectrum, 13(10):95–96, October 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Granger:1970:EPC

[Gra70a] J. V. N. Granger. The employment problem: Can IEEE help? IEEE Spectrum, 7(11): 32–33, November 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Granger:1970:IR

[Gra70b] J. V. N. Granger. IEEE reports for 1969. IEEE Spectrum, 7(6):74–86, June 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Granger:1970:PAC

[Gra70c] J. V. N. Granger. The President’s address at the 1970 Convention Banquet. IEEE Spectrum, 7(5):22–23, May 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Grayson:1974:MTT

[Gra74] L. P. Grayson. Manpower trends:

Tomorrow’s EE: The engineer of the future may well be an M.S., a Ms., black, or a combination of all three. IEEE Spectrum, 11(5):54–59, May 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gray:1979:MLL

[Gra79] I. Gray. Management: Learning to live with red tape: It can tie up projects unduly, yet the organization man (or woman) can use it wisely. IEEE Spectrum, 16(7):60–63, July 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Grebene:1971:MPL

[Gre71] A. B. Grebene. The monolithic phase-locked loop — a versatile building block. IEEE Spectrum, 8(3):38–49, March 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Grebene:1972:MWG

[Gre72] A. B. Grebene. Monolithic waveform generation. IEEE Spectrum, 9(4):34–40, April 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gregory:1974:APW

[Gre74] G. Gregory. Analysis/perspective: Why Japan succeeds: a unique capital investment technique, a dedicated labor force, and a knack for backing the right “technological horse” make Japan “go”. IEEE Spectrum, 11 (3):65–73, March 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Greening:1976:AAW

[Gre76a] W. C. Greening. Analyst to the analog world [spectrum analyser]. IEEE Spectrum, 13 (4):68–72, April 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gregory:1976:BOM

[Gre76b] G. Gregory. Business outlook: making it in the third world. IEEE Spectrum, 13 (3):74–77, March 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Greenberg:1979:PTN

[Gre79a] D. S. Greenberg. The politics of technology. National Academy of Engineering: Is it merely an honorary society or does it have clout on Capitol Hill? A report on what it is and what it could be. IEEE Spectrum, 16 (12):52–55, December 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Greenberg:1979:RDP

[Gre79b] D. S. Greenberg. Research and development: The President’s R&D budget: Basic is up, applied is not: President Carter asks for more real dollars for basic research, but holds the line in other categories. IEEE Spectrum, 16(3):62–64, March 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Greenhalgh:1979:NPE

[Gre79c] G. Greenhalgh. Nuclear power in Eastern Europe. IEEE Spectrum, 16(11):105–108, November 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gregory:1979:ITJ

[Gre79d] G. Gregory. International technology: Japan turns its ingenuity to the world computer market: Breakthroughs in component design and system performance, along with finely honed marketing skills, underlie the Japanese challenge. IEEE Spectrum, 16 (4):69–71, April 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gross:1972:IP

[Gro72] W. A. Gross. The innovation process. IEEE Spectrum, 9(3):53–57, March 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Gupta:1974:IRW

[Gup74] M. Gupta. Information retrieval, where to find it when you need it. IEEE Spectrum, 11(5):72–77, May 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Haggerty:1971:PS

[Hag71] P. E. Haggerty. The productive society. IEEE Spectrum, 8 (9):30–39, September 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hall:1977:ATM

[Hal77] H. P. Hall. Analog tests: the microprocessor scores. IEEE Spectrum, 14(4):37–42, April 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Handler:1979:SST

[Han79] P. Handler. Sociotechnology: Science, technology, and the human condition: The impact of science and engineering on citizens is great, despite their questionable impact on social institutions. IEEE Spectrum, 16 (8):49–52, August 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Happ:1973:PPS

[Hap73] H. H. Happ. Power pools and superpools. IEEE Spectrum, 10 (3):54–61, March 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Harsch:1972:DRC

[Har72] A. F. Harsch. ‘Dry runs’ by computer. IEEE Spectrum, 9 (9):43–46, September 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Harnden:1977:PSL

[Har77] J. D. Harnden. Power semiconductors: Looking ahead: Consumer applications will grow for moderate-current devices while large-diameter wafers await packaging innovations. IEEE Spectrum, 14(8):40–46, August 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Haviland:1970:WSB

[Hav70] R. P. Haviland. Why space broadcasting? IEEE Spectrum, 7(2):86–91, February 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Heinen:1977:BCA

[HB77] C. M. Heinen and E. W. Beckman. Balancing clean air against good mileage: Contradictory specifications for lower emissions and improved fuel economy require compromise solutions. IEEE Spectrum, 14(11):

47–50, November 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Healy:1974:TEC

[Hea74] T. J. Healy. Transportation: The electric cars will it really go? It may not save energy but it could be cheaper, less polluting, and in the long haul, petroleum conserving. IEEE Spectrum, 11 (4):50–55, April 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Heffner:1970:SSP

[Hef70] H. Heffner. Science in the seventies: the policy issues. IEEE Spectrum, 7(7):31–34, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Heilmeier:1979:GCM

[Hei79] G. H. Heilmeier. 100,000+ gates on a chip: mastering the minutia. II. Needed: a ‘miracle slice’ for VLSI fabrication. IEEE Spectrum, 16(3):45–48, March 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hellman:1979:WTI

[Hel79] M. E. Hellman. I. ‘DES will be totally insecure within ten years’. IEEE Spectrum, 16(7):32– 40, July 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hersch:1971:DC

[Her71] P. Hersch. Data communications. IEEE Spectrum, 8(2): 47–60, February 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hershberg:1972:EPP

[Her72] P. I. Hershberg. The EE’s place in public health. IEEE Spectrum, 9(9):63–66, September 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Herold:1973:KED

[Her73] E. W. Herold. Keeping up:

the engineer’s dilemma. IEEE Spectrum, 10(9):77–78, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Herold:1979:FHI

[Her79] E. W. Herold. Finances: How inflation affects your retirement dollar: At 13 percent annual inflation, an annual living expense of $24 000 in 1979 dollars will soar to $323 000 in 20 years. IEEE Spectrum, 16(9): 58–61, September 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hoffman:1974:MIK

[HFL74] A. A. J. Hoffman, R. L. French, and G. M. Lang. Minicomputer interfaces: know more, save more. IEEE Spectrum, 11 (2):64–68, February 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hodges:1979:CTE

[HGB79] D. A. Hodges, P. R. Gray, and R. W. Broderson. Circuit techniques: Enhancing MOS/LSI’s role in analog design: Highdensity amplifiers, A/D and D/A converters, filters, and coderdecoder circuits can now be fabricated. IEEE Spectrum, 16 (2):24–32, February 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hochman:1970:TAL

[HH70] H. T. Hochman and D. L. Hogan. Technological advances in largescale integration. IEEE Spectrum, 7(5):50–58, May 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hollomon:1971:TDU

[HH71] J. H. Hollomon and A. E. Harger. The technological dilemma of the United States. IEEE Spectrum, 8(10):55–62, October 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Heer:1974:ERT

[HH74] J. E. Heer and D. J. Hagerty. Energy: Refuse turns resource: Diverted from land fills to hammer mills, municipal waste becomes an economic energy resource. IEEE Spectrum, 11(9): 83–87, September 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hill:1974:ISL

[Hil74] J. C. Hill. Instrumentation: The state of logic analyzers: Simultaneous multichannel recording, pretrigger memory, and combinatorial triggering are now available for difficult digital measurements. IEEE Spectrum, 11(12): 63–70, December 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hinchman:1979:BTC

[Hin79] W. Hinchman. On Bell: I. Time for a change: a former FCC common carrier bureau chief believes there’s a better way to structure the communications industry. IEEE Spectrum, 16(12): 42–48, December 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Horelick:1976:CMS

[HL76] D. Horelick and R. S. Larsen.

CAMAC: a modular standard. IEEE Spectrum, 13(4):50–57, April 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Habermann:1979:CPM

[HL79] H. Habermann and G. L. Liard.

Control: Practical magnetic bearings: Electronically controlled electromagnets support spinning 1100-kg shafts to micrometer accuracy. IEEE Spectrum, 16(9):26–30, September 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Holland:1970:PEF

[Hol70] R. Holland. Piezoelectric effects in ferroelectric ceramics. IEEE Spectrum, 7(4):67–77, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hollomon:1972:TPU

[Hol72] J. H. Hollomon. Technology policy: the U.S. picks its way. IEEE Spectrum, 9(8):72–81, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Horn:1972:GNC

[Hor72] M. Horn. Good news for commuters: better and faster rapid transit. IEEE Spectrum, 9(2): 55–61, February 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Horan:1976:STL

[Hor76] J. J. Horan. Space technology: Landsat: Multispectral eye in the sky: Data collected by the Landsat satellites’ remote sensors are exploited in mapping, meteorology, and resource identification. IEEE Spectrum, 13 (3):59–63, March 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Horan:1978:ANV

[Hor78] J. J. Horan. Aerospace: Nimbus: The vanguard of remote sensing: Imaging and spectral analysis devices aboard Nimbus-7 measure the Earth’s radiation, wind speeds, and pollution, from space. IEEE Spectrum, 15(11): 36–43, November 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Howell:1970:ANH

[How70] R. P. Howell. Aerospace nomads: How professionals move. IEEE Spectrum, 7(1):32–41, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hsiao:1970:IMC

[HS70] H. S. Hsiao and C. Susskind. Infrared and microwave communication by moths. IEEE Spectrum, 7(3):69–76, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Heller:1977:F

[HS77] H. H. Heller and J. Snyder.

Forum. IEEE Spectrum, 14 (10):16–17, October 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Huang:1970:DDP

[Hua70] K. T. Huang. Desensitization of dc power supplies to momentary ac power fluctuations. IEEE Spectrum, 7(6):69–73, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Huse:1974:RDB

[Hus74] R. A. Huse. Research and development: Breaking the crisis: Options cover primary energy resource recovery, fuel conditioning, conversion, improved transmission, and storage. IEEE Spectrum, 11(6):40–47, June 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Hampel:1971:TL

[HW71] D. Hampel and R. O. Winder. Threshold logic. IEEE Spectrum, 8(5):32–39, May 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Iijima:1977:PCW

[IFSK77] S. Iijima, K. Fujii, H. Sugaya, and M. Kano. Pleasing the consumer while playing to win: Consumer electronics is Japan’s trump card, fine-tuned to woo customers, dispel adversity. IEEE Spectrum, 14(9): 45–50, September 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Inagaki:1973:CGB

[IKT73] J. Inagaki, M. Kuniyoshi, and S. Tadakuma. Commutators get the brushoff. IEEE Spectrum, 10(6):52–58, June 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ishihara:1977:RGL

[Ish77] T. Ishihara. Railroads: Green light for the bullets: Advances in vehicles, communications, and signals keep plans for the Shinkansen line on schedule. IEEE Spectrum, 14(9): 71–74, September 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ishii:1977:CJS

[IY77] Y. Ishii and H. Yamada. Computers in Japan: Super coordination: Together, designers, equipment manufacturers, and the government spur progress. IEEE Spectrum, 14(9):40–44, September 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jansen:1970:ELF

[Jan70] J. Jansen. EHV lines in the Federal Republic of Germany. IEEE Spectrum, 7(4):33–40, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jayadev:1976:PWS

[Jay76] T. S. Jayadev. Power: Windmills stage a comeback: Born again amid the search for fossil-fuel alternatives, wind-generated electricity now makes technical and economic sense. IEEE Spectrum, 13(11):45–49, November 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Johnson:1977:BR

[JE77] L. M. Johnson and P. Elias. Book reviews. IEEE Spectrum, 14(12): 68–77, December 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jedynak:1973:WSA

[Jed73] L. Jedynak. Where the (switch)

action is. IEEE Spectrum, 10 (10):56–62, October 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jenny:1978:EHR

[Jen78] H. K. Jenny. Education: Heavy readers are heavy hitters: a survey of 3000 RCA engineers not only reveals how they keep abreast, but correlates their information-retrieval habits with achievement. IEEE Spectrum, 15 (9):66–69, September 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jespersen:1972:TTT

[JF72] J. L. Jespersen and L. Fey. ‘Time-telling’ techniques. IEEE Spectrum, 9(5):51–58, May 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1974:PEO

[JF74] R. K. Jurgen and G. D. Friedlander. Power: The energy outlook: Ways to go the issues are complex; the solutions not obvious. But both conservation and inspiration will play key roles. IEEE Spectrum, 11 (3):83–87, March 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jacobs:1977:COT

[JM77] I. Jacobs and S. E. Miller. Communications: Optical transmission of voice and data: Bandwidth and noise immunity are up; size, weight, and costs are down. IEEE Spectrum, 14(2):

33–41, February 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Joel:1976:EM

[Joe76] A. E. Joel. ESS: minimonster. IEEE Spectrum, 13(8): 33–36, August 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Johnson:1970:BMD

[Joh70a] C. M. Johnson. Ballistic-missile defense radars. IEEE Spectrum, 7(3):32–41, March 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Johnson:1970:AIC

[Joh70b]  H. Johnson.  The anatomy

of integrated-circuit technology. IEEE Spectrum, 7(2): 56–66, February 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Johnson:1971:ENF

[Joh71] J. B. Johnson. Electronic noise:

the first two decades. IEEE Spectrum, 8(2):42–46, February 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Johnson:1979:PES

[Joh79] R. W. Johnson. Power/energy:

Solar power satellite: Putting it together: Enormous structures involving unique fabrication methods call for new materials, tools, and heavy-lift vehicles. IEEE Spectrum, 16(9):

37–41, September 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1977:ASD

[Jon77a] T. Jones. The automobile: It’s the super-driver of both technology and society. IEEE Spectrum, 14(11):26–30, November 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jones:1977:AES

[Jon77b] T. O. Jones. Automotive electronics i: Smaller and better: Electronics may be the only costeffective solution to the performance challenge of today’s autos. IEEE Spectrum, 14(11): 34–36, November 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jordan:1970:EEF

[Jor70] J. F. Jordan. From engineer to entrepreneur: financing the transition. IEEE Spectrum, 7 (2):40–41, February 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1970:MAS

[Jur70] R. K. Jurgen. Minicomputer applications in the seventies. IEEE Spectrum, 7(8):37–52, August 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1971:TWA

[Jur71] R. K. Jurgen. Two-way applications for cable television systems in the ’70s. IEEE Spectrum, 8 (11):39–54, November 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1972:CEN

[Jur72a] R. K. Jurgen. Catv enters a new era. IEEE Spectrum, 9(8):69–71, August 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1972:FCL

[Jur72b] R. K. Jurgen. Feedforward correction: a late-blooming design. IEEE Spectrum, 9(4):41–43, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1972:IDI

[Jur72c]  R. K. Jurgen.  Information

displays: innovation is the rule. IEEE Spectrum, 9(11): 18, November 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1972:UQC

[Jur72d] R. K. Jurgen. Untangling the ‘quad’ confusion. IEEE Spectrum, 9(7):55–62, July 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1972:WF

[Jur72e] R. K. Jurgen. Working for fun. IEEE Spectrum, 9(4):59–61, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1973:CWU

[Jur73a]  R. K. Jurgen.  Computer-

watching in the U.S.S.R. IEEE Spectrum, 10(10):37–39, October 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1973:HCD

[Jur73b] R. K. Jurgen. Health care delivery: a job for EEs? IEEE Spectrum, 10(4):34–39, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1973:MDL

[Jur73c] R. K. Jurgen. Medical device laws: a critical balancing act. IEEE Spectrum, 10(7):30– 36, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1974:PNU

[Jur74a] R. Jurgen. Perspective: Notes from the U.S.S.R.: Popov Congress delegates report on Soviet developments in biological effects of laser radiation, chargecoupled devices, and color TV. IEEE Spectrum, 11(12):71, December 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1974:CEC

[Jur74b]  R. K. Jurgen.  Consumer

electronics: From calculators and cameras galore to video recorders, the shopping list grows. IEEE Spectrum, 11 (1):78–79, January 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1974:EEE

[Jur74c] R. K. Jurgen. Education/energy:

Expanding power options: The power industry’s long-term challenges offer exciting opportunities to students. But do they know it? IEEE Spectrum, 11(5):68–71, May 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1974:ECE

[Jur74d] R. K. Jurgen. Energy conservation: Exploding some myths: When it comes to cutting energy expenditures, the seemingly obvious may prove fallacious under scrutiny. IEEE Spectrum, 11(6):74–75, June 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1974:ECW

[Jur74e] R. K. Jurgen. Energy conservation: what to tell your neighbors: the layman expects you to have the real facts. These tips may surprise him — and even you! IEEE Spectrum, 11(6):61–65, June 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1974:ICD

[Jur74f] R. K. Jurgen. Instrumentation:

Competing display technologies struggle for superiority: a comparison of display media, including light emitters and controllers and CRTs, reveals which are best for what. IEEE Spectrum, 11 (10):90–95, October 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1974:IVW

[Jur74g] R. K. Jurgen. Instrumentation VI: What users like and dislike; need and can’t get: Basic satisfaction with test equipment is tempered with a need for more self-checking, reliability, and human factors engineering. IEEE Spectrum, 11(11):69–72, November 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1974:MEU

[Jur74h] R. K. Jurgen. Medical electronics: Ultrasonics in medicine: Strong; getting stronger: Twoand three-dimensional imaging breakthroughs may spark the next generation of ultrasonic imaging systems. IEEE Spectrum, 11(8):62–66, August 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1975:AEF

[Jur75a] R. K. Jurgen. Applications: Electronic fuel injection: Mileage boost, pollution squelch: Electronic controls are designed to activate injectors, air valves, and fuel pump to improve engine performance. IEEE Spectrum, 12 (11):34–36, November 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1975:AEI

[Jur75b] R. K. Jurgen. Automotive electronics: Ignition systems go solid state: Standard on all 1975 U.S. cars, electronic ignitions reduce maintenance costs, improve cold starts, and keep engines tuned longer. IEEE Spectrum, 12(9): 49–51, September 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1975:AEM

[Jur75c] R. K. Jurgen. Automotive electronics: The microprocessor: In the driver’s seat? In the research laboratory, the microprocessor looks promising; in the real world, sensor/actuator problems may stall progress. IEEE Spectrum, 12(6):73–77, June 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1975:CEM

[Jur75d] R. K. Jurgen. Captain Eddy:

The man who launched a thousand EEs: His genius in getting things done meant success in training electronics technicians for the U.S. Navy in WW II. IEEE Spectrum, 12 (12):53–56, December 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1975:TFI

[Jur75e] R. K. Jurgen. Technology forecasting IV: Instrumentation: In the offing are low-cost sensors; voltage-tuned, wide-range signal sources; signal-sequence generators; CCD displays; and much more. IEEE Spectrum, 12 (4):52–55, April 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1976:BDD

[Jur76a] R. K. Jurgen. Biomedical: Devices for the disabled: a production dilemma: Manufacturers avoid producing technologically advanced devices for real and imagined reasons. IEEE Spectrum, 13(9):45–50, September 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1976:EAB

[Jur76b] R. K. Jurgen. Electronics aids the ‘biomedics’. IEEE Spectrum, 13(1):76–79, January 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1976:N

[Jur76c] R. K. Jurgen. It’s 75 for NBS. IEEE Spectrum, 13 (4):78–82, April 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1976:SWT

[Jur76d] R. K. Jurgen. Sensors: Where there’s smoke there’d better be a smoke detector. Inexpensive, life-saving systems vie for consumer interest. IEEE Spectrum, 13(8):24–28, August 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1976:SHM

[Jur76e] R. K. Jurgen. Software (and hardware) for the ‘medics’. IEEE Spectrum, 13(4):40–43, April 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1976:WWW

[Jur76f] R. K. Jurgen. What went wrong VI: Learning from (heart) failure: The implantable cardiac pacemaker is now beginning to reach the useful lifetime it was supposed to have had 15 years ago. IEEE Spectrum, 13 (10):71–74, October 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1977:ABW

[Jur77a] R. K. Jurgen. The automobile: For better or worse: Amid controversy about size, fuel consumption, and safety hazards, there is unanimity concerning the application of electronics to autos. IEEE Spectrum, 14(11): 31–33, November 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1977:EMM

[Jur77b] R. K. Jurgen. Electronics in medicine: a mixed blessing: Technology does lots more, but at a price. IEEE Spectrum, 14 (1):78–81, January 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1977:ST

[Jur77c] R. K. Jurgen. I the state of the technology. IEEE Spectrum, 14(5):51–53, May 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1978:AET

[Jur78a] R. K. Jurgen. Automotive electronics: Tuneup for Detroit: New auto radios are rolling along in GM, Ford, and Chrysler products as electronics replaces mechanical tuners. IEEE Spectrum, 15(5):32–35, May 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1978:EMD

[Jur78b] R. K. Jurgen. Electronics in medicine: Diagnostic imaging advances; microprocessors aid laser systems, prostheses. IEEE Spectrum, 15(1):68–72, January 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1978:MDS

[Jur78c]  R. K. Jurgen.  Microcom-

puters: For Detroit: Smaller and smarter chips: Semiconductor manufacturers are designing more complex microcomputer systems, while they, and the auto makers, worry about reliability. IEEE Spectrum, 15 (11):33–35, November 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Jurgen:1979:EMM

[Jur79] R. K. Jurgen. Electronics in medicine: The microprocessor takes over and the ultimate CAT scanner nears completion, but costs still soar. IEEE Spectrum, 16(1):76–80, January 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kalra:1975:TDB

[Kal75] P. Kalra. Transportation: Dynamic braking. IEEE Spectrum, 12(5):63–66, May 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kalra:1979:TGL

[Kal79] P. S. Kalra. Transportation: a green light for advanced train controls: Solid-state developments yield hands-off operation of second-generation systems. IEEE Spectrum, 16(2): 44–49, February 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1974:FSP

[Kap74a] Gadi Kaplan. Fax: Specs and projections. IEEE Spectrum, 11 (9):77–82, September 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1974:IVP

[Kap74b] Gadi Kaplan. Instrumentation VIII: Power plant controls: Displays, computers, and man: From channel redundancy to direct digital control, new concepts in design are aiding the man at the power switch. IEEE Spectrum, 11(11):78–82, November 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1975:BNF

[Kap75a] Gadi Kaplan. Bugs in the nuclear fuel cycle. IEEE Spectrum, 12 (9):56–64, September 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1975:SPS

[Kap75b] Gadi Kaplan. For solar power: Sunny days ahead? The future of sun-derived electric power is clouded by budgetary constraints, institutional and public inertia, and poor incentives. IEEE Spectrum, 12(12): 47–52, December 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1975:IEC

[Kap75c] Gadi Kaplan. Industrial electronics: Controllers get smarter, thyristors make inroads, hardware/software goes standard. IEEE Spectrum, 12 (1):74–79, January 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1975:SSS

[Kap75d] Gadi Kaplan. Seeking super safety [nuclear power plants]. IEEE Spectrum, 12(6):46–56, June 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1975:TDT

[Kap75e] Gadi Kaplan. Tools I/design techniques: Computer-aided design: No longer a mere infant, CAD now offers more sophisticated programs with Monte Carlo computations and tolerance assignment. IEEE Spectrum, 12(10):40–47, October 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1976:CVC

[Kap76a] Gadi Kaplan. Cities VIII:

Cities: Energy gluttons: a review of the energy demands and constraints that will be impressed upon our cities in future decades. IEEE Spectrum, 13(7):73–74, July 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1976:EET

[Kap76b] Gadi Kaplan. Energy: Europe:

Tilting toward fusion: Today the French Phenix breeder scores well; tomorrow, fusion seems the byword. IEEE Spectrum, 13(12): 37–41, December 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1976:EPS

[Kap76c] Gadi Kaplan. Energy: Planning solar’s future: Solar power struggles for economic viability; its proponents see it as an important, if small, complement to traditional sources. IEEE Spectrum, 13(6):55–59, June 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1976:IEB

[Kap76d] Gadi Kaplan. Industrial electronics: to boost productivity. IEEE Spectrum, 13(1): 87–90, January 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1976:WWW

[Kap76e] Gadi Kaplan. What went wrong IV: The Browns Ferry incident: Seeking air leaks with a candle at the TVA’s nuclear power plant, a worker found much more. IEEE Spectrum, 13(10):55–64, October 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1977:APT

[Kap77a] Gadi Kaplan. Autos and public transit: Friends or foes? A balanced transportation system would exploit the advantages of both private motor vehicles and public transportation systems. IEEE Spectrum, 14 (11):89–92, November 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1977:CEG

[Kap77b] Gadi Kaplan. Constraining the energy gobbler: Industry is the principal energy consumer; technology exists to save fuel and, of course, dollars. IEEE Spectrum, 14(12):27–33, December 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1977:ETC

[Kap77c] Gadi Kaplan. Electronics takes control in manufacturing: Computer numerical control excels, lasers invade, and energy is saved. IEEE Spectrum, 14(1): 58–61, January 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1977:PGC

[Kap77d] Gadi Kaplan. Power generation:

Constraints, yet progress: Technology helps power engineers circumvent problems: Land availability and population density. IEEE Spectrum, 14(9): 61–64, September 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1977:PEN

[Kap77e] Gadi Kaplan. Power/energy:

Nontraditional sources: a sampler: Engineers in Israel, France, and the U.K. choose their own ways to go in the energy crisis. IEEE Spectrum, 14(2): 49–56, February 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1977:TWC

[Kap77f] Gadi Kaplan. Two-way communication for load management: The results of system tests now underway will guide utilities in performance, cost, and legal aspects. IEEE Spectrum, 14 (8):47–51, August 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1978:PEP

[Kap78a] Gadi Kaplan. Power/energy progress report: a pastiche of developments promises near and long-term applications. IEEE Spectrum, 15(1):49–53, January 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1978:TOU

[Kap78b] Gadi Kaplan. Techniques that optimize the use of the world’s dwindling energy supply inevitably affect productivity as well. IEEE Spectrum, 15(10): 70–73, October 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1978:TEG

[Kap78c] Gadi Kaplan. Transportation: Electronics galore: Microprocessors broaden impact onland, sea, and air transport. IEEE Spectrum, 15(1):59–62, January 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1979:EWW

[Kap79a] Gadi Kaplan. Electronics on wings, wheels, and waves: More speed and more safety through microprocessors, sensors, and advanced communications. IEEE Spectrum, 16(1):67–70, January 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1979:NPJ

[Kap79b] Gadi Kaplan. Nuclear power in Japan. IEEE Spectrum, 16(11): 108–109, November 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1979:PEP

[Kap79c] Gadi Kaplan. Power/energy:

Progress in efficiency and reliability: Advances are defined in time-of-day rates, generation and transmission research, and protective systems. IEEE Spectrum, 16(1):58–61, January 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1979:TSD

[Kap79d] Gadi Kaplan. Three systems defined: Advanced satellite and microwave radio techniques will facilitate gigabit per second capacities. IEEE Spectrum, 16 (10):42–46, October 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Katz:1975:MPP

[Kat75] L. Katz. Management planning is for people: Painful, yet educational, surprises greet the engineering planner as he takes his ideas to the public. IEEE Spectrum, 12(2):57–60, February 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kauferle:1972:UDL

[Kau72] J. Kauferle. Using dc links to enhance ac system performance. IEEE Spectrum, 9 (6):31–37, June 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaya:1978:EIE

[Kay78] A. Kaya. Energy: Industrial energy control: The computer takes charge: Limited intelligence will suffice for local control, but computers are needed for overall system control. IEEE Spectrum, 15(7):48–53, July 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kenville:1971:NLR

[Ken71] R. F. Kenville. Noise in laser recording. IEEE Spectrum, 8 (3):50–57, March 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kertz:1976:HBP

[Ker76] H. L. Kertz. How Bell pinches power: Both design and good fortune play a role in conserving electricity, gasoline, and building fuels for the Bell System. IEEE Spectrum, 13(2):54–58, February 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kesner:1970:MVR

[Kes70] D. Kesner. Monolithic voltage regulators. IEEE Spectrum, 7(4):24–32, April 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Karady:1973:TVH

[KG73] G. Karady and T. Gilsig. The thyristor valve in HVDC transmission. IEEE Spectrum, 10(12): 36–43, December 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kock:1977:BR

[KH77] W. E. Kock and P. K. Houpt.

Book reviews. IEEE Spectrum, 14(6):66–67, June 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kasper:1979:MAO

[KH79] J. F. Kasper and C. E. Hutchinson. Marine/aviation: Omega: Global navigating by VLF fix: This internationally sponsored global VLF system is nearing full operational status. IEEE Spectrum, 16(5):59–63, May 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kitano:1977:AEB

[KI77] T. Kitano and S. Ito. Automotive electronics: Betting on the future: Dramatic experiments may help improve the already bright outlook for the Japanese automotive industry. IEEE Spectrum, 14 (9):75–77, September 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kinn:1971:PSE

[Kin71] J. M. Kinn. The present status of engineering employment. IEEE Spectrum, 8(11):37–38, November 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kinn:1972:IND

[Kin72] J. M. Kinn. For IEEE: new dimensions. IEEE Spectrum, 9 (12):72–73, December 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

King:1975:AAS

[Kin75] J. K. King. Avionics I: Air safety as seen from the tower: a veteran of more than 33 years in air traffic control says hazards lurk in proposed FAA procedures. IEEE Spectrum, 12 (8):67–71, August 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kaplan:1976:NPP

[KJ76] Gadi Kaplan and R. K. Jurgen. Nuclear power plant safetyi. IEEE Spectrum, 13(5):52–69, May 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kressel:1973:PLD

[KLE73] H. Kressel, H. F. Lockwood, and M. Ettenberg. Progress in laser diodes. IEEE Spectrum, 10(5):59–64, May 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kleitman:1977:INC

[Kle77] D. Kleitman. I2L: Novel circuits and formats. IEEE Spectrum, 14(6):35–38, June 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Klein:1974:PHI

[KLH74] M. Klein, J. Linnenkohl, and H. Heidenreich. Power: HVDC to illuminate darkest Africa: By 1975, electric energy will be produced in northwest Mozambique in one of the largest hydroelectric systems under construction. IEEE Spectrum, 11 (10):51–58, October 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kenney:1979:MOD

[KLM^{+}79] G. C. Kenney, D. Y. K. Lou, R. McFarlane, A. Y. Chan, J. S. Nadan, T. R. Kohler, J. G. Wagner, and F. Zernike. Memories: an optical disk replaces 25 mag tapes: This 30-cm disk features a long-life tellurium film upon which data is impressed by laser. IEEE Spectrum, 16 (2):33–38, February 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Knoblock:1975:III

[KLV75] D. E. Knoblock, D. C. Loughry, and C. A. Vissers. Instrumentation: Insight into interfacing: Thorough and unambiguous understanding of interface systems demands sound design concepts and a precise descriptive language. IEEE Spectrum, 12(5):50–57, May 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kino:1971:SPA

[KM71] G. S. Kino and H. Matthews. Signal processing in acoustic surface-wave devices. IEEE Spectrum, 8(8):22–35, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kohl:1971:ATF

[Koh71] W. H. Kohl. Assessment, transfer, and forecasting of technology. IEEE Spectrum, 8(1): 70–75, January 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kopelman:1977:PDC

[Kop77] R. Kopelman. Personal development: Career hurdles and how to clear them: Will an advanced degree and adherence to the work ethic suffice in moving ahead? IEEE Spectrum, 14 (2):66–69, February 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Koren:1977:CCB

[Kor77] Y. Koren. Computers: Computerbased machine-tool control: The best place to inject a computer in numerical control is locally, not as a controller of many systems. IEEE Spectrum, 14 (3):81–83, March 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Krause:1975:ATO

[Kra75] S. Krause. Applications: Troubleshooting with oscillographs:

Classic in concept, the lightbeam recording oscillograph preserves multichannel data for detailed examination. IEEE Spectrum, 12(11):30–33, November 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kelly:1978:ALA

[KRS78] R. J. Kelly, H. W. Redlien, and J. L. Shagena. Aerospace: Landing aircraft under poor conditions: Time-reference scanning beams characterize the allweather microwave landing system that will replace present systems. IEEE Spectrum, 15(9): 52–57, September 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Konnerth:1970:OTU

[KS70] K. L. Konnerth and B. R. Shah.

Optical transmission utilizing injection light sources. IEEE Spectrum, 7(9):37–46, September 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kritikos:1979:AMS

[KS79] H. N. Kritikos and J. Shiue.

Aerospace: Microwave sensing from orbit: Enhanced devices penetrate cloud cover to provide data for weather prediction and terrain analysis. IEEE Spectrum, 16(8):34–41, August 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kunze:1975:LSE

[Kun75]  H. Kunze.  Large systems:

EWS: Germany’s answer to ESS: Stored-program control is at the heart of the Deutsche Bundespost’s electronic telephone switching system. IEEE Spectrum, 12(11):51–54, November 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Kusko:1972:SSM

[Kus72] A. Kusko. Solid-state motorspeed controls. IEEE Spectrum, 9(10):50–55, October 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lapidus:1972:DMB

[Lap72a] G. Lapidus. The domain of magnetic bubbles. IEEE Spectrum, 9 (9):58–62, September 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lapidus:1972:MLL

[Lap72b] G. Lapidus. MOS/LSI launches the low-cost processor. IEEE Spectrum, 9(11):33–40, November 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lapidus:1973:ADT

[Lap73a] G. Lapidus. Amorphous devices: at the threshold. IEEE Spectrum, 10(1):44–52, January 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lapidus:1973:BLS

[Lap73b] G. Lapidus. Behind the lens of the SX-70. IEEE Spectrum, 10 (12):76–83, December 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lapidus:1973:CPT

[Lap73c] G. Lapidus. Cost-performance tradeoffs: a case history. IEEE Spectrum, 10(9):79–82, September 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lapidus:1973:EC

[Lap73d] G. Lapidus. Electronics in cameras. IEEE Spectrum, 10 (6):22–28, June 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lapidus:1973:FF

[Lap73e] G. Lapidus. Forum on FAMs.

IEEE Spectrum, 10(7):50–54, July 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lapidus:1973:SSE

[Lap73f] G. Lapidus. Solid state: it’s everywhere. IEEE Spectrum, 10 (3):26–30, March 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lapidus:1973:TFH

[Lap73g] G. Lapidus. Transistor family history. IEEE Spectrum, 10 (1):34–35, January 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lapidus:1973:WBB

[Lap73h] G. Lapidus. WESCON ’73: a broad-based, timely program. IEEE Spectrum, 10(8):51–56, August 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lapidus:1974:CSB

[Lap74a] G. Lapidus. Circuit/system building blocks: Recent new devices are faster, and cheaper, and provide a higher function density. IEEE Spectrum, 11 (1):54–60, January 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lapidus:1974:ECH

[Lap74b] G. Lapidus. The energy crisis: How manufacturers cope: Component and equipment makers strive to keep production up while holding tempers and temperatures down. IEEE Spectrum, 11(6):66–73, June 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lapidus:1974:NSS

[Lap74c] G. Lapidus. New solid-state circuits: a 500-kb bucket brigade scanner and a 4-watt audio amplifier are but two of several promising new developments. IEEE Spectrum, 11(1): 98–100, January 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lapidus:1974:PDC

[Lap74d] G. Lapidus. Product design: The computer: Now it’s a component! Large-scale integrated circuits spawned the CPU on a chip, but it raises questions of who designs what. IEEE Spectrum, 11(3):80–82, March 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lapidus:1974:SRI

[Lap74e] G. Lapidus. Special report: IN-TERCOM ’74: Getting down to business: Conference program organizers are serious about new products, new markets, and solving society’s problems, too. IEEE Spectrum, 11(2):90–93, February 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lavendel:1978:CRR

[Lav78] G. A. Lavendel. The computer in research: a real-time partner [information services]. IEEE Spectrum, 15(12):38–40, December 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Laaspere:1975:PEC

[LC75] T. Laaspere and A. O. Converse. Power/energy creative electric load management: New fools for the utilities include rate options, clock-actuated switches, radio control, and ripple control systems. IEEE Spectrum, 12 (2):46–50, February 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lee:1971:ESI

[Lee71] R. H. Lee. Electrical safety in industrial plants. IEEE Spectrum, 8(6):51–55, June 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Levi:1978:PMT

[Lev78] E. Levi. Power: MHD’s target: Payoff by 2000: Efficiencies over 50 percent and generating costs of 32 mills/kWh may make magnetohydrodynamic systems viable by 1995. IEEE Spectrum, 15(5):46–51, May 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Land:1970:EEF

[LH70] C. E. Land and R. Holland. Electrooptic effects in ferroelectric ceramics. IEEE Spectrum, 7 (2):71–78, February 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Liacco:1978:PES

[Lia78] T. E. D. Liacco. Power/energy: System security: The computer’s role: Several securityrelated functions can be aided by the digital computer, and linked together by a software scheme. IEEE Spectrum, 15 (6):43–50, June 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lindgren:1974:SRR

[Lin74a] N. Lindgren. Special report: a return visit with EE–astronaut Owen Garriott: seven years of preparation and a blast off like riding the rumble seat of a Model A … but the two-month trip was worth it! IEEE Spectrum, 11(7):50–60, July 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Linstone:1974:TFP

[Lin74b] H. A. Linstone. Technology forecasting: Planning: Toy or tool? Line managers frequently find little use for formal “forecasts.” Here are guidelines for planners who wish to rectify that situation. IEEE Spectrum, 11(4):42–49, April 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lindgren:1975:EWE

[Lin75a] N. Lindgren. Energy what EPRI is up to: The power utility industry is rooting for its new offspring to uphold its initial high promise. IEEE Spectrum, 12(4):66–69, April 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Linstone:1975:TFV

[Lin75b] H. A. Linstone. Technology forecasting VI: Where to look to look ahead: Articles outnumber books as recommended references on the complex, but intriguing, art of anticipation. IEEE Spectrum, 12(4):60–61, April 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lindgren:1976:CWC

[Lin76a] N. Lindgren. Circa I: What is a city? Each one is a mongrel, bred off unique circumstances, diverse human needs, and burgeoning technologies. IEEE Spectrum, 13(7):31–35, July 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lindgren:1976:CISa

[Lin76b] N. Lindgren. Circa III: Soleri’s arcologies: In Arizona, a visionary architect has decreed a stately pleasure-dome — the concept is a metropolis in miniature. IEEE Spectrum, 13(7):43– 46, July 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lindgren:1976:CISb

[Lin76c] N. Lindgren. Cities IV: Supurbia: The compact city: Layer upon layer, the wedding cake metropolis marries systems design and city planning. IEEE Spectrum, 13(7):47–50, July 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lindgren:1977:SFS

[Lin77] N. Lindgren. Semiconductors face the ’80s: Seven leaders of the boomingest business in the U.S. prognosticate the problems supergrowth could lead to. IEEE Spectrum, 14(10):42–49, October 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Litchford:1975:AMC

[Lit75] G. B. Litchford. Avoiding midair collisions. IEEE Spectrum, 12 (9):41–48, September 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

LD:1970:SSS

[LJ70] Lee D.Dickey and C. M. Jones.

Solid-state switching for aircraft electric systems. IEEE Spectrum, 7(11):73–79, November 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lombardo:1979:DMC

[Lom79] T. G. Lombardo. The decisionmakers: a cacophony of voices: The tangle of politics, finance, bureaucracy, and the persuasive arts yields “go no-go” answers for the nuclear industry. IEEE Spectrum, 16(11):81–91, November 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Loughry:1974:IIW

[Lou74] D. C. Loughry. Instrumentation III: What makes a good interface?: Flexibility, reasonable cost, and compatibility are musts when interconnecting independently manufactured devices into a workable system. IEEE Spectrum, 11(11):52–57, November 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lusignan:1976:CCO

[LPJ76] B. B. Lusignan, J. G. Potter, and J. M. Janky. Communications — the co-op in the sky. IEEE Spectrum, 13(2):59–65, February 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lusignan:1978:CSS

[Lus78] B. Lusignan. Communications: Single-sideband transmission for land mobile radio: From seven to ten times as many channels could be made available in existing VHF and UHF bands. IEEE Spectrum, 15(7):33–37, July 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Letzter:1970:NA

[LW70] S. Letzter and N. Webster. Noise in amplifiers. IEEE Spectrum, 7(8):67–75, August 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Lavi:1973:POD

[LZ73] A. Lavi and C. Zener. Plumbing the ocean depths: a new source of power. IEEE Spectrum, 10 (10):22–27, October 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Maccoby:1973:WLW

[Mac73] M. Maccoby. ‘Winning’ and ‘losing’ at work. IEEE Spectrum, 10(7):39–48, July 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

MacAdam:1975:MMW

[Mac75] W. K. MacAdam. Megawatts from municipal waste. IEEE Spectrum, 12(11):46–50, November 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Machol:1976:CEU

[Mac76] R. E. Machol. Communications: an ear to the Universe: Sensitive antennas, pattern recognizing computers, and informed conjecture are being tuned to possible signals from the stars. IEEE Spectrum, 13 (3):42–47, March 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mackintosh:1978:SSL

[Mac78] I. M. Mackintosh. Solid state:

Large-scale integration: Intercontinental aspects: Factors affecting the competitive positions of U.S., Japanese, and European firms include current business and economic climates. IEEE Spectrum, 15(6):51–56, June 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Madden:1974:ECA

[Mad74]  C. H. Madden.  Essay: a

challenge to the aging corporation. IEEE Spectrum, 11(9): 88–90, September 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Manasse:1970:GES

[Man70] F. K. Manasse. Getting engineering students involved: The Dartmouth/Tuskegee program. IEEE Spectrum, 7(12):59–64, December 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Manasse:1972:DCL

[Man72] F. K. Manasse. The Dartmouth concept: learning in the real world. IEEE Spectrum, 9 (2):36–40, February 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mandell:1978:EWC

[Man78] M. Mandell. Electronic watchdogs for calls and copies. IEEE Spectrum, 15(4):31–35, April 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mandey:1979:ANS

[Man79] T. Mandey. Assessing the new services: Political and legal issues will impact implementation of broadband digital transmission. IEEE Spectrum, 16 (10):42–50, October 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Martino:1972:TLA

[Mar72] J. P. Martino. Tools for looking ahead. IEEE Spectrum, 9 (10):32–40, October 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Martin:1977:COW

[Mar77] T. B. Martin. Communications: One way to talk to computers: Voice commands to computers may substitute in part for conventional input devices. IEEE Spectrum, 14 (5):35–40, May 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mason:1979:SAC

[Mas79a] J. Mason. Strategic arms control: SALT II: a state-of-the-art treaty: Compliance must be verifiable by technical means, so technology determines the treaty’s leeway and its restrictions. IEEE Spectrum, 16 (5):64–65, May 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mason:1979:AMP

[Mas79b] J. F. Mason. Aerospace and military: Progress in space structure research, aircraft landing systems, integrated optics, and digital communications. IEEE Spectrum, 16(1):71–75, January 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mason:1979:MPB

[Mas79c] J. F. Mason. Military: particlebeam weapons: a controversy: ‘highly questionable’ is the label applied by four M.I.T. scientists, but the U.S. Department of Defense is betting on the new scheme. IEEE Spectrum, 16(6):30–35, June 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mason:1979:PSS

[Mas79d] J. F. Mason. The profession: The secret society that never was: Founded in the fires of creative disagreement, Eta Kappa Nu soon found its niche, and this month celebrates its 75th anniversary. IEEE Spectrum, 16 (9):55–57, September 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mason:1979:TBB

[Mas79e] J. F. Mason. The technical blowby-blow: Details of the Three Mile Island accident as excerpted and edited from interviews with Nuclear Regulatory Commission investigators. IEEE Spectrum, 16 (11):33–58, November 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mason:1979:TLS

[Mas79f] J. F. Mason. Transportation. The LIRR: a system riding troubled tracks: Will the aggravations of its 100,000 commuters ever end? You are invited to help the railroad solve its problems. IEEE Spectrum, 16(12): 33–41, December 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mauchly:1975:RMT

[Mau75] J. W. Mauchly. Retrospective: Mauchly on the trials of building ENIAC: People could not see how something with 18,000 tubes and costing $500,000 could ever become practical! IEEE Spectrum, 12(4):70–76, April 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Maxwell:1977:LRP

[Max77] R. L. Maxwell. Looking down the road: a projection of how five major factors will impact the car of tomorrow. IEEE Spectrum, 14 (11):97–98, November 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McKevitt:1979:NOB

[MB79] J. McKevitt and J. Bayliss. New options from big chips. IEEE Spectrum, 16(3):28–34, March 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McAleer:1971:LAT

[McA71] H. T. McAleer. A look at automatic testing. IEEE Spectrum, 8(5):63–78, May 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McKenzie:1971:ES

[McK71]  A. A. McKenzie.  Engineer

salaries. IEEE Spectrum, 8 (8):36–40, August 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McKenzie:1972:SCP

[McK72a] A. McKenzie. Sarnoff: Controversial pioneer. IEEE Spectrum, 9(1):40–41, January 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McKenzie:1972:PEI

[McK72b] A. A. McKenzie. PABX: the epitome of ‘interconnection’. IEEE Spectrum, 9(12):42–52, December 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McKenzie:1972:WS

[McK72c] A. A. McKenzie. What’s up in satellites. IEEE Spectrum, 9(5):16–27, May 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McKenzie:1973:OWE

[McK73] A. A. McKenzie. One wheel is enough. IEEE Spectrum, 10 (4):55–56, April 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McKenzie:1974:TJM

[McK74] A. A. McKenzie. The three jewels of marconi: Most men achieve greatness with one idea; the father of radio produced two more. IEEE Spectrum, 11(12):46, December 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

McWilliams:1977:CAA

[McW77] G. I. McWilliams. The consumer and the automobile: Auto buyers don’t always choose what’s best for them, nor do makers always read consumers accurately. IEEE Spectrum, 14(11):93–96, November 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Meinert:1970:MEN

[ME70] J. Meinert and K. P. Ewelt.

A modernized electric network for the future needs of Hamburg. IEEE Spectrum, 7(6): 41–46, June 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Meyer-Ebrecht:1977:EMI

[ME77] D. Meyer-Ebrecht. Electronics in medicine: Imaging technology: a European survey: New devices and techniques provide more revealing images for medical diagnoses. IEEE Spectrum, 14(5):46–50, May 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1973:SFC

[Men73a] D. Mennie. Safety first for the consumer. IEEE Spectrum, 10 (11):50–51, November 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1973:WNC

[Men73b] D. Mennie. What’s new in consumer electronics circuits. IEEE Spectrum, 10(12):26–35, December 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1974:CRV

[Men74a] D. Mennie. Computer report V:

Power supplies: How to trim the budget: Orphaned in the world of decreasing cost-per-bit, power conditioning equipment cries for cost reduction. IEEE Spectrum, 11(2):57–61, February 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1974:DTB

[Men74b] D. Mennie. Designers’ tools: The big roundup of small calculators: Hand-held calculators are getting smaller and smarter; they make the slide rule a museum piece. IEEE Spectrum, 11(4):34–41, April 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1974:ETB

[Men74c] D. Mennie. Education trends: The B.E.T.: Engineer or supertechnician?: Is the bachelor of technology a needed link between engineer and technician … or simply the educator’s answer to sagging enrollments? IEEE Spectrum, 11(5):60–63, May 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1974:III

[Men74d] D. Mennie. Instrumentation IV: Have instrument, will travel: Be the readout analog, digital, CRT, or hard copy, “portable” and “performance” are now in the same package. IEEE Spectrum, 11(11):58–63, November 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1974:MAD

[Men74e] D. Mennie. Military and aerospace: Dollars for many projects may be grounded, but efforts to spend wisely are soaring. IEEE Spectrum, 11(1): 71–77, January 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1974:PCP

[Men74f] D. Mennie. Perspective: Consumer protection with teeth: a year old and fully staffed, the U.S. consumer product safety commission has become a watchdog with a bite. IEEE Spectrum, 11(10):96–97, October 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1974:SCT

[Men74g] D. Mennie. Systems and cybernetics: Tools of the discipline are progressing from the inspirational to the practical. IEEE Spectrum, 11(1):85, January 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1975:CMF

[Men75a] D. Mennie. Communications and microwave: Fiber optics, solid state, and digital techniques boost transmission efficiency. IEEE Spectrum, 12 (1):43–48, January 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1975:CET

[Men75b] D. Mennie. Consumer electronics: Television on a silver platter: a proliferation of options may spawn some orphans in the contest among developers of video disks. IEEE Spectrum, 12(8):34–39, August 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1975:MOW

[Men75c] D. Mennie. Microwave ovens: what’s cooking? IEEE Spectrum, 12(3):34–39, March 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1975:TVI

[Men75d] D. Mennie. Tools v/information retrieval: Engineering micrographics: Reduced from handmade originals or prepared directly by computer, microfilm data files claim size, access, and integrity advantages. IEEE Spectrum, 12(10):62–66, October 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1975:TCT

[Men75e] D. Mennie. ‘Troubleshooting’ color TV. IEEE Spectrum, 12 (6):57–62, June 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1976:CXP

[Men76a] D. Mennie. Cities x: People movers: The object is to make ‘easy riders’ of the majority of citizens as they pursue their daily chores and pleasures. IEEE Spectrum, 13(7):85– 90, July 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1976:CEO

[Men76b] D. Mennie. Communications:

Electronics and optics, too: Solid-state lasers, optical waveguide, and digital techniques for TV broadcasting come on line. IEEE Spectrum, 13(1): 57–63, January 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1976:CEM

[Men76c] D. Mennie. Communications:

The explosion at 27 mhz: Citizens band fever in epidemic proportions is straining technology, bandwidth, and the regulators, too. IEEE Spectrum, 13(5):36–41, May 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1976:CBT

[Men76d] D. Mennie. Componentsbatteries: today and tomorrow. IEEE Spectrum, 13(3):36–41, March 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1976:CEE

[Men76e] D. Mennie. Consumer electronics: Electronic gamesmanship: It’s man vs. microcircuit as video games entertain by challenging human skill (now) and intellect (soon). IEEE Spectrum, 13(12): 27–30, December 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1976:CEF

[Men76f] D. Mennie. Consumer electronics: Fun and function: John Q. Public goes digital with calculators, watches, sewing machines, and air conditioners. IEEE Spectrum, 13(1):84–86, January 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennle:1976:WWW

[Men76g] Don Mennle. What went wrong II: A try at automated ‘zip’. IEEE Spectrum, 13(10): 41–48, October 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1977:CMT

[Men77a] D. Mennie. Communications and microwave technology: in hardware, it’s fiber optics and digital radios; in “software,” AT&T and FCC trade parries. IEEE Spectrum, 14(1):43–47, January 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1977:CEE

[Men77b] D. Mennie. Consumer electronics: everybody’s doing it (‘computing’ at home): the demand for owner-operated microcomputers has blossomed, creating a brand-new market for kits, peripherals, and software. IEEE Spectrum, 14(5):29–34, May 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1977:MFC

[Men77c] D. Mennie. More fun for the consumer: Cb explodes and home video games capture the consumer’s fancy. IEEE Spectrum, 14(1):66–69, January 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1977:SCE

[Men77d] D. Mennie. Self-contained electronic games: New toys, from chess to football and all designed with dedicated microcircuits and LEDs make their debut this holiday season. IEEE Spectrum, 14 (12):21–26, December 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1977:WEW

[Men77e] D. Mennie. Where the EE works:

Is your cubbyhole, workbench, or office less than what you’d like it to be? Here are some suggestions. IEEE Spectrum, 14(7):40– 46, July 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1978:SFC

[Men78a] D. Mennie. Am stereo: five competing options. IEEE Spectrum, 15(6):24–31, June 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1978:CMT

[Men78b] D. Mennie. Communications and microwave technology: Optical fibers compete with premium coax; digital radio reaches cost effectiveness. IEEE Spectrum, 15 (1):31–35, January 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1978:CPC

[Men78c] D. Mennie. Computers: Personal computers for the entrepreneur: Microcomputers can help meet payrolls rather than play star wars availability of applications software packages is the pacing item. IEEE Spectrum, 15(9): 30–35, September 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1978:CEH

[Men78d] D. Mennie. Consumer electronics: Helping speeders beat the radar rap: Moving radar and dual-band detectors constitute a miniature ECM race between enforcement-minded authorities and motoring outlaws. IEEE Spectrum, 15(8):38–42, August 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1978:CVT

[Men78e] D. Mennie. Consumer: Video tape: Options for the buyer: Despite competing tape speeds, cassette sizes, and recording formats, the public is buying players, cameras, timers, and software. IEEE Spectrum, 15 (3):37–41, March 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1978:CPT

[Men78f] D. Mennie. Consumers as programmers: TV networks are upstaged by VCRs and video games; turnkey personal computers plunge in price. IEEE Spectrum, 15(1):54–58, January 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1978:MHS

[Men78g] D. Mennie. Most are hardened skeptics, yet production workers may be guided to higher productivity and greater job satisfaction. IEEE Spectrum, 15 (10):81–84, October 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1979:CMT

[Men79a] D. Mennie. Communications and microwave technology: The communications giants vie for interconnect turf; fiber optics go online; teletext is tested. IEEE Spectrum, 16(1):38–42, January 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mennie:1979:CEP

[Men79b] D. Mennie. Consumer electronics: Personal and plentiful: Optical video disks debut in market tests; in the wings are digital audio, AM stereo, and flat screen TV. IEEE Spectrum, 16 (1):62–66, January 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Malkin:1979:APS

[MF79] M. S. Malkin and R. F. Freitag. Aerospace: Planning Space Shuttle’s maiden voyage: In a matter of months space-oriented experiments will soar aloft aboard the Columbia. IEEE Spectrum, 16(7):42–48, July 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Michal:1978:PET

[Mic78] J. Michal. Power/energy: TVA’s Freeman settles in. IEEE Spectrum, 15(9):64–65, September 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Middendorf:1971:SEC

[Mid71]  W. H. Middendorf.  Stan-

dards: The evidence of concern. IEEE Spectrum, 8(8):70–73, August 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Middendorf:1974:PFC

[Mid74] W. H. Middendorf. Power: Fuses and circuit breakers: Dispelling some myths: Both have advantages and shortcomings. For home and light commercial use, a hybrid fuse-breaker combination is recommended. IEEE Spectrum, 11(8):67–72, August 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Midwinter:1979:IEB

[Mid79]  J. E. Midwinter.  III. Eu-

rope: Business benefits. IEEE Spectrum, 16(10):76–77, October 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mills:1971:PCF

[Mil71] R. G. Mills. The promise of controlled fusion. IEEE Spectrum, 8 (11):24–36, November 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Miller:1972:HER

[Mil72] G. H. Miller. How EEs rate math courses. IEEE Spectrum, 9(1):55–58, January 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mittelmann:1977:FBC

[Mit77] E. Mittelmann. Fast bonding cuts auto costs: an important development speeds assembly through welding of plastic to metal. IEEE Spectrum, 14 (11):73–76, November 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mulvey:1974:IGK

[MM74] J. Mulvey and J. Millay. Instrumentation: Get to know op amps; use a curve tracer: For the out of the ordinary application, it is not surprising to find op amp spec sheets wanting. IEEE Spectrum, 11(9):53–58, September 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Machover:1977:GDF

[MNS77a] C. Machover, M. Neighbors, and C. Stuart. Graphics displays: Factors in systems design. IEEE Spectrum, 14(10):23–27, October 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Machover:1977:GDU

[MNS77b] C. Machover, M. Neighbors, and C. Stuart. Graphics displays: Users are enticed by decreasing hardware costs and greater availability of software. IEEE Spectrum, 14(8):24–33, August 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mokhoff:1979:CSR

[Mok79a] N. Mokhoff. Communications: a special report. IEEE Spectrum, 16(10):28–29, October 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mokhoff:1979:DSD

[Mok79b] N. Mokhoff. Distributed software for distributed hardware. IEEE Spectrum, 16(10):57–60, October 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mokhoff:1979:OAC

[Mok79c] N. Mokhoff. Office automation: a challenge: Better tools simplify some office tasks, but serious problems block the integrated office of the future. IEEE Spectrum, 16(10):66–69, October 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Monahan:1976:WW

[Mon76] A. M. Monahan. The wide world of ‘128’. IEEE Spectrum, 13(5):94–95, May 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:1973:PLL

[Moo73] A. W. Moore. Phase-locked loops for motor-speed control. IEEE Spectrum, 10(4):61–67, April 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moore:1979:SSV

[Moo79] G. Moore. Solid state: VLSI:

Some fundamental challenges: Defining and designing the products made possible by very-large-scale integration are first on the list of priority tasks. IEEE Spectrum, 16(4):30, April 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Morrison:1978:CMI

[Mor78] R. L. Morrison. Computers: Microcomputers invade the linear world: Analog interfacing techniques are in demand as microcomputers are exploited in inexpensive equipment for linear applications. IEEE Spectrum, 15(7):38–41, July 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moschytz:1970:IFSa

[Mos70a] G. S. Moschytz. Inductorless filters: a survey I. Electromechanical filters. IEEE Spectrum, 7 (8):30–36, August 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moschytz:1970:IFSb

[Mos70b] G. S. Moschytz. Inductorless filters: a survey II. Linear active and digital filters. IEEE Spectrum, 7(9):63–75, September 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moschytz:1970:OAL

[Mos70c] G. S. Moschytz. The operational amplifier in linear active networks. IEEE Spectrum, 7 (1):42–50, January 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moss:1974:IFI

[Mos74] S. F. E. Moss. Impediments to freedom of information. IEEE Spectrum, 11(4):58–59, April 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Moskowitz:1975:TFL

[Mos75] R. Moskowitz. Technology ferrofluids: Liquid magnetics: a space-age research spin-off finds its way out of the laboratory and into a host of lubricating and damping applications. IEEE Spectrum, 12 (3):53–57, March 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mayfield:1977:IBB

[MRR^{+}77] E. Mayfield, E. Rechtin, J. D.

Ryder, D. Sinclair, R. Tannehill, D. Christiansen, and E. Rubinstein. The IEEE as a black box: Five members ranging in both experience and philosophy face off in a discussion of where we are and ought to be. IEEE Spectrum, 14(8):58–63, August 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mitroff:1973:WBH

[MT73] I. I. Mitroff and M. Turoff. The whys behind the hows. IEEE Spectrum, 10(3):62–71, March 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mulligan:1971:IR

[Mul71] J. H. Mulligan. IEEE reports for 1970. IEEE Spectrum, 8 (6):56–68, June 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Munson:1978:RQT

[Mun78] G. E. Munson. Robots quietly take their place alongside humans on the production line to raise productivity — and do the ‘dirty work’. IEEE Spectrum, 15 (10):66–70, October 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Mezrieh:1975:UIE

[MVE75] R. Mezrieh, D. Vilkomerson, and K. Etzold. Ultrasonic imaging exploits phase contrast: One way to ultrasonically image objects with negligible acoustic absorption is to use an optical microscopy technique. IEEE Spectrum, 12(12):35–39, December 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nagel:1974:EEP

[Nag74] T. J. Nagel. Energy: Electric power’s role in the U.S. energy crisis: a rapid increase in electrification is consistent with a necessary transition away from excessive dependence on oil and gas. IEEE Spectrum, 11(7):69– 72, July 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nash:1978:CES

[Nas78] R. Nash. Classic examples show how the government can influence, directly or indirectly, innovation and productivity. IEEE Spectrum, 15(10):85–86, October 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Naylor:1971:DASa

[Nay71a] J. R. Naylor. Digital and analog signal applications of operational amplifiers I — Multiplexers and converters. IEEE Spectrum, 8(5):79–87, May 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Naylor:1971:DASb

[Nay71b] J. R. Naylor. Digital and analog signal applications of operational amplifiers II — sample/hold modules, peak detectors, and comparators. IEEE Spectrum, 8(6):38–46, June 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nichol:1971:MEC

[NC71] M. Nichol and M. A. Cuevas.

Marketing as an engineering career. IEEE Spectrum, 8(9): 80–81, September 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Newcombe:1976:IHE

[New76] C. Newcombe. Instrumentation i: How to evaluate microprocessor instruments: On the plus side: Flexibility, reliability, interlace standardization; but some inferior designs confuse the user. IEEE Spectrum, 13(4):38–39, April 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nilles:1976:CXT

[Nil76] J. M. Nilles. Cities xi: Talk is cheaper: And so may be other forms of telecommuting, weighed against the time, energy, and expense of moving oneself. IEEE Spectrum, 13(7):91– 94, July 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Nuese:1972:FL

[NKL72] C. J. Nuese, H. Kressel, and I. Ladany. The future for LEDs. IEEE Spectrum, 9 (5):28–38, May 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Norris:1974:SEA

[NR74] P. M. Norris and C. W. Rea. Space electronics: Automatic checkout for missions to mars: Unmanned extraterrestrial landings call for reliable premission testing of complex guidance and communications systems. IEEE Spectrum, 11(10):42–50, October 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Niepoth:1977:CLE

[NS77] G. W. Niepoth and S. P. Stonestreet. Closed-loop engine control: Analog or digital circuits to aid emissions control may be the forerunner of onboard engine tuning. IEEE Spectrum, 14 (11):53–56, November 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Obermaier:1977:GWR

[Obe77] A. Obermaier. The ‘green wave’ rolls on [computerised road traffic control]. IEEE Spectrum, 14 (3):75–80, March 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Orr-Ewing:1970:I

[OE70]  I. Orr-Ewing.  Introduction.

IEEE Spectrum, 7(1):16c–16c, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Olyphant:1977:BR

[OGG77] M. Olyphant, N. Gorchow, and M. D. Godfrey. Book reviews. IEEE Spectrum, 14 (8):69–75, August 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ohigashi:1970:MMC

[Ohi70] H. Ohigashi. Man-machine conversational data communication systems. IEEE Spectrum, 7 (3):77–82, March 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Okress:1978:SEF

[Okr78] E. C. Okress. Satellite/energy:

The Franklin Institute has high hopes for its big balloon: a mile-long hot-air balloon is proposed for space research and solar power generation. IEEE Spectrum, 15(12):41–47, December 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oliver:1972:TNM

[Oli72] B. M. Oliver. Toward a new morality. IEEE Spectrum, 9 (1):52–54, January 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oliver:1974:ILA

[Oli74] B. M. Oliver. Instrumentation I: Looking ahead: Optical fibers, and tunable coherent sources and filters at infrared and optical frequencies, may key the future in test equipment. IEEE Spectrum, 11(11):44–45, November 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Olling:1978:ELA

[Oll78] G. Olling. Experts look ahead to the day of the full-blown computer-integrated automatic factory. IEEE Spectrum, 15 (10):60–66, October 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Olson:1971:EMS

[Ols71] H. F. Olson. Electronic music synthesis for recordings. IEEE Spectrum, 8(4):18–30, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Orchard:1976:SS

[OP76] L. Orchard and W. Peek. ‘Super’ scopes. IEEE Spectrum, 13(4):73–77, April 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Oppenheim:1970:SSU

[Opp70] A. V. Oppenheim. Speech spectrograms using the Fast Fourier Transform. IEEE Spectrum, 7 (8):57–62, August 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Park:1978:PHS

[Par78a] R. H. Park. Power: Hardware and software for system protection: a survey of control techniques that minimize both the occurrence and the duration of blackouts. IEEE Spectrum, 15(5):40–45, May 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Parsons:1978:PUL

[Par78b] G. P. Parsons. Patents: U.S.

lags in patent law reform: The employed inventor in countries other than the U.S. often is given better incentives to create, plus greater statutory protection. IEEE Spectrum, 15 (3):60–64, March 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patton:1979:EES

[Pat79a] C. Patton. Education: Equipment shortfall in EE labs: Undergraduate schools for the most part have substandard, obsolete equipment and it’s in short supply, too. IEEE Spectrum, 16 (3):65–69, March 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Patton:1979:ECK

[Pat79b] C. Patton. Education: The challenge of keeping current: How best to keep a step ahead of one’s job is debatable; the choices include several outlined herewith. IEEE Spectrum, 16 (8):53–59, August 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Payne:1975:CIT

[Pay75] M. Payne. Communications II: Telephones in the U.K.: Cautious progress: a few false starts, then a low-risk selection of staid switching technology. IEEE Spectrum, 12 (6):42–45, June 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pell:1970:L

[Pel70] A. R. Pell. Laid off! IEEE Spectrum, 7(9):34–36, September 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Petritz:1977:PMT

[Pet77]  R. L. Petritz.  The perva-

sive microprocessor: Trends and prospects: The progressive EE cannot afford to overlook new microprocessor technologies and applications. The author, an expert, tells why. IEEE Spectrum, 14(7):18–25, July 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pettit:1970:SIT

[PG70] J. M. Pettit and D. J. Grace. The Stanford Instructional Television Network. IEEE Spectrum, 7(5):73–80, May 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pensabene:1976:BUM

[PG76] S. F. Pensabene and J. W. Gould. Batteries: Unwanted memory spooks nickel-cadmium cells: New understanding of an undesirable phenomenon is the first step in avoiding it. IEEE Spectrum, 13(9):33–37, September 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Proxmire:1976:PIR

[PHC76] W. Proxmire, D. M. Heller, and M. R. Currie. Perspective: Independent R&D: Three viewpoints: a key question is to what extent iR&D costs should be reimbursed by government agencies and under what controls. IEEE Spectrum, 13(11):50–53, November 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pierce:1979:BIS

[Pie79] J. R. Pierce. On Bell: II.

The system is sound: ‘it ain’t broke, so don’t fix it,’ reflects the sentiments of an eminent Bell Labs emeritus executive. IEEE Spectrum, 16(12):48–51, December 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pitroda:1979:TGD

[Pit79] S. G. Pitroda. Telephones go digital: Lower cost, higher performance and greater flexibility favor the digital approach over analog switching. IEEE Spectrum, 16(10):51–55, October 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Platts:1979:PEL

[Pla79a] J. Platts. Power: Electrical load management: The British experience: Reduced utility rates and heaters with storage have been increasing winter off-peak loads. IEEE Spectrum, 16(2): 39–43, February 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Platts:1979:PEE

[Pla79b] J. Platts. Power/energy: Electrical load management: The British experience, part II: Local and remote controls coupled with off-peak rates have forced nighttime demand to record highs. IEEE Spectrum, 16(4):66–68, April 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Peil:1977:CES

[PM77] W. Peil and R. J. McFadyen. Consumer electronics: Singleslice superhet: New portable and table model AM/FM radios are designed with most active components on one 16-pin bipolar chip. IEEE Spectrum, 14(3):54–60, March 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Puckett:1977:AEI

[PMG77] G. Puckett, J. Marley, and J. Gragg. Automotive electronics II: The microprocessor is in: This tiny IC will satisfy the government’s requirements for energy and pollution control as well as the consumer’s preference for luxury. IEEE Spectrum, 14(11): 37–46, November 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pogust:1978:AML

[Pog78] F. Pogust. Avionics: Microwave landing systems: As MLS technology enters its third decade, competing approaches contend for standardization at the international level. IEEE Spectrum, 15(3):30–36, March 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Poloujadoff:1970:EDV

[Pol70] M. Poloujadoff. The Esclangon diagram for voltage and current along a transmission line. IEEE Spectrum, 7(2): 92–94, February 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Poloujadoff:1971:LIMa

[Pol71a] M. Poloujadoff. Linear induction machines I. History and theory of operation. IEEE Spectrum, 8 (2):72–80, February 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Poloujadoff:1971:LIMb

[Pol71b] M. Poloujadoff. Linear induction machines II. Applications. IEEE Spectrum, 8(3):79–86, March 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Pollack:1979:SS

[Pol79]  L. Pollack.  Switching in

space. IEEE Spectrum, 16(10): 32–37, October 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ponte:1970:REP

[Pon70] M. Ponte. The role of the engineer in public affairs. IEEE Spectrum, 7(7):51–58, July 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Prem:1970:OBP

[PP70] L. L. Prem and W. E. Parkins. Outlook for binary power plants using liquid-metal MHD. IEEE Spectrum, 7(5):35–44, May 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Punchard:1970:WAC

[Pun70] J. C. R. Punchard. What’s ahead in communications? IEEE Spectrum, 7(1):51–54, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Puetite:1971:DAS

[PW71] J. G. Puetite and A. M. Werth. Demand-assigned service for the Intelsat global network. IEEE Spectrum, 8(1):59–69, January 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Quaglione:1979:IGE

[Qua79] G. Quaglione. The Intelsat 5 generation: Electronics and advanced rocketry have kept pace with long-distance traffic demands to nurture new generation satellites. IEEE Spectrum, 16 (10):38–41, October 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Queyssac:1979:CPV

[Que79] D. Queyssac. Computers: Projecting VLSI’s impact on microprocessors: Single-chip microcomputers with 1 million bits of information, English-language programming, and canned software are seen ahead. IEEE Spectrum, 16(5):38–41, May 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rabinow:1978:HIU

[Rab78] J. Rabinow. How to improve the U.S. patent system, and encourage the ‘middle level’ inventor. IEEE Spectrum, 15 (10):48–50, October 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rackman:1978:IPT

[Rac78] M. I. Rackman. Inventors: ‘protect thyself’: Careful attention to the claims section will go far toward establishing patent validity and extending the scope of protection. IEEE Spectrum, 15 (2):54–60, February 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Raff:1978:PSC

[Raf78] S. Raff. In praise of the small company and the technical manager as innovators: They’re not afraid to take a chance. IEEE Spectrum, 15(10):80–81, October 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rainey:1976:ANT

[Rai76] F. Rainey. Archaeology: New tools for an old art: Thermoluminescence, magnetometers, satellite photography, and other techniques and instruments aid the archaeologist. IEEE Spectrum, 13(9):39–43, September 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rasmussen:1975:NPR

[Ras75]  N. C. Rasmussen.  Nuclear

power: Rasmussen on reactor safety: How nuclear power reactor risks are quantified; and nuclear sabotage, theft, shipping, and waste disposal risks put in perspective. IEEE Spectrum, 12 (8):46–57, August 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ray:1977:ICS

[Ray77] W. Ray. I2L: Complementing Schottky TTL. IEEE Spectrum, 14(6):35, June 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Reason:1979:ITJ

[Rea79] J. Reason. International trade:

Japan’s electronics markets: a pair of views: Some denigrate dumping practices; others say the U.S. and other countries must learn how to compete. IEEE Spectrum, 16 (6):50–53, June 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rechtin:1975:CEM

[Rec75] E. Rechtin. Careers equations for managers: Management “by the numbers” has been denigrated, yet the author, with tongue only partially in cheek, raises provocative possibilities. IEEE Spectrum, 12(3):58–61, March 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Reitman:1974:CAD

[Rei74] J. Reitman. Computer applications: Design decisions through simulation: Continuous and discrete-event techniques are not necessarily mutually exclusive; they may work well together. IEEE Spectrum, 11(3):76–79, March 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Reudink:1978:CSB

[Reu78] D. O. Reudink. Communications: Spot beams promise satellite communication breakthrough: Focused antenna beams with frequencies accessed by time division can mean higher uplink power and more powerful communication service. IEEE Spectrum, 15(9):36–43, September 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1975:PFC

[RF75] E. Rubinstein and G. D. Friedlander. Power/finance cash crunch dims power prospects: Money doesn’t buy what it used to, an inflationary fact that’s set back the utilities’ expansion programs by years. IEEE Spectrum, 12(3):40–44, March 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rippon:1979:NPA

[RG79] S. Rippon and G. Greenhalgh. Nuclear power around the world: Where the plants are located and a country-by-country projection showing nuclear’s part in total electric energy production. IEEE Spectrum, 16(11):97–98, November 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rumbaugh:1979:PET

[RGCS79] J. H. Rumbaugh, T. F. Garrity, R. Cohen, and R. L. Sullivan. Power/energy: Thermal energy conversion: Tapping the sea depths: Floating energy plants could convert heat stored at different ocean levels to electric power. IEEE Spectrum, 16 (8):42–48, August 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rieder:1970:CBPa

[Rie70a] W. Rieder. Circuit breakers:

Physical and engineering problems I — fundamentals. IEEE Spectrum, 7(7):35–43, July 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rieder:1970:CBPb

[Rie70b] W. Rieder. Circuit breakers physical and engineering problems II — design considerations. IEEE Spectrum, 7(8):90–94, August 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rieder:1970:CBPc

[Rie70c] W. Rieder. Circuit breakers physical and engineering problems III — arc-medium considerations. IEEE Spectrum, 7(9): 80–84, September 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rines:1970:PPB

[Rin70] R. H. Rines. A plea for a proper balance of proprietary rights. IEEE Spectrum, 7 (4):41–46, April 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rippon:1979:NPW

[Rip79] S. Rippon. Nuclear power in Western Europe. IEEE Spectrum, 16(11):98–105, November 1979. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Riter:1973:SDE

[RJD73] S. Riter, W. B. Jones, and H. Dozier. Speeding the deployment of emergency vehicles. IEEE Spectrum, 10(12): 56–62, December 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Reddoch:1979:PEN

[RK79] T. W. Reddoch and J. W. Klein. Power/energy: No ill winds for New Mexico utility: Under test in one municipality’s power system, an experimental machine supplies up to 15 percent of community needs. IEEE Spectrum, 16(3):57–61, March 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Russo:1974:CHS

[RL74] P. M. Russo and M. D. Lippman. Case history: Store and forward: Here, in one system, are interfaces for communications, a floppy disk, and a TV display. IEEE Spectrum, 11(9): 60–67, September 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1975:TIA

[RM75] E. Rubinstein and B. G. Music. Toys IV/avocation: The leisure bookshelf: Wherein a selection of topical books and useful magazines is offered for the hobbyist reader. IEEE Spectrum, 12 (10):81–82, October 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Richter:1976:PBT

[RO76] P. S. Richter and P. Opalack.

Personal business: Tax tips for EEs: Can you deduct residential expenses for doing business at home? New provisions of the tax law are explained. IEEE Spectrum, 13(12):51–54, December 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Richter:1977:CMT

[RO77] P. S. Richter and P. Opalack.

Career management: Taxes and the second income: For the U.S. engineer: Experts tell how to ease the tax bite on your alternate income. IEEE Spectrum, 14 (2):70–73, February 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Roberts:1974:CRI

[Rob74a] L. G. Roberts. Computer report III: Data by the packet: Because computing costs are now so low, an unusual new concept of data communications is feasible. IEEE Spectrum, 11 (2):46–51, February 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Roberts:1974:FSR

[Rob74b] R. W. Roberts. Freedom and social responsibilities. IEEE Spectrum, 11(4):59–60, April 1974.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Roizen:1978:VDH

[Roi78] J. Roizen. VTRs’ ‘double helix’: competing tape formats. IEEE Spectrum, 15(12):34–37, December 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Roth:1971:EMU

[Rot71] P. R. Roth. Effective measurements using digital signal analysis. IEEE Spectrum, 8(4):62–70, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rago:1977:F

[RPR^{+}77] J. J. Rago, K. S. Packard, E. Rechtin, R. Bliss, V. Ramamurti, K. Cogliatti, W. T. Dyall, and J. R. Pierce. Forum. IEEE Spectrum, 14(12):12–14, December 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rosenthal:1974:DCA

[RRS^{+}74] L. H. Rosenthal, L. R. Rabiner, R. W. Schafer, P. Cummiskey, and J. L. Flanagan. Data communications: Automatic voice response: Interfacing man with machine: This developing technique can relieve the human from repetitive instructional chores by piecing together his recorded words. IEEE Spectrum, 11(7):61–68, July 1974.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Redmond:1977:LPW

[RS77] Kent C. Redmond and Thomas M.

Smith. Lessons from Project Whirlwind: a talented team produced many computer “firsts”, and added new chapters to R&D management texts. IEEE Spectrum, 14(10):50–59, October 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubin:1973:WW

[Rub73a] M. D. Rubin. Waste not, want not. IEEE Spectrum, 10 (1):67–70, January 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1973:PP

[Rub73b] E. Rubinstein. Profiles in persistence. IEEE Spectrum, 10 (11):52–64, November 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1974:CES

[Rub74a] E. Rubinstein. Careers: EEs’ salaries: On the inflationary rack: With employees’ real disposable income plummeting, some corporations are reevaluating traditional wage policies. IEEE Spectrum, 11(7):78– 79, July 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1974:CLM

[Rub74b] E. Rubinstein. Careers: The latest manpower forecasts show EEs’ prospects improving: But supply/demand statistics vary with the source and efforts are underway to sort out the confusion. IEEE Spectrum, 11 (8):73–80, August 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1974:CUE

[Rub74c] E. Rubinstein. Careers: The under-30 EE speaks out: In firsthand interviews, young electrical engineers reveal their pleasures and displeasures with the Institute. IEEE Spectrum, 11 (4):71–75, April 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1974:CWN

[Rub74d] E. Rubinstein. Careers: Who needs EEs? And when, and where? IEEE is seeking answers that should help obviate mass dislocations of engineers. IEEE Spectrum, 11 (3):92–97, March 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1974:SAT

[Rub74e] E. Rubinstein. Societal aspects of technology: Embedded within the energy crisis are lessons applicable to other areas of sociotechnical impact. IEEE Spectrum, 11(1):86–90, January 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1974:SPA

[Rub74f] E. Rubinstein. Students and programs: the age of alternatives: who are the new EEs? There is great diversity in the hails of engineering academia … both in programs and people. IEEE Spectrum, 11(5):46–51, May 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1975:CAI

[Rub75a] E. Rubinstein. Careers: Age and inflation are conspiring against the EE: Further results from the 1975 IEEE U.S. Membership Survey permit the reader to place himself in terms of salary. IEEE Spectrum, 12 (8):62–66, August 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1975:CEM

[Rub75b] E. Rubinstein. Careers: EE minorities and discrimination: Survey results describe how women, Blacks, Oriental and Spanish-surnamed members perceive their status in the profession. IEEE Spectrum, 12(9): 72–75, September 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1975:CES

[Rub75c] E. Rubinstein. Careers: EEs’ salaries: Up but down: Returns from IEEE’s 1975 U.S. member survey suggest that you may or may not be better off today than you were in 1972. IEEE Spectrum, 12(7):46–49, July 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1975:CUE

[Rub75d] E. Rubinstein. Careers: The unemployed EE: At survey time, more than one quarter of those unemployed had been out of work for more than six months. IEEE Spectrum, 12(9): 68–71, September 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1975:ECC

[Rub75e] E. Rubinstein. Energy: Cash crunch: Its ripple effect: How the financial difficulties of U.S. electric utilities threaten to erode diverse segments of the economy. IEEE Spectrum, 12(5):41–43, May 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1975:TSE

[Rub75f] E. Rubinstein. Toys I/survey: The EE at leisure: Scores of readers responded to Spectrum’s mini-study of how they invest, or while away, their spare hours. IEEE Spectrum, 12(10): 72–75, October 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1976:CUS

[Rub76a] E. Rubinstein. Careers: Uncle Sam wants you! But do you want him? Not necessarily, says IEEE’s USAB, if the Government is offering you a service contract. Here’s why. IEEE Spectrum, 13(11):62–64, November 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1976:FIA

[Rub76b] E. Rubinstein. Financial: IEEE:

Another AMA? The inside story off how inflation and recession — those masters off extortion — are devouring professional society dues dollars. IEEE Spectrum, 13(3):64–69, March 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1976:FPD

[Rub76c] E. Rubinstein. Financial: Pres.

Dillard on the budget: Pulling no punches, IEEE President Dillard addresses the Institute’s financial problems and their implications on the member’s pocketbook. IEEE Spectrum, 13 (3):70–73, March 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1976:GIE

[Rub76d] E. Rubinstein. Government interface: EEs inside Congress: a close look at the much-maligned U.S. legislative process by five EEs who tried the System, liked it, but found it needs IEEE. IEEE Spectrum, 13(11): 58–61, November 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1976:OIT

[Rub76e] E. Rubinstein. Organization: IEEE today: Who can help you, and who does what at Headquarters — a quick reference guide identifying the staff that serves the Institute. IEEE Spectrum, 13(6):74–78, June 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1976:PIFa

[Rub76f] E. Rubinstein. Perspective: IEEE and the Founder Societies: Who are they, how did they develop, and what are they doing for their members? A comparison of five major engineering societies. IEEE Spectrum, 13(5):76–84, May 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1976:PIFb

[Rub76g] E. Rubinstein. Perspective: IEEE and the ‘Founders’ — II: AIChE is born, IRE’s roots are traced, the merger takes place creating IEEE, and today’s societies’ finances and futures are scanned. IEEE Spectrum, 13 (6):67–73, June 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1976:PVC

[Rub76h] E. Rubinstein. Policy: VPs Cotellessa and Saunders: How IEEE sets policy: Spectrum probes Institute policy-making, especially the IEEE position opposing the California Nuclear Initiative. IEEE Spectrum, 13 (5):85–88, May 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1976:SSS

[Rub76i] E. Rubinstein. Satellites, sensors, and society: Interface topics are: Satellite broadcasting, sensors in the Middle East, and energy legislation. IEEE Spectrum, 13(1):91–94, January 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1976:TSB

[Rub76j] E. Rubinstein. Treasurer Schneider on the budget: a tale of continuing reductions in IEEE’s 1976 operating budget and some dour implications for 1977. IEEE Spectrum, 13(7):95–99, July 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1976:WWW

[Rub76k] E. Rubinstein. What went wrong VII: Dollars vs. satellites: In 1974, fueled by budget constraints, NASA rocketed out off the communications satellite business; now it’s evaluating reentry. IEEE Spectrum, 13 (10):75–82, October 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1977:EAD

[Rub77a] E. Rubinstein. Education: Accreditation: a debatable issue: Key officers of IEEE express diverging views on the role the Institute should play in accrediting electrical engineering programs. IEEE Spectrum, 14 (3):84–88, March 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1977:FIL

[Rub77b] E. Rubinstein. Feedback from IEEE’s long-range planning report: While the report may have largely failed as a blueprint for change, it has confirmed the Institute’s need for advance planning. IEEE Spectrum, 14(7):47– 48, July 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1977:INS

[Rub77c] E. Rubinstein. IEEE: Is a new structure needed? The long range planning committee believes one is, and, in its most recent report, suggests one possible way of going about it. IEEE Spectrum, 14(6):60–61, June 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1977:ITL

[Rub77d] E. Rubinstein. Independent test labs: Caveat emptor: Although this industry is nearly 15 years old, is growing rapidly, and performs a needed service, it remains fraught with controversy. IEEE Spectrum, 14 (6):44–52, June 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1977:RI

[Rub77e] E. Rubinstein. Rent-an-instrument. IEEE Spectrum, 14(4):48–52, April 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1977:TSI

[Rub77f] E. Rubinstein. Technology and society: Indonesia is “wired,” while U.S. public interest groups “monitor” nuclear technology. IEEE Spectrum, 14 (1):82–86, January 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1977:USS

[Rub77g] E. Rubinstein. U.S. shows stage a comeback: a shell game played without a pea, the electronics exhibition is prey to economics and psychology. IEEE Spectrum, 14(4):53–55, April 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1978:I

[Rub78a] E. Rubinstein. Introduction. IEEE Spectrum, 15(10): 32–33, October 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1978:TSW

[Rub78b] E. Rubinstein. Technology and society: While politicians struggle with communications and energy policy, GE ‘engineers’ the perfect tomato. IEEE Spectrum, 15(1):73–78, January 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1979:ASH

[Rub79a] E. Rubinstein. The accident that shouldn’t have happened: a narrative account of what is believed to have occurred, based on reports from many experts. IEEE Spectrum, 16 (11):33, November 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1979:SIT

[Rub79b] E. Rubinstein. Special issue: Three Mile Island and the future of nuclear power [special issue intro.]. IEEE Spectrum, 16 (11):30–32, November 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rubinstein:1979:TSA

[Rub79c] E. Rubinstein. Technology and society: Action on the international telecommunications front; the latest in energy R&D; agriculture goes electronic. IEEE Spectrum, 16(1):81–86, January 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Rudin:1970:DTD

[Rud70] H. Rudin. Data transmission: a direction for future development. IEEE Spectrum, 7(2):

79–85, February 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Steele:1978:MUN

[SA78] R. Steele and R. Aurora. Military: The U.S. Navy bets on V/STOL: Integrating technological advances on a broad front is the key to successful vertical/short takeoff and landing aircraft. IEEE Spectrum, 15(9): 58–63, September 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sack:1977:IRR

[Sac77] E. Sack. I2L: Rugged and reliable. IEEE Spectrum, 14(6):34, June 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Salihi:1972:ECF

[Sal72a] J. T. Salihi. The electric car — fact and fancy. IEEE Spectrum, 9(6):44–48, June 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Salihi:1972:TR

[Sal72b] J. T. Salihi. Two for the road. IEEE Spectrum, 9(7):43– 47, July 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Saloschin:1973:SSA

[Sal73] R. L. Saloschin. Science and secrecy about EEs’ salaries. IEEE Spectrum, 10(11):72–75, November 1973. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Salihi:1975:KVL

[Sal75] J. T. Salihi. Kilowatthours vs. liters. IEEE Spectrum, 12 (3):62–67, March 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sallet:1977:MTH

[Sal77] H. W. Sallet. Magnetic tape: a high performer: Magnetic tape has evolved into a highly effective medium for high-density and low-cost-per-bit data recording. IEEE Spectrum, 14(7):26– 31, July 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Siegman:1975:BR

[SB75] A. E. Siegman and E. Brookner. Book reviews. IEEE Spectrum, 12(1):119–120, January 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schoenfeld:1977:BR

[SC77] R. L. Schoenfeld and E. R. Cacciamani. Book reviews. IEEE Spectrum, 14(11):106– 107, November 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schindler:1970:DM

[Sch70] H. R. Schindler. Delta modulation. IEEE Spectrum, 7 (10):69–78, October 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schick:1971:LCA

[Sch71] L. L. Schick. Linear circuit applications of operational amplifiers. IEEE Spectrum, 8 (4):36–50, April 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schneider:1977:RT

[Sch77a] B. H. Schneider. Report of the treasurer. IEEE Spectrum, 14(7):49–51, July 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schwartz:1977:CKW

[Sch77b] H. J. Schwartz. Cars and kilowatts: While Detroit remains uncommitted to electric and hybrid vehicles, private industry in Europe and Japan has already taken the plunge. IEEE Spectrum, 14 (11):65–68, November 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schweppe:1978:PSH

[Sch78] F. C. Schweppe. Power systems ‘2000’: hierarchical control strategies. IEEE Spectrum, 15(7):42–47, July 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Scott:1973:ES

[Sco73] N. R. Scott. Engineers are slipping. IEEE Spectrum, 10 (8):41–43, August 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Selzer:1971:SVR

[Sel71] A. Selzer. Switching in vacuum. A review. IEEE Spectrum, 8(6):26–37, June 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Smiley:1976:FG

[SESJ76] S. H. Smiley, M. L. Ernst, G. Sege, and R. T. Jaske. Feeding the glutton. IEEE Spectrum, 13(7):74–83, July 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sheingold:1972:UDD

[SF72a] D. H. Sheingold and R. A. Ferrero. Understanding A/D and D/A converters. IEEE Spectrum, 9(9):47–56, September 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sheingold:1972:UCC

[SF72b] D. H. Sheingold and R. A. Ferrero. Understanding converter circuitry. IEEE Spectrum, 9 (10):80–86, October 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sherman:1978:EC

[She78] H. Sherman. Electrotechnology in China. IEEE Spectrum, 15 (2):61–66, February 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sherwood:1979:CCS

[She79] B. A. Sherwood. Computers: The computer speaks: Rapid speech synthesis from printed text input could accommodate an unlimited vocabulary. IEEE Spectrum, 16(8):18–25, August 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shiers:1970:EST

[Shi70] G. Shiers. Early schemes for television. IEEE Spectrum, 7 (5):24–34, May 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shiers:1972:OED

[Shi72] G. Shiers. On the origins of electron devices. IEEE Spectrum, 9 (11):70–76, November 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shiromizu:1976:ETS

[Shi76] M. Shiromizu. Electronic telephone switching in Japan. IEEE Spectrum, 13(6):65–66, June 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shima:1979:SSC

[Shi79a]  M. Shima.  Solid state/

computers: Demystifying microprocessor design: Big chips require artful compromises, as demonstrated by this design case history. IEEE Spectrum, 16(7):22–30, July 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shimada:1979:IJU

[Shi79b] S. Shimada. II. Japan: Unusual applications. IEEE Spectrum, 16 (10):74–76, October 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Simrall:1971:IYC

[Sim71] H. C. Simrall. If you are a concerned engineer …. IEEE Spectrum, 8(2):69–71, February 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Simon:1972:SSB

[Sim72a] R. E. Simon. A solid-state boost for electron-emission devices. IEEE Spectrum, 9(12): 74–78, December 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Simpson:1972:CNE

[Sim72b] J. W. Simpson. The case for nuclear energy. IEEE Spectrum, 9 (9):71–73, September 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Shefer:1973:HRH

[SK73] J. Shefer and R. J. Klensch. Harmonic radar helps autos avoid collisions. IEEE Spectrum, 10 (5):38–45, May 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sanderson:1978:CMP

[SL78] L. B. Sanderson and J. C. Lord. Computers: Microcomputers promise less stop, more go: Designers off urban trafficcontrol systems propose hierarchical schemes to enhance local control, reliability, and maintainability. IEEE Spectrum, 15(11): 30–32, November 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Snelling:1972:FLAa

[Sne72a] E. C. Snelling. Ferrites for linear applications. I. Properties. IEEE Spectrum, 9(1):42–51, January 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Snelling:1972:FLAb

[Sne72b] E. C. Snelling. Ferrites for linear applications II. Performance requirements. IEEE Spectrum, 9 (2):26–32, February 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Soanes:1970:PHE

[Soa70] S. V. Soanes. A philatelic history of electrical science. IEEE Spectrum, 7(1):80–85, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sorenson:1970:LSE

[Sor70] H. W. Sorenson. Least-squares estimation: from Gauss to Kalman. IEEE Spectrum, 7 (7):63–68, July 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sugano:1977:SSS

[SOW77] T. Sugano, T. Okoshi, and M. Watanabe. Solid state: Synonym for progress: For Japan, semiconductor technology is where electronics really began, and where it is going. IEEE Spectrum, 14(9): 51–56H, September 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Spencer:1978:BDE

[Spe78] W. J. Spencer. Biomedical: For diabetics: an electronic pancreas: a bionic device implanted below the human diaphragm will deliver insulin into the peritoneum and be inductively programmable. IEEE Spectrum, 15(6):38–42, June 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Spitzer:1976:CVC

[Spi76a] C. R. Spitzer. Communications: The Vikings are coming: NASA’s first Viking Lander will touch down in just a few weeks, initiating a carefully planned search for Martian life. IEEE Spectrum, 13(6):49–54, June 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Spitzer:1976:WWW

[Spi76b] C. R. Spitzer. What went wrong X: Unlimbering Viking’s scoop: There was no assurance on this first of its kind mission that nothing would go wrong. When it did, the experts were ready. IEEE Spectrum, 13(10):92–94, October 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sporn:1971:OEO

[Spo71] P. Sporn. Our environment:

options on the way into the future. IEEE Spectrum, 8 (5):49–58, May 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stillings:1978:EMW

[SR78] D. Stillings and N. Roth. Electronics and medicine: When electroquackery thrived: The mystery and glamor of electricity were exploited by charlatans who bilked the public with promises to cure any ill. IEEE Spectrum, 15(11):56–61, November 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Staras:1971:SCL

[SS71] H. Staras and L. Schiff. Spectrum conservation in the land mobile radio services. IEEE Spectrum, 8(7):28–36, July 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sobol:1972:MPS

[SS72] H. Sobol and F. Sterzer. Microwave power sources. IEEE Spectrum, 9(4):20–33, April 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Schlicke:1973:GNI

[SS73] H. M. Schlicke and O. J. Struger.

Getting noise immunity in industrial controls. IEEE Spectrum, 10(6):30–35, June 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Starr:1972:MNP

[Sta72] E. A. Starr. Measuring noise pollution. IEEE Spectrum, 9 (6):18–25, June 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stevens:1973:KYS

[Ste73] R. T. Stevens. Keeping your system ‘up’. IEEE Spectrum, 10 (11):76–81, November 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stern:1975:INR

[Ste75a] A. P. Stern. Institute news: Report from the president: The news is largely good: Membership is up, professional and technical activities are growing, and the outlook is sound. IEEE Spectrum, 12(9):65–67, September 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stern:1975:SIC

[Ste75b] A. P. Stern. Sociotechnology IEEE Code of Ethics: In response to the members’ mandate, a set of guidelines on professional relationships is presented. IEEE Spectrum, 12(2):65, February 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Steiner:1977:NFF

[Ste77] D. Steiner. Nuclear fusion: Focus on tokamak: U.S. engineers and physicists will team to attempt commercial demonstration by the year 2000. IEEE Spectrum, 14(7):32–39, July 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stork:1977:EHI

[Sto77] E. O. Stork. EPA: How it fights pollution. IEEE Spectrum, 14 (11):51–52, November 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Strasser:1971:ISP

[Str71] G. Strasser. Impediments to societal problem solving what must happen before we can succeed? IEEE Spectrum, 8(7):43– 48, July 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Stroke:1972:OC

[Str72] G. W. Stroke. Optical computing. IEEE Spectrum, 9(12): 24–41, December 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sugarman:1978:CSC

[Sug78a] R. Sugarman. Computers: a second chance for computer-aided instruction: What networking promised but didn’t deliver may be developed at the local level as hardware costs drop. IEEE Spectrum, 15(8):29–37, August 1978.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sugarman:1978:CMG

[Sug78b] R. Sugarman. Computers are making giant strides in automating production — but human decision making is still crucial. IEEE Spectrum, 15(10): 53–60, October 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sugarman:1978:PEN

[Sug78c] R. Sugarman. Power/energy: New York City’s blackout: a $350 million drain: Ripple effects off the July 13, 1977, lightning stroke cost the public dearly in lost property, services, end income. IEEE Spectrum, 15(11): 44–46, November 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sugarman:1978:PIE

[Sug78d] R. Sugarman. Productivity ii: Electrotechnology to the rescue. IEEE Spectrum, 15(10):53, October 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sugarman:1978:WNT

[Sug78e] R. Sugarman. ‘What’s new, teacher?’ ask the computer. IEEE Spectrum, 15(9):44–51, September 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sugarman:1979:CFC

[Sug79a] R. Sugarman. Communications:

On foiling computer crime: Can a thief with lots of LSI chips crack present data scrambling schemes? The experts debate the matter. IEEE Spectrum, 16(7):31–32, July 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sugarman:1979:COM

[Sug79b] R. Sugarman. Computers: our ‘microuniverse’ expands. IEEE Spectrum, 16(1):32–37, January 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Sugarman:1979:NPP

[Sug79c] R. Sugarman. Nuclear power and the public risk: TMI underscored generic human factors and safety system problems, and challenged the complex art of risk assessment. IEEE Spectrum, 16 (11):59–80, November 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Suran:1970:PIE

[Sur70] J. J. Suran. A perspective on integrated electronics. IEEE Spectrum, 7(1):67–79, January 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Suran:1972:EMA

[Sur72] J. J. Suran. Econometric model aids decision makers. IEEE Spectrum, 9(10):87–94, October 1972.

CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Susskind:1970:EHEa

[Sus70a] C. Susskind. The early history of electronics V. Commercial beginnings. IEEE Spectrum, 7(4):78–83, April 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Susskind:1970:EHEb

[Sus70b] C. Susskind. The early history of electronics VI. Discovery of the electron. IEEE Spectrum, 7 (9):76–79, September 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Susskind:1973:DLN

[Sus73] A. K. Susskind. Diagnostics for logic networks. IEEE Spectrum, 10(10):40–47, October 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Svigals:1974:CMS

[SZ74] J. Svigals and H. A. Ziegler. Computers: Magnetic-stripe credit cards: Big business in the offing: a new type of credit card brings a banker, or his cashdispensing machine, on line and speeds checkout and payment for retail goods. IEEE Spectrum, 11 (12):56–62, December 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tanner:1972:FOF

[Tan72] R. H. Tanner. Financing our future. IEEE Spectrum, 9 (8):88–89, August 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tufte:1973:OMC

[TC73a] O. N. Tufte and D. Chen. Optical memories: controlling the beam. IEEE Spectrum, 10 (3):48–53, March 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tufte:1973:OTD

[TC73b] O. N. Tufte and D. Chen. Optical techniques for data storage. IEEE Spectrum, 10(2): 26–32, February 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Thompson:1974:CWY

[TDK74] P. Thompson, G. Dalton, and R. Kopelman. Careers: “But what have you done for me lately?” — the boss. IEEE Spectrum, 11(10):85–89, October 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tinney:1974:CCO

[TE74] W. F. Tinney and M. K. Enns.

Computers: Controlling and optimizing power systems: Economic dispatch, stability, security, and reliable system operation are principal tasks for the computer. IEEE Spectrum, 11(6):56–60, June 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Teague:1974:IEL

[Tea74]  O. E. Teague.  The infor-

mation explosion — a logistics problem. IEEE Spectrum, 11 (4):60–61, April 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tell:1972:BRH

[Tel72] R. A. Tell. Broadcast radiation: how safe is safe? IEEE Spectrum, 9(8):43–51, August 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Teller:1973:FIT

[Tel73] Edward Teller. A future ICE (thermonuclear, that is!). IEEE Spectrum, 10(1):60–64, January 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Terman:1972:CNP

[Ter72] F. E. Terman. Changing needs for Ph.D.’s. IEEE Spectrum, 9 (1):79–81, January 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Terry:1977:TJT

[Ter77] W. Terry. Tougher jobs, tougher instruments: More complex measurement problems are being solved by microprocessors, bussed systems, and software. IEEE Spectrum, 14(4):27–28, April 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tannas:1978:DFP

[TG78] L. E. Tannas and W. F. Goede.

Displays: Flat-panel displays: a critique: They face six hurdles in displacing the CRT: Matrix addressing, duty cycle, luminous efficiency, uniformity/gray scale, color, and cost. IEEE Spectrum, 15(7):26–32, July 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Turoff:1977:TMT

[TH77] M. Turoff and S. R. Hiltz.

Telecommunications: Meeting through your computer: Information exchange and engineering decision-making are made easy through computer-assisted conferencing. IEEE Spectrum, 14(5):58–64, May 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Thompson:1973:WGL

[Tho73a] P. H. Thompson. Who gets laid off? IEEE Spectrum, 10 (12):68–75, December 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Thornton:1973:FLM

[Tho73b] R. D. Thornton. Flying low with Maglev. IEEE Spectrum, 10(4):47–54, April 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tibbetts:1978:NTP

[Tib78] R. T. Tibbetts. NSF’s threephase program helps the smallbusiness innovator bootstrap an idea to commercial success. IEEE Spectrum, 15(10):86, October 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tilson:1970:PMN

[Til70] S. Tilson. Pulsars may be neutron stars. IEEE Spectrum, 7 (2):42–55, February 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Toombs:1978:SSU

[Too78a] D. Toombs. Solid state: an update: CCD and bubble memories: Commercial units include 64-kbit charge-coupled-device and 92-kbit magneticbubble memories. IEEE Spectrum, 15(4):22–30, April 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Toombs:1978:SSC

[Too78b] D. Toombs. Solid state: CCD and bubble memories system implications: Mainframe computers and intelligent terminals are predicted to be primary beneficiaries. IEEE Spectrum, 15 (5):36–39, May 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Torrero:1976:CDH

[Tor76a] E. A. Torrero. Circuits and devices: High-frequency components play catch-up: From RF to mm waves, improved circuit and packaging techniques have increased the density and performance of the latest components. IEEE Spectrum, 13(11): 31–36, November 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Torrero:1976:CWB

[Tor76b] E. A. Torrero. Convention:

WESCON’s 25th: It’s big. IEEE Spectrum, 13(9):54–58, September 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Torrero:1976:MBR

[Tor76c] E. A. Torrero. Memories: Bubbles rise from the lab: Developmental prototypes and even limited production of 65-kb and larger bubble memories are the result of advances on a broad front. IEEE Spectrum, 13(9): 29–32, September 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Torrero:1976:WWW

[Tor76d] E. A. Torrero. What went wrong III: The bypass of 3-D radar: It was already beset by technical and economic problems, and then advancing technology sealed its fate. IEEE Spectrum, 13 (10):49–54, October 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Torrero:1977:AEC

[Tor77a] E. A. Torrero. ATE: Not so easy: As components and subsystems become more complex, automatic testing grows difficult and costly. IEEE Spectrum, 14(4):29–36, April 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Torrero:1977:APL

[Tor77b] E. A. Torrero. Automating the production line: Henry Ford began it all when he designed the first car assembly line in 1914. IEEE Spectrum, 14(11): 71–72, November 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Torrero:1977:EJT

[Tor77c] E. A. Torrero. Electrotechnology in Japan today [special issue intro.]. IEEE Spectrum, 14 (9):34–35, September 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Torrero:1977:MAR

[Tor77d] E. A. Torrero. Military and aerospace: Radars see better; Vikings soar, roost, test, and tell. IEEE Spectrum, 14(1): 74–77, January 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Torrero:1977:MIE

[Tor77e] E. A. Torrero. The multifacets of i2l: Exploiting a technique once assumed merely to encourage annoying parasitics, integrated injection logic is burgeoning. IEEE Spectrum, 14(6):29, June 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Torrero:1977:SSD

[Tor77f] E. A. Torrero. Solid-state devices: The traditional bipolar vs. MOS battle lines are blurring. IEEE Spectrum, 14(1): 48–53, January 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Torrero:1977:TJF

[Tor77g] E. A. Torrero. Technological Japan: Formula for success: The ambition, dedication, and motivation of the Japanese people play an important role. IEEE Spectrum, 14(9):36–39, September 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Torrero:1977:UTC

[Tor77h] E. A. Torrero. Unjamming traffic congestion: Computer-based systems can speed traffic, increase safety, and improve gasoline mileage. IEEE Spectrum, 14 (11):77–80, November 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Torrero:1978:AMV

[Tor78a]  E. A. Torrero.  Aerospace

and military: Voyager begins planets rendezvous; improved night vision for the foot soldier. IEEE Spectrum, 15(1): 63–67, January 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Torrero:1978:SSD

[Tor78b] E. A. Torrero. Solid-state devices: Memories and microprocessors benefit from large-scale integration advances. IEEE Spectrum, 15(1):36–40, January 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Torrero:1979:VOS

[Tor79] E. A. Torrero. VLSI and other solid-state devices: New fabrication technologies put us on the threshold of a quantum jump in IC complexity. IEEE Spectrum, 16(1):43–47, January 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tribus:1974:DTM

[Tri74] M. Tribus. Design techniques: Making up the collective mind: Decision analysis, by combining inputs from diverse and often dissident sources, seeks the best risks for industry or government. IEEE Spectrum, 11 (10):59–68, October 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tribus:1978:EEP

[Tri78] M. Tribus. The engineer: The engineer and public policy-making: Like a gladiator with one arm bound behind him, the engineer enters the public arena handicapped by personality and training. IEEE Spectrum, 15 (4):48–51, April 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tropp:1974:CRV

[Tro74] H. Tropp. Computer report VII:

The effervescent years: a retrospective: Behind the computer revolution lay the vision and perseverance of a handful of pioneers committed to open information exchange. IEEE Spectrum, 11 (2):70–79, February 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tebo:1974:BR

[TSR74] J. D. Tebo, R. W. Schmitt, and B. Reich. Book reviews. IEEE Spectrum, 11(1):126–129, January 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

To:1974:IAT

[TT74] K. To and R. E. Tulloss. Instrumentation: Automatic test systems: Seven basic elements make up the ideal ATS; neglect of any one is an invitation to disaster. IEEE Spectrum, 11(9): 44–52, September 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tucker:1974:SEC

[Tuc74]  E. Tucker.  Sociotechnology:

an engineer in Congress tells us ‘how ii really is’: if you don’t like the direction events seem to be taking, step in and help with the steering, advises this Congressman-cum-engineer. IEEE Spectrum, 11 (7):80–81, July 1974. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tucker:1975:TVI

[Tuc75a] E. Tucker. Tools VI/information retrieval: Engineer’s source guide: Researching a complex problem or looking for the answer to a nagging technical question? This directory may help. IEEE Spectrum, 12(10):67, October 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tucker:1975:TIA

[Tuc75b] E. Tucker. Toys III/avocation: a collector’s guide to romance and nostalgia: Where to find antique vacuum tubes, TRF receivers, glass insulators, dc generators, and old trolley cars. IEEE Spectrum, 12(10):79–80, October 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Tuchman:1979:IHP

[Tuc79] W. Tuchman. IV. ‘Hellman presents no shortcut solutions to the DES’. IEEE Spectrum, 16(7):40–41, July 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Udolf:1973:BFP

[UG73] R. Udolf and I. Gilbert. Behind the front panel. IEEE Spectrum, 10(5):28–31, May 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ullmann:1970:CIP

[Ull70] J. E. Ullmann. Conversion and the import problem: a confluence of opportunities. IEEE Spectrum, 7(4):55–62, April 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Unger:1979:EEH

[Ung79] S. H. Unger. Engineering ethics:

How to be ethical and survive: Some engineering societies offer aid in a conflict, but competence and confidence can help mitigate trauma and duress. IEEE Spectrum, 16(12):56–57, December 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Uram:1977:CCC

[Ura77] R. Uram. Computer control of combined-cycle power plants: Complete automation and coordination of the plant is possible from startup through loading, and from full load to shutdown. IEEE Spectrum, 14(10): 36–41, October 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Utsunomiya:1977:MDV

[Uts77] T. Utsunomiya. Medical diagnostics: Visual/aural techniques: High-resolution scanning systems, a retinal camera, and a cardiac-output earpiece show promise. IEEE Spectrum, 14(9): 69–70, September 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

VanVeen:1971:IIT

[Van71] F. Van Veen. An introduction to IC testing. IEEE Spectrum, 8 (12):28–37, December 1971. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Vadasz:1971:SRA

[VCG71] L. L. Vadasz, H. T. Chua, and A. S. Grove. Semiconductor random-access memories. IEEE Spectrum, 8(5):40–48, May 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Verhofstadt:1977:ICN

[Ver77] P. Verhofstadt. I2L: Coexistence with NMOS. IEEE Spectrum, 14(6):33–34, June 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

VanTrees:1977:CSL

[VHM77] H. L. Van Trees, E. V. Hoversten, and T. P. McGarty. Communications satellites: Looking to the 1980s: It will be a decade of new and exciting applications, lower costs, and more reliable service. IEEE Spectrum, 14(12): 43–51, December 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

VanTuyl:1977:GAS

[VL77] R. Van Tuyl and C. Liechti. Gallium arsenide spawns speed. IEEE Spectrum, 14(3):41–48, March 1977. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

vonWillisen:1979:EWQ

[von79] F. K. von Willisen. Electronic watches: The quartz watch: Its life and times: Intervals between battery replacement will increase as the watch improves in accuracy and versatility. IEEE Spectrum, 16(6):18–23, June 1979. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

VanDuzer:1972:SNR

[VT72] T. Van Duzer and C. W. Turner.

Superconductivity: new roles for an old discovery. IEEE Spectrum, 9(12):53–63, December 1972. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wade:1976:TGA

[Wad76] R. M. Wade. Transportation:

Getting around town: Automatic urban transit: Prospects for personal and group rapid transit hinge on technology, safety, politics, and economics. IEEE Spectrum, 13(3):55–58, March 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Walker:1970:ADC

[Wal70] G. F. Walker. Applications of a dc constant-current source. IEEE Spectrum, 7(9):47–58, September 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wallenstein:1974:EES

[Wal74] G. D. Wallenstein. Essay: Engineers are supersnobs. IEEE Spectrum, 11(4):78–79, April 1974. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Walsh:1978:MEW

[Wal78] T. Walsh. Military: Electronic warfare: From RF to EO: Electrooptics gains on RF and acoustic technologies in electronic deception, jamming, and neutralization. IEEE Spectrum, 15 (2):30–38, February 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Ware:1972:UC

[War72] W. H. Ware. The ultimate computer. IEEE Spectrum, 9 (3):84–91, March 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Warner:1976:SSS

[War76] R. M. Warner. Solid state: Isquared L: a happy merger: Integrated injection logic is a fine example, if not the first, of how to gain high efficiency in semiconductor device design. IEEE Spectrum, 13(5):42–47, May 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Watt:1971:PAC

[Wat71] A. D. Watt. Placing atmospheric CO_{2} in perspective. IEEE Spectrum, 8(11):59–72, November 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Webster:1975:BR

[WB75] J. G. Webster and W. B. Bridges.

Book reviews. IEEE Spectrum, 12(12):74, December 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Weiner:1973:HTE

[Wei73] C. Weiner. How the transistor emerged. IEEE Spectrum, 10 (1):24–33, January 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Weitzman:1977:ACS

[Wei77] C. Weitzman. Aerospace: Conversations with a Space Shuttle. Keeping experimenters in touch with orbiter-launched payloads requires a new Earthbased communications support system. IEEE Spectrum, 14 (3):69–74, March 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Weinberg:1978:CFP

[Wei78a] E. Weinberg. A call for focusing on productivity: It directly affects inflation, cost of living, job stability — and the quality of life. IEEE Spectrum, 15 (10):34–39, October 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Weinschel:1978:UAB

[Wei78b] B. O. Weinschel. The U.S. Activities Board of IEEE recommends actions Congress could take on behalf of the U.S. technologist and businessman. IEEE Spectrum, 15(10):87, October 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wellenstein:1979:LYH

[Wel79]  N. Wellenstein.  LSI yields

high reliability: Cost and size of telephone-interface equipment shrinks, while IC designs displace electromechanical versions. IEEE Spectrum, 16(10): 78–81, October 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

West:1973:CJG

[Wes73] J. S. West. The Corporation: just a granfalloon? IEEE Spectrum, 10(6):44–47, June 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Whitlock:1972:MED

[Whi72]  R. E. Whitlock.  Manufac-

turing as an engineering discipline. IEEE Spectrum, 9 (5):45–46, May 1972. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wickham:1979:PNM

[Wic79] K. Wickham. Pascal is a ‘natural’ [for microprocessor software programs]. IEEE Spectrum, 16 (3):35–41, March 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wilmotte:1970:ETC

[Wil70] R. M. Wilmotte. Engineering truth in competitive environments. IEEE Spectrum, 7 (5):45–49, May 1970. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wilmotte:1971:MR

[Wil71] R. M. Wilmotte. The management and the risk. IEEE Spectrum, 8(4):31–35, April 1971. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wilson:1975:F

[Wil75] T. A. Wilson. Forum. IEEE Spectrum, 12(12):20–26, December 1975. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wilmotte:1976:TFT

[Wil76a]  R. Wilmotte.  Technology

forecasting-tv look-ahead. IEEE Spectrum, 13(2):34–40, February 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wilson:1976:CVD

[Wil76b] P. S. Wilson. Cities VI: Disposa-city: Unchecked abandonment off our older cities for sprawling Sunbelt suburb-cities will have a major impact on the urban future. IEEE Spectrum, 13(7):57– 64, July 1976. CODEN IEESAM.

ISSN 0018-9235 (print), 1939-9340 (electronic).

Withrow:1977:EOC

[Wit77] J. D. Withrow. Engine options for the car of tomorrow: Competing internal and external combustion engines, as well as electrics, are considered as contenders. IEEE Spectrum, 14(11): 57–62, November 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wolff:1975:IIB

[Wol75a] M. F. Wolff. Innovation: Inventing at breakfast. IEEE Spectrum, 12(5):44–49, May 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wolff:1975:IRB

[Wol75b] M. F. Wolff. Innovation: The R&D ‘bootleggers’: Inventing against odds. IEEE Spectrum, 12(7):38–45, July 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wolff:1976:IMA

[Wol76a] M. F. Wolff. Innovation: The marriage that almost was: Western Union has always been ridiculed for rejecting the telephone. But what actually happened wasn’t so ridiculous after all. IEEE Spectrum, 13 (2):41–53, February 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wolff:1976:RGI

[Wol76b] M. F. Wolff. Retrospective:

The genesis of the integrated circuit. IEEE Spectrum, 13 (8):45–53, August 1976. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wolff:1976:WWW

[Wol76c] M. F. Wolff. What went wrong IX: The Navy’s Big Dish. IEEE Spectrum, 13(10):89–91, October 1976. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wolff:1977:MCE

[Wol77] M. F. Wolff. Managing the creative engineer: a probe into the characteristics of the creative engineer, the sensitive manager, and the favorable environment. IEEE Spectrum, 14 (8):52–57, August 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wolff:1979:WHP

[Wol79] M. F. Wolff. Whatever happened to ‘participative’ management? IEEE Spectrum, 16 (2):60–68, February 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Woodson:1979:EAP

[Woo79] T. T. Woodson. Education: Accelerating the practice of engineering. IEEE Spectrum, 16(9): 50–54, September 1979. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wulfsberg:1978:MRB

[Wul78] A. H. Wulfsberg. Measuring risk, both technical and marketing, is an important step in launching successful innovative products. IEEE Spectrum, 15 (10):74–78, October 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Wilson:1978:ABH

[WZ78] G. L. Wilson and P. Zarakas. Anatomy of a blackout: How’s and why’s of the series of events that led to the shutdown of New York’s power in July 1977. IEEE Spectrum, 15(2):39–49, February 1978. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Yen:1978:PHW

[Yen78] J. T. Yen. Power: Harnessing the wind: New schemes may provide solutions to the age-old problem of building efficient wind machines. IEEE Spectrum, 15 (3):42–47, March 1978. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Young:1973:RP

[You73] L. Young. The road to professionalism. IEEE Spectrum, 10 (1):71–78, January 1973. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Yu:1970:MSC

[Yu70] A. Y. C. Yu. The metalsemiconductor contact: an old device with a new future. IEEE Spectrum, 7(3):83–89, March 1970. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Yu:1975:IGS

[Yu75] F. T. S. Yu. Instrumentation generating speech spectrograms optically: Coherent optics speeds analysis of speech samples; it’s five times faster than older methods. IEEE Spectrum, 12(2):51–56, February 1975. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zuch:1977:WWU

[Zuc77] E. L. Zuch. Where and when to use which data converter: a broad shopping list of monolithic, hybrid, and discretecomponent devices is available; the author helps select the most appropriate. IEEE Spectrum, 14(6):39–43, June 1977. CO-DEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).